A Complete Bibliography of Publications in the Journal of Cell Biology: 1985–1989

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
             WWW URL: http://www.math.utah.edu/~beebe/

04 June 2019 Version 1.00

Title word cross-reference

1 [SR89]. 100 [BPE^{+}86, NRL^{+}89b, Rob87]. 148 [BB88a]. 16 [GKBG87]. 17 [GKBG87]. 2 [GS86a, LSDP86, MFB^{+}89]. 20 [DFF85]. 3 [JBG^{+}87, KPK86, NB88b]. 300 [WTB87]. 5 [ODD85]. 6 [BNB^{+}88, BGHvF87, BGF86, DK88, EG89, GF86a, GF86b, GSS^{+}88, MGK89, Pfe87, SGG^{+}89, SE86]. 70 [MPB88]. 83 [YMM88]. ^{+} [MSQ^{+}89, SPO85]. ^{+2} [CWH86b]. ^{2+} [BH86b, BFJ^{+}85, BN85b, CHFA89, DMGS88, FC89b, HCG87, HML88, HWLS88, HB86, ITY^{+}88, KSH87, KS88, MMMP87, MSS^{+}87a, MSLR^{+}86, NRTE89, NGS85, PJW88, PW85, RT86, SS85, Sal86, SL85d, SDDZ^{+}89, SYT^{+}89, VAD^{+}85, VGKF86, dBBB^{+}86].

-adenosine [MEA89]. -ATPase [AMWN89, MSQ^{+}89, NH89, RYR^{+}89, SBC^{+}87, VSZ^{+}87]. -ATPases [vAAA86]. -azacytidine [ODD85]. -binding [HHMH87, HML88, HWLS88, KSH87, dBBB^{+}86]. -bp [BB88a]. -containing [FKP86]. -dependent [JY86, KOB89, NGS85, RT86, SL85d, SYT^{+}89, SRW^{+}89]. -free [CB88]. -Hydroxy- [KPK86, JBG^{+}87]. -hydroxyecdysone [DFF85]. -independent [PJW88]. -induced [HB86, Sal86]. -kD [MPB88, Rob87, WTB87]. -kilodalton [YMM88]. -like [GS86a, SR89, SCRA87]. -macroglobulin [MFB^{+}89]. -methylglutaryl [JBG^{+}87, KPK86]. -modulated [MSS^{+}87a]. -nm [BPE^{+}86, GKBG87]. -phosphate [BNB^{+}88, BGF86, DK88, EG89, GF86a, GF86b, GSS^{+}88, MGK89, NB88b, Pfe87, SGG^{+}89, SE86]. -phosphate-specific [BGHvF87]. -phosphate/insulin-like [EG89]. -regulated [FC89b, GG85, GTP87]. -residue [NRL^{+}89b]. -secreting [SS89b]. -stimulated [LSDP86].

/calmodulin [SS85]. /calmodulin-dependent [SS85]. /H [GWCW85, GCGR85, RT86]. /K [MSLR^{+}86]. /phospholipid [CWH86b]. /phospholipid-dependent [CWH86b].

000 [CFC89, KKB85]. 000- [SNNK85]. 000-D [KKB85]. 000-dalton [Fec87, SGRM85]. 000-mol-wt [LC85, MLM85, SNNK85, Wan85a, YLGG85, YSW^{+}85]. 000-Mr [CFC89].

1 [ABME88, ASJ^{+}88, AVP^{+}89, BMR85, BSK86, DMW88, FMJS89, HMP86, KJF^{+}88, KRN^{+}86, MMR88, MVSW86, NHKE88, PH89, PLA86, PSK^{+}87, RdRBR89, SG88b, SWTB88, SSMG86, SAGM88, WLG85, WPW^{+}85]. 100 [AHRM85, BNLC85, Rob89, WLGB85]. 100-kD [BNLC85, Rob89]. 100K [RP86]. 102.0- [BMDN87]. 110 [CCA^{+}88, CB88, CM87c, GCC^{+}89, MC89b, UK86]. 110-kD [CCA^{+}88, CB88, CM87c, MC89b, UK86]. 110-kD-calmodulin [GCC^{+}89]. 110K [MCC^{+}89, CCA^{+}88, CB87, CB88, VB85b]. 110K-calmodulin [MCC^{+}89, CCA^{+}88, CB87, CB88, VB85b]. 114 [Den88]. 115 [AY89]. 115-kD [AY89]. 12 [DJS^{+}88, MBLK88]. 12-O-tetradecanoylphorbol-13-acetate [SPO85]. 120 [HFG87]. 120K [CC85a]. 125I [BD86a, BD86b, CGV^{+}86, VS89]. 125I-labeled [BD86a, BD86b]. 125I-photoreactive [CGV^{+}86]. 125I-tissue [VS89]. 130 [Ran88, Rog87a]. 130-kD [Ran88]. 135 [VG86a, VG86b]. 135-kD [VG86a, VG86b]. 14 [YBSR87]. 140 [AYY86, SSSA87]. 140-kD [AYY86, SSSA87]. 140K [CCM86]. 14S [MRWJ88]. 160 [PSS89]. 17 [POW^{+}87, WL87]. 17-kD [POW^{+}87, WL87]. 18 [DRH^{+}87]. 180 [DLC88, PSS89, PSGS85]. 180-kD [DLC88, PSGS85]. 190 [MKA^{+}86]. 190-kD [MKA^{+}86]. 195 [MS86a]. 195-kD [MS86a]. 1A [SH87b]. 1B [DNSMA88]. 1C [PSV87, SPV88].

2’ [CMW87, AVP^{+}89, CF88a, GM85b, HMP86, HLU^{+}87, MMMP87, SVF^{+}89, SHSvdM89, TBK86, ZH86, vAAA86]. 2-dependent [SVF^{+}89]. 2-mediated [vAAA86]. 2’-O-ribose [CMW87]. 2/collagen [TH89a]. 20 [CPM86, CMK87, MBLK88, PHM85]. 21 [CMT^{+}88, TLD^{+}89]. 21-kD [TLD^{+}89]. 22S [Toy87a, Toy87b]. 230.0-nm [BMDN87]. 235E [RZG86]. 240/235 [RZG86]. 240/235E [RZG86]. 25 [OHH^{+}89]. 25-Dihydroxyvitamin [PBK88]. 252a [Has88]. 25K [OMS89]. 26 [TLD^{+}89]. 26-kD [TLD^{+}89]. 270K [Hir86]. 27S [HKA^{+}87]. 28 [DYT^{+}87, ZHE85]. 28-kD [DYT^{+}87, ZHE85]. 29 [TRC^{+}87]. 2A [PCBD89]. 2ar [NWE^{+}88]. 2H3 [BGM^{+}87].

3 [CCL^{+}86, KFW^{+}89]. 3-O-sulfate [CCL^{+}86]. 300K [LYG85]. 31P [JT85]. 32D [MMK^{+}89]. 34 [KRN^{+}86]. 34-kD [KRN^{+}86]. 350 [SKE86]. 350-kD [SKE86]. 36/32 [CCMB87]. 37 [MBKR89]. 37-kD [MBKR89]. 39 [KJW89]. 39-kD [KJW89]. 3H [NMKS85, NFN87, PSBN85]. 3T3 [BLSK88, AAM^{+}87, AR85, BSK89, BGL^{+}88, CM88a, CHS85, DBE^{+}87, DWFT88, FSM88, GPSC86, HP85a, HW86b, JS88a, LWT85, LPR^{+}89, LPLT85, MSLR^{+}86, MQ89, MDW86, OMY^{+}85, SPO85, TL89, TT86a, ZSSR86]. 3T3-adipocytes [CHS85]. 3T3-L1 [AR85, BGL^{+}88, TL89]. 3Y1 [SNNK85].

4 [DFSK87]. 4-dihydropyridine [JSA^{+}89]. 4.1-like [ABM86, BSK86]. 40 [DRD^{+}88, TYP^{+}87]. 400 [TIT89]. 400-kD [TIT89]. 40S [LAC^{+}85]. 43 [BKPG89, KCN^{+}89, MMS89, MFM89, SV89, YJL^{+}89]. 43-kD [BKPG89, KCN^{+}89, MMS89, MFM89, YJL^{+}89]. 431 [Bre89, CWOK87]. 43K [BF87, MCCM88, PF85a]. 450 [DFSK87, MYT87, YMT85].

5-bisphosphate [HMP86, MWS85]. 5-trisphosphate [TJF86]. 5-trisphosphate-induced [Miy88]. 50 [VHO^{+}89]. 52 [BCD86, CMB^{+}87]. 52-kD [BCD86, CMB^{+}87]. 53 [SFB^{+}88]. 53-kD [SFB^{+}88]. 54 [HPW89]. 54-kD [HPW89]. 55 [BBBM89, YMM86]. 55-kD [BBBM89, YMM86]. 58 [SPF87]. 58-kD [SPF87]. 58K [FMT^{+}87]. 5S [KW86b, SBH^{+}88].

62 [FS88a]. 62-kD [FS88a]. 64K [SGS86]. 65 [ZP85]. 65-kD [ZP85]. 66 [PL85a]. 66-kD [PL85a]. 69 [YBWLM88]. 69-kD [YBWLM88].

70 [HS89, KKB85]. 70-kD [HS89]. 73 [CRLW86]. 78 [CIK^{+}89, KTF87a]. 78-44-kD [KTF87a].

80 [UPL^{+}87]. 80-kD [UPL^{+}87]. 82 [THT89]. 82-kD [THT89]. 85 [JBdlTB87]. 85-95-kD [JBdlTB87]. 87 [CFC89]. 8Z30 [MNC86]. 8Z31 [MNC86].

9 [DG89]. 90 [LWMM^{+}85]. 937 [GKS87a]. 93D [BGT^{+}89]. 95K [CC85a].

A-1 [BMR85]. A-431 [Bre89, CWOK87]. A-activated [WGC85]. A-binding [KTF87a, SF88b]. A-CAM [VG86a, VG86b]. A-gold [BTR^{+}86, RFH85]. a/b [CHK^{+}86, DSA86, KHC^{+}86, LA87, LM85b, LU88]. A/cholesterol [CCC89]. A2 [BSSMF88, CKR^{+}85b]. A23187 [RTK^{+}86].

A431 [MBK^{+}86, WBD^{+}86, DAJJ^{+}88, GBH87, LWR85, SPLW89, WBW89, Wil88]. A549 [KOBLM88]. A59 [TTW88]. aberrant [FCH^{+}87, OMY^{+}85]. aberrations [MSC^{+}85]. Abnormal [FJR87, SKDF88, TIS^{+}86]. abnormalities [OB88]. aborted [GORW88]. Abrogation [SP85]. absence [BPC89, BP87b, CS86c, DDB^{+}88, DWO^{+}87, Kam88, Pfe87, SR85b, Wat85]. absent [TLC85]. Absorptive [HSMCL87, MBC86, RLB85, WWN87]. abundance [FCL86, Hol89a, MSS^{+}89, RJH^{+}87]. Abundant [SEG^{+}86, FBP89, LSC86, NB88a, PKS^{+}89, SBG^{+}85]. Acanthamoeba [VKP89, HKKP86, HBPK87, KSEP86, MBK89, RP89, SP89, SSP89]. Accelerated [DKH^{+}85, KS88]. acceptor [BD86b, RKFN88]. acceptor-mediated [BD86b]. acceptors [BD86a]. Access [CCG89, Col86]. Accessibility [MKK89, SM89a]. accessible [JFP89]. accessory [KOF88, VC89]. accompanied [GF86b, ICY^{+}88, SSW^{+}85]. accompanies [GK85]. accompany [Bre89, SBS87b, Wan85b]. accompanying [CPKJ87]. accordance [IKH87]. accounts [BKA^{+}89]. accumulates [MNC86]. Accumulation [CRF^{+}88, TB87, AM85, CM87b, DKH^{+}85, DMD^{+}86, GF86a, MP86a, PRS^{+}88, SE85, SSK^{+}88, SHR86, UF86, WPW^{+}85, ZDH88]. accuracy [FS86b]. acetate [CVWZ89, FTC87, LEFS^{+}87, LEG^{+}89, LMF^{+}86, MMT85, PKN^{+}89, SVG^{+}86, SPO85]. acetate-inducible [CVWZ89]. acetyl [TTW88]. acetylase [GMLR85, RCC^{+}88]. Acetylated [SGBG87, CDB87, DD87, LP86, LLCA89, PF85b, PLC87]. acetylation [BRP88, CRC^{+}87, LLCA89, MGR86, PFTV89]. Acetylcholine [ER87, GDM^{+}88, GRME^{+}89, HKPS89, MCCM88, Pum89, SF88b, AJCMR88, BB88a, BYB^{+}88, BGH^{+}87a, Blo86, BF87, BM89, BBMW85, CFC89, CAG89, CPG^{+}89, DLFE89, EZBW88, EZWB88, EMZ85, FG89, FKC87, FC89a, FMT^{+}87, HB88b, HB88c, HSS89, JLB86, JB87, KSSP86, KCN^{+}89, KRLU87, LWSF85, LP85b, MBKR89, MMS89, MFM89, PF85a, PB87, RRF87, RRP88, SCLG86, SMH88, SBS87a, SS89c, TTGFC85, TS89, UTK88, UF86, Wal86, Wal88, WWW^{+}85, WTB87]. acetylcholinesterase [AED^{+}89, AM85, BPJBF85, BMGI85, NPRR86, Wal86]. acetylglucosamine [HSS^{+}87]. acetyltransferase [AR88, CRC^{+}87]. Achlya [Kro86]. acid [ASA^{+}87, AGPT88, BHS88, BBP88, Bis86, BDWH88, BB88b, CHH^{+}88, DH87, DCB^{+}88, EBBW87, GF86a, GF86b, GF87, GDO^{+}85, GKZ89, GMCH87, GSTC87, HKPS89, KSSP86, KF89b, LLRR87, MGSG88, MRD^{+}89, NMKS85, RH86a, RBMM88, RMM89, SHC87, TCK87, VKP89]. acid-binding [Bis86, RMM89]. acid-containing [TCK87]. acid-induced [LLRR87]. acid-modulated [BBP88]. Acidic [LOH^{+}86, AO88, OCP86, PLS89, SS89b, SKK^{+}85, SQM89, WMM86]. Acidification [RYR^{+}89, SOPvD87, SFK^{+}89, YM87a, BGN^{+}88, BNLC85, CM88a, Heu89b, ORA^{+}86, YM87b]. acidification-defective [SFK^{+}89]. acidophilic [HFZM87]. acids [LWSF85, SSK89]. acinar [LMWG85]. acini [MBKO86]. acquire [BMVB89, SDDZ^{+}89, YBBS87]. acquires [FCK87]. Acquisition [BPMD87]. Acrosomal [TI85, AdS85, TI87]. acrosome [CDL^{+}85, LS89, LS87c, TES^{+}87]. across [CMB^{+}86, GW88a, HW88, Ibr87a, Ibr87b, KJW89, RG86b, SFJU86, SW85, SPL87, SSK89, WCB86]. act [CHG87, SGP88]. ACTH [TT86b]. ACTH-containing [TT86b]. Actin [Blo86, MFA89, Pal87, TTC88, TP89, AW89, AT86a, ABG^{+}89, AKVF88, APF^{+}88, BTL^{+}86, BBV^{+}89, BB87b, BTO88, BM86, BFPS85, BR85, BD89, BP87a, BC85, BH86b, Bry88, BDCT88, BBH87, CC85a, CWZ86, CCMB87, CJY86, CNB89, CCK^{+}88, CM85, CHM^{+}87, CB87, CB88, CM87c, Coo87, CF88b, DWFT88, DHH89, DI87, DT89, DF86b, DD88, DMB88, ES85, FL88, Fec87, FS88b, GTP87, GVGL85, HWDC89, HS86a, HCS89, HS86b, HZS88, HP85a, HF85b, HNH^{+}87, HO85, IF85, JBTA88, KR88, KS88, LU87, LG88, LFC^{+}88, LL89a, LL86, LJC^{+}87, MHK^{+}85, MRD^{+}89, MF87a, MSA88, MMM87, MGGP88, NRL^{+}89b, NJAS89, OH89, OPN^{+}85, OKLB86, OE85, PG86, PPG85, PM87b, Pol86, RP89, RS85, RG86b, RS88, SMVS^{+}86, SLB^{+}89, SJ86, SL87a, SL86a, SL88, SSS85, SSY^{+}86]. actin [SLL88, SB85d, SRT^{+}86, SOP85, SDPZ89, SD87, SB85f, TRCR89b, TN89, TI85, TFD87, TT88a, TT88b, TDR^{+}87b, TSS89, UK86, UW86, VKP89, VB85b, VS85, Wan85d, Wan87, WGS^{+}88, WSW85, WGPW89, WS85, WSV^{+}89, WL86, WL87, YMM86, YMM88, YSW^{+}85, YIJ88, ZV89a, Pol86]. actin- [DF86b, JBTA88]. actin-activated [CM87c, PG86]. actin-associated [SLL88]. Actin-binding [MFA89, AKVF88, BTL^{+}86, BP87a, BH86b, Bry88, BBH87, DMB88, GTP87, HS86a, OPN^{+}85, RP89, SD87, SB85f, UK86]. actin-bundling [MMM87, PM87b, SB85d, YMM86]. actin-capping [CCMB87]. actin-containing [DI87]. actin-myosin [CM85]. actin-nucleating [CWZ86]. actin-severing [APF^{+}88]. acting [BB88a, RIS87]. actinin [BL85, BHCD^{+}89, CC85a, DF86b, DBS^{+}89, HWSK89, MHK^{+}85, MMW85b, MJW86, MW86, SMPS86, SW87b, SW88, TM87b, WSLN86, YMM86]. actinins [NRL^{+}89b, SCBR85]. actins [HD85, MMW85a, SL89]. action [AS86, BNF^{+}87, CWH86b, MTA86]. Actions [FS88b, MBL89]. activate [EMZ85]. activated [Bro87, CM87c, HP87b, HMP89, JS86a, PG86, PSV87, SPV88, VRLG89, WGC85]. activates [SEB89, TS89]. activating [CR86, PUW^{+}88]. Activation [BMBL89, BIES89, BFJ^{+}85, Col86, GCR^{+}86, SHL^{+}89, SPT^{+}89, BDWH88, BA87, GCGR85, HPO88, HCHK89, LKOM88, NHKE88, OS88b, PON^{+}88, PW85, RS88, SR89, SBS87a, SWZ85, SMS^{+}85, THM86, VC85, VB88b, WG87b, WW87b, ZSSR86]. Activation-dependent [BMBL89]. activation-independent [PON^{+}88]. activator [BVD87, CSM^{+}89, COMS87, FHS^{+}86, GKS87a, HB88a, HBB^{+}87, KOBLM88, KNGH^{+}85, LSAKO86, LS87a, MVS88, PVMO87, PSS^{+}87, PHN^{+}88, PMR89, SMR87, SR88, SVG^{+}86, VBB85, VS89]. activator-inducing [PMR89]. Activators [BC89, DF87, DMW88, GDC^{+}86, ITY^{+}88, NRTE89, SHBV89]. Active [SCP85, AHM89, AWH^{+}85, DVVT88, ET87a, EKH89, FHE88, RH86a, Sal86, SEGS85, UF89, WDR89]. actively [LFS^{+}89, TACQ89]. activities [CSK^{+}88, Eva89, LBO^{+}85, PML^{+}89, SR88, SHGD85, WMSB^{+}89]. Activity [ACN^{+}89, AETT86, AR88, AC86b, BPC89, BSd^{+}88, BJMK89, CCC89, CMB^{+}87, CWZ86, CB86, CMP86, CCBKJ89, CB88, CS88d, CG86c, DAJJ^{+}88, DNSMA88, FKC87, FBCQPR87, GHW^{+}88, GSDR89, GTG85, GBB89, GMLR85, HWDC89, HCKJJ88, HW86b, JWH85, KOBLM88, Kol86, LCD^{+}89, MFB^{+}89, MSLR^{+}86, MHWB86a, MW88b, MEVM89, MVB87, MTA86, PG86, PCBD89, PMV^{+}89, Pra86, RCC^{+}88, SMR87, SB85c, SWZ85, SMA85, SE86]. actomyosin [IKH87]. acts [IS88, LCDRB89]. acute [BHSJ86, DWL86]. acute-phase [BHSJ86]. acyl [CCC89, ISL87, SL87b]. acyl-coenzyme [CCC89]. acylation [SV89]. acyltransferase [CCC89]. adaptation [DH89b, Van87]. adaptins [Rob89]. added [TIS^{+}86]. Addition [GHW^{+}88, HGS89, BM86, MLR^{+}86, Rap89, TTW88, WPBB^{+}85]. addressin [SRB88]. Adducin [KOB89, MMM87]. adenine [ACPFB^{+}85]. adenosine [DKHB87, MEA89, MM89, Pip88]. adenovirus [Ros89, SP85]. adenylate [GDC^{+}86, UMHI85, Van87]. adhere [BPL87]. adherence [CMB^{+}86, MS88b, NZ85, SDPZ89]. adherens [GVV85, GGJ^{+}87, OBH87, TT89a, THT89, TIT89, VGKF86, VG86a, VG86b]. adherens-type [GGJ^{+}87]. adherent [GBC^{+}86]. adherentes [Tok89]. adhering [CC85b]. adheron [TSK88]. adheron-containing [TSK88]. adherons [SL85c, SL85b]. Adhesion [DDH^{+}87, LEW^{+}88, AR88, AYCY89, AEB^{+}88, ABME88, APS87, AGPT88, Bec86, BMBL89, BMVB89, Bis86, BVK85, BSPdSG88, BRS87, BF85, CSK^{+}88, CGA89, CCM86, CTS^{+}87, CG89, CE85a, CE85b, CCE87, CSG85, CG86a, CLC^{+}86, DLG^{+}86a, DJ89, DCL^{+}87, DCC^{+}88, DLG^{+}89, DRC^{+}86, DS88, FGE86, FHE88, GR89, GCT86, GDvdM87, Gri86, GE88, GSG88, HR87, HNNT88, HFG87, HMTF88, HNH^{+}87, HFC^{+}86, HAK^{+}86, HAK^{+}88, JS88b, JAT86, KWLS88, KMR^{+}88, Knu85, KSM89, KFS87, KJF^{+}88, LCE87, LCR^{+}87, LHS^{+}89, LBEM89, MS86b, MS88a, MHF86, MP86a, MTK^{+}87, MHP^{+}86, MOP^{+}86, NT86, NAD^{+}85, PS87, PSS89, PNB89, PJW88, PRB^{+}89, PSGS85, PSK^{+}87, RDPR^{+}86, RVS88, RWS^{+}85, SL85c, SLE86, SS88a, SPS89a, SPS89b, SYT^{+}89, SCC87, SKC^{+}86, SRW^{+}89, SR87, TDGE85, TSK88, WDV87, WC87, WCPK88, WGPH^{+}89]. adhesion [vdPdPR86]. adhesion- [MHF86]. adhesion-mediating [SL85c]. adhesion-promoting [CSK^{+}88]. adhesions [MUH89, WRL86]. Adhesive [FCZ^{+}89, Col86, Fra87, GSS89, Gla85a, Gla85b, LH86, MLC^{+}88, RMP^{+}88, SSC86, TSK88]. Adhesomes [SSKK89]. Adipocyte [HSS85, AR85, CKR85a, NR88, NR89a, SDR^{+}88, TTLR89]. adipocytes [BGL^{+}88, JB86, TL89, CHS85]. adligin [AHC89]. administration [PAGPLM86]. adopt [PHL89]. ADP [AW89, Pol86]. ADP-actin [Pol86]. ADP-ribosylation [AW89]. adrenal [AETT86, CJO^{+}89, HB88b, HB88c, MKA^{+}86, PH89, ST86, vGRB86]. adrenergic [SFMT87]. adrenocorticotropin [RRB86, TB87]. adult [AMWN89, BPMD87, Bis86, BVK85, CE85b, CS86d, EPJBBW86, FCVDH89, FBE85, FS86a, GSB85, JSC85, JJ87, MCS85a, MS88a, RDB^{+}86, SSC86, TFS^{+}89, ZSB86]. advances [SRWdC87]. AEV [WBVL86]. AEV- [WBVL86]. affect [DRSG85, LBP^{+}86, RG86a]. affected [AY89]. affecting [KP85, RdLB87]. affects [BF85, BBH87, CKR^{+}85b, Col86, DH87, WTBBW85]. affinity [DAJJ^{+}88, DBM^{+}89, GRCG86, HSMM85, Kee87, Mas85, MFA89, PZBD87]. African [GWI^{+}87]. after [Abr85, AMWN89, BMVB89, BRR^{+}87, EMT88, GORW88, HMN87, JPH85, LS87c, Mac89a, MNC86, MG85, NH87, NZ85, NNJ^{+}87, PJFB89, PW86, SDDZ^{+}89, SR85c, STGK87, SMS^{+}85, SGS87, UPRS88, VJF^{+}88, WS86a, WIS89, WS85, WS86b, WHI86]. against [AIJ^{+}89, AB89, BKPG89, BFL88, BP87a, CRR87b, CBL85, DSTJ85, FDG88, GDO^{+}85, GAH86, HWSK89, HMdN^{+}88, HCKJJ88, ISM85, MBKR89, MVSW86, MAG88, PSP85, RWP^{+}89, RDB^{+}86, SRT^{+}86, TJ85, TLPTV85, WFW87, WKKS^{+}87, WLJM^{+}89]. agarose [KHKR86]. agent [MA86]. Agents [EMZ85, MA86, SDDZ^{+}89, YSW^{+}85]. agglutinability [Goo89]. agglutinin [COA85, GACOH85, HMHWS86, KdWB^{+}89, LWM85, MS88c]. agglutinin-associated [KdWB^{+}89]. agglutinin-binding [MS88c]. aggregate [LBO^{+}85]. aggregates [GW86, LP85b, MHWB86b, SW87b]. Aggregating [Wal86, GDM^{+}88]. Aggregation [DWO^{+}87, GBU^{+}86, LS89, SAA89, SF88b, SKC85a, Wal88]. aging [Wan85b]. agonist [IMP^{+}87]. agonist-induced [IMP^{+}87]. Agorins [AM86]. Agrin [FG89, RMM87, MSM88, NSM^{+}87, Wal88]. agrin-induced [Wal88]. Agrin-like [RMM87, MSM88]. Agrin-related [FG89]. airway [Gor85]. alanine [DCWJ89]. alanine/glyoxylate [DCWJ89]. albumin [GFA^{+}88, GFSS86, MWSP87, PBT^{+}86, RG86b]. albumin-binding [GFA^{+}88]. Alcohol [DVVT88]. Aldolase [PT88, Cla87]. aldosterone [VSZ^{+}87]. alga [HW85b]. algal [JIFH85, KR88]. alignment [MJW86]. alkaline [AC86b, NYRK87, dBBB^{+}86]. alkylation [NB89b]. all-or-nothing [Mac89a]. allele [CF88a, WVCR89]. Allium [Pal87]. Allosteric [Sch86b]. allow [HH86]. allows [VS88a]. along [AWH^{+}85, BNB^{+}88, GB87b, GSB87, HF85b, KBR87, KR88, KS88, KS86, LMSM89, ML85, NL86, RRF87, SM89a, SSRS88, TB87, VMH^{+}88, YMM88]. Alpha [GMLR85, HTRBP86, MHK^{+}85, SW87b, AGPT88, BGH^{+}87a, BTO88, BL85, BRP88, BDB^{+}87, BHCD^{+}89, CDB87, CC85a, CL86, CPG^{+}89, DD87, DF86b, DBS^{+}89, FKC87, FC89a, GW88a, GGR89, GKR87, GGN^{+}86, GZHM87, GLC88, GB86b, GKB87, HZS88, HERO^{+}87, HWSK89, HKPS89, LR85b, LRF88, LGW^{+}89, LCBS89, LP86, LL88, LSF89, LC88b, MGR86, Mas85, MB88, MFB^{+}89, MA85, MMW85b, MJW86, MW86, MND^{+}87, MCA^{+}89, MTA86, NL85, NRL^{+}89b, OLGT86, PL85a, PAHC89, PF85b, PLC87, PBT^{+}86, PRS^{+}88, RPW^{+}89, Rob89, RS88, STH^{+}89, SMPS86, SGBG87, SL89, SCBR85, SC86, SSS^{+}87a, SW87a, SDDZ^{+}89, SSS^{+}87b, SBS87a, SRT^{+}86, SMS^{+}85, SQM89, SW88, SS89d, SHSvdM89, SBC^{+}87, TH89a, TM87b, TDR88, WR88, WFW87, WSS^{+}89, WC87, WSLN86, WB85, WK85, YMM86, TSOC89]. alpha- [LC88b, SC86, SW87a]. alpha-actin [RS88]. Alpha-actinin [MHK^{+}85, BL85, CC85a, DBS^{+}89, HWSK89, MMW85b, MJW86, MW86, SMPS86, SW88, TM87b, WSLN86, YMM86]. alpha-actinin- [DF86b]. Alpha-actinin-containing [SW87b]. alpha-actinins [NRL^{+}89b, SCBR85]. alpha-adaptins [Rob89]. alpha-cardiac [HZS88]. alpha-enolase [WLW^{+}88]. alpha-fetoprotein [PBT^{+}86]. alpha-fodrin [GZHM87]. alpha-globin [SS89d]. alpha-granule [SMS^{+}85]. alpha-helical [SQM89]. alpha-human [WB85]. alpha-Internexin [PL85a]. alpha-isoactin [OLGT86, WR88]. alpha-latrotoxin [SDDZ^{+}89]. alpha-mannosidase [LSF89, WK85]. alpha-skeletal [HZS88]. alpha-smooth [SRT^{+}86]. alpha-spectrin [DBS^{+}89, WSS^{+}89]. alpha-spectrins [DBS^{+}89]. alpha-subunit [BGH^{+}87a, BDB^{+}87, FKC87, FC89a, GKR87, HKPS89, SBC^{+}87]. Alpha-thrombin [HTRBP86]. alpha-toxin [MA85]. Alpha-tubulin [GMLR85, CDB87, CL86, DD87, GGN^{+}86, GB86b, GKB87, LR85b, LRF88, LP86, MGR86, MND^{+}87, PF85b, PLC87, SGBG87, SSS^{+}87a, SSS^{+}87b, WFW87]. alpha-tubulins [GLC88]. alpha-vascular [SL89]. alpha/beta [TDR88]. Alphavirus [FKH88]. alphoid [MMM^{+}89]. also [NRL^{+}89b]. alter [EG89, VES85]. Alteration [RG86a, WSRT89, EOM86, MG85, RC85b, RWS^{+}85]. Alterations [BF85, BB88b, SSW^{+}85, HP85a, HTGP85, KH85b, ODD85, SC86, TT88b, TT89b]. Altered [DLG^{+}86a, RR88, POCM88, VB88b, WDV87]. alternative [BKA^{+}89, BCC^{+}87, CBS^{+}89, MOP^{+}86, PRB^{+}89, WGPH^{+}89]. Alternatively [MHP^{+}86, HAK^{+}86, HAK^{+}88, SSW89]. alters [HBC89]. alveolar [KWW85]. Alzheimer [RPCB88, WCSR85]. ameba [KS86]. amebae [CC85a, MLSW86, SW87a, VES85]. amebas [TRM88]. ameboid [GB85b, HWDC89]. amines [SDZB^{+}85]. Amino [ASA^{+}87, AGPT88, CHH^{+}88, DCB^{+}88, HSWH86, BDS87, BCMK87, BHCD^{+}89, DH87, EVG^{+}86, GDO^{+}85, GKZ89, LWSF85, LD86, MLM85, PL85b, RMP^{+}88, SHC87, SSK89]. Amino-terminal [HSWH86, BDS87, BHCD^{+}89, MLM85, RMP^{+}88]. Aminobutyric [TCK87]. aminopeptidase [GML^{+}89, RPRS89]. aminotransferase [DCWJ89, STS^{+}89]. amitotic [RSR^{+}88]. ammonia [BNS85]. amnion [RFV85]. amniotic [MTRR89]. amoeboid [SRR89]. among [GLC88, SW87a, vZC87]. amount [CPM86, vDSP^{+}88]. amounts [WBDR88, YSW^{+}85]. AMP [AR85, CWH86b, ECZ86, Eva89, Goo89, HJS^{+}85, HB88b, KSKW^{+}85, PG87, RRB86]. AMP- [CWH86b]. AMP-dependent [HB88b, KSKW^{+}85, RRB86]. AMP-independent [Eva89]. AMP-mediated [RRB86]. Amphibian [ASZ^{+}88, PJ87, DF86a, MABJ^{+}86, RBAF89]. amplification [MSC^{+}85, TASK86]. amplification-associated [MSC^{+}85]. amplified [HFF^{+}85, TASK86, Wri85]. amputated [BH86a]. amylase [LMWG85, MBKO86]. amyloid [RPCB88]. analogue [LPLT85, PSM89]. analogues [DWFT88, GMB89, HdAC89, JB86, KP89, LP85a, VRLG89]. Analyses [BS87, KPK86, PBB86]. Analysis [AYCY89, CWW89, DSG86, GM85b, IHJ88, JS86c, MACO88, MR85a, WS86a, ZV89b, AJRNBJ85, BB87a, BCC^{+}87, BC85, BFK87, BDB^{+}87, CM86a, CH87, CHLS88, CHS85, DCB89, DDR^{+}89, FMG89, FCH^{+}87, GFHDD89, GW89, GBM89, HDB88, HMHWS86, HKPS89, HGL88, JYG86, KLS85, KTT^{+}86, KHC^{+}86, KRLU87, KYS^{+}88, LC88a, LMW^{+}86, LHNF87, Mac89a, MMR88, MRWJ88, MGWM85, MVSW86, MAG88, MM89, MASA89, MG86b, OH88a, PN88a, PN88b, PBT^{+}86, POD89, RMP^{+}88, SG86, TNG89, VSKSP86, WVCR89, WBB89, vGRB86, Bro85a]. Analytical [ACPFB^{+}85, MSK^{+}87]. analyzed [FKP86, GH85a, GML^{+}89, WOP^{+}88]. anaphase [GSB87, GSB88, GB89a, HP86, HC87, HHY88, MMC88, Nic89, SM87, SPH87, TB89b, WSRR89]. anchor [CWW89, DH87, HGL88, LD88, LCDRB89, MBH87]. anchor-minus [MBH87]. Anchorage [BMGI85, CCK86, CW87, KGRP85]. anchored [ASCM89, GCR^{+}89, HD86, KSM89, NYRK87, WW88]. anchoring [CE89, KSL^{+}87, KEG88, NFW89, SKMB86, SVSH^{+}88, TSK88]. aneuploid [BPB^{+}88, MVB85]. aneural [FG89]. angiogenesis [IF89, MTRR89, RE86]. angiogenic [DF87]. animal [DCR89, MMMP87, SR86b, dSdSH^{+}88]. animals [KPK86, Ree85, SSSP85]. Anion [CML85, CSL87, OTS86, SNSS87, WBVL86]. Anionic [Ear87]. Anionized [GS85]. anions [GTG85]. Ankyrin [MCA^{+}89, BWSB86, KB88, MNWL85, NL85, NH89]. ankyrin-like [BWSB86]. annealing [MGGP88, RGM86, RGBM87]. Annulate [CM88b]. Anomalous [Wil88]. anoxia [FA85]. antagonism [RBKK85]. antagonist [SNC^{+}86]. anterior [FZ85, HFZM87, RRB86, RFZH85]. anterograde [ML85]. anterogradely [TB87]. Anti [WW87b, BBGI85, BWS86, FSM88, HLL88, Hir86, KL89b, WKRH86]. anti-arc-1 [BBGI85]. anti-carbohydrate [BWS86]. anti-fluorescyl [KL89b]. anti-MAP2 [Hir86]. Anti-peptide [WW87b]. anti-receptor [WKRH86]. anti-T [FSM88]. anti-tropomyosin [HLL88]. Antibodies [LLCA89, MH88, SPF87, SPS89a, TLPTV85, AK89, AIJ^{+}89, AYCY89, AKC^{+}86, BK86, BRR^{+}87, BCC^{+}87, Bro85a, Bro85b, BP87a, CP87b, CRR87b, CTSF86, CCBKJ89, DSA86, DFSK87, DHB86, ETT88, EDD^{+}86, FTC^{+}89, FPS^{+}85, FH86, FONW88, GDO^{+}85, GAH86, HKKP86, HLL89, HvdSdC^{+}89, HW86a, HMdN^{+}88, HCKJJ88, HBB85, ICS88, JBdlTB87, KKSK88, KL89b, KGHK^{+}87, LC87, MVSW86, MAG88, MBSMD86, OBCO85, PG86, PSP85, PWS^{+}89, PNB89, PF85b, RMP^{+}88, Rob87, RG87, SSG87, SHSvdM89, TH89b, TTSSH85, VMT^{+}88, WGS^{+}88, WRS^{+}86, WCPK88, WLJM^{+}89, WW87b, ZG87]. antibody [AB89, BCA^{+}86, BBGI85, BWS86, BPKS86, BF85, BFL88, CWR86, CMP86, CBL85, DHL85, DBWL87, DSTJ85, FDG88, FSM88, FR86, FTC87, GVGL85, HNR87, HLL88, Ip88, ISM85, IHJ88, JNR^{+}85, KBL^{+}86, KYS^{+}88, MNC86, MBKR89, MVSW86, MAG88, RE85, RWP^{+}89, RDB^{+}86, SSJ^{+}85, SFB^{+}88, SDMF85, SRT^{+}86, SKC^{+}86, TJ85, THFB87, TES^{+}87, WFW87, WFR86, WKKS^{+}87, WKRH86, ZRW85, VG86b]. Antibody-mediated [VG86b]. anticoagulant [SBHN86]. antigen [AYY86, BKA^{+}89, BGM^{+}87, BBGI85, BSLS^{+}89, BMB87, BSDH86, CVWZ89, CBYK88, CPM86, DKB^{+}85, DW89, DWT^{+}89, DHL85, DAJJ^{+}88, DRD^{+}88, DS88, ELKH88, FSM88, GAH86, HKH^{+}88, HDG^{+}85, JBdlTB87, KFHF85, MSQ^{+}89, MMM^{+}89, MGSG88, MHD^{+}87, OOCT89, PF85b, PHM85, RR87, SLG88, TYP^{+}87, TB88a, TLPTV85, WWN87, WHL^{+}86, YBSR87]. antigen-1 [DS88, RR87]. antigen-induced [BGM^{+}87]. antigen/cyclin [OOCT89]. Antigenic [GDD88, Bro85a, CM88b, GPJ88, HB88c, SSV^{+}85, SKK^{+}85, WRS^{+}86]. antigens [BKA^{+}89, BPMD87, BHB^{+}87, CHE87, FLM86, GK86, KHB88, LC87, MLPF89, MBLS87, NHH^{+}86, PW86, SS87, SP85, TBK86, WE86a, WSH85]. antiluciferase [NNJ^{+}87]. Antimalarials [KSG85]. antimitotic [MMY88]. antiparallel [PG86, SQM89]. antipeptide [RMP^{+}88]. antiport [OTS86, RT86]. antiproliferative [FRR85]. antisense [HW85a, RPW^{+}89]. Antisera [BKPG89, GLC88, HWSK89, SKDF88]. antiviral [GBB89]. aortas [FBCQPR87]. aortic [BTO88, CPC^{+}89, HBC89, KW86a, PMR89, ROCK86, WF86, YBWLM88]. aphidicolin [RG88]. apical [AFMD89, BFSH87, EKH89, KBL^{+}86, KDB87, LCDRB89, MG86b, OS88a, OE85, PPB^{+}89, RIPS85, UPL^{+}87, VSSGRB87, VSSRB87, VSSRB88, ZMH89]. apical-basal [MG86b, ZMH89]. apically [GM85c, MBH87]. Aplysia [EBGS86, GB86a, KSS86, SS85, SEG^{+}86]. apolipoprotein [BMR85, LFS^{+}89, MCG^{+}88a]. apparati [CJH85]. apparatus [AK86, BSR^{+}89, BB89, BCD86, DK88, GB85a, HHS^{+}85, KCAF89, LP85a, LPBW87, NBS88a, NM89, PSM89, PJFB89, SFB^{+}88, SDS87, TPBB85, VLD^{+}85, WMM86, YMT85]. Apparent [PMJ87]. apparently [GLD^{+}88]. appear [LLRR87, MLSW86, NRL^{+}89b]. appearance [AE88, PFS85, SCM87, TDGE85, VSSRB87, WS85]. apposed [TCK87]. apposition [GTB85]. approach [BSA87, GL85]. approximately [HLU^{+}87]. AR42J [LMWG85]. arabinogalactan [PKS^{+}89]. arachidonic [GMCH87, NMKS85]. Arbacia [WBGV85]. arc [BBGI85]. Architecture [WW88, BB88b, HS86a, HBV85, HTH85, HSK^{+}89a, HS87a, SCOE^{+}86, WHMA86]. area [DMD^{+}86]. areas [Ran88]. ARG [LWH88, BLK^{+}88, CPP^{+}87, DRP87, DAS^{+}87, DCC^{+}88, DLG^{+}89, GAPR88, GGA^{+}89, SKS^{+}87]. arginine [HFF^{+}85]. arise [KTF87a]. arises [CSFS89]. arm [AIJ^{+}89, CTSF86, KW88, NMM^{+}86]. arms [AIJ^{+}89, FS87, GH85b, Kam88, MR85b, OK86, SGH85]. Arrangement [DCB89, LCK88, UTK88, VS85, LP86, SDMF85]. arrangements [Kub87]. array [GKB89]. arrays [FCH^{+}88, Gor85, SFW87]. arrest [Ibr87b, PAF88, RGW^{+}89, SW85, WW89]. arrested [MHW85]. Arterial [CCK^{+}88, SHS^{+}88, HBF^{+}88]. arteries [MBM85]. articular [FCM85]. articulating [DB85]. artificial [VS88b]. arylsulfatase [dHTG88]. Ascaris [SRR89]. ascites [KWPH89]. ascorbic [HKPS89, KSSP86, EBBW87]. asexual [HLU^{+}87]. asialo [SR85c]. asialo-transferrin [SR85c]. asialoglycoprotein [AK89, ALL89, BLSL88, SFJU86, SGS87, SGG^{+}89, SDZB^{+}85, ZBSS^{+}87]. asialoglycoproteins [MH86b, WH85]. Asp [BLK^{+}88, CPP^{+}87, DAS^{+}87, GAPR88, GGA^{+}89, SKS^{+}87, LWH88, DRP87, DCC^{+}88, DLG^{+}89]. Asp-dependent [DLG^{+}89]. asparagine [FCCC^{+}89, GB85a, MB88, REB^{+}86]. asparagine-linked [FCCC^{+}89, GB85a, MB88, REB^{+}86]. aspects [BBGI85, ETA85, HS87a, IKH87]. Aspergillus [MGWM85, May89, MKH88, OR85, OMM87, WMGM85]. assay [BCA^{+}86, GGH89, GS86b, WE86b]. assemble [MMBA89, SSP89]. Assembled [GHL85, DVVT88, ZWE89]. assembles [SL88]. assemblies [JP85]. assembling [ATNH86]. Assembly [AST89, CHK^{+}86, CDB^{+}86, HBKK87, LL86, SCM87, SL86b, TL87a, WGBB87, WKS85, ASM^{+}87, AWJ85, AHM89, AYCY89, AF89, AVP^{+}89, ALWL89, AKVF88, BB87a, BRKE88, BS88b, BHAS85, BB87c, BFKF86, BSLS^{+}89, BP87b, CWR86, CTSF86, CA86, CF88a, CV87b, CMR^{+}87, CSL87, DKHB87, DFSK85, EBBW87, EHK86, FMJS89, FOW89, GK87a, GK87b, GB87b, GCK^{+}88, GHL^{+}89, HP87a, HK88a, Hol89b, Ip88, JFP89, JG85, Kee87, KCN^{+}89, LDA86, LHS^{+}89, LFHD89, MGR86, MB88, MSC^{+}89, MLM85, MLE87, NL85, OBH87, PG86, PN88a, PN88b, PN89, PHRM87, PBH^{+}87, PHR86, SCK^{+}89, SL86a, SMS^{+}88, SP89, SSP89, SB85e, SWTB88, SBH^{+}88, TI85, TC86, UMHI85, VB88a, WS87, dVOD^{+}89]. assembly-promoting [DFSK85, JFP89]. assisted [AJRNBJ85]. associate [BLSL88, GWM85, HBKK87]. associated [AC86a, AG87, ALCG88, AK86, BS88a, BF86, BS89, BM89, BWSB86, BCD86, BR85, BKR86, BM88, BJMK89, BSG85, CBS^{+}89, CFC89, CMM^{+}85, CK86, CRLW86, CHH^{+}88, CB88, CFB^{+}85, CS88d, DLH^{+}89, DNSMA88, DS88, EBCO88, FL88, FSM88, FLH^{+}88, GK87b, GM88, GS86a, GDM^{+}88, GLM86, GBBS89, GDC^{+}86, HvdSdC^{+}89, HBV85, HTH85, Hir86, HH87, HC86, HW85b, HML88, HWLS88, HABM85, JRSB87, JS86c, JPM85, JFP89, KFA89, KMM88, KTT^{+}86, KdWB^{+}89, KF89b, KRN^{+}86, LU87, LFR85, LR85a, LCBS89, LWM^{+}88, LYG85, LP85b, MTCE87, MCS85a, MLPF89, MS88a, MP86b, MBKO86, MOT^{+}85, MSC^{+}85, MMS89, MCC^{+}89, MKA^{+}86, MCCM88, MFM89, NPCL85, NR88, NB88a, NLC89, NT86, OHH^{+}89, PBBP85, PKS^{+}89, PSD^{+}85, PHN^{+}88, PSK^{+}87, RS85, RSF^{+}85, RR87, RR86, SVF^{+}89, SHC87, SNNK85, SS87, SFB^{+}88]. associated [SG88a, SRR89, SSY^{+}86, SLL88, SH87b, SLL89, SSMG86, TPBB85, Tra88, TACQ89, TTKM86, WBD^{+}86, YLGG85, YSW^{+}85]. associates [BF88, RJB86, TDR^{+}87b]. Association [HCS89, HCHK89, JLK89, JMW88, LWR85, LS87a, MABJ^{+}86, OMS89, PF85a, SLB^{+}89, BSG89, BHK86, BDS85, BMB87, Cha85, CWF^{+}89, DBK87, Gla85a, HFG87, ITY^{+}88, KOB89, KdS86, KWH89, KRN^{+}86, LN88, MWM89, MBK89, NYG87, NRL89a, OPN^{+}85, SE86, WWW^{+}85, WRR87, dSdSH^{+}88]. associations [CM86b, MKD^{+}89, PBBP85, YR87]. Aster [SFD^{+}86b, RDJ^{+}86]. Aster-free [SFD^{+}86b]. asters [SB85a]. astral [DCR89]. astrocyte [CFB^{+}85, FCMK^{+}86]. Astrocytes [TBK86, Fal85, NTLR88, RBAF89]. astroglia [GE88]. astroglial [ELKH88, FSFL88, GB85b, Hat85, Hat87, SFMT87]. Astrotactin [ELKH88]. Asymmetric [WIS89, NPRR86]. asymmetry [BN85a, Sal86]. Asynchronous [KCN^{+}89, ASLFP89, SMB^{+}88, WBVL86]. ATP [FC89b, GS86a, HC86, ISL87, MK85b, NT85, Pol86, PMV^{+}89, SM89b, SPH87, VSKSP86, WB86]. ATP- [Pol86]. ATP-binding [GS86a]. ATP-dependent [MK85b, PMV^{+}89, SM89b]. ATP-reactivated [NT85]. ATP-sensitive [HC86]. ATPase [AMWN89, BGH87b, CB88, DMF^{+}87, Den88, EVG^{+}86, GKR87, GTG85, HERO^{+}87, HS89, HHES88, JY86, MSQ^{+}89, MTG88, MSS87c, MCA^{+}89, NH89, PG86, PSV87, Pra86, RYR^{+}89, SPV88, SBC^{+}87, VSZ^{+}87]. ATPases [vAAA86]. atria [KF89a]. atrial [JSC85]. AtT [MBLK88]. AtT-20 [MBLK88]. AtT20 [Too85, TT86b, TB87, THFB87, TTF87]. Attachment [AT86b, CMNB88, APBP87, AAM^{+}87, BR89, BLK^{+}88, CWW89, CTS^{+}87, CK86, CPHM86, DKB^{+}85, FW85, GB87a, GB87b, HNR87, HDG^{+}85, LWH88, RSG87, RS85, Rog87a, SGRM85, SKS^{+}87, SV89, SCD88, TDR87a, WGPH^{+}89, YBWLM88]. attachments [MEA89]. aureus [MA85]. authentic [VS88a]. auto [CBL85]. auto-antibody [CBL85]. autoantibodies [JYG86]. autoantigen [ESM^{+}87]. Autocrine [HD88b, SR88, SC88]. autofluorescence [JIFH85]. autoimmune [HKH^{+}88]. autologous [HKH^{+}88]. autolysosomes [MYT87]. Autonomous [ZWE89, NMPG88, SS85]. autophosphorylation [LCD^{+}89]. Autoradiographic [BW86, BD86a, SMH88, BD86b]. autoradiography [CBIO85, KBB^{+}88, NMKS85]. Autoregulation [CJK85]. autoregulatory [PC85]. auxiliary [CSBS89]. available [LC87]. Avena [MP86c]. average [LGF86, MSW^{+}86]. avian [AYY86, BFBL88, CHMB86, DDH^{+}87, GCC^{+}89, GB87b, HSMM85, HMS^{+}89, KDBF86, LFS^{+}89, LHNF87, LGF86, MS86c, RS88, SL85a, SGM^{+}89]. axes [HW87]. Axial [SDMF85]. axis [Sal86]. Axokinin [THM86]. axolemma [TTKM86, VJF^{+}88]. axon [BBB89b, EBBW87, GB86a, HGL85, HBV85, Hir86, KTT^{+}86, TTKM86]. axon-related [BBB89b, EBBW87]. axonal [DLBH89, DFdN^{+}89, GAH86, HGL85, HBV85, HB87, Hol89b, KP85, SBK85, TJ85]. axonally [RSL^{+}89]. axonemal [GHL^{+}89, Sal86]. axonemes [HMN87, OK86, Pip88, SL85d]. axonin [RSL^{+}89]. axonin-1 [RSL^{+}89]. axons [CDB87, CRR87b, DFdN^{+}89, FL88, GB86a, HBV85, HTGP85, LN88, NL86, PBBP85]. axoplasm [AWH^{+}85, ML85, SBK85]. axoplasmic [GS86a, GS89, NB88a, Pra86]. axostyles [Heu86]. axotomy [JB87]. azacytidine [ODD85]. azide [PAF88]. azurophil [JGAR89].

B [MMM^{+}89, ADS88, CHK^{+}86, DSA86, HGL85, KHC^{+}86, LA87, LM85b, LU88, BBP88, BHB^{+}87, BD86a, BM86, BS85, BB88b, CCLC89, DW89, DSGS87, ESM^{+}87, GORW88, GB87a, GMB89, MDJ89, RH87, RPRS89, RWP^{+}89, RDB^{+}86, Rud88, SLG88, TT88b, VHO^{+}89, WSR89]. B-50 [VHO^{+}89]. b-deficient [ADS88]. B-induced [BB88b]. B1 [CSK^{+}88]. B16 [NZ85]. B2 [MG89, SALN88]. B23 [FSM88]. BA [BSY86]. baby [GBM89, NMPG88, TB85b, YLGG85]. Bacterial [RH86a, ES86, TP89]. BALB [GPSC86, HP85a, SPO85, OMY^{+}85]. Balb/3T3 [OMY^{+}85]. BALB/c [GPSC86, HP85a, SPO85]. BALB/c-3T3 [HP85a]. Balbiani [DLC88, KC85]. band [AKC^{+}86, BS89, KMKS87, MGW86, NFW89, NMM^{+}86, OO87, OOCT89, Pal87, SB85d, FW85, GCPE85]. band-associated [BS89]. banded [KSBB87]. bands [WE89, DJM^{+}89, SCBHS^{+}89]. barbed [BM86, CCMB87, THT89]. barley [CHK^{+}86, KMM88]. barrier [GS86b]. barriers [CMK87]. Basal [AM85, HW85b, MSS^{+}87a, And86, EBBW87, HML88, HWLS88, MBKO86, MOT^{+}85, MG86b, NPRR86, RNPR^{+}88, RdLB87, SFLM86, SH86, WSJ85, ZMH89]. base [ZBSS^{+}87]. base-induced [ZBSS^{+}87]. based [HMBV^{+}88, SBS88, TLZ88]. Basement [KKSK88, YR87, Abr85, AP86, AIJ^{+}89, AT86b, CTR86, Cou87, EHM86, GLK^{+}89, HvdSdC^{+}89, JPU89, JRB88, KBO^{+}88, KKML88, LHK^{+}89, LE86, MAC89b, STE88, SABVK89, SF87]. Basic [HLB^{+}88, JSCLP89, KF89a, LKF87, RKA^{+}89, DMG^{+}89, FBF^{+}89, KOF88, LOH^{+}86, MS86b, MTRR89, Mos88, MQ89, NMPG88, OMS89, OCP86, PMR89, RM89a, SMR87, SMSR88, SR88, SSK89]. basis [GB87b, Kub87, TSK88]. basolateral [BFSH87, BF88, PPB^{+}89, PFS85, RJB86, RIPS85, VSSGRB87]. basophilic [BPKS86, EMH^{+}87, LYE87, MHWB86a, MF87b, OSP^{+}85, PSD^{+}85]. BC3H1 [KST^{+}89, KSSAQC89, LOG85, TSI^{+}89, WR88]. be [BDS87, FZ85, GW88a, GORW88, HW88, HH86, HFF^{+}85, KSBB87, SCLG86, SSJ^{+}85, TCK87, VKP89]. beads [YSR87]. bearing [DSTJ85, GZO88]. become [ZMH89]. becomes [HLW^{+}89]. bee [CD89, DCB89]. before [BBBM89, DLC88, GGJ^{+}87, KM87, STGK87, THFB87]. begins [CL86, CG89, PW85]. behaves [PSSP88, SB88]. Behavior [LPLT85, DG88, FBP89, HMP89, Maj87, SM87, WIS89, WKR86]. Bending [HMN87, BN85a]. benzimidazole [DDR^{+}89]. benzo [PBS85, WB85]. bestows [CHM^{+}87]. Beta [CHM^{+}87, Maj87, ASCM89, AS86, BM89, BNT^{+}87, BHCD^{+}89, CRF^{+}88, CPG^{+}89, CMR^{+}87, CWF^{+}89, DNB88, DDR^{+}89, EVG^{+}86, FTC^{+}89, FCH^{+}87, FCH^{+}88, GF86b, GK85, HMF^{+}87, HBV85, HTB88, JYC87, JC89, KOBLM88, LSAKO86, LGW^{+}89, LL88, LC88b, LC87, LKOM88, MPT88, MH88, MSQ^{+}89, Mas85, MMY88, MSC^{+}89, MGWM85, May89, MFB^{+}89, MTG88, MGW86, MTA86, NPCL85, NL85, NV89, NRL^{+}89b, OSH^{+}86, OGYK88, PBBP85, PML^{+}89, PRS^{+}88, Rap89, RR88, RDSL89, RYF^{+}87, SMR87, SF88a, SR89, SC86, SFMT87, SW87a, SR86a, SRWdC87, TFS^{+}89, TDR88, TTLR89, TT88b, WSM^{+}87, WVLC86, WC87, WMGM85, WW87a, WMSB^{+}89, WMH^{+}88, YKCR88, YHS88]. beta- [CPG^{+}89]. beta’-actin [TT88b]. beta-adrenergic [SFMT87]. beta-crystallin [YKCR88]. beta-glucuronidase [GF86b]. beta-glucuronidase-egasyn [BNT^{+}87]. beta-heavy [SF88a]. beta’-iminodipropionitrile-intoxicated [HBV85]. beta’-iminodipropionitrile-treated [PBBP85]. beta-Internexin [NPCL85]. beta-nerve [SR86a]. beta-spectrin [BM89, NL85]. beta-subunit [EVG^{+}86]. beta-tubulin [DDR^{+}89, FCH^{+}87, GK85, HTB88, JYC87, JC89, LC88b, LC87, MMY88, MGWM85, SC86, SW87a, WVLC86, WMGM85, WW87a]. beta-tubulins [May89]. Beta-type [Maj87]. beta/alpha [MFB^{+}89]. betaglycan [ASCM89]. between [BS87, BEHJ^{+}88, BB89, BICFA88, BGF86, BBH87, BLK^{+}88, CM87a, CCK^{+}88, CHG87, CJH85, DTW^{+}85, DWS88, DCQB88, DBS^{+}89, FTLSS88, GGJ^{+}87, HLW^{+}89, HB88b, HFB87, HCG87, IF89, KSJV88, KW86b, KF89b, KP89, NRTE89, NBS88b, OSD85, OLGT86, PHL89, SSJ^{+}85, SMC89, SS88a, STF85, SGM^{+}86, SALN88, SGG^{+}89, SP86, TRM88, TI85, TTGFC85, TSS89, TL87b, UF89, VKP89, WKBR87, WRR87, YF87, YBSR87]. BHK [YLGG85]. BHK-21 [YLGG85]. Bidirectional [KS85c, AWH^{+}85, RH86b]. bifunctional [PSSP88]. bilayer [HG88]. bilayers [BGBB89]. bind [BBJG89, CB88, IF85, MCG88b, SHS^{+}88]. bindin [Gla85a, Gla85b, Gla85b]. Binding [CTYF85, OW86, PSSP88, SL86a, WMBR89, WLJM^{+}89, AYY86, AK86, AKVF88, ABM86, BSR^{+}89, BTL^{+}86, BCA^{+}86, BFG86, BS87, Bis86, BB87c, BHK86, BCD86, BGHvF87, BP87a, BH86b, Bry88, BBH87, CWR86, CCA^{+}88, CR88a, CS89, CSK^{+}88, CMP86, CHH^{+}88, CG86a, CLC^{+}86, CB87, CG86b, CG86c, CMG86, DBK87, DMB88, EB87, EHM86, EBGS86, FCCC^{+}89, FLH^{+}88, GVD^{+}87, GSDR89, GM85a, GFA^{+}88, GFSS86, GS86a, GTP87, GDvdM87, GM85b, GBU^{+}86, GE88, GLW^{+}89, HFHR88, HS86a, HCHK89, HBKK87, HS89, HVBB88, HSO88, HMHWS86, HCE88, HHMH87, HC86, HVR^{+}85, HMdN^{+}88, HML88, HWLS88, HBB^{+}87, HAK^{+}88, HBHL^{+}89, HP85b, IMP^{+}87, JRSB87, JBdlTB87, JFP89, JNA85, JJP^{+}89, KWLS88, KSH87, KSJV88, KTF87a, LH86, LCR^{+}87, LJC^{+}87, LBEM89, MvdMH^{+}88, MOU88, MS88c, MM89, MFA89, MMM87, MK85a, Miy88, MDW86]. binding [MKD^{+}89, MUM87, NSY86, NM89, NB89b, NLC89, NRL^{+}89b, OSD85, OPN^{+}85, PL85a, PG86, PZBD87, PV85, PHM85, QCKR85, RP89, RMM89, RE86, RBKK85, Rog87b, RLP^{+}87, SNCY87, SVF^{+}89, SN89, SOB^{+}89, SM89b, SLE86, SLB87, SCC87, SD87, SF88b, SB85f, SH86, TH89a, TJ85, TAHC^{+}89, TLZ88, TL87a, TT85, TJF86, UK86, VBB85, VS89, WGPW89, WW87a, WKS85, Wil88, YB87, YMM88, YBWLM88, YBSR87, ZSSR86, dBBB^{+}86]. binds [ASCM89, CRLW86, FNPF87, GST^{+}87, KB88, KRN^{+}86, LE86, OON89, SMSR88, SB88, SL88, SWTB88, WL87, YIJ88]. Biochemical [APS87, BPJBF85, LS87b, PBB86, PUW^{+}88, RC85a, RC85b, dBBB^{+}86, BOW85, CW87, CHLS88, GW89, HCH85, KPK86, LGBE89, MHK^{+}85, NTNT85, SWW85, SCOE^{+}86, TTS^{+}89, VG86a, WH85, WS86b, KTT^{+}86, PN88a]. biochemically [DWT^{+}89, NDJ^{+}89]. Biogenesis [ADS88, BFSH87, BVHF89, JHPS86, PLA86, SHP85a, SHP85b, BRKE88, CIK^{+}89, LLSRB89, MMR88, NH89, SMVS^{+}86]. biological [LWM85, WAF87, WS86c]. biology [RM89a, SRWdC87]. bioluminescence [JIFH85]. bioluminescent [BN87, Bas87, JIFH85, NNJ^{+}87]. biomatrix [LBO^{+}85]. Biosynthesis [BMR85, BFKF86, DBC^{+}87, HFG87, IKSF89, NAD^{+}85, PHR86, RFG^{+}85, SGBF87, DNB88, FOW^{+}85, GKS87a, Ibr87b, KMY88, MAG88, PHRM87, PBH^{+}87, RFH85, RPRS89, dHTG88]. Biotin [WSRR89, OH89]. biotin-labeled [OH89]. Biotin-tubulin [WSRR89]. biotinylated [BTR^{+}86, HVBB88, OH88a]. BiP [HBHL^{+}89, NRL89a]. biphasic [DFF85]. bipolar [PCP85]. bird [TT88a, TTC88]. bisphosphate [CWR86, Cla87, HMP86, MWS85]. bivalent [KHB88]. bivalents [CNL86]. bladder [BTV^{+}89]. blastocyst [SCD88]. blastomeres [HPM87]. block [CWZ89, HSWH86, HP86]. blockade [RRB86]. Blockage [MBKO86]. Blocked [Too85]. blocking [HA89]. blocks [CMP86, Has88, HHY88, LWM85, MDVG^{+}85, SKC^{+}86, VS88a]. blood [ABM86, CBCV85, DAJJ^{+}88, DRME89, GBC^{+}86, JS88b, MB85b, NYG87]. blood-derived [MB85b]. blot [HKPS89]. boar [AdS85]. bodies [BHRC85, GW87, HDP^{+}85, HW85b, HHS^{+}85, KCAF89, TACQ89]. body [EWHP87, HML88, HWLS88, KFH^{+}86, MSS^{+}87a, RdLB87, WSJ85]. body-associated [HML88, HWLS88]. bombesin [MSLR^{+}86, ZSSR86]. bonding [WMUP^{+}85]. bone [BNLC85, BKC^{+}86, CRF^{+}88, DWT^{+}89, GHA88, OJHAH89, WMH^{+}88]. bone-derived [GHA88]. border [AFMD89, BNLC85, BBH87, CCBKJ89, DD88, GG85, HVR^{+}85, HSMM85, MSQ^{+}89, MCC^{+}89, MC89b, SCM87]. borders [KCCM85]. both [AETT86, ALCG88, BS89, GST^{+}87, HKKP86, HKF^{+}87, JFAT89, LHM89, PSM89, PT88, PL85b, WMBR89, WLW^{+}88]. botulinum [BD86a, BD86b, NJAS89]. bound [BW86, CBIO85, CNHF85, EMH^{+}87, Gri86, JBG^{+}87, MHWB86b, MVSW86, Mos88, MCCM88, Oss88, PPG85, PS86, SPT^{+}89, UP86, WMV86, Wat85]. bovine [ABM86, CJO^{+}89, GM85c, HB88b, HB88c, KW86a, MGK89, MM87, Mos88, MKA^{+}86, PMR89, RFZH85, ST86, SKK^{+}85, WLGB85, vGRB86]. bp [BB88a]. Brain [GABT86, GY86, LVSC86, RZG86, EWAP85, HSK^{+}89b, LSC86, MMW85a, RMP^{+}88, RMM89, SSSP88, SKK^{+}85, SWW85, SPV88, TBK86, UF86, VB88a, WSS^{+}89, WLGB85]. brain-abundant [LSC86]. brain-derived [SKK^{+}85]. Brain-specific [LVSC86]. branched [VS88b]. break [VCM88]. Breakdown [WD86a, LM85a]. Brefeldin [DRY89]. brevin [BH86b, HB86]. bridge [CS86a]. bridged [DB85]. Bridgelike [SP86]. bridges [ML85, VS88a]. bridging [Wan85b]. brief [TB89b]. broken [SPLW89]. brucei [Cla87, DIF^{+}88, RSP^{+}89b, SDS89, SSS^{+}87a, SSS^{+}87b, WG88a]. brush [AFMD89, BBH87, CCBKJ89, DD88, GG85, HSMM85, KCCM85, MSQ^{+}89, MCC^{+}89, MC89b, SCM87]. bud [ETA85, ET87a, VBWC88]. budding [Bäc88, PKK89, Too85]. bulk [DVHC89]. bull [dCFB86]. bundle [BT86, BDCT88, HF85b, KTO^{+}89, TFD87, Wan85b]. bundles [BAD86, HNH^{+}87, KS88, TIT89, VS85]. bundling [Fec87, MMM87, PM87b, SB85d, YMM86]. burst [NSF^{+}89]. button [HH87]. button-shaped [HH87]. Buttonin [HH87]. butyrylcholinesterase [Wal86].

c [JYC87, KBPW89, VME85, BC89, BH86b, BSB^{+}86, CPC^{+}89, CKR^{+}85b, DJM^{+}89, DF87, DMW88, EMZ85, FBF^{+}89, GWI^{+}87, GLK^{+}89, HS85a, HRP87, ICP^{+}86, ITY^{+}88, JLK89, KC88, LPR^{+}89, LLRR87, LRF89, NR88, NAD^{+}85, PH89, PMR89, SBHN86, VHO^{+}89, WG86, ZSSR86, BS86, GPSC86, SPO85]. c-3T3 [HP85a]. C-bands [DJM^{+}89]. C-dependent [CPC^{+}89]. C-mediated [NR88]. C-propeptide [HRP87]. C-protein [SBHN86]. C-terminal [BH86b]. C2 [BGH^{+}87a, GZO88, GRME^{+}89, NJAS89, PJW88]. C3bi [SSKK89]. C3bi/fibrinogen [SSKK89]. C3d [LE86]. C6 [ZV89b]. Ca [MSW^{+}86, BH86b, BFJ^{+}85, BN85b, CHFA89, CWH86b, DMF^{+}87, DMGS88, FC89b, GG85, GTP87, HHMH87, HCG87, HML88, HWLS88, HB86, ITY^{+}88, KOB89, KSH87, KS88, Mac89a, MMMP87, MSS^{+}87a, MSLR^{+}86, NRTE89, NGS85, PJW88, PW85, RT86, SVF^{+}89, SS85, Sal86, SL85d, SDDZ^{+}89, SYT^{+}89, SOP85, VAD^{+}85, VGKF86, Wal88, dBBB^{+}86, vAAA86]. cables [ABC^{+}88, SSZ88]. Cachectin [SC88]. Cachectin/ [SC88]. Caco [EKH89, AVP^{+}89]. Caco-2 [EKH89, AVP^{+}89]. cadherin [HNNT88, NTLR88, NT86, SYT^{+}89, CG89]. cadherins [HNH^{+}87]. Caenorhabditis [AC86a, AHC89, AE87, DDR^{+}89, EOM86, EBCO88, FW85, HW87, Hym89, Str86, WRS^{+}86]. cage [SCFP86]. calcification [HRP87]. calcineurin [Cha85, MLK^{+}87]. calciosomes [HBL^{+}88]. Calcitonin [FKC87]. Calcium [BB89, BA87, Cha85, CB87, EB87, Hep85, Hep89, HNH^{+}87, MW88a, NT85, OBH87, OHK86, RTK^{+}86, ACN^{+}89, CJO^{+}89, CR86, DSGS87, DMW88, DF88a, ER85, Eva89, FCCC^{+}89, GAB86, HNNT88, HC87, ICP^{+}86, Knu85, KM87, KWPH89, LL89a, LH86, LWH88, LMW^{+}86, MMT85, Miy88, MF87b, OSD85, OH88c, PF87, PBH^{+}87, PBK88, RMS88, RBKK85, Rog87b, RG86b, SIR^{+}88, SNNK85, Sch85, SSY^{+}86, SG88b, SE85, SW86, TB89b, VB88b, WKS85, YKSR89, ZV89b]. calcium- [ACN^{+}89]. calcium-binding [FCCC^{+}89, LH86, RBKK85, Rog87b]. Calcium-dependent [Cha85, HNH^{+}87, HNNT88, Knu85, SNNK85, WKS85]. calcium-depleted [KWPH89]. Calcium-independent [BA87]. Calcium-induced [OBH87]. calcium-ion [Sch85]. calcium-modulated [ZV89b]. Calcium-regulated [CB87]. calcium-selective [Sch85]. Calcium-sensitive [EB87]. calcium-stimulated [Eva89]. caldesmon [BL85, DHB86, YMM88]. calf [AETT86, CPC^{+}89, JVG^{+}85, KTP^{+}89]. caliber [HTGP85]. Calmodulin [CMTC85, SRW88, BCD86, BN85a, CCA^{+}88, Cha85, CB88, CM87c, HML88, HP85b, LPLT85, MMM87, MCC^{+}89, MC89b, SW87a, SNC^{+}86, TKP^{+}88, TT85, VLD^{+}85, VB85b, ZV89b, CCA^{+}88, CB87, CB88, GCC^{+}89]. calmodulin-associated [CB88]. calmodulin-binding [BCD86, CCA^{+}88, HP85b, TT85]. calmodulin-calcineurin [Cha85].

Calmodulin-dependent [CMTC85, SS85, TKP^{+}88]. calmodulin-regulated [MMM87]. calmodulins [FCL86]. calpactin [ZG87]. Calpactins [GTP87]. Calretinin [Rog87b]. calsequestrin [FAKVM87, HBB^{+}89, JSC85]. calsequestrin-like [HBB^{+}89]. caltractin [HML88]. calvarial [WMH^{+}88]. CAM [MTK^{+}87, MHP^{+}86, CG86a, CLC^{+}86, CS86d, CMGS86, DMW88, DBD^{+}89, FR86, FBE85, GSS89, HFG87, HFC^{+}86, MS86b, MS88a, PS87, PSGS85, RGE85, RNPR^{+}88, RDB^{+}86, SSC86, VG86a, VG86b, WFR86, WGA85]. cAMP [BN88, Bro87, BA87, DMD^{+}86, DNB88, EBGS86, GAB86, GDF86, KSJV88, KLHS85, LFC^{+}88, LFFW89, LWM^{+}88, PFD^{+}89, RLJ86, THM86, Van87, VES85, WSV^{+}89, YBBS87]. cAMP-binding [EBGS86]. cAMP-dependent [Bro87, LFC^{+}88, LFFW89, LWM^{+}88, PFD^{+}89, THM86, YBBS87]. cAMP-induced [Van87]. cAMP-mediated [WSV^{+}89]. cAMP-regulated [GDF86]. can [BDS87, CC89, CNL86, FZ85, GHW^{+}88, GW88a, GORW88, GDvdM87, HW88, HSWH86, HJV^{+}89, KSBB87, KS85c, MMBA89, NHOY85, PHL89, SCLG86, TCK87, VKP89, VS88b, WB86]. canalicular [SBC^{+}87]. Candida [SL87b]. candidate [HS89]. canine [ASJ^{+}88, BAD^{+}89, BBGI85, BPP^{+}89, BKK87, BHM89a, FCM85, FGF88, FS86b, GDRB89, HM89c, JY86, JJ87, LLSRB89, MFKVK87, MCA^{+}89, MBH87, NV86, NV87, NH89, OS88a, PPB^{+}89, PN88a, PN88b, PBH^{+}87, PFS85, RIPS85, RIS87, RG86a, SGRM85, SG88b, SAGM88, UPL^{+}87, VSSGRB87, VSSRB87, ZMH89, vMSvRS87]. cannibalistic [WD86a]. cannot [AJCMR88]. Cap [CCMB87]. capabilities [HVR^{+}85]. capacitation [WHI86]. capacity [CCL^{+}86, GTN86, SR85a]. capillaries [FS86a, MBM85, SSSP85]. capillary [DF87, GFA^{+}88, GFSS86, GM85c, KKML88, MWSP87, Mos88, OD87, PVMO87, PSC^{+}89, SMR87]. capillary-like [KKML88]. capping [CCMB87, SD88, THT89]. caps [WSW85]. capture [HK88b, MK85b, MTS89]. captured [Heu86]. carbamyl [NAS^{+}87]. carbamylcholine [RRP88]. carbocyanine [HH86]. carbohydrate [BWS86, BRS87, GHW^{+}88, KKSK86, KFS87, KJF^{+}88, MS86b, MDW86, YBSR87]. carbohydrate-binding [MDW86, YBSR87]. carbohydrates [YBBS87]. Carboxy [ZR89, APF^{+}88, GKS87b, GKS88, KWH89, MMY88, SL87b, VS88b]. Carboxy-terminal [ZR89, APF^{+}88, KWH89, MMY88, SL87b, VS88b]. carboxylase [CWR86, HMP86, MWS85]. carboxypeptidase [BDS87, HW88, SRPS86]. Carcinoembryonic [BKA^{+}89]. carcinoma [AF87, BS87, BHSJ86, BTV^{+}89, KOBLM88, KWS^{+}89, LMSM89, LLRR87, MBK^{+}86, SVG^{+}86, WBW89, WBD^{+}86]. cardiac [BOGK88, DBWL87, FM86, GSTS86, HS85a, HZS88, KFHF85, KDIG89, LHS^{+}89, MACO88, MMW85b, MMW85a, TOOK^{+}85, TM87a, TM87b, Tok89, UPRS88, WGS^{+}88, WE87, WTBBW85]. cardiogenesis [RS88]. cardiomyocyte [RS88]. cardiomyocytes [ABC^{+}88]. carrier [GGH89, KSC^{+}89]. carrot [TDR^{+}87b]. carteri [EMW^{+}89]. Cartilage [MEBV^{+}89, DKH^{+}85, FCM85, GF85, GSTC87, GHA88, HRP87, KIKS87, KEG88, KKG86, KDFG89, MGHG^{+}86, SL85a, TQN^{+}87, TIS^{+}86, VMH^{+}88, dBBB^{+}86]. cartilage-derived [DKH^{+}85]. cartilage-matrix [KIKS87]. casein [DNSMA88, SDNWA87]. cat [CS88c, LD88]. catalase [BVHF89]. catalase-negative [BVHF89]. catalytic [FBF^{+}89, GKR87]. catalytically [RH86a]. catalyzes [CRC^{+}87]. catecholamines [Owe86]. Cathepsin [SLHvF85, FWP^{+}87]. cation [BNB^{+}88, CPP^{+}87, MCG88b]. cation-dependent [CPP^{+}87, MCG88b]. cation-independent [BNB^{+}88]. cationized [GS85]. Caulobacter [Smi87]. cause [MA86]. caused [MSD^{+}88]. causes [BBP88, FCH^{+}87, Ibr87b, MMY88, MQ89]. caveatum [WD86a]. CCl4 [CMW87]. CD11 [CGA89, NSF^{+}89]. CD11/CD18 [CGA89, NSF^{+}89]. CD18 [CGA89, NSF^{+}89]. CD3 [ADS88]. CD8 [RH89]. CDC28 [WRR87]. CDC31 [HML88]. cDNA [BFK87, BDB^{+}87, DNM86, DMG^{+}89, ESM^{+}87, FTLSS88, HOPM^{+}87, HNNT88, HNP^{+}86, HML88, KTO^{+}89, LC85, LVSC86, LSC86, MGM87, NMPG88, Pau86, RKWS86, SSSA87, TGNG87, YHS88, ZN89]. cDNAs [AF87, AF89, CPG^{+}89, DBD^{+}89, KG86, Rob89]. Cell [BNLC85, BTO88, BVK85, BSPdSG88, CSG85, CM87b, DRC^{+}86, HD86, HL88, JRSB87, KGRP85, LWH88, LS85, LBEM89, MGD88, MWM89, MMRR86, MOP^{+}86, MTA86, OSS^{+}89, OH88b, QCKR85, RJB86, RBKK87, SB88, SSK^{+}88, AR88, ASM^{+}87, AYY86, AYCY89, AF87, AAM^{+}87, AE87, ABM86, AR85, BTL^{+}86, BPB86, BMB86, BWK87, BHS88, Bec86, BMVB89, BSY86, BHB^{+}87, BFA88, BSBZ88, BDS88, BPC89, BGF88, BPB^{+}88, BR89, BMB87, BR85, BF85, BJ86, BIES89, BBB89b, BCK85, BA87, CCC89, CM88a, CSH89, CVWZ89, CTR86, CL85a, CL86, CHMB86, CPC^{+}89, CRR87b, CSK^{+}88, CHH^{+}85, CCM86, CBYK88, CS87, CG89, CE85a, CE85b, CG86a, CLC^{+}86, CdCP^{+}89, CKF85, CML85, DKB^{+}85, DLG^{+}86a, DJ89, DKH^{+}85, DNB88, DRP87, DAJJ^{+}88, DLG^{+}89, DNSMA88, DF87, DMW88, DMH^{+}85, DRME89, ELKH88, EHM86, FSFL88, FBPS87]. cell [FSM88, FCM85, FGF88, FJY86, FMG89, FGA86, FFH88, FGE86, FHE88, FKM87, FZ85, FM86, GCR^{+}86, GK85, GLL^{+}86, GAPR88, GDD88, GCR^{+}89, GS85, GAB86, GTB85, GFHDD89, GBC^{+}86, GH85a, GGMH86, GDvdM87, GBU^{+}86, GRCG86, GHA88, GE88, GRME^{+}89, GSG88, HBSQ^{+}85, HAB^{+}88, HNR87, HR87, HDP^{+}85, HD88b, HP87a, HCS89, HNNT88, Hat87, HKH^{+}88, HFG87, HGS89, HHR^{+}89, HB88a, HE86, HS85b, HBB^{+}89, HMTF88, HMP86, HNH^{+}87, HMP85, HFC^{+}86, HFB87, HHD^{+}86, HDG^{+}85, HM89b, HLU^{+}87, HBB^{+}87, HSMCL87, HAK^{+}86, HAK^{+}88, HW87, IBSS85, IS88, JWH85, JASP88, JNR^{+}85, JRB88, JS88a, JFAT89, JS88b, JGYG85, KOB89, KWLS88, KWW85, KH85a, KFW^{+}89, KSJF87, Knu85, KSM89, KS85c, KF89b, KBK86, KDBF86, KFS87, KJF^{+}88, KYS^{+}88, LRF88, LMSM89, LOG85, LTG85, LEW^{+}88, LM87, LCE87, LHFB88]. cell [LN88, LLSRB89, LM85a, LR89, MHK^{+}85, Mac89a, MLR^{+}86, Maj87, MS86b, MS88a, MSY^{+}87, MvdMH^{+}88, MBKO86, MOU88, MAG88, MMK^{+}89, MCS85b, MS86c, MG86a, MASA89, MW88b, MTS89, MK85c, MTK^{+}87, MB85b, MKD^{+}89, MEVM89, MIHS86, MBvS^{+}89, MHP^{+}86, MB87, NHOY85, NB89a, NB88a, NV87, NH89, NMPG88, NRL89a, NMM^{+}86, NL86, NT86, NGI^{+}89, NAD^{+}85, OS88a, OOCT89, ODD85, OMY^{+}85, OH88c, OD87, Oss88, PH89, PPB^{+}89, PN88a, PN88b, PJ87, PS87, PSS89, PM87a, PNB89, PRB^{+}89, PFPM86, PSGS85, PSK^{+}87, Qua85a, Rap89, RR88, RBG85, REB^{+}86, RBGF89, RDPR^{+}86, RM89a, RIS87, RSF86, RFMR86, RLP^{+}87, RF87, RLB^{+}89, RC85a, RC85b, RSF^{+}85, RRP88, RdLB87, RBAF89, RMS86, RVS88, RWS^{+}85, ST86, SGRM85, SC85a, SVF^{+}89, SMSR88, SMVS^{+}86, SVSH^{+}88, SME^{+}89, SR88, SR89, SMC89, SKK^{+}85, SBD^{+}89]. cell [SL85b, SK85, SPS89a, SPS89b, SK88, SP85, SLL88, SMS^{+}88, SSS^{+}87b, SYT^{+}89, SG88b, SNC^{+}86, SKS^{+}87, SCC87, SR85c, Sny89, SYH^{+}86, SMB^{+}88, SGP88, SF87, SG86, SPO85, SCD88, SE86, SHP85b, TRL87, TDL^{+}85, TDGE85, TT88a, TTBG87, TRB86, TDR87a, TDR88, TB88a, TT88b, TDR^{+}87b, TNG89, TSS89, TT89a, THT89, TSM87, UPL^{+}87, VSSGRB87, VSSRB87, VSSRB88, WDV87, WD86b, Wan85b, Wan85c, WB86, WC87, WCPK88, WGPH^{+}89, WBW87, WM88, WE86a, WGA85, WSH85, WS89, WRR87, WDR89, WSM86, WE86b, WMC86, WMH^{+}88, WJSB88, YBWLM88, ZV89b, ZH86, dSdSH^{+}88, vAAA86, vdPdPR86, CG86c, BSLS^{+}89, RR87]. Cell-adhesion [BVK85, CE85a, CE85b]. cell-attachment [AAM^{+}87]. cell-binding [GDvdM87, HAK^{+}88, KWLS88, RLP^{+}87]. cell-derived [MEVM89, SMSR88]. cell-free [BWK87, BGF88, LM85a, LR89, MBvS^{+}89, SMS^{+}88, SG86, WB86, WE86b]. cell-glass [GTB85]. Cell-mediated [BNLC85, MTA86]. cell-neuron [SPS89b]. cell-specific [CVWZ89, Wan85c]. Cell-substrate [BSPdSG88, BPB86, Bec86, FHE88]. Cell-substratum [CSG85, CG86a]. cell-surface [HGS89, HHR^{+}89, SL85b]. cell-to-cell [AR85, MBKO86, SC85a, TT89a, THT89]. cell-type-specific [CML85]. Cells [CL85b, AK89, AJAW86, AETT86, ACN^{+}89, AF89, AAM^{+}87, AU89, ACPFB^{+}85, AMWN89, ASJ^{+}88, AVP^{+}89, AJCMR88, AHRM85, AM86, AS86, AT86b, BAD^{+}89, Bäc88, BTL^{+}86, BMB86, BCA^{+}86, BFG86, BSSMF88, BGN^{+}88, BS87, BEHJ^{+}88, BS88a, BOW85, BOW86, BHSJ86, BGM^{+}87, BBGI85, BMVB89, BRR^{+}87, BCKJW88, BBP^{+}85, BLSK88, BSK89, BF88, BPL87, BF86, BGS87, Bis86, BLSL88, BLR88, BGH^{+}87a, BTO88, BVK85, BAD86, BPP^{+}89, BPKS86, BCC^{+}87, BTV^{+}89, BCD86, BGK^{+}89, BKK87, BHM89a, BS86, BNF^{+}87, BL85, Bre89, BSd^{+}88, BKR86, BS85, BB87d, BGH87b, BM88, BP87b, BK85, BRP88, BJMK89, BLK^{+}88, CSM^{+}89, COMS87, CMB^{+}87, CRT87, CE89, CRR87a, CR87, CWOK87, CTJG88, CPS88, CSR^{+}86, CCL^{+}86, CPC^{+}89, CW88, CJO^{+}89, CCM86, CTS^{+}87, CK86, CPHM86, CPP^{+}87, CKR^{+}85b, CBL85, CCK^{+}88, CHG87]. cells [CSD89, CS86c, CBS^{+}87, CLF^{+}88, CMR^{+}87, CML85, CDTS85, DG89, DKB^{+}85, DWL86, DTW^{+}85, DHL85, DBE^{+}87, DDB^{+}88, DAS^{+}87, DCL^{+}87, DCC^{+}88, DCR89, DSTJ85, DBD^{+}89, DSHW85, DF86b, DFSK85, DK86, DKKK88, DNI^{+}88, DF88a, DS88, ES89, EKH89, EBBW87, EMH^{+}87, Eva89, EWHP87, Fal85, FGW85, FC86, FCZ^{+}89, FGA86, Fra87, FGE86, FRR85, FS86a, FS86b, GKS87a, GGN^{+}86, GFA^{+}88, GWCW85, GCT86, GBH87, GY86, GF88, GJD^{+}88, GZ89, GFS^{+}89, GBC^{+}86, GTN86, GDF86, GSB88, GSF^{+}88, GGR^{+}86, GDRB89, GPSC86, GBM89, GHA88, GLC88, HAB^{+}88, HdAC89, HWDC89, HBC89, HTRBP86, HSS85, HFZM87, Has88, HSB^{+}85, HNP^{+}86, HPR85, HBF^{+}88, HLL88, HJS^{+}85, Hep85, HC87, HvdSdC^{+}89, HP85a, HKBT87, HB88b, HB88c, HVL87, HGMS88, HSK^{+}89b, HMHWS86, HVR^{+}85, HKPS89, HSS89, HHD^{+}86, HFF^{+}85, HCG87]. cells [HM89c, HMS^{+}89, ICP^{+}86, IKSF89, ITY^{+}88, JPU89, JS86a, JNR^{+}85, JRSB87, JBdlTB87, JRB88, JY86, JMBM85, JFAT89, JS88b, JMW88, JNA85, JSC85, JSCLP89, KR88, KB88, KCAF89, KFHF85, KDHS87, KST^{+}89, KSSAQC89, KOBLM88, KLHK88, KW86a, KKGJ89, KBL^{+}86, KBB^{+}88, KDB87, KWS^{+}89, KTF87b, KF89b, KNGH^{+}85, KKML88, KWPH89, LFC^{+}88, LWR85, LBO^{+}85, LCL^{+}89, LSP86, LMD^{+}86, LGW^{+}89, LDKG87, LPR^{+}89, LLRR87, LSF89, LS87a, LC88b, LSB89, LHHC85, LL86, LHL88, LEG^{+}89, LCDRB89, LYE87, LMWG85, LL89b, LPLT85, LPBW87, LBW89, LOH^{+}86, Mac89a, MBC86, MPT88, MNC86, MP89, MLPF89, MB85a, MCB85, MCG^{+}88a, MMMP87, MACO88, MFKVK87, MSQ^{+}89, MBD89, Mat86, MBLK88, MER^{+}86, MH86a, MHF86, MGSG88, MGM87, MMW85b, MLE87, MMT85, MMWD89, MOT^{+}85, MW86, MSLR^{+}86, MHWB86a, MRD^{+}89, MVSW86, MSC^{+}85, MBK^{+}86]. cells [MSS87b, MF87b, MLK^{+}87, MNWL85, MCA^{+}89, Mos88, MQ89, MBH87, MDW86, MDVG^{+}85, MLC^{+}88, MG86b, MUM87, MUH89, MFM89, MA86, MTA86, NJD^{+}86, NWM88, NV86, NV87, NH89, NZ85, NMKS85, NMPG88, NKMK86, NV89, NR89b, NGS85, NHKE88, OSD85, OS88a, OH88a, OH89, OSP^{+}85, OSS^{+}89, OTS86, OD87, Oss88, OKLB86, OE85, Owe85, Owe86, OLGT86, OGYK88, PSBN85, PSM89, PCK^{+}87, PN88a, PN88b, PN89, PL86, PL87, PLA86, PW87, PP88a, PF85a, PHRM87, PBH^{+}87, PVMO87, PSC^{+}89, PHR86, PFS85, PSD^{+}85, PWS^{+}89, PKN^{+}89, PSK^{+}89, PLC87, PBMB85, PSS^{+}87, PS86, PMR89, QCKR85, RJB86, RR88, RMP^{+}88, RFV85, RRB86, RTK^{+}86, RJH^{+}87, RH87, RLJ86, RDSL89, RZG86, RNPR^{+}88, RIPS85, RIS87, RT88, RPW^{+}89, RSG87, Rob85, ROCK86, RKA^{+}89, RLP^{+}87, RG86a, RC85a, RR87, RHL^{+}85]. cells [RT86, RRP88, RLB85, RGPRB87, ST86, SGRM85, SMR87, SBS88, SM89a, SMPS86, SSJ^{+}85, SS89b, SNNK85, SR89, SB88, SIYT86, Sch85, SGS86, SC86, SFK^{+}89, SL87a, SK86, SK87, SFW87, SS88a, SK88, SFMT87, SDDZ^{+}89, SAA89, SG88b, SHS^{+}88, SRM89, STF85, SF89, SPLW89, SR85c, SR86b, SGM^{+}89, SYH^{+}85, SMW89, SWTB88, SBH^{+}88, SPT^{+}89, SBHN86, SAGM88, SW87b, SF88b, SS88c, SS89d, SGS87, SGG^{+}89, SVG^{+}86, SDZB^{+}85, SPO85, TS88, TKK^{+}87, TSI^{+}89, TBK86, TASK86, TYP^{+}87, TRT^{+}88, TDL^{+}85, TT86a, TFS^{+}89, TCBF^{+}85, TS86, TT88a, TTC88, TTS^{+}89, TB89b, Too85, TT86b, TB87, THFB87, TTF87, TKFH89, TB85b, TLPTV85, TAMR85, TDR^{+}87b, TNG89, TAHC^{+}89, TB88b, TRC^{+}87, UP86, VHO^{+}89, VLD^{+}85, VSSGRB87, VSSRB87, VSSRB88, VSZ^{+}87, VAD^{+}85, VSKSP86, VS85, VSR87, VB88b]. cells [Wan85a, WR88, WSW85, WC87, WKRH86, WS86b, WLJM^{+}89, WBW89, WBD^{+}86, Wil88, WRL86, WWN87, WFL87, WB85, WBVL86, WDF86, WKR86, YMM86, YMM88, YM87a, YM87b, YLGG85, YBWLM88, YMD87, ZSSR86, ZR87, ZP85, ZBSS^{+}87, ZMH89, dGRS87, vMSvRS87, vGRB86, GF86a]. Cellular [GDF86, HZS88, PBS85, PBT^{+}86, Ros89, SS89b, SL89, SHSvdM89, WS86b, ABME88, AYG85, BB87b, DPM88, DKKK88, EMW^{+}89, GH88, HCKJJ88, KHB88, LHK^{+}89, MBL89, MWM89, OGYK88, PSBN85, PMR86, SVF^{+}89, SR86b, VBB85, WSM^{+}87, WS86c]. CENP [MMM^{+}89, POW^{+}87, ESM^{+}87]. centers [DMD^{+}86, MHW85, TMB85, TPBB85]. central [BFR85, CS88b, Fal85, GDD88, GABT86, GH85b, HMN87, HAK^{+}88, MS88c, PL85a, PCS88, RLP^{+}87, TCP^{+}85, TCK87, vdPdPR86]. central-pair [HMN87]. centric [WMC86]. Centrin [SS89a, SBS88]. centrin-based [SBS88]. Centrin-mediated [SS89a]. Centriole [SR85a]. centromere [CHE87, ESM^{+}87, KAB88, LNH86, MMM^{+}89, POW^{+}87]. centromeric [MMM^{+}89]. Centrosome [Hym89, BS88a, EMK85, KFA89, MP88b]. centrosome-associated [BS88a]. centrosome-free [MP88b]. centrosomes [ES85, KTP^{+}89, SR85b, SMR86, SB85e, WG88b]. centrosphere [KB85]. cepa [Pal87]. ceramide [PSM89]. cercarial [SC85b]. Cerebellar [MBD89, CCE87, ELKH88, VS89]. cerebellum [CDB87, PS87, RBMM88]. cerebral [CS88c]. cerevisiae [HPW89, BMEG89, CSBS89, HW88, JASP88, KAB88, LHM89, MN87, MDJ89, PKK89, PMJ87, SS88b, UMHI85, WRR87]. certain [REB^{+}86]. ceruloplasmin [TKK86]. cessation [CJH85]. cGMP [BN88]. Ch [CMT^{+}88]. chain [AE87, BF86, BB87c, BHK86, BB87d, BA87, CB86, CSK^{+}88, CF88a, DBWL87, DSTJ85, DBK87, EPJBBW86, FL89, GSB85, GDS87, HBPK87, HBKK87, KKT^{+}86, KW88, KTP87, KWPH89, KRN^{+}86, LFC^{+}88, LD86, MPBZBE89, MCS85b, MTG88, NB89b, OB88, PFRS89, PBvTS88, SF88a, STH^{+}89, SL85a, SSV^{+}85, SSS85, SP89, SHS^{+}88, TOOK^{+}85, TGNG87, TB89a, TMP87, VS88b, WPBB^{+}85, ZSB86, ZG87]. chain-binding [BB87c, DBK87, HBKK87]. chain-deficient [PBvTS88]. chains [BHK86, BSLS^{+}89, CBYK88, GHW^{+}88, HBKK87, ICS88, KL89b, KKSK86, KGHK^{+}87, MHP^{+}86, PWS^{+}89, Rap89, SHSvdM89, TLRW86, Toy87b, WE87]. chamber [FMG89]. change [BBP88, Eck87, GML^{+}89, HO85, NYG87, RBKK85, WW87b]. Changes [BBB89a, BS86, CDMC88, DRH^{+}87, GZHM87, HKBT87, Heu89a, HTGP85, JS86a, LM87, RF87, SA85, TB88a, VGKF86, WLG85, WPW^{+}85, WHI86, AR88, Bre89, CC85a, CRR87a, CPKJ87, DTW^{+}85, EMT88, FJMW87, FC86, FC89b, HNR87, Has88, HB86, ICY^{+}88, JPH85, KBR87, LHK^{+}89, LHHC85, MS86a, Mac89a, MWSF87, MSC^{+}85, MTK^{+}87, OSS^{+}89, OH88c, PSD^{+}85, RMS88, SDSO86, SLL88, SKDF88, SPO85, TRM88, TS86, TTM88, VHO^{+}89, VC85, WS85, WHMA86, YMD87, dGRS87, MSLR^{+}86]. channel [ICY^{+}88, MG85, SIR^{+}88, ZKR^{+}88]. channels [AELE88, EZBW88, RT87, SC85a]. Characeae [KS88]. characean [KR88]. characteristic [BPMD87, KFH^{+}86, PCP85]. characteristics [FHS^{+}86, ZKR^{+}88]. Characterization [AK87, AYY86, ASJ^{+}88, ASLFP89, BBV^{+}89, BSK89, BHAS85, BNL89, COA85, CCS^{+}86, ET87a, GST^{+}87, GSB85, GM85c, GGH89, HKKP86, HKH^{+}88, HAM86, MW88b, MCC^{+}89, MVB87, NFN87, NNJ^{+}87, RO85, RSR^{+}88, SGRM85, SCBHS^{+}89, SL87a, SHN87, SPV88, TJ85, TES^{+}87, VLD^{+}85, VS89, WM88, WS89, YMM88, APS87, AB88, BB88a, BRKE88, BOW85, CCA^{+}88, CW87, DHB86, FL89, FCDF87, FCL86, GDD88, GLM86, GBU^{+}86, HFHR88, HOPM^{+}87, HCH85, HHS^{+}85, HWLS88, HABM85, IBSS85, ISM85, JVG^{+}85, KSEP86, KKSK88, KBKC86, KR86, KKGJ89, KG86, LBB^{+}88, LN88, LHHC85, LS87b, MMWM^{+}89, MAC89b, MGK89, MAG88, MKA^{+}86, NAD^{+}85, PH89, PL86, PCC85, RP89, RPCB88, RHD^{+}85, SGBG87, SIYT86, SBD^{+}89, SB85d, SRB88, TKP^{+}88, TGNG87, TASK86, TTS^{+}89, TMP87, TDR88, TLD^{+}89, THT89, TSM87]. characterization [TSOC89, UK86, UF86, VB85a, WH85, WN87, WGPH^{+}89, WS85, WSRT89, dCFB86, BFJ^{+}85, DBB^{+}87]. characterized [Bro85a, Bro85b]. characterizing [HMBV^{+}88]. charge [GS85, HKF^{+}87, MBC86]. charged [DH87, LG88]. chartin [AG87]. CHEF [HSS85]. Chemical [JS86d, BB87a, FMG89, NB89b, RPCB88, SAB^{+}89]. chemically [PBMB85]. chemistry [SRWdC87]. chemoattractant [JTBA89]. chemoattractants [SDPZ89]. chemotactic [BSBZ88, CWZ86, HWDC89, HO85, JBTA88, KLHS85, SIYT86, SHGD85, SWZ85, TDL^{+}85, YSW^{+}85, ZR87]. Chemotaxis [AAM^{+}87, WBGV85, ABG^{+}89, BKN^{+}85, FMG89, PRS^{+}88, SFG87, TRL87, TLZ88]. chick [AR88, ACN^{+}89, BPJBF85, CMDCC86, CB86, CCM86, CSG85, CBCV85, CMTC85, DLFE89, ETA85, ET87a, EZBW88, EZWB88, FGO^{+}85, FC89a, FBCQPR87, GF85, HFHR88, HNP^{+}86, JB87, JSCLP89, KSM89, MP88a, MJW86, MG86a, NSY86, OSD85, PCC85, PN85, RF87, RRF87, RRP88, SL85c, SL85b, SLE86, SCM87, SBS87a, SE85, TQN^{+}87, TS88, TDGE85, TTS^{+}89, TMS85, TM87a, TM87b, Tok89, UF86]. chicken [BMR85, CS86b, CHLS88, CHG^{+}86, ET87b, FS86a, GCR^{+}89, GCK^{+}88, GSB85, GCBH86, GDRB89, HS86b, HZS88, HAK^{+}88, Hut87, ISM85, KKT^{+}86, KMKS87, KDB87, KDFG89, LSEN87, LHHC85, MACO88, MOP^{+}85, MGW86, MHP^{+}86, OE85, SB85b, SL86b, SALN88, TOOK^{+}85, TB89a, ZSB86]. chickens [MVB85]. chimeric [Gor89, KYS^{+}88, MSC^{+}89, RDSG86, RGPRB87, VS88a]. chinese [LEW^{+}88, BHS88, CCC89, CSD89, CDTS85, FBPS87, GWCW85, HHR^{+}89, MGM87, MSC^{+}85, RFMR86, SC86, SFK^{+}89, SBD^{+}89, SK88, TASK86, YM87a, YM87b]. Chironomus [DLC88, KC85]. Chitin [CSBS89]. Chlamydomonas [ASM^{+}87, AWJ85, BIES89, COA85, GH85a, GACOH85, GGMH86, Goo89, GMLR85, HMP86, HMdN^{+}88, HMN87, IBSS85, IS88, KSFL89, KW88, KdWB^{+}89, LR85b, LP86, MMR88, MMWM^{+}89, MSY^{+}87, MWS85, MR85b, OK86, PG87, Pip88, SSH87, SBS88, SS89a, SDSO86, SL85d, SKC^{+}86, SEB89, WMR86, WVCR89, WSJ85, dHTG88, dVOD^{+}89]. chloride [BGN^{+}88]. Chlorogonium [HW85b]. chlorophyll [ADS88, CHK^{+}86, DSA86, KHC^{+}86, LA87, LM85b, LU88]. chloroplast [CA86, HRR89, KMM88, MT86]. chloroplasts [DJBD85, LA87, MWS85]. chloroquine [CM88a]. chloroquine-resistant [CM88a]. CHO [CSFS89]. cholecystokinin [CJ86, MLR^{+}86]. cholera [DSGS87]. cholesterol [CCC89, KS85b, LRF89]. cholesteryl [PSBN85]. cholestyramine [KPK86]. cholestyramine-treated [KPK86]. choline [AR88]. cholinergic [JPH85, SSR89, SMA85]. chondrocalcin [HRP87]. chondrocyte [BB88b, BHM^{+}89b, CDMC88, HRP87, KIKS87, VVL^{+}89]. chondrocytes [APBP87, BBP88, BB87d, CMT^{+}88, CMDCC86, HNP^{+}86, ISM85, KG85a, KG85b, RFH85, SHSvdM89, TIS^{+}86, ZRW85]. chondrogenesis [MTCE87]. chondrogenic [MTCE87]. chondroitin [CS88a, CW87, DLH^{+}89, FM86, MAC89b, PJ87, Rap89, RFH85, VVL^{+}89]. chorioallantoic [Oss88]. chorionic [MB88, MSC^{+}89, RPW^{+}89, WB85]. chromaffin [AETT86, BTL^{+}86, CJO^{+}89, HB88b, HB88c, PW86, ST86]. chromatid [LNH86]. chromatin [BK86, KAB88, MVB87, SSW^{+}85]. chromatography [MFA89]. chromatophores [MP86d]. chromogranin [RWP^{+}89]. chromosomal [GK86, HSK88]. chromosome [AY89, BSA87, DG88, EHC^{+}85, FA85, GSB88, Kub87, LM85a, NMM^{+}86, Nic89, OON89, PKK89, SFD^{+}86b, HS87b]. chromosome-induced [SFD^{+}86b]. Chromosomes [GSB87, AJRNBJ85, CM86c, CM87a, DJM^{+}89, EH85, FS88a, HMG^{+}86, Hut87, LLAS85, NG85, RDJ^{+}86, RG87, SR85b, VB85a]. Chronic [CMM^{+}85]. chronically [AC86b]. Chylamydomonas [Kam88]. Chymotrypsin [JB86]. Chymotryptic [Toy87a, Toy87b]. cilia [GH85b, MRWJ88, PF85b, SD88, VT89]. Ciliary [GG87a, SSR89, BN88, Den88, JB87, NT85, RF87, RRF87, Ste85, Toy87b]. ciliate [OOCT89]. ciliated [GG87a, KBL^{+}86]. cilium [HFB87]. circumferential [TIT89]. cis [BB88a, NVB^{+}88, SPF87, SFB^{+}88, WDF86, YBBS87]. cis-acting [BB88a]. cis-Golgi [SPF87]. cis-middle [YBBS87]. cis-side [SFB^{+}88]. cisterna [GPSM85]. cisternae [CCS^{+}86, WDF86]. Cl [SPO85]. Cl- [SPO85]. clam [SBG^{+}85, SHR86]. class [BHB^{+}87, CVWZ89, LKF87, PNB89, SP85, SBG89, WFE89]. classes [FL88]. Classification [WFR86]. Clathrin [BB87c, Bro85a, Bro85b, Kee87, SCFP86, TT86b, VB88a, GHL85, HKA^{+}87, HK88a, Heu89b, HA89, HS89, KGHK^{+}87, LDA86, MMBA89, MPBZBE89, PBvTS88, SLHvF85, TSBW86]. Clathrin-coated [TT86b, VB88a, HA89, SLHvF85]. Cleavage [MB87, ALL89, BBV^{+}89, BC89, BA86, GORW88, HTSH85, HSWH86, JRSB87, MHK^{+}85, MTG88, THFB87]. cleaved [FRG^{+}87]. cleaving [Sch85]. Clonal [ZH86, GHA88, HMBV^{+}88]. clone [LSDP86, SSSA87]. Cloned [MLR^{+}86, DNM86, HKPS89, LC85, LVSC86, LSC86]. clones [BDB^{+}87, DRH^{+}87, HOPM^{+}87, RBGF89]. Cloning [BFK87, HNNT88, KG86, Rob89, TGNG87, ESM^{+}87, HML88, Pau86, SJOB89, SYT^{+}89, WMR86, WVCR89]. clonogenicity [ODD85]. close [FONW88]. closely [ABM86, WSH85]. clostridial [AW89]. clots [KCR86]. cluster [AJCMR88, MVB85]. clustered [BM89, BBMW85]. Clustering [MHWB86a, ER87, IMP^{+}87, MK87]. clusters [ABC^{+}88, Blo86, BF87, DLFE89, FG89, MBKR89, PF85a, PB87, Pum89, RRF87]. cmd [TIS^{+}86]. cmd/cmd [TIS^{+}86]. Co [FGA86, WPBB^{+}85, BNB^{+}88, DTW^{+}85, MTA86, SR89, WRL86]. co-action [MTA86]. co-culture [DTW^{+}85, SR89]. co-distribution [BNB^{+}88]. Co-expression [FGA86, WPBB^{+}85]. co-purify [WRL86]. CoA [ISL87, SL87b, WBDR88]. coagulation [GCR^{+}86]. coassembly [GLC88, JYC87, MC89a]. Coat [MMBA89, Kee87]. coated [BPJBF85, DM87, GPSC86, HA89, HM89c, KHKR86, KB86, LDA86, LHvF^{+}87, MMBA89, MDVG^{+}85, RP86, Rob87, Rob89, RSF86, SOPvD87, SOB^{+}89, SLHvF85, SPLW89, Too85, TT86b, TH89c, VP85, VB88a, WLGB85, DF88b]. coating [BGH87b, OMS89]. coats [CV87a]. cochlea [TT88a, TTC88, TTS^{+}89]. coculture [RBGF89]. cocultured [CJO^{+}89]. code [BKA^{+}89, MHP^{+}86]. coding [GKG88, GLM86, HOPM^{+}87, KG86, NMM^{+}86, STH^{+}89, SSSA87, WVCR89]. coenzyme [CCC89, JBG^{+}87, KPK86]. coexistence [FJMW87]. coexpress [JPU89]. cohesion [CS87, WDV87]. coil [QS87, SQM89]. coiled [QS87, SQM89]. coiled-coil [QS87, SQM89]. coincide [HTGP85]. coincidence [TTGFC85]. coincides [MSS^{+}89]. cold [AY89]. cold-sensitive [AY89]. coli [DBLS87, SCK^{+}89, WGPW89]. Collagen [BFBL88, GCK^{+}88, KKG86, NHOY85, AT86b, BBP88, BT86, BMGI85, BB87d, BPE^{+}86, CMDCC86, CDMC88, CTYF85, CTSF86, CL85b, DKH^{+}85, DRP87, EHK86, GVD^{+}87, GF85, GLK^{+}89, GG87b, HNP^{+}86, HMTF88, HRP87, ISM85, KSF^{+}88, KSBB87, KSL^{+}87, KEG88, KDFG89, LHK^{+}89, LGF86, MB85a, MvdMH^{+}88, MEBV^{+}89, MGHG^{+}86, NSY86, PP89, SNCY87, SKMB86, SV87, STH^{+}89, SL85a, SRW^{+}89, SH86, SGS^{+}89, SHSvdM89, TH89a, TDR87a, TDR88, TLC85, TC86, UW86, VMH^{+}88, WC87, YR87, ZMH89]. collagen-binding [MvdMH^{+}88, NSY86, SNCY87, TH89a]. collagen-tailed [BMGI85]. collagenase [VKB89, WHMA86, WTB^{+}89]. collagens [JPU89]. collapse [SAA89]. collecting [BRD^{+}88, SS89b]. Colocalization [SH87b, CCM86, GGN^{+}86, HB88a, LOH^{+}86, WG88a]. colocalized [RDPR^{+}86]. colonocytes [GSF^{+}88]. comb [NT85]. combination [FTC87, TKK^{+}87]. combine [FGA86]. combined [SCOE^{+}86]. Commitment [WHMA86, CLOH87, LMSM89, TSI^{+}89]. Common [DMF^{+}87, CCC86, LMK88, PSR^{+}88, RHL^{+}85, TOOK^{+}85, WKKS^{+}87, WC87]. Communication [KL88, KL89a, KMR^{+}88, MBKO86, MBL89, PSC^{+}89, PGH^{+}88]. compaction [Tra88]. Comparative [KDFG89, MMR88, TLD^{+}89]. compared [Fal85, SR87]. Comparison [DMW88, GG85, BFSH87, BL85, CM88b, KDHS87, STH^{+}89]. compartment [CBYK88, EMW^{+}89, LBFS88, RKFN88, SFB^{+}88, SSS88, TLPTV85, VSSRB87, VSSRB88, WM85]. Compartmentalization [JIFH85, CR85, GB85a, TNG89]. compartmentation [CRF^{+}88, EOM86]. compartments [AK87, AO88, BEHJ^{+}88, BT86, BGL^{+}88, CIK^{+}89, GH88, KL88, KP89, LMW^{+}86, SM88, SM89a, WH85]. competent [RBGF89, SMS^{+}88, ZP85]. complement [AGPT88, CKF85, DWO^{+}87, LE86, WLG85, WPW^{+}85]. complementation [ZWE89]. complements [LM87]. Complete [HBPK87, HFK^{+}89, DBS^{+}89, HOPM^{+}87, SSSA87]. Complex [LC88b, Ake89, AG89, AB89, BWK87, BMEG89, BN87, BSA87, BFL88, BGF86, BNT^{+}87, CCA^{+}88, CC89, Cha85, CPP^{+}87, CHK^{+}86, CB87, CB88, CM87c, CKF89, DSA86, DMD^{+}86, DRP87, DRD^{+}88, DRC^{+}86, FGW85, FDG88, GCC^{+}89, GH85b, GPSM85, GPSC86, GFB^{+}89, GSG88, HAB^{+}88, HM89a, HSS^{+}87, IHJ88, JWH85, KOF88, KAB88, LKHR87, LM85b, LU88, MFB^{+}89, MFA89, MHD^{+}87, MC89b, NH89, PMC^{+}88, RPW86, RKFN88, SHL^{+}89, SSG87, SRW^{+}89, SBH^{+}88, Ste85, TLRW86, TT89b, UPL^{+}87, VP85, VB85b, WFE89, dVOD^{+}89, vDTP^{+}86]. complexed [GH88]. complexes [BTR^{+}86, BC85, CHH^{+}85, EMH^{+}87, FLM86, HF85a, JHPS86, KHB88, Pra86, SLB^{+}89, SCGS85, WM88]. complexity [HMS^{+}89]. component [ASJ^{+}88, BBGI85, CCMB87, EHC^{+}85, FS88a, GBBS89, HHSF85, MB87, OD88, PJW88, PSGS85, RYFM89, RKA^{+}89, SROB86, SKE86, SKMB86, SFJU86, SDS89, SK88]. Components [PN85, BB89, BKC^{+}86, BICFA88, CS88a, CC85b, CMB^{+}86, CIK^{+}89, GBC^{+}86, HKM^{+}87, HGL85, HB87, JPU89, KKGJ89, KB85, LW86, MP86a, PHRM87, PBH^{+}87, PJ87, Pum89, SCBHS^{+}89, STF85, WS87, WBVL86]. composed [DSTJ85]. Composition [EWHP87, BAG86, MPT88, MM87, Pra86, RSF86, RC85a, SSK^{+}88, ZHES89]. compression [DJS^{+}88, JCBH85]. computer [AJRNBJ85]. computer-assisted [AJRNBJ85]. computerized [BSK89]. concanavalin [BWS86, BMP^{+}86, KTF87a, SF88b, WGC85]. concentrated [FG89, FMT^{+}87, Ran88, SKU^{+}85, TS88, WTB87]. Concentration [TH89c, CR86, ET87a, Pfe87, Sch85, SSY^{+}86, VAD^{+}85]. concentrations [DF88a, JHP85b, YKSR89]. concert [SGP88]. concomitant [SDSO86]. Concurrent [EMT88, SAA89]. Condensation [TKFH89, FA85, LM85a, OON89]. Condensation-sorting [TKFH89]. condensed [TTF87]. conditional [ADS88, HGS89, HHR^{+}89, KP85]. conditioned [DWL86, LKOM88]. Conditioning [FM86]. conditions [CR87, DVVT88, SP89, ZKR^{+}88]. conductance [BGN^{+}88, LSDP86, MG85]. conduction [GSTS86]. conductive [GSB85]. Cone [Eck87, FS88b, GBBS89, GP86, LB89, SV89, TSK88, vdPdPR86]. cones [BD89, BA87, CR85, EWAP85, GAB86, GBBS89, HP85b]. conferred [LCL^{+}89]. confers [BB88a, MP86d]. configuration [CS86a, KTP^{+}89]. confluent [KG85b]. confocal [WAF87]. conformation [AHM89, CCBKJ89, NRL^{+}89b, SSW^{+}85]. Conformational [GML^{+}89, HB86, TH89b, WW87b]. conidial [WMGM85]. conjugated [PMJ87]. conjugates [vDTP^{+}86]. conjugating [PFTV89]. connect [AC86a]. connected [RBAF89, SGGS88]. Connectin [MYH^{+}85, HMP89]. connecting [HFB87, HAK^{+}88]. connective [CBCV85, HvdSdC^{+}89, KVB^{+}86]. connector [WSJ85]. connexin32 [GPJ88]. connexin43 [BKPG89, BPG87]. consequences [SMC89]. conservation [DBS^{+}89, GG87a, MHD^{+}87, WJSB88]. conserved [GGR89, GKH^{+}89, KWH89, MH88, BDB^{+}87]. consists [DB85]. constant [TT88a]. constants [BBP^{+}85, Pol86, WOP^{+}88]. constituent [GH85a]. constituents [JGAR89]. constitutes [GB87a]. Constitutive [CGA89, UPL^{+}87, RH87, RM89b, SMW89, WN87]. constitutively [DRME89, FFH88, YHS88]. constrained [WFE89]. Construction [LFC89]. Contact [MP86a, RGS88, WO86, AR88, BPB86, CS87, DI87, KOB89, KFW^{+}89, NV87, PN88a, PN88b, PMV^{+}89, Ran88, RGW^{+}89, ST86, SHN87, SG88b, SKS^{+}87, SCC87, VSSRB88]. Contact-induced [MP86a]. contact-mediated [AR88]. contactin [GCR^{+}89, Ran88]. contacts [BSG89, DNI^{+}88, JLK89, PHN^{+}88, RRF87, SW88]. contain [BPJBF85, Gla85b, GSTC87, HPW89, HD85, HSS^{+}87, HHD^{+}86, KB88, KR86, KS85c, LSF89, MSM88, MBD89, RCC^{+}88, RBAF89, SSS^{+}87a, SCRA87, SSW^{+}85, TL89]. containing [BJMK89, CR88b, DI87, DF86b, HB87, JRB88, LFFW89, LP86, MWK89, NH89, NRL^{+}89b, PLC87, SSKK89, SHL87, Ste85, SW87b, TCK87, TSK88, VS88b, Wal86, DAS^{+}87, FKP86, GAPR88, TT86b]. contains [CCL^{+}86, LHM89, LGF86, MHF86, MEBV^{+}89, MCCM88, NLC89, SSSA87, TLRW86, YPG86]. content [CDTS85, HO85, MCS85b, SAGM88, TT88a, WMSB^{+}89, vZC87]. contents [FBPS87, OE85]. continue [RG88]. continuity [BEHJ^{+}88, SME^{+}89]. continuous [GFSS86, SFLM86]. Contractile [CB86, JHP85a, JHP85b, BCKJW88, BGF88, HG89, IKH87, KCAF89, LWMM^{+}85, MSS^{+}87a, SE85, TSI^{+}89]. contractility [DW86, KDHS87]. contracting [BCKJW88]. contraction [BC89, BA87, GG87b, IKH87, KCCM85, SSH87]. contraction-relaxation [IKH87]. contrast [GB86a]. contrast-differential [GB86a]. contributions [HFC^{+}86, RGBM87]. Control [BKK87, KCK^{+}87, LOG85, LTG85, LR89, WBVL86, AWJ85, CLOH87, CHG87, CHS85, EZBW88, FMJS89, HW86b, JS86a, JS86b, JTBA89, KPK86, LL89a, NT85, NL85, PM87b, PCBD89, PC85, SN89, SJ86, SC88, SBS87a, SAB^{+}89, TYP^{+}87, TT86a, VN88, VSSRB87, RC85b, VN88]. controlled [ADS88, CF88b, MCG^{+}88a, SFMT87, VSSRB88]. controlling [CHG87]. controls [Koz88, SMR86, WR88]. conventional [WAF87]. Conversion [FBE85, ORA^{+}86, HSS85]. conveys [DDR^{+}89]. COOH [CHH^{+}88, POD89, SSS^{+}87b]. COOH-terminal [CHH^{+}88, POD89]. cooperation [DTW^{+}85]. cooperative [BDWH88, CB87]. Cooperativity [STF85]. coordinate [DFSK85]. coordinated [OGF87]. coordinately [GSB87, MLSW86, ORA^{+}86, SW87a, SYH^{+}86]. copolymers [LC87]. coprophila [Kub87]. copurifies [POW^{+}87, WTB87]. copy [LSC86]. cord [HBSQ^{+}85, SMA85]. core [CS88a, CB89, EMOB85, EBCO88, GCBH86, HvdSdC^{+}89, JRB88, KSS86, LM87, POW^{+}87, PN89, RFH85, RPCB88, SCM87]. cores [SCGS85]. cornea [LGF86, STE88]. corneal [BFBL88, HNP^{+}86, SGS86, SH86]. coronavirus [MR87, TTF87]. correct [RGPRB87]. corrected [IHJ88]. Correction [TIS^{+}86]. correlate [MBL89]. correlated [CS88c, Heu89a]. correlates [SCBR85, TCBF^{+}85]. correlation [CDTS85, HLW^{+}89, HTTFC85, OSD85, TKP^{+}88, VRLG89]. corresponding [COMS87, SQM89]. cortex [CS88c, MKA^{+}86]. cortical [BC89, BPC89, CWZ89, CCE87, Hym89, LOH^{+}86, TJF86, ZSE85]. cortices [OHK86]. corticostatin [OGJ^{+}89]. corticostatin/defensin [OGJ^{+}89]. corticotrophs [RFZH85]. costal [KG85a, KG85b, KIKS87]. costameres [Tok89]. Cotranslational [NADS86]. cotransport [SPO85]. cotyledons [BMP^{+}86, TB85a]. Coupled [CCM86, CGV^{+}86, ZKR^{+}88]. Coupling [ABC^{+}88, SSZ88, Bas87, CJ86, CHG87, GM86, WLG85, WPW^{+}85]. covalent [SB87]. covalently [CGV^{+}86, MCCM88]. cow [FZ85, HFZM87]. cranial [BFL88]. crassus [RLP85]. crawling [SRR89]. crayfish [BFAI87, Hir86, JS88b]. creatine [MAK^{+}86, WWW^{+}85]. crescentus [Smi87]. crest [BF85, BFL88, PJ87, RMS86]. critical [KJY89, Ris85]. critically [SMVS^{+}86]. crm1 [AY89]. cRNA [SKDF88]. Cross [CP87b, KHB88, MHWB86b, ML85, AC86a, AHC89, AB89, BB87a, BS88b, BHB^{+}87, CS86a, CM85, DHL85, DCB89, HKKP86, Heu86, Hir86, HMHWS86, KdS86, KdWB^{+}89, MSA88, NRL^{+}89b, PE85, SKK^{+}85, TASK86, VKP89, Wan85b, WKBR87]. cross-beta [NRL^{+}89b]. cross-bridge [CS86a]. Cross-bridges [ML85]. cross-bridging [Wan85b]. cross-fractured [KdS86]. cross-linked [VKP89]. Cross-linking [KHB88, MHWB86b, AHC89, BB87a, BS88b, BHB^{+}87, CM85, Heu86, KdWB^{+}89, MSA88, PE85, WKBR87]. cross-links [AC86a, DCB89]. cross-react [HKKP86]. cross-reacting [Hir86]. cross-reactive [HMHWS86]. Cross-reactivity [CP87b, SKK^{+}85]. cross-reacts [DHL85]. cross-resistant [TASK86]. crossbridges [BFAI87]. cruzi [BSPdSG88]. cryo [Ake89, MF87a]. cryo-electron [Ake89, MF87a]. cryoelectron [LU88, UTK88]. cryostat [CCS87]. Crypt [Qua85a, BJMK89, GSF^{+}88, OGJ^{+}89]. cryptdin [OGJ^{+}89]. cryptic [LCR^{+}87]. crystal [STGK87]. crystallin [YKCR88]. crystallin/alpha [WLW^{+}88]. Crystalline [QS87]. Crystallization [LM85b]. crystallized [UTK88]. crystalloid [JBG^{+}87, PLA86]. Crystals [GGMH86, KRLU87]. CS [WGPH^{+}89]. CS-1 [WGPH^{+}89]. CSAT [DKB^{+}85, HNR87, HDG^{+}85, BSDH86]. ctenophores [NT85]. cubes [SCFP86]. cucumber [TB85a]. culture [AW88, BBB89a, BHM^{+}89b, CHMB86, CNHF85, DKB^{+}85, DTW^{+}85, EPJBBW86, ER87, GB86a, GB89b, HBC89, HPR85, HLL88, HRP87, KLHK88, KDIG89, Kol86, LBO^{+}85, OW86, OH88b, PCK^{+}87, PLC87, Qua85b, RJH^{+}87, RF87, RRF87, SSR89, SR89, SGS86, SS88a, SPS89b, SHL87, TIS^{+}86, TTBG87]. cultured [AMG85, BCA^{+}86, BBBM89, BTO88, BF87, BTV^{+}89, BKR86, BRP88, BBMW85, CR88b, CRT87, CW88, CHH^{+}85, CS86c, DCC^{+}88, DDBHL86, DS88, EKH89, EMZ85, EWHP87, FGA86, GLC88, HTRBP86, HBF^{+}88, HMHWS86, HW86b, HFB^{+}89, IKSF89, IGSGD87, JPH85, KG85a, KKGJ89, KP89, LSAKO86, LMD^{+}86, LEFS^{+}87, LHL88, LP85a, LWMM^{+}85, MD85, MP89, MCG^{+}88a, MAK^{+}86, MBD89, MDW86, MG86b, NWM88, NKMK86, OSS^{+}89, OKLB86, OLGT86, OGYK88, PSBN85, PF85a, PHRM87, PAF88, PWS^{+}89, PSS^{+}87, PB87, RFV85, RT88, RRP88, RG86b, STS^{+}89, SF89, SF88b, SF87, SHSvdM89, TFBW85, TLPTV85, TLD^{+}89, UW86, VSZ^{+}87, Wal86, WE86a, YMM86, YMM88, YMD87, vGRB86]. cultures [BPB86, BSB^{+}86, CB86, DMW88, FCM85, GCK^{+}88, GJD^{+}88, HZS88, HH86, KG85b, KIKS87, KFW^{+}89, Maj87, MOP^{+}85, MAG88, MCS85b, RR87, Ros89, SR86a, SE85, SFS^{+}87]. culturing [TRT^{+}88]. current [CS86c]. currents [BR85, ICY^{+}88, Kro86]. cutaneous [DMG^{+}89]. cuticle [DFF85, LOH^{+}86, WFF86]. cuticular [DT89, TTS^{+}89]. CuZn [CSGC88]. Cx26 [ZN89]. Cyanobacterial [SSZ88]. cyclase [GDC^{+}86, UMHI85, Van87, WMV86]. cyclase-deficient [UMHI85]. cycle [BTO88, BMB87, BGF86, CL85a, CL86, CPC^{+}89, HBB^{+}89, HMP86, JASP88, LTG85, LM87, MSV^{+}88, MWM89, OSS^{+}89, QCKR85, SK88, SSS^{+}87b, TDR^{+}87b, WBW87, WRR87, WMC86, vGRB86]. cycle-dependent [MWM89, OSS^{+}89]. cycles [RG88]. Cyclic [AR85, Goo89, HB88b, PG87, CS89, CWH86b, ECZ86, Eva89, HJS^{+}85, KSKW^{+}85, RRB86, IKH87]. cyclin [BMB87, LR89, OOCT89, TB88a, WSR89]. cyclin/proliferating [BMB87, TB88a]. cycling [Ros89, Wan85c]. Cyclosporine [RT86, SNC^{+}86]. cysteine [BBV^{+}89, KWH89, WMSB^{+}89]. cysteine-specific [BBV^{+}89]. cytoadherent [WS89]. cytoadhesins [GLW^{+}89]. cytoarchitecture [SP88]. Cytochalasin [BM86, BR85, BS85, Coo87, MBC86, TT88b]. cytochalasins [FS88b]. cytochemical [AMG85, BPJBF85, CMHC85, HSK88]. cytochemistry [JPH85]. Cytochrome [MYT87, DFSK87, KBPW89, MVK^{+}88, SSK89, VME85, YMT85, MYT87]. cytodifferentiation [CCM86, OLGT86]. cytogenetic [MSC^{+}85, TASK86]. Cytokeratin [KFH^{+}86, BBBM89, BAD86, RBAF89]. cytokeratins [BLA^{+}88, FGA86, HF85a, LBB^{+}88, RFV85]. Cytokine [NSF^{+}89]. Cytokine-induced [NSF^{+}89]. Cytokinesis [WDV87, DCR89, HG89, SL87a, WMC86]. cytological [GC87]. cytomechanics [DLBH89]. cytometry [CDTS85, GML^{+}89, JIFH85]. cytoplasm [BKR86, CNL86, CM86c, DF86a, GORW88, LPTL86, MP88b, PT88, RCC^{+}88, SROB86, TDR^{+}87b, Wan87]. Cytoplasmic [KF86, LP86, PMR86, VA85, AWH^{+}85, AKVF88, BEHJ^{+}88, BKBR88, BRHR^{+}86, BDS85, BCPC87, BFRT87, CVWZ89, CHD^{+}86, CRJM88, CM87a, CMW87, CH85, CdCP^{+}89, DI87, DSGS87, DRSG85, DSG86, ES86, FBP89, FKH88, FWF88, GK87a, Heu89a, Heu89b, HL88, HTH85, HLU^{+}87, IKH87, KBR87, KR88, KMY88, LWSF85, LBP^{+}86, MKK89, MNC86, MH88, MBLK88, MKH88, NB88a, PC85, RG88, Ris85, RPW86, RG86a, RDSG86, RNW89, SME^{+}89, SL87a, SB85a, SPV88, SMR86, SNSS87, SSS88, TS86, WFE89, WKKS^{+}87, WS85, WRR87, WMC86, WL87, ZH86, BKR86]. cytoplasmically [HSS^{+}87, LD86]. cytoplasts [AJAW86, PC85]. Cytoskeletal [HBV85, HTH85, HFB87, KKB89, NH87, AYCY89, BCA^{+}86, BS88a, BDS85, BSNG85, BTV^{+}89, BB88b, BJMK89, EB87, HBC89, HSK^{+}89a, HB87, HSMM85, JJP^{+}89, KP85, KS86, LB89, LEG^{+}89, LKF87, LWMM^{+}85, MEA89, MMS89, NH89, OS88b, OH88c, PBB86, PHR86, PSD^{+}85, Pri87, Pum89, SSY^{+}86, SLL89, WHMA86, WBD^{+}86, ZG87]. cytoskeletal-associated [BJMK89, SSY^{+}86]. Cytoskeletal-membrane [HFB87]. cytoskeletally [LR85a]. Cytoskeleton [FLH^{+}88, BSK86, BFPS85, CSL87, DJS^{+}88, DD88, FL88, GBBS89, HB88a, HBW^{+}87, Hol89b, JS86a, JCBH85, KA86, KTT^{+}86, KCN^{+}86, LWR85, MBC86, MRD^{+}89, MKD^{+}89, NYG87, OS88a, PZBD87, PH89, PN85, RJB86, RFV85, RBAF89, SBS88, SNNK85, SRR89, TRM88, TTKM86, TTM88, WS85, WGC85, YSW^{+}85]. Cytoskeleton-associated [FLH^{+}88, SNNK85]. cytoskeleton-integrin [MKD^{+}89]. cytoskeletons [RPCB88, VB85b]. cytosol [SOPvD87]. Cytosolic [KM87, Sch85, BMEG89, BN85b, CJO^{+}89, DMGS88, LMW^{+}86, MMMP87, MMT85, MPB88, RG86b, VAD^{+}85]. cytotactin [CHG^{+}86, DCPR^{+}89, FHE88, HCE88, RDPR^{+}86]. cytotactin-binding [HCE88]. cytotoxicity [GWCW85, MDVG^{+}85]. cytotrophoblast [FCZ^{+}89].

D [TSOC89, VMH^{+}88, FWP^{+}87, HMP86, KKB85, MBC86, SLHvF85]. D-1 [HMP86]. D-2 [HMP86]. D-periodic [VMH^{+}88]. D3 [PBK88]. dalton [Fec87, SGRM85]. daltons [GAH86]. damage [BFPS85]. damaged [RMM87]. Darby [ASJ^{+}88, BAD^{+}89, BBGI85, BPP^{+}89, BKK87, BHM89a, FGF88, FS86b, GDRB89, HM89c, JY86, LLSRB89, MFKVK87, MCA^{+}89, MBH87, NV86, NV87, NH89, OS88a, PPB^{+}89, PN88a, PN88b, PBH^{+}87, PFS85, RIPS85, RIS87, RG86a, SGRM85, SG88b, SAGM88, UPL^{+}87, VSSGRB87, VSSRB87, ZMH89, vMSvRS87]. dark [KMM88]. dark-grown [KMM88]. deadaptation [Van87]. deal [KG89]. death [MSD^{+}88]. decarboxylase [SHL^{+}89, SPO85]. Decomposition [Toy87a]. decorate [HvdSdC^{+}89]. decorated [CB87]. decorates [HH87]. decrease [Wan85b]. decreased [MSS^{+}89]. decreases [LSC86]. decreasing [VES85]. deduced [FTLSS88, GCC^{+}89, HBPK87, PAHC89, ZN89]. Deep [HKA^{+}87, HK88a, ASF89, BO85, FAKVM87, GH85a, LC88a, NH87]. Deep-etch [HKA^{+}87, HK88a, ASF89, BO85, FAKVM87, GH85a, LC88a, NH87]. defect [AJCMR88, CM88a, CMW87, DCWJ89]. Defective [AWJ85, BSR^{+}89, BRKE88, BGH^{+}87a, DS87, EFR^{+}89, KSJV88, LRF89, MvdMH^{+}88, RKFN88, SFK^{+}89, SPO85, TSI^{+}89, TSM87, UMHI85, WDV87, WMUP^{+}85]. defects [CNB89, KKSK86, KBK86, OB88, RFMR86]. defensin [OGJ^{+}89]. deficiency [AMG85]. deficient [ADS88, ABG^{+}89, KKSK86, LEW^{+}88, MRO^{+}86, MGM87, PBvTS88, RYR^{+}89, SBD^{+}89, SYH^{+}85, UMHI85, RdRBR89]. define [MSC^{+}89, PNB89]. defined [KMP88]. defines [GD89, MCS85b, Ran88]. defining [LCBS89]. Definition [BH86b]. deformability [CM86b]. degeneration [IKH87]. Degradation [SCLG86, SS89c, ALL89, BSLS^{+}89, CSM^{+}89, CRR87a, CBYK88, DCQB88, FGO^{+}85, HVR^{+}85, JBG^{+}87, KTF87a, LL88, MUM87, SMSR88, TRM88, Toy87b, WKRH86]. degradative [FCZ^{+}89]. degraded [MYT87]. degrades [GGR89]. degrading [BIES89, GSDR89, IS88, SEB89]. degranulation [AS86, EMH^{+}87, KHB88]. degree [HLW^{+}89]. degrees [GLK^{+}89]. dehydrogenase [NB88b, SE86]. delay [SEGS85]. deletion [HSWH86, JBG^{+}87, WGPW89]. Deletions [PTB^{+}85, DSG86, LBP^{+}86, POD89, SS88b]. delivery [AFMD89, BP87b, CH85, EKH89, GSS^{+}85, LLSRB89, RRB86, RIPS85, RIS87, WG87a]. delta [BB88a, CPG^{+}89]. delta-subunit [CPG^{+}89]. demonstrated [BK86, RdRBR89]. Demonstration [GR89, AMG85, GF86a, KNGH^{+}85, LWSF85, MGGP88, PC85, RLB85, WLJM^{+}89]. demyelinating [GFHDD89]. denervated [CM87b, CMGS86, CCS87, GSS89, RMM87, SSC86, SBS87a, SS89c]. denervation [CMM^{+}85, FHS^{+}86, JB87, TS89]. dense [KCAF89, KSS86, Rog87a]. densities [SB85b, Wil88]. density [AK87, BMR85, BHB^{+}87, BTO88, DTW^{+}85, DCB^{+}88, GW86, KKSK88, KKSK86, KBK86, LRF89, MM87, PSBN85, PMC^{+}88, PBMB85, TFBW85, WE86a]. dependence [AE88, BTO88, HLL88, NSF^{+}89, SSS^{+}87b, SOP85]. dependency [LMW^{+}86]. dependent [ACN^{+}89, AELE88, BMBL89, BR89, Bro87, CR88a, CPC^{+}89, Cha85, CKR85a, CPP^{+}87, CM87a, CV87b, CMTC85, CWH86b, DLG^{+}89, GK85, GFS^{+}89, GZO88, HNNT88, HB88b, HNH^{+}87, JY86, KOB89, Knu85, KSKW^{+}85, LFC^{+}88, LFFW89, LDA86, LWM^{+}88, LD86, MS86a, MER^{+}86, MWM89, MK85b, MCG88b, NGS85, OSS^{+}89, OH88c, PFD^{+}89, PMV^{+}89, RRB86, RO85, RT86, SNCY87, SVF^{+}89, SS85, SM89b, SNNK85, SC86, SMC89, SL85d, SYT^{+}89, SHS^{+}88, SRW^{+}89, SRPS86, THM86, TKP^{+}88, WESP85, WKS85, WHL^{+}86, YBBS87]. depending [TT88a]. depends [CSG85, GBBS89, HP87b, MPT88, MK87, NG85, PB86, SSK^{+}88]. dephosphorylation [EG89, GF86b, ZP85]. depleted [JBTA88, KWPH89, FKP86]. depletion [GWCW85, HTTFC85, JM86]. deployment [TMH^{+}85]. Depolarization [MF87b, JPH85]. Depolymerization [AYG85, BC85, CWS86, GGMH86, SOP85]. depolymerizing [BB87b, JS86b]. deposition [CMHC85, CRC^{+}87, EHK86, HBC89, LSAKO86, LBO^{+}85, LLCA89, MCB85, SABVK89, STF85]. deposition- [LLCA89]. deposition-related [CRC^{+}87]. depress [HKM^{+}87]. depressed [DM87, PRS^{+}85]. deprivation [MSD^{+}88]. deprived [SR86a]. derepressible [dHTG88]. derivatives [VME85]. derivatized [YSR87]. derived [BSBZ88, CR88a, DKH^{+}85, GDD88, GM85c, GHA88, GAH86, HKM^{+}87, HTRBP86, HP85a, KC88, KBO^{+}88, LHvF^{+}87, LRF89, MCB85, MB85b, MEVM89, NMPG88, PML^{+}89, RH89, RYFM89, SMSR88, SKK^{+}85, SHGD85, SLG88, SHS^{+}88, TRT^{+}88, ZR87, AdS85]. dermatan [LCR^{+}87, SRH^{+}87]. dermatosparactic [MvdMH^{+}88]. dermis [DDB^{+}88]. desensitized [UTK88]. desmin [FS86a, HDL^{+}86, Ip88, PSK^{+}89, TMS85]. Desmocalmin [TT85]. desmoglein [PN89]. desmoplakin [PN88a, PN88b]. desmosomal [BTV^{+}89, CKF85, DF88a, HTT89, KOF88, KTF87a]. desmosome [JG85, PN88a, PN88b, PN89, PHRM87, PBH^{+}87]. desmosomes [BAD86, GGJ^{+}87, JYG86, RBAF89, TT85]. destined [DMM89]. destruction [DSTJ85, LR89]. Detachment [HPR85, DAS^{+}87]. detect [RMP^{+}88, TH89b, WW87b, YSR87]. detectable [FBP89]. detected [BLA^{+}88, LVSC86, MFA89, TTM88]. Detection [CBIO85, HWSK89, ZSB86]. detects [LSC86]. detergent [CE89, CCG89, GLH85, LWR85, MCC^{+}89, WFL87]. detergent-insoluble [LWR85]. detergent-resistant [MCC^{+}89]. detergent-solubilized [CE89, CCG89, GLH85]. determinant [WRS^{+}86]. determinants [CCL^{+}86, FR86, MSC^{+}89, MWS85, WFR86]. Determination [DF86a, DFSK87, HW87]. determine [CMR^{+}87, HFGD89]. determined [MMS89, UTK88]. determines [KGRP85]. determining [Hym89]. detyrosinated [BKK87, CDB87, GGN^{+}86, KGB88, WGBB87]. detyrosination [BRP88, GKB87, KGB88]. Developing [LSF89, ACEE87, CV87a, CJ86, CR85, CS88c, CS86d, CMGS86, DM87, FG89, FLBP^{+}86, GABT86, GSB85, KCN^{+}89, LHK^{+}89, LSB89, MS86b, MCG88b, MAE88, MGW86, NFI^{+}87, OD88, RBMM88, SC85a, SV87, SCM87, SABVK89, TB89a, TMS85, ZSB86]. Development [CHH^{+}85, FMJS89, GJD^{+}88, KMKS87, Qua85b, AP86, AMWN89, AE88, BBB89a, BYB^{+}88, BSB^{+}86, CRF^{+}88, CMDCC86, CRLW86, CHG^{+}86, DHK^{+}85, FC89a, GF85, GZHM87, GB89b, GG87a, GCPE85, HD88a, HBSQ^{+}85, HMF^{+}87, HPM87, JS86a, JFAT89, KSH87, LSEN87, LCE87, LLC85, LSC86, MP88a, MP86a, NV86, NB88b, OE85, PS87, PSGS85, PGH^{+}88, Qua85a, RB85, RLP85, SLB87, SA85, TS88, WMGM85, WG85, WSM86, YKCR88]. Developmental [MS86c, NWE^{+}88, OGJ^{+}89, YMD87, CKR85a, GKG88, HNR87, HCE88, LMSM89, LHNF87, MOP^{+}86, PL86]. Developmentally [CMT^{+}88, CHS85, DLC88, AE87, CHMB86, CSL87, FL89, GLD^{+}88, KFHF85]. developments [RM89a]. devoid [OK86]. Dexamethasone [GZO88, DNB88, GHA88, MLE87, OGF87]. Dexamethasone-dependent [GZO88]. dextran [MM87, PMJ87]. diacylglycerol [MSE89]. diameter [KSBB87]. diaminobenzidine [AK87]. diaphragmed [LMF^{+}86]. diaphragms [BO85]. diatom [WMC86]. Dictyostelium [ABG^{+}89, BHRC85, BP87a, CC85a, CGD86, CS87, DBLS87, EFR^{+}89, Fec87, FMG89, FWF88, GDF86, GW86, GDS87, JT85, KWLS88, KSJV88, KLHS85, KTP87, LSF89, MCD85, NRL^{+}89b, PLS89, PG86, PFRS89, PSP85, RWE^{+}88, SL86a, SL88, SFG87, SCC87, SB85f, Van87, VES85, WD86a, WDV87, WSV^{+}89, WSLN86, WK85, WL87]. differ [HM89c, SAGM88, WSH85, WMSB^{+}89]. Differences [AED^{+}89, HDP^{+}85, MM87]. Different [GM88, Heu86, BPL87, CCC86, CGD86, CSR^{+}86, CKR^{+}85b, CRLW86, CKF85, DRSG85, FTC^{+}89, FS86a, FCH^{+}88, GE88, HFC^{+}86, IGSGD87, JYC87, LH86, LL86, MP88a, MSV^{+}88, MMK^{+}89, MHP^{+}86, OR85, RZG86, RSR^{+}88, STE88, SDDZ^{+}89, TRL87, YBBS87]. Differential [BN88, DFdN^{+}89, DRM^{+}88, FBF^{+}89, GK86, HNP^{+}86, HMTF88, HFC^{+}86, JC89, LSEN87, LHL88, LEG^{+}89, NR89a, PFD^{+}89, PAF88, RKWS86, SRM89, SHR86, TAHC^{+}89, Van87, VRLG89, WG87a, WG85, WMH^{+}88, CCE87, FOW89, GF88, GB86a, HVR^{+}85, LL88, LN88, MK85c, OB88, OS88b, RH86b, SS87, WLG85, WPW^{+}85, WFF86]. differentially [HD88a, HERO^{+}87, JBdlTB87, MND^{+}87, OHH^{+}89]. differentiate [BRP88, RBGF89, SW87a]. Differentiated [SSSP85, BBP88, BJMK89, GS85, KMKS87, LSB89, ZRW85]. differentiating [ACN^{+}89, BPMD87, CMT^{+}88, JC89, LMSM89, LL89b, SL89, SDDZ^{+}89, TS86, VSR87]. Differentiation [EBBW87, GHA88, MS86a, SGS86, UPRS88, AW88, AR85, BSPdSG88, BGK^{+}89, BS86, CSH89, CRT87, CDMC88, CKR85a, CLOH87, CH87, CFB^{+}85, CKF89, DG89, DDB^{+}88, DNSMA88, DRH^{+}87, FSFL88, FCL86, GK85, Gor89, GZO88, GW87, GKB87, GKB89, HBSQ^{+}85, HRBR89, HM89b, HSMCL87, IF89, JS86a, KMR^{+}88, KST^{+}89, KSSAQC89, KSJF87, KTF87a, KFH^{+}86, KDIG89, KWS^{+}89, KTF87b, KKG86, KKML88, LRF88, LOG85, LTG85, LLRR87, LL88, LS85, LC88b, LL86, MTCE87, MLSW86, MER^{+}86, MMK^{+}89, MSS^{+}89, NR88, NR89a, NB89b, OSH^{+}86, OH88b, PCK^{+}87, PL87, RSP^{+}89b, RKA^{+}89, RPCB88, RS88, SSR89, STS^{+}89, SRT^{+}86, SDR^{+}88, SL86b, SKDF88, TSI^{+}89, TTLR89, TNG89, TRC^{+}87, TMH^{+}85, VBWC88, WS86c, WBVL86, YKSR89, ZV89b]. Differentiation-dependent [MS86a, CKR85a]. Differentiation-related [SGS86]. differently [HGL85, PS86, VES85, WTBBW85]. Differing [SMW89]. differs [SDPZ89]. diffraction [BMDN87, KLS85, MMS89, STGK87]. diffuse [BBMW85]. diffusibility [WHI86, WSM86]. diffusion [BLSK88, CMK87, DCQB88, FMK^{+}86, HRR89, RG86b, SC85a, SHH85, SHP^{+}87, SF88b, SGP88, TLP^{+}87, WFE89, WE86a, WHL^{+}86]. digested [VT89]. digestion [JWH85, OCP86, Toy87a, Toy87b]. Digital [BBP^{+}85, PBS85]. dihydrocytochalasin [BBP88, BB88b]. dihydropyridine [JSA^{+}89]. Dihydroxyvitamin [PBK88]. dimensional [BSA87, BSK89, BDCT88, CD89, HS87a, HS87b, LU88, MASA89, MMS89, MSS87c, SKL85, TRCR89a, TRCR89b, YF87]. dimensionality [MP86b]. dimer [CMR^{+}87, UK86]. dimeric [CHFA89]. dimerization [SSP89]. dimers [MHWB86b]. dimethylguanosine [LHM89]. dinoflagellate [JIFH85]. dinoflagellates [NNJ^{+}87]. dinucleotide [ACPFB^{+}85]. dipeptidyl [RPRS89]. diphtheria [MOU88, MDVG^{+}85]. diploid [WO86]. Direct [Bäc88, CWOK87, KMK86, MASA89, MGGP88, RGM85, WKBR87, BICFA88, CS86c, GVD^{+}87, KGB88, KW86b, RDSG86, SB85e, TSS89]. directed [AC86b, BKPG89, DW89, HWSK89, HMdN^{+}88, MVSW86, MAG88, MBSMD86, OH88c, TJ85, BLK^{+}88, CPP^{+}87]. Direction [FS87, MP86d]. directly [BKR86, GR89, HJV^{+}89]. directs [AM85, ES86, HKF^{+}87, PL85b]. disappearance [CJH85, Wan85c]. disassemble [GSB87]. disassembly [AFSL85, CWZ89, LEFS^{+}87, LHS^{+}89, MGR86, SPH87, SG86]. Discharge [PW85]. discoid [BFPS85]. discoideum [BHRC85, CC85a, CGD86, CS87, EFR^{+}89, Fec87, FMG89, FWF88, GDF86, GW86, JT85, KWLS88, KLHS85, MCD85, PLS89, PG86, RWE^{+}88, SL86a, SL88, SFG87, SCC87, SB85f, Van87, WSLN86, WK85, WL87]. Discoidin [BHRC85]. discrete [AIJ^{+}89, CCA^{+}88]. discriminate [GLC88]. discriminates [HB88b]. discs [Pri87]. disease [GFHDD89, RPCB88, WCSR85]. Disialoganglioside [CK86, CPP^{+}87]. Disialogangliosides [CPHM86]. disk [CD89, DCB89, MM87, PAHC89]. disks [MCC^{+}89]. dismutase [CSGC88]. disorders [AMG85]. displaced [VP85]. displacement [OH88c]. display [DAJJ^{+}88, GSS87, YBBS87, ZH86]. displaying [KFH^{+}86]. displays [BPC89]. disposed [HSS^{+}87]. disrupt [SWTB88]. Disrupted [HJV^{+}89, GW88b]. Disruption [PB87, BFG86, GLC88]. disrupts [BR85, CF88a]. Dissecting [BMVB89]. Dissection [SMB^{+}88, WGPW89]. dissociated [SGM^{+}86, TAMR85]. dissociates [NYG87]. Dissociation [BBGI85, MSE89, SHGD85, MFB^{+}89, SW88]. distal [Eck87, MSQ^{+}89, NMM^{+}86]. Distinct [PSS^{+}87, SSS^{+}87b, TRB86, WBW89, BSG89, BBJG89, BD86a, BNS85, CM86b, CML85, FKC87, GB87b, GTP87, GDvdM87, HDP^{+}85, HFZM87, HSK^{+}89b, HHS^{+}85, HM89b, JBdlTB87, JRB88, MHF86, MM89, MG86a, MNWL85, MK85c, MPB88, NB89b, OW86, PB87, RKWS86, RZG86, RLP^{+}87, RVS88, SCRA87, TLRW86, WSW85, WGA85, YM87a, vZC87]. Distinctive [SF87, Pum89, RSF86]. distinctly [CGD86]. distinguish [CHG87, HW86a]. distinguished [KSSAQC89]. distinguishes [BBH87, Rog87a, RDB^{+}86]. distributed [CBCV85, HGL85, NL86]. distributes [CK86]. Distribution [AELE88, BB87b, CS86d, DKB^{+}85, DD87, GB86b, HTSH85, KDBF86, RNPR^{+}88, SMH88, SK88, TB89a, Tra88, TCP^{+}85, WSM^{+}87, YBWLM88, AED^{+}89, AMWN89, BNB^{+}88, BFR85, BKBR88, CR87, CWH^{+}86a, CHH^{+}88, Cou87, DH89a, DWFT88, DMD^{+}86, DRH^{+}87, GSS87, GK86, GKZ89, GDF86, GPJ88, GW89, HRR89, HP85a, Heu89a, Hol89a, HO85, HHSF85, JLB86, JSA^{+}89, KSF^{+}88, KSSP86, KCN^{+}86, KKB89, KYS^{+}88, LFFW89, LCE87, LLSRB89, MHK^{+}85, MMW85a, NB88a, NMKS85, OD88, OS88a, PRSGD89, RH89, Rob87, RRF87, RG86a, SBS88, SL89, SSK^{+}88, SA85, Str86, TDGE85, TB85b, TB88a, VLD^{+}85, VMH^{+}88, WM88, WF86, WMC86, ZN89]. Distributions [TMS85, BL85, HDP^{+}85, SS87, WSW85, YKCR88]. disulfide [VS88a, WMUP^{+}85]. disulfonate [VS88b]. Dithiothreitol [Ibr87a]. divalent [CPP^{+}87, MCG88b]. divergent [JYC87, May89]. Diverse [HBF^{+}88, HTB88, AE87, CIK^{+}89, GM85c, PNB89]. diversity [GP86]. dividing [Hep85, TDR^{+}87b]. division [HW87]. divisions [Hym89, WSW85]. dl312 [Ros89]. Dm0 [GLD^{+}88]. DNA [AJRNBJ85, BBBM89, BMB87, CW88, CDTS85, EMG86, FTC87, HKTY89, HVBB88, HFF^{+}85, KW86b, LLAS85, LNH86, MMM^{+}89, MMWM^{+}89, OON89, OO87, Owe85, PV85, RG88, Ros89, SR88, TASK86, TB88a, VP85, WLG85, WPW^{+}85, WDR89]. DNA-binding [PV85]. DNA-protein [VP85]. DnaJ [SCK^{+}89]. do [AW88, CSR^{+}86, HSWH86, HBKK87, LBP^{+}86, NVB^{+}88, SR85b, VFB^{+}89, ZR89]. document [LLCA89]. dodecylimidazole [WFL87]. Does [IF85, BM86, FRG^{+}87, RG86a]. dollar [HTSH85]. domain [APF^{+}88, BBH85, BH86b, BHCD^{+}89, CVWZ89, CJY86, CTSF86, CHH^{+}88, CG86a, CLC^{+}86, DMH^{+}85, DRSG85, DCB^{+}88, GFS^{+}89, GM85b, HG88, HGL88, HAK^{+}88, JRSB87, JBG^{+}87, KWLS88, LCBS89, LCR^{+}87, LBP^{+}86, MR87, MKK89, MH88, MMY88, NFW89, NAS^{+}87, NBS88b, NLC89, NRL^{+}89b, PN89, PTB^{+}85, POCM88, PMR86, RG86a, RGPRB87, SVSH^{+}88, SKU^{+}85, TH89a, TC86, VSSRB88, WGPH^{+}89, WW87a, YIJ88]. domains [AMWN89, BSG89, Blo86, BKR86, BBMW85, CCA^{+}88, CRJM88, CMK87, DMF^{+}87, DRM^{+}88, DSG86, DBS^{+}89, EDD^{+}86, GWM85, GDvdM87, GML^{+}89, HMTF88, HM89c, KKB85, KGHK^{+}87, KJY89, MBLK88, MHF86, MM89, MBSMD86, PHL89, PBB86, PL85b, PA88, PMR86, Pum89, RLP^{+}87, RSF^{+}85, RDSG86, STE88, SSV^{+}85, SHP^{+}87, SSW^{+}85, TRL87, TTKM86, WFE89, YJL^{+}89, YE87, dCFB86]. domesticus [Hut87]. dominance [Ree85]. dorsal [RBG85]. dosage [MVB85, MVB87, PB86, SRPS86]. dosage-dependent [SRPS86]. Dose [CM87a, MMS89]. Dose-dependent [CM87a]. dots [TM87b]. double [BFAI87, SH87b, SLL89, TB85a]. double-headed [BFAI87]. double-label [SH87b, SLL89, TB85a]. doublet [HFB87, LS87b]. down [BSd^{+}88, LDKG87]. down-regulates [BSd^{+}88]. down-regulation [LDKG87]. downregulation [LCW^{+}89, MQ89]. dried [Ris85]. driven [ICS88, VH88]. droebachiensis [OHK86]. droplets [GMF85]. Drosophila [BDB^{+}87, FCDF87, GLD^{+}88, MG89, PBD^{+}89, ASLFP89, BGT^{+}89, BHCD^{+}89, CF88a, CNL86, DM87, DFF85, DBS^{+}89, FA85, GST^{+}87, GSDR89, GK86, GLM86, HMG^{+}86, HS87a, HS87b, JFAT89, KG89, KA86, KFA89, KF86, MFA89, MASA89, MTS89, OB88, RG88, SCBHS^{+}89, SGBF87, SF89, TMP87, WSW85, WFW87, WG85, WFF86, WSRT89]. Drosophilia [DBB^{+}87]. drug [EKH89, JS86b]. drugs [MMY88, RIS87]. duct [BRD^{+}88, SS89b]. due [Wil88]. Dunaliella [EMT88, MT86]. duplicated [APF^{+}88]. duplication [RdLB87]. duplications [ETA85]. during [AP86, AMWN89, And86, AVP^{+}89, AS86, AE88, AR85, BH86a, BBB89a, BAD^{+}89, BYB^{+}88, BAG86, BSPdSG88, BTV^{+}89, BMB87, BB88b, BM88, BIES89, BBB89b, BA86, CGV^{+}86, CWOK87, CRF^{+}88, CMDCC86, CC89, CCM86, CK86, CA86, CCE87, CG86b, CHE87, CHG^{+}86, DG89, DHK^{+}85, DNSMA88, DKKK88, DDH^{+}87, DRH^{+}87, DH89b, DN88, ER85, FCVDH89, FC89a, FCL86, FM86, GKR87, GHL^{+}89, GZHM87, GFHDD89, GM86, GGJ^{+}87, GG87a, GCPE85, GB86b, HD88a, HCHK89, HS86b, HC87, Hep89, HMP86, HF85b, HDL^{+}86, HVL87, HMP89, HPM87, HM89b, HHSF85, IF89, JS86a, JFAT89, JGYG85, KSH87, KTF87a, KW86a, KW86b, KKG86, KKB89, KM87, LR85b, LRF88, LCD^{+}89, LLRR87, LSDP86, LSEN87, LL88, LM87, LCE87, LLC85, LSC86, LC88b, LL86, LBL^{+}85, MHK^{+}85, MKK89, MP88a, MLSW86, MBLS87, MMC88]. during [MNM87, MND^{+}87, MSS^{+}89, MG89, MTK^{+}87, MAE88, MIHS86, NV86, OSP^{+}85, OH88b, OE85, PFD^{+}89, PS87, PBD^{+}89, PBT^{+}86, PSGS85, PRS^{+}85, RMS88, RGS88, RPW86, RB85, RSP^{+}89b, RSR^{+}88, RS88, SROB86, SSH87, SS89a, SFLM86, SR89, SL89, SFJU86, SL87a, SDSO86, SK88, SB85c, SAA89, SKS^{+}87, SRM89, SGBF87, SF89, SEB89, SA85, SL86b, SALN88, SKDF88, TS88, TSBW86, TMB85, TPBB85, TT86a, TI85, TI87, TTBG87, TT86b, TB88a, Tra88, TMH^{+}85, TF86, UMHI85, VBWC88, VB88b, WS86a, WSRR89, WSW85, WBW87, WS86b, WG85, WLG85, WPW^{+}85, WHI86, WSM86, WG88b, WK85, WMC86, YBWLM88, ZV89b]. dwarfism [WTBBW85]. dyadic [BN87]. dye [CH85]. dyes [HH86, SNSS87, SSS88]. Dynamic [BEHJ^{+}88, IKH87, KBR87, MKD^{+}89, SK87, WOP^{+}88, WKR86, AST89, CPS88, CS86b, DWFT88, GFB^{+}89, JPM85, SFW87, SV89, SW87b]. Dynamics [CWS86, NV86, FJMW87, GHL^{+}89, GSB88, HBB^{+}89, KMK86, LSF86, MRD^{+}89, NJAS89, NL88, OH88a, OSS^{+}89, PKK89, RBKK87, RGBM87, SGB87, SL87a, SK86]. dynein [FS87, GTG85, GS89, GH85b, HTH85, Kam88, KW88, MRWJ88, MR85b, MTG88, NB88a, OK86, PSV87, Pip88, SGH85, SF88a, SPV88, Toy87a, Toy87b, WMR86]. dynein-like [PSV87]. dyneins [VT89]. dysenteriae [SOB^{+}89]. dystrophin [BHCD^{+}89, HWSK89, NRL^{+}89b].

E-cadherin [CG89]. E.B [Pol89]. E1 [MR87, TTW88]. E2 [CG86c, CMG86]. E8 [GRvdM89]. each [LC87]. ear [BNS85]. Early [CMW87, MSLR^{+}86, TS86, ZSSR86, AKVF88, BS88b, DS87, DRC^{+}86, DMM89, ETA85, FMJS89, GM86, GGH89, GKB89, HVL87, HW87, Hym89, KA86, KFA89, MCS85a, MCS85b, NKA^{+}87, NL85, NFN87, PRS^{+}85, RWE^{+}88, SKS^{+}87, SF89, Str86, TT86a, TM87a, TM87b, Tok89, WSRR89]. ecdysis [DLC88]. Echinonectin [AEB^{+}88]. ECM [SH86, SH86]. ecto [RLB85]. ecto-galactosyltransferase [RLB85]. ectoderm [CG89]. ectodomain [JRSB87]. ectopically [RPW^{+}89]. ectoplasmic [HGL88, VS85]. edge [DWFT88, Wan85d]. edges [PAHC89]. Effect [BDS85, BB87d, CHFA89, CMB^{+}86, FY86, KG85a, KIKS87, KTP87, Owe85, Owe86, SNNK85, SDZB^{+}85, TCBF^{+}85, BHB^{+}87, CL85a, FGO^{+}85, FHS^{+}86, GHA88, MLPF89, MCC86, PAGPLM86, SP89, SVG^{+}86, SH86, SRW88, Wil88]. Effects [BNS85, CRT87, CM85, Coo87, Heu89b, JB87, KLHK88, Kol86, MBC86, NTNT85, SHP^{+}87, WE86a, YSW^{+}85, dLG86b, AR85, CSM^{+}89, CC88, DMW88, DRM^{+}88, DLF88, FBF^{+}89, FOW^{+}85, FCH^{+}88, GDS87, HNR87, HBF^{+}88, HJS^{+}85, HMTF88, HTB88, ITY^{+}88, JM86, JS88a, KSSP86, LWM85, MWSF87, NR89a, NCB88, OB88, OCP86, PG86, PRB^{+}89, SMR87, SMA85, TL87a, Van87, VB85b, WMH^{+}88, MW88b, WM88]. Efficient [CH85, POCM88, DSHW85, TKK^{+}87]. egasyn [BNT^{+}87]. EGF [MBK^{+}86, WG86, MBK^{+}86, WG86]. EGF-stimulated [MBK^{+}86]. egg [BS88b, BW86, HBB^{+}89, OHK86, PHM85, PHTK87, SB85c, TSOC89, WBGV85, YB87, ZSE85]. eggs [BC89, BG86, BFJ^{+}85, BN85b, CWZ89, ER85, GK87a, GK87b, HTSH85, JWH85, KB85, LC88a, MHK^{+}85, Miy88, SB85a, SMS^{+}88, SR85b, SBG^{+}85, SW86, SE86, TJF86]. Ehrlich [KWPH89]. either [GLD^{+}88]. ejaculated [dCFB86]. ejection [RDJ^{+}86]. elaborate [SHL87]. elastase [SHGD85]. Elastic [HMP89]. elastin [BSBZ88, FBCQPR87]. elastin-derived [BSBZ88]. elastin/glycosaminoglycan [FBCQPR87]. Electric [GBH87, OH88c, TFBW85, CCK86, CFC89, CS86c, LP85b, MMM^{+}85, NSM^{+}87, Sow86, Wal86]. electrical [Rob85]. electrocyte [CCLC89, KCN^{+}89]. electrocytes [KCN^{+}86, WWW^{+}85]. electrolytes [NHKE88]. Electron [DMRB87, HVBB88, KLS85, PTW^{+}85, PCC85, Ake89, BTR^{+}86, CW87, FYRK88, FKP86, GDO^{+}85, GVD^{+}87, HR87, JS86a, JS86b, KRLU87, LGBE89, MP86c, MF87a, MMS89, MSS87c, NMKS85, PSM89, RBJ89, RBMM88, RGM85, SMH88, STGK87]. electropermeabilized [TN89]. electrophoresis [KHKR86, MSK^{+}87, SF88b]. Electrophysiological [Kro86, SGM^{+}86, TCBF^{+}85]. electroplaques [FMT^{+}87]. Electrostatic [BGBB89]. elegans [AC86a, AHC89, AE87, DDR^{+}89, EOM86, EBCO88, FW85, HW87, Hym89, Str86, WRS^{+}86]. elements [LB89, LKF87, MWSF87, SJ86, SALN88, SE86, WBD^{+}86]. elevated [BN85b, PRS^{+}85, SB85a]. elevations [Mac89a, VB88b]. elicited [HKM^{+}87, LC87, MSLR^{+}86, ZSSR86]. eliminates [JBG^{+}87]. elongating [CM86a]. Elongation [SW85, DLBH89, GAB86, HHY88, MMC88, OE85, WW89]. elongatum [HW85b]. embedded [LCBS89, Ris85, BTR^{+}86]. embryo [BPJBF85, CMDCC86, CS86b, CHLS88, CHG^{+}86, DRC^{+}86, DDH^{+}87, DN88, FGO^{+}85, FCCC^{+}89, FMJS89, FOW89, GCK^{+}88, HMF^{+}87, HBB^{+}89, HM89b, JLK89, JSCLP89, KL88, KL89a, KFA89, KSM89, KDBF86, LHHC85, MP88a, MS86c, NSY86, OD88, OSD85, SA85, TQN^{+}87, TOOK^{+}85, TDGE85]. embryo-specific [TOOK^{+}85]. embryogenesis [AKVF88, BR85, HW87, MG89, PBD^{+}89, SL89, SF89]. embryoid [GW87]. embryonal [BS87, KWS^{+}89, LMSM89, LLRR87]. Embryonic [HNR87, LGF86, TOOK^{+}85, AR88, AEB^{+}88, CHMB86, CB86, CL85b, CE85a, CE85b, DNI^{+}88, ET87b, FCVDH89, FBE85, GZHM87, JVG^{+}85, KSH87, LCE87, LHNF87, MCS85a, MACO88, MOP^{+}85, NKA^{+}87, PCC85, RE86, RDB^{+}86, SSC86, SIYT86, Sch85, SH86, SGS^{+}89, VJT89, VN88, WG85]. embryos [BBW86, CG89, FA85, GM86, KA86, MFA89, MASA89, MWSD86, NKA^{+}87, NB88b, NT86, RG88, SHBV89, Str86, TS88, TM87a, TM87b, Tok89, WSW85, WFW87, WG87b]. emerge [CL85b]. Emergence [DDBHL86, LB89]. emerges [PLA86]. EMP [HLU^{+}87]. encapsulation [UMHI85]. encode [FCH^{+}88]. encoded [CA86, DMC85, GM88, GLD^{+}88, HOPM^{+}87, SDS89, SBG^{+}85]. encodes [AY89, BMEG89, DS89, HHY88, LGC^{+}88, WLW^{+}88]. encoding [BS86, BFK87, CHH^{+}88, FTLSS88, HNNT88, HMP86, JPU89, LSC86, Rob89, SR86a]. encompasses [ABME88]. End [RGM86, AIJ^{+}89, BM86, CCMB87, GK87b, THT89]. end-capping [THT89]. End-to-end [RGM86]. endochondral [CRF^{+}88, GF85, PP89]. Endocrine [LMK88, BK85, BCK85, KNGH^{+}85, SC88]. endocytic [AK87, CM88a, GSS^{+}88, GBM89, GGH89, LHvF^{+}87, PPB^{+}89, RDSG86, SM88, SM89a, TLPTV85, Wil88]. endocytosed [EG89, JNA85, SGS87, SGGS88]. Endocytosis [BPP^{+}89, CVWZ89, HVR^{+}85, KSH89, SOB^{+}89, BLSK88, BD86b, CdCP^{+}89, DZ85, DCH86, EFR^{+}89, GF86b, GGR^{+}86, GDRB89, HA89, ICP^{+}86, JB86, JNA85, MH86b, PP88a, PBvTS88, PMJ87, RFMR86, SOPvD87, SR85c, Too85, WH85, WM85, Wat85, vGRB86]. endocytotic [WBW89]. endoderm [GW87]. Endogenous [BBH85, HMHWS86, MDW86, BHRC85, BR85, BSB^{+}86, CMNB88, LLSRB89, MGM87, NYRK87, OCP86, PSM89, VN88]. endogenously [LP85a]. Endolyn [CIK^{+}89]. Endolyn-78 [CIK^{+}89]. endoneuraminidase [RWS^{+}85]. endoplasmic [BWK87, BB89, BHAS85, BSH^{+}85, CC89, DS87, DS89, DRY89, FRG^{+}87, GORW88, GRME^{+}89, HFGD89, HOPM^{+}87, HBB^{+}89, HAM86, HGL88, Ibr87a, Ibr87b, JBG^{+}87, KA85, KR88, KS85b, KSH89, KJW89, LFC89, LKHR87, MVK^{+}88, NBS88a, NM89, NRL89a, PLA86, PTB^{+}85, PA88, RDS^{+}89, RKFN88, SCK^{+}89, SPL87, SALN88, SWC^{+}87, TCF86, TKFH89, VVL^{+}89, WCMG88, YMT85, ZR89]. endoproteolytic [ALL89, THFB87]. endosomal [CIK^{+}89, Pfe87, WH85, WWN87]. endosome [BSK89, DMM89, GGH89, MBK^{+}86, MBvS^{+}89, SFK^{+}89, YM87b]. endosome-lysosome [MBvS^{+}89]. endosomes [BGF86, FKH88, MN87, MSK^{+}87, MK87, MH86b, PPB^{+}89, RSF86, YM87a]. endosperm [BMB86, SL87a, VLD^{+}85]. endothelia [GM85c]. Endothelial [BO85, HHD^{+}86, LGW^{+}89, LMF^{+}86, PCP85, SMSR88, BSd^{+}88, CTS^{+}87, DTW^{+}85, DCL^{+}87, DCC^{+}88, DLG^{+}89, DF87, DS88, EWHP87, FKM87, FS86a, GCR^{+}86, GR89, GS85, GFA^{+}88, HTRBP86, HS85b, HMTF88, HBW^{+}87, JBdlTB87, KLHK88, KW86a, KKML88, LSAKO86, LS87a, MPT88, MMWD89, Mos88, MG86b, MUM87, MUH89, NKMK86, OD87, PVMO87, PSC^{+}89, PBMB85, PMR89, RJH^{+}87, RKA^{+}89, RG86b, SMR87, SR88, SR89, SKK^{+}85, SMW89, SBHN86, SGP88, TS88, TDL^{+}85, TAHC^{+}89, WD86b, WJSB88, YBWLM88, ZR87, dGRS87]. endothelial-type [LSAKO86]. endothelium [DW86, GFSS86, JBdlTB87, MBM85, MWSP87, PSSP88, SAB^{+}89, TKK86, WF86]. endplate [FC89a]. ends [BG86, FJMW87, GSB87, IF85, Pol86, SB85e, TWB88, WIS89]. energies [SKC85a]. energy [ABC^{+}88, GML^{+}89, SSZ88, TT89b, UP86]. energy-transmitting [ABC^{+}88, SSZ88]. Engelbreth [KKSK88]. engineered [PL85b]. enhance [PML^{+}89]. Enhanced [KGB88, LSAKO86, GB86a, HPO88, JBG^{+}87, MP86c, WKRH86]. Enhancement [GWCW85, ZBG89, SAB88]. enhances [FGE86, Goo89, NV89, TBK86]. enolase [WLW^{+}88]. enrich [GL85]. enriched [CS87, KGB88, LL86, MGK89, MCC^{+}89, PMV^{+}89, TACQ89]. ensheathment [SPS89a]. entactin [DCB^{+}88]. enteric [SL89, WLJM^{+}89]. enterocyte [GML^{+}89]. enterocytes [BJMK89, SCM87]. enterocytic [TRC^{+}87]. entirely [DRME89]. entry [SF89]. enucleated [VSR87]. enucleation [KKB89]. envelope [AB89, BS88b, BF86, CM88b, DH87, DJBD85, GB87a, GB87b, HDB88, KWH89, LPR^{+}89, LM85a, MT86, PLA86, RFG^{+}85, WKKS^{+}87, WS87, WN88, ZDH88]. enveloped [Bäc88, BS85]. environment [DMG^{+}89, SS89b]. enzymatic [RMS86]. enzymatically [AHM89, DVVT88]. enzyme [ACPFB^{+}85, BKN^{+}85, BNS85, BIES89, CSBS89, CCBKJ89, DCWJ89, EFR^{+}89, GF86a, GGR89, GSS^{+}85, GHL85, HFF^{+}85, HBB^{+}87, IS88, LA87, LHvF^{+}87, LSF89, LBR86, MSY^{+}87, MCD85, SS85, SEGS85, WW87a, WLW^{+}88]. enzymes [BNB^{+}88, BGHvF87, BGF86, CGD86, RWE^{+}88, WFL87]. Epidermal [BBBM89, SJ87, WR88, WG86, WBD^{+}86, AF87, BLSK88, BSK89, Bre89, CWOK87, CS86c, DHK^{+}85, DAJJ^{+}88, DBM^{+}89, DMH^{+}85, DCH86, DCB^{+}88, ESA86, GST^{+}87, GSDR89, GBH87, GGJ^{+}87, HGMS88, JFAT89, KSJF87, KFH^{+}86, Kol86, KTF87b, KF89b, LCD^{+}89, LCW^{+}89, LWR85, LR85a, LDKG87, LBP^{+}86, MS86a, MFKVK87, Mas85, MBK^{+}86, MBSMD86, NHKE88, OH88b, PS86, RFV85, Sch86b, SKDF88, WBW89, Wil88, YKSR89, ZSSR86]. epidermis [CKF89, KTF87a, NCB88, OMS89, TF86, YPG86]. epidermoid [MBK^{+}86, WBD^{+}86]. epilation [FJR87]. epiphyseal [KDFG89]. epithelia [BLA^{+}88, KOF88, Rap89, SSMG86]. Epithelial [ACEE87, AF87, AF89, AU89, AMWN89, AE88, BOW86, BBGI85, BMVB89, BF88, BVK85, BLA^{+}88, BTV^{+}89, BGH87b, CTJG88, DDB^{+}88, DF86b, EKH89, FGF88, FGA86, GF88, GJD^{+}88, Gor89, GGR^{+}86, GM85c, GS86b, GSG88, HSB^{+}85, HKH^{+}88, HHD^{+}86, JNR^{+}85, JRSB87, JRB88, JGYG85, KCCM85, KKGJ89, KBL^{+}86, KBB^{+}88, LS85, LCDRB89, MD85, MSQ^{+}89, Mat86, MH86a, MCC86, MG86a, NV86, NV87, NH89, OE85, PCK^{+}87, PN88a, PN88b, PN89, PHRM87, RJB86, RR88, RFV85, RBGF89, RJH^{+}87, RT88, SMVS^{+}86, SVSH^{+}88, SB88, SGS86, STE88, SAA89, SG88b, SRM89, SH86, TS88, TNG89, TB88b, TTSSH85, UPL^{+}87, VJT89, VSSGRB87, VSSRB87, VSSRB88, vMSvRS87]. Epithelial-mesenchymal [ACEE87, AE88, VJT89]. epithelium [AFMD89, BAD^{+}89, BOW85, CMB^{+}86, Gor85, Kol86, KMP88, MSV^{+}88, OGJ^{+}89, PN85, RFV85, RKWS86, RBAF89, SH86, ZBG89]. epithelium-type [RBAF89]. Epitope [DW89, DBWL87, KFS87, MS86b, MAG88]. Epitope-directed [DW89]. epitopes [CM88b, FTC^{+}89, FR86, FONW88, GDO^{+}85, JBdlTB87, MAG88, RMM89, WFR86]. equal [ZP85]. equivalent [BGS87]. Er/Er [FJR87]. erythroblasts [KKB89, MCG88b, MGW86]. Erythrocyte [CM86b, DH89a, MMM87, CKR^{+}85b, FDB85, GM85a, HdAC89, HSMM85, KMKS87, LDP87, SJS86, Sow86, SKC85a, SKC85b]. erythrocytes [CC85b, GB87b, HUA^{+}86, LHNF87, SYSP88, SB85d, WE89, WS89]. Erythroid [CSL87, CML85, HVL87, LL88, MNWL85, PL86, SL86b, WBVL86]. erythroleukemia [KWS^{+}89, PL87, PHR86, SR85c, VSR87]. erythropoiesis [HVL87]. escaping [HGMS88]. Escherichia [DBLS87, SCK^{+}89, WGPW89]. esophageal [BOW85]. Essential [HKTY89, HCG87, BB89, HPW89, KWH89, MGD88, NAS^{+}87]. established [RG86a]. establishing [GS86b]. establishment [VSSRB88, WLG85]. ester [BSB^{+}86, DJ89, GKS87a, GCGR85, LPR^{+}89, MDM88, MW86, PKN^{+}89, RRP88, Wal88]. ester-induced [GCGR85]. ester-treated [LPR^{+}89]. esters [RBKK85, RBKK87, SPO85, SYS85]. Estimation [vDSP^{+}88]. Estramustine [SWTB88]. estrogen [CMB^{+}87]. estrogen-induced [CMB^{+}87]. etch [ASF89, BO85, BCPC87, FAKVM87, GH85a, HKA^{+}87, HK88a, LC88a, NH87]. etched [SDS87]. ethyl [ATNH86]. ethylmaleimide [JWH85, WMBR89]. etiochloroplasts [CHK^{+}86]. eucaryotic [NB89a]. euchromatin [KdS86]. eugametos [HMdN^{+}88, KdWB^{+}89]. eukaryotic [BF86, SB87]. Euplotes [AKC^{+}86, OO87, OOCT89, RLP85]. eurystomus [AKC^{+}86, OO87, OOCT89]. evaluated [SMH88]. evaluation [CHD^{+}86, DCQB88, PW86, WBW87, WAF87]. even [TT88a]. event [DJ89, GKB89, WS86c]. events [GG85, HFB^{+}89, IGSGD87, MSLR^{+}86, RWE^{+}88, SM89b, SME^{+}89, TKFH89, WS86b, WHMA86, ZSSR86, dVOD^{+}89]. Evidence [AP86, ALWL89, CR88a, CA86, DI87, ES85, GTN86, KTF87a, KCN^{+}86, LL89a, LS89, MG86a, RFSL88, RMS86, Ste85, TSS89, TF86, UF89, VFB^{+}89, AWH^{+}85, AHRM85, BNLC85, BGM^{+}87, BSA87, BCKJW88, BPJBF85, BD86b, BICFA88, BMDN87, BNS85, CCA^{+}88, CMK87, CNHF85, DMD^{+}86, DMB88, EBCO88, GB85a, GF87, GWCW85, HDP^{+}85, HBPK87, HP87b, ITY^{+}88, JHP85b, KWW85, LCR^{+}87, LN88, LLSRB89, MBC86, Mit89, MVB85, NL86, PTW^{+}85, PLA86, PFPM86, PSSP88, SBS87a, SE85, SB85e, SPH87, SFD^{+}86b, SQM89, SALN88, TKK^{+}87, TDR88, TB87, THFB87, WW88, WRR87, YF87, YR87, dBBB^{+}86, CCL^{+}86]. evolution [LC85]. evolutionarily [GGR89]. Evolutionary [WJSB88]. evolving [SDS87]. Examination [AKC^{+}86, MGW86, WL86]. examining [KTF87b]. exceedingly [DMGS88]. excess [GKS87a]. Exchange [Wan85d, AST89, FBPS87, GWCW85, GCGR85, KW86b, Owe86, RGBM87, TL87b]. Exchangeability [MMW85b]. excision [KAB88, SS89a]. excitation [Van87]. exclusion [NFF86]. exerts [SC88]. exhibit [MG86b, TWB88]. exhibits [DNI^{+}88, SVF^{+}89, UPRS88]. Existence [BMB87, BMDN87, LLCA89]. exists [CPP^{+}87, PT88]. Exit [GPSM85, TTF87]. exocrine [AC87, BCK85, CCC86, CJ86, TKFH89, vZC87]. exocytic [BNB^{+}88, LHvF^{+}87, MDM88]. Exocytosis [MLK^{+}87, VSSRB88, ZSE85, BGM^{+}87, BC89, CWZ89, CHG87, CPM86, CJ85, GH88, LMW^{+}86, MF87b, NL88, STGK87, TJF86, VSKSP86, ZP85, vGRB86]. exocytosis-competent [ZP85]. exocytosis/endocytosis [vGRB86]. Exocytotic [PW86, HCG87, JWH85, MLK^{+}87, vAAA86]. exogenous [BBH85, CTR86, CWH^{+}86a, KG85a, MLK^{+}87, OCP86, SHS^{+}88]. exogenously [TIS^{+}86]. exon [PRB^{+}89]. expansion [MT86]. Experimental [GB89b, JS86b, SR85b, HKH^{+}88, MG85]. explains [Bas87]. explant [SR86a]. explants [CL85b]. export [vZC87]. exposed [LD86, MT86, Ros89]. Exposure [WKRH86, ATNH86, PW86, SDDZ^{+}89]. express [BM88, CPG^{+}89, MLR^{+}86, SHS^{+}88, TOOK^{+}85, WGA85]. expressed [AU89, CAG89, DVVT88, DRME89, EWAP85, GM85c, HERO^{+}87, HVL87, HSK^{+}89b, JMBM85, JFAT89, LRF88, LFS^{+}89, MG86a, MFM89, OHH^{+}89, Rog87b, SYH^{+}86, SMS^{+}85, TGNG87, TDR88, TF86, WS89, YHS88]. expressing [CCC89, CSM^{+}89, MNC86]. Expression [AF89, BLA^{+}88, BGK^{+}89, BJ86, CG89, CE85a, CE85b, CKF89, DHK^{+}85, DCPR^{+}89, DBLS87, FWP^{+}87, GFS^{+}89, HSB^{+}85, HFF^{+}85, HM89c, KTO^{+}89, KFHF85, KSH87, KMP88, LCE87, LL89b, MCG^{+}88a, MAK^{+}86, MC89a, MIHS86, NMPG88, NGI^{+}89, OLGT86, PB86, POD89, STH^{+}89, SSC86, SGM^{+}89, SCD88, TFS^{+}89, WGPW89, WF86, YKCR88, YKSR89, APBP87, AF87, AVP^{+}89, ACEE87, BB88a, BYB^{+}88, BTO88, BRD^{+}88, BS86, BP87b, CCP88, CL85a, CDMC88, CB86, CCM86, CCE87, CCK^{+}88, CML85, CHG^{+}86, DLG^{+}86a, DH87, DNM86, DMW88, DLC88, ES89, FJR87, FGA86, FGE86, FOW89, GLL^{+}86, GCK^{+}88, GZHM87, GKG88, GKF85, GF88, GCPE85, HD86, HNNT88, HS86b, HZS88, HGS89, HHR^{+}89, HL88, HCE88, HHMH87, KKT^{+}86, KLHK88, KSJF87, KKGJ89, KTF87b, KKG86, KDFG89, LCL^{+}89, LHK^{+}89, LSEN87, LBB^{+}88, LLC85, LC85]. expression [LVSC86, LHNF87, MACO88, MGM87, MVS88, MMRR86, MW88b, MG89, MNWL85, MAE88, MGW86, MVB85, NV89, NWE^{+}88, NT86, PFD^{+}89, PSK^{+}89, PC85, PBT^{+}86, PSGS85, QCKR85, RFV85, RTK^{+}86, RSP^{+}89b, SL89, SCBR85, SGS86, SBD^{+}89, SPS89b, SP85, STS^{+}89, SR86a, SBS87a, SYT^{+}89, SDR^{+}88, SKDF88, SS88c, SPO85, SEG^{+}86, TSI^{+}89, TBK86, TRT^{+}88, TT86a, TTBG87, VJT89, VKB89, VSSRB87, VBWC88, WVLC86, WR88, WHMA86, WTB^{+}89, WPBB^{+}85, WMR86, WB85, WBVL86, Wri85]. extended [KSL^{+}87]. extends [FONW88]. extensin [CV87a, MAG88]. Extension [CRR87b, HAK^{+}88, TI85, TI87, WMSB^{+}89]. extensive [NRL89a, SHL87]. extent [EG89]. external [GW86, RGPRB87]. externalization [PTW^{+}85]. externalize [HHES88]. externalized [BHRC85]. extra [GFS^{+}89]. Extracellular [BT86, HBSQ^{+}85, BNLC85, BS88b, BMGI85, BPE^{+}86, CSM^{+}89, CTR86, CRJM88, CTS^{+}87, CPHM86, CHH^{+}88, DW86, EMW^{+}89, FGO^{+}85, GDM^{+}88, GRvdM89, GW87, HNR87, HBC89, HB88a, HCE88, IF89, KSF^{+}88, KG85a, LSAKO86, LC88a, MPT88, MGHG^{+}86, MBSMD86, NSF^{+}89, NHKE88, OD88, PCK^{+}87, PJ87, RYFM89, RKA^{+}89, RLB^{+}89, RMS86, SVF^{+}89, SH87a, SKE86, SB85b, SG88b, SSK^{+}88, STF85, SEB89, SHL87, SH86, SCD88, TRB86, VGKF86, WCPK88, WS87, WE86a, WMH^{+}88, YKSR89, dGRS87]. extractability [SRM89]. extraction [AHRM85]. extracts [ASZ^{+}88, GK87a, LJC^{+}87, LM85a, SMA85, YB87]. extraembryonic [KL89a]. extrajunctional [SMH88]. extraocular [WPBB^{+}85]. extrasynaptic [CS86d]. extreme [SCFP86]. extrinsic [RO85]. extrusion [HMN87]. eye [DMRB87, OE85, SKK^{+}85]. eye-derived [SKK^{+}85]. ezrin [Bre89].

F [ABG^{+}89, BBV^{+}89, BR85, BC85, CB87, CB88, DG89, HF85b, HO85, MFA89, NJAS89, RG86b, SL87a, VB85b, WSW85, WL87, ZV89a]. F-9 [DG89]. F-actin [ABG^{+}89, BBV^{+}89, BR85, BC85, CB87, CB88, HF85b, HO85, MFA89, NJAS89, RG86b, SL87a, VB85b, WSW85, WL87, ZV89a]. F1 [EVG^{+}86]. F1-ATPase [EVG^{+}86]. F3 [GCR^{+}89]. F9 [KWS^{+}89, BGK^{+}89]. face [CCS^{+}86, HLU^{+}87]. facilitation [Bas87]. factor [ALCG88, ASCM89, AS86, BB87b, BMEG89, BBBM89, BLSK88, BSK89, BAG86, BRD^{+}88, BGK^{+}89, Bre89, BM88, BSG85, CWOK87, CRF^{+}88, CMP86, CLOH87, CR86, CWH86b, CWF^{+}89, DWL86, DKH^{+}85, DNB88, DAJJ^{+}88, DBM^{+}89, DLG^{+}89, DSTJ85, DMG^{+}89, DRME89, DFSK85, DKKK88, DCH86, DCB^{+}88, EG89, EWHP87, FBF^{+}89, FSM88, FTC^{+}89, FGE86, FOW^{+}85, GST^{+}87, GW88a, GGR89, GSDR89, GBH87, GBU^{+}86, GRCG86, GDC^{+}86, HLB^{+}88, HKM^{+}87, HTRBP86, Has88, HJS^{+}85, HMF^{+}87, HP85a, HKBT87, HGMS88, HSK^{+}89b, IF89, JS86a, JVG^{+}85, JS88b, JSCLP89, KF89a, KG85b, KC88, KOBLM88, KSJF87, KW86b, LCD^{+}89, LCW^{+}89, LSAKO86, LFR85, LWR85, LR85a, LWM85, LOG85, LTG85, LDKG87, LGC^{+}88, LBP^{+}86, LKOM88, MPT88, MCB85, Maj87, MFKVK87, MSD^{+}88, Mas85, MER^{+}86, MFB^{+}89, MTRR89, MBK^{+}86, Mos88, MQ89, MBSMD86]. factor [NR88, NR89a, NMPG88, NV89, NRL^{+}89b, NHKE88, OSH^{+}86, OH88b, OGF87, OGYK88, PCBD89, PML^{+}89, PS86, PMR89, PUW^{+}88, Rap89, RR88, RLJ86, RM89a, RE86, RYF^{+}87, RKA^{+}89, SSR89, SMR87, SMSR88, SR88, SR89, SJ87, Sch86b, SKK^{+}85, SLB87, SHGD85, SR86a, SC88, SGB^{+}86, SHL^{+}89, SHS^{+}88, SCP85, SRWdC87, SMW89, SGP88, TJ85, TL89, TRT^{+}88, TDL^{+}85, TFS^{+}89, TS86, TTLR89, VHO^{+}89, VRLG89, WMM86, WSM^{+}87, Wal86, WR88, WCB86, WBW89, WG86, WBD^{+}86, Wil88, WMH^{+}88, ZSSR86, ZR87, SKK^{+}85]. factor- [LR85a]. factor-activated [JS86a, VRLG89]. factor-alpha [GGR89, Mas85]. factor-beta [ASCM89, CWF^{+}89, FTC^{+}89, HMF^{+}87, KOBLM88, LSAKO86, LKOM88, MPT88, Mas85, MFB^{+}89, PML^{+}89, RR88, SR89, SRWdC87, TFS^{+}89, WSM^{+}87, WMH^{+}88]. factor-beta-induced [OGYK88]. factor-beta/alpha [MFB^{+}89]. factor-binding [GSDR89]. factor-induced [BBBM89, BM88, BSG85, DFSK85, FSM88, Has88, HP85a, KC88, MTRR89, VHO^{+}89]. factor-like [HTRBP86, ZR87]. factor-mediated [HGMS88]. factor-modulated [BSG85]. factor-producing [DRME89]. factor-promoted [BAG86, RLJ86]. factor-receptor [DAJJ^{+}88]. factor-regulated [LGC^{+}88]. factor-repressible [NR88]. factor-stimulated [IF89, LR85a, WBW89]. factor-treated [DKKK88]. factor-type [RYF^{+}87]. factors [BHSJ86, DFSK85, DN88, EG89, GST^{+}87, GABT86, GY86, HMBV^{+}88, JS86d, JS88a, KST^{+}89, KSSAQC89, MMT85, MMWD89, MR85a, MPB88, MTA86, RDSL89, SSRS88, Wri85, YSW^{+}85]. factors-beta [RDSL89]. fail [CC85b, PMC^{+}88, RYR^{+}89]. falciparum [HdAC89, HLU^{+}87, KSG85, SYSP88, WS89, HUA^{+}86]. falciparum-infected [HdAC89, HUA^{+}86]. family [BKA^{+}89, EBGS86, GCC^{+}89, GKZ89, GGA^{+}89, HNNT88, LRF88, MWSD86, PKS^{+}89, PSK^{+}87, Ran88, SKK^{+}85]. Farquhar [Rud88]. fasciae [Tok89]. fasciculation [RWF^{+}87, RSL^{+}89]. Fast [SBK85, BFAI87, BN87, MCS85b, NNJ^{+}87, SS87, TB89a]. fast-freeze [BN87, NNJ^{+}87]. fat [GHA88]. Fate [CNHF85, TMB85, JGYG85, KBB^{+}88]. fates [NL86, SALN88]. fatty [SV89]. Fc [DMGS88, HM89c, JNA85, MSW^{+}86, TSBW86]. Fc-receptor-mediated [MSW^{+}86]. feathers [CE85a, CE85b]. features [BDB^{+}87, FGA86, NB88a, SR87]. Female [DM87]. fenestrae [LMF^{+}86]. fenestral [BO85]. fertilization [BS88b, BN85b, ER85, HTSH85, LC88a, LBL^{+}85, MHK^{+}85, SB85c, SEB89, SW86, WKS85, WS87]. fertilized [SR85b]. fetal [AMWN89, BPMD87, DHK^{+}85, EWAP85, HSK^{+}89b, MCG88b, PMR89, Qua85b, STH^{+}89, STS^{+}89, WMH^{+}88]. fetoprotein [PBT^{+}86]. fetus [FJR87]. few [MNC86]. fgd [KSJV88]. fgf [MQ89]. fgr [NGI^{+}89]. fiber [ATNH86, CMM^{+}85, GB89a, GSTS86, GKBG87, RSL^{+}89, SSJ^{+}85, WSRR89, ZHES89]. fiber-like [ATNH86]. fibers [AT86a, Bis86, CL85b, DW86, DMRB87, ER87, HTGP85, LMD^{+}86, SS87, SP86, TIT89, WF86]. fibril [BT86, BFBL88, KSBB87]. Fibrillar [SYH^{+}86, DNI^{+}88, SCBHS^{+}89]. fibrillary [SQM89]. Fibrillin [SKE86]. fibrillogenesis [PM87a]. fibrils [BPE^{+}86, IKH87, KSBB87, KSL^{+}87, MEBV^{+}89, MGHG^{+}86, SKMB86, VMH^{+}88]. fibrin [SHL87]. Fibrinogen [DCL^{+}87, ABME88, FOW^{+}85, GLW^{+}89, IMP^{+}87, OGF87, SSKK89, TAHC^{+}89]. Fibroblast [NV89, SR87, BGK^{+}89, CS88b, CLOH87, DMG^{+}89, FBF^{+}89, HLB^{+}88, HB88a, IF89, JAT86, JSCLP89, KF89a, KG85b, LOG85, LTG85, LKOM88, MTRR89, Mos88, MQ89, NR89a, NMPG88, OH88b, PRS^{+}88, PMR89, RM89a, RGS88, RKA^{+}89, SMR87, SMSR88, SR88, SKK^{+}85, SKS^{+}87, SAB88, SYH^{+}86, YE87]. fibroblast-conditioned [LKOM88]. fibroblastic [DKB^{+}85, OH89, RBGF89]. fibroblastoid [BTV^{+}89]. Fibroblasts [Bre88, CPG^{+}89, GSS89, SFLM86, ABC^{+}88, AMG85, AR85, CJK85, CHH^{+}85, CS86b, CHLS88, DLH^{+}89, DWFT88, DNM86, FGO^{+}85, FSM88, GG87b, HNP^{+}86, HW86b, JLK89, JS88a, KTO^{+}89, KCR86, KP89, LSAKO86, LFFW89, LWT85, LCR^{+}87, LHHC85, LP85a, LRF89, LWMM^{+}85, MSE89, MNM87, MvdMH^{+}88, MMW85b, MMW85a, MWK89, MLM85, MC89a, MDW86, NSY86, PSS^{+}87, SRH^{+}87, SLHvF85, SRM89, TFBW85, TRT^{+}88, TLC85, UW86, VKB89, Wan85d, Wan85a, WS85, WE86a, WO86, YHS88]. fibroin [TKK^{+}87]. Fibronectin [DCC^{+}88, DNI^{+}88, GCT86, JAT86, LCR^{+}87, TTBG87, APBP87, AYY86, AYCY89, AAM^{+}87, ASA^{+}87, BPL87, BCC^{+}87, BJ86, CP87a, CCP88, CBS^{+}89, CSR^{+}86, CCM86, CWH^{+}86a, CK86, CS88c, DNB88, DAS^{+}87, DMH^{+}85, DRC^{+}86, FBF^{+}89, FCVDH89, GSS89, GFS^{+}89, GPSC86, Gri86, HPR85, HBF^{+}88, HDG^{+}85, HAK^{+}86, HAK^{+}88, KH85a, KCR86, KDBF86, LEW^{+}88, LBEM89, MHF86, MLM85, MLE87, MKD^{+}89, NHOY85, PL86, PL87, PJ87, PM87a, PON^{+}88, RLP^{+}87, RLB^{+}89, SSC86, SB88, SRH^{+}87, SBD^{+}89, SSW89, SKS^{+}87, SSK^{+}88, SSKK89, SYH^{+}85, SYH^{+}86, SHL87, SR87, TRM88, TDR87a, WC87, WCPK88, WGPH^{+}89, WTB^{+}89]. fibronectin-binding [AYY86]. fibronectin-coated [GPSC86]. fibronectin-like [CS88c]. Fibronectin-mediated [LCR^{+}87]. Fibronectin-plasma [GCT86]. fibronectins [MLC^{+}88]. fibrosarcoma [MLE87, NMKS85, WC87]. fibrosis [CWF^{+}89]. field [GBH87, OH88c, TFBW85]. field-directed [OH88c]. field-induced [GBH87, TFBW85]. fields [CS86c, LP85b, Rob85]. filaggrin [DHK^{+}85]. filament [AST89, BBBM89, BM86, CP87b, CJY86, CRLW86, CM85, DHL85, DD88, Fec87, FLH^{+}88, FW85, GM85a, GWM85, GB87a, HCHK89, HMP89, Ip88, JP85, JGYG85, KSJF87, Kol86, LFFW89, LGC^{+}88, LYG85, MGGP88, NPCL85, OB88, PL85a, PSK^{+}89, RFV85, Ris85, RPCB88, RSF^{+}85, RR86, SQM89, TFD87, TT88a, VS85, YLGG85, dSdSH^{+}88]. filament-associated [CRLW86, LYG85, NPCL85, RR86, YLGG85]. filament-binding [PL85a]. filament-bundling [Fec87]. filament-free [HCHK89]. filament-membrane [GM85a]. filament-severing [CJY86]. Filamentous [Try89, KEG88, Wan87, ABC^{+}88]. filaments [AF89, BOW85, BOW86, BFPS85, BPE^{+}86, CS86a, CB87, CPKJ87, DCB89, ESA86, EBCO88, FS88b, FC89b, FONW88, GB87b, GGJ^{+}87, HS86a, HCS89, HF85a, HF85b, HP87b, HMP89, IF85, JG85, KR88, KLS85, KTP87, LU87, LL89a, LCK88, LYG85, LAA85, MYH^{+}85, Pol86, RBAF89, SMVS^{+}86, SLB^{+}89, SL86a, SSS85, SDMF85, SAA89, SKL85, SP86, TB89a, TTC88, TP89, TL87b, TSK88, VC85, VC89, Wan85b, WW88, WS85, WM88, WE87, WCSR85, YIJ88]. filamin [SFD86a]. Filopodia [CS87, Fec87]. filter [BPP^{+}89]. filter-grown [BPP^{+}89]. filtration [DF86a]. fimbrin [YMM86]. firefly [GKS87b]. first [BA86, HBB^{+}89, TM87b]. fish [CS86c]. fission [CW89, HKTY89]. Five [DBE^{+}87, LLC85, DBD^{+}89]. Five-parameter [DBE^{+}87]. fix [PGH^{+}88]. fixation [BN87, BFPS85, NNJ^{+}87]. fixed [PSM89]. flagella [GTG85, GH85b, Goo89, GMLR85, KW88, LWM^{+}88, LAA85, MTG88, PF85b, SGH85]. Flagellar [SSH87, TI87, AWJ85, BWS86, BN85a, Bro87, CP87b, FS87, GACOH85, HW85b, KdWB^{+}89, LR85b, LS87b, MR85b, SS89a, SSS^{+}87a, SW87a, THM86, WMR86, WVCR89]. flagellates [FCL86, MLSW86, SW87a, UF89]. flanking [PW87, PL85b]. Flexural [WE89]. flight [ASLFP89, CD89, CF88a, DCB89, SCBHS^{+}89, TRCR89a, TRCR89b]. flow [CDTS85, DVHC89, GML^{+}89, JIFH85, MSK^{+}87, SYS85]. flowering [PKS^{+}89]. fluctuations [TLZ88]. fluid [DZ85, PT88, SYS85]. fluid-phase [DZ85]. fluorescein [PMJ87]. Fluorescence [BFRT87, Str86, BBP^{+}85, BSK89, CS89, DBE^{+}87, GTB85, GML^{+}89, JS86d, LWT85, Mit89, PBS85, WS86a, WAF87]. Fluorescent [HH86, KT85, SYH^{+}85, VCM88, CJO^{+}89, CH85, DWFT88, GMF85, HdAC89, KP89, LP85a, LPLT85, MN87, NFF86, PSM89, SC85a, SME^{+}89, SNSS87, SSS88, UP86, VME85, WG87a]. fluorescently [AT86a, HTSH85, MMW85a, SMPS86, WFW87]. fluorescyl [KL89b]. flux [Hep89, LSP86, Mit89]. fluxes [MSLR^{+}86, Owe85]. Focal [BRD^{+}88, Gri86, BSG89, DI87, DNI^{+}88, JLK89, MUH89, PHN^{+}88, SKS^{+}87, SW88, WRL86]. fodrin [BSNG85, GZHM87, JBTA88, MCA^{+}89, NV86, NV87, NH89]. fodrin-rich [JBTA88]. folding [DRM^{+}88, GHW^{+}88, RGW^{+}89]. follicle [CKF89, FH86, LOH^{+}86, RR86]. follicles [JP85]. followed [CGV^{+}86, CKR^{+}85b, KBB^{+}88, PSR^{+}88]. follows [BN85b, MMT85]. force [FS87, SPH87]. foreign [MK85c, SBK85]. Forest [CG86c, CMG86, KH85b, RG86a]. form [BLSL88, DN88, FBE85, HW88, HFG87, ISM85, KBPW89, LDA86, LPBW87, MSY^{+}87, MBH87, PTW^{+}85, PF85b, RBGF89, VBB85, ZG87]. Formation [CG86b, DLFE89, MJW86, MWK89, SPLW89, VH88, VSSGRB87, BAD^{+}89, BC89, BSK89, BT86, CS87, CCE87, CNL86, CB87, ESA86, EBBW87, FCCC^{+}89, FCH^{+}87, GB86a, GGJ^{+}87, GSG88, HBSQ^{+}85, HG89, HF85a, HG88, HA89, HKTY89, HTB88, JBG^{+}87, KTO^{+}89, KMKS87, LM85a, MBKR89, MDVG^{+}85, MGW86, RGS88, SFLM86, SKS^{+}87, SFD^{+}86b, TM87b, TT86b, VG86b, Wan85b, WN88]. formed [CF88b, HF85b, PLA86, PF85a, SM88, TIS^{+}86]. former [PL85b]. forming [JMBM85]. forms [AED^{+}89, ASCM89, AE88, BL85, Bro85b, CDB87, CMGS86, FBE85, GB85a, GM88, HL88, HCE88, KSL^{+}87, KBK86, KMY88, RWE^{+}88, SHP85a, SHP85b, WGA85]. formyl [NJAS89]. formylated [PZBD87]. fos [SHL^{+}89, SDR^{+}88]. fos/jun [SHL^{+}89]. found [ACPFB^{+}85, ASLFP89, MBM85, SDNWA87]. four [GKS88]. fraction [PMV^{+}89, ZHE85]. fractionated [SG86]. Fractionation [Den88, VB85a, BFSH87, DMM89, FKP86, SHP85b]. fractions [CS88b]. fracture [CBIO85, KdS86, KBB^{+}88, STGK87, SP86]. fracture-label [KdS86]. fractured [KdS86, TB85b]. fragment [DBLS87, GBU^{+}86, GW88b, IKH87, JFP89, LE86, MLM85, SQM89]. fragmenting [ABG^{+}89]. fragments [BH86a, BBV^{+}89, BMB86, BGBB89, CB88, FHE88, LBW89, PJFB89]. framework [IBSS85, IS88, KS86]. Free [HC87, TLP^{+}87, BWK87, BGF88, BN85b, CRC^{+}87, CB88, DSGS87, HSS85, HCHK89, KM87, LMW^{+}86, LM85a, LR89, MMMP87, MSK^{+}87, MP88b, MBvS^{+}89, RCC^{+}88, Sch85, SSY^{+}86, SMS^{+}88, SHL87, SFD^{+}86b, SG86, TWB88, TB89b, VB88b, WB86, WE86b]. freedom [AYG85]. freeze [ASF89, BN87, BO85, BCPC87, CBIO85, GH85a, KBB^{+}88, LC88a, NH87, NNJ^{+}87, SDS87, STGK87, SP86, TB85b, TTM88]. freeze-etch [BCPC87]. freeze-etched [SDS87]. freeze-fracture [CBIO85, STGK87, SP86]. freeze-fractured [TB85b]. freezing [Heu86]. frequencies [WOP^{+}88]. freshwater [KS86]. Friend [PHR86]. frog [BFJ^{+}85, BN85b, CS86a, KSH87, MYH^{+}85, NPRR86, VJF^{+}88]. frozen [BKR86, BD89, ROCK86]. fructose [Cla87]. frustrated [KM87, TSBW86]. fucoid [BR85]. Full [GW88a, HW88]. Full-length [GW88a, HW88]. function [AYCY89, BGK^{+}89, BJ86, CCE87, CNB89, DK86, DS88, GDS87, GLC88, HMBV^{+}88, HB88b, HTB88, JYG86, JYC87, KEG88, KGB88, KKB89, LCBS89, LBL^{+}85, MD85, NGI^{+}89, OB88, RFMR86, RR87, SNCY87, SN89, SFS^{+}87, SCD88, TDR88, TNG89, TRC^{+}87, VS85, WCPK88, WSJ85, ZDH88]. function-associated [DS88, LCBS89, RR87]. Functional [DHH89, DNM86, DRME89, FHE88, KHC^{+}86, MGM87, MWS85, TT88b, WG87b, ALWL89, APS87, BSG89, CPP^{+}87, CPG^{+}89, CG86b, DWT^{+}89, DBLS87, FR86, FCH^{+}88, GDD88, GJD^{+}88, GDRB89, GS86b, HERO^{+}87, KGHK^{+}87, KJY89, LBO^{+}85, LC88b, MSK^{+}87, MGWM85, PUW^{+}88, RDSG86, STE88, SRB88, TKP^{+}88, TCBF^{+}85, WFR86, YJL^{+}89]. functionality [BSDH86]. Functionally [RVS88, BGS87, MGD88, May89, RBGF89, WMUP^{+}85]. Functions [JASP88, EVG^{+}86, HCE88, LD88, OSP^{+}85, PG87, PON^{+}88, PRS^{+}88, Sch86a, WFW87]. fungi [MH88]. fura [CJO^{+}89, MMMP87]. fura-2 [CJO^{+}89, MMMP87]. furrow [BC89]. Further [JYG86, TKK^{+}87, GBU^{+}86, MBC86]. fused [EMH^{+}87]. fusiform [ZMH89]. Fusion [KBB^{+}88, SM89a, AJAW86, Bäc88, BDB^{+}87, ECZ86, EZBW88, EZWB88, GDYW86, Gla85b, GGH89, KWW85, Knu85, MLK^{+}87, PHTK87, RFSL88, RC85a, RC85b, SM88, SME^{+}89, SPL87, TTGFC85, VSKSP86, WMBR89, ZSE85]. fusion-resistant [RC85a]. fusogenic [CG86c, KH85b, Sow86].

G [BWK87, BF88, CHG87, DJM^{+}89, DKHB87, DRM^{+}88, FGW85, GB85a, MKK89, NRTE89, PFS85, SLB^{+}89, SBG89]. G-protein [NRTE89, SLB^{+}89]. G-proteins [CHG87]. G/Q [DJM^{+}89]. G/Q- [DJM^{+}89]. G1 [CLOH87, LTG85, PRS^{+}85]. G2 [GL85, OMM87]. G2-M [OMM87]. G2-phase-specific [GL85]. Galactocerebroside [JMBM85]. galactolipid [JM86, SCGS85]. galactosamine [TTW88]. galactose [OW86]. galactosyl [AK89]. galactosyltransferase [CW89, ES89, LBL^{+}85, LS87c, RLB85, RVS88]. Gallus [Hut87]. gamete [MSY^{+}87, RFSL88]. gametes [HMdN^{+}88, PG87, SKC^{+}86]. gamma [CPG^{+}89, FKM87, FTC87, GBB89, KW88, LWSF85, LD88, OKLB86, Ros89, SL89, TCK87]. gamma- [CPG^{+}89]. gamma-Aminobutyric [TCK87]. gamma-enteric [SL89]. gamma-interferon [FKM87]. gamma-subunit [LWSF85]. gamma-thrombin [FTC87]. ganglion [JB87, NL86, RBG85, RF87, RRF87]. ganglioside [SYH^{+}85]. ganglioside-deficient [SYH^{+}85]. gangliosides [BLK^{+}88, FSFL88, GPSC86, SYH^{+}85, SYH^{+}86]. Gap [KL89a, MG85, BPG87, BKPG89, DYT^{+}87, GKG88, GPJ88, KMR^{+}88, KG86, MBL89, MG86a, Pau86, SGM^{+}86, SFS^{+}87, TLD^{+}89, WG87b, YJL^{+}89, ZKR^{+}88, ZHE85, ZN89, SV89]. GAP-43 [SV89]. gastric [MRD^{+}89]. gastrulating [KL88, KL89a]. gastrulation [CG89]. gated [LSDP86]. GD2 [CK86, CPHM86, CPP^{+}87]. GD3 [CPHM86]. gel [DF86a, DT89, Gla85a, GG87b, KHKR86, ZMH89]. gel-phase [Gla85a]. gelation [NRL^{+}89b]. gels [CF88b, GG87b, UW86]. Gelsolin [Bry88, HVL87, AKVF88, BTL^{+}86, BC85, CHD^{+}86, CJY86, CBS^{+}87, CLF^{+}88, HCS89, HB86, KMY88, KJY89, LJC^{+}87, WGPW89, YIJ88]. gelsolin-actin [BC85]. gelsolin-like [BTL^{+}86]. Gene [MSC^{+}85, SRPS86, AY89, APBP87, ASLFP89, BB88a, BMEG89, BHAS85, BCK85, CSH89, CKR85a, CNB89, DLC88, FKC87, GD89, GF88, GLD^{+}88, HBPK87, HD86, HD88b, HMBV^{+}88, HNNT88, HZS88, HHY88, HS87b, HML88, JFAT89, KC85, KG89, KSFL89, KW86b, KKG86, KBK86, KMP88, LRF88, LCL^{+}89, LFS^{+}89, LSC86, LHNF87, MACO88, MAK^{+}86, MVS88, MMRR86, MWSD86, MVB85, MVB87, NR88, NMM^{+}86, NV89, OMM87, OGF87, PB86, PBT^{+}86, RTK^{+}86, RE85, RSP^{+}89b, Rog87b, RS88, SCK^{+}89, SHC87, SR86a, SS88b, SDR^{+}88, SPO85, TGNG87, TYP^{+}87, TMP87, WG87b, WTB^{+}89, WVCR89, WLW^{+}88, WRR87, WB85, YKCR88]. gene-related [FKC87, KMP88]. General [LAC^{+}85, PV85]. generated [BCC^{+}87, CBS^{+}89, FS87, TTC88]. Generation [GLC88, GKB89, JPM85, IKH87, SPH87, TM87b, Tok89]. generation-degeneration [IKH87]. genes [CL86, CDMC88, HS86b, HMP86, IGSGD87, JPU89, MSC^{+}89, MGWM85, PB86, PGH^{+}88, RDS^{+}89, RS88, SDS89, TSI^{+}89, TS89, WPBB^{+}85, WMR86]. Genetic [BGH^{+}87a, DDR^{+}89, FCH^{+}87, FL89]. Genetics [LC85]. genital [HBB^{+}87]. Genomic [KMY88]. geometry [QS87]. germ [HBSQ^{+}85, KdWB^{+}89, Kub87, LWM85, WSM86]. germinating [TB85a]. GH3 [SS88c]. GH4C1 [PS86]. ghosts [Sow86]. Giant [KP85, KC85, KTT^{+}86, KS86, TTKM86]. giardin [PAHC89]. giardins [PAHC89]. Giarida [PAHC89]. Gist [Rud88]. gland [HMG^{+}86, KNGH^{+}85, TKK^{+}87]. glands [AC86b, CCC86, DLC88, FLBP^{+}86, KC85]. glass [GTB85]. glia [APS87, FGE86, GY86, GE88, PP88b, TDGE85, GABT86]. glia-promoting [GY86, GABT86]. glial [CR88a, GABT86, GY86, GFHDD89, RBAF89, SQM89]. glial-derived [CR88a]. Gliding [AWH^{+}85]. Glioblastoma [PCS88]. gliogenesis [GABT86]. glioma [ZV89b]. globin [SPL87, SS89d, VSR87]. Globular [NPRR86, NFW89]. Glomerular [KBO^{+}88, Abr85, AP86, HHD^{+}86]. glomerulus [LE86]. glucocerebroside [LP85a]. Glucocorticoid [CKR85a, FJY86, HAB^{+}88, MVS88]. Glucocorticoid-modulated [MVS88]. Glucocorticoid-regulated [FJY86, HAB^{+}88]. Glucocorticoids [LMWG85]. glucose [BGL^{+}88, SHL^{+}89, SE86, TL89]. glucose- [SE86]. glucosidase [LBR86]. glucuronidase [BNT^{+}87, GF86b]. glutamic [VKP89]. glutathione [BSY86, HKM^{+}87]. GLY [LWH88, BLK^{+}88, CPP^{+}87, DRP87, DAS^{+}87, DCC^{+}88, DLG^{+}89, GAPR88, GGA^{+}89, SKS^{+}87]. glyceraldehyde [NB88b]. glyceraldehyde- [NB88b]. glycine [TCP^{+}85, TCK87]. glycoalbumin [PSSP88]. glycocalyx [SC85b]. glycoconjugate [RMS86]. glycoconjugates [AdS85, HFB87]. glycogen [CMHC85]. glycohydrolase [ACPFB^{+}85]. glycolipid [SOB^{+}89]. glycolipid-binding [SOB^{+}89]. glycophorin [CRJM88]. glycophospholipid [CWW89, LCDRB89]. glycopolypeptides [KTF87a]. Glycoprotein [CBCV85, PON^{+}88, BAHE86, BSNG85, BSK86, BP87b, BK85, BLK^{+}88, CTS^{+}87, CIK^{+}89, DH87, DIF^{+}88, EWAP85, FTLSS88, FCK87, FJY86, Fra87, GWG85, GGMH86, GSF^{+}88, GWI^{+}87, GDRB89, GM85c, HDB88, HSWH86, IHJ88, JBdlTB87, JFAT89, KB88, KJW89, LH86, LSF86, MR87, MS88a, MS88c, NBS88a, NRL89a, OS88a, OPN^{+}85, PPG85, PHRM87, PSK^{+}87, PTB^{+}85, PMR86, Ran88, RJH^{+}87, RFG^{+}85, RPRS89, RSP^{+}89b, RS85, RG86a, SNCY87, SKE86, SHP^{+}87, SCC87, SRW^{+}89, SWC^{+}87, TTW88, Tra88, TACQ89, UPL^{+}87, VMT^{+}88, WESP85, WG88a, WE86b, WBB89, WL87, dBBB^{+}86]. glycoprotein-monoclonal [IHJ88]. Glycoproteins [LGM^{+}85, BWS86, CRR87b, CMW^{+}85, Col86, DBK87, GH85a, GGR^{+}86, GP86, HAB^{+}88, HT86, HHR^{+}89, HSS^{+}87, KGRP85, KA85, KdS86, KTF87a, Knu85, KSM89, KdWB^{+}89, LLSRB89, NVB^{+}88, NKMK86, PKS^{+}89, RWF^{+}87, RIPS85, RSF^{+}85, SHC87, SR86b, SSG87, TDR87a, TB85b, TLP^{+}87, TSM87, WCPK88, WO86, WSH85]. Glycosaminoglycan [PRS^{+}85, FBCQPR87, LEW^{+}88, LCR^{+}87]. glycosaminoglycan-binding [LCR^{+}87]. glycosaminoglycan-deficient [LEW^{+}88]. glycosaminoglycans [MCB85, SFS^{+}87]. glycosomes [Cla87]. glycosylated [BDS87, GKS87a, HL88]. glycosylation [CMB^{+}87, CSFS89, DBK87, ES86, HD86, JAT86, SSK89, WMM86]. glycosyltransferase [CSFS89]. glyoxylate [DCWJ89]. glyoxysome [TB85a]. glyoxysome-peroxisome [TB85a]. GMP [SMS^{+}85]. GMP-140 [SMS^{+}85]. Gn [BBB88]. gold [DF88b, HVBB88, JSA^{+}89, KSKW^{+}85, MM87, PSBN85, WG88a, BTR^{+}86, RFH85]. gold- [PSBN85]. gold-labeling [JSA^{+}89]. golden [Miy88]. goldfish [CAG89]. Golgi [AdS85, AK86, ALL89, BSR^{+}89, BH87, BWK87, BMEG89, BB89, BGL^{+}88, BSLS^{+}89, BGF86, CC89, CSFS89, CBYK88, CNB89, CJH85, CdCP^{+}89, DMD^{+}86, DRY89, DK88, FGW85, FFH88, GB85a, GPSM85, GFB^{+}89, HAB^{+}88, HHS^{+}85, LP85a, LKHR87, LPBW87, LBW89, MR87, NBS88a, NM89, NVB^{+}88, PSM89, PJFB89, PMC^{+}88, RFH85, RKFN88, SPF87, SFB^{+}88, SR86b, SGGS88, TPBB85, TTF87, TT89b, WMM86, WMBR89, WDF86, YMT85, YBBS87, vDTP^{+}86, vDSP^{+}88]. Golgi-derived [AdS85]. gonadotrophin [MB88]. gonadotrophs [RFZH85]. Gonadotropin [CMR^{+}87, MSC^{+}89, RPW^{+}89, WB85]. gonads [Str86]. Gonyaulax [JIFH85, NNJ^{+}87]. governing [MACO88]. GP140 [TLC85]. gp160 [HDB88]. gp180 [BSNG85]. gp80 [KWLS88]. GP85 [KB88]. gp90MEL [GR89]. GPIa [TH89a]. GPIIb [DBC^{+}87]. graded [JHP85b, Mac89a]. gradients [ET87a, FMG89, VES85]. granular [BGN^{+}88, FLBP^{+}86, HHSF85]. granule [BKN^{+}85, BC89, CCC86, CWZ89, CHLS88, HLL89, JGAR89, LSF89, MBD89, RH86b, SMS^{+}85, VS89]. granules [AC86b, BGN^{+}88, BPKS86, BCK85, BCMK87, FZ85, HDP^{+}85, HFZM87, LMK88, NRTE89, OMS89, PW86, POCM88, RWP^{+}89, RNW89, SSKK89, TT86b, TB87, THFB87, vZC87]. granulosa [COMS87, PSBN85]. greater [KHB88, LGF86]. green [HW85b]. GRGDS [SW88]. ground [VS85]. group [CCL^{+}86, DAJJ^{+}88, DJM^{+}89, RDB^{+}86, SCRA87, SSG87]. growing [DF88a, FBCQPR87, RCC^{+}88]. grown [BPP^{+}89, KG85b, KMM88]. Growth [CLOH87, HW86b, KST^{+}89, KSSAQC89, KDIG89, MMWD89, vdPdPR86, ALCG88, ASCM89, AS86, BBBM89, BLSK88, BSK89, BAG86, BDS88, BRD^{+}88, BGK^{+}89, Bre89, BD89, BM88, BSG85, CS88b, CWOK87, CRF^{+}88, CR85, CWH86b, CWF^{+}89, DKH^{+}85, DDB^{+}88, DNB88, DAJJ^{+}88, DBM^{+}89, DSTJ85, DF87, DMG^{+}89, DFSK85, DKKK88, DCH86, DCB^{+}88, EG89, EWAP85, FC86, FBF^{+}89, FSM88, FTC^{+}89, FS88b, FGE86, FZ85, GST^{+}87, GGR89, GSDR89, GBH87, GABT86, GY86, GB86a, GBB89, GBBS89, GRCG86, GP86, GDC^{+}86, HLB^{+}88, HKM^{+}87, HPW89, HTRBP86, Has88, HJS^{+}85, HS85b, HMF^{+}87, HP85a, HKBT87, HGMS88, HSK^{+}89b, HB87, HP85b, IF89, JS86a, JS88a, JSCLP89, KF89a, KG85b, KC88, KOBLM88, KSJF87, LCD^{+}89, LCW^{+}89, LSAKO86, LWR85, LR85a, LWM85, LOG85, LTG85, LDKG87, LB89, LGC^{+}88, LS87a, LBP^{+}86, LKOM88]. growth [MPT88, MCB85, Maj87, MCG^{+}88a, MFKVK87, MSD^{+}88, Mas85, MMY88, MER^{+}86, MNM87, MFB^{+}89, MMT85, MBL89, MTRR89, MBK^{+}86, Mos88, MQ89, MIHS86, MBSMD86, MTA86, NR88, NR89a, NMPG88, NV89, NHKE88, OSH^{+}86, ODD85, OH88b, OD87, OLGT86, OGYK88, PML^{+}89, PS86, PMR89, Rap89, RR88, RLJ86, RDSL89, RM89a, RFG^{+}85, RYF^{+}87, RKA^{+}89, SMR87, SMSR88, SR88, SR89, SJ87, SMC89, Sch86b, SKK^{+}85, SLB87, SK85, SS88a, SHGD85, SR86a, SGB^{+}86, SAB88, SHS^{+}88, SV89, SGBF87, Sny89, SRWdC87, TJ85, TL89, TRT^{+}88, TDL^{+}85, TFS^{+}89, TS86, TP89, TTLR89, TNG89, TSK88, UPRS88, VHO^{+}89, VRLG89, WSM^{+}87, WESP85, WR88, WBW89, WG86, WBD^{+}86, WO86, Wil88, WMH^{+}88, ZSSR86, ZR87]. growth-dependent [WESP85]. GRP78 [NRL89a]. GRP78-BiP [NRL89a]. gruberi [LRF88, MLSW86, SW87a]. GTP [BSR^{+}89, BB89, CG86b, KSJV88, NM89, SN89, TJF86, WCMG88]. GTP-binding [BSR^{+}89, CG86b, KSJV88, NM89, TJF86]. Guanine [TN89, HCG87, Miy88, PZBD87]. guanyl [JBTA88]. guanylate [Van87, WMV86]. guinea [CPM86, CMK87, PHM85]. gut [AE88]. gyration [TI87].

H [GWCW85, GCGR85, RT86, AC86b, BGH87b, Owe86, RYR^{+}89, TBK86, WLG85, WPW^{+}85, ZH86, vAAA86]. H-1 [WLG85, WPW^{+}85]. H-2 [TBK86, ZH86]. H1 [BLSL88, RSR^{+}88, VMT^{+}88, VB88b, WKR86]. H2 [BLSL88]. H3 [SS88b]. H4 [CL86, RCC^{+}88, SS88b]. Haemanthus [BMB86]. hair [CKF89, Hep85, HC87, JP85, LOH^{+}86, RR86, TT88a, TTC88, TTS^{+}89]. hair/nail [LOH^{+}86]. half [FONW88, RDJ^{+}86]. half-sarcomere [FONW88]. half-spindle [RDJ^{+}86]. halt [FRG^{+}87]. hamster [BHS88, BJ86, CCC89, CSD89, CDTS85, FBPS87, GWCW85, GBM89, HHR^{+}89, KBK86, LEW^{+}88, MGM87, MSC^{+}85, Miy88, NMPG88, RFMR86, SC86, SFK^{+}89, SBD^{+}89, SK88, TASK86, TB85b, YM87a, YM87b, YLGG85, YHS88]. handedness [MSR^{+}86]. Hansenula [DVVT88]. haptotaxis [TRL87]. hard [LOH^{+}86]. harvesting [CHK^{+}86, DSA86, HM89a, KHC^{+}86, LA87, LM85b, LU88, SDSO86, SCGS85]. HCl [KDB87]. HCO3 [SS89b]. head [CCA^{+}88, GWM85, NFW89, PSP85, RSP89a]. head-tail [RSP89a]. headed [BFAI87, Toy87a]. heads [LCK88]. healing [DBE^{+}87, FCVDH89, MHL88]. heart [BPG87, GSB85, GCPE85, HW86a, KF89a, YJL^{+}89, ZHE85, ZSB86]. Heat [LCL^{+}89, MP89, MWSF87, Sch86a, VKB89, dJHMB86, BGT^{+}89, BFA88, CRR87a, CL85a, CS86b, DH89b, HPO88, KLHK88, MLPF89, MW88b, MPB88, NSY86, NCB88, PBK88, RL89, SCK^{+}89, SAA89, SGBF87, WS85, WM88, WSRT89]. heat-shock [CRR87a, HPO88, MW88b, RL89, WS85, WM88]. heavy [AE87, BF86, BHK86, CB86, CF88a, DBWL87, DBK87, EPJBBW86, FL89, GSB85, HBPK87, HBKK87, ICS88, KKT^{+}86, KW88, KTP87, MPBZBE89, MCS85b, MTG88, OB88, PFRS89, PBvTS88, PWS^{+}89, SF88a, SSV^{+}85, SP89, TB89a, Toy87b, WE87, WPBB^{+}85, ZSB86]. heavy-chain-like [BF86]. HeLa [CRR87a, CR87, HFF^{+}85, LAC^{+}85, LPBW87, LBW89, MLPF89, VB88b, WKR86]. helical [RPCB88, SDS89, SQM89, WCSR85]. helix [MSR^{+}86, BOW85]. helper [LSDP86]. hemagglutinin [BDWH88, CDB^{+}86, DRSG85, DSG86, FRG^{+}87, GHW^{+}88, GDYW86, HGS89, HBHL^{+}89, MCG88b, NRL89a, RIS87, SME^{+}89, SRM89, WW87b]. hemagglutinin-induced [SME^{+}89]. hemagglutinin-neuraminidase [NRL89a]. hemagglutinins [RDSG86, RGPRB87]. hematopoietic [MMK^{+}89, WVLC86]. hemeundecapeptides [GS85]. hemidesmosome [KKGJ89]. hemidiaphragm [CMM^{+}85]. Hemodynamics [HBW^{+}87]. hemopoietic [GCT86, SIYT86]. Hep2 [MDVG^{+}85]. Heparan [JNR^{+}85, SKU^{+}85, BMGI85, BFL88, CE89, CSG85, Cou87, FC86, FRR85, GSS89, GLK^{+}89, HvdSdC^{+}89, JRB88, KKSK88, MOT^{+}85, RBG85, SMSR88, SSC86, STE88, SABVK89, SF87]. Heparin [CPC^{+}89, GG87b, MB85a, BB87d, CCL^{+}86, CSK^{+}88, CCK^{+}88, CG86a, CLC^{+}86, FBF^{+}89, MCB85, MFB^{+}89, RE86, SLB87, TDL^{+}85]. heparin-binding [CSK^{+}88, CG86a, CLC^{+}86, RE86, SLB87]. heparin-like [MCB85]. hepatic [CMHC85, CWF^{+}89, ZN89]. hepatitis [GORW88, SLG88, TTW88]. hepatoblasts [BPMD87]. Hepatocyte [STS^{+}89, BBBM89, BBH85, BFSH87, BHSJ86, CMW87, FOW^{+}85, GPJ88, HCH85, HBB85, MB87, OGF87, RR87, SFJU86, TKK86]. hepatocyte-stimulating [BHSJ86, OGF87]. hepatocytes [BPMD87, CJK85, CSGC88, DYT^{+}87, DCH86, FC86, HHS^{+}85, IGSGD87, KSC^{+}89, MA85, MH86b, OW86, PAF88, STS^{+}89, SGM^{+}86, SHL87, SKU^{+}85, SHP85a, SHP85b, SBC^{+}87, TLD^{+}89, WH85, YMT85]. hepatoma [BLSL88, DWL86, HAB^{+}88, LSP86, LKHR87, SGG^{+}89, SDZB^{+}85, ZBSS^{+}87]. HepG2 [SGS87]. herbimycin [KWS^{+}89]. Hereditary [WTBBW85]. herpes [MBLS87]. heterodimer [HMS^{+}89]. heterodimeric [TDR88]. heterodimers [DBWL87]. Heterogeneity [BKBR88, BBP^{+}85, CMM^{+}85, GVV85, GSS87, JYG86, Rob87, SHSvdM89]. Heterogeneous [Cou87, DMD^{+}86, HLL88, LM87, MLR^{+}86, MLPF89, MSS^{+}89, PRSGD89, YF87]. Heterokaryon [CH87]. heterokaryons [PB86]. Heterologous [RDSG86, HF85a, LDKG87, WVLC86]. heterotypic [HF85a]. heterovalent [SFJU86]. Hexabrachion [ET87b, LGBE89]. hexagonal [LDP87]. hexagonally [HH87]. hierarchical [BSA87]. Hierarchies [BS88b]. High [CWZ89, DJM^{+}89, GLW^{+}89, NYRK87, STE88, BMR85, BJMK89, BA87, DAJJ^{+}88, DBM^{+}89, DNI^{+}88, GDO^{+}85, GKS87a, GR89, GRCG86, HTT89, JBdlTB87, KTT^{+}86, LN88, MCS85a, Mas85, NB88a, SCRA87, SSMG86, SBG89, TT85, Wil88]. high-affinity [GRCG86, Mas85]. High-mobility [DJM^{+}89]. high-molecular-weight [KTT^{+}86]. Higher [AY89, BMB86, BF86, BIES89, CBL85, DHL85, LA87, SL87a, SFW87, VLD^{+}85]. Highly [RE85, CML85, May89, MGK89, RCC^{+}88, SW87b, TCF86]. hinge [HBPK87]. hippocampal [BBB89a, GB89b]. Histamine [RG86b, BGM^{+}87, VB88b]. histidine [HUA^{+}86]. histidine-rich [HUA^{+}86]. histochemical [FTC^{+}89]. histocompatibility [BHB^{+}87, CVWZ89, WFE89, WE86a, WSH85]. Histone [PFTV89, SS88b, CL86, CRC^{+}87, LLCA89, RCC^{+}88, RSR^{+}88, VMT^{+}88, WLG85, WPW^{+}85, WKR86]. histones [CRC^{+}87, LLCA89, POW^{+}87]. histotypic [LBO^{+}85]. HL [BS86]. HL60 [ICP^{+}86]. HMG [DJM^{+}89, SCRA87, WBDR88]. HMG-I [DJM^{+}89]. HN [PHL89]. HNK-1 [KJF^{+}88, PSK^{+}87]. Holm [KKSK88]. Homing [SIYT86, BNL89, PNB89, SRB88, WJSB88, YSR87, YBSR87]. homodimers [DBWL87]. homologies [LH86, MMRR86]. homologous [APF^{+}88, BPG87, GCBH86, HF85a, SYT^{+}89, WG86, ZHE85]. homologue [BNL89, HPW89, RSL^{+}89]. homology [HML88, SCK^{+}89, SHC87, TH89a, DCB^{+}88]. homophilic [SCC87]. hormonal [CKR85a]. hormone [CMR^{+}87, FZ85, NR89b, PAGPLM86, RFG^{+}85, SK85, TNG89, VB88b]. hormone-viral [RFG^{+}85]. hormones [MLR^{+}86, OJHAH89]. horseradish [AK87, SYS85, TAMR85]. host [BDS85, HMP85, HLU^{+}87, SYSP88]. host-cell [HMP85]. hours [MNC86]. however [TWB88]. HRP [HUA^{+}86]. HSP27 [LCL^{+}89]. hsp70 [KCK^{+}87, KMP88, MWM89, MM89]. hsr [BGT^{+}89, FBP89]. hst [MQ89]. hst/K [MQ89]. hst/K-fgf [MQ89]. HT [HSMCL87, TRC^{+}87]. HT-29 [HSMCL87, TRC^{+}87]. HT29 [DRH^{+}87]. HT29-18 [DRH^{+}87]. Human [CC85b, DDB^{+}88, DAS^{+}87, FDB85, MHF86, PP88a, SFJU86, SDDZ^{+}89, TDL^{+}85, YHS88, AJAW86, AW88, ASA^{+}87, AGPT88, APF^{+}88, AS86, BHSJ86, BMBL89, BOGK88, BLSL88, BCC^{+}87, BPB^{+}88, BNL89, BJ86, BH86b, CCC89, CSM^{+}89, CSR^{+}86, CTS^{+}87, CK86, CPHM86, CPP^{+}87, CMB^{+}86, DHK^{+}85, DWL86, DLH^{+}89, DNB88, DCL^{+}87, DCC^{+}88, DWO^{+}87, DHB86, DJM^{+}89, DBD^{+}89, DRME89, DBC^{+}87, ESM^{+}87, EB87, ESA86, EDD^{+}86, ET87b, EWHP87, FWP^{+}87, FL89, FCZ^{+}89, GDO^{+}85, GVD^{+}87, GKS87a, GM85a, GKF85, GF88, GMCH87, GBC^{+}86, HDB88, HPO88, HLB^{+}88, HDP^{+}85, HTRBP86, HCHK89, HSB^{+}85, HGMS88, HSK^{+}89b, HAK^{+}86, HB86, ICY^{+}88, ITY^{+}88, JTBA89, JNA85, JP85, KCK^{+}87, KFW^{+}89, KEG88, KOBLM88, KSJF87, KDIG89, KKML88, KG86, LSAKO86, LCL^{+}89, LH86, LS87a, LMW^{+}86, LT89, LGBE89, LJC^{+}87, LOH^{+}86, MS86a, Mac89a, MNC86, MAK^{+}86, MBM85]. human [MMM^{+}89, MER^{+}86, MNM87, MB88, MSC^{+}89, MGM87, MLE87, MSC^{+}85, MTRR89, MM89, MB85b, MLC^{+}88, NH87, NSF^{+}89, NMPG88, NKMK86, NMM^{+}86, NJAS89, NL88, OMS89, OPN^{+}85, ODD85, PSBN85, PW87, PNB89, PON^{+}88, PRB^{+}89, PBK88, PSS^{+}87, PRS^{+}88, POCM88, POD89, RJH^{+}87, RH86a, RPW^{+}89, RSG87, RHD^{+}85, RT86, Ros89, RNW89, SV87, STH^{+}89, SLB^{+}89, SLHvF85, SK85, SHGD85, SP85, SJS86, SYT^{+}89, SAB^{+}89, SB85d, SNC^{+}86, SSKK89, SAB88, SOP85, SR85c, SPT^{+}89, SGP88, SGG^{+}89, SVG^{+}86, SDZB^{+}85, SHSvdM89, TFBW85, TN89, TTBG87, TNG89, TH89c, TF86, VBB85, Wan85a, WGPW89, WC87, WBW89, WO86, WS89, WB85, YPG86, YSW^{+}85, dGRS87]. Hyaluronate [MNM87, AU89, KT85, LU87]. hyaluronic [GSTC87, RBMM88, RMM89]. hybrid [DSTJ85, RR88]. hybridization [BTR^{+}86, BLA^{+}88, HNP^{+}86, HHMH87, HKPS89, LHK^{+}89, MAE88, NWE^{+}88, PBT^{+}86, RKWS86, SV87, SLL89]. hybridoma [RR87]. hybridomas [BHK86]. hybrids [RFG^{+}85]. Hydra [HKM^{+}87]. hydrated [GG87b]. hydrogen [PBK88]. hydrolases [GF86a, GF86b, GF87, HSB^{+}85, RH86a, SMB^{+}88]. hydrolysis [ISL87, VAD^{+}85, WB86]. hydrophobic [BGBB89, PTB^{+}85]. Hydroxy [KPK86, JBG^{+}87]. hydroxyecdysone [DFF85]. hydroxylase [AETT86, NGS85, ST86]. hyperactivation [WHI86]. hyperosmotic [EMT88, ZSE85]. hyperoxaluria [DCWJ89]. hyperproducing [HGMS88]. hyperproliferation [KF89b, TF86]. hyperproliferation-associated [KF89b]. hypersensitivity [MMY88]. Hypertonic [HA89]. hypertrophic [TQN^{+}87]. hypertrophied [KKT^{+}86]. hypertrophy [BHM^{+}89b, OGYK88]. hyphae [Kro86]. hypomethylation [CMW87]. hypoosmotic [MT86]. hypothesis [Cle87].

I-band [FW85]. Ia [SRW^{+}89]. Ib [OPN^{+}85]. Ic [PON^{+}88, WCPK88]. ICAM [DS88]. ICAM-1 [DS88]. Identical [MMW85a, SDS89, SAGM88]. Identification [AK89, AB88, BS88a, Bec86, BSG89, BLR88, BSBZ88, BCD86, BL85, BWL^{+}86, BK85, CP87a, CAG89, CMW^{+}85, DHB86, DF86b, FCM85, FBB^{+}88, GFA^{+}88, GS86a, GKS87b, GKS88, GBU^{+}86, HWDC89, HBB85, HAK^{+}86, JJP^{+}89, KOF88, KFA89, KKB85, KB85, KRN^{+}86, KJY89, LW86, LT89, MGWM85, MOU88, NSM^{+}87, OBCO85, PHTK87, SS87, SFB^{+}88, SLL88, SSSA87, SSMG86, TKP^{+}88, TSM87, WC87, WGPH^{+}89, WWN87, YIJ88, BTL^{+}86, BB88a, BBGI85, BKBR88, BGH87b, BPE^{+}86, Den88, EWAP85, FDB85, GLM86, GS86b, HB86, JFP89, KTP^{+}89, KJW89, LGBE89, MP86c, MAG88, OHH^{+}89, SPF87, SH86, TT86a, TDR87a, TTKM86, WRL86, Bro85b, GY86]. identified [CBL85, EMOB85, HH86, LB89, SHSvdM89, ZRW85]. identify [BP87a, SSG87]. identity [Cha85, HNNT88]. IEF [MNC86]. if [GW88a]. Ifm [CF88a]. Ig [WDF86]. Ig-secreting [WDF86]. IgA [HCH85, SFJU86, SHP85a, SHP85b, WLJM^{+}89]. IgE [EMH^{+}87, KHB88, Mac89a, MHWB86b, PSD^{+}85]. IgE-mediated [Mac89a, PSD^{+}85]. IgE-receptor [KHB88]. IGF [HSK^{+}89b]. IgG [Hir86, JNA85]. II [HUA^{+}86, BD86b, BKR86, Bro85b, BDB^{+}87, CCL^{+}86, CHK^{+}86, CE85b, CMG86, DSA86, DJS^{+}88, DNSMA88, EH85, EHC^{+}85, EG89, FR86, GY86, Gla85b, GKF85, HKKP86, HNP^{+}86, HE86, HMP86, HRP87, HM89a, HSK^{+}89b, HS87b, JHPS86, JHP85b, KSSAQC89, LWM^{+}88, LD88, LBR86, MMR88, MHW85, MER^{+}86, MEBV^{+}89, MSLR^{+}86, MK85b, NRL89a, PN88b, RC85b, RHL^{+}85, SV87, STH^{+}89, SKK^{+}85, SDSO86, SDNWA87, SSZ88, SP89, SSP89, SCGS85, SHP85b, TL89, TRCR89b, TMS85, TM87b, Toy87b, TTKM86, VG86b, WFE89, WM88, WPW^{+}85, dVOD^{+}89]. II-arrested [MHW85]. II-related [DNSMA88]. IIa [PON^{+}88, SRW^{+}89, WCPK88]. IIb [CTS^{+}87, NKMK86, PPG85]. IIb/IIIa [CTS^{+}87]. IIb/IIIa-like [CTS^{+}87]. III [GSTS86, HAK^{+}88, IKH87, KSBB87, PPG85, SV87, STH^{+}89, TI85, Tok89]. IIIa [DBC^{+}87, NKMK86]. IIIa-like [CTS^{+}87]. ileum [SGB^{+}86, WWN87]. illuminated [KMM88]. illumination [VCM88]. image [AJRNBJ85, BSK89, DCB89, FMG89, JIFH85, KRLU87, TRCR89a, TRCR89b]. image-intensified [BSK89, JIFH85]. imaging [BBP^{+}85, BFRT87, CJO^{+}89, DBE^{+}87, PBS85, RBJ89, WAF87]. iminodipropionitrile [HBV85, PBBP85]. immature [CCP88, GJD^{+}88]. immediately [SGS87]. immobile [DNI^{+}88]. immobilization [MHWB86b]. immobilize [GBC^{+}86]. immobilized [BRS87, FLH^{+}88, MKD^{+}89, WO86]. Immortalization [AJAW86]. immortalized [BPB^{+}88, GDD88, MMK^{+}89]. immune [SGP88]. Immuno [HB86]. Immuno-identification [HB86]. immunoassay [GHL85]. Immunochemical [AE87, KKGJ89, LWSF85]. immunocolloidal [JSA^{+}89, KSKW^{+}85]. Immunocytochemical [FS86a, ITY^{+}88, KNGH^{+}85, PMC^{+}88, RLB85, TM87a, TM87b, Tok89, TB85b, BVHF89, CMTC85, HBV85, VLD^{+}85, JHP85b]. Immunocytochemistry [HBL^{+}88, CSGC88]. Immunocytolocalization [CV87a]. immunodeficiency [HDB88, PP88a]. Immunoelectron [FLM86, JJ87, MS86b, MS88a, PS87, PP89, RMM89, BCPC87, BPE^{+}86, DD88, FONW88, KL89b, KPK86, LAA85, MYH^{+}85, OH89, STE88, SDMF85, SH87b, TB85a, TCP^{+}85, WW88, TMS85, VG86a]. Immunofluorescence [AJRNBJ85, MGW86, JSA^{+}89]. Immunofluorescent [RP86, MBSMD86, TM87a]. immunoglobulin [BHK86, BHM89a, DNM86, GCR^{+}89, HBKK87, MHWB86a, MBH87, Ran88, SHC87]. immunoglobulins [WLJM^{+}89]. Immunogold [LWM^{+}88, MP86c, CV87a, FYRK88, IMP^{+}87, NNJ^{+}87, PAGPLM86]. immunogold-silver [CV87a]. immunogold-surface [IMP^{+}87]. Immunohistochemical [SL85a, dBBB^{+}86]. Immunohistologic [SRB88].

Immunoidentification [SGS^{+}89]. immunolabeling [PW86]. immunoliposomes [CH85]. Immunolocalization [BOGK88, GKBG87, LBR86, OKLB86, WESP85, BLA^{+}88, JSCLP89, KVB^{+}86, MABJ^{+}86, WFF86, YJL^{+}89]. Immunological [MAC89b, CCA^{+}88, GL85, PBB86, SGBG87, SWW85, WSRT89]. immunologically [CCLC89, HMHWS86, MSQ^{+}89, NKMK86, WRL86, WGA85]. immunoperoxidase [PBT^{+}86, JHP85a]. immunoreactive [BL85]. immunoreactivities [AIJ^{+}89]. Impaired [CSD89, CRR87b]. impermeant [SSS88]. implantation [NT86]. implanting [SHBV89]. implicate [BBB89b]. implicated [DWT^{+}89, FW85]. implications [BPC89, CIK^{+}89, HERO^{+}87, JPM85, KC85, NH89, RLB^{+}89, SWC^{+}87]. Import [Cla87, PSR^{+}88, SL87b, SHK^{+}87, AG89, DS87, EVG^{+}86, HFF^{+}85, HKF^{+}87, HJV^{+}89, KBPW89, MWS85, MPB88, NAS^{+}87, PMV^{+}89, SHN87, VS88a, VS88b]. importance [HAK^{+}88, PSBN85, TKK^{+}87]. important [KTF87b, SCK^{+}89]. imported [DVVT88, RGW^{+}89]. impurities [PMJ87]. IMR [LWMM^{+}85]. IMR-90 [LWMM^{+}85]. inactivated [JWH85]. inactivation [SDSO86]. inactive [BIES89, MFB^{+}89]. INCENP [CHE87]. includes [HFK^{+}89]. including [HKM^{+}87]. incorporated [DH89a, HVBB88, KBB^{+}88]. incorporates [WSRR89]. Incorporation [BHS88, OH89, WE87, AT86a, BGF88, DI87, JGAR89, MMC88, MVK^{+}88, VJF^{+}88]. Increase [OE85, SVG^{+}86, GK85, KWW85, KSG85, LMWG85, MSW^{+}86]. Increased [LSDP86, WBDR88, DAJJ^{+}88, DSGS87, HKPS89, MKK89, MBKO86, MVB85, PRS^{+}88, SR86a]. increases [AETT86, AYG85, CPM86, HC87, KM87, MDM88, PBK88, RG86b, VKB89]. increasing [BGN^{+}88, CMM^{+}85, VES85]. incubation [GLK^{+}89, MTA86]. independent [BNB^{+}88, BA87, HGMS88, Kam88, MDM88, PON^{+}88, VSSGRB87, Eva89, PJW88]. independently [BF88, CM86b, ECZ86, GDvdM87, HZS88]. indicate [MTS89, SS88a]. indicates [BPJBF85]. indicator [SWZ85]. indirect [CF88a]. Individual [WGA85, BBP^{+}85, KTF87b, WOP^{+}88, WIS89]. individuality [TWB88]. induce [DDB^{+}88, SFS^{+}87, SSK89, VMT^{+}88, WBDR88]. induced [ASZ^{+}88, And86, BBBM89, BGM^{+}87, BGL^{+}88, Bre89, BM88, BSG85, CMB^{+}87, CC85a, CHMB86, CGA89, CCBKJ89, DLG^{+}86a, DFSK85, ETA85, FSM88, FA85, FC89b, GBH87, GCGR85, HTTFC85, Has88, HP85a, HO85, IMP^{+}87, KWW85, KC88, KH85b, LLRR87, LEFS^{+}87, LP85b, MB85a, MP86a, MWSF87, MW86, MTRR89, Miy88, MKD^{+}89, NSF^{+}89, OBH87, OGYK88, PVMO87, PSC^{+}89, RB85, SME^{+}89, SIYT86, Sow86, SFD^{+}86b, SKC85a, TFBW85, TRL87, TN89, TS86, TTM88, VJT89, VHO^{+}89, VAD^{+}85, VB88b, WD86b, Wal88, YMD87, ZBSS^{+}87, BB88b, HB86, Sal86, Van87]. inducer [BGT^{+}89]. inducer-specific [BGT^{+}89]. inducers [BGT^{+}89]. induces [BGN^{+}88, DCL^{+}87, ES89, GKS87a, Gla85b, HPO88, HTRBP86, KLHS85, KdWB^{+}89, MP89, MPBZBE89, MS88b, SSR89, SNC^{+}86, SHL^{+}89, SDPZ89, SW88, TES^{+}87, Wal86, WTB^{+}89, ZMH89]. inducible [ABME88, CVWZ89, DFSK87, FBP89, GZO88]. inducing [DN88, HG89, PMR89, WHMA86]. Induction [BHM^{+}89b, KWS^{+}89, LM85a, MHL88, MMWM^{+}89, MLE87, RRP88, TRT^{+}88, CTR86, CKR85a, CE85a, CE85b, DFSK85, GMCH87, HZS88, MLK^{+}87, ST86, SB85a, SPO85, UW86, WG87b, WB85]. inextensible [WW88]. infected [CPG^{+}89, HdAC89, HUA^{+}86, RIPS85, TB85b, WS89]. infection [BDS85, MBLS87, PP88a]. infiltration [PCS88]. inflammatory [SC88]. inflow [DCQB88]. Influence [EMK85, NJAS89, JS86d, MBC86]. influences [JGAR89, PCK^{+}87, SLB87]. influenza [BDWH88, CDB^{+}86, DRSG85, DSG86, FRG^{+}87, GHW^{+}88, GDYW86, HGS89, HGL88, HBHL^{+}89, RIS87, RDSG86, RGPRB87, SME^{+}89, SRM89, WW87b]. influx [PW85, TI85]. information [DVHC89]. infusion [OS88b]. inheritance [WBW87]. inhibit [CR88b, CCL^{+}86, HA89, JBdlTB87, JB86, SPS89a, SWTB88, WCPK88]. inhibitable [MMT85, SNSS87]. inhibited [HW85a, MKH88, RG88]. Inhibition [BRR^{+}87, CWR86, CTSF86, DZ85, FNPF87, FFH88, FKM87, GAPR88, GLH85, HBC89, ICS88, MDVG^{+}85, OD87, RWE^{+}88, SR89, WMUP^{+}85, BKN^{+}85, BA87, CRT87, ETT88, FPS^{+}85, FRH85, GZO88, GDC^{+}86, Hat87, HGMS88, KWPH89, LCR^{+}87, MNM87, NR88, OGYK88, SAA89, VRLG89, WSV^{+}89, WO86, ZSSR86]. inhibitor [GKS87a, HWDC89, Has88, KWS^{+}89, LSAKO86, LS87a, MVS88, MEVM89, PSS^{+}87, RRB86]. Inhibitors [MSD^{+}88, BVD87, CC88, HP86, RBKK85, dLG86b]. inhibitory [CS88b, HW86b]. inhibits [BSG85, CSH89, CPC^{+}89, CdCP^{+}89, FDG88, GG87b, GLW^{+}89, MF87b, PJ87, RT86, SVF^{+}89, SOPvD87]. Initial [SME^{+}89, TDGE85, WMM86, CMHC85, CTS^{+}87, HFB^{+}89, LBEM89, PJW88, VB85a]. initiates [CMNB88]. initiation [AF89, JG85, KWPH89, NR89a, STS^{+}89, WRR87]. injected [HTSH85, HW85a]. injection [WFW87]. injured [GABT86, Gor85]. injury [DLG^{+}86a]. inner [ASF89, BBB89b, CHE87, FL88, FS87, GH85b, HG88, HJV^{+}89, LBFS88, OK86, SCOE^{+}86, SG88a, SALN88, TLP^{+}87]. innervated [CCS87, MAK^{+}86]. innervation [SR86a]. inorganic [BKC^{+}86]. Inositol [Miy88, PF87, BFJ^{+}85, MSE89, SW86, TJF86]. insect [SFD^{+}86b, TRCR89a, TRCR89b]. inserted [SS89c]. Insertion [BBMW85, DLFE89, FRG^{+}87, HFGD89, HOPM^{+}87, LD86, PSR^{+}88, RDS^{+}89, WE86b, vAAA86]. inside-out [LHFB88]. insights [CKF89, Gor89]. insoluble [LWR85, WRR87]. instability [CPS88, JPM85, MRO^{+}86, WOP^{+}88]. Insulin [BGL^{+}88, BRD^{+}88, CGV^{+}86, GST^{+}87, GGR89, GSDR89, ICP^{+}86, JB86, MER^{+}86, NHKE88, ORA^{+}86, RH87, SJ87, TL89]. insulin-degrading [GSDR89]. Insulin-induced [BGL^{+}88]. insulin-like [BRD^{+}88, EG89, MER^{+}86, NHKE88, TL89]. insulin-related [GST^{+}87]. intact [BMBL89, CCG89, GPJ88, HMTF88, ISM85, KBPW89, KB86, LMW^{+}86, MCA^{+}89, NADS86, RRB86, SM89a, SJS86, SD88, VT89].

Integral [SG88a, BPKS86, DRRB88, HHR^{+}89, HGL88, KJW89, LMK88, NJD^{+}86, PMR86, RPRS89, RS85, Rog87a, RSL^{+}89, SJOB89, WMBR89, WBB89, WL87, YBBS87]. Integration [HGL88, WCMG88]. integrative [MGWM85]. Integrin [BSDH86, DG89, HMS^{+}89, DJ89, GGA^{+}89, LGW^{+}89, LCBS89, LWH88, MH88, MKD^{+}89, SGM^{+}89, TDR87a, KFW^{+}89]. integrin-mediated [DJ89]. integrin-related [TDR87a]. integrins [NSF^{+}89, NTLR88, TH89a, TDR88]. integrity [BSDH86, LFC^{+}88]. intense [VJF^{+}88]. intensified [BSK89, JIFH85]. intensity [ADS88]. interacting [MFA89]. Interaction [BD86a, BD86b, GVD^{+}87, Gla85a, Gla85b, MLM85, SRH^{+}87, SKC85a, SKC85b, TRM88, BB87a, BFG86, BICFA88, CTSF86, CM85, DS88, GCT86, GSTC87, HD88b, HMP85, HFB87, HAM86, JTBA89, KSJV88, MHK^{+}85, MBvS^{+}89, NRTE89, NMM^{+}86, PSM89, PSK^{+}87, QS87, RLP^{+}87, TSBW86, TSS89, VB85b, VGKF86, WKBR87, WL86]. Interactions [Ake89, BBJG89, HBHL^{+}89, MUM87, RNW89, TDR87a, ACEE87, AE88, BBV^{+}89, BS87, BGBB89, CSG85, CHM^{+}87, ELKH88, FBCQPR87, GM85a, GWM85, HS85b, HFB87, KJF^{+}88, LS85, MS88b, MMRR86, NHOY85, OS88a, PFRS89, REB^{+}86, SS88a, SPS89b, SQM89, TRB86, TSOC89, UF89, VJT89, VSSGRB87, VSSRB87, WGC85]. interactive [HCE88]. interacts [GE88, LE86, LT89, MMM^{+}89]. Intercellular [PRB^{+}89, DS88, GKBG87, JLH^{+}86, PJW88, VGKF86, VG86a, VG86b]. interchangeable [May89]. interchromatinic [BSY86]. intercompartmental [SEGS85]. interconnections [SP86, YF87]. Interconversion [SF89, SGBF87]. interdependent [TCF86]. interference [GB86a]. interferes [EKH89, MGSG88]. interferon [FKM87, GBB89, HMP85, SGP88, TBK86, YPG86]. interferon-gamma [GBB89]. interferon-like [YPG86]. interleukin [DWL86, LSDP86, PRS^{+}88, SAB88, YHS88]. interleukin-1 [DWL86, SAB88]. Intermediate [BOW85, BOW86, JG85, RFV85, AF89, AST89, BVK85, CP87b, CRLW86, DHL85, ESA86, FLH^{+}88, GM85a, GWM85, GB87a, GB87b, GGJ^{+}87, HF85a, HS89, Ip88, JP85, JGYG85, KSJF87, Kol86, LFFW89, LGC^{+}88, LYG85, LKHR87, NPCL85, PL85a, PSK^{+}89, RR86, SBS87a, SQM89, Wan85b, WM88, YLGG85, dSdSH^{+}88]. intermediate-sized [HF85a]. intermediates [VP85]. intermembrane [KBPW89]. Intermolecular [PFRS89, MSA88]. internal [AC86b, GTB85, HGL88, KWW85, LWT85, PL85b, VS88a]. internalization [BSK89, BBMW85, CGV^{+}86, CWOK87, CNHF85, GF86b, LCD^{+}89, LCW^{+}89, NRL89a, SGS87]. internalize [GRCG86]. internalized [vDTP^{+}86, vDSP^{+}88]. interneuronal [Ran88]. Internexin [NPCL85, PL85a]. interorganellar [LSF86]. interphase [BRP88, DN88, FA85, GGN^{+}86, HS87a, HS87b, HHSF85, LWT85, SBS88, SL87a, SK86, VB88b]. interphotoreceptor [HVR^{+}85]. intersections [MGHG^{+}86]. interstitial [HKH^{+}88, HVR^{+}85, JPU89, SB88]. intervals [PP88b]. Interzone [SM87]. Intestinal [BJMK89, Gor89, AFMD89, AMWN89, BVK85, BNF^{+}87, CCA^{+}88, CM87c, DD88, DRH^{+}87, EKH89, GCC^{+}89, GG85, GSF^{+}88, HSB^{+}85, HSMCL87, KCCM85, MBC86, MSQ^{+}89, MCC^{+}89, MC89b, OGJ^{+}89, RT88, RLB85, SCM87, SMB^{+}88, TNG89, WLJM^{+}89]. intestine [AMWN89, GSF^{+}88, MSQ^{+}89, Qua85a, Qua85b, SABVK89]. intoxicated [HBV85]. Intracellular [ALL89, CGV^{+}86, DK88, DIF^{+}88, HP85b, JT85, KS85a, LP85a, NRL89a, PFS85, RH89, SM88, SP88, SB85c, SR85c, TB89b, VB88b, WG88a, WSH85, YMM86, AO88, BMR85, BAHE86, Bas87, BB85, BCMK87, BFJ^{+}85, CS88a, CR87, CR86, CDB^{+}86, DF86a, DH87, DMD^{+}86, DRM^{+}88, DRSG85, FWP^{+}87, FKC87, FCL86, GB85a, GKR87, GWCW85, GZGM87, GMF85, HSWH86, HRBR89, HSB^{+}85, HLL89, ICP^{+}86, JLB86, KP89, Kro86, LLSRB89, LRF89, MKK89, MP86c, MSLR^{+}86, PBS85, RFG^{+}85, Rob87, RL89, SNCY87, SSY^{+}86, SB85b, SRM89, SOP85, SS88c, SS89d, SGG^{+}89, SKU^{+}85, SHP85a, SHP85b, TL89, TP89, TSM87, VHO^{+}89, VSSRB87, WM88, WSV^{+}89, WF86]. intraerythrocytic [HLU^{+}87]. intramembrane [SDSO86]. intramolecular [PFRS89]. Intranuclear [NHH^{+}86, FS88a, WS85]. intrinsic [BSLS^{+}89, DMRB87, YKCR88]. introducing [DSHW85]. introduction [GLC88]. invaginations [Eck87]. invariant [LD86]. invasion [BMVB89, CSM^{+}89, GAPR88, Oss88, PW85, RR87, RdRBR89, SHBV89]. invasive [BMVB89, SIYT86]. Invertase [SEGS85, HW88]. invertebrates [BOW85, MH88]. investigated [GHL^{+}89]. Investigation [TB85a, ASF89, BVHF89, WFW87]. involved [ALWL89, APS87, BSG89, BWL^{+}86, CPHM86, CG86a, DNSMA88, DMH^{+}85, FCCC^{+}89, GH88, GS86b, HK88a, KST^{+}89, KJF^{+}88, MCD85, NM89, PSBN85, PM87a, PSK^{+}87, RWF^{+}87, RSL^{+}89, SGRM85, SMVS^{+}86, SHN87, SRB88, TDR87a, TLPTV85, WM85, Wan85b, WJSB88, YSR87, YBSR87]. Involvement [BK86, CDL^{+}85, KSM89, PCBD89, RR87, RdRBR89, WMGM85, BS88b, ES85, FOW89, HNH^{+}87, MLK^{+}87, PCS88, RGS88, RFSL88, SE85]. involves [BB87c, DFSK85, ZP85]. involving [CKR^{+}85b]. inward [AC86b]. inward-directed [AC86b]. iodination [MG86b]. ion [BSG85, CR86, EZBW88, Owe85, Sch85]. Ionic [RT87, ICY^{+}88, Kro86, YB87]. ionophore [DMW88, NT85, RTK^{+}86]. ionophore-treated [NT85]. ionophores [OCP86]. ions [GD89, VGKF86]. iris [SR86a]. irises [SR86a]. irradiation [WIS89]. islet [MLR^{+}86, NADS86]. isoactin [OLGT86, WR88]. isoenzyme [DFSK87]. isoform [CBS^{+}89, CCK^{+}88, EOM86, GS89, HWSK89, OKLB86, SL89]. isoforms [AE87, BCC^{+}87, CMM^{+}85, DHH89, DBD^{+}89, EPJBBW86, HLL89, HERO^{+}87, HM89c, KSEP86, KTO^{+}89, KKT^{+}86, LHHC85, LL86, LHL88, SSS^{+}87a, SGBF87, SF89, TB89a]. Isolated [AC86b, BKC^{+}86, SF88a, ACPFB^{+}85, Blo86, BF87, CBCV85, CJ85, EWAP85, EHM86, EWHP87, GS86a, GTP87, GMLR85, HTH85, HSMCL87, KCCM85, MMBA89, MM87, NFW89, NFN87, OHK86, SGH85, SJ87, SL85d, TT85, TIT89, VB85b, WD86b, WE89, WS87]. Isolation [BB88a, CCC89, GF88, GKZ89, HFHR88, HOPM^{+}87, HHR^{+}89, HHS^{+}85, JVG^{+}85, JS88b, JP85, KBKC86, KR86, MGK89, MAG88, PPG85, PH89, Pip88, RPCB88, RHD^{+}85, SBD^{+}89, SL85c, SL85b, SB85f, TMP87, TT89a, UK86, AHC89, AMWN89, BOW85, GLM86, HMHWS86, KTP^{+}89, LFR85, LC85, MSK^{+}87, WKS85]. isomaltase [TRC^{+}87]. isomorphy [GC87]. isomyosins [JHP85b]. isoproteins [SP88]. isothiocyanate [PMJ87]. isothiocyanate-conjugated [PMJ87]. isotubulins [DFdN^{+}89]. isotype [SGBG87, WVLC86]. Isotypes [MND^{+}87, GLC88, JC89, LLC85, LC88b, LC87]. isozyme [WTBBW85]. isozymes [GVGL85, MAK^{+}86]. Iterative [DMM89]. itinerant [DRY89]. IV [AT86b, BFKF86, CTYF85, CTSF86, CL85b, EWAP85, GLK^{+}89, HMTF88, LHK^{+}89, PW85, TDR87a, TDR88, TC86, YR87]. IX [ACPFB^{+}85, ISM85, MEBV^{+}89, MGHG^{+}86, VMH^{+}88].

J1 [GSS89, PSS89, SSC86]. J1-160 [PSS89]. J1-180 [PSS89]. J774 [SSS88]. jackbean [BMP^{+}86]. japonicum [HMHWS86]. jelly [WBGV85]. jun [SHL^{+}89]. junction [ASJ^{+}88, AVP^{+}89, BPG87, BICFA88, CG86b, DYT^{+}87, FMJS89, GKG88, GAH86, JLH^{+}86, KG86, MD85, MG85, Pau86, RGE85, RSP89a, SCLG86, SG88b, SFS^{+}87, SSMG86, SAGM88, TLD^{+}89, THT89, VJF^{+}88, VG86b, ZHE85]. junction-specific [FMJS89]. Junctional [PSC^{+}89, AR85, BN87, CM87b, CCS^{+}86, GM86, GSG88, KL89a, KMR^{+}88, KKB85, MBC86, MBL89, MG85]. Junctions [SSJ^{+}85, BKPG89, BVK85, FMJS89, GVV85, GPJ88, GGJ^{+}87, GKBG87, MBC86, MG86a, OBH87, PHRM87, PBH^{+}87, SS89c, SGM^{+}86, TOB86, TT89a, TIT89, VSSGRB87, VGKF86, VG86a, VG86b, WG87b, YJL^{+}89, ZKR^{+}88, ZHES89, ZHE85, ZN89]. juvenile [BM88].

K-252a [Has88]. K-fgf [MQ89]. K14 [CKF89]. K4M [BTR^{+}86]. K4M-embedded [BTR^{+}86]. K562 [WKRH86]. K7 [GF88]. karmellae [WBDR88]. karyoskeletal [MWSF87]. KB [TT88b]. kD [GCC^{+}89, AY89, AYY86, BBBM89, BNLC85, BKPG89, BCD86, CMB^{+}87, CCA^{+}88, CB88, CM87c, DYT^{+}87, DLC88, FS88a, HPW89, HS89, HH87, JBdlTB87, KTF87a, KCN^{+}89, KJW89, KRN^{+}86, MS86a, MBKR89, MMS89, MC89b, MPB88, MKA^{+}86, MFM89, PL85a, POW^{+}87, PSGS85, Ran88, Rob87, Rob89, SKE86, SPF87, SFB^{+}88, SSSA87, TLD^{+}89, THT89, TIT89, UK86, UPL^{+}87, VG86a, VG86b, WTB87, WL87, YMM86, YJL^{+}89, YBWLM88, ZHE85, ZP85]. keeps [VSKSP86]. keratan [ZRW85]. keratin [AF87, AF89, CKF89, ESA86, FH86, GKF85, KFH^{+}86, KTF87b, RKWS86, SGS86, SAA89, SKDF88]. keratinization [BPB^{+}88, KTF87b]. keratinocyte [BPB^{+}88, KFH^{+}86]. keratinocytes [AW88, DDB^{+}88, GGJ^{+}87, HLB^{+}88, MS86a, OBH87, TTBG87]. keratins [DHK^{+}85, FJR87, KF89b, LOH^{+}86, MWSD86, TF86]. kidney [AIJ^{+}89, ASJ^{+}88, ACEE87, BAD^{+}89, BBGI85, BPP^{+}89, BRD^{+}88, BKK87, BHM89a, CW88, FGF88, FS86b, GDRB89, GBM89, HSK^{+}89b, HM89c, JY86, LHK^{+}89, LLSRB89, MFKVK87, MCA^{+}89, MBH87, MTA86, NV86, NV87, NH89, NMPG88, NTNT85, OS88a, PPB^{+}89, PN88a, PN88b, PBH^{+}87, PFS85, RJH^{+}87, RIPS85, RIS87, RE86, RSF86, RG86a, RGPRB87, SGRM85, SG88b, SAGM88, TB85b, UPL^{+}87, VSSGRB87, VSSRB87, VSZ^{+}87, YLGG85, ZMH89, vMSvRS87]. kidney-derived [NMPG88]. kidneys [Abr85, AP86]. killing [WFL87]. kilodalton [YMM88]. kinase [ACN^{+}89, BC89, BSB^{+}86, BJMK89, BA87, CPC^{+}89, CS88d, DAJJ^{+}88, DNSMA88, DF87, DMW88, EMZ85, Eva89, FBF^{+}89, GZ89, GCBH86, GDS87, Has88, HL88, ICP^{+}86, ITY^{+}88, JLK89, KC88, KSKW^{+}85, LFC^{+}88, LFFW89, LWR85, LPR^{+}89, LWM^{+}88, LBP^{+}86, MAK^{+}86, MBSMD86, NR88, NHKE88, PH89, PFD^{+}89, PMR89, RRB86, SS85, Sch86b, SDSO86, SDNWA87, SHL^{+}89, THM86, VHO^{+}89, VRLG89, WWW^{+}85, WG86, WRR87, ZSSR86]. kinases [CWH86b]. kinesin [Hol89a, ICS88, NWM88, PWS^{+}89, SSRS88, VH88]. kinesin-driven [ICS88, VH88]. Kinetic [BC85, vGRB86, SHP85a]. Kinetics [CS88a, GZGM87, PN88a, PN88b, YM87b, CE89, HO85, PN89, ZSE85]. Kinetochore [RBJ89, BB87a, GHL^{+}89, GSB87, GB89a, HK88b, KFA89, MK85a, MK85b, Nic89, SRW88, VB85a, WSRR89]. kininogen [GLW^{+}89]. Kupffer [ACPFB^{+}85].

L [AJAW86, BS85, CSM^{+}89, GF86a, MC89a]. L-cells [GF86a]. L1 [AR85, BGL^{+}88, MS86b, MS88a, PS87, SSC86, SS88a, SPS89a, TL89]. L2 [KJF^{+}88, PSK^{+}87]. L2/HNK [KJF^{+}88, PSK^{+}87]. L2/HNK-1 [KJF^{+}88, PSK^{+}87]. L3 [KFS87]. L5 [HKPS89, SBH^{+}88]. L6 [Wri85]. label [KdS86, SH87b, SLL89, TB85a]. labeled [AT86a, HTSH85, MMW85a, MJW86, OH89, PSBN85, SMPS86, SSJ^{+}85, WFW87, BD86a, BD86b]. labeling [Abr85, GS85, JSA^{+}89]. lack [GHL^{+}89]. lacking [DAJJ^{+}88, GPSC86, ZH86]. lacks [MR85b]. lacunae [BNLC85]. laevis [AKVF88, BYB^{+}88, BC89, BRHR^{+}86, BFJ^{+}85, BN85b, FLBP^{+}86, GZHM87, KSH87, LC88a, LCE87, MWSD86, NCB88, NB88b, SP85, SPL87]. lamblia [PAHC89]. lamellae [CM88b]. lamellipodia [ES89]. Lamin [GB87a, GMB89, BK86, CCLC89, GLD^{+}88, KWH89, LSEN87, LL89b, SGBF87, SF89, SALN88]. lamina [And86, AM85, EBBW87, MOT^{+}85, NPRR86, SFLM86, SG88a]. laminae [RNPR^{+}88]. Laminin [AETT86, ES89, Abr85, AIJ^{+}89, BFL88, CTYF85, CTSF86, CSK^{+}88, CCM86, CMNB88, CHH^{+}88, ETT88, EDD^{+}86, FBF^{+}89, FCDF87, GDvdM87, GRvdM89, GLK^{+}89, HFHR88, HBF^{+}88, HMTF88, HDG^{+}85, KDBF86, KKML88, LFR85, LGW^{+}89, LHK^{+}89, LE86, MS88b, MG89, MEVM89, RVS88, SSC86, STE88, SSKK89, TDR87a, TDR88, YBWLM88]. laminin-binding [HFHR88, YBWLM88]. laminin-heparan [BFL88]. laminin-like [CMNB88]. Lamins [LLRR87, BK86, KWH89]. Lampbrush [Hut87, RG87]. lapse [MASA89]. Large [LBP^{+}86, MMM^{+}85, RLP85, APF^{+}88, DRD^{+}88, GSF^{+}88, HMP86, KR86, KBO^{+}88, MSQ^{+}89, RGPRB87, SHR86, TYP^{+}87]. Large-scale [MMM^{+}85]. larger [DRD^{+}88, MHWB86b]. largest [HKTY89]. larvae [NCB88]. larval [DLC88, WG85]. larval/pupal [DLC88]. Laser [TWB88]. Laser-transected [TWB88]. lasting [RMS88]. late [CM88a, PPB^{+}89, RWE^{+}88, SM87, WSRR89]. latency [GTG85]. latent [DN88, LKOM88, SR89]. Lateral [BHB^{+}87, CMK87, HRR89, JBTA88, SHH85, WHL^{+}86, AELE88, BLSK88, DCQB88, DF86b, DNI^{+}88, FMK^{+}86, GSS87, IHJ88, JY86, MVSW86, NYRK87, SHP^{+}87, SGP88, SBC^{+}87, WE86a, YR87]. laternal [LBP^{+}86]. latipes [Sch85]. latrotoxin [SDDZ^{+}89]. lattice [Heu89b, LDP87, MSR^{+}86, STGK87]. lattices [LDA86]. layer [TRCR89a, TRCR89b, WBGV85]. layers [CCE87, LGF86]. lead [ACEE87, RIS87]. leader [HKF^{+}87, SHK^{+}87]. leading [DWFT88, Wan85d, dVOD^{+}89]. leads [MHWB86b, RWE^{+}88]. leaflets [HG88]. leaky [FS86b]. learned [Cle87]. least [DS88, MPB88, Pum89, PGH^{+}88]. Lec4A [CSFS89]. lectin [BHRC85, BRHR^{+}86, FNPF87, GR89, SKC85a]. lectin-induced [SKC85a]. lectin-like [GR89]. lectins [FSM88, HMHWS86, MDW86, SKC85b]. leech [CMNB88]. legumes [BWL^{+}86]. Lemna [CHK^{+}86]. length [CM85, FJMW87, GW88a, HW88, HP87b, KHB88, NG85, TT88a]. lengths [TTC88]. lens [ABM86, CSH89, DMRB87, GKBG87, JLH^{+}86, KKB85, LG89, MG86a, SSJ^{+}85, WLW^{+}88, YKCR88, ZHES89, dJHMB86]. LEP100 [FTLSS88]. less [ZP85]. leukemia [BPKS86, EMH^{+}87, LYE87, MHWB86a, MF87b, OSP^{+}85, PSD^{+}85]. leukemic [ITY^{+}88, QCKR85]. leukocyte [AGPT88, CGA89, HCHK89, LCBS89, RR87, SDPZ89, TLZ88]. leukocytes [CWZ86, DZ85, SSKK89]. level [BTR^{+}86, BP87b, CWZ89, FWF88, HSS89, KGB88, LSC86, PC85, SR86a, VN88, WDR89]. levels [AETT86, ALCG88, BJMK89, DMGS88, DKKK88, FKC87, KCK^{+}87, LMWG85, OMM87, RB85]. LFA [DS88, RR87, RdRBR89]. LFA-1 [DS88, RR87, RdRBR89]. LFA-1-deficient [RdRBR89]. Ligand [CC85a, LCD^{+}89, LCW^{+}89, ZBSS^{+}87, BHM89a, CR88a, DWO^{+}87, HD88b, IMP^{+}87, JGAR89, OSD85, Pfe87, PSSP88, SLB^{+}89, SFJU86, TSBW86]. Ligand- [ZBSS^{+}87]. Ligand-induced [CC85a]. Ligand-mediated [LCD^{+}89, LCW^{+}89]. ligand-receptor [HD88b, SLB^{+}89]. ligands [CBIO85, DMM89, EG89]. ligase [CA87, SWW85, WW87a]. Light [KMM88, RBMM88, SSS85, TTM88, ADS88, BB87c, BA87, CHK^{+}86, DSA86, FWF88, GDS87, HM89a, KL89b, KHC^{+}86, KGHK^{+}87, KRN^{+}86, LFC^{+}88, LA87, LGBE89, LM85b, LU88, PSM89, PWS^{+}89, RB85, SDSO86, SDMF85, SOP85, SCGS85, TOOK^{+}85, TGNG87, TMP87, WOP^{+}88, WAF87, ZSE85, ZG87]. light-harvesting [CHK^{+}86, DSA86, HM89a, KHC^{+}86, LA87, LM85b, LU88, SDSO86, SCGS85]. Light-induced [TTM88, RB85]. Light-regulated [KMM88]. light-scattering [ZSE85]. like [ATNH86, ABM86, BTL^{+}86, BF86, BRD^{+}88, BWSB86, BSK86, CCA^{+}88, CR88b, CTR86, CTS^{+}87, CMNB88, CS88c, EG89, FHS^{+}86, GR89, GS86a, GLK^{+}89, GS86b, HTRBP86, HBB^{+}89, HHY88, KKML88, MSM88, MCB85, MER^{+}86, NV89, NR89b, NHKE88, PSV87, PWS^{+}89, RMM87, RYR^{+}89, SR89, SCRA87, SHS^{+}88, SA85, SWTB88, TL89, YPG86, YMM88, ZR87]. limb [CHMB86, ETA85, ET87a, KKG86, VBWC88]. limbal [SGS86]. Limited [HT86, GF86b]. Limulus [LCK88, PF87, SKL85]. line [BPB^{+}88, BCK85, CM88a, CCMB87, DRME89, FGA86, FONW88, GK85, GDD88, HKH^{+}88, HSMCL87, Kub87, LMSM89, LOG85, MMK^{+}89, MK85c, PBK88, SK85, SPO85, TDR87a, TDR88, UPL^{+}87, WW88, SCBR85]. lineage [MMK^{+}89, MASA89]. lineage-restricted [MMK^{+}89]. lineages [MS86c]. linear [CF88b]. lines [AF87, BJ86, CCC89, DNB88, FJY86, HHR^{+}89, MLR^{+}86, MCS85a, MYH^{+}85, MJW86, MMK^{+}89, MTK^{+}87, MB85b, ODD85, RG86a, SF87, WRL86]. link [CJH85, FCM85, GSTC87, MYH^{+}85, RMM89]. Linkage [HB88a, CHH^{+}85, SMC89]. linked [BSNG85, BSK86, FCCC^{+}89, GB85a, HFG87, HDL^{+}86, KKSK86, MB88, MS88c, REB^{+}86, SDS89, VKP89, ZG87, DBK87, HSS^{+}87, KKSK86, YBBS87]. linking [AHC89, BB87a, BS88b, BHB^{+}87, CM85, Heu86, KHB88, KdWB^{+}89, KJW89, MHWB86b, MSA88, PE85, WKBR87]. links [AC86a, DCB89, MCA^{+}89]. linoleate [PSBN85]. linoleate-labeled [PSBN85]. Lipid [CKR^{+}85b, KP89, EB87, FMK^{+}86, GMF85, HDP^{+}85, KKSK86, RC85b, SME^{+}89, SB87, Ste85, WHI86, WSM86, RC85a]. lipid-binding [EB87]. lipid-linked [KKSK86]. lipid-protein [Ste85]. lipids [DWS88, KBB^{+}88, PSM89, SGP88]. lipopolysaccharide [DRME89, RH86a]. lipopolysaccharide-stimulated [DRME89]. lipoprotein [BMR85, DTW^{+}85, DCB^{+}88, KKSK86, KBK86, LRF89, PSBN85, PMC^{+}88, TFBW85]. lipoprotein-derived [LRF89]. lipoproteins [PBMB85]. Liposome [RRB86]. liposomes [KBB^{+}88, LG88, SCGS85]. Listeria [TP89]. Lithium [BSG85, Bro87]. lithium-sensitive [Bro87]. lived [Sow86]. Liver [TKK86, AK89, ACPFB^{+}85, ASJ^{+}88, AB88, BMR85, BVHF89, BPG87, DFSK87, GDRB89, GLH85, HBL^{+}88, IGSGD87, KHKR86, KPK86, KG86, LBO^{+}85, MYT87, MGK89, MCG88b, NTNT85, PJFB89, Pau86, PV85, PBT^{+}86, SJ87, SCOE^{+}86, SFS^{+}87, TSS89, TLD^{+}89, TT89a, VP85, WH85, YF87, ZHE85]. liver-specific [IGSGD87]. living [CPS88, CBS^{+}87, DBE^{+}87, DSHW85, DK86, GSB88, HH86, JMW88, LFC^{+}88, LFFW89, LPLT85, MMW85b, MMW85a, MJW86, MWK89, MW86, MASA89, MSS87b, PP88b, ROCK86, SMPS86, UP86, Wan85d, Wan87, WFW87]. lobulations [SNC^{+}86]. local [MP86a]. Localization [BSSMF88, CLF^{+}88, DJBD85, EH85, FYRK88, FC89a, FH86, HM89a, HHSF85, KSS86, KSKW^{+}85, LAA85, LG89, MSV^{+}88, MBK^{+}86, NWM88, NMM^{+}86, PM87a, RSF^{+}85, SV87, SSV^{+}85, SG88b, SFD86a, SBC^{+}87, TL89, AK89, AIJ^{+}89, AE87, AJRNBJ85, BB87b, BNLC85, BS88a, BTR^{+}86, BD86a, BBB88, BL85, CRF^{+}88, CHD^{+}86, CMP86, CA87, CLC^{+}86, CMTC85, EDD^{+}86, FCDF87, FTC^{+}89, FLH^{+}88, FW85, GWG85, GBBS89, GWI^{+}87, HAB^{+}88, HNP^{+}86, HZS88, HBB^{+}89, HTT89, HBV85, Hol89a, HM89c, JNR^{+}85, JSC85, JJ87, JHP85a, KFW^{+}89, KBKC86, KPK86, KSC^{+}89, LW86, LPR^{+}89, LT89, LWM^{+}88, LGBE89, LHL88, LC87, MS86b, MS88a, MM89, MDW86, NHH^{+}86, OOCT89, PFD^{+}89, PMC^{+}88, PAHC89, PAGPLM86, PS87, PBS85, PHN^{+}88, RFH85, RKWS86, RBMM88, RMM89, RPRS89, RP86, SHBV89, STE88, SL85a, SDS87, SB85b]. localization [SSS^{+}87b, SSO89, TASK86, TIT89, VHO^{+}89, VLD^{+}85, VG86a, WWW^{+}85, WAG^{+}89, WRR87, WSJ85, YMM86, YMM88, vdPdPR86, SHP85b]. localizations [PSS^{+}87]. localize [SKDF88]. localized [AY89, Bec86, BSY86, BKPG89, BVK85, MGHG^{+}86, THT89]. localizes [DJM^{+}89, Sny89]. Localizing [Smi87]. located [FKH88, GKS88, Pri87, YBBS87]. locates [OON89]. Location [DMH^{+}85, KRLU87, RSP89a, AHC89, BNT^{+}87, CC85a, DRD^{+}88, DRC^{+}86, HS86a, JFP89, MSY^{+}87, MMS89, MHD^{+}87, MBSMD86, PL85b, RL89, SNCY87, SGS86, TT88a]. Locations [LLAS85, RZG86]. loci [Kam88, TSK88]. locomoting [DNI^{+}88]. locomotion [GRvdM89, PJ87, RGS88, Wan85b]. locomotory [HCKJJ88]. locus [BGT^{+}89]. Long [LDKG87, RMS88, AIJ^{+}89, CTSF86, HLW^{+}89, HH86, NFW89, NMM^{+}86, Sow86]. Long-lasting [RMS88]. long-lived [Sow86]. Long-term [LDKG87, HH86]. loops [BVHF89]. lose [ZMH89]. Loss [MSS^{+}89, RPW^{+}89, BMVB89, BA87, DYT^{+}87, GM86, RSP^{+}89b, WS86c, RFMR86]. lost [OH88b]. Low [CdCP^{+}89, YB87, BNLC85, CMT^{+}88, DTW^{+}85, DMGS88, DF88a, DCB^{+}88, FH86, FWF88, GTN86, GRCG86, HKKP86, KKSK88, KKSK86, KBK86, LRF89, MMS89, PSBN85, PMC^{+}88, PBH^{+}87, PBMB85, PMJ87, SYT^{+}89, TFBW85, WW87b]. low- [GRCG86]. low-pH [WW87b]. low-sulfur [FH86]. Lowering [NYG87]. Lowicryl [BTR^{+}86]. luciferase [GKS87b]. Lumenal [BNT^{+}87, LG89, SDS87]. luminal [SSSP85]. lung [BFKF86, CCM86, DLH^{+}89, GTP87, HDP^{+}85, HS86a, KOBLM88, LSAKO86, Mac89a, RPW^{+}89, WB85]. lutropin [MSC^{+}89]. lymph [BNL89, GR89, JBdlTB87, SRB88]. lymphatics [MBM85]. lymphoblastoid [BHB^{+}87]. lymphoblasts [BPC89]. Lymphocyte [DS88, JBdlTB87, PE85, BKBR88, Cha85, FSM88, GCGR85, GLH85, HM89c, LSDP86, PNB89, SRB88, WGPH^{+}89, WJSB88, YSR87, YBSR87]. lymphocyte-endothelial [WJSB88]. lymphocytes [AJAW86, BRS87, CP87a, DW89, DWS88, DSGS87, GR89, HPO88, KCK^{+}87, SNC^{+}86]. Lymphoid [BPL87, GTN86]. Lymphoma [BSK86, BWSB86, GKS87a, HMP85, KB88, RR87, RdRBR89, BSNG85]. lysate [WW89]. lyse [CC85b, GBC^{+}86]. lysed [BA87, GAB86]. lysin [BIES89, IS88]. lysine [VKP89]. lysophosphatidylcholine [GBC^{+}86]. Lysosomal [BP87b, CGD86, GF86a, LHvF^{+}87, LSF86, BNLC85, BNB^{+}88, BGHvF87, BNS85, BGF86, CMW^{+}85, CIK^{+}89, EFR^{+}89, GWG85, GSS^{+}85, GSS^{+}88, GZGM87, LSF89, LGM^{+}85, MCD85, RWE^{+}88, RH86a, SM89a, TLPTV85, WG87a, WFL87]. lysosomally [DMM89]. lysosome [CIK^{+}89, Heu89a, MBvS^{+}89, RYR^{+}89, WM85]. lysosome-like [RYR^{+}89]. lysosomes [CGD86, CSD89, DCQB88, FTLSS88, FBPS87, FKH88, HT86, MK87, ROCK86, RSF86, SE85, SBS87b]. lysosomotropic [SDZB^{+}85, WFL87]. Lysyl [FBCQPR87, KVB^{+}86]. lytic [MSY^{+}87].

M [BF88, FONW88, GCPE85, MC89a, NFW89, OMM87, PCBD89, Pri87, WLJM^{+}89]. M-band [GCPE85]. M-discs [Pri87]. M-protein [GCPE85]. M2 [HGL88, PHL89]. Mac [ABME88]. Mac-1 [ABME88]. macroaggregate [BT86]. macroglobulin [MFB^{+}89]. macromolecular [SJ87]. macromolecule [MN87]. macromolecules [DF86a, DSHW85, SSC86]. macronuclear [AKC^{+}86, OO87, RLP85]. macronuclei [CRC^{+}87, WAG^{+}89]. macronucleus [RSR^{+}88]. Macrophage [MS88b, GBB89, HM89c, KWW85, SKAW86, TSBW86]. macrophage-lymphocyte [HM89c]. Macrophages [SNSS87, BNS85, CR86, DMGS88, HS86a, HCS89, ICY^{+}88, KM87, MSW^{+}86, MS88b, MCG88b, OCP86, PUW^{+}88, RT87, RH86a, SSS88, SYS85, SBS87b, WG87a]. Madin [ASJ^{+}88, BAD^{+}89, BBGI85, BPP^{+}89, BKK87, BHM89a, FGF88, FS86b, GDRB89, HM89c, JY86, LLSRB89, MFKVK87, MCA^{+}89, MBH87, NV86, NV87, NH89, OS88a, PPB^{+}89, PN88a, PN88b, PBH^{+}87, PFS85, RIPS85, RIS87, RG86a, SGRM85, SG88b, SAGM88, UPL^{+}87, VSSGRB87, VSSRB87, ZMH89, vMSvRS87]. MAG [MS86b]. magnesium [BA87]. magnetic [JT85, VA85, ZV89a]. main [TC86, YKCR88]. maintain [CMK87]. maintained [CR87, OS88a, PCK^{+}87]. maintenance [CMM^{+}85, GSG88, LBO^{+}85, NR89a]. Maize [RB85]. Major [DYT^{+}87, GM86, AKVF88, AM86, BHSJ86, BLR88, BBB88, BP87b, BFK87, CVWZ89, CHH^{+}88, DRRB88, DMRB87, ESM^{+}87, FS88a, HSK88, HHES88, HHD^{+}86, KTF87a, LSF86, MMS89, NSY86, PN89, PHRM87, PBH^{+}87, RFZH85, SNCY87, SKMB86, SGS86, SDS89, SS88a, SKDF88, WFE89, WE86a]. making [CB89]. malaria [WS89]. malarial [HUA^{+}86]. male [HBB^{+}87]. males [Kub87]. malignant [KF89b, TF86]. mammalian [AB89, BFR85, BGS87, BCD86, CMK87, ES86, FHS^{+}86, GW88a, GLC88, HP87a, HSK88, HC86, HMS^{+}89, Ibr87a, Ibr87b, KKB89, LC88b, LWM^{+}88, LKF87, MFM89, OTS86, PL85a, PL86, PLC87, Pri87, RTK^{+}86, RZG86, SL89, SBH^{+}88, SS89d, TGNG87, TDR88, TB89b, VB85a, VBWC88, WVLC86, WS85, WS86b, WFL87, WDR89, dLG86b]. mammals [MIHS86]. mammary [BAD86, CW88, JNR^{+}85, JRSB87, JRB88, KGRP85, LS85, PCK^{+}87, RJB86, Rap89, RR88, RBGF89, SB88, SAA89]. mammotrophs [RFZH85]. manner [SHS^{+}88]. Mannose [BGF86, GF86b, BNB^{+}88, BGHvF87, DK88, EG89, GF86a, GSS^{+}88, MGK89, Pfe87, SGG^{+}89, SBG89, NFN87]. Mannose- [BGF86, BNB^{+}88]. mannosidase [LSF89, SR86b, WK85, TSOC89]. mansoni [CC85b, FMK^{+}86, GBC^{+}86, SC85b]. many [GDM^{+}88]. MAP [KRN^{+}86, FONW88, MTG88, DNSMA88, GS86a, GM85b, KRN^{+}86, OSS^{+}89, PSV87, SPV88, SWTB88]. MAP-1 [KRN^{+}86, KRN^{+}86]. MAP-1-like [SWTB88]. MAP-1B [DNSMA88]. MAP-2 [GM85b]. MAP1 [LSC86]. MAP1.2 [ALCG88]. MAP1B [NLC89]. MAP2 [Hir86, LVSC86, NLC89]. MAP3 [HABM85]. mapped [NMM^{+}86]. Mapping [CB88, EDD^{+}86, GDO^{+}85, KWLS88, KL89b, KGHK^{+}87, NB89a, DBWL87, FHE88, RB85]. marginal [BS89, KMKS87, MGW86, WE89]. Marilyn [Rud88]. mark [RS88]. marker [GSF^{+}88, HE86, PSM89, VSSGRB87]. markers [BBH85, GTB85, YKSR89]. marmorata [CCLC89, KCN^{+}86, KCN^{+}89, KRLU87, MMM^{+}85]. Mass [KTP^{+}89, BIES89, HKKP86, HTT89, LN88, MRWJ88]. mast [CHG87, HDP^{+}85, HCG87, Mac89a]. material [BGH87b, CBL85, HPM87]. maternal [NKA^{+}87, SBG^{+}85]. mating [KdWB^{+}89, RLP85]. matrices [FKP86, KSF^{+}88, KG85a, RLB^{+}89, RMS86, TRB86]. Matrix [HvdSdC^{+}89, AYCY89, BS88b, BBW86, BMGI85, BB85, BLK^{+}88, CSM^{+}89, CS88a, CR88a, CTR86, CWZ89, CTS^{+}87, CPHM86, CHH^{+}88, DW86, EHK86, EMW^{+}89, FGO^{+}85, FSM88, FS88a, GDM^{+}88, GRvdM89, GW87, GPSC86, HBSQ^{+}85, HNR87, HBC89, HRP87, HCE88, HVR^{+}85, IF89, JRSB87, KIKS87, LC88a, LBFS88, MPT88, MCB85, MLM85, MLE87, Mos88, MLC^{+}88, MGHG^{+}86, NSF^{+}89, OD88, PCK^{+}87, PL87, PJ87, RYFM89, RKA^{+}89, RLB^{+}89, SVF^{+}89, SB88, SK88, SSK^{+}88, STF85, SEB89, SHL87, SNSS87, SSS88, SH86, SCD88, TIS^{+}86, WW88, WCPK88, WS87, WE86a, WRR87, WMH^{+}88, dGRS87]. matrix-assembly [MLM85]. Matrix-associated [HvdSdC^{+}89, CHH^{+}88]. matrix-binding [JRSB87]. matrix-bound [Mos88]. matrix-degrading [SEB89]. Maturation [GKR87, ASZ^{+}88, BS85, CJH85, FJY86, GF87, KKB89, MND^{+}87, MWS85, NRL89a, NGI^{+}89, RO85, SCP85, WHL^{+}86, WK85]. maturation-dependent [RO85, WHL^{+}86]. maturation-promoting [SCP85]. mature [BNS85, SEGS85, SCP85, TAMR85]. maturing [CM86c, ORA^{+}86, vZC87]. may [HFF^{+}85, LL89a, MBC86, ODD85, SL85d, WSJ85]. MCF7 [CMB^{+}87]. MDCK [NH89]. means [RGW^{+}89]. measured [NMKS85, Sch85, ZV89a]. Measurement [FWF88, HTTFC85, MMMP87]. measurements [BFRT87, CJO^{+}89, DJS^{+}88, GW86, LBEM89]. Mechanical [SSSP88, Kol86, PE85]. mechanically [MMWD89]. Mechanism [KWPH89, LHFB88, WB85, AT86a, ALWL89, BGM^{+}87, CPC^{+}89, DLG^{+}89, ETA85, GDYW86, GKB87, HB88b, KR88, LHNF87, MT86, MMC88, PC85, RMS86, SIYT86, SSP89, SPH87, TFD87, VSSRB88]. Mechanisms [HSMM85, Ree85, AF89, DS88, GE88, HDP^{+}85, HG89, HGMS88, OR85, PSBN85, PML^{+}89, SGB87, WS86c, WJSB88]. mechanistic [BBGI85]. Mechano [SAB^{+}89]. Mechano-chemical [SAB^{+}89]. Mechanochemical [IF89]. mechanoenzyme [GCC^{+}89, MC89b]. medalist [Rud88, Pol89]. media [HA89]. mediate [BNF^{+}87, EMH^{+}87, ML85, PGH^{+}88, RVS88, SL85d, TRB86, VFB^{+}89]. mediated [AR88, AAM^{+}87, BNLC85, BMVB89, Bis86, BD86b, BMGI85, CSL87, CWH86b, DJ89, DZ85, DMGS88, DCH86, FPS^{+}85, GF86b, GFSS86, HAB^{+}88, Hat87, HA89, HGMS88, JNA85, Knu85, LCD^{+}89, LCW^{+}89, LTG85, LCR^{+}87, MFKVK87, MSS^{+}87a, MSW^{+}86, Miy88, MH86b, MUM87, MTA86, PP88a, PBvTS88, PBMB85, PRB^{+}89, PMR89, SS89a, SJ87, SS88c, TRL87, WH85, WSV^{+}89, YB87, KMR^{+}88, Mac89a, NR88, PSD^{+}85, RRB86, RWS^{+}85, VG86b, vAAA86]. mediates [BR89, CR88a, ELKH88, GR89, GGA^{+}89, HAK^{+}86, Ibr87b, KC88, RSG87, SCC87, SRW^{+}89, SS89d, TKK86, WW89]. mediating [SL85c, SS88a, WCPK88, WS89]. medium [DWL86, HSS85, LKOM88, PBH^{+}87]. medullary [vGRB86]. Meeting [PPB^{+}89]. megakaryoblastic [ITY^{+}88]. megakaryocytes [DBC^{+}87, MB85b]. meiosis [Kub87, WSR89]. Meiotic [CJH85, DG88, CC89]. MEL [KWS^{+}89, YBSR87]. MEL-14 [YBSR87]. melanocyte [DDB^{+}88]. melanocytes [DMG^{+}89, HLB^{+}88, JVG^{+}85]. melanogaster [CF88a, CNL86, DM87, GLM86, HMG^{+}86, HS87a, HS87b, KF86, MASA89, SCBHS^{+}89, TMP87]. melanoma [BLK^{+}88, CK86, CPHM86, CPP^{+}87, GLL^{+}86, MHF86, NZ85, RSG87]. melanophore [RH86b]. member [GCC^{+}89, GCR^{+}89, GKZ89, GGA^{+}89, PSK^{+}87, Ran88]. members [BKA^{+}89]. Membrane [AMWN89, ASCM89, BA86, CE89, DLH^{+}89, EMH^{+}87, FRH85, GP86, LP85b, MRD^{+}89, PSD^{+}85, RWF^{+}87, RPW86, SR86b, Abr85, AP86, AFMD89, AdS85, AHM89, ACPFB^{+}85, AHRM85, AM86, AT86b, AR85, BGN^{+}88, BEHJ^{+}88, BNLC85, BBH85, BFSH87, BVHF89, BD86a, BFA88, BGL^{+}88, BWS86, BPKS86, BWSB86, BSH^{+}85, BGH87b, BP87a, BLK^{+}88, CWW89, CR88b, CTR86, CS88b, CCLC89, CC85b, Cha85, CM86b, CMW^{+}85, CA86, CG86b, CCS^{+}86, Cou87, CMK87, CSL87, CJ85, CIK^{+}89, CG86c, CMG86, DFSK87, DWS88, Den88, DS89, DJBD85, DCQB88, DF86b, DB85, DRRB88, DMRB87, EMT88, EWAP85, EHM86, FTLSS88, FRG^{+}87, FY86, FMK^{+}86, FBB^{+}88, GW88a, GORW88, GM85a, GWM85, GB87b, GDYW86, GSS^{+}88, GCT86, GZ89, GBC^{+}86, GSF^{+}88, GML^{+}89, GLK^{+}89, GZGM87, GM85c, GPSM85, HFGD89, HdAC89, HD86]. membrane [HP87a, HCHK89, Hat87, HSB^{+}85, HT86, HFG87, HHR^{+}89, HS85b, HHES88, HB87, HFB87, HAM86, HLU^{+}87, HSMM85, HBB85, HFB^{+}89, HGL88, HJV^{+}89, ISL87, IHJ88, ITY^{+}88, JPU89, JM86, JRSB87, JRB88, JBTA88, JTBA89, JBG^{+}87, KA85, KS85a, KS85b, KKSK88, KB86, KKB85, KBO^{+}88, KBB^{+}88, KdWB^{+}89, KCN^{+}89, KP89, KJW89, KKML88, LHK^{+}89, LS87a, LGM^{+}85, LD86, LD88, LFHD89, LSF86, LP85a, LCDRB89, LDP87, LBFS88, LMK88, LG89, MT86, MSQ^{+}89, MBD89, Mat86, MAC89b, MP86a, MP86b, MOT^{+}85, MOU88, MVSW86, MAG88, MTRR89, MBK89, MLK^{+}87, MVK^{+}88, MCC^{+}89, MB87, NBS88a, NJD^{+}86, NL85, NV86, NV87, NH89, NKMK86, NYRK87, OMS89, Oss88, PPG85, PHL89, PCK^{+}87, PN89, PW86, PKS^{+}89, PBD^{+}89, PSR^{+}88, PHN^{+}88, PMV^{+}89, PTB^{+}85, PHTK87, PMR86, RH89, RGS88, RPRS89, RSF86, RS85]. membrane [Rog87a, RG86a, RSF^{+}85, RSL^{+}89, SC85a, SVSH^{+}88, SCLG86, SYSP88, SM89b, SJOB89, SMC89, STE88, SCOE^{+}86, SHP^{+}87, SG88a, SJS86, SW85, SABVK89, SPL87, SV89, SRW^{+}89, SSS88, SMS^{+}85, SALN88, SF87, SKC85b, SSK89, TLP^{+}87, TLPTV85, TTSSH85, TT89b, UP86, VMT^{+}88, VH88, VSSRB87, VSSRB88, VSKSP86, VGKF86, WESP85, WMV86, WCB86, WKS85, WMBR89, WO86, WSH85, WCMG88, WN88, WS89, WHI86, WSM86, WHL^{+}86, WE86b, WBVL86, WBB89, WBDR88, YBBS87, YR87, dCFB86]. Membrane-anchored [ASCM89, HD86]. Membrane-associated [DLH^{+}89, BWSB86, JRSB87, MP86b, MOT^{+}85, PKS^{+}89, PHN^{+}88]. membrane-binding [CG86c, CMG86]. Membrane-bound [EMH^{+}87, JBG^{+}87, MVSW86, UP86, WMV86]. membrane-coating [BGH87b]. membrane-cytoskeletal [NH89]. Membrane-cytoskeleton [MRD^{+}89]. membrane-impermeant [SSS88]. membrane-like [CTR86, GLK^{+}89]. membrane-membrane [HS85b]. membrane-microfilament [VGKF86]. Membrane-related [LP85b]. membrane-skeleton [NL85, NV86]. membrane-spanning [LD86, RG86a]. membrane-specific [MAC89b]. Membranes [GSS87, AIJ^{+}89, AK86, ADS88, ASF89, ACPFB^{+}85, ABC^{+}88, AYG85, BF88, BHB^{+}87, BCPC87, BP87a, CCC86, CM88b, CCG89, Den88, ES86, GM85a, GLH85, HDB88, HW88, HOPM^{+}87, HCS89, HCHK89, HvdSdC^{+}89, HM89a, HHD^{+}86, KGRP85, KDB87, KHC^{+}86, LE86, LHFB88, LM85b, MGK89, MM87, MMM^{+}85, NRTE89, NBS88b, PPG85, RIPS85, ST86, SHH85, SL86a, SL88, SDS87, SSZ88, Ste85, SGP88, SB85f, SBC^{+}87, TS88, TSBW86, TB85b, TSS89, TACQ89, TSOC89, UTK88, UP86, VB88a, WWW^{+}85, WMBR89, WTB87, WL87, YE87]. membranous [Pum89]. meningococcus [RDB^{+}86]. meromyosin [SDMF85]. mesenchymal [ACEE87, AE88, FM86, VJT89]. mesenchyme [ACEE87, AE88, KH85a, VJT89]. Mesosecrin [RJH^{+}87]. mesothelial [GKF85, KSJF87, RJH^{+}87]. messenger [CBS^{+}89, CS88d, HD88a, IGSGD87, LW86, LGC^{+}88, PRS^{+}88]. meta [BOGK88, SC87]. meta-vinculin [BOGK88, SC87]. Metabolic [DTW^{+}85, HP86, HB88c, TT89b]. metabolically [WDR89]. Metabolism [Mos88, BHS88, CR87, DTW^{+}85, EMT88, GMCH87, KP89, MRO^{+}86, PSBN85, Ros89, RL89]. metalloendopeptidases [WHMA86]. metalloendoproteases [RFSL88]. metalloprotease [PCS88, SEB89]. metalloproteinases [VKB89]. metallothionein [TYP^{+}87]. metaphase [CC89, CM86c, CM87a, FS88a, GB89a, Hep85, MHW85, VB85a, WS86a]. metastasis [RdRBR89]. Metastatic [NZ85, MHF86]. methanesulfonate [ATNH86]. method [DSHW85, KSKW^{+}85, MM87, PAGPLM86, SP86, TTM88, VA85]. methodology [KYS^{+}88]. methods [SWW85, WBW87]. methylglutaryl [JBG^{+}87, KPK86]. methyltransferase [LHM89]. mevalonic [BHS88]. Mg [DMF^{+}87, SRW^{+}89]. MgATPase [CM87c]. mice [BNS85, CSH89, FJR87, GFHDD89, Gor89, MRO^{+}86, PP88b, TFS^{+}89, TNG89, WTBBW85]. microbeam [WIS89]. microcultures [ELKH88]. microdomain [GLL^{+}86]. microdomains [AIJ^{+}89, JBTA88, SSSP85]. microelectrodes [Sch85]. microfibrils [SKE86]. Microfilament [BBP88, BS89, BPC89, DCL^{+}87, GBBS89, Kol86, LFC^{+}88, Rog87a, TTKM86, TIT89, VGKF86]. microfilament-associated [BS89, GBBS89, TTKM86]. Microfilaments [SFW87, FL88, GGJ^{+}87, KTT^{+}86, LL86, RS85, Str86, UF89, WG88b, YMM88]. microglia [GB85b]. Microheterogeneity [CF88b]. Microinjected [BG86, JMW88, MJW86, OH89, SMPS86, WKR86]. Microinjection [CRR87a, CR87, CBS^{+}87, MBKR89, BK86, BRR^{+}87, GHL^{+}89, HLL89, MNC86, OH88a, PJFB89]. Micromanipulated [CNL86]. Micrometer [YE87]. Micrometer-scale [YE87]. micronuclear [TMH^{+}85]. Micronuclei [RCC^{+}88, OOCT89, WAG^{+}89]. microparticles [ZV89a]. microphotometry [AJRNBJ85]. microprocesses [CK86]. microribbons [PAHC89]. microscope [BTR^{+}86, DD88, FLM86, NMKS85, SMH88]. microscopic [CW87, DMRB87, HVBB88, JS86a, JS86b, KLS85, MS86b, MS88a, OH89, PTW^{+}85, PCC85, PS87, PP89, RBMM88, RMM89, STE88, TMS85, VG86a, WW88]. microscopical [JJ87, KPK86]. microscopy [Ake89, Bäc88, BBP^{+}85, BSK89, BCPC87, BFRT87, BPE^{+}86, CS89, FYRK88, FKP86, FONW88, GDO^{+}85, GVD^{+}87, GB86a, HR87, JIFH85, KL89b, LWT85, LGBE89, LAA85, LU88, MYH^{+}85, MP86c, MF87a, MASA89, MMS89, MSS87c, PSM89, PBS85, RGM85, SDMF85, SH87b, STGK87, TB85a, TCP^{+}85, UTK88, UP86, WOP^{+}88, WAF87, ZSE85]. microsomal [BNT^{+}87, ES86, GW88a, HDB88, HW88, SW85, SSK89, WCB86]. microsomes [ACPFB^{+}85, KPK86, OHK86]. microspectrofluorometer [FWF88]. microspikes [GLL^{+}86]. microsporidian [PW85]. microtubular [SGBG87]. Microtubule [BH86a, CBL85, CFB^{+}85, GK87a, GHL^{+}89, GSB88, MP86d, MTS89, OH88a, OK86, PBBP85, RIS87, SGB87, SK86, VT89, AC86a, AHC89, AG87, ALCG88, AK86, BBB89a, BS89, BAG86, BKK87, BM88, BSG85, CWS86, CPS88, DMD^{+}86, DNSMA88, DFSK85, DKKK88, EKH89, FJMW87, FCH^{+}87, FCH^{+}88, GK87b, GM88, GLM86, GM85b, GDC^{+}86, GLC88, GKB89, HJS^{+}85, Heu86, HBV85, HTH85, Hir86, HH87, HC86, HM89b, HABM85, HTB88, HK88b, ICS88, JS86b, JS86c, JPM85, JFP89, KTO^{+}89, KFA89, KMKS87, KMK86, KRN^{+}86, MHW85, MMC88, Mit89, MKA^{+}86, NB88a, NLC89, PSV87, RGM85, RGBM87, Sal86, SS89a, SM87, SABK87, SH87b, SPV88, SB85e, SPH87, SWTB88, SD88, SRW88, TMB85, TPBB85, TB88b, TTKM86, TMH^{+}85, TT89b, VH88, WIS89, WFW87, MK85a, MK85b]. microtubule- [BS89, TTKM86]. Microtubule-acting [RIS87]. microtubule-activated [PSV87, SPV88]. microtubule-active [EKH89]. Microtubule-associated [CFB^{+}85, PBBP85, AC86a, AG87, ALCG88, BM88, BSG85, DNSMA88, GK87b, GM88, GLM86, GDC^{+}86, HBV85, HTH85, Hir86, HH87, HC86, HABM85, JS86c, JPM85, JFP89, KFA89, KRN^{+}86, MKA^{+}86, NB88a, NLC89, SH87b]. microtubule-binding [AK86, GM85b, JFP89]. microtubule-capping [SD88]. microtubule-depolymerizing [JS86b]. microtubule-organizing [DMD^{+}86, TMB85]. Microtubules [AC86a, GB89a, PLC87, SMVS^{+}86, TCF86, AFMD89, AG87, AWH^{+}85, AYG85, BMB86, BPC89, BKK87, BB85, BRP88, CV87b, DMD^{+}86, EMK85, FS88b, FS87, FCH^{+}87, GHL^{+}89, GS86a, GSB87, GBBS89, GKB87, GW88b, HSO88, HC86, HFB87, HMN87, HPM87, JASP88, JYC87, KBR87, KGB88, KTT^{+}86, KS85c, KRN^{+}86, LP86, LSB89, LAA85, LS87b, LC87, MSR^{+}86, MGR86, MK87, MP88b, MKH88, ML85, MTS89, NG85, PBBP85, PSV87, PWS^{+}89, Pra86, RGS88, RGM86, RNW89, SF88a, SSSP88, SSS^{+}87a, SSRS88, SK87, SSS^{+}87b, SH87b, TWB88, TB87, UF89, VFB^{+}89, VCM88, WSRR89, WOP^{+}88, WIS89, WG88b, WMC86]. Microvascular [HD85, KDHS87, DHH89, PSSP88, YBWLM88]. Microvesicles [NDJ^{+}89]. microvillar [CCA^{+}88, CB87, CB88, CB89, EB87, GCBH86, MCC^{+}89, SCM87, SMB^{+}88]. microvilli [RSF86, SH87a, TTM88]. microvillus [CM87c, GCC^{+}89, GSF^{+}88, HSB^{+}85, MC89b, VB85b]. midbody [CHE87, SK88, TB87]. middle [WDF86, YBBS87]. midpiece [TI87]. midportion [HKF^{+}87]. migrating [DWFT88, ES89, PVMO87, PSC^{+}89, PHM85, ZBG89]. migration [AYCY89, APS87, BF85, BFL88, CMB^{+}86, DMH^{+}85, DRC^{+}86, DLFE89, FM86, GW87, HMTF88, KH85a, KW86a, KCR86, MKH88, MCC86, PJ87, RMS86, TRL87, WMC86, YBWLM88]. Mild [JWH85]. milk [GM85c]. mimicked [ST86]. mineralization [GCK^{+}88]. mineralized [RYFM89]. Mineralocorticoid [CW88]. mini [CNL86]. mini-spindle [CNL86]. minifilaments [SSP89, TL87a]. minus [COA85, MBH87]. MIP [YKCR88, YKCR88]. miscompartmentalization [CSFS89]. misfolded [HBHL^{+}89]. missorting [RWE^{+}88]. missorts [EFR^{+}89]. Mitochondria [OR85, VS88b, ABC^{+}88, DCWJ89, HJV^{+}89, KBPW89, MPB88, PV85, PSR^{+}88, PMV^{+}89, RGW^{+}89, SCOE^{+}86, SHK^{+}87, VME85, VS88a]. mitochondrial [ASF89, ABC^{+}88, EVG^{+}86, HFF^{+}85, HKF^{+}87, LHM89, LBFS88, MMWM^{+}89, NAS^{+}87, NBS88b, SHN87, VP85, VS88a]. Mitogen [HPO88, HLB^{+}88, MBD89, SHS^{+}88, UPRS88, WD86b]. mitogenesis [CCK^{+}88, FSM88]. Mitogenic [MER^{+}86, PMR89, SNNK85, SHGD85, TT86a]. mitogens [DF87, GCR^{+}86, RDSL89, SKK^{+}85]. Mitosis [DN88, ASZ^{+}88, CHE87, GM86, GB86b, Hep89, HHSF85, JGYG85, OSP^{+}85, PRS^{+}85, RMS88, SL87a, SF89, SALN88, TB89b, TMH^{+}85, VB88b, WS86a]. Mitosis-inducing [DN88]. Mitotic [LBW89, BSA87, BRR^{+}87, BCD86, CM86a, EH85, EHC^{+}85, FGW85, GGN^{+}86, GL85, HHY88, HTH85, KG89, KB85, LPBW87, Mit89, RG88, Ros89, SROB86, SBS88, VLD^{+}85, VB88b, WMC86]. mitotically [Eva89]. mixed [BDWH88, Gla85b, MEBV^{+}89]. mixed-phase [Gla85b]. mixing [SME^{+}89]. Mo1 [AGPT88]. Mobility [VME85, Wan87, AELE88, BHB^{+}87, CLF^{+}88, DJM^{+}89, DNI^{+}88, JY86, LSP86, LBP^{+}86, MHWB86a, MVSW86, NYRK87, SCRA87]. mobilization [MRD^{+}89, OSD85]. Mode [SAB88, HSO88]. model [DCR89, DCQB88, GM85a, HBV85, JPM85, Kee87, Koz89, TLZ88, UP86, WS86a]. models [BA87, GAB86, HSMCL87]. modification [BH87, BAHE86, BD85, BBP88, HM89b, JS86b, KLHS85, LR85b, LC88a, SABK87, SB87]. modifications [BRP88, CDB87]. modified [BFA88, BFPS85, GKB89, Oss88, PBMB85, SCLG86, SLL88, TS88, WS89]. modifies [SKAW86]. modular [Fra87]. modulatable [CRJM88]. modulate [GHW^{+}88, NHOY85]. modulated [BBP88, BSG85, CWZ86, DG89, MVS88, SG88b, TB89b, YIJ88, ZV89b, MSS^{+}87a]. modulates [GG87b, JAT86, Mas85, MUH89]. Modulation [APBP87, BN85a, HCKJJ88, Ip88, KW86a, LFFW89, MLC^{+}88, NV87, PRS^{+}88, TTLR89, VSSRB87, BHB^{+}87, BB88b, CCBKJ89, CCK^{+}88, DJ89, HB88c, HSMM85, LMF^{+}86, MPT88, MOP^{+}86, SD87, VG86b, WSM^{+}87, YBWLM88]. moieties [CMW87]. mol [LC85, MLM85, SNNK85, Wan85a, YLGG85, YSW^{+}85]. mold [WD86a]. Molecular [CSGC88, CMGS86, ESM^{+}87, EHK86, FL89, FCMK^{+}86, GVV85, HCE88, HML88, LSP86, LBB^{+}88, OB88, PSM89, Pau86, SJOB89, SCOE^{+}86, SYT^{+}89, SQM89, TSK88, WMR86, WVCR89, AED^{+}89, AJCMR88, AYG85, AE88, BICFA88, BIES89, CMT^{+}88, CWZ89, CKR^{+}85b, DW86, DDR^{+}89, EMW^{+}89, GLW^{+}89, HKKP86, HTT89, HSO88, HSK^{+}89a, KTT^{+}86, KB85, LMD^{+}86, LN88, MCS85a, NB88a, PMJ87, QCKR85, QS87, Rob87, SSMG86, TRL87, TASK86, TRB86, TT85]. molecularly [Rog87a]. molecule [AR88, AYY86, APS87, BVK85, CS87, CG89, CS88c, CLC^{+}86, DS88, FGE86, GHW^{+}88, GE88, GSG88, HR87, HTRBP86, HNNT88, HFG87, Heu86, KWLS88, LCBS89, MOU88, MTK^{+}87, MHP^{+}86, MOP^{+}86, NMM^{+}86, NT86, NAD^{+}85, PRB^{+}89, PSGS85, POD89, RMP^{+}88, SR89, SLE86, SPS89a, SPS89b, SYT^{+}89, TDGE85, vdPdPR86]. molecule-1 [DS88, LCBS89]. molecules [CGA89, CCM86, CE85a, CE85b, CCE87, DLG^{+}86a, DDH^{+}87, FG89, GSS89, HNH^{+}87, HCE88, KFS87, KJF^{+}88, LFR85, LCE87, MSM88, MS86b, MS88a, NFW89, PS87, PNB89, PSK^{+}87, RMM87, RE85, RDPR^{+}86, SC85a, SS88a, SQM89, WFE89]. mollusc [BOW85]. Monensin [OMY^{+}85, WMUP^{+}85]. Monensin-resistant [OMY^{+}85]. monitored [SME^{+}89, ZSE85]. monkey [RGPRB87]. Monoclonal [BCC^{+}87, DSA86, DHL85, FTC87, HW86a, HMdN^{+}88, ISM85, MNC86, MVSW86, PG86, PSP85, PWS^{+}89, PNB89, PF85b, RG87, SSG87, TH89b, TTSSH85, VMT^{+}88, WRS^{+}86, AIJ^{+}89, AYCY89, AKC^{+}86, BCA^{+}86, BBGI85, BWS86, BPKS86, Bro85a, Bro85b, BF85, BFL88, CMP86, CCBKJ89, DBWL87, DSTJ85, DHB86, EDD^{+}86, FDG88, FPS^{+}85, FR86, FH86, FONW88, GDO^{+}85, GVGL85, GAH86, HKKP86, HLL88, HLL89, HvdSdC^{+}89, HBB85, ICS88, Ip88, IHJ88, JNR^{+}85, KKSK88, KBL^{+}86, KGHK^{+}87, MBKR89, MAG88, OBCO85, Rob87, RDB^{+}86, SSJ^{+}85, SFB^{+}88, SDMF85, SRT^{+}86, SKC^{+}86, TJ85, TES^{+}87, WFR86, WKKS^{+}87, WCPK88, WKRH86, WLJM^{+}89, ZRW85, ZG87]. Monocyte [DWL86]. Monocyte-conditioned [DWL86]. monocytelike [PKN^{+}89]. monocytes [CWS86, DRME89, JNA85, PP88a, SNC^{+}86, SSKK89]. monocytic [BS86, PBK88]. monocytogenes [TP89]. monoglucosylated [SBG89]. monolayer [FGF88, MD85, Qua85b]. monolayers [DBD^{+}89, FOW^{+}85, MMMP87, MB87, RG86b]. monomer [BM86]. monomeric [JNA85]. monomers [GGMH86]. mononuclear [JNA85]. mononucleated [HDL^{+}86]. monooriented [RDJ^{+}86]. monospecific [SKDF88]. months [PP88b]. morphogenesis [BT86, CCM86, DKKK88, DMB88, Eck87, FRH85, TQN^{+}87, CE85a, CE85b]. morphogenetically [ET87a]. morphologic [SMA85]. Morphological [WS85, Bre89, CTYF85, FLBP^{+}86, GJD^{+}88, HCH85, KT85, KKML88, MWSF87, MQ89, NTNT85, TCBF^{+}85, PN88b, TTKM86]. morphologically [CIK^{+}89, YM87a]. Morphology [SFMT87, CKR^{+}85b, DH89a, Hat85, Heu89b, HCKJJ88, NCB88, RBKK85, WBW87]. most [SBG^{+}85, TWB88]. motif [NLC89]. motifs [NRL^{+}89b]. motile [Heu86, MR85b]. Motility [CS86c, HLL88, ABG^{+}89, BN88, CDL^{+}85, ES85, FMG89, FPS^{+}85, GTG85, ICS88, MHF86, RH86b, THM86, TLZ88, VES85, dLG86b]. Motility-dependence [HLL88]. motility-promoting [MHF86]. motion [GSB88, HW85b, HMN87, KCAF89]. motional [AYG85]. motions [VA85]. Motor [MSM88, BD86a, BD86b, DFdN^{+}89, FC89a, HTGP85, Nic89, PBBP85]. mounts [Ris85]. Mouse [CSM^{+}89, COMS87, JRB88, KB88, LNH86, AJAW86, ASJ^{+}88, BB88a, BW86, CW88, CM86c, DJM^{+}89, FMJS89, FOW89, GF86a, GK85, GCR^{+}89, GM86, GGJ^{+}87, GLH85, HMF^{+}87, HMP85, HPM87, HM89b, HBB^{+}87, JNR^{+}85, JRSB87, KL88, KL89a, KKSK88, KWS^{+}89, KYS^{+}88, LMSM89, LLRR87, LSDP86, LLC85, LSC86, LS89, LD86, LBL^{+}85, LS87c, MHW85, MS88a, MS88b, MSC^{+}85, MAE88, NT86, OD88, OCP86, OMY^{+}85, OGJ^{+}89, PS87, RJB86, Rap89, RRB86, RKWS86, SHBV89, SB88, STE88, SYT^{+}89, SA85, SCP85, SYH^{+}85, SCD88, TT86a, TTW88, Tra88, TSM87, WG87a, WHI86, WSM86]. move [Bre88, GSB87, OR85, TB87]. Movement [TFD87, AWH^{+}85, BGHvF87, CHE87, DMC85, DK88, ER87, HLL89, HMP89, Hym89, LSF86, MWK89, Nic89, PKK89, SSH87, SR88, SR89, SSRS88, SSS85, SR85c, TP89, VH88, WSV^{+}89]. Movements [BBB89b, BKR86, KS86, RDJ^{+}86]. moving [Kol86, LN88]. mp20 [ASLFP89]. MP26 [JLH^{+}86, SSJ^{+}85]. MP70 [GKBG87]. Mr [CFC89, CS87, FMT^{+}87, HLU^{+}87, HW86b, KOF88, MB85a, VBB85]. mRNA [AVP^{+}89, BDS85, BCC^{+}87, CHH^{+}88, FKC87, FCL86, GM88, GKG88, GLD^{+}88, HW85a, HZS88, HFK^{+}89, HHMH87, HKPS89, LHK^{+}89, LMWG85, MSV^{+}88, MTK^{+}87, MOP^{+}86, OJHAH89, OMM87, OGJ^{+}89, RKWS86, SHBV89, SR86a, SHS^{+}88, SHR86, SVG^{+}86, TT86a, TLC85, WPW^{+}85]. mRNA-encoding [SR86a]. mRNA-species [RKWS86]. mRNAs [BKA^{+}89, CHS85, DMC85, HNP^{+}86, HZS88, HSK^{+}89b, MHP^{+}86, SV87, STH^{+}89, SW87a, SLL89, VKB89, VSZ^{+}87, WG85]. mt [SKC^{+}86]. mucosal [JBdlTB87]. mucus [HSMCL87]. mucus-secreting [HSMCL87]. multicatalytic [GW89]. multidrug [TASK86]. multilamellar [BHRC85]. multimeric [SDS87]. multimerization [WMM86]. multimers [WMUP^{+}85]. multinucleated [RH89]. Multiple [BGT^{+}89, BRS87, HD88a, LN88, NL86, RDS^{+}89, SLB87, WBW87, BKA^{+}89, Bro85a, CCE87, EBCO88, HKM^{+}87, KTO^{+}89, KSS86, LH86, MS86c, PSR^{+}88, WRS^{+}86, WC87, WCPK88, WPBB^{+}85]. multipotent [HSMCL87]. multistep [AG89]. multitubulin [Cle87]. multivesicular [HHS^{+}85, TACQ89]. Murine [MCG88b, BNL89, DMG^{+}89, LL88, MMK^{+}89, NMKS85, PL87, PSSP88, PUW^{+}88, RT87, SSSP85, SYH^{+}86, TLD^{+}89, VSR87, YKSR89]. Muscarinic [VAD^{+}85, OSD85]. Muscle [FW85, And86, AM85, AJCMR88, AE87, AS86, ASLFP89, BFAI87, BOW86, BCKJW88, Bis86, BGH^{+}87a, BTO88, BICFA88, BFA88, BMDN87, BB87d, CS86a, CFC89, CCMB87, CCL^{+}86, CPC^{+}89, CB86, CD89, CF88a, CLOH87, CH87, CHH^{+}88, CCK^{+}88, CM87b, CS86d, CMGS86, CCS87, CPKJ87, DTW^{+}85, DCB89, EPJBBW86, ER87, FG89, FYRK88, FHS^{+}86, FKC87, FAKVM87, FRR85, FMT^{+}87, GSS89, GFS^{+}89, GSTS86, GVGL85, GHA88, GCPE85, GRME^{+}89, GAH86, HFHR88, HBC89, HS85a, HZS88, HBF^{+}88, HD85, HW86a, HKPS89, HSS89, HP87b, HMP89, IKSF89, IF85, JMW88, JSC85, JJ87, JSA^{+}89, JSCLP89, KCAF89, KFHF85, KBKC86, KST^{+}89, KSSAQC89, KKT^{+}86, KLS85, LOG85, LC88b, LL86, LL89b, MCS85a, MB85a, MCB85, Maj87, MGD88, MCG^{+}88a, MACO88, MAK^{+}86, MYH^{+}85, MMW85b, MMW85a, MCS85b, MTK^{+}87, NFI^{+}87]. muscle [NPRR86, OB88, OH88b, OD87, Owe85, Owe86, OLGT86, OGYK88, PB86, PCC85, PF85a, PSK^{+}89, PRB^{+}89, Pri87, RNPR^{+}88, ROCK86, RRP88, RS88, SCBHS^{+}89, SMH88, SSC86, SR89, SL89, SCBR85, SS87, SP88, SSV^{+}85, SSS85, SB85b, SBS87a, SC87, SE85, SSSP85, SHS^{+}88, SRT^{+}86, SFD86a, SMA85, SKL85, SF88b, SP86, TSI^{+}89, TB89a, TRCR89a, TRCR89b, TRT^{+}88, TL87a, TL87b, Try89, TH89b, TS89, VN88, WW88, WR88, WG87b, WRL86]. muscle-derived [GAH86]. muscle-gene-specific [MAK^{+}86]. muscle-specific [GVGL85, LL89b, OLGT86, PSK^{+}89, TSI^{+}89]. muscles [BOGK88, RMM87, TOOK^{+}85]. musculature [WPBB^{+}85]. Mutagenesis [MSC^{+}89, BCMK87, GDYW86, WGPW89]. mutant [AWJ85, AF87, AF89, ADS88, ABG^{+}89, CSFS89, CSD89, DS87, EFR^{+}89, FJY86, FJR87, KKSK86, MMR88, MRO^{+}86, MKH88, MR85b, MVB87, OB88, OK86, OMY^{+}85, PMC^{+}88, RKFN88, SGM^{+}89, SBG89, SPO85, TT88b, TSM87, YM87a, YM87b]. Mutants [CG86c, CMG86, BSR^{+}89, BRKE88, EOM86, GRCG86, HSWH86, KSJV88, KBK86, LEW^{+}88, MMWM^{+}89, NFN87, RFMR86, RdRBR89, RYR^{+}89, SC86, SFK^{+}89, SFG87, UMHI85, WDV87, dVOD^{+}89]. mutation [DM87, FCH^{+}87, HHY88, KP85, RdLB87, WVCR89]. Mutational [MM89, KHC^{+}86]. Mutations [CNB89, DRSG85, FCH^{+}88, Kam88, AY89, DRM^{+}88, HTB88, SHP^{+}87].

MVS [MVSW86]. MVS-1 [MVSW86]. myac [TRCR89a]. myc [MAE88, HSK^{+}89b]. Myelin [MRO^{+}86, PSK^{+}87, BP87b, CR88b, CS88b, DDBHL86, EBBW87, HFB^{+}89, JMBM85, MS86b, MS88a, MS88c, SHC87, Tra88, TACQ89, YMD87]. Myelin-associated [PSK^{+}87, MS88a, SHC87, Tra88, TACQ89]. myelin-deficient [MRO^{+}86]. myelin-induced [YMD87]. myelin-like [CR88b]. myelinating [LFS^{+}89, TACQ89]. myelination [BBB89b, CTR86]. myeloma [WDF86]. myelomonocytic [NGI^{+}89]. Myoblast [ECZ86, EZBW88, Knu85]. myoblasts [EZWB88, GRvdM89, GZO88, HDL^{+}86, KSM89, KDIG89, PJW88, TSI^{+}89, Wri85]. myocardium [BKPG89, PSSP88]. myocytes [LHS^{+}89, MMW85a, SL89, UPRS88, WE87]. myofibril [FOW89, TM87b]. myofibrillar [CF88a, LEG^{+}89, SE85]. Myofibrillogenesis [SMPS86, HDL^{+}86, TM87a, TM87b, Tok89, WGS^{+}88]. myofibrils [ATNH86, BGF88, JMW88, LEFS^{+}87, LHS^{+}89]. Myogenesis [FOW89, FC89a, HS86b, MTK^{+}87, NL85, SFLM86, TMB85, TPBB85, WLG85, WPW^{+}85, Wri85]. myogenic [ACN^{+}89, DKB^{+}85, FOW89, GKB89, HZS88, LTG85, MS86c, OSH^{+}86]. Myomesin [GCPE85]. myosheets [ATNH86]. Myosin [EMOB85, GSTS86, GDS87, KBL^{+}86, AE87, BCKJW88, BF86, BD89, BA87, CS86a, CC85a, CCA^{+}88, CMM^{+}85, CB86, CF88a, CCBKJ89, CM85, CPKJ87, DWFT88, DBLS87, DBWL87, EPJBBW86, EOM86, FL89, FPS^{+}85, FDB85, FC89b, GCC^{+}89, GSB85, HBPK87, HF85b, HDL^{+}86, HCKJJ88, JMW88, KL89b, KCCM85, KKT^{+}86, KF86, KTP87, LFC^{+}88, LMD^{+}86, LCK88, MWK89, MCS85b, MF87a, MSS87b, MBK89, MCC^{+}89, MC89b, NYG87, NFI^{+}87, OB88, PG86, PFRS89, PSP85, QS87, RSP89a, SJ86, SSV^{+}85, SSS85, SDMF85, TOOK^{+}85, TGNG87, TB89a, TMP87, TL87a, TL87b, Try89, TH89b, VC85, VC89, VBWC88, WGS^{+}88, WE87, WTBBW85, WPBB^{+}85, YB87, ZSB86, HKKP86, SP89, SSP89]. myosin-binding [YB87]. myosin-containing [MWK89]. myosin-like [CCA^{+}88]. myosin-mediated [FPS^{+}85]. myotendinous [TOB86]. myotube [BB88a, RRF87]. myotube-specific [BB88a]. myotubes [BPB86, BPJBF85, BF87, BBMW85, DLFE89, EMZ85, LEFS^{+}87, MJW86, MCS85b, PB87, Pum89, RH89, TMS85, UF86, Wal86]. myristate [CVWZ89, FTC87, LEFS^{+}87, LEG^{+}89, LMF^{+}86, MMT85, MCCM88, PKN^{+}89, SVG^{+}86]. Mytilus [TFD87]. myxomycete [UF89].

N [MTK^{+}87, MAE88, MHP^{+}86, RGE85, BH86b, CG86a, CLC^{+}86, CS86d, CMGS86, DMW88, DBD^{+}89, DBK87, FR86, FBE85, GSS89, GML^{+}89, GZO88, HFG87, HSK^{+}89b, HFC^{+}86, HSS^{+}87, JWH85, KKSK86, MS86b, MS88a, NTLR88, NJAS89, PZBD87, PS87, PSGS85, RNPR^{+}88, RDB^{+}86, SSC86, TTW88, WFR86, WMBR89, WGA85, WFL87, YBBS87]. N-acetyl-galactosamine [TTW88]. N-acetylglucosamine [HSS^{+}87]. N-cadherin [NTLR88]. N-CAM [MTK^{+}87, MHP^{+}86, RGE85, CG86a, CLC^{+}86, CS86d, CMGS86, DMW88, DBD^{+}89, FR86, FBE85, GSS89, HFG87, HFC^{+}86, MS86b, MS88a, PS87, PSGS85, RNPR^{+}88, RDB^{+}86, SSC86, WFR86, WGA85]. N-dodecylimidazole [WFL87]. N-ethylmaleimide-inactivated [JWH85]. N-ethylmaleimide-sensitive [WMBR89]. N-formyl [NJAS89]. N-formylated [PZBD87]. N-linked [DBK87, KKSK86, YBBS87]. N-myc [MAE88, HSK^{+}89b]. N-ras [GZO88]. N-terminal [BH86b]. N2 [LHM89]. N2-dimethylguanosine-specific [LHM89]. Na/H [Owe86]. NADPH [MYT87]. NADPH-cytochrome [MYT87]. Naegleria [FCL86, LRF88, MLSW86, SW87a]. nail [LOH^{+}86]. NAPA [CRLW86]. NAPA-73 [CRLW86]. Nascent [NB89b, ATNH86, BMR85, BHK86, CCG89, HVBB88, PRSGD89, WE87, WKBR87]. native [AWH^{+}85, CKR^{+}85b, FM86]. natural [HLB^{+}88]. nature [GFB^{+}89]. NC1 [TC86]. NCAM [KMR^{+}88, NTLR88, NMM^{+}86, RWS^{+}85, SSSA87]. NCAM-mediated [KMR^{+}88, RWS^{+}85]. near [GSS89, Nic89]. nearly [TT88a]. nebulin [WW88]. necessary [SSY^{+}86]. necrosis [DWL86, SC88, SGP88, TTLR89]. needed [SPH87]. negative [BVHF89, MH86b]. negatively [HVL87]. neighboring [MLC^{+}88]. neither [GRCG86]. nematode [AC86a, AE87, EBCO88, RPW86, SRR89, Str86]. nematodes [BOW86]. neomycin [CHG87]. Neonatal [EPJBBW86, BFKF86, CB86, TYP^{+}87, TFS^{+}89]. neoplastic [DRME89]. nephritis [HKH^{+}88]. Nerve [ALCG88, And86, CW87, CWH86b, DFSK85, EWAP85, FGE86, LR85a, VHO^{+}89, Bis86, BD86a, BD86b, BAG86, BD89, BM88, BSG85, CCK86, DLG^{+}86a, DFdN^{+}89, DSTJ85, DKKK88, EPJBBW86, FCMK^{+}86, GRCG86, GP86, GDC^{+}86, Has88, HJS^{+}85, HKBT87, Hir86, HC86, HP85b, JS86a, LWM85, LDKG87, LFS^{+}89, LGC^{+}88, MSD^{+}88, MS86b, MS88a, MIHS86, NDJ^{+}89, OH88a, RLJ86, RSL^{+}89, SR86a, SGB^{+}86, TJ85, TS86, Tra88, VRLG89]. Nerve-induced [And86]. nerve-muscle [EPJBBW86]. nerves [BD86a, BD86b, RBAF89]. nervous [BFR85, CRLW86, DMD^{+}86, EBGS86, Fal85, GDD88, GABT86, HAK^{+}88, MS88c, PL85a, PCS88, Ran88, RLP^{+}87, SS88a, vdPdPR86]. nestling [CE85a]. net [HKF^{+}87]. network [CdCP^{+}89, KSL^{+}87, KEG88, MFA89, NL86, RSF^{+}85, SKAW86, TTF87, TDR^{+}87b, YR87, vDSP^{+}88]. networks [JGYG85, VH88, dSdSH^{+}88]. neu [SHL^{+}89]. Neural [AR88, KMR^{+}88, SPS89b, BF85, BFL88, BK85, CRR87b, CSG85, CG86a, CLC^{+}86, HNR87, HR87, HNNT88, HFG87, HFC^{+}86, KFS87, KJF^{+}88, LMSM89, MS86b, MS88a, MTK^{+}87, MHP^{+}86, MOP^{+}86, NAD^{+}85, PJ87, PS87, PRB^{+}89, PSGS85, PSK^{+}87, RZG86, RMS86, SL85c, SL85b, SLE86, WGA85, vdPdPR86]. neuraminidase [NRL89a]. Neurite [CCS87, HAK^{+}88, WD86b, WMSB^{+}89, BAG86, BM88, BSG85, CS88b, CMP86, DFSK85, ETT88, EDD^{+}86, Has88, HJS^{+}85, JS86a, JS86b, LFR85, LL89a, MEVM89, NTLR88, RWF^{+}87, RMP^{+}88, RLJ86, RRF87, RSL^{+}89]. neurite-myotube [RRF87]. neurite-promoting [EDD^{+}86, MEVM89, RMP^{+}88]. neurites [BH86a, CRR87b, DJS^{+}88, Fal85, JCBH85, JC89]. neuritic [JS86a]. neuroblastoma [CPHM86, DNSMA88, GK85, MER^{+}86, MLC^{+}88, RMP^{+}88, SDDZ^{+}89]. neuroectoderm [KMR^{+}88]. neuroendocrine [NJD^{+}86]. neurofilament [BD85, FGA86, HTGP85, LC85, MC89a, NL86]. neurofilaments [AST89, LN88, PBBP85]. neurohypophysis [NDJ^{+}89]. neuromuscular [DCPR^{+}89, GAH86, RGE85, SCLG86, SS89c, VJF^{+}88, WTB87]. Neuron [GE88, APS87, ELKH88, FGE86, LB89, PSK^{+}87, RBG85, REB^{+}86, RF87, SS88a, SPS89b, TDGE85, PP88b, REB^{+}86]. neuron-astroglial [ELKH88]. Neuron-glia [GE88, APS87, FGE86, TDGE85]. neuron-neuron [REB^{+}86]. neuron-oligodendrocyte [PSK^{+}87]. Neuron/ [PP88b]. Neuronal [DBD^{+}89, Hat85, Hat87, RDPR^{+}86, APS87, BOW85, BLR88, BSB^{+}86, DLG^{+}86a, ELKH88, FS88b, FGA86, GCR^{+}89, HD88a, HP85b, JLB86, JFAT89, LSB89, LMK88, MSD^{+}88, MND^{+}87, OHH^{+}89, RBG85, SLB87, SK85, SH87b, SV89, TRB86, TDR87a, TDR88, WESP85, WMSB^{+}89]. neurons [AR88, AELE88, AST89, BBB89a, CMNB88, CS88c, CNHF85, DBD^{+}89, DDBHL86, EWAP85, GB89b, GE88, HH86, HAK^{+}88, JPH85, KB86, LN88, MSM88, MOP^{+}85, NJD^{+}86, NL86, PBBP85, PBB86, RMP^{+}88, Rog87b, SSR89, SS85, SPS89a, SMA85, SEG^{+}86, VS89, vdPdPR86]. neuropathy [KP85]. neurosecretory [KSS86]. neurotoxins [BD86a, BD86b]. neurotrophic [BNF^{+}87, SSR89]. neutrophil [BKN^{+}85, BBB88, CMB^{+}86, GGA^{+}89, JBTA88, JGAR89, MCC86, SHGD85, SWZ85]. neutrophils [CC88, DWO^{+}87, GMCH87, GLW^{+}89, HO85, JTBA89, LMW^{+}86, NSF^{+}89, NJAS89, NL88, OS88b, PZBD87, RBKK85, RBKK87, RNW89, SLB^{+}89, SOP85, TN89, YSW^{+}85]. newborn [Abr85, AP86, Qua85a]. Newly [FGW85, RH87, AC87, BPJBF85, BSLS^{+}89, EMOB85, GPSM85, NL86, PLA86, PF85a, RT88, TLP^{+}87]. newt [DMH^{+}85, EMG86, RG87, WE89]. NF [MC89a]. Ng [HFC^{+}86]. NGF [DMW88]. NH2 [ES86, HGL88, PTB^{+}85, SSK89]. NH2-terminal [PTB^{+}85, SSK89]. Nicotinamide [ACPFB^{+}85]. nicotinic [CAG89, FC89a, HB88c, JLB86, MMS89]. nidogen [STE88]. nidulans [MGWM85, May89, MKH88, OR85, OMM87, WMGM85]. Niemann [LRF89]. NIH [BLSK88, BSK89, LPR^{+}89, MQ89]. NIH-3T3 [BSK89]. NIH-3T3-transfected [BLSK88]. NILE [SSC86, GMF85]. nimA [OMM87]. nitrate [KSFL89]. nitrocellulose [CV87a]. nm [BMDN87, BPE^{+}86, GKBG87]. no [FBP89, KFH^{+}86]. Nocodazole [EKH89, SRW88, JS86b]. Nodal [GSTS86]. node [BNL89, JBdlTB87, RDPR^{+}86]. nodes [FCMK^{+}86, GR89, SRB88]. nodule [PGH^{+}88]. Non [MHW85, BOW85, JMBM85, NFF86, SE85, TRB86]. non-myelin-forming [JMBM85]. non-neuronal [BOW85, TRB86]. non-nuclear [NFF86]. Non-spindle [MHW85]. non-uniform [SE85]. noncentrosomal [BKK87]. nonchromatin [FKP86, SSW^{+}85]. noncollagenous [TC86]. nonerythroid [CML85, MNWL85]. nonglial [Fal85]. nonglycosylated [GKS87a, HL88]. Nonhistone [BSY86]. nonjunctional [Rog87a]. nonlysosomal [ALL89]. nonmethylotrophic [DVVT88]. nonmuscle [BL85, HBPK87, HZS88, HLL89, HD85, LFC^{+}88, LHL88, MFM89, OKLB86, PSK^{+}89, TGNG87, WRL86, YMM88]. Nonneuronal [BNF^{+}87, GDvdM87, HKBT87]. nonpermissive [PSS89, RFMR86]. nonpolar [SRM89]. nonpolarized [RIS87]. nonproliferating [Wan85a, Wan85c]. nonproteinous [MOU88]. Nonrandom [Kub87]. nonrepetitive [FONW88]. nonsecreting [BHK86]. nonself [WD86a]. nontyrosinated [GB86b]. nonuniformly [NL86]. nor [GRCG86]. Norepinephrine [BSd^{+}88]. Normal [BPB^{+}88, CDTS85, YPG86, ABG^{+}89, BSSMF88, BCC^{+}87, BNS85, CSH89, CR87, CSR^{+}86, DCPR^{+}89, FGO^{+}85, FG89, Gor85, GLC88, HW85b, KKT^{+}86, KF89b, KDFG89, MNM87, MKH88, MTA86, PBBP85, RMM87, SGBF87, SGM^{+}89, SCFP86, SKDF88, WMH^{+}88]. normally [GPSC86]. Northern [HKPS89]. notch [JFAT89]. nothing [Mac89a]. Novel [GZ89, MB85b, TSOC89, APS87, BKA^{+}89, BS88a, BFA88, BGH87b, CR88a, CFC89, CAG89, CSK^{+}88, CRLW86, ELKH88, FMG89, GHW^{+}88, GD89, GSF^{+}88, GGA^{+}89, HC86, HABM85, KC85, KR86, KFS87, KYS^{+}88, LKF87, NSY86, NM89, NT86, OHH^{+}89, PNB89, PRSGD89, RE85, Rog87b, RMS86, SROB86, SR87, VA85, VSSRB87, WVLC86]. novo [FMJS89, FBE85]. nu1 [KCN^{+}89]. nuc2 [HHY88]. Nuclear [HSS^{+}87, JS88a, MKH88, RG88, SJ86, Ake89, AG89, AB89, BK86, BF86, BMB87, BS86, CM88b, CA86, DRD^{+}88, DLF88, DF88b, Ear87, ER87, FLM86, FDG88, FSM88, FNPF87, FA85, GB87a, GB87b, GLD^{+}88, HKKP86, HSK88, HHY88, HS87a, JT85, KSFL89, KWH89, KSKW^{+}85, LW86, LPR^{+}89, LSEN87, LM87, LT89, LHM89, LM85a, LAC^{+}85, LL89b, MNC86, MLPF89, MBLS87, NFF86, OOCT89, PLA86, PRSGD89, RP89, SSH87, SHH85, SJ87, SG88a, SSO89, SNC^{+}86, SMR86, SSW^{+}85, SGBF87, SF89, SSG87, SALN88, SG86, TS86, TLP^{+}87, TB88a, WKKS^{+}87, WM88, WN88, WMC86, WBB89, WSRT89, ZDH88]. nuclear-encoded [CA86]. Nucleated [ASM^{+}87, EMK85, Fra87, OTS86]. nucleating [BH86a, CWZ86, CBL85]. nucleation [BKK87, GHL^{+}89, HWDC89, MK85a, MTS89]. nuclei [BHS88, CM86c, CM87a, FC86, GK86, HMG^{+}86, HS87a, HS87b, KdS86, Kub87, OR85, SJ87, SMS^{+}88, WDR89]. nucleocytoplasmic [FDG88, JS86d, LSP86, SJ86]. nucleolar [AB88, BRR^{+}87, MM89, Ree85]. nucleoli [WS85]. nucleolus [HKTY89, HHSF85]. nucleoplasmin [DRD^{+}88, MABJ^{+}86]. nucleosomal [CRC^{+}87]. nucleosome [POW^{+}87]. nucleotide [CS89, FTLSS88, HBPK87, JBTA88, Miy88, PZBD87, PAHC89, RH86b, TN89, TSS89, TL87a]. nucleotide-binding [Miy88, PZBD87]. nucleotide-induced [TN89]. nucleotide-sensitive [TSS89]. nucleotides [HVBB88, HCG87, VB85b]. Nucleus [WAG^{+}89, AY89, BSY86, BBB89b, FKP86, GD89, MNC86, NB89a, OON89, RH89, SCK^{+}89, TDR^{+}87b, WKKS^{+}87, WSJ85, WBDR88]. nucleus-basal [WSJ85]. Nucleus-specific [WAG^{+}89]. nud [MKH88]. numatrin [FSM88]. number [AWJ85, DAJJ^{+}88, DLF88, KSSP86, MSR^{+}86, MQ89, NG85, RF87, SR85a, TT88a, ZKR^{+}88]. numbers [DM87].

O [CCL^{+}86, CMW87, SPO85, HSS^{+}87, KKSK86]. O-linked [HSS^{+}87, KKSK86]. O2 [SDS87]. O2-evolving [SDS87]. oblique [DCB89]. observation [Bäc88, KMK86, RGM85]. observations [CPS88, DMRB87, GB86a, GB89b, Hep89, UP86]. observed [PP88b, SP86]. obtained [MRWJ88]. occluded [HM89a]. occludens [SSMG86]. Occluding [MD85, GS86b, MBC86]. occupancy [RG86b, SOP85]. occur [CC89, KS85c, OW86, PB87, VC85, WSW85, WB86]. occurrence [BKK87]. occurs [CCC86, CLOH87, DBM^{+}89, DMGS88, LR85b, MSW^{+}86, MTS89, MB87, ORA^{+}86, SGS87, TI87, vZC87]. ochre [WVCR89]. OKT9 [WKRH86]. old [AP86, HB88b, SSSP85]. oligodendrocyte [HRBR89, MRO^{+}86, PSS89, PSK^{+}87]. oligodendrocyte-oligodendrocyte [PSK^{+}87]. oligodendrocyte-secreted [PSS89]. oligodendrocytes [CR88a, CR88b, DDBHL86, HFB^{+}89, MS88c, TACQ89]. oligomeric [BSDH86]. oligomerization [BDWH88]. oligomers [MHWB86a]. oligosaccharide [CMR^{+}87, FCCC^{+}89, LBR86]. oligosaccharides [GB85a, GP86, HGS89, MB88, REB^{+}86, SBG89]. Oligospecificity [ABME88]. omega [BGT^{+}89, FBP89]. oncogene [HSK^{+}89b, KSSAQC89, MQ89, MAE88]. oncogenes [BS86, DMG^{+}89, GZO88]. oncogenic [RLB^{+}89]. one [DS88, ECZ86, MPB88, PHL89, SBG^{+}85, Toy87a, WPBB^{+}85, WLW^{+}88]. one- [Toy87a]. only [CF88a, DDB^{+}88, FS86a, GW88a]. onset [FOW89, OON89, RS88, TB89b]. onto [CSL87, SL88, SB85e, TSBW86]. ontogeny [GDF86]. oocyte [ASZ^{+}88, CM87a, MABJ^{+}86]. oocytes [BC89, CC89, CM86c, CJH85, CS88d, DF86a, DM87, DMC85, DF88b, FWP^{+}87, GK87a, HW85a, MHW85, PJFB89, SP85, SPL87, SCP85, SHR86, WM85]. oogenesis [AKVF88, SF89]. operated [EZBW88]. opiomelanocortin [THFB87]. opposing [KSJF87, SMR87]. opsin [VFB^{+}89]. opsonization [CC88]. opsonizing [JGAR89]. optic [CR85, FCMK^{+}86, RBAF89]. optical [GW86, KLS85, MASA89]. optional [BN87]. order [AY89]. Ordered [SH87a, NB88a]. organ [CCK86, CFC89, MMM^{+}85, NZ85, NSM^{+}87, Qua85b, Wal86]. organ-specific [NZ85]. Organelle [BRKE88, BKR86, EWHP87, HHS^{+}85, KS85c, KS86, RPW86, SSRS88]. organelles [CR85, GSS87, HDP^{+}85, HB87, KBR87, KS88, LB89, LMWG85, NNJ^{+}87, SGBG87, SLHvF85, SAA89, SKU^{+}85, WS85]. Organic [GTG85, BBW86, BKC^{+}86, SNSS87]. organism [DB85, DRRB88]. organisms [PF85b]. Organization [BAD86, BFPS85, DD88, DF88a, KA86, LB89, MP88b, AYCY89, BAD^{+}89, BSA87, BTV^{+}89, BD89, DCL^{+}87, DCC^{+}88, DWS88, FA85, FS88b, FW85, FONW88, GG87b, HBC89, HMG^{+}86, HS87a, HS87b, JS86a, JS86c, KCAF89, KP85, Kol86, KMY88, LMD^{+}86, LWT85, LAC^{+}85, MPT88, NV86, PN88a, PN88b, SYH^{+}85, SYH^{+}86, SWTB88, TRM88, TT89b, WFW87, ZHES89]. organizational [JTBA89, JGYG85, Maj87]. organize [SR85b]. organized [EMOB85, RLB^{+}89]. organizing [DMD^{+}86, GBBS89, MHW85, NSM^{+}87, TMB85, TPBB85]. organogenesis [MAE88, NT86]. organs [KYS^{+}88]. orientation [BBB89a, HFB^{+}89, MVK^{+}88, PHL89, SAB^{+}89]. orientations [STE88]. Origin [Abr85, SABVK89, HH86, QCKR85]. originates [SROB86]. ornithine [HFF^{+}85, HKF^{+}87, NAS^{+}87, SHL^{+}89, SPO85, SHK^{+}87]. Orthogonal [Gor85]. Oryzias [Sch85]. oscillates [KM87]. oscillation [Bro87]. oscillations [GW86]. Oscillatory [RDJ^{+}86, ASLFP89, BPC89]. osmium [BFPS85]. Osmotic [GCGR85, LHNF87]. ossification [PP89]. osteoblast [NV89, NR89b, RYFM89]. osteoblast-derived [RYFM89]. osteoblast-like [NV89, NR89b]. Osteoblasts [RYF^{+}87, GCK^{+}88]. osteoclast [BNLC85, BNB^{+}88, DWT^{+}89, OBCO85]. osteoclast-specific [OBCO85]. osteoclasts [BKC^{+}86]. osteogenic [MTCE87]. osteonectin [HHMH87, NWE^{+}88]. osteopontin [NR89b, NWE^{+}88]. osteosarcoma [DAS^{+}87]. other [APF^{+}88, BKPG89, BWL^{+}86, BBH87, CS88a, KDHS87, PRS^{+}88, SSRS88, TH89a, YSW^{+}85]. ouabain [HTTFC85]. Outer [UMHI85, DJBD85, Eck87, FRH85, HG88, Kam88, KW88, MR85b, MM87, OK86, SGH85]. outgrowth [BLR88, BAG86, BM88, BSG85, CCS87, DBD^{+}89, DFSK85, ETT88, Fal85, Has88, HJS^{+}85, LFR85, LL89a, NTLR88, RLJ86, SCD88, TRB86, TDR87a, WD86b]. ovarian [COMS87]. ovaries [SHBV89]. ovary [BHS88, CCC89, CSD89, CDTS85, FBPS87, GWCW85, HHR^{+}89, LEW^{+}88, MGM87, RFMR86, SC86, SFK^{+}89, SBD^{+}89, SK88, TASK86, YM87a, YM87b]. Overexpression [BLSK88, CSH89]. overlap [LD88]. overproduce [DAS^{+}87]. oversecretes [EFR^{+}89]. oversecretion [RWE^{+}88]. ovo [LL89b]. ovoperoxidase [WKS85]. oxidant [OS88b]. oxidase [DVVT88, FBCQPR87, ISL87, KVB^{+}86, SL87b]. oxide [KWW85]. oxide-induced [KWW85]. oxyntic [KDB87]. Oxytricha [GG87a].

P [DFSK87, MYT87, YMT85]. P-450 [DFSK87, MYT87, YMT85]. P0 [Tra88]. p107 [SSW^{+}85]. P16 [PV85]. p28 [SSW^{+}85]. p30 [RMP^{+}88]. p36 [CHMB86]. p38 [BFK87, NJD^{+}86, VJF^{+}88]. p42 [KC88]. P450 [MVK^{+}88]. P450IIC2 [SSK89]. p60c [GH88]. p60c-src [GH88]. p62 [RG86a]. p81 [GCBH86]. P815 [AHRM85, AM86]. packaged [RWP^{+}89, THFB87]. packaging [SK85]. packed [FZ85]. Packing [BB85, SQM89]. PAI [LS87a]. PAI-1 [LS87a]. Pair [HF85a, GH85b, HCE88, HMN87]. pair/projection [GH85b]. paired [RPCB88, WCSR85]. pairing [LNH86]. pairs [ESA86, SGM^{+}86, WBDR88]. Palade [EWHP87]. pale [BNS85]. pancreas [AC87, CJ86, HBL^{+}88, SPF87]. pancreatic [LMWG85, MBKO86, NRTE89, RH87, TKFH89]. paper [CV87a]. paracellular [MBC86]. paracrine [SC88]. paracrystalline [HLW^{+}89]. paracrystals [SQM89]. paradox [GC87]. paraflagellar [SDS89]. parafollicular [BGN^{+}88]. Parallel [CCBKJ89, PG86, WW88]. Paramecium [BN88, GHL^{+}89, MLK^{+}87, RdLB87, STGK87, TMH^{+}85, VSKSP86, ZP85]. parameter [DBE^{+}87]. paramyosin [AE87, EMOB85]. parasite [CC85b, HdAC89, HLU^{+}87, TP89, WS89]. parathyroid [NR89b]. parietal [GW87, MRD^{+}89]. parotid [AC86b]. part [FRG^{+}87, KJW89, NFN87, SW86, YBBS87]. Partial [AFSL85, GCC^{+}89, JBG^{+}87, SB85d, KDIG89, MWS85, RHD^{+}85, UK86, WS87]. partially [CCG89, FCH^{+}88]. participate [CS87]. Participation [SBHN86, CG86b]. particle [ALWL89, HPW89, Ibr87b, KR86, LD86, SDSO86, TLRW86, VA85, WSV^{+}89, WKBR87, WW89]. particle-dependent [LD86]. particles [CS88d, DPM88, DF88b, JIFH85, LAC^{+}85, POW^{+}87, SDS87, Toy87a]. particular [SGBG87, WMGM85]. partition [TB85b]. parvalbumin [KSH87]. pass [LP85a]. patches [Wal86]. pathogenic [TRM88]. pathway [BNB^{+}88, CVWZ89, CGV^{+}86, DCH86, GBM89, HCH85, KST^{+}89, LMSM89, LBW89, MDM88, MDJ89, NFN87, PMR89, RH87, RHL^{+}85, RDSG86, SDPZ89, SBHN86, SDZB^{+}85, TLPTV85, Toy87b]. Pathways [MCD85, BFSH87, BNS85, FKC87, GSS^{+}85, KST^{+}89, KSSAQC89, LHvF^{+}87, MK85c, PPB^{+}89, PSR^{+}88, RM89b, SM89a, SMW89, SGGS88, SGG^{+}89, WBW89, vZC87]. patients [AMG85, DCWJ89]. pattern [BCKJW88, ETA85, FCVDH89, TTC88]. patterns [BGT^{+}89, HCE88, NT86]. PB [DFSK87]. PB-4 [DFSK87]. PC [DJS^{+}88, JCBH85, MBLK88]. PC-12 [DJS^{+}88, MBLK88]. PC12 [SK85, TDR87a, TDR88, BM88, DMW88, DFSK85, DKKK88, FGA86, FGE86, GRCG86, HJS^{+}85, JS86a, JS86b, LR85a, LDKG87, NGS85, OH88a, RLJ86, RKA^{+}89, TS86, VHO^{+}89, VAD^{+}85]. PC12h [Has88]. PDGF [SHS^{+}88, HRBR89, SHS^{+}88]. PDGF-like [SHS^{+}88]. peanut [MS88c]. pectoral [CB86, GCPE85, TB89a]. pelletability [MP86c]. pellucida [PHM85]. pemphigus [JYG86]. Penetration [FGF88, OMY^{+}85]. pentasaccharide [CCL^{+}86]. peptide [BSBZ88, BDS88, BNF^{+}87, CSK^{+}88, DAS^{+}87, FKC87, GORW88, HSWH86, HKF^{+}87, HO85, JFP89, KRN^{+}86, LC87, NJAS89, PZBD87, SWZ85, SW88, SHK^{+}87, WBGV85, WW87b]. peptide-induced [HO85]. Peptides [GB85b, BKPG89, CR88b, CWZ86, CWH^{+}86a, FLBP^{+}86, GAPR88, GABT86, GY86, HPR85, KSS86, MSLR^{+}86, RHL^{+}85, SIYT86, ZSSR86]. periaxonal [TACQ89]. Pericellular [CC88, HvdSdC^{+}89]. pericentral [BPMD87]. pericentriolar [CBL85, HPM87]. perichromosomal [WKKS^{+}87]. pericyte [KDHS87]. pericytes [DHH89, HD85, JHP85a, JHP85b, OD87, SR89]. perinatal [FL89, PBT^{+}86]. Perineurial [JPU89]. Perinuclear [MBSMD86, LKF87]. period [PBT^{+}86]. Periodic [KCAF89, AC86a, BPC89, BPE^{+}86, Miy88, VMH^{+}88, YMM88]. periodicities [BMDN87]. peripheral [BNL89, CE89, Hir86, HAK^{+}88, LFS^{+}89, MB85b, MIHS86, RLP^{+}87, SS88a, SRB88, Tra88, WBVL86]. periphery [AY89, HS86a, Pri87]. periplasm [BIES89]. periplasmic [dHTG88]. peritoneal [OCP86, PUW^{+}88, WG87a]. peritubular [STF85]. permanent [FGA86]. permanently [OH88b]. permeability [AR85, GS85, MBC86, MCC86]. Permeabilization [MA85]. permeabilized [BFG86, HCG87, WDR89]. permit [WBW87]. peroxidase [AK87, KBPW89, SYS85, TAMR85]. peroxidase-diaminobenzidine [AK87]. peroxide [PBK88]. peroxisomal [AMG85, BVHF89, DCWJ89, GKS87b, GKS88, YF87]. peroxisome [TB85a]. peroxisomes [AFSL85, AMG85, BVHF89, DVVT88, GKH^{+}89, ISL87, KPK86, KSC^{+}89, SL87b]. Persistence [IGSGD87]. perturbants [WLGB85]. Perturbation [dGRS87, MM87]. perturbed [CDTS85]. perturbs [BFL88]. pertussis [BKN^{+}85, BBB88, KST^{+}89, KSSAQC89]. Pf [HUA^{+}86, FJR87, WVCR89, HLU^{+}87]. pf-14 [WVCR89]. pf/pf [FJR87]. PH [CPM86, CMK87, PHM85, AC86b, BNLC85, BFRT87, CH85, CdCP^{+}89, FWF88, GW86, Heu89a, JT85, KH85b, KSG85, MSLR^{+}86, NYG87, OTS86, PW85, SNCY87, WMM86, WW87b, vAAA86]. PH-20 [CPM86, CMK87, PHM85]. pH-dependent [SNCY87]. pH-induced [KH85b]. pH-regulated [OTS86]. pH-sensitive [CH85]. pH-triggered [PW85]. phagocytes [JNA85]. phagocytic [WDV87]. Phagocytosis [ICY^{+}88, DMGS88, GGA^{+}89, KM87, MH86a, MSW^{+}86, TSBW86]. phagosomes [JGAR89, SKAW86, WG87a]. phalloidin [Coo87]. Phase [FJMW87, BHSJ86, CLOH87, DWL86, DZ85, Gla85a, Gla85b, GL85, LTG85, PCBD89]. phase-promoting [PCBD89]. phases [PT88]. Phasic [AC87]. phenobarbital [DFSK87]. phenobarbital-inducible [DFSK87]. phenomenon [CVWZ89]. phenotype [BBP88, DMG^{+}89, GFS^{+}89, MB85a, SHS^{+}88]. phenotype- [SHS^{+}88]. phenotypes [GFHDD89]. Phenotypic [MPT88, BB88b, HBF^{+}88]. Phenylarsine [KWW85]. pheochromocytoma [VHO^{+}89]. pHi [CDL^{+}85]. Phorbol [BSB^{+}86, DJ89, GKS87a, MDM88, SPO85, SYS85, CVWZ89, FTC87, GCGR85, LPR^{+}89, LEG^{+}89, LMF^{+}86, MMT85, MW86, PKN^{+}89, RBKK85, RBKK87, RRP88, SVG^{+}86, Wal88]. phosphatase [CMTC85, Eva89, GDS87, MLK^{+}87, NYRK87, TKP^{+}88, dBBB^{+}86]. phosphatases [PCBD89]. phosphate [BNB^{+}88, BGHvF87, BGF86, DK88, GF86a, GF86b, GSS^{+}88, MGK89, NB88b, Pfe87, SGG^{+}89, SE86]. phosphate/insulin [EG89]. phosphatidylcholine [CKR^{+}85b, DJBD85, Gla85b, KS85a]. phosphatidylinositol [CE89, GCR^{+}89, GWI^{+}87, HFG87, KSM89, MS88c, NYRK87]. phosphatidylinositol-anchored [GCR^{+}89, KSM89, NYRK87]. phosphatidylinositol-linked [HFG87, MS88c]. phosphatidylinositol-specific [GWI^{+}87]. phosphatidylserine [Gla85b]. phosphodiesterase [CHG87]. phosphoinositide [VAD^{+}85]. phospholamban [FYRK88, JJ87]. phospholipase [BSSMF88, CKR^{+}85b, GWI^{+}87]. phospholipid [ACN^{+}89, BGBB89, EMT88, Gla85a, Gla85b, GTP87, GMCH87, HdAC89, HG88, SCGS85]. phospholipid- [GTP87]. phospholipid-dependent [ACN^{+}89, CWH86b]. phospholipids [BBJG89, DH89a]. Phosphomannosyl [YSR87, FCK87]. Phosphomannosyl-derivatized [YSR87]. phosphoprotein [JLH^{+}86, MLK^{+}87, SL85d, ZP85]. phosphoproteins [BS86].

Phosphorylase [SWZ85, MAK^{+}86]. phosphorylated [CWOK87, GF86a, LN88, MP88a]. Phosphorylation [CMB^{+}87, Eva89, SS85, SL85d, WMV86, AG87, ALCG88, BBBM89, BCKJW88, Bre89, Bro87, BSB^{+}86, BSG85, CHMB86, CGA89, CPKJ87, DG89, DNSMA88, DH89b, FWP^{+}87, FBF^{+}89, GK85, GDC^{+}86, GDS87, GCGR85, HLW^{+}89, Has88, KC88, KB86, KWS^{+}89, KTP87, LFFW89, LR85a, MP89, MPBZBE89, MDM88, NYG87, NGS85, PFRS89, PM87b, RRP88, RSR^{+}88, SNNK85, SSS85, SDNWA87, SP89, THM86, TL87a, Try89, VMT^{+}88, WG86, YBBS87]. Phosphorylation-dependent [SS85]. Phosphotyrosine [TS88, BJMK89]. phosphotyrosine-containing [BJMK89]. Phosphotyrosine-modified [TS88]. photoactivation [Mit89]. photobleaching [BBP^{+}85, JS86d, WS86a]. photocross [KJW89]. photocross-linking [KJW89]. photogenes [RB85]. photoinhibition [SDSO86]. photolytic [MTG88]. photophobic [SL85d]. photoreactive [CGV^{+}86]. photoreceptor [HFB87, SH87a, TTM88]. photoreceptors [MEA89, PF87, VFB^{+}89]. photoregulation [RB85]. photosynthesis [dVOD^{+}89]. photosynthetic [HM89a, JM86, SDS87]. photosystem [DSA86, HMP86, JHPS86, MMR88, SDSO86, SCGS85, dVOD^{+}89]. phototaxis [HW85b]. phycoerythrin [LG89]. Physarum [ASZ^{+}88, CL85a, CL86, DD87, GL85, IKH87, SGBG87]. physical [VB88a]. Physiological [CM86a, CR87, RSR^{+}88, WS86b]. physiologically [MG86a]. phytochrome [MP86c]. phytohemagglutinin [TC87]. Pick [LRF89]. pig [CPM86, CMK87, KTF87a, PHM85]. pigment [MH86a, MP86d, MW88a, RH86b, TB88b]. pigmentation [JVG^{+}85]. pigmented [OE85, PN85]. pili [Smi87]. pinocytosed [SYS85]. pinocytosis [GTN86, SYS85, SBS87b]. pit [HA89, HM89c, MDVG^{+}85]. pits [DM87, LDA86, MMBA89, RSF86, SOPvD87, SOB^{+}89, SPLW89, Too85]. pituitary [FZ85, HFZM87, KNGH^{+}85, MK85c, PS86, RRB86, RFZH85, SS88c, WTBBW85]. placed [SR86a]. Placement [FR86]. placenta [GTP87, MBM85]. placental [EB87, FCZ^{+}89, SYT^{+}89, TH89c]. planar [LDA86]. plane [HGL85]. plant [BMB86, BEHJ^{+}88, CBL85, DHL85, DVHC89, LA87, MAG88, MWS85, SMC89, SL87a, SFW87, TC87, VLD^{+}85]. plants [PKS^{+}89, Ree85]. plaque [BMBL89, CKF85, DF88a, HTT89, KOF88, Rog87a, RPCB88, UMHI85]. plaques [DF86b, LHS^{+}89]. Plasma [MBK89, ACPFB^{+}85, AHRM85, AM86, AYG85, BBH85, BFSH87, BHSJ86, BGL^{+}88, BPKS86, BWSB86, Cha85, CJY86, CMK87, CJ85, DCQB88, DRRB88, FTLSS88, GWM85, GB87b, GCT86, GML^{+}89, GZGM87, GPSM85, HT86, HHES88, HBB85, HFB^{+}89, HAK^{+}86, IHJ88, ITY^{+}88, JBTA88, JTBA89, KGRP85, KS85a, KS85b, KBB^{+}88, KCR86, KP89, LP85a, MT86, MSQ^{+}89, Mat86, MVSW86, MAG88, MM87, MMM^{+}85, NRTE89, PKS^{+}89, PHN^{+}88, PMR86, RIPS85, Rog87a, ST86, SVSH^{+}88, SCLG86, SYSP88, SMC89, SL86a, SL88, SHP^{+}87, SMS^{+}85, TSBW86, TLP^{+}87, TSOC89, VSSRB87, VSSRB88, WGPW89, WGPH^{+}89, WO86, WHI86, WSM86, WHL^{+}86, WE86b, WL87, YE87, dCFB86]. plasmalemma [SSSP85]. Plasmalemmal [MG86b, GS85, GFSS86, PCP85, SKU^{+}85]. plasmid [BWL^{+}86]. plasmin [SR89]. Plasminogen [FHS^{+}86, HBB^{+}87, SHBV89, BVD87, CSM^{+}89, COMS87, GKS87a, HB88a, KOBLM88, KNGH^{+}85, LSAKO86, LS87a, MVS88, PVMO87, PFPM86, PSS^{+}87, PHN^{+}88, PMR89, SMR87, SR88, SPT^{+}89, SVG^{+}86, VBB85, VS89]. plasmodial [IKH87]. plasmodium [IKH87, HdAC89, HUA^{+}86, HLU^{+}87, KSG85, SYSP88, WS89]. plastic [KBB^{+}88, Ris85]. plastic-embedded [Ris85]. plastid [RB85]. plastocyanin [HRR89]. plate [DT89, TTS^{+}89]. Platelet [CR86, HKM^{+}87, HP85a, MCB85, PML^{+}89, RSG87, AS86, BC85, CTS^{+}87, Col86, DHB86, DBC^{+}87, GDO^{+}85, HTRBP86, HCHK89, IMP^{+}87, KC88, LJC^{+}87, NKMK86, PPG85, PUW^{+}88, RHD^{+}85, SHGD85, SHS^{+}88, SMS^{+}85, TH89a, TRT^{+}88, WCPK88, ZR87]. platelet-activating [PUW^{+}88]. Platelet-derived [HP85a, MCB85, PML^{+}89, HKM^{+}87, HTRBP86, KC88, SHGD85, SHS^{+}88, TRT^{+}88, ZR87]. platelets [AS86, BMBL89, BFPS85, FBB^{+}88, Fra87, GLW^{+}89, HCS89, NYG87, NH87, OPN^{+}85, PON^{+}88, SRW^{+}89, TAHC^{+}89, WGC85, dGRS87]. plates [NT85]. played [SW86]. plectin [FLH^{+}88]. pleiotropic [KKSK86]. plus [FS86a, GK87b]. plus-end [GK87b]. PMA [MS88b, dGRS87]. podocalyxin [HHD^{+}86]. point [CL86, Ris85]. point-dried [Ris85]. polar [RG86a, Smi87, WSW85]. polarities [MTS89]. polarity [BBB89a, BPC89, FS86b, GJD^{+}88, GB89b, HPM87, HSMCL87, HK88b, HM89c, KCN^{+}86, MEA89, MSQ^{+}89, MP86d, MG86b, NFW89, NV86, NH89, PCK^{+}87, SMVS^{+}86, SVSH^{+}88, SMW89, TB88b, VSSGRB87, ZMH89]. polarization [BR85, HM89b, TTSSH85]. Polarized [BNB^{+}88, CR85, GDRB89, RIPS85, ZR87, AFMD89, BAD^{+}89, BF88, CTJG88, CJH85, FGF88, GGR^{+}86, LCDRB89, LLSRB89, OS88a, PCK^{+}87, RIS87, RT88, RGPRB87, SRM89, UPL^{+}87]. pole [KBL^{+}86, KDB87, RDJ^{+}86, SROB86, VSSGRB87, WSW85]. poles [SR85a, SFD^{+}86b]. poleward [GSB87, Nic89]. Polewards [Mit89, SPH87]. pollen [KS88]. poly [MLSW86]. polyaspartate [CHH^{+}88]. polycephalum [ASZ^{+}88, CL85a, DD87, GL85]. polyclonal [LC87]. polyedra [NNJ^{+}87]. Polygons [LHS^{+}89]. polymer [GK85, RGBM87]. polymer-dependent [GK85]. polymerase [BRR^{+}87, HKTY89]. polymeric [BHM89a, DNM86, HCH85, MBH87, SHP85a, SHP85b]. Polymerization [LG88, SB85e, GGMH86, HWDC89, KTP87, MSA88, SOP85, SDPZ89, TN89]. polymerized [UW86]. polymers [CWZ89]. polymorpha [DVVT88]. polymorphic [ASCM89]. polymorphism [SQM89]. Polymorphonuclear [SDPZ89, CWZ86, DZ85, SSKK89]. polypeptide [BF86, CB88, DLC88, FS88a, FRG^{+}87, GS86b, HBPK87, HW86b, KC85, KWPH89, KRN^{+}86, LM87, MHP^{+}86, NAD^{+}85, SKK^{+}85, SSSA87, SHR86, SSMG86, TLRW86, UF86, VS88b, WRS^{+}86, YKCR88, YJL^{+}89]. polypeptides [BLSL88, CCC86, CCG89, CML85, FY86, HF85a, HMP86, PHL89, SNNK85, SLL88, SLG88, VB88a, ZHE85]. polyphosphoinositide [CHG87, YIJ88]. polyphosphoinositide-modulated [YIJ88]. polysaccharides [BRS87, RDB^{+}86]. polysomes [KMM88]. polytene [HS87a, LLAS85]. polytopic [AHM89]. pomatia [BOW85]. pombe [AY89, CW89, HPW89, HHY88]. ponticulin [WL87]. pool [Smi87]. pools [GHL85, KWW85]. poor [AdS85]. population [GHA88, MKH88, RE85]. populations [BMB87, BRP88, KSS86, SK87, SF87, WMH^{+}88]. pore [Ake89, AG89, AB89, FDG88, HSS^{+}87, SJ87, SSG87, WBB89]. pore-mediated [SJ87]. pores [DF88b, FNPF87]. portion [NMM^{+}86, SL87b, SQM89]. portions [CS86d]. position [Hym89, LTG85, NBS88b, RDJ^{+}86, WMSB^{+}89]. positional [HP87b]. positioning [ES85, WMC86]. positive [DVHC89, GF86a, HKF^{+}87, KW86b, MH86b]. positively [LG88]. possess [AC86b, BJMK89, GRCG86, SNSS87]. possessing [WC87]. Possible [GSS^{+}85, AF89, DRC^{+}86, GCGR85, HNH^{+}87, MLK^{+}87, PFRS89, SLB^{+}89, TH89a, UW86, VS85, Wan85d, WG86, WS89]. post [BGL^{+}88]. post-Golgi [BGL^{+}88]. postconfluent [FRR85]. Postendocytic [GF87]. Postendocytotic [BHM89a]. posterior [TKK^{+}87]. postfield [GBH87, TFBW85]. postganglionic [JB87]. postjunctional [SCLG86]. postmitotic [BK86, CH87, HDL^{+}86]. postnatal [PS87]. Postpolymerization [GKB87]. postreceptor [DJ89]. postsynaptic [BF87, BCPC87, CCLC89, FMT^{+}87, KCN^{+}89, MR85a, PF85a, UTK88]. Posttranscriptional [VN88, TF86]. Posttranslational [BAHE86, BHK86, BGF88, BDWH88, BMP^{+}86, BRP88, CDB87, HM89b, NL85, SABK87, SV89, dVOD^{+}89, BD85, BP87b, JHPS86, LR85b, NADS86, SS88c, TRC^{+}87]. posttranslationally [GKB89, HW88, HBKK87, WB86]. Potassium [LDA86, LSDP86]. Potassium-dependent [LDA86]. potent [Has88, SVG^{+}86]. potential [BB85, HFK^{+}89, ISL87, KSSAQC89, Kro86, MSS^{+}89, NZ85, PHTK87, RPW^{+}89, TSOC89, WS86c]. potentiated [MFB^{+}89]. pp60c [RE85]. pp60c-src [RE85]. pp60v [RE85]. pp60v-src [RE85]. Pre [BSLS^{+}89, ALL89, CBYK88, HW88, LW86, MR85a, NAS^{+}87]. pre- [MR85a]. Pre-Golgi [BSLS^{+}89, ALL89, CBYK88]. pre-invertase [HW88]. pre-messenger [LW86]. pre-ornithine [NAS^{+}87]. preassembled [CSL87]. precedes [Knu85]. preceding [LB89]. precore [GORW88]. precursor [BPL87, BNS85, BIES89, CHK^{+}86, DI87, DCB^{+}88, EVG^{+}86, FLBP^{+}86, GLD^{+}88, HFF^{+}85, HJV^{+}89, JFAT89, KBO^{+}88, LA87, OGJ^{+}89, PL86, PSR^{+}88, RGW^{+}89, RWE^{+}88, SIYT86, SLG88, SBH^{+}88, SHK^{+}87, VS88a, VS88b]. Precursors [VVL^{+}89, BCC^{+}87, BMP^{+}86, CBS^{+}89, DS87, EFR^{+}89, HP87a, JSCLP89, LHvF^{+}87, PW87, SLHvF85, VS88a]. predatory [WD86a]. predominantly [MBH87, RH87]. predominates [CG89]. Preferential [CCP88, Fal85, KdS86]. preferentially [AY89, AU89, CK86]. preganglionic [JB87]. pregnant [ZHE85]. preimplantation [HM89b, SA85]. Preliminary [TTS^{+}89]. Prelysosomal [PLS89, GF87, SSS88, TLPTV85]. premyofibril [TM87a]. Preparation [ZHE85, ISM85, WCPK88, EWAP85]. preparations [DG88]. preparative [KHKR86, MSK^{+}87]. Prepro [HW88, GW88a]. prepro-alpha-factor [GW88a]. Prepro-carboxypeptidase [HW88]. preproinsulin [ES86]. preprolactin [Ibr87a, WKBR87, WW89]. prepromelittin [Ibr87b]. preprophase [Pal87]. preprosomatostatin [SS88c, SS89d, NADS86]. preprotein [SM89b]. prerequisite [SW85]. Presence [CCLC89, AG87, BC85, CC88, CS86c, FS86a, GSF^{+}88, JHP85b, KG85b, MSQ^{+}89, MKH88, RPW^{+}89, ROCK86, SCBR85, TH89a, WKKS^{+}87, TM87a]. present [BFBL88, CS88c, CIK^{+}89, DN88, Fec87, GZ89, GHA88, KSBB87, NG85, RFZH85, SCP85, TDR^{+}87b, Wan85a]. presented [FRG^{+}87]. preservation [BFPS85]. prespore [GDF86]. prestalk [GDF86]. presumptive [KDIG89]. presynaptic [HSK^{+}89a, MMM^{+}85, NDJ^{+}89]. prevent [BM86, MSD^{+}88]. prevention [BHM^{+}89b]. prevents [BFPS85, JBG^{+}87, MBKR89]. Primary [LCBS89, PA88, WSS^{+}89, WBB89, BSB^{+}86, CNHF85, DCWJ89, GJD^{+}88, HZS88, HSS89, KH85a, KSC^{+}89, LCE87, PG87, SFS^{+}87, SHL87, SVG^{+}86, TH89a]. primed [VSKSP86]. principally [Rog87b]. printing [CV87a]. pro [SPT^{+}89, THFB87]. pro-opiomelanocortin [THFB87]. pro-urokinase [SPT^{+}89]. probable [Cha85]. probe [KT85, KF89b, LC85, LVSC86, SRT^{+}86]. probed [VA85]. probenecid [SNSS87]. probenecid-inhibitable [SNSS87]. probes [AYG85, BTR^{+}86, FMK^{+}86, GS85, HNP^{+}86, KL89b, RKWS86, SYH^{+}85, SKDF88, TTSSH85, UP86, WG87a]. Probing [AT86a, HLL89, LPTL86, MSA88, WKBR87]. procedure [SMH88]. proceeds [PP88a]. process [AG89, BLR88, BBB89b, CRJM88, DBD^{+}89, MAG88, TI85, TI87, TRB86, TDR87a, Wan85c]. processed [CHK^{+}86, KGRP85, KBPW89, PS86]. processes [DPM88, FCMK^{+}86, KS85c, NB88a, NB89b, RF87]. Processing [GB85a, GWG85, KA85, LA87, MDJ89, SWC^{+}87, BMR85, BDS88, BMP^{+}86, BNS85, CS88a, CA86, CMR^{+}87, DW89, DBC^{+}87, FJR87, FMG89, FLBP^{+}86, GF87, LW86, MGSG88, NAS^{+}87, NADS86, PHRM87, RWE^{+}88, SS88c, SHK^{+}87, TRC^{+}87]. procollagen [BFKF86, PRS^{+}88]. procollagenase [UW86]. Procyclin [RSP^{+}89b]. produce [COMS87, CL85b, JRB88, TBK86, WSW85]. produced [BBV^{+}89, CMW87, DKH^{+}85, FRR85, MA86, OB88, RJH^{+}87, SF87]. produces [MTA86, SS85]. producing [DRME89]. product [BHS88, BHAS85, FBP89, GD89, GORW88, HD88b, HOPM^{+}87, HML88, JFAT89, NGI^{+}89, RE85, WRR87]. Production [RE86, BFA88, LSAKO86, MSE89, NR89b, OS88b, RMP^{+}88, RH86a, STF85, SH86]. products [HD86, MABJ^{+}86]. proenzyme [BVD87]. profilin [HCHK89, KSEP86, LJC^{+}87, VKP89]. profound [ODD85]. profusion [Koz88]. progenitor [GHA88, HVL87]. progeny [TTF87]. progesterone [PAGPLM86]. program [CKR85a, KFH^{+}86]. programmed [FCL86, LR89]. programs [CKF89]. progression [BBB89b, CPC^{+}89]. Progressive [LSB89, DSG86]. proinsulin [ORA^{+}86, RH87]. proinsulin/insulin [RH87]. proinsulins [POCM88]. projection [GH85b]. projections [OE85]. prolactin [FZ85, TLPTV85]. proliferate [GSS89]. Proliferating [OOCT89, AU89, BVHF89, BMB87, DYT^{+}87, TB88a]. proliferation [CCL^{+}86, DKH^{+}85, FSM88, FKM87, GB85b, GLH85, HBC89, Hat85, Hat87, HE86, KST^{+}89, KSSAQC89, KF89b, LSDP86, MGD88, MMRR86, RBG85, RKA^{+}89, SVF^{+}89, WBDR88, YMD87]. proliferative [MSS^{+}89, WS86c]. prolongs [Hep85, MTA86]. prometaphase [GHL^{+}89]. prominently [ACPFB^{+}85]. promiscuity [HF85a]. promote [BLR88, GDvdM87, LS85, NTLR88, SFLM86]. promoted [BAG86, RLJ86]. promoter [MMRR86, SVG^{+}86, TYP^{+}87]. promoters [NR89a]. promotes [CTS^{+}87, DLG^{+}89, GK87b, GS89, GRvdM89, JVG^{+}85, LFR85, MTCE87, Rap89, RG86b, SKS^{+}87, UF86, WG86]. promoting [CSK^{+}88, CMP86, DFSK85, EDD^{+}86, GY86, JFP89, MHF86, MEVM89, PCBD89, RMP^{+}88, SCP85, GABT86]. pronuclei [SR85b]. propagation [SW86]. propeptide [POCM88, SS89d, HRP87]. proper [DDB^{+}88, KG89, RDS^{+}89]. Properties [DRRB88, MK85a, MK85b, SC87, BMVB89, BS89, BWSB86, BBH87, CS88b, CM86b, FCZ^{+}89, HBF^{+}88, HDG^{+}85, Kro86, LP86, MHK^{+}85, MB85b, PSV87, PV85, RDJ^{+}86, SSSP88, SDDZ^{+}89, SGM^{+}86, TJ85, VT89, VB88a, WDV87]. prophase [SROB86, SG86]. proposed [TFD87]. prosomes [dSdSH^{+}88]. prostaglandins [EZBW88]. prostanoid [ECZ86]. protamine [MSV^{+}88]. protease [BIES89, GW89, JFP89, MGSG88, MDJ89, SEB89, dLG86b]. protease-accessible [JFP89]. proteases [MKK89]. protects [SMSR88]. Protein [AG89, CC89, CS88d, DF87, DH89b, KC88, KJW89, NJD^{+}86, NBS88b, PBvTS88, PMV^{+}89, WCB86, vZC87, AHC89, AY89, ACN^{+}89, AHM89, ALCG88, AK86, AEB^{+}88, ASJ^{+}88, AVP^{+}89, ABG^{+}89, AKVF88, AB88, ABM86, ASLFP89, BSR^{+}89, BTL^{+}86, BH87, BWK87, BRKE88, BNLC85, BS88a, BS88b, Bec86, BMBL89, BC89, BD85, BSY86, BF88, BHAS85, BPG87, BKPG89, BTR^{+}86, BS89, BF87, BHK86, BPKS86, BDS85, BSNG85, BWSB86, BSK86, BCD86, BSd^{+}88, BP87a, BFK87, BCK85, BSB^{+}86, BCMK87, BA86, CS88a, CMT^{+}88, CWR86, CMB^{+}87, CC85a, CCK86, CRR87a, CW87, CS88b, CFC89, CCLC89, CCMB87, CPC^{+}89, Cha85, CM86b, CTS^{+}87, CMNB88, CHK^{+}86, CRLW86, CMW87, CHH^{+}88, CSG85, CHM^{+}87, CJH85, CG86b, CM87c, CHE87, CMTC85, CMK87, CWH86b, CDTS85, CG86c, CMG86, DSA86, DMD^{+}86, DAJJ^{+}88]. protein [DYT^{+}87, DS87, DS89, DNSMA88, DJM^{+}89, DMW88, DKHB87, DRM^{+}88, DBK87, DK86, DKKK88, DBB^{+}87, DRRB88, DMRB87, EKH89, ES86, EMZ85, FGW85, FDG88, Fec87, FBF^{+}89, FSM88, FBP89, FMJS89, FFH88, FMT^{+}87, GB85a, GST^{+}87, GORW88, GSDR89, GK87b, GM88, GCR^{+}89, GS86a, GKG88, Gla85a, Gla85b, GKZ89, GSTC87, GLM86, GDF86, GCBH86, GH88, GDC^{+}86, GG87a, GCGR85, HFGD89, HLW^{+}89, HFHR88, HS86a, HS85a, Has88, HSK88, HBKK87, HBB^{+}89, HvdSdC^{+}89, HL88, HTT89, HHY88, Hir86, HH87, HMHWS86, HHMH87, HC86, HVR^{+}85, HAM86, HUA^{+}86, HLU^{+}87, HML88, HWLS88, HABM85, HFB^{+}89, HHSF85, HGL88, HBHL^{+}89, ICP^{+}86, ITY^{+}88, JLK89, JLH^{+}86, JFP89, JJP^{+}89, KOF88, KBKC86, KSH87, KSC^{+}89, KSJV88, KH85b, KKB85, KHC^{+}86, KWS^{+}89, KCN^{+}89, KSKW^{+}85, KG86, KRN^{+}86, LFC^{+}88, LFFW89, LA87]. protein [LR85a, LCBS89, LPR^{+}89, LGC^{+}88, LC85, LT89, LWM^{+}88, LYG85, LD86, LD88, LBFS88, LAC^{+}85, MS86a, MKK89, MCS85a, MP89, MBKR89, MSD^{+}88, MS86b, MvdMH^{+}88, MSS^{+}87a, MSC^{+}85, MM89, MMM87, MMRR86, MMS89, Miy88, MW88b, MM87, MK85c, MTK^{+}87, MC89b, MDW86, MPB88, MKA^{+}86, MCCM88, MFM89, NSY86, NBS88a, NM89, NPCL85, NR88, NB88a, NFF86, NCB88, NSM^{+}87, NLC89, NYRK87, NHKE88, OMS89, OON89, OPN^{+}85, OHH^{+}89, PL85a, POW^{+}87, PH89, PBBP85, PFD^{+}89, PN89, Pau86, PV85, PAHC89, PF85a, PHRM87, PL85b, PAF88, PM87b, PFS85, PCBD89, PSK^{+}89, PRSGD89, Pra86, PMR89, PHTK87, RH89, RGW^{+}89, RFH85, RRB86, RP89, RMM89, Rog87b, RPCB88, RFZH85, RL89, RR86, RSL^{+}89, RKFN88, SGRM85, SCK^{+}89, SVF^{+}89, SS85, SN89, SYSP88, SOB^{+}89, SPF87, SDS89, SL85c, SLE86]. protein [SDSO86, SFB^{+}88, SP85, SH87b, SAA89, SW85, SB85d, SG88b, SV89, SSW^{+}85, SMA85, Sny89, SA85, SL86b, SWTB88, SMS^{+}85, Ste85, SBHN86, SQM89, SWC^{+}87, SGGS88, SB85f, TC87, THM86, TKP^{+}88, TSI^{+}89, Tra88, TT85, THT89, TIT89, TJF86, UK86, VHO^{+}89, VP85, VN88, VSSGRB87, VSSRB87, VSZ^{+}87, VS88a, VRLG89, WMUP^{+}85, Wan85a, Wan85c, WKKS^{+}87, WB86, WKS85, WMBR89, WM88, WW87b, WG86, WLGB85, WVCR89, WS89, WLW^{+}88, WRR87, WTB87, YPG86, YB87, YMM86, YLGG85, YBWLM88, YSW^{+}85, ZSSR86, ZHES89, ZN89, ZDH88, ZH86, dVOD^{+}89, GCPE85, NRTE89, SLB^{+}89, SBHN86, VP85]. protein-1 [KRN^{+}86]. protein-2 [JFP89, KSC^{+}89]. protein-calmodulin [CM87c, MC89b]. protein-heparan [CSG85]. protein-mediated [Miy88]. protein-specific [HvdSdC^{+}89]. protein-tyrosine [DAJJ^{+}88, GCBH86]. proteinaceous [MWSF87]. proteinase [CC88, CCG89]. proteinases [MTRR89]. Proteins [AKVF88, HGL85, AC86a, AFMD89, AG87, AST89, AM86, AB89, APF^{+}88, AC87, ABM86, BTL^{+}86, BB87a, BMEG89, BS87, BFSH87, BHSJ86, BVHF89, BHS88, BBJG89, BS89, BLR88, BGF88, BTV^{+}89, BSH^{+}85, BM88, BBH87, BJMK89, BLK^{+}88, BSG85, BDB^{+}87, CPHM86, CA86, CS86b, CB89, CFB^{+}85, CKF85, DFF85, DRY89, DVHC89, DMB88, DMC85, DDBHL86, DF88a, Ear87, EB87, EBGS86, EBCO88, ET87b, FCM85, FLH^{+}88, FW85, FGA86, FH86, FOW89, GD89, GSS^{+}88, GFA^{+}88, GTP87, GDM^{+}88, GACOH85, GKS88, GKH^{+}89, GZGM87, GPSM85, GL85, GCPE85, GS86b, HKKP86, HPO88, HNR87, HKM^{+}87, HP87a, HFZM87, Has88, HK88a, HBV85, HTH85, HSO88, HBB85, HJV^{+}89, HP85b, JRB88, JS86c, JBTA88, JTBA89, JPM85, JHP85a, JHP85b, KB86, Kee87, KFA89, KSJF87, KMM88, KTT^{+}86, KCN^{+}86, LW86, LH86, LSEN87]. proteins [LEG^{+}89, LFHD89, LKHR87, LKF87, LL89b, LWMM^{+}85, LMK88, MNC86, MMBA89, MP88a, MH88, Mat86, MWM89, MFA89, MSS^{+}89, MK85c, MIHS86, MG86b, NSF^{+}89, NKMK86, NFF86, NL86, OCP86, PZBD87, PHL89, PKS^{+}89, PSR^{+}88, PMV^{+}89, Pri87, PMR86, RO85, RT88, RPW86, RBKK85, RP86, Rob87, Rob89, RHL^{+}85, RG87, RDS^{+}89, RYR^{+}89, Sch86a, SCRA87, SHN87, SB87, SG88a, SCM87, SKAW86, SE85, SSO89, SPL87, SLL89, SMA85, SALN88, SD87, SGP88, SBG89, SEG^{+}86, TS88, TT86a, TS86, TT86b, TTF87, TNG89, TLD^{+}89, VS88b, VC89, WRS^{+}86, WLJM^{+}89, WAG^{+}89, WLGB85, WRL86, WCMG88, WMSB^{+}89, WFF86, WMH^{+}88, YBBS87, ZR89, ZV89b, CHG87]. Proteoglycan [RFH85, ASCM89, BFL88, CS88a, CRT87, CW87, Cou87, GLL^{+}86, GSS89, GLK^{+}89, HBC89, HvdSdC^{+}89, HCE88, JRSB87, KG85a, KG85b, KIKS87, KKSK88, KW86a, LCR^{+}87, MAC89b, PJ87, RJB86, Rap89, RR88, RBG85, SSC86, SB88, SJOB89, STE88, SRH^{+}87, SABVK89, TIS^{+}86, VVL^{+}89]. Proteoglycans [SFS^{+}87, BMGI85, CE89, CTJG88, DLH^{+}89, FM86, JNR^{+}85, JRB88, KBO^{+}88, MOT^{+}85, PCC85, SKU^{+}85, SF87]. proteoliaisin [WKS85]. proteolipid [HFB^{+}89]. proteoliposomes [DMRB87]. proteolyses [KW86b]. proteolysis [CC88]. Proteolytic [LKOM88, BMR85, CMB^{+}87, FGO^{+}85, FHE88, JWH85, NADS86, RWE^{+}88, SMSR88, SPT^{+}89]. proteolytically [VT89]. proto [HSK^{+}89b, MAE88]. proto-oncogene [HSK^{+}89b, MAE88]. protofilament [MSR^{+}86]. proton [BGH87b]. proton-secreting [BGH87b]. protooncogene [NGI^{+}89, SDR^{+}88]. protoplasts [MVSW86]. protozoa [Heu86]. protozoan [GG87a]. protrusions [DWFT88]. protrusive [Kol86]. prourokinase [CSM^{+}89]. prourokinase-type [CSM^{+}89]. provide [HSMCL87]. provides [PSM89, THFB87]. providing [FMG89]. proximal [GHL^{+}89, RSF86]. psaA [KMM88]. psaA-psaB [KMM88]. psaB [KMM88]. psbA [KMM88]. Psophocarpus [BWL^{+}86]. PtK2 [BS88a]. pulses [Sow86]. pump [AC86b, MSLR^{+}86]. punctulata [WBGV85]. pupal [DFF85, DLC88, WFF86]. pupoid [FJR87]. Purification [HWLS88, KSEP86, LFR85, LYG85, MKA^{+}86, PV85, RP89, SWW85, TDR88, UF86, WN87, dCFB86, dHTG88, AK89, FDB85, GGMH86, GW89, JS88b, Kee87, LHHC85, MMM^{+}85, PHTK87, SNCY87, SB85d, THT89, HW86b]. Purified [EBCO88, COA85, DBB^{+}87, DMRB87, HS85a, HF85a, PJ87, VB85b]. purify [WRL86]. purine [VRLG89]. purpuratus [SB85a]. putative [BJ86, DS89, DRC^{+}86, GGH89, JLB86, KJW89, KDBF86, NRL^{+}89b]. pyrene [PBS85, WB85]. pyruvate [HL88].

Q- [DJM^{+}89]. q23 [NMM^{+}86]. quaking [Tra88]. qualitative [DD88]. quantal [HTTFC85, TTGFC85, VJF^{+}88]. quantitate [RGW^{+}89]. Quantitation [SBS87a, AJRNBJ85, BD86a, BBB88, VS88a]. Quantitative [FMG89, KDB87, LMW^{+}86, DCQB88, DD88, GHL85, GBM89, KPK86, LBEM89, ZSE85, DJS^{+}88]. quantitatively [MKH88]. quaternary [DRM^{+}88]. quick [BO85, GH85a, Heu86, LC88a, NH87]. quick-freeze [BO85, GH85a, LC88a, NH87]. quick-freezing [Heu86]. quiescent [MMT85, MSLR^{+}86, SNNK85, ZSSR86]. rabbit [GML^{+}89, JSA^{+}89, KG85a, KG85b, KIKS87, MR85a, SS89b, SSK89, VKB89, YSW^{+}85]. rabbits [WE87]. rachitic [KDFG89]. radial [GH85b, WMR86, WVCR89]. radiation [Ros89]. radioautographic [CMHC85]. radiolabeled [BCA^{+}86]. radixin [THT89]. raises [CR86]. ram [PFD^{+}89, WHL^{+}86]. random [TLZ88]. Ranvier [FCMK^{+}86, RDPR^{+}86]. Rapid [Bis86, Bre89, JPH85, MLPF89, MSK^{+}87, TSBW86, Wan85c, Wat85, ASF89, BD89, CNHF85, HBV85, KLHS85, MT86, MHWB86b, RMS88, SPO85, TTM88, WKRH86, ZP85]. rapid-freeze [ASF89, TTM88]. Rapidly [HB87, FBPS87, LFS^{+}89, ROCK86]. RAPsyn [MFM89]. ras [BSSMF88, KSSAQC89, SEG^{+}86, BS86, GZO88]. ras-transformed [BSSMF88]. Rat [LYE87, MH86a, SHL87, AK89, Abr85, AP86, ACPFB^{+}85, AMWN89, AB88, AC87, BPB86, BBBM89, BBH85, BFSH87, BVHF89, BPG87, BTO88, BF87, BAD86, BPKS86, BRD^{+}88, BTV^{+}89, CDB87, COMS87, CMHC85, CRT87, CS88b, CMM^{+}85, CTJG88, CSGC88, Cou87, DFSK87, DFdN^{+}89, DSGS87, DCH86, EWAP85, EMH^{+}87, FCMK^{+}86, FCVDH89, FJY86, FBCQPR87, FGA86, GDD88, GJD^{+}88, GSF^{+}88, HAB^{+}88, HFHR88, HBC89, HOPM^{+}87, HCH85, HHS^{+}85, HSS89, HCG87, HBB85, JLK89, JSC85, KPK86, KSC^{+}89, KMP88, KNGH^{+}85, KG86, LBO^{+}85, MSV^{+}88, MYT87, MA85, MHWB86a, MRD^{+}89, MF87b, MH86b, MB87, MTA86, NFI^{+}87, NR89b, NTNT85, OSP^{+}85, PSBN85, PJFB89, Pau86, PV85, PAF88, PSD^{+}85, PBT^{+}86, PS86, PB87, Pum89, Qua85a, Qua85b, RDSL89, RO85, RBMM88, RSF86, SSR89, SPF87, SFJU86, SJ87, SK85]. rat [SCOE^{+}86, STS^{+}89, SR86a, SGB^{+}86, SSSA87, SGM^{+}86, SHP85a, SHP85b, SHK^{+}87, SBC^{+}87, TJ85, TYP^{+}87, TS86, TDR88, TT89a, TSOC89, VP85, VN88, WH85, WS85, WWN87, WMH^{+}88, YMT85, YF87, YMM88, YKCR88, dJHMB86]. Rate [Pol86, BBP^{+}85, BLSK88, CMR^{+}87, MDM88, ODD85, SS89c, WOP^{+}88]. Rates [MP86b, SC85a, CGD86, SCLG86, WSH85]. rather [RH87]. ratio [BFRT87]. rats [AC86b, Bis86, ZHE85]. ray [BMDN87, MMS89, STGK87]. rbcL [KMM88]. RCC1 [OON89]. rDNA [BG86]. re [CHD^{+}86, WL86]. re-evaluation [CHD^{+}86]. re-examination [WL86]. reach [PMC^{+}88]. reaches [vDSP^{+}88]. react [AB89, BKPG89, HKKP86, MH88, WCPK88]. reactants [SPT^{+}89]. reacting [Hir86]. reaction [HS89, LS89, LS87c, TI85, TES^{+}87]. reactions [CDL^{+}85, KdWB^{+}89, MP86b, Pol86]. Reactivated [RH86b, CM86a, GTG85, GAB86, NT85]. Reactivation [KS86, MMC88]. reactive [HMHWS86]. reactivity [CP87b, SKK^{+}85, dGRS87]. reacts [DHL85]. Real [CPS88]. Real-time [CPS88]. Reappearance [FCVDH89]. Rearrangements [BTV^{+}89]. reassembly [BH86a, LBW89]. Reassociation [CB89]. Receptor [DCH86, LBL^{+}85, MFKVK87, MH86b, PS86, WH85, AK89, AYCY89, AAM^{+}87, AU89, ABME88, ALL89, ALWL89, ASA^{+}87, AGPT88, BB88a, BYB^{+}88, BLSK88, Bis86, BLSL88, BGH^{+}87a, BVD87, BW86, Blo86, BF87, BSBZ88, BSLS^{+}89, BR89, BNL89, BGHvF87, BHM89a, BJ86, CR88a, CGV^{+}86, CWOK87, CAG89, CWH^{+}86a, CBYK88, CPP^{+}87, CPG^{+}89, Col86, DZ85, DWT^{+}89, DMD^{+}86, DRP87, DAS^{+}87, DAJJ^{+}88, DNM86, DSTJ85, DMGS88, DRC^{+}86, DNI^{+}88, DLFE89, DCB^{+}88, EG89, ER87, EMH^{+}87, FG89, FSM88, FKC87, FC89a, FMT^{+}87, FTC87, FS86b, GF86a, GF86b, GR89, GMB89, GSS87, GFSS86, GDM^{+}88, GDRB89, GGA^{+}89, GRME^{+}89, HD88b, HA89, HB88b, HGMS88, HKPS89, HSS89, HDG^{+}85, HMS^{+}89, IMP^{+}87, JLB86, JTBA89, JNA85, JSA^{+}89, KWW85, KKSK86, KLHS85, KSSP86, KCN^{+}89, KBK86, KDBF86, KRLU87, LWSF85, LU87]. receptor [LCD^{+}89, LCW^{+}89, LWR85, LGW^{+}89, LHK^{+}89, LDKG87, LBP^{+}86, LP85b, MBKR89, MOP^{+}85, MGM87, MDM88, MLM85, MSW^{+}86, MOU88, MHWB86b, MBK^{+}86, MMRR86, MMS89, Mos88, MQ89, MBH87, MUM87, MBSMD86, MCCM88, MFM89, NJAS89, OSD85, PZBD87, PTW^{+}85, PL86, PP88a, PBvTS88, PF85a, PAGPLM86, PSR^{+}88, Pfe87, PON^{+}88, PB87, Pum89, RRF87, RLB^{+}89, RG86b, SM89b, SLB^{+}89, SB88, SFJU86, Sch86b, SKK^{+}85, SBD^{+}89, SL85b, SBS87a, SOP85, SR85c, SDZB^{+}85, SHP85a, SHP85b, TLRW86, TH89a, TSBW86, TJ85, TRT^{+}88, TLZ88, TS89, TH89c, UTK88, VAD^{+}85, VG86a, VG86b, WSM^{+}87, Wal88, WWW^{+}85, Wat85, WCPK88, WGPH^{+}89, WMBR89, WKRH86, WTB^{+}89, WGC85, WG86, Wil88, WN88, WTB87, YSR87, YBSR87, KHB88]. receptor-aggregating [GDM^{+}88]. receptor-associated [MCCM88, MFM89]. receptor-bearing [DSTJ85]. Receptor-bound [PS86, MHWB86b, Mos88]. receptor-cytoskeleton [WGC85]. receptor-deficient [KKSK86]. receptor-hyperproducing [HGMS88]. receptor-induced [VAD^{+}85]. receptor-ligand [TSBW86]. Receptor-mediated [DCH86, MFKVK87, MH86b, WH85, DZ85, DMGS88, GF86b, GFSS86, HA89, JNA85, MUM87, PP88a, PBvTS88]. receptor-negative [MH86b]. receptor-positive [GF86a, MH86b]. receptor-rich [Blo86, FMT^{+}87, WWW^{+}85, WTB87]. Receptors [YBSR87, AJCMR88, BNB^{+}88, BF87, BM89, BRS87, BGF86, BSB^{+}86, BLK^{+}88, BBMW85, CP87a, CCP88, CFC89, CPG^{+}89, DBM^{+}89, DCC^{+}88, DWO^{+}87, DK88, DMM89, EZWB88, GSS^{+}88, GBH87, GRCG86, HB88b, HB88c, HM89c, JB87, JBTA88, JB86, KWW85, KSJV88, KBK86, LWH88, LS89, Mas85, MGK89, OH88b, PE85, PMC^{+}88, PNB89, PKN^{+}89, PFPM86, RRP88, RVS88, SCLG86, SMH88, SFMT87, SS89c, SSK^{+}88, SSKK89, SR85c, SF88b, SGG^{+}89, SKC85b, SCD88, TFBW85, TL89, TTBG87, TDR88, TCP^{+}85, TCK87, UF86, VB88b, Wal86, WC87, WCPK88, WBD^{+}86, WDF86, ZBSS^{+}87]. recessive [DM87]. Reciprocal [WGC85]. recognition [ABME88, ALWL89, BWL^{+}86, HPW89, Ibr87b, ISM85, JBdlTB87, LD86, RLB^{+}89, TLRW86, WD86a, WDV87, WKBR87, WW89, WJSB88]. recognize [PF85b, RG87, SSO89, TLPTV85, WRS^{+}86]. recognized [BG86]. recognizes [DRP87, GVGL85, MAG88, RE85]. recombinant [CPG^{+}89, PRS^{+}88, WMSB^{+}89]. recombination [LYG85]. recombinations [HF85a]. reconstituted [DMG^{+}89, DMRB87]. Reconstitution [BWK87, MBvS^{+}89, PJFB89, RKFN88, SCGS85, BFG86, CJ85, FLH^{+}88, VB88a, WS87]. reconstruction [BDCT88, CD89, SKL85, TRCR89a, TRCR89b, YF87]. recording [Kro86]. recovering [ATNH86]. recovery [DH89b, WS86b]. recruitment [CSL87]. recycle [WDF86]. Recycling [NVB^{+}88, CGV^{+}86, DMM89, FS86b, KP89, LCW^{+}89, MDM88, MBSMD86, PW86, SM89a, SDZB^{+}85, VJF^{+}88]. red [ABM86, GBC^{+}86, GMF85, HP87a, LHFB88, WS89]. redirect [SYS85]. redistributes [DRY89]. Redistribution [BWS86, HPM87, LS87c, MBLS87, BMBL89, BA86, GBH87, IHJ88, MP86a, RRP88, SCM87, SALN88, TSBW86, TFBW85, WS86a, WKRH86, ZBSS^{+}87]. reduce [EMZ85]. reduces [RG86b]. reductase [JBG^{+}87, KPK86, KSFL89, MYT87, SBG^{+}85, SHR86, WBDR88]. reduction [MNM87, MP86b]. reendothelialization [WG88b]. reentering [Wan85c]. reexpress [BBP88]. reexpression [BB88b]. reference [HTH85]. reflection [GTB85, LWT85]. reflects [DH89a, PMJ87]. reformation [BRR^{+}87, PB87]. regardless [KSBB87]. regenerating [AM85, DCPR^{+}89, FLBP^{+}86, HTGP85, MS88a, YF87]. regeneration [CMP86, MIHS86, RNPR^{+}88, SSH87]. regeneration-promoting [CMP86]. region [BB88a, BB87c, CM87b, CG86c, CMG86, EB87, HFK^{+}89, MR87, RMM89, WKKS^{+}87]. regional [TDGE85]. regionalization [WHL^{+}86]. regionalized [WSM86]. regions [BSY86, Ear87, HCHK89, HBV85, HFC^{+}86, KdS86, MACO88]. regrowth [BH86a, TWB88]. regulate [DDB^{+}88, DCC^{+}88, FKC87, GB85b, MCB85]. Regulated [SEB89, AG87, CMT^{+}88, CHMB86, CM86b, CB87, CHS85, CSL87, DLC88, ECZ86, FL89, FJY86, GLD^{+}88, HD88a, HAB^{+}88, HDP^{+}85, HZS88, HVL87, HMP85, KFHF85, KMM88, KS88, LGC^{+}88, LLC85, MND^{+}87, MMM87, OTS86, RH87, RM89b, RHL^{+}85, SPS89b, SW87a, SDDZ^{+}89, Smi87, SMW89, SBG^{+}85, Try89, WDR89, YKSR89, FC89b, GG85, GTP87, GDF86]. regulates [AR88, ALCG88, BSd^{+}88, EBBW87, HBSQ^{+}85, HB88b, MB85a, PC85, SSS85]. regulating [AKVF88, DKHB87, LL89a, Wri85]. Regulation [BYB^{+}88, BHSJ86, BAG86, CWH^{+}86a, DNB88, DKKK88, GAB86, GBB89, HMP86, HSS89, JTBA89, KOBLM88, KSJF87, LFC^{+}88, OJHAH89, OSH^{+}86, OMM87, PZBD87, PCK^{+}87, PN89, PKN^{+}89, SE85, SE86, TJF86, VSZ^{+}87, Wal88, WS87, WS86c, vAAA86, ACN^{+}89, BN88, BB85, BP87b, CL85a, CW88, CKR85a, CDL^{+}85, CH87, DWT^{+}89, FCL86, GP86, GCGR85, HWDC89, Hat85, HSK88, HS85b, HS87b, HSMM85, JHPS86, KST^{+}89, KSSAQC89, LDKG87, LC88b, MW88a, MS86c, MW88b, MVB85, NR89b, OGF87, OGJ^{+}89, PN88a, PN88b, PFRS89, PL86, RH86b, SMR87, SR88, Sch86b, SR86a, SDR^{+}88, TF86, WG86, WPW^{+}85]. regulator [Bro87, OON89, RTK^{+}86]. regulators [GABT86, HP85b, KTF87b]. regulatory [BH86b, HMBV^{+}88, HFK^{+}89, IGSGD87, JBTA88, KSJV88, KJY89, LWM^{+}88, SLB^{+}89, TGNG87]. reinhardtii [AWJ85, GGMH86, HMP86, Kam88, LP86, MMR88, MMWM^{+}89, MSY^{+}87, PG87, SSH87, SBS88, SS89a, SDSO86, WSJ85, dHTG88, dVOD^{+}89]. related [APF^{+}88, ABM86, BBGI85, BBB89b, CCLC89, CRC^{+}87, DWT^{+}89, EBBW87, FG89, FKC87, GST^{+}87, GCR^{+}89, GACOH85, KMP88, LBB^{+}88, LLCA89, LP85b, MSQ^{+}89, MSLR^{+}86, MPB88, NDJ^{+}89, NKMK86, PKS^{+}89, Rob89, SGS86, SQM89, TDR87a, WRL86, WSH85, ZSSR86, DNSMA88]. Relation [DWS88, HB87, KF89b]. Relations [SGG^{+}89]. Relationship [SOP85, BF87, BS86, BB85, CM87a, CCK^{+}88, DBK87, DW86, GF85, GW89, GGJ^{+}87, LOH^{+}86, MGR86, MVB87, OLGT86, SB85b, TI85, WHMA86, YBSR87]. relationships [MD85, MF87a, PP88b, YMM86]. relative [BIES89, BJMK89, HAK^{+}88, RDJ^{+}86, RGBM87]. relaxation [GBH87, IKH87, TFBW85]. Release [RBKK85, SD88, AC87, AS86, BGM^{+}87, BFJ^{+}85, CSR^{+}86, HTRBP86, HFG87, KBPW89, MBKO86, Miy88, OHK86, PSD^{+}85, RRB86, RRF87, SIR^{+}88, SBHN86]. released [ER85, GORW88, GB85b, MMWD89, RH87]. releases [GCR^{+}86, PF87]. remains [TPBB85]. remodeling [And86, SHBV89, SS89b]. removal [Gri86, VGKF86]. remove [SNSS87]. renal [HKH^{+}88, MCA^{+}89, SS89b]. Renin [FCK87]. Reorganization [BMB86, MW86, UW86, BBBM89, BK86, Bre89, LBO^{+}85, NH87, OH88c, SSW^{+}85, WSV^{+}89, WG88b]. reorientation [MSS^{+}87a]. repair [DLG^{+}86a, DKH^{+}85, PML^{+}89, TB88a]. repeated [FJR87, GSTC87, NLC89]. repertoire [WVLC86]. replacement [CKR^{+}85b]. replica [IMP^{+}87]. replicase [FKH88]. replication [AKC^{+}86, BMB87, OO87, OOCT89, SMS^{+}88, WLG85, WPW^{+}85]. replication-competent [SMS^{+}88]. replicational [NB89a]. replicative [VP85]. repressed [CLOH87]. repressible [NR88]. reproduction [SMR86]. reproductive [SR85a]. required [BSR^{+}89, BMEG89, BGM^{+}87, BHAS85, BDS88, CB86, DS89, HD86, HW85b, HFF^{+}85, KG89, MR87, NBS88a, OMM87, PL87, PA88, RBG85, REB^{+}86, RT88, RKA^{+}89, RDS^{+}89, VSKSP86, WMM86, WG87b, WN88, WRR87]. requirement [MTRR89, RVS88]. Requirements [PJW88, ETT88, FLBP^{+}86, LFHD89, RH86b, TAHC^{+}89]. requires [BSDH86, CVWZ89, CG86b, DVHC89, HD88b, ISL87, VSSGRB87, WB86, WMBR89, WCMG88]. rerouted [DCWJ89]. Rerouting [LBFS88]. resact [WBGV85]. resemble [TSI^{+}89]. resialylation [SR85c]. resident [DRY89, PTB^{+}85]. residue [NRL^{+}89b]. residues [FCK87, HFF^{+}85, WMSB^{+}89]. resinless [FKP86]. resistance [LCL^{+}89, SAGM88]. resistant [CM88a, DAS^{+}87, MSC^{+}85, MCC^{+}89, OMY^{+}85, RC85a, TASK86, TT88b]. resolution [STE88]. resolved [KHKR86]. Resonance [UP86, JT85]. resorb [BKC^{+}86]. resorbing [BNLC85]. resorption [BNLC85, DWT^{+}89, LR85b]. respective [DCC^{+}88]. respiratory [Gor85, NSF^{+}89]. respond [GRCG86, RYF^{+}87]. responds [SN89]. Response [OS88b, DWL86, DFF85, HNR87, HKM^{+}87, HKBT87, LEG^{+}89, LBEM89, MER^{+}86, MTA86, PE85, RR88, RBKK85, RC85b, SS89b, SL85d, SAA89, SGBF87, TSBW86, TT86a, TT89b, UPRS88, VES85, WS85, dJHMB86]. responses [Bas87, GDC^{+}86, MLC^{+}88, OW86, PUW^{+}88, Rob85, SC88, VRLG89]. responsible [HW86b, MP86c]. responsiveness [CJ86]. Resting [CS86a, BFPS85, KDB87, UTK88, VAD^{+}85]. restoration [MQ89]. restores [JWH85]. restricted [CHG^{+}86, GLL^{+}86, GFSS86, KSF^{+}88, MMK^{+}89, WAG^{+}89, YKSR89]. restriction [Hep85]. result [Kam88, ODD85, PTB^{+}85, RDJ^{+}86, Too85]. resulting [MOP^{+}86]. results [AHM89, FBE85, GLC88, HGS89, NMPG88, PHL89, WMUP^{+}85, WKRH86]. retain [CB88, GPSC86, ZR89]. retained [CBYK88, GRME^{+}89]. retard [ZR89]. retarded [ZSE85]. retards [FFH88]. Retention [BSH^{+}85, PC85, MR87, PA88, SYH^{+}85, SWC^{+}87, SBG89]. reticulocyte [WW89]. reticulocytes [PTW^{+}85, PL87]. reticulum [BWK87, BB89, BHAS85, BICFA88, BSH^{+}85, CHFA89, CC89, CCS^{+}86, DMF^{+}87, DS87, DS89, DRY89, FRG^{+}87, GORW88, GRME^{+}89, HFGD89, HOPM^{+}87, HBB^{+}89, HAM86, HGL88, Ibr87a, Ibr87b, JBG^{+}87, KA85, KR88, KS85b, KSH89, KJW89, LFC89, LKHR87, MVK^{+}88, NBS88a, NM89, NRL89a, PLA86, PTB^{+}85, PA88, RDS^{+}89, RKFN88, SCK^{+}89, SIR^{+}88, SPL87, SALN88, SWC^{+}87, SGGS88, TCF86, TKFH89, VVL^{+}89, WCMG88, YMT85, YF87, ZR89]. retina [CS89, CAG89, CSG85, CMTC85, MR85a, SL85c, SL85b, SLE86, TAMR85, WD86b]. retinal [BA87, FRH85, GAB86, HNR87, LN88, MEA89, MH86a, MM87, NTLR88, NL86, OE85, PN85, TB88b]. retinoic [BBP88, BB88b, KF89b, LLRR87]. retinoid [BS87, ETA85, ET87a]. retinoid-binding [BS87]. retinoid-induced [ETA85]. Retinoids [KTF87b, KSJF87, MBL89, RBKK87]. retinol [HVR^{+}85, SLE86]. retinol-binding [HVR^{+}85, SLE86]. retraction [DLBH89]. retrograde [ML85, TJ85]. retroviral [CPG^{+}89, GDD88]. retrovirions [Too85]. Retrovirus [SS88c]. Retrovirus-mediated [SS88c]. return [DK88, NVB^{+}88, SR86b]. reveal [GL85]. revealed [Ake89, BN87, BRS87, BDB^{+}87, FONW88, GTB85, HHMH87, KdS86, KCAF89, KBL^{+}86, MYH^{+}85, MF87a, MABJ^{+}86, MAE88, NH87, NWE^{+}88, SH87b]. reveals [AF89, LHK^{+}89, MC89a, STE88, WKKS^{+}87]. Reversal [LR85b, MQ89, NT85]. Reversible [FA85, LJC^{+}87, CV87b, GDS87, HCHK89, KLHS85, RGW^{+}89, SALN88, SE86, ZP85]. review [Rob85]. revisited [Cle87, MP86b]. RGD [BR89, CR88a, CR88b, SMC89]. RGD-containing [CR88b]. RGD-dependent [BR89, CR88a, SMC89]. RGDS [SR87]. rheology [VA85]. Rhizobium [PGH^{+}88, BWL^{+}86, HMHWS86]. rhodamine [MJW86]. rhodamine-labeled [MJW86]. Rhodomonas [LG89]. rhoptry [SYSP88]. ribonuclear [MSS^{+}89]. Ribonucleoprotein [DPM88, CS88d, FLM86, FKP86, GK86, KR86, LAC^{+}85, MLPF89, MBLS87, WM88]. ribonucleoprotein-associated [MLPF89]. ribonucleoproteins [WSRT89]. ribonucleotide [SBG^{+}85, SHR86]. ribophorin [HOPM^{+}87]. ribose [CMW87]. ribosomal [CMW87, HHSF85, MVB85, MVB87, NHKE88]. ribosome [CG86b, HAM86, SBH^{+}88]. ribosome-membrane [CG86b, HAM86]. ribosylation [AW89]. ribulose [CWR86, HMP86, MWS85]. ribulose-1 [HMP86, MWS85]. rich [BKBR88, Blo86, FMT^{+}87, HUA^{+}86, JBTA88, SKAW86, WWW^{+}85, WTB87]. ricin [DSTJ85, GWCW85, MM87, MDVG^{+}85, vDTP^{+}86, vDSP^{+}88]. ricin-gold-dextran [MM87]. rigid [KHB88]. rigidity [WE89]. Rigor [BFAI87, TRCR89a, TRCR89b]. RII [DMD^{+}86]. ring [DLC88, HG89, KC85]. rise [ECZ86, MMT85, SSY^{+}86]. rises [Miy88]. RLC [TGNG87, TGNG87]. RNA [BRR^{+}87, CBS^{+}89, CMW87, CMGS86, CDTS85, DF88b, FDG88, FBP89, FKH88, HW85a, HKTY89, IGSGD87, KCK^{+}87, KR86, KW86b, KMP88, LW86, LGC^{+}88, LSC86, MSD^{+}88, MVB85, MVB87, NWE^{+}88, RPW^{+}89, VN88, VSR87]. RNA-coated [DF88b]. RNA2 [LW86]. RNA3 [LW86]. RNAs [HD88a, MLSW86, NKA^{+}87, PRS^{+}88, SBG^{+}85]. RNP [FKP86, FKP86, LM87, PRSGD89]. RNP-depleted [FKP86]. Rod [TAMR85, FRH85, MM87, NFW89, NRL^{+}89b, SDS89, SDMF85, SQM89]. rodent [GPJ88, LCL^{+}89, SGM^{+}89]. rodlike [HBPK87]. Role [AFMD89, ATNH86, DKHB87, HMF^{+}87, ICP^{+}86, KH85a, KCCM85, KCR86, KKML88, NHKE88, AWH^{+}85, BSLS^{+}89, BR85, BCMK87, CRJM88, CDB^{+}86, DMB88, DRC^{+}86, ESA86, EZWB88, FSFL88, GBBS89, GW87, GCGR85, GSG88, HLL89, HS85b, HMHWS86, HP87b, IF89, KSSAQC89, LWH88, LBW89, MEA89, MB88, MGHG^{+}86, NB88b, Oss88, PZBD87, PBK88, PHM85, PHTK87, RLJ86, RNPR^{+}88, SVSH^{+}88, SLB^{+}89, SSRS88, SS88a, SPO85, TC86, TSOC89, Wan85d, WSR89, WG86, WFL87, WS89, WMC86]. roles [AYCY89, HKTY89, KSJF87, TT89b]. root [RBG85, SSH87]. rotation [SB85a, TI87]. rotavirus [KA85, PTB^{+}85, PA88, SWC^{+}87]. Rough [KG89, KA85, LKHR87, TKFH89]. Rous [DH87, GZ89, HSWH86, LHHC85, MPBZBE89, MSE89, RE85]. route [LP85a]. Routing [vDTP^{+}86, TNG89]. row [FS87]. rRNA [SBH^{+}88]. rRNA/L5 [SBH^{+}88]. ruffled [BNLC85]. ruffles [RGS88]. ruffling [MA86, MTA86].

s [HTRBP86, WMBR89, ZP85, BSY86, BSd^{+}88, HPW89, SBHN86]. S-transferase [BSY86]. S1 [HHSF85]. S100 [GKZ89, WMSB^{+}89, ZV89b]. S13 [WBVL86]. S13-transformed [WBVL86]. S49 [HMP85]. S6 [NHKE88]. SA11 [KA85]. SAC1 [CNB89]. Saccharomyces [HPW89, BMEG89, CSBS89, HW88, JASP88, KAB88, LHM89, MN87, MDJ89, PKK89, PMJ87, SS88b, UMHI85, WRR87]. salamander [TAMR85]. salina [EMT88, MT86]. salivary [DLC88, HMG^{+}86, KC85]. Salmonella [FGF88, JGAR89]. same [BFBL88, CL86, FZ85, PFPM86, RPW86, SSS^{+}87a]. sand [HTSH85]. Sar1p [NM89]. sarcolemma [SS87]. sarcolemma-associated [SS87]. sarcoma [DH87, GZ89, HSWH86, LHHC85, MPBZBE89, MSE89, PSS^{+}87, RE85, SPT^{+}89]. sarcomere [FONW88, HP87b, WW88]. sarcomeric [TGNG87, TSI^{+}89]. sarcoplasmic [BICFA88, CHFA89, CCS^{+}86, DMF^{+}87, SIR^{+}88]. satellite [EMG86, LNH86, MMM^{+}89]. saturable [Wil88]. scaffold [HHY88]. scaffold-like [HHY88]. scaffolds [EHC^{+}85]. scale [BN87, Bas87, MMM^{+}85, RLP85, YE87]. scallop [CHFA89, VC85, VC89]. scanning [AJRNBJ85, Koz89, SMH88]. scattering [SOP85, ZSE85]. Schistosoma [CC85b, FMK^{+}86, GBC^{+}86, SC85b]. Schistosomula [GBC^{+}86, CC85b]. Schizosaccharomyces [AY89, CW89, HPW89, HHY88]. Schwann [CTR86, BLR88, BP87b, BBB89b, CRT87, CE89, EBBW87, Fal85, JMBM85, MBD89, MOT^{+}85, RBG85, REB^{+}86, RDSL89, RNPR^{+}88, SS88a, SPS89a, SPS89b, TYP^{+}87, YMD87]. Schwannoma [MEVM89]. Sciara [Kub87]. sciatic [DFdN^{+}89, HKBT87, MS86b, MS88a]. scorpion [KLS85, SKL85]. Sea [NKA^{+}87, BBW86, BN85a, CWZ89, CDL^{+}85, ER85, FCCC^{+}89, GW89, HBB^{+}89, JWH85, KH85a, KB85, LAA85, MHK^{+}85, MTG88, OHK86, RFSL88, SGH85, Sal86, SB85a, SB85c, SR85b, SBG^{+}85, SW86, SE86, TES^{+}87, TJF86, VMT^{+}88, WMV86, WKS85, YB87, ZSE85]. seam [MSR^{+}86]. search [Sch86a]. SEC11 [BDS88]. Sec12p [NBS88a]. SEC14 [BMEG89]. Sec15 [SN89]. sec23 [RKFN88]. Sec4 [SN89]. Sec53p [BHAS85]. SEC62 [DS89]. second [ZN89]. secondary [DCQB88, HT86]. secrete [GY86, LYE87, SHS^{+}88]. Secreted [SLG88, CMB^{+}87, FFH88, GB85a, ISM85, KMY88, MGSG88, MBH87, PSS89, RJH^{+}87, RGPRB87, RSL^{+}89, SVF^{+}89, YHS88]. secretes [HKH^{+}88]. secreting [BHK86, BPKS86, BGH87b, HSMCL87, RHL^{+}85, SS89b, WDF86]. Secretion [BRHR^{+}86, HUA^{+}86, SYSP88, SM89b, AC86b, BNB^{+}88, BKN^{+}85, BNS85, CCC86, CTJG88, CJ86, CJO^{+}89, CHG87, CJH85, DBK87, EHK86, GWG85, HTTFC85, HBKK87, HCG87, KGRP85, KC85, LMWG85, MB88, MSC^{+}89, MRD^{+}89, MCD85, PCK^{+}87, PTB^{+}85, RT88, SSW89, SKAW86, SEB89, SRPS86, SS88c, SS89d, TKK^{+}87, TTGFC85, UPL^{+}87, VJF^{+}88, ZR87]. secretogranin [RWP^{+}89]. Secretogranins [RHL^{+}85]. Secretory [BTL^{+}86, HHES88, MK85c, WB86, AC87, BMEG89, BGN^{+}88, BPKS86, BK85, BCK85, BCMK87, CJ86, CJ85, DS87, DVHC89, DLC88, FCK87, HFZM87, HHS^{+}85, HAM86, KDB87, LSF89, LKHR87, LMWG85, LMK88, MBLK88, MB87, NFN87, NB89b, ORA^{+}86, PAF88, RM89b, RFZH85, RHL^{+}85, RWP^{+}89, RDS^{+}89, SFJU86, SDDZ^{+}89, SW85, SMW89, SGGS88, SDZB^{+}85, TCBF^{+}85, TT86b, TB87, THFB87, TTF87, TNG89, WN87, ZR89, vGRB86]. section [FKP86]. sectioned [CD89]. sections [BTR^{+}86, CCS87, DCB89, KBB^{+}88, Ris85]. sedimented [BB85]. seed [CV87a, HMHWS86]. seedlings [KMM88]. segment [Eck87, HAK^{+}88, MM87]. segmental [SSSP85, WGPW89]. segments [FRH85, LBFS88, YF87]. segregated [NKA^{+}87, VVL^{+}89]. segregation [FCL86, JBTA88, KG89, Kub87, RM89b, VSSGRB87, WSJ85]. SEKDEL [ZR89]. selected [Eva89]. Selection [MMK^{+}89, SFG87, NZ85, NFN87]. Selective [AIJ^{+}89, CBYK88, DSTJ85, GDC^{+}86, KAB88, KSSP86, PSGS85, RLP^{+}87, ST86, SVSH^{+}88, SSW89, SK85, SBG89, BBBM89, DMG^{+}89, GMF85, HT86, MG86b, Sch85]. selectively [CPC^{+}89, Fec87, MS88b]. selectivity [MBC86]. self [CTSF86, TC86, WD86a]. self-assembly [CTSF86, TC86]. self/nonself [WD86a]. seminiferous [KMP88, MSV^{+}88]. Semliki [CG86c, CMG86, KH85b, RG86a]. senescent [Wan85a, Wan85c]. sensitive [AY89, Bro87, CSD89, EB87, HHY88, KST^{+}89, MMY88, RIPS85, RFMR86, SFK^{+}89, TSS89, WN88, CH85, HC86, WMBR89]. sensitivity [BSLS^{+}89, CHM^{+}87, DDR^{+}89, EMZ85, GDO^{+}85, OS88b, SDDZ^{+}89, YBBS87]. sensory [DBD^{+}89, SS88a, SR86a]. separable [OW86, SM89b]. separate [FZ85, GM88, YKCR88]. separation [SR85b, SMA85]. Sequence [APF^{+}88, BHCD^{+}89, GKG88, GKF85, Ran88, ZN89, ASA^{+}87, AGPT88, BDS87, BFK87, CCL^{+}86, CHH^{+}88, DRP87, DRD^{+}88, DBS^{+}89, DCB^{+}88, DLF88, FTLSS88, GCC^{+}89, GF88, GKZ89, GFS^{+}89, HBPK87, HOPM^{+}87, HFK^{+}89, HML88, MMM^{+}89, MMRR86, MG89, MOP^{+}86, NAS^{+}87, NLC89, PAHC89, PL85b, PA88, RMP^{+}88, SEGS85, SKS^{+}87, SSSA87, WKBR87, WVCR89]. sequences [BGS87, BWL^{+}86, GSTC87, GLM86, HFF^{+}85, LCE87, MWS85, MLC^{+}88, PW87, SHC87, SSO89, ZR89]. sequencing [GDO^{+}85]. Sequential [AG87, CCE87, HS86b, LEFS^{+}87, RS88, FKP86, SM89b, UPRS88, WG87a]. sequentially [SM88]. sequestering [MP86c]. sequesters [SSS88]. Ser [SKS^{+}87]. SERGE [CMHC85]. serial [JS86a, JS86b]. series [CHG87, ECZ86]. serine [MDM88, SEB89]. serine-24 [MDM88]. serotonin [MR85a, PSD^{+}85]. Sertoli [EHM86, NB88a, HBSQ^{+}85, STF85, VS85]. Serum [EG89, CMB^{+}86, HSS85, JVG^{+}85, SHL87]. serum-free [HSS85, SHL87]. set [GD89, WW88]. seven [PSP85]. several [LH86, LFHD89, PP88b, SSC86]. severed [WIS89]. severin [ABG^{+}89]. severing [APF^{+}88, CJY86, SS89a]. sexual [BIES89, Goo89, HMdN^{+}88, PG87]. Sgp [Rog87a]. SH [MER^{+}86]. shape [BBP88, CKR^{+}85b, Heu89a, HO85, JS86a, JS86b, JS88a, KBR87, NYG87, OH88c, RdLB87, SLL88, ZMH89]. shaped [HH87]. shared [KFS87, MS86b, WRS^{+}86]. shares [NRL^{+}89b]. shear [CF88b]. sheath [MIHS86]. sheaths [NPRR86]. shed [JRSB87]. shedding [BWS86]. sheep [MvdMH^{+}88, PTW^{+}85]. sheets [MCC^{+}89]. shift [AK87, AE88]. shifting [RFMR86]. Shiga [SOB^{+}89]. Shigella [SOB^{+}89]. shock [BGT^{+}89, CRR87a, CL85a, CS86b, EMT88, HPO88, KLHK88, LCL^{+}89, MT86, MP89, MLPF89, MWSF87, MW88b, MPB88, NSY86, NCB88, PBK88, RL89, SCK^{+}89, Sch86a, SAA89, SGBF87, VKB89, WS85, WM88, WSRT89, dJHMB86]. shock-induced [MWSF87]. shock-related [MPB88]. short [AIJ^{+}89, BB87d, MMM^{+}89, SL85a]. show [CDMC88, MH86a]. showing [HTT89, QS87]. shown [SSJ^{+}85]. shows [ABG^{+}89, GCC^{+}89, HC86, SR87]. shuttles [FTLSS88]. sialic [Bis86]. sialoglycoprotein [Rog87a]. sialoprotein [HHD^{+}86, OJHAH89]. sialylated [YBBS87]. side [GHW^{+}88, HvdSdC^{+}89, SFB^{+}88]. sides [YIJ88]. Signal [CRJM88, DBM^{+}89, Ibr87b, KSJV88, LD86, LD88, WTB^{+}89, WW89, ALWL89, BGM^{+}87, BGS87, BDS87, BDS88, CWW89, EVG^{+}86, FRG^{+}87, GORW88, GKS87b, HPW89, HSWH86, HGL88, LCDRB89, NAS^{+}87, PG87, PL85b, RT88, SEGS85, TLRW86, VB88b, WKBR87]. signal-anchor [HGL88]. signaling [DSGS87, HRBR89, KST^{+}89, KSSAQC89, PMR89]. signalling [BIES89]. Signals [MVK^{+}88, BGM^{+}87, DLF88, GKS88, HFGD89, LT89, LHM89]. silk [TKK^{+}87]. silver [CV87a]. simian [KSH89]. similar [AB88, OB88, SDNWA87]. similarity [BHCD^{+}89]. simple [AF87, BLA^{+}88, GF88, RFV85]. simplex [MBLS87]. Simulation [HG89]. Simultaneous [CJO^{+}89]. simultaneously [SME^{+}89]. Sindbis [TB85b]. Single [Mac89a, AWH^{+}85, CJO^{+}89, CRC^{+}87, ER85, GLD^{+}88, KTO^{+}89, KCAF89, KR86, KS85c, LSC86, NFW89, RH89, VP85]. Single-cell [Mac89a]. sinks [ER85]. sinus [DW86]. sinusoidal [SKU^{+}85, SBC^{+}87]. sis [HD86, HD88b]. sister [LNH86]. Site [BS85, CHG^{+}86, FPS^{+}85, GWM85, TTW88, AAM^{+}87, BFG86, CMHC85, CMP86, DFSK87, EDD^{+}86, GB87a, GDYW86, HMdN^{+}88, HAK^{+}86, Hym89, JFP89, JJP^{+}89, KFH^{+}86, MGWM85, NRL^{+}89b, PSR^{+}88, SCC87, SR86b, THFB87, VBB85, WW87a]. Site-restricted [CHG^{+}86]. Site-specific [BS85, FPS^{+}85, DFSK87, GDYW86]. sites [AM85, BPB86, Bec86, BBJG89, BPL87, BMB87, Bro85a, Bry88, BBH87, CS89, CBL85, FMT^{+}87, GHW^{+}88, GSS89, GB87b, GFSS86, GSTC87, GPJ88, Gri86, HFK^{+}89, HVBB88, HAK^{+}88, KOB89, KFW^{+}89, KW86b, LH86, MP86c, MTG88, NB89a, PF87, PSP85, PM87a, PSR^{+}88, PMV^{+}89, RGW^{+}89, RMM87, RS85, Rog87a, RRF87, RSR^{+}88, SHN87, Sny89, SW88, SF88b, SH86, VS88a, VSKSP86, WGPW89]. situ [AHC89, BTR^{+}86, BLA^{+}88, DW86, FLH^{+}88, GACOH85, GW86, HNP^{+}86, HHMH87, HKPS89, JMBM85, JSA^{+}89, LR85a, LHK^{+}89, LB89, MWSF87, MAE88, NWE^{+}88, PBT^{+}86, RKWS86, SGH85, SV87, SH87b, SLL89, YR87]. six [LC87, NRL^{+}89b]. sixth [Pip88]. size [HDP^{+}85, SDSO86, SAB^{+}89]. sized [HF85a]. Skelemins [Pri87]. Skeletal [MTK^{+}87, TS89, BBW86, BICFA88, CFC89, CCMB87, CM86b, CLOH87, CHH^{+}88, CS86d, CMGS86, CCS87, FHS^{+}86, FAKVM87, FMT^{+}87, GSS89, HP87a, HZS88, HW86a, HP87b, HMP89, IKSF89, IF85, JSA^{+}89, KFHF85, KBKC86, LL86, MACO88, MYH^{+}85, NFI^{+}87, OH88b, PCC85, RMM87, SV87, SSC86, SCBR85, SS87, SB85b, SBS87a, SMA85, SP86, TOOK^{+}85, TSI^{+}89, VN88, WW88, WTBBW85]. skeleton [AMWN89, AHRM85, AM86, DB85, FCCC^{+}89, FBB^{+}88, GZ89, LDP87, NL85, NV86, NV87, PBD^{+}89, SJS86, WBVL86]. Skin [FLBP^{+}86, BRHR^{+}86, DHK^{+}85, DDB^{+}88, KEG88, LGBE89]. sliding [HF85b, KR88, KS88, MMC88, OK86, Sal86]. slime [WD86a]. Slow [BD85, MCS85b, NFI^{+}87, HGL85, KKT^{+}86, SS87, SGP88, VBWC88]. slows [BLSK88, BM86]. small [APF^{+}88, BBJG89, GSF^{+}88, HGL88, KR86, MHWB86a, MWS85, NJD^{+}86, NDJ^{+}89, OGJ^{+}89, Qua85a, SRH^{+}87, SBG^{+}85, SHR86, WMUP^{+}85, WM88, WG88b, WSRT89]. small-intestinal [GSF^{+}88]. smooth [AS86, BOGK88, BCKJW88, BTO88, BL85, BB87d, CCL^{+}86, CPC^{+}89, CCK^{+}88, DTW^{+}85, FYRK88, FRR85, GFS^{+}89, GVGL85, HBC89, HBF^{+}88, KCAF89, MB85a, MCB85, Maj87, MGD88, MCG^{+}88a, OD87, Owe85, Owe86, OLGT86, OGYK88, ROCK86, RS88, SR89, SL89, SSV^{+}85, SSS85, SC87, SHS^{+}88, SRT^{+}86, SFD86a, TOOK^{+}85, TGNG87, TRT^{+}88, TL87a, TL87b, Try89, TH89b, WR88, WRL86]. SNAP [OHH^{+}89]. SNAP-25 [OHH^{+}89]. snRNP [NHH^{+}86, SSW^{+}85]. snRNPs [GW89]. sodium [AELE88, SB85c, VSZ^{+}87]. soft [LOH^{+}86]. solid [PT88]. solubility [MS86a, YB87]. Solubilization [WLGB85]. solubilized [CE89, CCG89, GLH85]. soluble [ASCM89, DF86a, HPR85, LA87, LBFS88, MB87, MTA86, SSRS88, WCB86, WMBR89]. solution [GC87, SP89]. somatostatin [PS86]. Some [SRWdC87, CMB^{+}86, KFS87, MFM89, dSdSH^{+}88]. somitogenesis [DDH^{+}87, VBWC88]. sort [RYR^{+}89]. Sorting [GSS^{+}88, GGR^{+}86, HFZM87, SGS87, TTF87, vMSvRS87, BRKE88, BHM89a, DHH89, DVHC89, DMM89, GSS87, GZGM87, LLSRB89, Mat86, NBS88b, PFS85, RT88, RFG^{+}85, RGPRB87, TKFH89, vZC87]. sorts [GKH^{+}89]. Source [ER85]. soybean [CV87a, HMHWS86, MVSW86]. SPA2 [Sny89]. space [KBPW89]. spanning [KBKC86, LD86, LFHD89, RG86a]. SPARC [HHMH87, NWE^{+}88, SVF^{+}89]. Spatial [BCKJW88, HMG^{+}86, RM89b, AWJ85, BBP^{+}85, BFRT87, ETA85, HS87b, JS86c, LHK^{+}89, LP86, PN88a, PN88b, YMM86, YKCR88]. spatially [LSB89, NKA^{+}87]. special [HTH85, KFH^{+}86]. specialization [FCMK^{+}86]. specializations [LP85b, VS85]. specialized [GLC88, KSF^{+}88, PF87]. species [CKR^{+}85b, FRR85, GLD^{+}88, KR86, MTK^{+}87, WMGM85, WRL86, WLG85, RKWS86]. Specific [GFSS86, REB^{+}86, RWS^{+}85, BH87, BBV^{+}89, BB88a, BGT^{+}89, BPKS86, BGHvF87, BS85, BK85, CVWZ89, CP87b, CRR87b, CML85, DW89, DFSK87, EOM86, FMJS89, FPS^{+}85, GLL^{+}86, GDYW86, GMCH87, GVGL85, GWI^{+}87, GBU^{+}86, HMBV^{+}88, HE86, HMTF88, HvdSdC^{+}89, HWSK89, HAK^{+}86, HAK^{+}88, IGSGD87, JGAR89, KTP^{+}89, KMP88, KYS^{+}88, LVSC86, LHM89, LLCA89, LL89b, MR87, MNC86, MAK^{+}86, MMM^{+}89, MGWM85, MAC89b, MP86a, MWM89, MWSD86, MNWL85, MIHS86, MOP^{+}86, NRTE89, NJD^{+}86, NZ85, NRL89a, OBCO85, OLGT86, PSK^{+}89, PF85b, PBMB85, PMV^{+}89, RRB86, RKWS86, RE85, RCC^{+}88, RT88, RPW^{+}89, SVSH^{+}88, SSJ^{+}85, SG88a, SSKK89, SHSvdM89, TOOK^{+}85, TSI^{+}89, TT86a, TFS^{+}89, THFB87, Wan85c, WRS^{+}86, WAG^{+}89, WPBB^{+}85, WKBR87, WWN87, WJSB88, ZSB86, Gla85a, GL85, HS87a]. specifically [BF88, GK87b, HKM^{+}87, HW85a, WCPK88]. specificities [HNH^{+}87]. specificity [ABME88, APF^{+}88, GWM85, KFH^{+}86, MH86a]. specified [NBS88b]. specifies [Maj87]. spectrin [ABM86, BKBR88, BM89, Bre89, BDB^{+}87, BHCD^{+}89, CHM^{+}87, DWS88, DBB^{+}87, DBS^{+}89, LL88, MRD^{+}89, MMM87, MNWL85, NL85, PBD^{+}89, PHR86, RZG86, SA85, WSS^{+}89, RZG86]. spectrin-actin [CHM^{+}87, MMM87]. spectrin-binding [ABM86]. spectrin-like [SA85]. spectrin-rich [BKBR88]. spectrins [DBS^{+}89, HSMM85]. spectroscopic [RBJ89]. spectrum [CCC86]. Sperm [CPM86, AK89, AdS85, BW86, BN85a, Bro87, CDL^{+}85, CM86c, CM87a, CMK87, GTG85, Gla85a, Gla85b, HBB^{+}87, LS89, LWM^{+}88, LAA85, LS87b, LBL^{+}85, LS87c, MTG88, PHM85, PHTK87, RO85, SGH85, Sal86, SRR89, THM86, TKP^{+}88, TI85, TI87, TFD87, TES^{+}87, TSOC89, WRS^{+}86, WHI86, WHL^{+}86]. sperm-egg [PHTK87, TSOC89]. sperm-specific [WRS^{+}86]. spermatids [HSK88, LKF87]. spermatocyte [CNL86]. spermatocytes [GK86, SFD^{+}86b]. spermatogenesis [PFD^{+}89, RPW86, WSM86]. spermatogenic [AK89]. spermatozoa [HBB^{+}87, LKF87, VMT^{+}88, WBGV85, WMV86, dCFB86, dLG86b]. Spheroidal [LBO^{+}85]. sphingolipids [LP85a, vMSvRS87]. sphingomyelin [KP89, LP85a]. sphingomyelinase [CKR^{+}85b]. spicule [BBW86]. spike [KH85b]. spinach [DJBD85]. spinal [MOP^{+}85, PBBP85, SMA85]. Spindle [TMH^{+}85, BCD86, CNL86, HHY88, HTH85, HH87, KFA89, Kub87, KB85, LM85a, MHW85, MMC88, Mit89, NG85, RDJ^{+}86, SROB86, SR85a, SR85b, SFD^{+}86b, WMC86]. spindles [CM86a, SM87]. Spiral [SB85a]. splenic [DW86]. spliced [HAK^{+}86, HAK^{+}88, MHP^{+}86, SSW89]. spliceosomes [ZWE89]. splicing [AP86, BKA^{+}89, BCC^{+}87, CBS^{+}89, FCVDH89, MOP^{+}86, SBS87a, ZWE89]. split [GPJ88]. splitting [FY86]. spoke [WMR86, WVCR89]. spokes [GH85b]. sponge [GBU^{+}86]. spontaneous [LHFB88]. spontaneously [BPB^{+}88, HF85b]. spore [PW85, UMHI85]. sporulation [UMHI85]. spots [TM87a, TM87b]. spread [DG88, FLM86, GPSC86, TP89]. spreading [CS88b, DCL^{+}87, GDvdM87, JAT86, MS88b, OW86, RSG87, RVS88, SGRM85, SVF^{+}89, SKS^{+}87, YBWLM88]. sprouting [CMNB88, EWAP85, GAH86, HP85b]. squamous [AF87, BHSJ86, ZBG89]. squid [AWH^{+}85, GS89, KTT^{+}86, ML85, SH87a, SBK85, TTKM86, TTM88]. squirrel [VS85]. src [GH88, RE85]. Stability [DMC85, BAG86, BKK87, CM86b, HJS^{+}85, HB88c, HP87b, HK88b, JPM85, KGB88, LL89a, LSB89, LGF86, MWSF87, NV87, SABK87, WG85]. stabilization [PN88a, PN88b, PN89, SRW88]. stabilize [GTG85, TH89b]. stabilized [GD89]. Stable [KSFL89, Pra86, BKK87, ES89, GKB89, HFB87, SK87, TWB88]. stacked [WBDR88]. Stage [MWSD86, DS87, MER^{+}86]. Stage-specific [MWSD86]. Stages [GB86a, BS88b, CDMC88, CRLW86, DRSG85, KTF87b, KMP88, MSV^{+}88, MMK^{+}89, PB87, SME^{+}89, TM87a]. stain [GMF85, PWS^{+}89]. staining [CV87a, HLL88, NNJ^{+}87]. staircase [TTC88]. stamen [Hep85, HC87]. Staphylococcus [MA85]. starfish [PCBD89]. starting [FONW88]. starts [MSR^{+}86]. state [ALCG88, CH87, CS86b, DF86a, FC86, GF86a, KDB87, KMK86, LLSRB89, MCG^{+}88a, PZBD87, PRS^{+}88, SHS^{+}88, Sow86, TRC^{+}87, VSKSP86]. state-dependent [SHS^{+}88]. States [LMSM89, Heu86, RSR^{+}88, TH89b, UPRS88, UTK88]. statin [Wan85c]. stationary [DNI^{+}88, FMG89, GSB87, LN88, NL86]. Steady [LLSRB89, ALCG88, GF86a, KMK86, PRS^{+}88]. Steady-state [LLSRB89, ALCG88, KMK86, PRS^{+}88]. stem [BJMK89, DG89, SGS86]. step [ALWL89, CA86, RFSL88, SHK^{+}87]. Stephanopyxis [WMC86]. Steps [SMS^{+}88, PGH^{+}88, SSP89, WW87b]. stereocilia [TT88a, TTC88, TTS^{+}89]. stereociliary [TTC88]. stereological [SCOE^{+}86]. sterile [DM87]. sternal [GF85]. steroid [GZO88, OJHAH89]. steroid-inducible [GZO88]. steroidogenesis [PSBN85]. sterol [JBG^{+}87, KSC^{+}89]. sterol-enhanced [JBG^{+}87]. stiff [HLW^{+}89]. stiffen [LYE87]. stilbene [VS88b]. stimulate [DWL86, SJ87, SYS85]. stimulated [AC86b, DRME89, Eva89, FSM88, GWG85, IF89, LR85a, LSDP86, PUW^{+}88, RBKK87, SSY^{+}86, SBS87b, WCB86, WBW89, GGA^{+}89, MBK^{+}86]. stimulates [FTC87, MMM87]. stimulating [BHSJ86, FOW^{+}85, MPB88, OGF87]. stimulation [BSB^{+}86, DCR89, DF87, ETT88, HWDC89, HD88b, Mac89a, MMT85, MRD^{+}89, NRTE89, NH87, PRS^{+}88, RBG85, RT86, SNNK85, YMM88]. stimulator [GL85]. Stimulus [CJ86, GMCH87, CGA89, CHG87, OS88b]. stimulus-induced [CGA89]. Stimulus-secretion [CJ86, CHG87]. Stimulus-specific [GMCH87]. stochastic [TLZ88]. stomatitis [BWK87, BF88, DKHB87, DRM^{+}88, FGW85, GB85a, MKK89, OMY^{+}85, PFS85, PMR86, SHP^{+}87, SBG89]. stop [NBS88b]. stop-transfer [NBS88b]. storage [EWHP87, POCM88, SS88c, VSSRB87, vZC87]. stored [BIES89, CS88d, MSY^{+}87, RH87, SHR86]. stores [BFJ^{+}85]. strain [AW88, KYS^{+}88, WDR89]. strains [FS86b, SAGM88, WD86a, ZP85]. strands [VP85]. stratification [LBB^{+}88]. stratification-related [LBB^{+}88]. stratified [BLA^{+}88, KOF88, ZBG89]. stratify [AW88]. streaks [DNI^{+}88]. streaming [KR88]. strength [YB87]. strengthening [LBEM89]. streptavidin [HVBB88]. streptavidin-gold [HVBB88]. Stress [DW86, AT86a, ATNH86, CHLS88, DH89b, FBP89, LMD^{+}86, TIT89, WS85, WS86b, WF86]. stress-inducible [FBP89]. stretched [SP86]. striated [BMDN87, JSCLP89, MCS85a, Pri87, SL89]. striped [PCP85]. strips [DB85]. stroma [BFBL88]. stromal [HvdSdC^{+}89]. stromelysin [VKB89, WHMA86, WTB^{+}89]. Strongylocentrotus [OHK86, SB85a]. Structural [BICFA88, CCA^{+}88, CCL^{+}86, CPKJ87, ETT88, EBGS86, FC89b, LWT85, LFHD89, MF87a, SDS87, VC85, ZKR^{+}88, AM86, CAG89, DMF^{+}87, DBS^{+}89, EHC^{+}85, FC86, FR86, HERO^{+}87, Heu86, KL89b, KBK86, MWSF87, SKMB86, Ste85, SE86, TAHC^{+}89, TTM88, WFR86, WLW^{+}88, ZHES89]. structurally [AE87, FCH^{+}88, MSLR^{+}86, NKMK86, ZSSR86]. Structure [BGH87b, FTLSS88, GACOH85, HS85a, KW88, MRWJ88, PHRM87, PBH^{+}87, RPRS89, AED^{+}89, AY89, Ake89, AHRM85, ABM86, BSA87, BKBR88, BFA88, BDS85, BKR86, Bro85a, Bro85b, CHFA89, DT89, DRM^{+}88, DW86, DCB^{+}88, EMOB85, EMW^{+}89, EMK85, FAKVM87, FCH^{+}87, GDS87, HFGD89, HLW^{+}89, HSO88, HSK^{+}89a, HKF^{+}87, JM86, JYG86, KLS85, KMKS87, KTP87, LCBS89, LH86, LNH86, LSF86, LPTL86, LU88, MD85, MMS89, MSS87c, MVB87, PCP85, QCKR85, RR88, RBJ89, SH87a, SAGM88, SQM89, SRW88, TH89a, VA85, WSS^{+}89, WW87a, WCSR85, WSLN86, WBB89, YR87]. structure-function [MD85]. structures [AST89, ATNH86, EBCO88, GLK^{+}89, HMP85, HW85b, KR86, KKML88, MWK89, MG85, PWS^{+}89, RBGF89, SB85b, SW87b, SD88, TCF86, Wan85b, WAF87]. studied [HH86, JIFH85, Kro86, LWT85, Rob87]. Studies [ETA85, GDYW86, SS88a, SR86a, WGS^{+}88, WRL86, WSLN86, BR85, BCK85, CMP86, EBGS86, FS86a, GS85, HRP87, JYG86, KDFG89, NTNT85, PP89, RBMM88, RC85a, RC85b, SHP85a, SHP85b, TM87a, TM87b, Tok89, VB88a, VG86a, ZG87]. Study [BPKS86, ACPFB^{+}85, BBV^{+}89, BO85, CMHC85, CTYF85, CKR^{+}85b, DD88, FAKVM87, GM85a, GDF86, GHL85, HSK88, IMP^{+}87, JS86a, JS86b, JT85, JS86d, JSA^{+}89, KMR^{+}88, KDB87, OH89, RLP85, SCOE^{+}86, TMS85, TB85b, TNG89, TCP^{+}85, WS85, dVOD^{+}89]. studying [LHNF87]. Subaxolemmal [KTT^{+}86, TTKM86]. Subcellular [BBB88, GWI^{+}87, JSA^{+}89, KSC^{+}89, WWW^{+}85, WRR87, ACPFB^{+}85, BAD^{+}89, BFSH87, CMHC85, DRH^{+}87, GKZ89, GDF86, GW89, GH88, GKB87, Hol89a, KDB87, LMW^{+}86, MDW86, NMKS85, PFD^{+}89, SS85]. subclass [DBM^{+}89, HNH^{+}87]. subclasses [JPM85]. subclones [HSMCL87]. subcompartment [VVL^{+}89]. subcompartments [RFH85]. subfamilies [ESA86]. subfraction [EWAP85]. subfractions [SPF87]. subfragment [GRvdM89, QS87]. subfragment-2 [QS87]. subfragments [JMW88]. subjected [CRR87a]. submitochondrial [PMV^{+}89]. subnuclear [CA87]. Subpellicular [SSS^{+}87a]. subplate [CS88c]. subpopulation [PPG85, TBK86]. Subpopulations [KHKR86, HM89b, OHH^{+}89, SFK^{+}89, ZRW85]. subsarcolemmal [SB85b]. subsequent [LBEM89, WG85]. subset [HKM^{+}87, MSS^{+}89]. substances [OCP86]. substitute [PMR86]. substitution [DH87]. substitutions [SEGS85, SSK89]. substrata [LEW^{+}88]. substrate [APBP87, AEB^{+}88, BPB86, Bec86, BBB88, BSPdSG88, CC88, CHMB86, CK86, FHE88, GCBH86, HPR85, HDG^{+}85, JB86, SSK^{+}88, TRM88, VHO^{+}89]. substrate-associated [CK86]. substrates [FM86, GZ89, GPSC86, IS88, PJ87, PSS89, dLG86b]. substratum [CSG85, CG86a, DKB^{+}85, Gri86, HNR87, LS87a, MS88b]. substratum-bound [Gri86]. Substructure [GH85b, VC89, SGH85]. substructures [FKP86]. subsynaptic [WTB87]. subtypes [RZG86]. Subunit [TL87b, WCSR85, AGPT88, BB88a, BGH^{+}87a, BSLS^{+}89, BDB^{+}87, CCA^{+}88, CAG89, CPG^{+}89, DSGS87, EVG^{+}86, FBF^{+}89, FKC87, FC89a, GKR87, HPW89, HERO^{+}87, HMP86, HKTY89, HKPS89, JYC87, LWSF85, LCBS89, LN88, LWM^{+}88, MH88, MSQ^{+}89, MB88, MWS85, MG89, MCA^{+}89, Pip88, RPW^{+}89, RGBM87, SF88a, SBS87a, Smi87, SBG^{+}85, SHR86, SBC^{+}87, TH89a, TKK^{+}87, TLC85]. subunits [ALL89, CP87b, CMR^{+}87, FCH^{+}88, GWM85, KRLU87, KSKW^{+}85, MSC^{+}89, MC89a, MG89, PFD^{+}89, SGM^{+}89, SHR86, UTK88, Wan85d, WC87, ZG87]. successive [SSP89]. such [TIT89]. suckling [SGB^{+}86, WWN87]. sucrase [TRC^{+}87]. sucrase-isomaltase [TRC^{+}87]. sucrose [ZSE85]. sufficient [RGPRB87, SSY^{+}86, THM86]. suggests [KEG88, SGS86, WW88]. suicide [NFN87]. sulfate [BMGI85, BFL88, CS88a, CE89, CW87, CSG85, Cou87, DLH^{+}89, FC86, FRR85, FM86, GSS89, GLK^{+}89, HvdSdC^{+}89, JNR^{+}85, JRB88, KKSK88, LCR^{+}87, MAC89b, MOT^{+}85, PJ87, Rap89, RFH85, RBG85, SMSR88, SSC86, STE88, SRH^{+}87, SABVK89, SKU^{+}85, SF87, VVL^{+}89, ZRW85, CCL^{+}86]. Sulfated [KG85b, KW86a, RFZH85, RHL^{+}85, UPL^{+}87]. sulfation [BH87, FFH88]. sulfur [FH86]. superfamily [GCR^{+}89, LCBS89, SHC87]. supergene [Ran88]. superoxide [CSGC88, RBKK85]. superprecipitation [HF85b]. supporting [RKA^{+}89]. suppress [CNB89, DF87]. suppresses [RH86a]. Suppression [GAH86, DMG^{+}89]. supralethal [Ros89]. supramolecular [BFKF86]. suramin [MQ89]. Surface [JLB86, OSP^{+}85, AYY86, AAM^{+}87, BAHE86, BS88b, BCPC87, BJ86, CBIO85, CRR87b, CHH^{+}85, CPP^{+}87, CdCP^{+}89, CPM86, CNHF85, CG86c, DH87, DRP87, DRME89, DIF^{+}88, ES89, ELKH88, FMK^{+}86, FFH88, FKH88, GLL^{+}86, GCR^{+}89, GS85, GACOH85, GSF^{+}88, GWI^{+}87, HAB^{+}88, HD86, HD88b, HGS89, HHR^{+}89, HB88a, HL88, HH87, HFB87, HBB^{+}87, IMP^{+}87, JWH85, JNR^{+}85, JRSB87, JRB88, JFAT89, KWW85, KFHF85, KSJV88, KSM89, KDBF86, LLSRB89, LBL^{+}85, LS87c, LG89, MGD88, MSR^{+}86, MvdMH^{+}88, MG86b, MOP^{+}86, MB87, NRL89a, OS88a, OH88b, Oss88, PE85, PCP85, PM87a, PHM85, PHTK87, QCKR85, RH89, RJB86, Rap89, RR88, RBG85, RMP^{+}88, RO85, RIS87, RBKK87, RSP^{+}89b, RSF^{+}85, RRP88, RVS88, SGRM85, SMVS^{+}86, SB88, SL85b, SDS87, SP85, SKAW86, SLG88, SSK^{+}88, SR85c, SYH^{+}86, SPT^{+}89]. surface [SMB^{+}88, SH86, SCD88, WSW85, WG88a, WSH85, WL87, ZH86]. surface-bound [CBIO85, Oss88, SPT^{+}89]. surface-selective [MG86b]. surfaces [AK89, HM89a, MTS89, OW86, PFPM86, SKC85b, TRB86]. surrounding [WBDR88]. survival [SLE86, WMSB^{+}89]. suspensions [MMMP87]. suum [SRR89]. SV [DRD^{+}88, TYP^{+}87]. SV-40 [DRD^{+}88, TYP^{+}87]. SV3T3 [AAM^{+}87]. SV40 [TLC85]. SV5 [NRL89a]. swainsonine [NTNT85]. Swarm [KKSK88]. Swiss [CM88a, DBE^{+}87, DWFT88, FSM88, MSLR^{+}86, TT86a, ZSSR86]. switching [IF89, SFJU86]. SY5Y [MER^{+}86]. symbiotic [PGH^{+}88]. sympathetic [AR88, JPH85, SSR89, SR86a]. synapses [TCP^{+}85, TCK87, WTB87]. synapsin [BBV^{+}89, BGBB89, BBJG89, HD88a, HSK^{+}89a, PM87b]. synaptic [AM85, BBJG89, BFK87, CS86d, GSS89, HTTFC85, JPH85, LMK88, NDJ^{+}89, NSM^{+}87, RMM87, SBK85, SK85, TTGFC85]. synaptogenesis [And86]. Synaptonemal [MHD^{+}87]. Synaptophysin [VJF^{+}88]. synaptosomal [OHH^{+}89]. synaptosomal-associated [OHH^{+}89]. Synchronous [ZP85, RLP85]. syncytiotrophoblast [MBM85]. Syndecan [VJT89, Rap89, SJOB89]. Synergism [BLK^{+}88]. Synergistic [BGM^{+}87, HJS^{+}85, ITY^{+}88]. synergy [HCG87]. synovial [FCM85, JBdlTB87, VKB89]. synthase [CSBS89]. Synthesis [HP87a, LWMM^{+}85, MOP^{+}85, MOT^{+}85, NKMK86, VSR87, BBBM89, BGH^{+}87a, BDS85, BB87d, CSBS89, CMT^{+}88, CR88b, CRT87, CE89, CJK85, CMDCC86, DFF85, DRH^{+}87, DH89b, FBE85, FTC87, GF85, HPO88, HBC89, HOPM^{+}87, HKBT87, HVBB88, HSS89, HBB^{+}87, HFB^{+}89, KG85a, KG85b, KIKS87, KOBLM88, KKSK86, KW86a, LL88, MB85a, MCB85, MSD^{+}88, MW88b, MVB87, NCB88, OO87, Owe85, PN89, Pfe87, PC85, PRS^{+}85, RG88, SR88, SAA89, SFS^{+}87, SBHN86, SVG^{+}86, TS86, TB88a, VN88, VSZ^{+}87, WLG85, WMH^{+}88, ZBG89]. synthesize [GSS89, RYF^{+}87]. synthesized [AC87, BPJBF85, BSLS^{+}89, FGW85, GPSM85, LP85a, NL86, RH87, RT88, RPW^{+}89, TLP^{+}87]. synthetase [MNM87]. Synthetic [SW88, CSK^{+}88, CWH^{+}86a, CMW87, GAPR88, LC87, OW86]. system [BWK87, BN87, Bas87, BFR85, BPC89, BGF88, CRLW86, DCPR^{+}89, DMD^{+}86, DVHC89, EPJBBW86, EBGS86, Fal85, FMG89, GDD88, GM85a, GWCW85, GABT86, GSB85, HMBV^{+}88, HBV85, HAK^{+}88, IKH87, LHNF87, LR89, MS88c, MR85a, MBvS^{+}89, PCS88, PHR86, PFPM86, Ran88, Ris85, RLP^{+}87, SDSO86, SS88a, SMS^{+}88, SPT^{+}89, SG86, WB86, Wil88, YBBS87, vdPdPR86]. system-derived [GDD88]. systems [BFA88].

T [BSLS^{+}89, BSNG85, CP87a, CBYK88, DRD^{+}88, FSM88, KB88, LSDP86, MACO88, PP88a, RR87, RT86, SNC^{+}86, TYP^{+}87, WGPH^{+}89]. T-cell [BSLS^{+}89, RR87]. T-lymphoma [BSNG85]. tadpoles [MWSD86]. tail [DBLS87, HBPK87, PG86, PSP85, RSP89a, ZH86]. tailed [BMGI85]. tails [Sal86]. taken [BFSH87]. Talin [TOB86, BMBL89, JJP^{+}89]. tandemly [GSTC87, SDS89]. TAP [CW87]. TAP-1 [CW87]. taper [Eck87]. tarantula [CPKJ87]. target [HKH^{+}88, ZV89b]. targeted [BCK85]. Targeting [GORW88, MBLK88, BCMK87, DLF88, FWP^{+}87, GKS87b, GKS88, LHM89, LCDRB89, MCD85, MK85c, POCM88, SP88]. tau [BFR85, KTO^{+}89, NLC89, DK86, HLW^{+}89, HSO88, WLW^{+}88]. Tau-crystallin [WLW^{+}88]. Tau-crystallin/alpha-enolase [WLW^{+}88]. Taxol [SC86, CV87b, TCBF^{+}85]. Taxol-dependent [SC86]. taxol-treated [CV87b]. technique [AK87, GH85a, NH87]. techniques [ASF89]. tectum [CR85]. tektin [CP87b, LAA85]. tektins [LS87b]. teleost [BA87, GAB86, TB88b]. telomeres [BG86]. telophase [SM87, TB87]. Temperature [CV87b, CSD89, HGS89, HHR^{+}89, HHY88, MMY88, RIPS85, RFMR86, SFK^{+}89, SB85a, Van87]. temperature-conditional [HGS89]. Temperature-dependent [CV87b]. temperature-sensitive [CSD89, HHY88, MMY88, RIPS85, RFMR86, SFK^{+}89]. Temporal [TTGFC85, AWJ85, BFRT87, ETA85, LHK^{+}89, PN88a, PN88b, VES85]. temporally [HTGP85, Smi87, WAG^{+}89]. ten [FONW88]. Tenascin [AE88, BR89, MTCE87, ACEE87, LBEM89, MHL88, VJT89, LGBE89, GSS89]. Tenascin/ [LGBE89]. tendon [BT86]. Tension [DJS^{+}88, JCBH85, Kol86]. tentans [DLC88]. teratocarcinoma [BGK^{+}89, DG89, GW87]. term [HH86, LDKG87]. terminal [AW88, APF^{+}88, BDS87, BHCD^{+}89, CCK86, CW87, CTSF86, CLOH87, CCS^{+}86, GAH86, HSWH86, HSK^{+}89a, HSMCL87, KCCM85, KTF87a, KTF87b, KWH89, MMY88, MLM85, MSS^{+}89, OH88b, PL85b, RMP^{+}88, SL87b, SSK89, TSI^{+}89, VS88b, WS86c, ZR89, BH86b, CHH^{+}88, PTB^{+}85, POD89]. terminally [SSS^{+}87b]. terminals [BD86a, BD86b, NDJ^{+}89, TCK87]. terminated [GW88a]. terminus [BCMK87, EVG^{+}86, ES86, GKS87b, GKS88, HGL88, LD86]. test [SGB87]. testicular [EHM86, HBSQ^{+}85, STF85]. testing [HG89]. testis [AK89, KMP88, LC88b]. testis-specific [KMP88]. tetanus [CNHF85]. tetradecanoylphorbol [SPO85]. tetrahedral [HKA^{+}87]. Tetrahymena [BG86, CRC^{+}87, Den88, LLCA89, MRWJ88, PFTV89, RCC^{+}88, RSR^{+}88, SCRA87, SD88, Toy87a, Toy87b, VT89, WAG^{+}89]. tetramer [HKA^{+}87]. tetrameric [SDS87]. tetraurelia [MLK^{+}87]. TGF [RYF^{+}87]. TGF-beta [RYF^{+}87]. theca [LKF87]. their [BBV^{+}89, CK86, CCK^{+}88, DCC^{+}88, DMC85, ETT88, GWM85, GSB87, GLC88, HWDC89, HB87, HJV^{+}89, JYG86, JSCLP89, KKGJ89, KTT^{+}86, LLC85, LN88, LBW89, LOH^{+}86, MS86b, MBL89, PSBN85, PG86, RDJ^{+}86, SLL89, TTBG87, TRC^{+}87, WFE89, WSH85, WMC86, ZV89b, dGRS87]. thermophila [PFTV89]. thermotolerance [NB88b]. thermotolerant [MW88b, WM88]. these [HDP^{+}85, MNC86]. thick [CS86a, EBCO88, HP87b, HMP89, KLS85, LCK88, MYH^{+}85, OB88, SDMF85, SKL85, SP86, TB89a, WE87]. thin [FW85, IF85, NFW89]. third [NMM^{+}86]. those [SW87a]. though [TT88a]. Three [BDCT88, CD89, HERO^{+}87, HS87a, HS87b, KKSK86, MMS89, SKL85, TRCR89a, TRCR89b, WSW85, YF87, BSA87, BGT^{+}89, BSK89, Bry88, DRM^{+}88, DDBHL86, DS88, HFZM87, KKGJ89, LMW^{+}86, LMK88, MCS85b, MASA89, MSS87c, MG89, Pum89, SHP^{+}87, SSP89, Toy87a, UPRS88]. Three-dimensional [BDCT88, CD89, HS87a, HS87b, MMS89, SKL85, TRCR89a, TRCR89b, YF87, BSA87, BSK89, MASA89, MSS87c]. three-headed [Toy87a]. threonine [WG86]. threonine-654 [WG86]. threshold [OW86]. thrombin [FTC87, HTRBP86, dGRS87]. Thrombomodulin [MBM85]. Thrombospondin [Fra87, MUH89, RYFM89, TRL87, GDO^{+}85, GVD^{+}87, HFK^{+}89, KLHK88, LH86, LWH88, MCB85, MGD88, MUM87, OD88, POD89, RSG87]. Thrombospondin-induced [TRL87]. throughout [TDR^{+}87b, WMC86]. Thy [BSK86, DMW88]. Thy-1 [BSK86, DMW88]. thylakoid [ADS88, KHC^{+}86, LG89, LM85b, SDSO86]. thylakoids [HRR89]. thymocytes [CCP88]. thymus [KTP^{+}89, SIYT86]. Thyone [TI85, TI87]. thyroid [WE87]. thyroid-treated [WE87]. thyrotrophs [RFZH85]. Thyrotropin [BGN^{+}88]. tibial [CMDCC86]. Tight [SAGM88, ASJ^{+}88, AVP^{+}89, FMJS89, FS86b, SG88b, SSMG86, VSSGRB87]. tightly [LWM^{+}88, YKSR89, ZG87]. time [CPS88, MASA89, RF87, TM87b]. time-lapse [MASA89]. times [LFHD89]. timing [DI87, HRBR89]. Tissue [MNWL85, BTR^{+}86, COMS87, CV87a, CBCV85, DRME89, FHS^{+}86, GKZ89, GSTS86, HMBV^{+}88, HvdSdC^{+}89, KVB^{+}86, KNGH^{+}85, MVS88, MOP^{+}86, NADS86, PCS88, PML^{+}89, SC85a, SHBV89, TS88, TRT^{+}88, WPBB^{+}85, WJSB88, ZN89, HS87a, VS89]. tissue- [MVS88]. Tissue-specific [MNWL85, HMBV^{+}88, MOP^{+}86, WPBB^{+}85, WJSB88, HS87a]. tissue-type [COMS87, KNGH^{+}85]. tissues [BKPG89, Cou87, ET87b, GDM^{+}88, GM85c, KL89a, KKSK88, MP88a, MAG88, PL85a, SV87, STH^{+}89, SL85a, SYT^{+}89, SGS^{+}89, TGNG87, TFS^{+}89]. Titin [HDL^{+}86, FONW88, FOW89, HP87b, IKSF89, NFW89, TM87a, TM87b, WGS^{+}88]. titins [HW86a]. tongue [RKWS86]. tooth [VJT89]. Topogenesis [LBFS88]. Topogenic [HDB88]. Topographic [CLC^{+}86]. Topography [FR86, GTB85, MSY^{+}87, WFR86, DRRB88, SSV^{+}85]. Topoisomerase [EHC^{+}85, EH85, MVB87, HE86]. topoisomerases [HKTY89]. Topological [GPJ88, STE88]. topology [DFSK87, YJL^{+}89, CMG86]. Torpedo [BCPC87, CFC89, CCLC89, CPG^{+}89, FMT^{+}87, GDM^{+}88, KCN^{+}86, KCN^{+}89, KRLU87, MMM^{+}85, NSM^{+}87, UTK88, Wal86, WWW^{+}85]. torsional [WDR89]. total [GTB85, LWT85, NG85]. toxicity [PBK88]. toxin [BKN^{+}85, BBB88, CNHF85, DSTJ85, DSGS87, KST^{+}89, KSSAQC89, MA85, MOU88, MDVG^{+}85, NJAS89, SOB^{+}89]. toxin-binding [MOU88]. toxin-sensitive [KST^{+}89]. toxins [AW89]. Toxoplasma [SKAW86]. TP1 [HSK88]. TPA [Owe85]. tract [HBB^{+}87]. tracts [RSL^{+}89]. Tradescantia [Hep85, HC87]. traffic [DCQB88, SN89, SR86b, TT89b]. trafficking [BB89, DCWJ89, GF86a]. trans [BH87, CdCP^{+}89, FFH88, GPSM85, PSM89, SGGS88, TTF87, WDF86, YBBS87, vDSP^{+}88]. trans-Golgi [CdCP^{+}89, FFH88, PSM89, SGGS88, TTF87, vDSP^{+}88]. trans-Golgi-specific [BH87]. trans-part [YBBS87]. transbilayer [DH89a]. transcarbamylase [HFF^{+}85, HKF^{+}87, SHK^{+}87]. Transcellular [HCH85]. transcribed [PW87]. transcript [BYB^{+}88, FBP89, RB85]. Transcription [CL86, EMG86, CW88, CRC^{+}87, FLM86, KW86b, LLCA89, MABJ^{+}86, OGF87, RG87, SHL^{+}89, WSLN86]. transcription-related [CRC^{+}87, LLCA89]. Transcriptional [NR89b, CHS85, HSK88, JHPS86, MVB87, SBS87a]. transcripts [BGT^{+}89, BTR^{+}86, CML85, FC89a, GM88, MNWL85, PRSGD89, RS88, YKCR88, KMM88]. Transcytosis [MWSP87, GFSS86, HM89c, MFKVK87, PSSP88]. transducing [JTBA89]. transduction [CRJM88, DBM^{+}89, KSJV88, SDPZ89, VB88b, WTB^{+}89]. transected [TWB88]. transection [HKBT87]. transepithelial [SGB^{+}86, SAGM88, WLJM^{+}89]. transfected [AF87, BSK89, CW88, CPG^{+}89, CMR^{+}87, DBD^{+}89, GGR^{+}86, GDRB89, KTO^{+}89, NYRK87, YHS88, BLSK88]. Transfection [TYP^{+}87, LHNF87]. transfer [BCK85, CC85b, GML^{+}89, HT86, NBS88b, UP86]. transferase [NAS^{+}87, BSY86]. Transferrin [FS86b, WDF86, ICP^{+}86, MOP^{+}85, MGM87, MDM88, MMRR86, PTW^{+}85, SR85c, SGS87, SGGS88, SGG^{+}89, TL89, TH89c, Wat85, WG88a, WKRH86]. transferrin-gold [WG88a]. transformants [FGO^{+}85]. Transformation [CM86c, DMG^{+}89, MPBZBE89, MQ89, APBP87, CWH^{+}86a, CM87a, GWG85, JS88a, KSFL89, MGWM85, RLB^{+}89, SLL88, TF86]. Transformed [CSR^{+}86, AJAW86, AJCMR88, BSSMF88, BCC^{+}87, DMG^{+}89, FGO^{+}85, GZ89, JVG^{+}85, LHHC85, MSE89, MNM87, MGSG88, MA86, SW87b, TLC85, WBVL86]. Transforming [FTC^{+}89, Mas85, MFB^{+}89, OGYK88, Rap89, RDSL89, ASCM89, AS86, CRF^{+}88, CWF^{+}89, DNB88, GGR89, HMF^{+}87, KOBLM88, LSAKO86, LKOM88, MPT88, Maj87, MTA86, NV89, OSH^{+}86, PML^{+}89, RR88, RYF^{+}87, SMR87, SR89, SRWdC87, TFS^{+}89, TTLR89, WSM^{+}87, WMH^{+}88]. Transgenic [PSK^{+}89, CSH89, Gor89, TC87, TNG89]. transglutaminase [BS88b]. Transient [SDSO86, Bas87, Ibr87b, MMT85, NRL89a, TB89b, WW89]. transients [Hep89]. Transit [WK85, BPKS86, MWS85, WSH85]. transition [BTV^{+}89, HSK88, OMM87, TB85a, WOP^{+}88]. transitions [WTBBW85]. Translatable [TLC85]. translated [FBP89, GLD^{+}88]. Translation [HW85a, GW88a, HSS89, KMM88, Koz89, VN88, WSLN86]. Translational [TT86a, WFE89, HSK88, JHPS86]. translationally [SBG^{+}85]. translocate [GS86a]. translocated [BDS87, CCG89, GW88a, HW88]. translocates [PSV87, SF88a]. translocating [FRG^{+}87, KBR87]. Translocation [BGS87, DF88b, ISL87, MK87, RGW^{+}89, SPL87, BGL^{+}88, CG86b, DS89, ES86, FRG^{+}87, GORW88, GS89, HAM86, Ibr87a, Ibr87b, ITY^{+}88, KJW89, LP85a, MK85b, NB89b, PL85b, PMV^{+}89, SS85, SM89b, SHN87, SW85, SSK89, VT89, WCB86, WB86]. Transmembrane [DSGS87, AHM89, BSNG85, BK85, DH87, DSG86, FY86, KB88, LBFS88, MR87, RDSG86, SH87a, SLG88]. transmembranous [WGC85]. transmitter [RRF87]. transmitting [ABC^{+}88, SSZ88]. transplantable [MLR^{+}86]. transplantation [SP85]. Transport [DVHC89, HdAC89, HLU^{+}87, KS85b, MN87, AWH^{+}85, BSR^{+}89, BWK87, BMR85, BAHE86, BMEG89, CC89, CDB^{+}86, CdCP^{+}89, CG86c, DH87, DFdN^{+}89, DKHB87, DRM^{+}88, DRSG85, DIF^{+}88, FDG88, FNPF87, FFH88, GHW^{+}88, GKR87, GZGM87, HD86, HSWH86, HSB^{+}85, HGL85, HBV85, HTGP85, Hol89b, HCH85, JS86d, JS88a, KS85a, KS85c, KP89, LRF89, LKHR87, MKK89, MP86d, MW88a, ML85, MK85c, NBS88a, NM89, NFF86, NRL89a, PN89, PBvTS88, PHRM87, PAF88, PMR86, RH86b, RKFN88, SEGS85, SFJU86, SJ86, SJ87, SBK85, SGB^{+}86, SRM89, SMB^{+}88, SS89d, SPO85, TJ85, TT86b, TSM87, VFB^{+}89, VSZ^{+}87, WLJM^{+}89, WSH85]. transported [CGD86, FGW85, HGL85, HB87, RPW86, TC87, YMT85]. transporter [CML85, CSL87, SHL^{+}89, WBVL86]. transporters [BGL^{+}88, SNSS87, TL89]. Transposition [PHL89]. transverse [BICFA88, BMDN87, DMF^{+}87, HGL85]. trapping [PSM89]. treadmilling [FJMW87, RGM85, WS86a, Wan85d]. treated [BS85, CV87b, DKKK88, KPK86, LPR^{+}89, LEG^{+}89, NT85, PBBP85, WE87].

Treatment [ODD85, BR85, CKR^{+}85b, HMP85, MDM88, PAF88, WS85]. triads [FAKVM87, KBKC86]. trichocysts [STGK87]. Trichohyalin [RR86]. trichomes [SSZ88]. trigger [BC89, CNL86, UW86]. triggered [BGHvF87, PE85, TB89b, vGRB86, PW85]. triggering [CKR85a, MHWB86a]. triggers [KHB88, LS89]. Trimeric [HS89]. trimers [BDWH88, CDB^{+}86, DKHB87]. trimming [LBR86]. tripartite [HFGD89]. tripeptide [GKH^{+}89]. triphosphatase [MEA89, Pip88]. triphosphate [DKHB87, MM89]. triskelia [HS89]. trisphosphate [BFJ^{+}85, MSE89, Miy88, PF87, SW86, TJF86]. Triton [AHRM85, Den88, WLGB85]. tRNA [CA87, LHM89]. Trophic [SMA85]. trophoblast [LE86]. tropicalis [SL87b]. Tropomyosin [BBH87, LHHC85, BL85, HLL88, HLL89, JHP85a, LL86, LHL88, MBKR89, MF87a, SCBR85, WGS^{+}88, YMM86, YMM88]. tropomyosin-enriched [LL86]. troponin [MACO88, SCBR85]. troponin-tropomyosin [SCBR85]. truncated [HW88, HGS89, HL88]. Truncation [MMY88]. Trypanosoma [BSPdSG88, Cla87, DIF^{+}88, RSP^{+}89b, SDS89, SSS^{+}87a, SSS^{+}87b, WG88a]. trypanosome [BAHE86]. trypanosomes [GWI^{+}87]. trypsin [WN88]. trypsin-sensitive [WN88]. trypsinogen [BCK85, BCMK87]. tryptic [MTG88]. tube [KS88, PW85]. tubes [QS87]. Tubular [HMP85, SBS87b, HKH^{+}88, KRLU87]. tubule [BICFA88, DMF^{+}87, MCA^{+}89, RSF86]. tubule/sarcoplasmic [BICFA88]. Tubulin [BB87a, SDNWA87, WW87a, BCA^{+}86, BRP88, CDB87, CJK85, CL85a, CL86, DD87, DDR^{+}89, FCH^{+}87, FCH^{+}88, GK85, GGN^{+}86, GHL^{+}89, GMLR85, GB86b, GKB87, HTSH85, HTB88, JYC87, JC89, KGB88, LR85b, LRF88, LP86, LLC85, LC88b, LC87, MGR86, MMC88, MMY88, MGWM85, MND^{+}87, MK85a, MGW86, OH88a, PF85b, PLC87, PC85, QCKR85, SGBG87, SSSP88, SC86, SSS^{+}87a, SWW85, SW87a, SSS^{+}87b, SB85e, Ste85, WSRR89, WVLC86, WFW87, WMGM85, WGBB87]. tubulin-containing [Ste85]. Tubulin-tyrosine [WW87a, SWW85]. tubulins [GLC88, May89]. tubulo [SFB^{+}88]. tubulo-vesicular [SFB^{+}88]. Tumor [SGP88, AHRM85, AM86, BMVB89, BSBZ88, CBS^{+}89, CW88, DWL86, GAPR88, HGMS88, KGRP85, KKSK88, KWPH89, MLR^{+}86, NR89a, Oss88, RRB86, RPW^{+}89, SC88, SVG^{+}86, TRL87, TTLR89, WB85]. tumor-associated [CBS^{+}89]. tumorigenic [RPW^{+}89]. tumors [ET87b]. tunicamycin [FRH85]. tunnel [KJW89]. turn [GB89a]. turnover [DAJJ^{+}88, OH89, PBH^{+}87, QCKR85, SGB87, SE85, WGBB87]. turnovers [YBBS87]. turris [WMC86]. twelve [Kam88]. twist [SD87]. Two [CS88b, CHG87, FL88, FCL86, GB87b, GDvdM87, HGMS88, LU88, YBBS87, ZRW85, ALWL89, BSG89, BPL87, BMB87, BNS85, BBH87, CDMC88, CM86b, CMW^{+}85, CA86, DCWJ89, DK88, ETT88, GTP87, GPJ88, GCPE85, GLD^{+}88, HPO88, IS88, JGYG85, JHP85b, KSEP86, KGHK^{+}87, LCE87, MLPF89, MG86a, MPB88, NKMK86, PB87, PGH^{+}88, RZG86, Rob89, RLP^{+}87, RHL^{+}85, RYR^{+}89, SME^{+}89, SCBR85, SDS89, SCRA87, SAGM88, SHK^{+}87, TLRW86, TMP87, Toy87a, VKB89, WW87a, WSH85, ZR89, dCFB86, vZC87, KSSAQC89]. Two-dimensional [LU88]. two-domain [WW87a]. two-headed [Toy87a]. two-step [ALWL89, CA86, SHK^{+}87]. Type [AS86, BPE^{+}86, CMDCC86, GF85, KSF^{+}88, KSBB87, KSL^{+}87, LPR^{+}89, SKMB86, ZMH89, AGPT88, AT86b, BFBL88, BVD87, BLA^{+}88, CSM^{+}89, COMS87, CTYF85, CTSF86, CL85b, CML85, DCWJ89, DRP87, EHK86, FJY86, GKS87a, GKF85, GGJ^{+}87, HDB88, HMTF88, HRP87, HAK^{+}86, HAK^{+}88, ISM85, JLK89, KEG88, KFH^{+}86, KNGH^{+}85, KDFG89, LSAKO86, LHK^{+}89, LWM^{+}88, LD88, LRF89, Maj87, MMR88, MVS88, MGHG^{+}86, MTA86, NRL89a, NV89, OSH^{+}86, PVMO87, PFPM86, PSS^{+}87, PHN^{+}88, PP89, PRS^{+}88, Rap89, RYF^{+}87, RG86b, RBAF89, STH^{+}89, SL85a, SFK^{+}89, SGS^{+}89, TDR88, TNG89, TLC85, TC86, VMH^{+}88, YM87a, YM87b, YR87]. type-specific [HAK^{+}86]. types [AE87, BD86a, CRLW86, CKF85, GVD^{+}87, GSTS86, HNP^{+}86, JGYG85, KKSK86, MMMP87, MEBV^{+}89, MCS85b, PNB89, SV87, STH^{+}89, WGA85]. typhimurium [JGAR89]. tyrosinated [GGN^{+}86, GB86b, WFW87]. Tyrosine [BH87, MP88a, AETT86, BJMK89, CHMB86, DAJJ^{+}88, FFH88, GZ89, GCBH86, KC88, MP89, NGS85, RFZH85, RHL^{+}85, ST86, SWW85, STS^{+}89, SHL^{+}89, WW87a]. tyrosine-sulfated [RFZH85, RHL^{+}85]. tyrosinolated [SSS^{+}87b].

U [GKS87a, SSW^{+}85]. U-937 [GKS87a]. U-snRNP [SSW^{+}85]. U1 [NKA^{+}87, PW87]. U937 [JNA85, PKN^{+}89, PBK88]. ubiquitin [CRR87a, CR87]. ubiquitous [NPCL85]. Ultrastructural [AMG85, CHLS88, GGN^{+}86, JSC85, KVB^{+}86, PAHC89, PAGPLM86, PHN^{+}88, SLL89, BKBR88, CM86a, GCK^{+}88, HS86a, KFW^{+}89, OB88, PW86, WFF86, BD86a]. Ultrastructure [KEG88, SIR^{+}88, SJS86, EMT88, TAMR85]. ultrathin [BTR^{+}86]. ultraviolet [WIS89]. ultraviolet-microbeam [WIS89]. umbilical [EWHP87, NKMK86]. unassembled [ALL89, GHL85]. unchanged [Pfe87]. uncoating [HS89]. uncoupled [ZKR^{+}88]. uncoupling [LBFS88, MMC88]. undercoat [TIT89]. undergo [AW88]. underlies [CMW87]. underlying [RSF^{+}85]. Unequal [LL88]. unfertilized [OHK86]. unicellular [DB85, DRRB88]. unidentified [BGM^{+}87]. Unidirectional [HF85b]. uniform [SE85]. uniformity [CMM^{+}85]. Unique [OD88, APF^{+}88, FC86, HTT89, HH87, PKS^{+}89, PML^{+}89, PRSGD89, SKK^{+}85, SRR89, WC87]. uniquely [Wan85a]. unit [RG87]. unregulated [BA87, MA86]. unrelated [NLC89]. unstimulated [AC87]. untransformed [MA86]. untranslated [HFK^{+}89, SHR86]. Unusual [FBP89, KBK86, ROCK86, AW88, BM89, CHH^{+}88, SCRA87, TB88b, WM85]. update [Koz89]. upon [CSG85, GBBS89, HLW^{+}89, MPT88, MRD^{+}89, PN88a, PN88b, SNNK85, SF89, TRT^{+}88, VC85, VGKF86, Wan85c]. upper [LOH^{+}86]. upstream [MACO88]. Uptake [NMKS85, PSBN85, PBMB85, SGB^{+}86, BD86b, DLF88, MF87b, OHK86, PBS85, PMJ87, TKK86, TAMR85]. urchin [BBW86, BN85a, CWZ89, CDL^{+}85, ER85, FCCC^{+}89, GW89, HBB^{+}89, JWH85, KH85a, KB85, LAA85, MHK^{+}85, MTG88, NKA^{+}87, OHK86, RFSL88, SGH85, Sal86, SB85a, SB85c, SR85b, SBG^{+}85, SW86, SE86, TES^{+}87, TJF86, VMT^{+}88, WMV86, WKS85, YB87, ZSE85]. urinary [MVS88]. urinary-type [MVS88]. Urokinase [BVD87, PVMO87, COMS87, FHS^{+}86, GKS87a, HB88a, HBB^{+}87, KOBLM88, Oss88, PKN^{+}89, PFPM86, PSS^{+}87, PHN^{+}88, SPT^{+}89, SVG^{+}86, VBB85]. urokinase-like [FHS^{+}86]. Urokinase-type [BVD87, PVMO87, COMS87, GKS87a, PSS^{+}87, PHN^{+}88]. usage [SGBG87]. Use [SKDF88, TNG89, CHG87, GBC^{+}86, JYG86, KF89b, LGW^{+}89, PRB^{+}89, KL89b]. used [SDPZ89]. Using [BP87a, AK89, AK87, BWK87, BFSH87, BTR^{+}86, BWS86, BPKS86, DFSK87, DSTJ85, EDD^{+}86, FYRK88, FMG89, FH86, FWF88, GB86a, GHL85, HNP^{+}86, HBB85, KKSK88, KSFL89, KYS^{+}88, LVSC86, LC87, MBSMD86, OH88a, SMH88, SLL89, TB85a, WS86a, WLJM^{+}89, ZG87, dVOD^{+}89]. UT [PLA86]. UT-1 [PLA86]. uteri [ZHE85]. uterine [CTJG88, GJD^{+}88]. uterus [ZHE85]. utilization [JC89]. uvomorulin [BBGI85, BMVB89, BVK85, GS86b, GSG88, SSC86]. uvomorulin-like [GS86b]. uvomorulin-mediated [BMVB89].

v [HD86, HD88b, BFBL88, GVD^{+}87, HW86b, TTC88]. v-sis [HD86, HD88b]. VAC [VSSRB88]. vacuolar [BRKE88, RYR^{+}89, SRPS86, TC87, VSSRB88]. vacuole [DVHC89, MN87, PBvTS88, RPRS89, RYR^{+}89, TC87, WBW87]. vacuoles [GW88b, PLS89]. vanadate [KIKS87]. variable [AWJ85, GWI^{+}87, SD87]. variant [BAHE86, DIF^{+}88, GRME^{+}89, PW87, RSP^{+}89b, SCFP86, SYH^{+}86, WG88a, WPW^{+}85]. Variants [GPSC86, BGH^{+}87a, CSR^{+}86, DAJJ^{+}88, ETT88, LN88, SBD^{+}89, SSW89]. variation [HMBV^{+}88, ZH86]. variations [DLF88, SSSP85]. varies [RL89]. variety [PF85b, RHL^{+}85, SSMG86]. various [KKSK88, KTF87b, MMMP87]. vary [TT88a]. vascular [AS86, CCL^{+}86, DTW^{+}85, GFS^{+}89, HBW^{+}87, KVB^{+}86, KDHS87, MB85a, Maj87, MGD88, MCG^{+}88a, MWS85, MG86b, Owe85, Owe86, OLGT86, OGYK88, RKA^{+}89, RS88, SL89, SRB88]. vasoactive [BNF^{+}87]. vector [GDD88]. Vectorial [CTJG88, GB87b, MFKVK87]. vegetative [MSY^{+}87, WDV87]. vein [EWHP87, NKMK86]. veins [MBM85]. ventral [SMA85]. ventricles [KF89a]. ventricular [DBWL87, JSC85, JJ87, UPRS88, ZSB86]. ventricular-specific [ZSB86]. venules [GR89]. Vero [MOU88]. versatility [LC88b]. versus [BTO88, PFRS89, SMR86, VES85, WAF87]. vertebrate [CS89, CFC89, DRC^{+}86, FAKVM87, KF89a, LC87, RGE85, SCLG86, SMH88, SDMF85, SS89c]. vertebrates [MH88]. vertices [HS89]. very [ZP85]. vesicle [BBJG89, BFK87, GS89, GGH89, HW88, HS89, JPH85, KSG85, ML85, PWS^{+}89, RP86, Rob87, Rob89, TTGFC85, TJF86, VFB^{+}89, VB88a]. vesicle-like [PWS^{+}89]. vesicles [BBJG89, BPJBF85, BK85, CJ85, DM87, GSS^{+}88, GS85, GFSS86, GS86a, Gla85a, Gla85b, GGH89, HTTFC85, HHES88, HHS^{+}85, KHKR86, KSS86, LHvF^{+}87, LMK88, MMBA89, MBLK88, NJD^{+}86, NDJ^{+}89, ORA^{+}86, PCP85, Pra86, SBK85, SK85, SPLW89, TL89, TH89c, WN87, WKKS^{+}87, WLGB85, WN88, ZSE85]. vesicular [BWK87, BB89, BF88, DKHB87, DRM^{+}88, FGW85, GB85a, LKHR87, MKK89, OMY^{+}85, PTW^{+}85, PFS85, PMR86, SN89, SFB^{+}88, SHP^{+}87, SKAW86, SBG89, TT86b]. vesiculation [LHFB88]. vessels [CBCV85]. VGVAPG [BSBZ88]. VI [BPE^{+}86, EHK86, KEG88, TLC85]. via [BSK86, CVWZ89, DLG^{+}89, EG89, GWM85, GE88, MN87, MFB^{+}89, PP88a, RH87, SC85a]. viable [Bäc88]. video [Bäc88, GB86a, JIFH85, WOP^{+}88]. video-enhanced [GB86a]. view [AO88]. Vigna [BWL^{+}86]. VII [KSL^{+}87, SKMB86]. VIII [KSF^{+}88]. villin [APF^{+}88, BBH87, DRH^{+}87]. vimentin [CSH89, Eva89, FS86a, GM85a, GWM85, GB87b, LFFW89, MC89a, TMS85]. vimentin-membrane [GWM85]. vincristine [MSC^{+}85]. vincristine-resistant [MSC^{+}85]. Vinculin [BFG86, SB85b, BOGK88, BSG89, DI87, DF86b, HB88a, HP85a, JJP^{+}89, MW86, RHD^{+}85, SC87, SW88, WL86, WRL86, ZBG89]. vinculin-containing [DF86b]. viral [APBP87, BDS85, GFHDD89, GGR^{+}86, PHL89, PMR86, RIPS85, RFG^{+}85, SHP^{+}87, SLL88, WLJM^{+}89, WE86b]. virally [AJCMR88]. virus [Bäc88, BWK87, BF88, CW88, CG86c, CMG86, DH87, DKHB87, DRM^{+}88, DRSG85, FGW85, GB85a, GHW^{+}88, GORW88, GDYW86, GZ89, HDB88, HSWH86, HGS89, HGL88, HBHL^{+}89, KGRP85, KSH89, KH85b, LHHC85, MKK89, MPBZBE89, MBLS87, MSE89, MA86, OMY^{+}85, PP88a, PFS85, PMR86, RE85, RG86a, RDSG86, RGPRB87, SRM89, SBG89, TB85b, WW87b, TTW88]. virus-transformed [GZ89, LHHC85, MSE89, MA86]. viruses [BS85, RIPS85]. visceral [HHD^{+}86]. Viscoelasticity [ZV89a]. Visualization [BCPC87, CS89, DBWL87, HR87, HG88, LDP87, MSS87b, NFW89, BW86, CWOK87, FLM86, FA85, HKA^{+}87, HK88a, HVBB88, MBSMD86, SLL89, Str86]. visualized [Abr85, BWS86, GSB88]. visualizing [WBW87]. vitellogenin [WM85]. vitro [AC86a, AHM89, BSR^{+}89, BB87a, BPMD87, BLR88, BCMK87, CM86a, CMDCC86, CDMC88, CB89, CFB^{+}85, CJ85, CWF^{+}89, DWL86, DDB^{+}88, DCL^{+}87, DSTJ85, DFF85, EBBW87, Eva89, FSFL88, Fal85, FNPF87, FCZ^{+}89, FPS^{+}85, FLH^{+}88, FBE85, FCH^{+}87, GAPR88, Gla85b, GACOH85, GGH89, HBSQ^{+}85, HOPM^{+}87, Hat85, HF85a, HMP85, IF89, KBPW89, KMR^{+}88, KDHS87, KWS^{+}89, LYG85, LL86, LMF^{+}86, LL89b, MTCE87, MMC88, MH86a, MWSF87, MW88a, MTRR89, MK85a, MK85b, MLK^{+}87, MGGP88, NRTE89, NZ85, NTLR88, NFF86, OO87, PCS88, PSV87, RYF^{+}87, RGM86, RGBM87, SVF^{+}89, SF88a, SFLM86, SM89b, SPS89a, STS^{+}89, SR86a, SNC^{+}86, SMA85, SWTB88, TQN^{+}87, TMB85, VH88, VSKSP86, WIS89, WS87, WN88, WHI86, WG88b, YKSR89, ZR87, ZSE85, vdPdPR86]. Vitronectin [BPB86, DWT^{+}89, DCC^{+}88, SSK^{+}88, SSKK89]. vivo [Abr85, BMB86, BFSH87, Bro85b, BFL88, CMP86, CCK^{+}88, CWF^{+}89, DF86a, DT89, DFF85, FDG88, FS86a, FCH^{+}87, GB85a, GFHDD89, GGH89, HMBV^{+}88, HP86, HVBB88, HMP85, HHMH87, JYC87, KMKS87, LCW^{+}89, LC87, MTCE87, MLK^{+}87, Oss88, PBMB85, SGB87, SGS86, SDNWA87, SR86a, SB85e, TMS85, VSKSP86, WGBB87, ZWE89]. vivo-assembled [ZWE89]. VLA [SRW^{+}89, KFW^{+}89, TH89a]. VLA-2 [SRW^{+}89, TH89a]. VLA-2/collagen [TH89a]. VLA-3 [KFW^{+}89]. voltage [AELE88, LSDP86]. voltage-dependent [AELE88]. voltage-gated [LSDP86]. volume [BB85, CM87a, GTB85, GCGR85, HDP^{+}85]. Volvox [ASM^{+}87, EMW^{+}89]. VP7 [PTB^{+}85, PA88, SWC^{+}87]. vs [Mac89a].

wall [BIES89, GH85a, GGMH86, IBSS85, IS88, KDHS87, MSY^{+}87, MAG88, SMC89]. walls [ASM^{+}87]. water [TI85]. wave [BN85b, SW86]. weak [ZBSS^{+}87]. web [KCCM85]. Weibel [EWHP87]. weight [CMT^{+}88, CWZ89, GLW^{+}89, KTT^{+}86, MCS85a, NB88a, PMJ87, SSMG86, TT85]. well [GGR89, WDF86]. Wheat [KdWB^{+}89, LWM85, ADS88, LA87]. whereas [COMS87]. whether [KGRP85]. which [AY89, AW88, GM85c, HKH^{+}88, MGK89, PHL89, PSV87, RG88, RCC^{+}88, SDPZ89, SAGM88, YIJ88]. while [VSSGRB87]. whole [Ris85]. whose [GWG85, LSC86]. widely [CBCV85]. width [SCBR85, TT88a]. wild [FJY86, MMR88, SFK^{+}89, YM87a, YM87b]. wild-type [FJY86, MMR88, SFK^{+}89, YM87a, YM87b]. Willebrand [DLG^{+}89, EWHP87, SMW89, WMUP^{+}85, WMM86]. Wilson [Pol89, Rud88]. withdrawal [UPRS88]. within [DH87, DM87, FLM86, HAB^{+}88, IS88, KRLU87, LM85b, MBKO86, MBK^{+}86, PBBP85, PF87, RFH85, RZG86, SSV^{+}85, TM87b, VVL^{+}89]. without [BBP88, BDS87, DDBHL86, GLC88, HM89c, MSW^{+}86]. wool [FH86]. worm [BN87]. worms [Bas87]. wound [DKH^{+}85, DBE^{+}87, FCVDH89, WG88b]. wounded [MMWD89]. wounds [MHL88]. wt [LC85, MLM85, SNNK85, Wan85a, YLGG85, YSW^{+}85].

X [BMDN87, AHRM85, CMDCC86, Den88, GF85, KDFG89, MMS89, PP89, SL85a, STGK87, WLGB85]. X-100 [AHRM85, WLGB85]. X-114 [Den88]. X-ray [BMDN87, MMS89, STGK87]. Xenopus [AKVF88, BYB^{+}88, BC89, BG86, BRHR^{+}86, BFJ^{+}85, BN85b, BA86, CC89, CG89, CJH85, CS88d, DMC85, DN88, FWP^{+}87, FLBP^{+}86, GK87a, GK87b, GZHM87, GKG88, HW85a, KSH87, LC88a, LCE87, MWSD86, NCB88, NB88b, PJFB89, SP85, SMS^{+}88, SPL87, SF88b, WM85, WG87b]. XI [SHSvdM89, MEBV^{+}89]. XII [SGS^{+}89]. XIII [STH^{+}89]. XK81 [MWSD86]. XLR [GD89].

Y-1 [MMR88, PH89]. Yeast [AB89, BDS87, DMB88, GW88b, SSO89, AB88, BMEG89, BRKE88, BHAS85, BGS87, BDS88, CW89, CA87, CNB89, DS87, DS89, DG88, EVG^{+}86, GMB89, HKTY89, HHES88, HML88, HJV^{+}89, LW86, MMY88, NBS88a, NFN87, PKK89, PBvTS88, RPRS89, RL89, RDS^{+}89, RYR^{+}89, RKFN88, SCK^{+}89, SN89, SM89b, Sny89, SRPS86, TC87, WN87, WCB86, WB86, WBW87, WBDR88]. yields [FBP89, GLK^{+}89]. yolkless [DM87]. young [SSSP85]. ypt1 [BSR^{+}89, BSR^{+}89].

Z [CCMB87, AJRNBJ85, CCMB87, CD89, DCB89, FONW88, LLAS85, MCS85a, MYH^{+}85, MJW86, SCBHS^{+}89, SCBR85, WW88, WRL86, WDR89]. Z-bands [SCBHS^{+}89]. Z-line [SCBR85, CCMB87]. Z-lines [WRL86]. Zeugmatin [MCS85a]. zinc [CDL^{+}85, GD89]. ZO [ASJ^{+}88, AVP^{+}89, FMJS89, SG88b, SSMG86, SAGM88]. ZO-1 [ASJ^{+}88, AVP^{+}89, FMJS89, SG88b, SSMG86, SAGM88]. zona [PHM85]. zone [JPU89]. zones [BCKJW88]. zonula [SSMG86]. ZP3 [LS89]. ZR [NHKE88]. ZR-75-1 [NHKE88]. zygotes [BR85]. zymogen [NRTE89].

References

Albini:1987:CSC

[AAM^{+}87] A. Albini, G. Allavena, A. Melchiori, F. Giancotti, H. Richter, P. M. Comoglio, S. Parodi, G. R. Martin, and G. Tarone. Chemotaxis of 3T3 and SV3T3 cells to fibronectin is mediated through the cell-attachment site in fibronectin and a fibronectin cell surface receptor. Journal of Cell Biology, 105(4):1867–??, October 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/105/4/
1867.

Aris:1988:ICY

[AB88] J. P. Aris and G. Blobel. Identification and characterization of a yeast nucleolar protein that is similar to a rat liver nucleolar protein. Journal of Cell Biology, 107(1):17–??, July 1988. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/107/1/17.

Aris:1989:YNE

[AB89] J. P. Aris and G. Blobel. Yeast nuclear envelope proteins cross react with an antibody against mammalian pore complex proteins. Journal of Cell Biology, 108(6):2059–??, June 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/108/6/2059.

Amchenkova:1988:CME

[ABC^{+}88] A. A. Amchenkova, L. E. Bakeeva, Y. S. Chentsov, V. P. Skulachev, and D. B. Zorov. Coupling membranes as energytransmitting cables. I. Filamentous mitochondria in fibroblasts and mitochondrial clusters in cardiomyocytes. Journal of Cell Biology, 107(2):481–??, August 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/107/2/481.

Andre:1989:DMD

[ABG^{+}89] E. André, M. Brink, G. Gerisch, G. Isenberg, A. Noegel, M. Schleicher, J. E. Segall, and E. Wallraff. A Dictyostelium mutant deficient in severin, an F-actin fragmenting protein, shows normal motility and chemotaxis. Journal of Cell Biology, 108 (3):985–??, March 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/108/3/985.

Aster:1986:LPB

[ABM86] J. C. Aster, G. J. Brewer, and H. Maisel. The 4.1-like proteins of the bovine lens: spectrin-binding proteins closely related in structure to red blood cell protein 4.1. Journal of Cell Biology, 103(1):115–??, July 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/103/1/115.

Altieri:1988:OCA

[ABME88] D. C. Altieri, R. Bader, P. M. Mannucci, and T. S. Edgington. Oligospecificity of the cellular adhesion receptor Mac-1 encompasses an inducible recognition specificity for fibrinogen. Journal of Cell Biology, 107(5):1893–??, November 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/107/5/1893.

Abrahamson:1985:OGB

[Abr85] D. R. Abrahamson. Origin of the glomerular basement membrane visualized after in vivo labeling of laminin in newborn rat kidneys. Journal of Cell Biology, 100(6):1988–??, June 1985. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/100/6/1988.

Aamodt:1986:MMA

[AC86a] E. J. Aamodt and J. G. Culotti. Microtubules and microtubuleassociated proteins from the nematode Caenorhabditis elegans: periodic cross-links connect microtubules in vitro. Journal of Cell Biology, 103(1):23–??, July 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/103/1/23.

Arvan:1986:ISG

[AC86b] P. Arvan and J. D. Castle. Isolated secretion granules from parotid glands of chronically stimulated rats possess an alkaline internal pH and inward-directed H+ pump activity. Journal of Cell Biology, 103(4):1257–??, October 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/103/4/1257.

Arvan:1987:PRN

[AC87] P. Arvan and J. D. Castle. Phasic release of newly synthesized secretory proteins in the unstimulated rat exocrine pancreas. Journal of Cell Biology, 104(2):243–??, February 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/104/2/243.

Aufderheide:1987:EMI

[ACEE87] E. Aufderheide, R. Chiquet-Ehrismann, and P. Ekblom. Epithelial-mesenchymal interactions in the developing kidney lead to expression of tenascin in the mesenchyme. Journal of Cell Biology, 105(1):599–??, July 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/105/1/599.

Adamo:1989:ARC

[ACN^{+}89] S. Adamo, C. Caporale, C. Nervi, R. Ceci, and M. Molinaro. Activity and regulation of calcium-, phospholipid-dependent protein kinase in differentiating chick myogenic cells. Journal of Cell Biology, 108(1):153–??, January 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/108/1/153.

Amar-Costesec:1985:ASM

[ACPFB^{+}85] A. Amar-Costesec, M. Prado-Figueroa, H. Beaufay, J. F. Nagelkerke, and T. J. van Berkel. Analytical study of microsomes and isolated subcellular membranes from rat liver. IX. Nicotinamide adenine dinucleotide glycohydrolase: a plasma membrane enzyme prominently found in Kupffer cells. Journal of Cell Biology, 100(1):189–??, January 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/100/1/189.

Aguas:1985:AMB

[AdS85] A. P. Aguas and P. Pinto da Silva. The acrosomal membrane of boar sperm: a Golgi-derived membrane poor in glycoconjugates. Journal of Cell Biology, 100(2):528–??, February 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/100/2/528.

Allen:1988:BTM

[ADS88] K. D. Allen, M. E. Duysen, and L. A. Staehelin. Biogenesis of thylakoid membranes is controlled by light intensity in the conditional chlorophyll b-deficient CD3 mutant of wheat. Journal of Cell Biology, 107(3):907–??, September 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/107/3/907.

Ardizzi:1987:ILM

[AE87] J. P. Ardizzi and H. F. Epstein. Immunochemical localization of myosin heavy chain isoforms and paramyosin in developmentally and structurally diverse muscle cell types of the nematode Caenorhabditis elegans. Journal of Cell Biology, 105(6): 2763–??, December 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/6/2763.

Aufderheide:1988:TDG

[AE88] E. Aufderheide and P. Ekblom. Tenascin during gut development: appearance in the mesenchyme, shift in molecular forms, and dependence on epithelial-mesenchymal interactions. Journal of Cell Biology, 107(6):2341–??, December 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/107/6/2341.

Alliegro:1988:ENE

[AEB^{+}88] M. C. Alliegro, C. A. Ettensohn, C. A. Burdsal, H. P. Erickson, and D. R. McClay. Echinonectin: a new embryonic substrate adhesion protein. Journal of Cell Biology, 107(6):2319–??, December 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/107/6/
2319.

Abramson:1989:DSD

[AED^{+}89] S. N. Abramson, M. H. Ellisman, T. J. Deerinck, Y. Maulet, M. K. Gentry, B. P. Doctor, and P. Taylor. Differences in structure and distribution of the molecular forms of acetylcholinesterase. Journal of Cell Biology, 108(6):2301–??, June 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/108/6/
2301.

Angelides:1988:DLM

[AELE88] K. J. Angelides, L. W. Elmer, D. Loftus, and E. Elson. Distribution and lateral mobility of voltage-dependent sodium channels in neurons. Journal of Cell Biology, 106(6):1911–??, June 1988. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/106/6/1911.

Acheson:1986:LIB

[AETT86] A. Acheson, D. Edgar, R. Timpl, and H. Thoenen. Laminin increases both levels and activity of tyrosine hydroxylase in calf adrenal chromaffin cells. Journal of Cell Biology, 102(1):151–??, January 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
102/1/151.

Albers:1987:EME

[AF87] K. Albers and E. Fuchs. The expression of mutant epidermal keratin cDNAs transfected in simple epithelial and squamous cell carcinoma lines. Journal of Cell Biology, 105(2):791–??, August 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/105/2/
791.

Albers:1989:EMK

[AF89] K. Albers and E. Fuchs. Expression of mutant keratin cDNAs in epithelial cells reveals possible mechanisms for initiation and assembly of intermediate filaments. Journal of Cell Biology, 108 (4):1477–??, April 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/108/4/1477.

Achler:1989:RMP

[AFMD89] C. Achler, D. Filmer, C. Merte, and D. Drenckhahn. Role of microtubules in polarized delivery of apical membrane proteins to the brush border of the intestinal epithelium. Journal of Cell Biology, 109(1):179–??, July 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/109/1/179.

Alexson:1985:PDP

[AFSL85] S. E. Alexson, Y. Fujiki, H. Shio, and P. B. Lazarow. Partial disassembly of peroxisomes. Journal of Cell Biology, 101(1):294–??, July 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/101/1/
294.

Aletta:1987:SPC

[AG87] J. M. Aletta and L. A. Greene. Sequential phosphorylation of chartin microtubule-associated proteins is regulated by the presence of microtubules. Journal of Cell Biology, 105(1):277–??, July 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/105/1/
277.

Akey:1989:PIT

[AG89] C. W. Akey and D. S. Goldfarb. Protein import through the nuclear pore complex is a multistep process. Journal of Cell Biology, 109(3):971–??, September 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/109/3/971.

Arnaout:1988:AAS

[AGPT88] M. A. Arnaout, S. K. Gupta, M. W. Pierce, and D. G. Tenen. Amino acid sequence of the alpha subunit of human leukocyte adhesion receptor Mo1 (complement receptor type 3). Journal of Cell Biology, 106(6):2153–??, June 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/106/6/2153.

Aamodt:1989:ISL

[AHC89] E. Aamodt, R. Holmgren, and J. Culotti. The isolation and in situ location of adligin: the microtubule cross-linking protein from Caenorhabditis elegans. Journal of Cell Biology, 108 (3):955–??, March 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/108/3/955.

Ahrem:1989:VMA

[AHM89] B. Ahrem, H. K. Hoffschulte, and M. Müller. In vitro membrane assembly of a polytopic, transmembrane protein results in an enzymatically active conformation. Journal of Cell Biology, 108(5):1637–??, May 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/108/5/1637.

Apgar:1985:TXE

[AHRM85] J. R. Apgar, S. H. Herrmann, J. M. Robinson, and M. F. Mescher. Triton X-100 extraction of P815 tumor cells: evidence for a plasma membrane skeleton structure. Journal of Cell Biology, 100(5):1369–??, May 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/100/5/1369.

Abrahamson:1989:SIK

[AIJ^{+}89] D. R. Abrahamson, M. H. Irwin, P. L. St John, E. W. Perry, M. A. Accavitti, L. W. Heck, and J. R. Couchman. Selective immunoreactivities of kidney basement membranes to monoclonal antibodies against laminin: localization of the end of the long arm and the short arms to discrete microdomains. Journal of Cell Biology, 109(6):3477–??, December 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/109/6/3477.

Abken:1986:IHL

[AJAW86] H. Abken, H. Jungfer, W. H. Albert, and K. Willecke. Immortalization of human lymphocytes by fusion with cytoplasts of transformed mouse L cells. Journal of Cell Biology, 103(3): 795–??, September 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/103/3/795.

Anthony:1988:MDV

[AJCMR88] D. T. Anthony, R. J. Jacobs-Cohen, G. Marazzi, and L. L. Rubin. A molecular defect in virally transformed muscle cells that cannot cluster acetylcholine receptors. Journal of Cell Biology, 106(5):1713–??, May 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/106/5/1713.

Arndt-Jovin:1985:ILZ

[AJRNBJ85] D. J. Arndt-Jovin, M. Robert-Nicoud, P. Baurschmidt, and T. M. Jovin. Immunofluorescence localization of Z–DNA in chromosomes: quantitation by scanning microphotometry and computer-assisted image analysis. Journal of Cell Biology, 101 (4):1422–??, October 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/101/4/1422.

Allan:1986:MBP

[AK86] V. J. Allan and T. E. Kreis. A microtubule-binding protein associated with membranes of the Golgi apparatus. Journal of Cell Biology, 103(6):2229–??, December 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/103/6/2229.

Ajioka:1987:CEC

[AK87] R. S. Ajioka and J. Kaplan. Characterization of endocytic compartments using the horseradish peroxidase-diaminobenzidine density shift technique. Journal of Cell Biology, 104(1):77–??, January 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
104/1/77.

Abdullah:1989:IRT

[AK89] M. Abdullah and A. L. Kierszenbaum. Identification of rat testis galactosyl receptor using antibodies to liver asialoglycoprotein receptor: purification and localization on surfaces of spermatogenic cells and sperm. Journal of Cell Biology, 108(2):367–??, February 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/108/2/
367.

Allen:1986:EMR

[AKC^{+}86] R. L. Allen, S. J. Kennel, L. Cacheiro, A. L. Olins, and D. E. Olins. Examination of the macronuclear replication band in Euplotes eurystomus with monoclonal antibodies. Journal of Cell Biology, 102(1):131–??, January 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/102/1/131.

Akey:1989:ISN

[Ake89] C. W. Akey. Interactions and structure of the nuclear pore complex revealed by cryo-electron microscopy. Journal of Cell Biology, 109(3):955–??, September 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/109/3/955.

Ankenbauer:1988:PRA

[AKVF88] T. Ankenbauer, J. A. Kleinschmidt, J. Vandekerckhove, and W. W. Franke. Proteins regulating actin assembly in oogenesis and early embryogenesis of Xenopus laevis: gelsolin is the major cytoplasmic actin-binding protein. Journal of Cell Biology, 107 (4):1489–??, October 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/107/4/1489.

Aletta:1988:NGF

[ALCG88] J. M. Aletta, S. A. Lewis, N. J. Cowan, and L. A. Greene. Nerve growth factor regulates both the phosphorylation and steadystate levels of microtubule-associated protein 1.2 (MAP1.2). Journal of Cell Biology, 106(5):1573–??, May 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/106/5/1573.

Amara:1989:IDU

[ALL89] J. F. Amara, G. Lederkremer, and H. F. Lodish. Intracellular degradation of unassembled asialoglycoprotein receptor subunits: a pre-Golgi, nonlysosomal endoproteolytic cleavage. Journal of Cell Biology, 109(6):3315–??, December 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/109/6/3315.

Andrews:1989:ETS

[ALWL89] D. W. Andrews, L. Lauffer, P. Walter, and V. R. Lingappa. Evidence for a two-step mechanism involved in assembly of functional signal recognition particle receptor. Journal of Cell Biology, 108(3):797–??, March 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/108/3/797.

Anglister:1985:BLD

[AM85] L. Anglister and U. J. McMahan. Basal lamina directs acetylcholinesterase accumulation at synaptic sites in regenerating muscle. Journal of Cell Biology, 101(3):735–??, September 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/101/3/735.

Apgar:1986:AMS

[AM86] J. R. Apgar and M. F. Mescher. Agorins: major structural proteins of the plasma membrane skeleton of P815 tumor cells.

Journal of Cell Biology, 103(2):351–??, August 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/103/2/351.

Arias:1985:UCD

[AMG85] J. A. Arias, A. B. Moser, and S. L. Goldfischer. Ultrastructural and cytochemical demonstration of peroxisomes in cultured fibroblasts from patients with peroxisomal deficiency disorders. Journal of Cell Biology, 100(5):1789–??, May 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/100/5/1789.

Amerongen:1989:MDI

[AMWN89] H. M. Amerongen, J. A. Mack, J. M. Wilson, and M. R. Neutra. Membrane domains of intestinal epithelial cells: distribution of Na+,K+-ATPase and the membrane skeleton in adult rat intestine during fetal development and after epithelial isolation. Journal of Cell Biology, 109(5):2129–??, November 1989. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/109/5/2129.

Anderson:1986:NIR

[And86] M. J. Anderson. Nerve-induced remodeling of muscle basal lamina during synaptogenesis. Journal of Cell Biology, 102 (3):863–??, March 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/102/3/863.

Anderson:1988:VAI

[AO88] R. G. Anderson and L. Orci. A view of acidic intracellular compartments. Journal of Cell Biology, 106(3):539–??, March 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/106/3/
539.

Abrahamson:1986:ESN

[AP86] D. R. Abrahamson and E. W. Perry. Evidence for splicing new basement membrane into old during glomerular development in newborn rat kidneys. Journal of Cell Biology, 103(6): 2489–??, December 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/103/6/2489.

Adams:1987:MFG

[APBP87] S. L. Adams, M. Pacifici, D. Boettiger, and K. M. Pallante. Modulation of fibronectin gene expression in chondrocytes by viral transformation and substrate attachment. Journal of Cell Biology, 105(1):483–??, July 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/105/1/483.

Arpin:1988:SHV

[APF^{+}88] M. Arpin, E. Pringault, J. Finidori, A. Garcia, J. M. Jeltsch, J. Vandekerckhove, and D. Louvard. Sequence of human villin: a large duplicated domain homologous with other actin-severing proteins and a unique small carboxy-terminal domain related to villin specificity. Journal of Cell Biology, 107(5):1759–??, November 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/107/5/
1759.

Antonicek:1987:BFC

[APS87] H. Antonicek, E. Persohn, and M. Schachner. Biochemical and functional characterization of a novel neuron-glia adhesion molecule that is involved in neuronal migration. Journal of Cell Biology, 104(6):1587–??, June 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/104/6/1587.

Azarnia:1985:CAE

[AR85] R. Azarnia and T. R. Russell. Cyclic AMP effects on cell-to-cell junctional membrane permeability during adipocyte differentiation of 3T3-L1 fibroblasts. Journal of Cell Biology, 100 (1):265–??, January 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/100/1/265.

Acheson:1988:NCA

[AR88] A. Acheson and U. Rutishauser. Neural cell adhesion molecule regulates cell contact-mediated changes in choline acetyltransferase activity of embryonic chick sympathetic neurons. Journal of Cell Biology, 106(2):479–??, February 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/106/2/479.

Assoian:1986:TBT

[AS86] R. K. Assoian and M. B. Sporn. Type beta transforming growth factor in human platelets: release during platelet degranulation and action on vascular smooth muscle cells. Journal of Cell Biology, 102(4):1217–??, April 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/102/4/1217.

Argraves:1987:AAS

[ASA^{+}87] W. S. Argraves, S. Suzuki, H. Arai, K. Thompson, M. D. Pierschbacher, and E. Ruoslahti. Amino acid sequence of the human fibronectin receptor. Journal of Cell Biology, 105(3): 1183–??, September 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/3/1183.

Andres:1989:MAS

[ASCM89] J. L. Andres, K. Stanley, S. Cheifetz, and J. Massagué. Membrane-anchored and soluble forms of betaglycan, a polymorphic proteoglycan that binds transforming growth factor-beta. Journal of Cell Biology, 109(6):3137–??, December 1989. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/109/6/3137.

Allen:1989:IMI

[ASF89] R. D. Allen, C. C. Schroeder, and A. K. Fok. An investigation of mitochondrial inner membranes by rapid-freeze deep-etch techniques. Journal of Cell Biology, 108(6):2233–??, June 1989. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/108/6/2233.

Anderson:1988:CZP

[ASJ^{+}88] J. M. Anderson, B. R. Stevenson, L. A. Jesaitis, D. A. Goodenough, and M. S. Mooseker. Characterization of ZO-1, a protein component of the tight junction from mouse liver and Madin–Darby canine kidney cells. Journal of Cell Biology, 106 (4):1141–??, April 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/106/4/1141.

Ayme-Southgate:1989:CGM

[ASLFP89] A. Ayme-Southgate, P. Lasko, C. French, and M. L. Pardue. Characterization of the gene for mp20: a Drosophila muscle protein that is not found in asynchronous oscillatory flight muscle. Journal of Cell Biology, 108(2):521–??, February 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/108/2/521.

Adair:1987:NAC

[ASM^{+}87] W. S. Adair, S. A. Steinmetz, D. M. Mattson, U. W. Goodenough, and J. E. Heuser. Nucleated assembly of Chlamydomonas and Volvox cell walls. Journal of Cell Biology, 105 (5):2373–??, November 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/5/2373.

Angelides:1989:AEI

[AST89] K. J. Angelides, K. E. Smith, and M. Takeda. Assembly and exchange of intermediate filament proteins of neurons: neurofilaments are dynamic structures. Journal of Cell Biology, 108 (4):1495–??, April 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/108/4/1495.

Adlakha:1988:AOM

[ASZ^{+}88] R. C. Adlakha, G. L. Shipley, J. Y. Zhao, K. B. Jones, D. A. Wright, P. N. Rao, and H. W. Sauer. Amphibian oocyte maturation induced by extracts of Physarum polycephalum in mitosis. Journal of Cell Biology, 106(5):1445–??, May 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/106/5/1445.

Amato:1986:PMI

[AT86a] P. A. Amato and D. L. Taylor. Probing the mechanism of incorporation of fluorescently labeled actin into stress fibers. Journal of Cell Biology, 102(3):1074–??, March 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/102/3/1074.

Aumailley:1986:ACB

[AT86b] M. Aumailley and R. Timpl. Attachment of cells to basement membrane collagen type IV. Journal of Cell Biology, 103(4): 1569–??, October 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/103/4/1569.

Antin:1986:RSF

[ATNH86] P. B. Antin, S. Tokunaka, V. T. Nachmias, and H. Holtzer. Role of stress fiber-like structures in assembling nascent myofibrils in myosheets recovering from exposure to ethyl methanesulfonate. Journal of Cell Biology, 102(4):1464–??, April 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/102/4/1464.

Alho:1989:HRP

[AU89] A. M. Alho and C. B. Underhill. The hyaluronate receptor is preferentially expressed on proliferating epithelial cells. Journal of Cell Biology, 108(4):1557–??, April 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/108/4/1557.

Anderson:1989:ZMP

[AVP^{+}89] J. M. Anderson, C. M. Van Itallie, M. D. Peterson, B. R. Stevenson, E. A. Carew, and M. S. Mooseker. ZO-1 mRNA and protein expression during tight junction assembly in Caco-2 cells. Journal of Cell Biology, 109(3):1047–??, September 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/109/3/1047.

Adams:1988:USH

[AW88] J. C. Adams and F. M. Watt. An unusual strain of human keratinocytes which do not stratify or undergo terminal differentiation in culture. Journal of Cell Biology, 107(5):1927–??, November 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/107/5/
1927.

Aktories:1989:ARA

[AW89] K. Aktories and A. Wegner. ADP-ribosylation of actin by clostridial toxins. Journal of Cell Biology, 109(4):1385–??, October 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/109/4/
1385.

Allen:1985:GMB

[AWH^{+}85] R. D. Allen, D. G. Weiss, J. H. Hayden, D. T. Brown, H. Fujiwake, and M. Simpson. Gliding movement of and bidirectional transport along single native microtubules from squid axoplasm: evidence for an active role of microtubules in cytoplasmic transport. Journal of Cell Biology, 100(5):1736–??, May 1985. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/100/5/1736.

Adams:1985:DTS

[AWJ85] G. M. Adams, R. L. Wright, and J. W. Jarvik. Defective temporal and spatial control of flagellar assembly in a mutant of Chlamydomonas reinhardtii with variable flagellar number. Journal of Cell Biology, 100(3):955–??, March 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/100/3/955.

Adachi:1989:HOC

[AY89] Y. Adachi and M. Yanagida. Higher order chromosome structure is affected by cold-sensitive mutations in a Schizosaccharomyces pombe gene crm1+ which encodes a 115-kD protein preferentially localized in the nucleus and its periphery. Journal of Cell Biology, 108(4):1195–??, April 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/108/4/1195.

Akiyama:1989:AFR

[AYCY89] S. K. Akiyama, S. S. Yamada, W. T. Chen, and K. M. Yamada. Analysis of fibronectin receptor function with monoclonal antibodies: roles in cell adhesion, migration, matrix assembly, and cytoskeletal organization. Journal of Cell Biology, 109(2):863–??, August 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/2/863.

Aszalos:1985:DMI

[AYG85] A. Aszalos, G. C. Yang, and M. M. Gottesman. Depolymerization of microtubules increases the motional freedom of molecular probes in cellular plasma membranes. Journal of Cell Biology, 100(5):1357–??, May 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/100/5/1357.

Akiyama:1986:CKA

[AYY86] S. K. Akiyama, S. S. Yamada, and K. M. Yamada. Characterization of a 140-kD avian cell surface antigen as a fibronectinbinding molecule. Journal of Cell Biology, 102(2):442–??, February 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/102/2/
442.

Byers:1986:MPR

[BA86] T. J. Byers and P. B. Armstrong. Membrane protein redistribution during Xenopus first cleavage. Journal of Cell Biology, 102(6):2176–??, June 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/102/6/2176.

Burnside:1987:CIC

[BA87] B. Burnside and N. Ackland. Calcium-independent contraction in lysed cell models of teleost retinal cones: activation by unregulated myosin light chain kinase or high magnesium and loss of cAMP inhibition. Journal of Cell Biology, 105(1):397–??, July 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/105/1/
397.

Bachi:1988:DOB

[Bäc88] T. Bächi. Direct observation of the budding and fusion of an enveloped virus by video microscopy of viable cells. Journal of Cell Biology, 107(5):1689–??, November 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/107/5/1689.

Bologna:1986:OCB

[BAD86] M. Bologna, R. Allen, and R. Dulbecco. Organization of cytokeratin bundles by desmosomes in rat mammary cells. Journal of Cell Biology, 102(2):560–??, February 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/102/2/560.

Bacallao:1989:SOM

[BAD^{+}89] R. Bacallao, C. Antony, C. Dotti, E. Karsenti, E. H. Stelzer, and K. Simons. The subcellular organization of Madin–Darby canine kidney cells during the formation of a polarized epithelium. Journal of Cell Biology, 109(6):2817–??, December 1989. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/109/6/2817.

Black:1986:RMC

[BAG86] M. M. Black, J. M. Aletta, and L. A. Greene. Regulation of microtubule composition and stability during nerve growth factor-promoted neurite outgrowth. Journal of Cell Biology, 103 (2):545–??, August 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/103/2/545.

Bangs:1986:PMI

[BAHE86] J. D. Bangs, N. W. Andrews, G. W. Hart, and P. T. Englund. Posttranslational modification and intracellular transport of a trypanosome variant surface glycoprotein. Journal of Cell Biology, 103(1):255–??, July 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/103/1/255.

Bassot:1987:TIC

[Bas87] J. M. Bassot. A transient intracellular coupling explains the facilitation of responses in the bioluminescent system of scale worms. Journal of Cell Biology, 105(5):2235–??, November 1987. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/105/5/2235.

Brown:1985:PVS

[BB85] P. A. Brown and R. D. Berlin. Packing volume of sedimented microtubules: regulation and potential relationship to an intracellular matrix. Journal of Cell Biology, 101(4):1492–??, October 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/101/4/
1492.

Balczon:1987:TIK

[BB87a] R. D. Balczon and B. R. Brinkley. Tubulin interaction with kinetochore proteins: analysis by in vitro assembly and chemical cross-linking. Journal of Cell Biology, 105(2):855–??, August 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/105/2/
855.

Bamburg:1987:DCL

[BB87b] J. R. Bamburg and D. Bray. Distribution and cellular localization of actin depolymerizing factor. Journal of Cell Biology, 105(6): 2817–??, December 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/6/2817.

Blank:1987:CAI

[BB87c] G. S. Blank and F. M. Brodsky. Clathrin assembly involves a light chain-binding region. Journal of Cell Biology, 105(5): 2011–??, November 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/5/2011.

Brown:1987:EHS

[BB87d] C. C. Brown and G. Balian. Effect of heparin on synthesis of short chain collagen by chondrocytes and smooth muscle cells. Journal of Cell Biology, 105(2):1007–??, August 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/105/2/1007.

Baldwin:1988:ICM

[BB88a] T. J. Baldwin and S. J. Burden. Isolation and characterization of the mouse acetylcholine receptor delta subunit gene: identification of a 148-bp cis-acting region that confers myotube-specific expression. Journal of Cell Biology, 107(6):2271–??, December 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/107/6/
2271.

Brown:1988:ACC

[BB88b] P. D. Brown and P. D. Benya. Alterations in chondrocyte cytoskeletal architecture during phenotypic modulation by retinoic acid and dihydrocytochalasin B-induced reexpression. Journal of Cell Biology, 106(1):171–??, January 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/106/1/171.

Beckers:1989:CGE

[BB89] C. J. Beckers and W. E. Balch. Calcium and GTP: essential components in vesicular trafficking between the endoplasmic reticulum and Golgi apparatus. Journal of Cell Biology, 108(4):1245–??, April 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/108/4/1245.

Bokoch:1988:SLQ

[BBB88] G. M. Bokoch, K. Bickford, and B. P. Bohl. Subcellular localization and quantitation of the major neutrophil pertussis toxin substrate, Gn. Journal of Cell Biology, 106(6):1927–??, June 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/106/6/
1927.

Baas:1989:CMP

[BBB89a] P. W. Baas, M. M. Black, and G. A. Banker. Changes in microtubule polarity orientation during the development of hippocampal neurons in culture. Journal of Cell Biology, 109(6): 3085–??, December 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/6/3085.

Bunge:1989:MSC

[BBB89b] R. P. Bunge, M. B. Bunge, and M. Bates. Movements of the Schwann cell nucleus implicate progression of the inner (axonrelated) Schwann cell process during myelination. Journal of Cell Biology, 109(1):273–??, July 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/109/1/273.

Baribault:1989:EGF

[BBBM89] H. Baribault, R. Blouin, L. Bourgon, and N. Marceau. Epidermal growth factor-induced selective phosphorylation of cultured rat hepatocyte 55-kD cytokeratin before filament reorganization and DNA synthesis. Journal of Cell Biology, 109(4):1665–??, October 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/109/4/
1665.

Behrens:1985:DMD

[BBGI85] J. Behrens, W. Birchmeier, S. L. Goodman, and B. A. Imhof. Dissociation of Madin–Darby canine kidney epithelial cells by the monoclonal antibody anti-arc-1: mechanistic aspects and identification of the antigen as a component related to uvomorulin. Journal of Cell Biology, 101(4):1307–??, October 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/101/4/1307.

Bartles:1985:EED

[BBH85] J. R. Bartles, L. T. Braiterman, and A. L. Hubbard. Endogenous and exogenous domain markers of the rat hepatocyte plasma membrane. Journal of Cell Biology, 100(4):1126–??, April 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/100/4/1126.

Burgess:1987:TDB

[BBH87] D. R. Burgess, K. O. Broschat, and J. M. Hayden. Tropomyosin distinguishes between the two actin-binding sites of villin and affects actin-binding properties of other brush border proteins. Journal of Cell Biology, 104(1):29–??, January 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/104/1/29.

Benfenati:1989:ISS

[BBJG89] F. Benfenati, M. Bähler, R. Jahn, and P. Greengard. Interactions of synapsin I with small synaptic vesicles: distinct sites in synapsin I bind to vesicle phospholipids and vesicle proteins. Journal of Cell Biology, 108(5):1863–??, May 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/108/5/1863.

Bursztajn:1985:IIA

[BBMW85] S. Bursztajn, S. A. Berman, J. L. McManaman, and M. L. Watson. Insertion and internalization of acetylcholine receptors at clustered and diffuse domains on cultured myotubes. Journal of Cell Biology, 101(1):104–??, July 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/101/1/104.

Benson:1985:DIF

[BBP^{+}85] D. M. Benson, J. Bryan, A. L. Plant, A. M. Gotto, and L. C. Smith. Digital imaging fluorescence microscopy: spatial heterogeneity of photobleaching rate constants in individual cells. Journal of Cell Biology, 100(4):1309–??, April 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/100/4/1309.

Benya:1988:MMD

[BBP88] P. D. Benya, P. D. Brown, and S. R. Padilla. Microfilament modification by dihydrocytochalasin B causes retinoic acid-modulated chondrocytes to reexpress the differentiated collagen phenotype without a change in shape. Journal of Cell Biology, 106(1): 161–??, January 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/106/1/161.

Bahler:1989:CSF

[BBV^{+}89] M. Bähler, F. Benfenati, F. Valtorta, A. J. Czernik, and P. Greengard. Characterization of synapsin I fragments produced by cysteine-specific cleavage: a study of their interactions with Factin. Journal of Cell Biology, 108(5):1841–??, May 1989. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/108/5/1841.

Benson:1986:OMS

[BBW86] S. C. Benson, N. C. Benson, and F. Wilt. The organic matrix of the skeletal spicule of sea urchin embryos. Journal of Cell Biology, 102(5):1878–??, May 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/102/5/1878.

Bryan:1985:KAF

[BC85] J. Bryan and L. M. Coluccio. Kinetic analysis of F-actin depolymerization in the presence of platelet gelsolin and gelsolinactin complexes. Journal of Cell Biology, 101(4):1236–??, October 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/101/4/
1236.

Bement:1989:APK

[BC89] W. M. Bement and D. G. Capco. Activators of protein kinase C trigger cortical granule exocytosis, cortical contraction, and cleavage furrow formation in Xenopus laevis oocytes and eggs. Journal of Cell Biology, 108(3):885–??, March 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/108/3/885.

Ball:1986:RMA

[BCA^{+}86] R. L. Ball, D. H. Carney, T. Albrecht, D. J. Asai, and W. C. Thompson. A radiolabeled monoclonal antibody binding assay for cytoskeletal tubulin in cultured cells. Journal of Cell Biology, 103(3):1033–??, September 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/103/3/1033.

Borsi:1987:MAA

[BCC^{+}87] L. Borsi, B. Carnemolla, P. Castellani, C. Rosellini, D. Vecchio, G. Allemanni, S. E. Chang, J. Taylor-Papadimitriou, H. Pande, and L. Zardi. Monoclonal antibodies in the analysis of fibronectin isoforms generated by alternative splicing of mRNA precursors in normal and transformed human cells. Journal of Cell Biology, 104(3):595–??, March 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/104/3/595.

Brady:1986:IKC

[BCD86] R. C. Brady, F. Cabral, and J. R. Dedman. Identification of a 52kD calmodulin-binding protein associated with the mitotic spindle apparatus in mammalian cells. Journal of Cell Biology, 103 (5):1855–??, November 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/103/5/1855.

Burgess:1985:EPT

[BCK85] T. L. Burgess, C. S. Craik, and R. B. Kelly. The exocrine protein trypsinogen is targeted into the secretory granules of an endocrine cell line: studies by gene transfer. Journal of Cell Biology, 101(2):639–??, August 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/101/2/639.

Bennett:1988:SPM

[BCKJW88] J. P. Bennett, R. A. Cross, J. Kendrick-Jones, and A. G. Weeds. Spatial pattern of myosin phosphorylation in contracting smooth muscle cells: evidence for contractile zones. Journal of Cell Biology, 107(6):2623–??, December 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/107/6/2623.

Burgess:1987:VMT

[BCMK87] T. L. Burgess, C. S. Craik, L. Matsuuchi, and R. B. Kelly. In vitro mutagenesis of trypsinogen: role of the amino terminus in intracellular protein targeting to secretory granules. Journal of Cell Biology, 105(2):659–??, August 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/105/2/659.

Bridgman:1987:VCS

[BCPC87] P. C. Bridgman, C. Carr, S. E. Pedersen, and J. B. Cohen. Visualization of the cytoplasmic surface of Torpedo postsynaptic membranes by freeze-etch and immunoelectron microscopy.

Journal of Cell Biology, 105(4):1829–??, October 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/105/4/1829.

Bennett:1985:SPM

[BD85] G. S. Bennett and C. DiLullo. Slow posttranslational modification of a neurofilament protein. Journal of Cell Biology, 100 (5):1799–??, May 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/100/5/1799.

Black:1986:ILBa

[BD86a] J. D. Black and J. O. Dolly. Interaction of 125I-labeled botulinum neurotoxins with nerve terminals. I. Ultrastructural autoradiographic localization and quantitation of distinct membrane acceptors for types A and B on motor nerves. Journal of Cell Biology, 103(2):521–??, August 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/103/2/521.

Black:1986:ILBb

[BD86b] J. D. Black and J. O. Dolly. Interaction of 125I-labeled botulinum neurotoxins with nerve terminals. II. Autoradiographic evidence for its uptake into motor nerves by acceptor-mediated endocytosis. Journal of Cell Biology, 103(2):535–??, August 1986. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/103/2/535.

Bridgman:1989:OMA

[BD89] P. C. Bridgman and M. E. Dailey. The organization of myosin and actin in rapid frozen nerve growth cones. Journal of Cell Biology, 108(1):95–??, January 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/108/1/95.

Byers:1987:DSI

[BDB^{+}87] T. J. Byers, R. Dubreuil, D. Branton, D. P. Kiehart, and L. S. Goldstein. Drosophila spectrin. II. Conserved features of the alpha-subunit are revealed by analysis of cDNA clones and fusion proteins. Journal of Cell Biology, 105(5):2103–??, November 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/105/5/

Bullitt:1988:TDR

[BDCT88] E. S. Bullitt, D. J. DeRosier, L. M. Coluccio, and L. G. Tilney. Three-dimensional reconstruction of an actin bundle. Journal of Cell Biology, 107(2):597–??, August 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/107/2/597.

Bonneau:1985:EVI

[BDS85] A. M. Bonneau, A. Darveau, and N. Sonenberg. Effect of viral infection on host protein synthesis and mRNA association with the cytoplasmic cytoskeletal structure. Journal of Cell Biology, 100(4):1209–??, April 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/100/4/1209.

Blachly-Dyson:1987:YCC

[BDS87] E. Blachly-Dyson and T. H. Stevens. Yeast carboxypeptidase Y can be translocated and glycosylated without its amino-terminal signal sequence. Journal of Cell Biology, 104(5):1183–??, May 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/104/5/
1183.

Bohni:1988:SRS

[BDS88] P. C. Böhni, R. J. Deshaies, and R. W. Schekman. SEC11 is required for signal peptide processing and yeast cell growth. Journal of Cell Biology, 106(4):1035–??, April 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/106/4/1035.

Boulay:1988:POC

[BDWH88] F. Boulay, R. W. Doms, R. G. Webster, and A. Helenius. Posttranslational oligomerization and cooperative acid activation of mixed influenza hemagglutinin trimers. Journal of Cell Biology, 106(3):629–??, March 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/106/3/629.

Beckerle:1986:INP

[Bec86] M. C. Beckerle. Identification of a new protein localized at sites of cell-substrate adhesion. Journal of Cell Biology, 103(5): 1679–??, November 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/103/5/1679.

Baron-Epel:1988:DCC

[BEHJ^{+}88] O. Baron-Epel, D. Hernandez, L. W. Jiang, S. Meiners, and M. Schindler. Dynamic continuity of cytoplasmic and membrane compartments between plant cells. Journal of Cell Biology, 106 (3):715–??, March 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/106/3/715.

Bronner-Fraser:1985:ANC

[BF85] M. Bronner-Fraser. Alterations in neural crest migration by a monoclonal antibody that affects cell adhesion. Journal of Cell Biology, 101(2):610–??, August 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/101/2/610.

Berrios:1986:MHC

[BF86] M. Berrios and P. A. Fisher. A myosin heavy-chain-like polypeptide is associated with the nuclear envelope in higher eukaryotic cells. Journal of Cell Biology, 103(3):711–??, September 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/103/3/
711.

Bloch:1987:RPP

[BF87] R. J. Bloch and S. C. Froehner. The relationship of the postsynaptic 43K protein to acetylcholine receptors in receptor clusters isolated from cultured rat myotubes. Journal of Cell Biology, 104(3):645–??, March 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/104/3/645.

Bergmann:1988:MPV

[BF88] J. E. Bergmann and P. J. Fusco. The M protein of vesicular stomatitis virus associates specifically with the basolateral membranes of polarized epithelial cells independently of the G protein. Journal of Cell Biology, 107(5):1707–??, November 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/107/5/1707.

Block:1988:SMS

[BFA88] B. A. Block and C. Franzini-Armstrong. The structure of the membrane systems in a novel muscle cell modified for heat production. Journal of Cell Biology, 107(3):1099–??, September 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/107/3/
1099.

Bard:1987:RCD

[BFAI87] F. Bard, C. Franzini-Armstrong, and W. Ip. Rigor crossbridges are double-headed in fast muscle from crayfish. Journal of Cell Biology, 105(5):2225–??, November 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/105/5/2225.

Birk:1988:CTT

[BFBL88] D. E. Birk, J. M. Fitch, J. P. Babiarz, and T. F. Linsenmayer. Collagen type I and type V are present in the same fibril in the avian corneal stroma. Journal of Cell Biology, 106 (3):999–??, March 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/106/3/999.

Ball:1986:VIP

[BFG86] E. H. Ball, C. Freitag, and S. Gurofsky. Vinculin interaction with permeabilized cells: disruption and reconstitution of a binding site. Journal of Cell Biology, 103(2):641–??, August 1986. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/103/2/641.

Busa:1985:AFX

[BFJ^{+}85] W. B. Busa, J. E. Ferguson, S. K. Joseph, J. R. Williamson, and R. Nuccitelli. Activation of frog (Xenopus laevis) eggs by inositol trisphosphate. I. Characterization of Ca^{2+} release from intracellular stores. Journal of Cell Biology, 101(2):677–??, August 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/101/2/
677.

Buckley:1987:CSA

[BFK87] K. M. Buckley, E. Floor, and R. B. Kelly. Cloning and sequence analysis of cDNA encoding p38, a major synaptic vesicle protein.

Journal of Cell Biology, 105(6):2447–??, December 1987. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/105/6/2447.

Blumberg:1986:BSA

[BFKF86] B. Blumberg, L. I. Fessler, M. Kurkinen, and J. H. Fessler. Biosynthesis and supramolecular assembly of procollagen IV in neonatal lung. Journal of Cell Biology, 103(5):1711–??, November 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/103/5/
1711.

Bronner-Fraser:1988:MAA

[BFL88] M. Bronner-Fraser and T. Lallier. A monoclonal antibody against a laminin-heparan sulfate proteoglycan complex perturbs cranial neural crest migration in vivo. Journal of Cell Biology, 106 (4):1321–??, April 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/106/4/1321.

Boyles:1985:OCR

[BFPS85] J. Boyles, J. E. Fox, D. R. Phillips, and P. E. Stenberg. Organization of the cytoskeleton in resting, discoid platelets: preservation of actin filaments by a modified fixation that prevents osmium damage. Journal of Cell Biology, 101(4):1463–??, October 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/101/4/
1463.

Binder:1985:DTM

[BFR85] L. I. Binder, A. Frankfurter, and L. I. Rebhun. The distribution of tau in the mammalian central nervous system. Journal of Cell Biology, 101(4):1371–??, October 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/101/4/1371.

Bright:1987:FRI

[BFRT87] G. R. Bright, G. W. Fisher, J. Rogowska, and D. L. Taylor. Fluorescence ratio imaging microscopy: temporal and spatial measurements of cytoplasmic pH. Journal of Cell Biology, 104 (4):1019–??, January 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/104/4/1019.

Bartles:1987:BRH

[BFSH87] J. R. Bartles, H. M. Feracci, B. Stieger, and A. L. Hubbard. Biogenesis of the rat hepatocyte plasma membrane in vivo: comparison of the pathways taken by apical and basolateral proteins using subcellular fractionation. Journal of Cell Biology, 105(3): 1241–??, September 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/3/1241.

Berg:1986:MTR

[BG86] C. A. Berg and J. G. Gall. Microinjected Tetrahymena rDNA ends are not recognized as telomeres in Xenopus eggs. Journal of Cell Biology, 103(3):691–??, September 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/103/3/691.

Benfenati:1989:EHI

[BGBB89] F. Benfenati, P. Greengard, J. Brunner, and M. Bähler. Electrostatic and hydrophobic interactions of synapsin I and synapsin I fragments with phospholipid bilayers. Journal of Cell Biology, 108(5):1851–??, May 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/108/5/1851.

Brown:1986:MPR

[BGF86] W. J. Brown, J. Goodhouse, and M. G. Farquhar. Mannose-6-phosphate receptors for lysosomal enzymes cycle between the Golgi complex and endosomes. Journal of Cell Biology, 103(4): 1235–??, October 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/103/4/1235.

Bouche:1988:PIC

[BGF88] M. Bouché, S. M. Goldfine, and D. A. Fischman. Posttranslational incorporation of contractile proteins into myofibrils in a cell-free system. Journal of Cell Biology, 107(2):587–??, August 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/107/2/
587.

Black:1987:GVC

[BGH^{+}87a] R. Black, D. Goldman, S. Hochschwender, J. Lindstrom, and Z. W. Hall. Genetic variants of C2 muscle cells that are defective in synthesis of the alpha-subunit of the acetylcholine receptor. Journal of Cell Biology, 105(3):1329–??, September 1987. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/105/3/1329.

Brown:1987:SNM

[BGH87b] D. Brown, S. Gluck, and J. Hartwig. Structure of the novel membrane-coating material in proton-secreting epithelial cells and identification as an H+ATPase. Journal of Cell Biology, 105 (4):1637–??, October 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/4/1637.

Braulke:1987:MMP

[BGHvF87] T. Braulke, C. Gartung, A. Hasilik, and K. von Figura. Is movement of mannose 6-phosphate-specific receptor triggered by binding of lysosomal enzymes? Journal of Cell Biology, 104 (6):1735–??, June 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/104/6/1735.

Braunhut:1989:EFG

[BGK^{+}89] S. J. Braunhut, L. J. Gudas, T. Kurokawa, J. Sasse, and P. A. D’Amore. Expression of fibroblast growth factor by F9 teratocarcinoma cells as a function of differentiation. Journal of Cell Biology, 108(6):2467–??, June 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/108/6/2467.

Blok:1988:IIT

[BGL^{+}88] J. Blok, E. M. Gibbs, G. E. Lienhard, J. W. Slot, and H. J. Geuze. Insulin-induced translocation of glucose transporters from post-Golgi compartments to the plasma membrane of 3T3-L1 adipocytes. Journal of Cell Biology, 106(1):69–??, January 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/106/1/
69.

Beaven:1987:SSM

[BGM^{+}87] M. A. Beaven, D. F. Guthrie, J. P. Moore, G. A. Smith, T. R. Hesketh, and J. C. Metcalfe. Synergistic signals in the mechanism of antigen-induced exocytosis in 2H3 cells: evidence for an unidentified signal required for histamine release. Journal of Cell Biology, 105(3):1129–??, September 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/105/3/1129.

Barasch:1988:TIA

[BGN^{+}88] J. Barasch, M. D. Gershon, E. A. Nunez, H. Tamir, and Q. al Awqati. Thyrotropin induces the acidification of the secretory granules of parafollicular cells by increasing the chloride conductance of the granular membrane. Journal of Cell Biology, 107 (6):2137–??, December 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/107/6/2137.

Bird:1987:TYM

[BGS87] P. Bird, M. J. Gething, and J. Sambrook. Translocation in yeast and mammalian cells: not all signal sequences are functionally equivalent. Journal of Cell Biology, 105(6):2905–??, December 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/105/6/
2905.

Bendena:1989:MID

[BGT^{+}89] W. G. Bendena, J. C. Garbe, K. L. Traverse, S. C. Lakhotia, and M. L. Pardue. Multiple inducers of the Drosophila heat shock locus 93D (hsr omega): inducer-specific patterns of the three transcripts. Journal of Cell Biology, 108(6):2017–??, June 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/108/6/
2017.

Baas:1986:MRN

[BH86a] P. W. Baas and S. R. Heidemann. Microtubule reassembly from nucleating fragments during the regrowth of amputated neurites. Journal of Cell Biology, 103(3):917–??, September 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/103/3/917.

Bryan:1986:DTA

[BH86b] J. Bryan and S. Hwo. Definition of an N-terminal actin-binding domain and a C-terminal Ca^{2+} regulatory domain in human brevin. Journal of Cell Biology, 102(4):1439–??, April 1986. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/102/4/1439.

Baeuerle:1987:TST

[BH87] P. A. Baeuerle and W. B. Huttner. Tyrosine sulfation is a trans-Golgi-specific protein modification. Journal of Cell Biology, 105 (6):2655–??, December 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/6/2655.

Bernstein:1985:CGP

[BHAS85] M. Bernstein, W. Hoffmann, G. Ammerer, and R. Schekman. Characterization of a gene product (Sec53p) required for protein assembly in the yeast endoplasmic reticulum. Journal of Cell Biology, 101(6):2374–??, December 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/101/6/2374.

Bierer:1987:LMC

[BHB^{+}87] B. E. Bierer, S. H. Herrmann, C. S. Brown, S. J. Burakoff, and D. E. Golan. Lateral mobility of class I histocompatibility antigens in B lymphoblastoid cell membranes: modulation by crosslinking and effect of cell density. Journal of Cell Biology, 105(3): 1147–??, September 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/3/1147.

Byers:1989:SSA

[BHCD^{+}89] T. J. Byers, A. Husain-Chishti, R. R. Dubreuil, D. Branton, and L. S. Goldstein. Sequence similarity of the amino-terminal domain of Drosophila beta spectrin to alpha actinin and dystrophin. Journal of Cell Biology, 109(4):1633–??, October 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/109/4/1633.

Bole:1986:PAI

[BHK86] D. G. Bole, L. M. Hendershot, and J. F. Kearney. Posttranslational association of immunoglobulin heavy chain binding protein with nascent heavy chains in nonsecreting and secreting hybridomas. Journal of Cell Biology, 102(5):1558–??, May 1986. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/102/5/1558.

Breitfeld:1989:PSL

[BHM89a] P. P. Breitfeld, J. M. Harris, and K. E. Mostov. Postendocytotic sorting of the ligand for the polymeric immunoglobulin receptor in Madin–Darby canine kidney cells. Journal of Cell Biology, 109(2):475–??, August 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/2/475.

Bruckner:1989:IPC

[BHM^{+}89b] P. Bruckner, I. Hörler, M. Mendler, Y. Houze, K. H. Winterhalter, S. G. Eich-Bender, and M. A. Spycher. Induction and prevention of chondrocyte hypertrophy in culture. Journal of Cell Biology, 109(5):2537–??, November 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/109/5/2537.

Barondes:1985:DEL

[BHRC85] S. H. Barondes, P. L. Haywood-Reid, and D. N. Cooper. Discoidin I, an endogenous lectin, is externalized from Dictyostelium discoideum in multilamellar bodies. Journal of Cell Biology, 100 (6):1825–??, June 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/100/6/1825.

Beck:1988:IPM

[BHS88] L. A. Beck, T. J. Hosick, and M. Sinensky. Incorporation of a product of mevalonic acid metabolism into proteins of Chinese hamster ovary cell nuclei. Journal of Cell Biology, 107 (4):1307–??, October 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/107/4/1307.

Baumann:1986:RMA

[BHSJ86] H. Baumann, R. E. Hill, D. N. Sauder, and G. P. Jahreis. Regulation of major acute-phase plasma proteins by hepatocytestimulating factors of human squamous carcinoma cells. Journal of Cell Biology, 102(2):370–??, February 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/102/2/370.

Block:1988:SED

[BICFA88] B. A. Block, T. Imagawa, K. P. Campbell, and C. Franzini-Armstrong. Structural evidence for direct interaction between the molecular components of the transverse tubule/sarcoplasmic reticulum junction in skeletal muscle. Journal of Cell Biology, 107(6):2587–??, December 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/107/6/2587.

Buchanan:1989:ACW

[BIES89] M. J. Buchanan, S. H. Imam, W. A. Eskue, and W. J. Snell. Activation of the cell wall degrading protease, lysin, during sexual signalling in Chlamydomonas: the enzyme is stored as an inactive, higher relative molecular mass precursor in the periplasm. Journal of Cell Biology, 108(1):199–??, January 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/108/1/199.

Bischoff:1986:RAN

[Bis86] R. Bischoff. Rapid adhesion of nerve cells to muscle fibers from adult rats is mediated by a sialic acid-binding receptor. Journal of Cell Biology, 102(6):2273–??, June 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/102/6/2273.

Brown:1986:EFP

[BJ86] P. J. Brown and R. L. Juliano. Expression and function of a putative cell surface receptor for fibronectin in hamster and human cell lines. Journal of Cell Biology, 103(4):1595–??, October 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/103/4/
1595.

Burgess:1989:ICS

[BJMK89] D. R. Burgess, W. P. Jiang, S. Mamajiwalla, and W. Kinsey. Intestinal crypt stem cells possess high levels of cytoskeletalassociated phosphotyrosine-containing proteins and tyrosine kinase activity relative to differentiated enterocytes. Journal of Cell Biology, 109(5):2139–??, November 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/109/5/2139.

Buckley:1985:ITG

[BK85] K. Buckley and R. B. Kelly. Identification of a transmembrane glycoprotein specific for secretory vesicles of neural and endocrine cells. Journal of Cell Biology, 100(4):1284–??, April 1985. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/100/4/1284.

Benavente:1986:INL

[BK86] R. Benavente and G. Krohne. Involvement of nuclear lamins in postmitotic reorganization of chromatin as demonstrated by microinjection of lamin antibodies. Journal of Cell Biology, 103 (5):1847–??, November 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/103/5/1847.

Barnett:1989:CAA

[BKA^{+}89] T. R. Barnett, A. Kretschmer, D. A. Austen, S. J. Goebel, J. T. Hart, J. J. Elting, and M. E. Kamarck. Carcinoembryonic antigens: alternative splicing accounts for the multiple mRNAs that code for novel members of the carcinoembryonic antigen family. Journal of Cell Biology, 108(2):267–??, February 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/108/2/267.

Black:1988:HLS

[BKBR88] J. D. Black, S. T. Koury, R. B. Bankert, and E. A. Repasky. Heterogeneity in lymphocyte spectrin distribution: ultrastructural identification of a new spectrin-rich cytoplasmic structure. Journal of Cell Biology, 106(1):97–??, January 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/106/1/97.

Blair:1986:IOR

[BKC^{+}86] H. C. Blair, A. J. Kahn, E. C. Crouch, J. J. Jeffrey, and S. L. Teitelbaum. Isolated osteoclasts resorb the organic and inorganic components of bone. Journal of Cell Biology, 102(4):1164–??, April 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
102/4/1164.

Bre:1987:CMN

[BKK87] M. H. Bré, T. E. Kreis, and E. Karsenti. Control of microtubule nucleation and stability in Madin–Darby canine kidney cells: the occurrence of noncentrosomal, stable detyrosinated microtubules. Journal of Cell Biology, 105(3):1283–??, September 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/105/3/

Becker:1985:ING

[BKN^{+}85] E. L. Becker, J. C. Kermode, P. H. Naccache, R. Yassin, M. L. Marsh, J. J. Munoz, and R. I. Sha’afi. The inhibition of neutrophil granule enzyme secretion and chemotaxis by pertussis toxin. Journal of Cell Biology, 100(5):1641–??, May 1985. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/100/5/1641.

Beyer:1989:ADA

[BKPG89] E. C. Beyer, J. Kistler, D. L. Paul, and D. A. Goodenough. Antisera directed against connexin43 peptides react with a 43-kD protein localized to gap junctions in myocardium and other tissues. Journal of Cell Biology, 108(2):595–??, February 1989. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/108/2/595.

Bridgman:1986:SCD

[BKR86] P. C. Bridgman, B. Kachar, and T. S. Reese. The structure of cytoplasm in directly frozen cultured cells. II. Cytoplasmic domains associated with organelle movements. Journal of Cell Biology, 102(4):1510–??, April 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/102/4/1510.

Bretscher:1985:ILI

[BL85] A. Bretscher and W. Lynch. Identification and localization of immunoreactive forms of caldesmon in smooth and nonmuscle cells: a comparison with the distributions of tropomyosin and alphaactinin. Journal of Cell Biology, 100(5):1656–??, May 1985. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/100/5/1656.

Bosch:1988:ESE

[BLA^{+}88] F. X. Bosch, R. E. Leube, T. Achtstätter, R. Moll, and W. W. Franke. Expression of simple epithelial type cytokeratins in stratified epithelia as detected by immunolocalization and hybridization in situ. Journal of Cell Biology, 106(5):1635–??, May 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/106/5/1635.

Burns:1988:SBM

[BLK^{+}88] G. F. Burns, C. M. Lucas, G. W. Krissansen, J. A. Werkmeister, D. B. Scanlon, R. J. Simpson, and M. A. Vadas. Synergism between membrane gangliosides and Arg–Gly–Asp–directed glycoprotein receptors in attachment to matrix proteins by melanoma cells. Journal of Cell Biology, 107(3):1225–??, September 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/107/3/1225.

Bloch:1986:ARR

[Blo86] R. J. Bloch. Actin at receptor-rich domains of isolated acetylcholine receptor clusters. Journal of Cell Biology, 102(4):1447–??, April 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
102/4/1447.

Bixby:1988:IMP

[BLR88] J. L. Bixby, J. Lilien, and L. F. Reichardt. Identification of the major proteins that promote neuronal process outgrowth on Schwann cells in vitro. Journal of Cell Biology, 107(1):353–??, July 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/107/1/
353.

Benveniste:1988:OEG

[BLSK88] M. Benveniste, E. Livneh, J. Schlessinger, and Z. Kam. Overexpression of epidermal growth factor receptor in NIH-3T3-transfected cells slows its lateral diffusion and rate of endocytosis. Journal of Cell Biology, 106(6):1903–??, June 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/106/6/1903.

Bischoff:1988:HHP

[BLSL88] J. Bischoff, S. Libresco, M. A. Shia, and H. F. Lodish. The H1 and H2 polypeptides associate to form the asialoglycoprotein receptor in human hepatoma cells. Journal of Cell Biology, 106 (4):1067–??, April 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/106/4/1067.

Bonder:1986:CBS

[BM86] E. M. Bonder and M. S. Mooseker. Cytochalasin B slows but does not prevent monomer addition at the barbed end of the actin filament. Journal of Cell Biology, 102(1):282–??, January 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/102/1/

Brugg:1988:PCE

[BM88] B. Brugg and A. Matus. PC12 cells express juvenile microtubuleassociated proteins during nerve growth factor-induced neurite outgrowth. Journal of Cell Biology, 107(2):643–??, August 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/107/2/643.

Bloch:1989:UBS

[BM89] R. J. Bloch and J. S. Morrow. An unusual beta-spectrin associated with clustered acetylcholine receptors. Journal of Cell Biology, 108(2):481–??, February 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/108/2/481.

Bajer:1986:RME

[BMB86] A. S. Bajer and J. Molè-Bajer. Reorganization of microtubules in endosperm cells and cell fragments of the higher plant Haemanthus in vivo. Journal of Cell Biology, 102(1):263–??, January 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/102/1/
263.

Bravo:1987:ETP

[BMB87] R. Bravo and H. Macdonald-Bravo. Existence of two populations of cyclin/proliferating cell nuclear antigen during the cell cycle: association with DNA replication sites. Journal of Cell Biology, 105(4):1549–??, October 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/105/4/1549.

Beckerle:1989:ADR

[BMBL89] M. C. Beckerle, D. E. Miller, M. E. Bertagnolli, and S. J. Locke. Activation-dependent redistribution of the adhesion plaque protein, talin, in intact human platelets. Journal of Cell Biology, 109 (6):3333–??, December 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/6/3333.

Bordas:1987:XRD

[BMDN87] J. Bordas, G. R. Mant, G. P. Diakun, and C. Nave. X-ray diffraction evidence for the existence of 102.0- and 230.0-nm transverse periodicities in striated muscle. Journal of Cell Biology, 105(3):

1311–??, September 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/3/1311.

Bankaitis:1989:SCS

[BMEG89] V. A. Bankaitis, D. E. Malehorn, S. D. Emr, and R. Greene. The Saccharomyces cerevisiae SEC14 gene encodes a cytosolic factor that is required for transport of secretory proteins from the yeast Golgi complex. Journal of Cell Biology, 108(4):1271–??, April 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/108/4/
1271.

Brandan:1985:ACT

[BMGI85] E. Brandan, M. Maldonado, J. Garrido, and N. C. Inestrosa. Anchorage of collagen-tailed acetylcholinesterase to the extracellular matrix is mediated by heparan sulfate proteoglycans. Journal of Cell Biology, 101(3):985–??, September 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/101/3/985.

Bowles:1986:PPC

[BMP^{+}86] D. J. Bowles, S. E. Marcus, D. J. Pappin, J. B. Findlay, E. Eliopoulos, P. R. Maycox, and J. Burgess. Posttranslational processing of concanavalin A precursors in jackbean cotyledons. Journal of Cell Biology, 102(4):1284–??, April 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/102/4/1284.

Banerjee:1985:BHD

[BMR85] D. Banerjee, T. K. Mukherjee, and C. M. Redman. Biosynthesis of high density lipoprotein by chicken liver: intracellular transport and proteolytic processing of nascent apolipoprotein A-1. Journal of Cell Biology, 101(4):1219–??, October 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/101/4/1219.

Behrens:1989:DTC

[BMVB89] J. Behrens, M. M. Mareel, F. M. Van Roy, and W. Birchmeier. Dissecting tumor cell invasion: epithelial cells acquire invasive properties after the loss of uvomorulin-mediated cell–cell adhesion. Journal of Cell Biology, 108(6):2435–??, June 1989. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/108/6/2435.

Brokaw:1985:MAS

[BN85a] C. J. Brokaw and S. M. Nagayama. Modulation of the asymmetry of sea urchin sperm flagellar bending by calmodulin. Journal of Cell Biology, 100(6):1875–??, June 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/100/6/1875.

Busa:1985:EFC

[BN85b] W. B. Busa and R. Nuccitelli. An elevated free cytosolic Ca^{2+} wave follows fertilization in eggs of the frog, Xenopus laevis. Journal of Cell Biology, 100(4):1325–??, April 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/100/4/1325.

Bassot:1987:ODJ

[BN87] J. M. Bassot and G. Nicolas. An optional dyadic junctional complex revealed by fast-freeze fixation in the bioluminescent system of the scale worm. Journal of Cell Biology, 105(5): 2245–??, November 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/5/2245.

Bonini:1988:DRP

[BN88] N. M. Bonini and D. L. Nelson. Differential regulation of Paramecium ciliary motility by cAMP and cGMP. Journal of Cell Biology, 106(5):1615–??, May 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/106/5/1615.

Baron:1988:PSL

[BNB^{+}88] R. Baron, L. Neff, W. Brown, P. J. Courtoy, D. Louvard, and M. G. Farquhar. Polarized secretion of lysosomal enzymes: codistribution of cation-independent mannose-6-phosphate receptors and lysosomal enzymes along the osteoclast exocytic pathway. Journal of Cell Biology, 106(6):1863–??, June 1988. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/106/6/1863.

Brenneman:1987:NCM

[BNF^{+}87] D. E. Brenneman, E. A. Neale, G. A. Foster, S. W. d’Autremont, and G. L. Westbrook. Nonneuronal cells mediate neurotrophic action of vasoactive intestinal peptide. Journal of Cell Biology, 104(6):1603–??, June 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/104/6/1603.

Bowen:1989:CHH

[BNL89] B. R. Bowen, T. Nguyen, and L. A. Lasky. Characterization of a human homologue of the murine peripheral lymph node homing receptor. Journal of Cell Biology, 109(1):421–??, July 1989. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/109/1/421.

Baron:1985:CME

[BNLC85] R. Baron, L. Neff, D. Louvard, and P. J. Courtoy. Cell-mediated extracellular acidification and bone resorption: evidence for a low pH in resorbing lacunae and localization of a 100-kD lysosomal membrane protein at the osteoclast ruffled border. Journal of Cell Biology, 101(6):2210–??, December 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/101/6/2210.

Brown:1985:EAP

[BNS85] J. A. Brown, E. K. Novak, and R. T. Swank. Effects of ammonia on processing and secretion of precursor and mature lysosomal enzyme from macrophages of normal and pale ear mice: evidence for two distinct pathways. Journal of Cell Biology, 100 (6):1894–??, June 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/100/6/1894.

Brown:1987:LLM

[BNT^{+}87] J. Brown, E. K. Novak, K. Takeuchi, K. Moore, S. Medda, and R. T. Swank. Lumenal location of the microsomal betaglucuronidase-egasyn complex. Journal of Cell Biology, 105(4): 1571–??, October 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/4/1571.

Bearer:1985:EFD

[BO85] E. L. Bearer and L. Orci. Endothelial fenestral diaphragms: a quick-freeze, deep-etch study. Journal of Cell Biology, 100 (2):418–??, February 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/100/2/418.

Belkin:1988:IMV

[BOGK88] A. M. Belkin, O. I. Ornatsky, M. A. Glukhova, and V. E. Koteliansky. Immunolocalization of meta-vinculin in human smooth and cardiac muscles. Journal of Cell Biology, 107(2): 545–??, August 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/107/2/545.

Bartnik:1985:IFN

[BOW85] E. Bartnik, M. Osborn, and K. Weber. Intermediate filaments in non-neuronal cells of invertebrates: isolation and biochemical characterization of intermediate filaments from the esophageal epithelium of the mollusc Helix pomatia. Journal of Cell Biology, 101(2):427–??, August 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/101/2/427.

Bartnik:1986:IFM

[BOW86] E. Bartnik, M. Osborn, and K. Weber. Intermediate filaments in muscle and epithelial cells of nematodes. Journal of Cell Biology, 102(6):2033–??, June 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/102/6/2033.

Brown:1987:UAA

[BP87a] S. S. Brown and A. S. Petzold. Using antibodies against Dictyostelium membranes to identify an actin-binding membrane protein. Journal of Cell Biology, 104(3):513–??, March 1987. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/104/3/513.

Brunden:1987:LDM

[BP87b] K. R. Brunden and J. F. Poduslo. Lysosomal delivery of the major myelin glycoprotein in the absence of myelin assembly: posttranslational regulation of the level of expression by Schwann cells. Journal of Cell Biology, 104(3):661–??, March 1987. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/104/3/661.

Baetscher:1986:VSC

[BPB86] M. Baetscher, D. W. Pumplin, and R. J. Bloch. Vitronectin at sites of cell-substrate contact in cultures of rat myotubes. Journal of Cell Biology, 103(2):369–??, August 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/103/2/369.

Boukamp:1988:NKS

[BPB^{+}88] P. Boukamp, R. T. Petrussevska, D. Breitkreutz, J. Hornung, A. Markham, and N. E. Fusenig. Normal keratinization in a spontaneously immortalized aneuploid human keratinocyte cell line. Journal of Cell Biology, 106(3):761–??, March 1988. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/106/3/761.

Bornens:1989:CMS

[BPC89] M. Bornens, M. Paintrand, and C. Celati. The cortical microfilament system of lymphoblasts displays a periodic oscillatory activity in the absence of microtubules: implications for cell polarity. Journal of Cell Biology, 109(3):1071–??, September 1989. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/109/3/1071.

Bruns:1986:TVC

[BPE^{+}86] R. R. Bruns, W. Press, E. Engvall, R. Timpl, and J. Gross. Type VI collagen in extracellular, 100-nm periodic filaments and fibrils: identification by immunoelectron microscopy. Journal of Cell Biology, 103(2):393–??, August 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/103/2/393.

Beyer:1987:CPR

[BPG87] E. C. Beyer, D. L. Paul, and D. A. Goodenough. Connexin43: a protein from rat heart homologous to a gap junction protein from liver. Journal of Cell Biology, 105(6):2621–??, December 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/105/6/
2621.

Benson:1985:BCE

[BPJBF85] R. J. Benson, K. Porter-Jordan, P. Buoniconti, and R. E. Fine. Biochemical and cytochemical evidence indicates that coated vesicles in chick embryo myotubes contain newly synthesized acetylcholinesterase. Journal of Cell Biology, 101(5): 1930–??, November 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/101/5/1930.

Bonifacino:1986:STI

[BPKS86] J. S. Bonifacino, P. Perez, R. D. Klausner, and I. V. Sandoval. Study of the transit of an integral membrane protein from secretory granules through the plasma membrane of secreting rat basophilic leukemia cells using a specific monoclonal antibody. Journal of Cell Biology, 102(2):516–??, February 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/102/2/516.

Bernardi:1987:LPC

[BPL87] P. Bernardi, V. P. Patel, and H. F. Lodish. Lymphoid precursor cells adhere to two different sites on fibronectin. Journal of Cell Biology, 105(1):489–??, July 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/105/1/489.

Bennett:1987:AAC

[BPMD87] A. L. Bennett, K. E. Paulson, R. E. Miller, and J. E. Darnell. Acquisition of antigens characteristic of adult pericentral hepatocytes by differentiating fetal hepatoblasts in vitro. Journal of Cell Biology, 105(3):1073–??, September 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/105/3/1073.

Bomsel:1989:EFG

[BPP^{+}89] M. Bomsel, K. Prydz, R. G. Parton, J. Gruenberg, and K. Simons. Endocytosis in filter-grown Madin–Darby canine kidney cells. Journal of Cell Biology, 109(6):3243–??, December 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/109/6/3243.

Brawley:1985:CTD

[BR85] S. H. Brawley and K. R. Robinson. Cytochalasin treatment disrupts the endogenous currents associated with cell polarization in fucoid zygotes: studies of the role of F-actin in embryogenesis. Journal of Cell Biology, 100(4):1173–??, April 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/100/4/1173.

Bourdon:1989:TMC

[BR89] M. A. Bourdon and E. Ruoslahti. Tenascin mediates cell attachment through an RGD-dependent receptor. Journal of Cell Biology, 108(3):1149–??, March 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/108/3/1149.

Bortz:1988:FEI

[BRD^{+}88] J. D. Bortz, P. Rotwein, D. DeVol, P. J. Bechtel, V. A. Hansen, and M. R. Hammerman. Focal expression of insulin-like growth factor I in rat kidney collecting duct. Journal of Cell Biology, 107(2):811–??, August 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/107/2/811.

Bretcher:1988:FM

[Bre88] M. S. Bretcher. Fibroblasts on the move. Journal of Cell Biology, 106(2):235–??, February 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/106/2/235.

Bretscher:1989:RPR

[Bre89] A. Bretscher. Rapid phosphorylation and reorganization of ezrin and spectrin accompany morphological changes induced in A-431 cells by epidermal growth factor. Journal of Cell Biology, 108(3):921–??, March 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/108/3/921.

Bols:1986:SCL

[BRHR^{+}86] N. C. Bols, M. M. Roberson, P. L. Haywood-Reid, R. F. Cerra, and S. H. Barondes. Secretion of a cytoplasmic lectin from Xenopus laevis skin. Journal of Cell Biology, 102(2):492–??, February 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/102/2/
492.

Banta:1988:OAY

[BRKE88] L. M. Banta, J. S. Robinson, D. J. Klionsky, and S. D. Emr. Organelle assembly in yeast: characterization of yeast mutants defective in vacuolar biogenesis and protein sorting. Journal of Cell Biology, 107(4):1369–??, October 1988. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/107/4/1369.

Brodsky:1985:CSCa

[Bro85a] F. M. Brodsky. Clathrin structure characterized with monoclonal antibodies. I. Analysis of multiple antigenic sites. Journal of Cell Biology, 101(6):2047–??, December 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/101/6/2047.

Brodsky:1985:CSCb

[Bro85b] F. M. Brodsky. Clathrin structure characterized with monoclonal antibodies. II. Identification of in vivo forms of clathrin. Journal of Cell Biology, 101(6):2055–??, December 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/101/6/2055.

Brokaw:1987:LSR

[Bro87] C. J. Brokaw. A lithium-sensitive regulator of sperm flagellar oscillation is activated by cAMP-dependent phosphorylation. Journal of Cell Biology, 105(4):1789–??, October 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/105/4/1789.

Bulinski:1988:PMA

[BRP88] J. C. Bulinski, J. E. Richards, and G. Piperno. Posttranslational modifications of alpha tubulin: detyrosination and acetylation differentiate populations of interphase microtubules in cultured cells. Journal of Cell Biology, 106(4):1213–??, April 1988. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/106/4/1213.

Benavente:1987:INR

[BRR^{+}87] R. Benavente, K. M. Rose, G. Reimer, B. Hügle-Dörr, and U. Scheer. Inhibition of nucleolar reformation after microinjection of antibodies to RNA polymerase I into mitotic cells. Journal of Cell Biology, 105(4):1483–??, October 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/105/4/1483.

Brandley:1987:MCR

[BRS87] B. K. Brandley, T. S. Ross, and R. L. Schnaar. Multiple carbohydrate receptors on lymphocytes revealed by adhesion to immobilized polysaccharides. Journal of Cell Biology, 105(2):991–??, August 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
105/2/991.

Bryan:1988:GTA

[Bry88] J. Bryan. Gelsolin has three actin-binding sites. Journal of Cell Biology, 106(5):1553–??, May 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/106/5/1553.

Brown:1985:SSM

[BS85] J. C. Brown and N. L. Salomonsky. Site-specific maturation of enveloped viruses in L cells treated with cytochalasin B. Journal of Cell Biology, 100(2):357–??, February 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/100/2/357.

Brelvi:1986:CEO

[BS86] Z. S. Brelvi and G. P. Studzinski. Changes in the expression of oncogenes encoding nuclear phosphoproteins but not c–Ha–ras have a relationship to monocytic differentiation of HL 60 cells. Journal of Cell Biology, 102(6):2234–??, June 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/102/6/2234.

Barkai:1987:AIB

[BS87] U. Barkai and M. I. Sherman. Analyses of the interactions between retinoid-binding proteins and embryonal carcinoma cells. Journal of Cell Biology, 104(3):671–??, March 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/104/3/671.

Baron:1988:ILN

[BS88a] A. T. Baron and J. L. Salisbury. Identification and localization of a novel, cytoskeletal, centrosome-associated protein in PtK2 cells. Journal of Cell Biology, 107(6):2669–??, December 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/107/6/
2669.

Battaglia:1988:HPC

[BS88b] D. E. Battaglia and B. M. Shapiro. Hierarchies of protein crosslinking in the extracellular matrix: involvement of an egg surface transglutaminase in early stages of fertilization envelope assembly. Journal of Cell Biology, 107(6):2447–??, December 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/107/6/2447.

Birgbauer:1989:MBA

[BS89] E. Birgbauer and F. Solomon. A marginal band-associated protein has properties of both microtubule- and microfilamentassociated proteins. Journal of Cell Biology, 109(4):1609–??, October 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
109/4/1609.

Belmont:1987:TDA

[BSA87] A. S. Belmont, J. W. Sedat, and D. A. Agard. A threedimensional approach to mitotic chromosome structure: evidence for a complex hierarchical organization. Journal of Cell Biology, 105(1):77–??, July 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/1/77.

Burgess:1986:PER

[BSB^{+}86] S. K. Burgess, N. Sahyoun, S. G. Blanchard, H. LeVine, K. J. Chang, and P. Cuatrecasas. Phorbol ester receptors and protein kinase C in primary neuronal cultures: development and stimulation of endogenous phosphorylation. Journal of Cell Biology, 102(1):312–??, January 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/102/1/312.

Blood:1988:ITC

[BSBZ88] C. H. Blood, J. Sasse, P. Brodt, and B. R. Zetter. Identification of a tumor cell receptor for VGVAPG, an elastin-derived chemotactic peptide. Journal of Cell Biology, 107(5):1987–??, November 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/107/5/
1987.

Brett:1988:NRA

[BSd^{+}88] J. G. Brett, S. F. Steinberg, P. G. deGroot, P. P. Nawroth, and D. M. Stern. Norepinephrine down-regulates the activity of protein S on endothelial cells. Journal of Cell Biology, 106 (6):2109–??, June 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/106/6/2109.

Buck:1986:ICA

[BSDH86] C. A. Buck, E. Shea, K. Duggan, and A. F. Horwitz. Integrin (the CSAT antigen): functionality requires oligomeric integrity. Journal of Cell Biology, 103(6):2421–??, December 1986. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/103/6/2421.

Burstein:1985:LII

[BSG85] D. E. Burstein, P. J. Seeley, and L. A. Greene. Lithium ion inhibits nerve growth factor-induced neurite outgrowth and phosphorylation of nerve growth factor-modulated microtubuleassociated proteins. Journal of Cell Biology, 101(3):862– ??, September 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/101/3/862.

Bendori:1989:ITD

[BSG89] R. Bendori, D. Salomon, and B. Geiger. Identification of two distinct functional domains on vinculin involved in its association with focal contacts. Journal of Cell Biology, 108(6):2383–??, June 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/108/6/
2383.

Brands:1985:RMP

[BSH^{+}85] R. Brands, M. D. Snider, Y. Hino, S. S. Park, H. V. Gelboin, and J. E. Rothman. Retention of membrane proteins by the endoplasmic reticulum. Journal of Cell Biology, 101(5):1724–??, November 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/101/5/
1724.

Bourguignon:1986:LTG

[BSK86] L. Y. Bourguignon, S. J. Suchard, and E. L. Kalomiris. Lymphoma Thy-1 glycoprotein is linked to the cytoskeleton via a 4.1-like protein. Journal of Cell Biology, 103(6):2529–??, December 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/103/6/

Benveniste:1989:CIE

[BSK89] M. Benveniste, J. Schlessinger, and Z. Kam. Characterization of internalization and endosome formation of epidermal growth factor in transfected NIH-3T3 cells by computerized imageintensified three-dimensional fluorescence microscopy. Journal of Cell Biology, 109(5):2105–??, November 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/109/5/2105.

Bonifacino:1989:PGD

[BSLS^{+}89] J. S. Bonifacino, C. K. Suzuki, J. Lippincott-Schwartz, A. M. Weissman, and R. D. Klausner. Pre-Golgi degradation of newly synthesized T-cell antigen receptor chains: intrinsic sensitivity and the role of subunit assembly. Journal of Cell Biology, 109(1):73–??, July 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/1/73.

Bourguignon:1985:LTG

[BSNG85] L. Y. Bourguignon, S. J. Suchard, M. L. Nagpal, and J. R. Glenney. A T-lymphoma transmembrane glycoprotein (gp180) is linked to the cytoskeletal protein, fodrin. Journal of Cell Biology, 101(2):477–??, August 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/101/2/477.

Bonaldo:1988:CSA

[BSPdSG88] M. C. Bonaldo, T. Souto-Padron, W. de Souza, and S. Goldenberg. Cell-substrate adhesion during Trypanosoma cruzi differentiation. Journal of Cell Biology, 106(4):1349–??, April 1988. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/106/4/1349.

Bacon:1989:GBP

[BSR^{+}89] R. A. Bacon, A. Salminen, H. Ruohola, P. Novick, and S. Ferro-Novick. The GTP-binding protein Ypt1 is required for transport in vitro: the Golgi apparatus is defective in ypt1 mutants. Journal of Cell Biology, 109(3):1015–??, September 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/109/3/1015.

Bar-Sagi:1988:LPA

[BSSMF88] D. Bar-Sagi, J. P. Suhan, F. McCormick, and J. R. Feramisco. Localization of phospholipase A2 in normal and ras-transformed cells. Journal of Cell Biology, 106(5):1649–??, May 1988. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/106/5/1649.

Bennett:1986:NPB

[BSY86] C. F. Bennett, D. L. Spector, and L. C. Yeoman. Nonhistone protein BA is a glutathione S-transferase localized to interchromatinic regions of the cell nucleus. Journal of Cell Biology, 102 (2):600–??, February 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/102/2/600.

Birk:1986:ECT

[BT86] D. E. Birk and R. L. Trelstad. Extracellular compartments in tendon morphogenesis: collagen fibril, bundle, and macroaggregate formation. Journal of Cell Biology, 103(1):231–??, July 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/103/1/
231.

Bader:1986:SCA

[BTL^{+}86] M. F. Bader, J. M. Trifaró, O. K. Langley, D. Thiersé, and D. Aunis. Secretory cell actin-binding proteins: identification of a gelsolin-like protein in chromaffin cells. Journal of Cell Biology, 102(2):636–??, February 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/102/2/636.

Blank:1988:CCV

[BTO88] R. S. Blank, M. M. Thompson, and G. K. Owens. Cell cycle versus density dependence of smooth muscle alpha actin expression in cultured rat aortic smooth muscle cells. Journal of Cell Biology, 107(1):299–??, July 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/107/1/299.

Binder:1986:SHE

[BTR^{+}86] M. Binder, S. Tourmente, J. Roth, M. Renaud, and W. J. Gehring. In situ hybridization at the electron microscope level: localization of transcripts on ultrathin sections of Lowicryl K4Membedded tissue using biotinylated probes and protein A-gold complexes. Journal of Cell Biology, 102(5):1646–??, May 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/102/5/1646.

Boyer:1989:RDC

[BTV^{+}89] B. Boyer, G. C. Tucker, A. M. Vallés, W. W. Franke, and J. P. Thiery. Rearrangements of desmosomal and cytoskeletal proteins during the transition from epithelial to fibroblastoid organization in cultured rat bladder carcinoma cells. Journal of Cell Biology, 109(4):1495–??, October 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/109/4/1495.

Blasi:1987:UTP

[BVD87] F. Blasi, J. D. Vassalli, and K. Danø. Urokinase-type plasminogen activator: proenzyme, receptor, and inhibitors. Journal of Cell Biology, 104(4):801–??, January 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/104/4/801.

Baumgart:1989:BPI

[BVHF89] E. Baumgart, A. Völkl, T. Hashimoto, and H. D. Fahimi. Biogenesis of peroxisomes: immunocytochemical investigation of peroxisomal membrane proteins in proliferating rat liver peroxisomes and in catalase-negative membrane loops. Journal of Cell Biology, 108(6):2221–??, June 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/108/6/2221.

Boller:1985:CAM

[BVK85] K. Boller, D. Vestweber, and R. Kemler. Cell-adhesion molecule uvomorulin is localized in the intermediate junctions of adult intestinal epithelial cells. Journal of Cell Biology, 100(1):327–??, January 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
100/1/327.

Bleil:1986:AVM

[BW86] J. D. Bleil and P. M. Wassarman. Autoradiographic visualization of the mouse egg’s sperm receptor bound to sperm. Journal of Cell Biology, 102(4):1363–??, April 1986. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/102/4/1363.

Balch:1987:RTV

[BWK87] W. E. Balch, K. R. Wagner, and D. S. Keller. Reconstitution of transport of vesicular stomatitis virus G protein from the endoplasmic reticulum to the Golgi complex using a cell-free system. Journal of Cell Biology, 104(3):749–??, March 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/104/3/749.

Broughton:1986:IRP

[BWL^{+}86] W. J. Broughton, C. H. Wong, A. Lewin, U. Samrey, H. Myint, H. Meyer, D. N. Dowling, and R. Simon. Identification of Rhizobium plasmid sequences involved in recognition of Psophocarpus, Vigna, and other legumes. Journal of Cell Biology, 102 (4):1173–??, April 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/102/4/1173.

Bloodgood:1986:RSF

[BWS86] R. A. Bloodgood, M. P. Woodward, and N. L. Salomonsky. Redistribution and shedding of flagellar membrane glycoproteins visualized using an anti-carbohydrate monoclonal antibody and concanavalin A. Journal of Cell Biology, 102(5):1797–??, May 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/102/5/
1797.

Bourguignon:1986:LPM

[BWSB86] L. Y. Bourguignon, G. Walker, S. J. Suchard, and K. Balazovich. A lymphoma plasma membrane-associated protein with ankyrinlike properties. Journal of Cell Biology, 102(6):2115–??, June 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/102/6/
2115.

Baldwin:1988:RAR

[BYB^{+}88] T. J. Baldwin, C. M. Yoshihara, K. Blackmer, C. R. Kintner, and S. J. Burden. Regulation of acetylcholine receptor transcript expression during development in Xenopus laevis. Journal of Cell Biology, 106(2):469–??, February 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/106/2/469.

Chia:1986:ETS

[CA86] C. P. Chia and C. J. Arntzen. Evidence for two-step processing of nuclear-encoded chloroplast proteins during membrane assembly. Journal of Cell Biology, 103(3):725–??, September 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/103/3/725.

Clark:1987:SLT

[CA87] M. W. Clark and J. Abelson. The subnuclear localization of tRNA ligase in yeast. Journal of Cell Biology, 105(4):1515–??, October 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
105/4/1515.

Cauley:1989:INN

[CAG89] K. Cauley, B. W. Agranoff, and D. Goldman. Identification of a novel nicotinic acetylcholine receptor structural subunit expressed in goldfish retina. Journal of Cell Biology, 108(2): 637–??, February 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/108/2/637.

Cerny:1986:CAR

[CB86] L. C. Cerny and E. Bandman. Contractile activity is required for the expression of neonatal myosin heavy chain in embryonic chick pectoral muscle cultures. Journal of Cell Biology, 103(6): 2153–??, December 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/103/6/2153.

Coluccio:1987:CRC

[CB87] L. M. Coluccio and A. Bretscher. Calcium-regulated cooperative binding of the microvillar 110K-calmodulin complex to Factin: formation of decorated filaments. Journal of Cell Biology, 105(1):325–??, July 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/1/325.

Coluccio:1988:MMC

[CB88] L. M. Coluccio and A. Bretscher. Mapping of the microvillar 110K-calmodulin complex: calmodulin-associated or -free fragments of the 110-kD polypeptide bind F-actin and retain AT-Pase activity. Journal of Cell Biology, 106(2):367–??, February 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/106/2/
367.

Coluccio:1989:RMC

[CB89] L. M. Coluccio and A. Bretscher. Reassociation of microvillar core proteins: making a microvillar core in vitro. Journal of Cell Biology, 108(2):495–??, February 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/108/2/495.

Colombatti:1985:GGI

[CBCV85] A. Colombatti, G. M. Bressan, I. Castellani, and D. Volpin. Glycoprotein 115, a glycoprotein isolated from chick blood vessels, is widely distributed in connective tissue. Journal of Cell Biology, 100(1):18–??, January 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/100/1/18.

Carpentier:1985:DSB

[CBIO85] J. L. Carpentier, D. Brown, B. Iacopetta, and L. Orci. Detection of surface-bound ligands by freeze-fracture autoradiography. Journal of Cell Biology, 101(3):887–??, September 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/101/3/887.

Clayton:1985:MNS

[CBL85] L. Clayton, C. M. Black, and C. W. Lloyd. Microtubule nucleating sites in higher plant cells identified by an auto-antibody against pericentriolar material. Journal of Cell Biology, 101 (1):319–??, July 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/101/1/319.

Cooper:1987:MGL

[CBS^{+}87] J. A. Cooper, J. Bryan, B. Schwab, C. Frieden, D. J. Loftus, and E. L. Elson. Microinjection of gelsolin into living cells. Journal of Cell Biology, 104(3):491–??, March 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/104/3/491.

Carnemolla:1989:TAF

[CBS^{+}89] B. Carnemolla, E. Balza, A. Siri, L. Zardi, M. R. Nicotra, A. Bigotti, and P. G. Natali. A tumor-associated fibronectin isoform generated by alternative splicing of messenger RNA precursors. Journal of Cell Biology, 108(3):1139–??, March 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/108/3/1139.

Chen:1988:SDC

[CBYK88] C. Chen, J. S. Bonifacino, L. C. Yuan, and R. D. Klausner. Selective degradation of T cell antigen receptor chains retained in a pre-Golgi compartment. Journal of Cell Biology, 107(6): 2149–??, December 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/107/6/2149.

Carboni:1985:LIC

[CC85a] J. M. Carboni and J. S. Condeelis. Ligand-induced changes in the location of actin, myosin, 95K (alpha-actinin), and 120K protein in amebae of Dictyostelium discoideum. Journal of Cell Biology, 100(6):1884–??, June 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/100/6/1884.

Caulfield:1985:HEA

[CC85b] J. P. Caulfield and C. M. Cianci. Human erythrocytes adhering to schistosomula of Schistosoma mansoni lyse and fail to transfer membrane components to the parasite. Journal of Cell Biology, 101(1):158–??, July 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/101/1/158.

Campbell:1988:PPN

[CC88] E. J. Campbell and M. A. Campbell. Pericellular proteolysis by neutrophils in the presence of proteinase inhibitors: effects of substrate opsonization. Journal of Cell Biology, 106 (3):667–??, March 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/106/3/667.

Ceriotti:1989:PTE

[CC89] A. Ceriotti and A. Colman. Protein transport from endoplasmic reticulum to the Golgi complex can occur during meiotic metaphase in Xenopus oocytes. Journal of Cell Biology, 109 (4):1439–??, October 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/4/1439.

Carboni:1988:SIC

[CCA^{+}88] J. M. Carboni, K. A. Conzelman, R. A. Adams, D. A. Kaiser, T. D. Pollard, and M. S. Mooseker. Structural and immunological characterization of the myosin-like 110-kD subunit of the intestinal microvillar 110K-calmodulin complex: evidence for discrete myosin head and calmodulin-binding domains. Journal of Cell Biology, 107(5):1749–??, November 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/107/5/1749.

Citi:1989:PMB

[CCBKJ89] S. Citi, R. A. Cross, C. R. Bagshaw, and J. Kendrick-Jones. Parallel modulation of brush border myosin conformation and enzyme activity induced by monoclonal antibodies. Journal of Cell Biology, 109(2):549–??, August 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/109/2/549.

Cameron:1986:CSP

[CCC86] R. S. Cameron, P. L. Cameron, and J. D. Castle. A common spectrum of polypeptides occurs in secretion granule membranes of different exocrine glands. Journal of Cell Biology, 103(4): 1299–??, October 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/103/4/1299.

Cadigan:1989:ICH

[CCC89] K. M. Cadigan, C. C. Chang, and T. Y. Chang. Isolation of Chinese hamster ovary cell lines expressing human acyl-coenzyme A/cholesterol acyltransferase activity. Journal of Cell Biology, 108(6):2201–??, June 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/108/6/2201.

Chuong:1987:SED

[CCE87] C. M. Chuong, K. L. Crossin, and G. M. Edelman. Sequential expression and differential function of multiple adhesion molecules during the formation of cerebellar cortical layers. Journal of Cell Biology, 104(2):331–??, February 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/104/2/331.

Connolly:1989:APK

[CCG89] T. Connolly, P. Collins, and R. Gilmore. Access of proteinase K to partially translocated nascent polypeptides in intact and detergent-solubilized membranes. Journal of Cell Biology, 108 (2):299–??, February 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/108/2/299.

Carlson:1986:NTA

[CCK86] S. S. Carlson, P. Caroni, and R. B. Kelly. A nerve terminal anchorage protein from electric organ. Journal of Cell Biology, 103(2):509–??, August 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/103/2/509.

Clowes:1988:ASM

[CCK^{+}88] A. W. Clowes, M. M. Clowes, O. Kocher, P. Ropraz, C. Chaponnier, and G. Gabbiani. Arterial smooth muscle cells in vivo: relationship between actin isoform expression and mitogenesis and their modulation by heparin. Journal of Cell Biology, 107 (5):1939–??, November 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/107/5/1939.

Castellot:1986:SDC

[CCL^{+}86] J. J. Castellot, J. Choay, J. C. Lormeau, M. Petitou, E. Sache, and M. J. Karnovsky. Structural determinants of the capacity of heparin to inhibit the proliferation of vascular smooth muscle cells. II. Evidence for a pentasaccharide sequence that contains a 3-O-sulfate group. Journal of Cell Biology, 102(5):1979–??, May 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/102/5/
1979.

Cartaud:1989:PPI

[CCLC89] A. Cartaud, J. C. Courvalin, M. A. Ludosky, and J. Cartaud. Presence of a protein immunologically related to lamin B in the postsynaptic membrane of Torpedo marmorata electrocyte. Journal of Cell Biology, 109(4):1745–??, October 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/109/4/1745.

Chen:1986:CEC

[CCM86] W. T. Chen, J. M. Chen, and S. C. Mueller. Coupled expression and colocalization of 140K cell adhesion molecules, fibronectin, and laminin during morphogenesis and cytodifferentiation of chick lung cells. Journal of Cell Biology, 103(3):1073–??, September 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/103/3/
1073.

Casella:1987:CZB

[CCMB87] J. F. Casella, S. W. Craig, D. J. Maack, and A. E. Brown. Cap Z(36/32), a barbed end actin-capping protein, is a component of the Z-line of skeletal muscle. Journal of Cell Biology, 105(1):371–??, July 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/1/371.

Cardarelli:1988:PEF

[CCP88] P. M. Cardarelli, I. N. Crispe, and M. D. Pierschbacher. Preferential expression of fibronectin receptors on immature thymocytes. Journal of Cell Biology, 106(6):2183–??, June 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/106/6/2183.

Costello:1986:CJF

[CCS^{+}86] B. Costello, C. Chadwick, A. Saito, A. Chu, A. Maurer, and S. Fleischer. Characterization of the junctional face membrane from terminal cisternae of sarcoplasmic reticulum. Journal of Cell Biology, 103(3):741–??, September 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/103/3/741.

Covault:1987:NOC

[CCS87] J. Covault, J. M. Cunningham, and J. R. Sanes. Neurite outgrowth on cryostat sections of innervated and denervated skeletal muscle. Journal of Cell Biology, 105(6):2479–??, December 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/105/6/
2479.

Cheng:1989:TDR

[CD89] N. Q. Cheng and J. F. Deatherage. Three-dimensional reconstruction of the Z disk of sectioned bee flight muscle. Journal of Cell Biology, 108(5):1761–??, May 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/108/5/1761.

Copeland:1986:AIH

[CDB^{+}86] C. S. Copeland, R. W. Doms, E. M. Bolzau, R. G. Webster, and A. Helenius. Assembly of influenza hemagglutinin trimers and its role in intracellular transport. Journal of Cell Biology, 103 (4):1179–??, October 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/103/4/1179.

Cambray-Deakin:1987:PMA

[CDB87] M. A. Cambray-Deakin and R. D. Burgoyne. Posttranslational modifications of alpha-tubulin: acetylated and detyrosinated forms in axons of rat cerebellum. Journal of Cell Biology, 104 (6):1569–??, June 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/104/6/1569.

Cosson:1989:LCP

[CdCP^{+}89] P. Cosson, I. de Curtis, J. Pouysségur, G. Griffiths, and J. Davoust. Low cytoplasmic pH inhibits endocytosis and transport from the trans-Golgi network to the cell surface. Journal of Cell Biology, 108(2):377–??, February 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/108/2/377.

Clapper:1985:IZR

[CDL^{+}85] D. L. Clapper, J. A. Davis, P. J. Lamothe, C. Patton, and D. Epel. Involvement of zinc in the regulation of pHi, motility, and acrosome reactions in sea urchin sperm. Journal of Cell Biology, 100(6):1817–??, June 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/100/6/1817.

Castagnola:1988:CEC

[CDMC88] P. Castagnola, B. Dozin, G. Moro, and R. Cancedda. Changes in the expression of collagen genes show two stages in chondrocyte differentiation in vitro. Journal of Cell Biology, 106(2): 461–??, February 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/106/2/461.

Crissman:1985:NPC

[CDTS85] H. A. Crissman, Z. Darzynkiewicz, R. A. Tobey, and J. A. Steinkamp. Normal and perturbed Chinese hamster ovary cells: correlation of DNA, RNA, and protein content by flow cytometry. Journal of Cell Biology, 101(1):141–??, July 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/101/1/141.

Chuong:1985:ECAa

[CE85a] C. M. Chuong and G. M. Edelman. Expression of cell-adhesion molecules in embryonic induction. I. Morphogenesis of nestling feathers. Journal of Cell Biology, 101(3):1009–??, September 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/101/3/
1009.

Chuong:1985:ECAb

[CE85b] C. M. Chuong and G. M. Edelman. Expression of cell-adhesion molecules in embryonic induction. II. Morphogenesis of adult feathers. Journal of Cell Biology, 101(3):1027–??, September 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/101/3/
1027.

Carey:1989:MAH

[CE89] D. J. Carey and D. M. Evans. Membrane anchoring of heparan sulfate proteoglycans by phosphatidylinositol and kinetics of synthesis of peripheral and detergent-solubilized proteoglycans in Schwann cells. Journal of Cell Biology, 108(5):1891–??, May 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/108/5/
1891.

Chun:1988:IMH

[CF88a] M. Chun and S. Falkenthal. Ifm(2)2 is a myosin heavy chain allele that disrupts myofibrillar assembly only in the indirect flight muscle of Drosophila melanogaster. Journal of Cell Biology, 107 (6):2613–??, December 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/107/6/2613.

Cortese:1988:MAG

[CF88b] J. D. Cortese and C. Frieden. Microheterogeneity of actin gels formed under controlled linear shear. Journal of Cell Biology, 107 (4):1477–??, October 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/107/4/1477.

Couchie:1985:MAP

[CFB^{+}85] D. Couchie, C. Fages, A. M. Bridoux, B. Rolland, M. Tardy, and J. Nunez. Microtubule-associated proteins and in vitro astrocyte differentiation. Journal of Cell Biology, 101(6):2095–??, December 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/101/6/
2095.

Carr:1989:NMP

[CFC89] C. Carr, G. D. Fischbach, and J. B. Cohen. A novel 87,000-Mr protein associated with acetylcholine receptors in Torpedo electric organ and vertebrate skeletal muscle. Journal of Cell Biology, 109(4):1753–??, October 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/109/4/1753.

Cole:1986:HBD

[CG86a] G. J. Cole and L. Glaser. A heparin-binding domain from N-CAM is involved in neural cell-substratum adhesion. Journal of Cell Biology, 102(2):403–??, February 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/102/2/403.

Connolly:1986:FFR

[CG86b] T. Connolly and R. Gilmore. Formation of a functional ribosomemembrane junction during translocation requires the participation of a GTP-binding protein. Journal of Cell Biology, 103(6): 2253–??, December 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/103/6/2253.

Cutler:1986:MMBa

[CG86c] D. F. Cutler and H. Garoff. Mutants of the membrane-binding region of Semliki Forest virus E2 protein. I. Cell surface transport and fusogenic activity. Journal of Cell Biology, 102(3):889–??, March 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
102/3/889.

Choi:1989:ECA

[CG89] Y. S. Choi and B. Gumbiner. Expression of cell adhesion molecule E-cadherin in Xenopus embryos begins at gastrulation and predominates in the ectoderm. Journal of Cell Biology, 108(6):2449– ??, June 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
108/6/2449.

Chatila:1989:CSI

[CGA89] T. A. Chatila, R. S. Geha, and M. A. Arnaout. Constitutive and stimulus-induced phosphorylation of CD11/CD18 leukocyte adhesion molecules. Journal of Cell Biology, 109(6):3435–??, December 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
109/6/3435.

Cardelli:1986:LED

[CGD86] J. A. Cardelli, G. S. Golumbeski, and R. L. Dimond. Lysosomal enzymes in Dictyostelium discoideum are transported to lysosomes at distinctly different rates. Journal of Cell Biology, 102(4):1264–??, April 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/102/4/1264.

Carpentier:1986:IPF

[CGV^{+}86] J. L. Carpentier, H. Gazzano, E. Van Obberghen, M. Fehlmann, P. Freychet, and L. Orci. Intracellular pathway followed by the insulin receptor covalently coupled to 125I-photoreactive insulin during internalization and recycling. Journal of Cell Biology, 102(3):989–??, March 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/102/3/989.

Connor:1985:ECD

[CH85] J. Connor and L. Huang. Efficient cytoplasmic delivery of a fluorescent dye by pH-sensitive immunoliposomes. Journal of Cell Biology, 101(2):582–??, August 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/101/2/582.

Clegg:1987:HAM

[CH87] C. H. Clegg and S. D. Hauschka. Heterokaryon analysis of muscle differentiation: regulation of the postmitotic state. Journal of Cell Biology, 105(2):937–??, August 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/105/2/937.

Chantler:1985:CDA

[Cha85] P. D. Chantler. Calcium-dependent association of a protein complex with the lymphocyte plasma membrane: probable identity with calmodulin-calcineurin. Journal of Cell Biology, 101 (1):207–??, July 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/101/1/207.

Carron:1986:REC

[CHD^{+}86] C. P. Carron, S. Y. Hwo, J. Dingus, D. M. Benson, I. Meza, and J. Bryan. A re-evaluation of cytoplasmic gelsolin localization. Journal of Cell Biology, 102(1):237–??, January 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/102/1/237.

Cooke:1987:ICP

[CHE87] C. A. Cooke, M. M. Heck, and W. C. Earnshaw. The inner centromere protein (INCENP) antigens: movement from inner centromere to midbody during mitosis. Journal of Cell Biology, 105 (5):2053–??, November 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/5/2053.

Castellani:1989:ECD

[CHFA89] L. Castellani, P. M. Hardwicke, and C. Franzini-Armstrong. Effect of Ca^{2+} on the dimeric structure of scallop sarcoplasmic reticulum. Journal of Cell Biology, 108(2):511–??, February 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/108/2/
511.

Crossin:1986:SRE

[CHG^{+}86] K. L. Crossin, S. Hoffman, M. Grumet, J. P. Thiery, and G. M. Edelman. Site-restricted expression of cytotactin during development of the chicken embryo. Journal of Cell Biology, 102 (5):1917–??, May 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/102/5/1917.

Cockcroft:1987:TGP

[CHG87] S. Cockcroft, T. W. Howell, and B. D. Gomperts. Two G-proteins act in series to control stimulus-secretion coupling in mast cells: use of neomycin to distinguish between G-proteins controlling polyphosphoinositide phosphodiesterase and exocytosis. Journal of Cell Biology, 105(6):2745–??, December 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/105/6/2745.

Chen:1985:DCS

[CHH^{+}85] W. T. Chen, E. Hasegawa, T. Hasegawa, C. Weinstock, and K. M. Yamada. Development of cell surface linkage complexes in cultured fibroblasts. Journal of Cell Biology, 100(4):1103–??, April 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
100/4/1103.

Clegg:1988:AAS

[CHH^{+}88] D. O. Clegg, J. C. Helder, B. C. Hann, D. E. Hall, and L. F. Reichardt. Amino acid sequence and distribution of mRNA encoding a major skeletal muscle laminin binding protein: an extracellular matrix-associated protein with an unusual COOHterminal polyaspartate domain. Journal of Cell Biology, 107 (2):699–??, August 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/107/2/699.

Chitnis:1986:APP

[CHK^{+}86] P. R. Chitnis, E. Harel, B. D. Kohorn, E. M. Tobin, and J. P. Thornber. Assembly of the precursor and processed lightharvesting chlorophyll a/b protein of Lemna into the lightharvesting complex II of barley etiochloroplasts. Journal of Cell Biology, 102(3):982–??, March 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/102/3/982.

Collier:1988:UBA

[CHLS88] N. C. Collier, J. Heuser, M. A. Levy, and M. J. Schlesinger. Ultrastructural and biochemical analysis of the stress granule in chicken embryo fibroblasts. Journal of Cell Biology, 106 (4):1131–??, April 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/106/4/1131.

Coleman:1987:BSB

[CHM^{+}87] T. R. Coleman, A. S. Harris, S. M. Mische, M. S. Mooseker, and J. S. Morrow. Beta spectrin bestows protein 4.1 sensitivity on spectrin-actin interactions. Journal of Cell Biology, 104 (3):519–??, March 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/104/3/519.

Carter:1986:TPS

[CHMB86] C. Carter, A. R. Howlett, G. S. Martin, and M. J. Bissell. The tyrosine phosphorylation substrate p36 is developmentally regulated in embryonic avian limb and is induced in cell culture. Journal of Cell Biology, 103(5):2017–??, November 1986. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/103/5/2017.

Cook:1985:DRM

[CHS85] K. S. Cook, C. R. Hunt, and B. M. Spiegelman. Developmentally regulated mRNAs in 3T3-adipocytes: analysis of transcriptional control. Journal of Cell Biology, 100(2):514–??, February 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/100/2/
514.

Croze:1989:EMG

[CIK^{+}89] E. Croze, I. E. Ivanov, G. Kreibich, M. Adesnik, D. D. Sabatini, and M. G. Rosenfeld. Endolyn-78, a membrane glycoprotein present in morphologically diverse components of the endosomal and lysosomal compartments: implications for lysosome biogenesis. Journal of Cell Biology, 108(5):1597–??, May 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/108/5/1597.

Crabb:1985:VRE

[CJ85] J. H. Crabb and R. C. Jackson. In vitro reconstitution of exocytosis from plasma membrane and isolated secretory vesicles. Journal of Cell Biology, 101(6):2263–??, December 1985. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/101/6/2263.

Chang:1986:SSC

[CJ86] A. Chang and J. D. Jamieson. Stimulus-secretion coupling in the developing exocrine pancreas: secretory responsiveness to cholecystokinin. Journal of Cell Biology, 103(6):2353–??, December 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/103/6/
2353.

Colman:1985:MMX

[CJH85] A. Colman, E. A. Jones, and J. Heasman. Meiotic maturation in Xenopus oocytes: a link between the cessation of protein secretion and the polarized disappearance of Golgi apparati. Journal of Cell Biology, 101(1):313–??, July 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/101/1/313.

Caron:1985:ATS

[CJK85] J. M. Caron, A. L. Jones, and M. W. Kirschner. Autoregulation of tubulin synthesis in hepatocytes and fibroblasts. Journal of Cell Biology, 101(5):1763–??, November 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/101/5/1763.

Cheek:1989:SMC

[CJO^{+}89] T. R. Cheek, T. R. Jackson, A. J. O’Sullivan, R. B. Moreton, M. J. Berridge, and R. D. Burgoyne. Simultaneous measurements of cytosolic calcium and secretion in single bovine adrenal chromaffin cells by fluorescent imaging of fura-2 in cocultured cells. Journal of Cell Biology, 109(3):1219–??, September 1989. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/109/3/1219.

Chaponnier:1986:AFS

[CJY86] C. Chaponnier, P. A. Janmey, and H. L. Yin. The actin filamentsevering domain of plasma gelsolin. Journal of Cell Biology, 103 (4):1473–??, October 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/103/4/1473.

Cheresh:1986:DGD

[CK86] D. A. Cheresh and F. G. Klier. Disialoganglioside GD2 distributes preferentially into substrate-associated microprocesses on human melanoma cells during their attachment to fibronectin. Journal of Cell Biology, 102(5):1887–??, May 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/102/5/1887.

Cowin:1985:CDP

[CKF85] P. Cowin, H. P. Kapprell, and W. W. Franke. The complement of desmosomal plaque proteins in different cell types. Journal of Cell Biology, 101(4):1442–??, October 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/101/4/1442.

Coulombe:1989:EKK

[CKF89] P. A. Coulombe, R. Kopan, and E. Fuchs. Expression of keratin K14 in the epidermis and hair follicle: insights into complex programs of differentiation. Journal of Cell Biology, 109(5): 2295–??, November 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/5/2295.

Chapman:1985:GRA

[CKR85a] A. B. Chapman, D. M. Knight, and G. M. Ringold. Glucocorticoid regulation of adipocyte differentiation: hormonal triggering of the developmental program and induction of a differentiationdependent gene. Journal of Cell Biology, 101(4):1227–??, October 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/101/4/
1227.

Christiansson:1985:LMS

[CKR^{+}85b] A. Christiansson, F. A. Kuypers, B. Roelofsen, J. A. Op den Kamp, and L. L. van Deenen. Lipid molecular shape affects erythrocyte morphology: a study involving replacement of native phosphatidylcholine with different species followed by treatment of cells with sphingomyelinase C or phospholipase A2. Journal of Cell Biology, 101(4):1455–??, October 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/101/4/1455.

Carrino:1985:EHS

[CL85a] J. J. Carrino and T. G. Laffler. The effect of heat shock on the cell cycle regulation of tubulin expression in Physarum polycephalum. Journal of Cell Biology, 100(2):642–??, February 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/100/2/
642.

Chen:1985:CEE

[CL85b] J. M. Chen and C. D. Little. Cells that emerge from embryonic explants produce fibers of type IV collagen. Journal of Cell Biology, 101(4):1175–??, October 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/101/4/1175.

Carrino:1986:TAT

[CL86] J. J. Carrino and T. G. Laffler. Transcription of alpha-tubulin and histone H4 genes begins at the same point in the Physarum cell cycle. Journal of Cell Biology, 102(5):1666–??, May 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/102/5/1666.

Clayton:1987:IFB

[Cla87] C. E. Clayton. Import of fructose bisphosphate aldolase into the glycosomes of Trypanosoma brucei. Journal of Cell Biology, 105 (6):2649–??, December 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/6/2649.

Cole:1986:TLH

[CLC^{+}86] G. J. Cole, A. Loewy, N. V. Cross, R. Akeson, and L. Glaser. Topographic localization of the heparin-binding domain of the neural cell adhesion molecule N-CAM. Journal of Cell Biology, 103 (5):1739–??, November 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/103/5/1739.

Cleveland:1987:MHR

[Cle87] D. W. Cleveland. The multitubulin hypothesis revisited: what have we learned? Journal of Cell Biology, 104(3):381–??, March 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/104/3/
381.

Cooper:1988:LMG

[CLF^{+}88] J. A. Cooper, D. J. Loftus, C. Frieden, J. Bryan, and E. L. Elson. Localization and mobility of gelsolin in cells. Journal of Cell Biology, 106(4):1229–??, April 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/106/4/1229.

Clegg:1987:GFC

[CLOH87] C. H. Clegg, T. A. Linkhart, B. B. Olwin, and S. D. Hauschka. Growth factor control of skeletal muscle differentiation: commitment to terminal differentiation occurs in G1 phase and is repressed by fibroblast growth factor. Journal of Cell Biology, 105(2):949–??, August 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/2/949.

Coleman:1985:EAF

[CM85] T. R. Coleman and M. S. Mooseker. Effects of actin filament cross-linking and filament length on actin-myosin interaction. Journal of Cell Biology, 101(5):1850–??, November 1985. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/101/5/1850.

Cande:1986:PUA

[CM86a] W. Z. Cande and K. McDonald. Physiological and ultrastructural analysis of elongating mitotic spindles reactivated in vitro. Journal of Cell Biology, 103(2):593–??, August 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/103/2/593.

Chasis:1986:EMD

[CM86b] J. A. Chasis and N. Mohandas. Erythrocyte membrane deformability and stability: two distinct membrane properties that are independently regulated by skeletal protein associations. Journal of Cell Biology, 103(2):343–??, August 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/103/2/343.

Clarke:1986:TSN

[CM86c] H. J. Clarke and Y. Masui. Transformation of sperm nuclei to metaphase chromosomes in the cytoplasm of maturing oocytes of the mouse. Journal of Cell Biology, 102(3):1039–??, March 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/102/3/

Clarke:1987:DDR

[CM87a] H. J. Clarke and Y. Masui. Dose-dependent relationship between oocyte cytoplasmic volume and transformation of sperm nuclei to metaphase chromosomes. Journal of Cell Biology, 104 (4):831–??, January 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/104/4/831.

Connor:1987:CAJ

[CM87b] E. A. Connor and U. J. McMahan. Cell accumulation in the junctional region of denervated muscle. Journal of Cell Biology, 104(1):109–??, January 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/104/1/109.

Conzelman:1987:KPC

[CM87c] K. A. Conzelman and M. S. Mooseker. The 110-kD proteincalmodulin complex of the intestinal microvillus is an actinactivated MgATPase. Journal of Cell Biology, 105(1):313–??, July 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/105/1/
313.

Cain:1988:CRS

[CM88a] C. C. Cain and R. F. Murphy. A chloroquine-resistant Swiss 3T3 cell line with a defect in late endocytic acidification. Journal of Cell Biology, 106(2):269–??, February 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/106/2/269.

Chen:1988:ALC

[CM88b] T. Y. Chen and E. M. Merisko. Annulate lamellae: comparison of antigenic epitopes of annulate lamellae membranes with the nuclear envelope. Journal of Cell Biology, 107(4):1299–??, October 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/107/4/
1299.

Cramer:1986:EHS

[CMB^{+}86] E. B. Cramer, L. C. Milks, M. J. Brontoli, G. K. Ojakian, S. D. Wright, and H. J. Showell. Effect of human serum and some of its components on neutrophil adherence and migration across an epithelium. Journal of Cell Biology, 102(5):1868–??, May 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/102/5/
1868.

Capony:1987:PGP

[CMB^{+}87] F. Capony, M. Morisset, A. J. Barrett, J. P. Capony, P. Broquet, F. Vignon, M. Chambon, P. Louisot, and H. Rochefort. Phosphorylation, glycosylation, and proteolytic activity of the 52-kD estrogen-induced protein secreted by MCF7 cells. Journal of Cell Biology, 104(2):253–??, February 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/104/2/253.

Castagnola:1986:TXC

[CMDCC86] P. Castagnola, G. Moro, F. Descalzi-Cancedda, and R. Cancedda. Type X collagen synthesis during in vitro development of chick embryo tibial chondrocytes. Journal of Cell Biology, 102(6):2310–??, June 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/102/6/2310.

Cutler:1986:MMBb

[CMG86] D. F. Cutler, P. Melancon, and H. Garoff. Mutants of the membrane-binding region of Semliki Forest virus E2 protein. II. Topology and membrane binding. Journal of Cell Biology, 102 (3):902–??, March 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/102/3/902.

Covault:1986:MFC

[CMGS86] J. Covault, J. P. Merlie, C. Goridis, and J. R. Sanes. Molecular forms of N-CAM and its RNA in developing and denervated skeletal muscle. Journal of Cell Biology, 102(3):731–??, March 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/102/3/
731.

Cardell:1985:SSS

[CMHC85] R. R. Cardell, J. E. Michaels, J. T. Hung, and E. L. Cardell. SERGE, the subcellular site of initial hepatic glycogen deposition in the rat: a radioautographic and cytochemical study. Journal of Cell Biology, 101(1):201–??, July 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/101/1/201.

Cowan:1987:LDP

[CMK87] A. E. Cowan, D. G. Myles, and D. E. Koppel. Lateral diffusion of the PH-20 protein on guinea pig sperm: evidence that barriers to diffusion maintain plasma membrane domains in mammalian sperm. Journal of Cell Biology, 104(4):917–??, January 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/104/4/
917.

Cox:1985:ATH

[CML85] J. V. Cox, R. T. Moon, and E. Lazarides. Anion transporter: highly cell-type-specific expression of distinct polypeptides and transcripts in erythroid and nonerythroid cells. Journal of Cell Biology, 100(5):1548–??, May 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/100/5/1548.

Carraro:1985:CDR

[CMM^{+}85] U. Carraro, D. Morale, I. Mussini, S. Lucke, M. Cantini, R. Betto, C. Catani, L. Dalla Libera, D. Danieli Betto, and D. Noventa. Chronic denervation of rat hemidiaphragm: maintenance of fiber heterogeneity with associated increasing uniformity of myosin isoforms. Journal of Cell Biology, 100(1):161–??, January 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/100/1/161.

Chiquet:1988:AEL

[CMNB88] M. Chiquet, L. Masuda-Nakagawa, and K. Beck. Attachment to an endogenous laminin-like protein initiates sprouting by leech neurons. Journal of Cell Biology, 107(3):1189–??, September 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/107/3/
1189.

Chiu:1986:MAB

[CMP86] A. Y. Chiu, W. D. Matthew, and P. H. Patterson. A monoclonal antibody that blocks the activity of a neurite regenerationpromoting factor: studies on the binding site and its localization in vivo. Journal of Cell Biology, 103(4):1383–??, October 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/103/4/

Corless:1987:GBS

[CMR^{+}87] C. L. Corless, M. M. Matzuk, T. V. Ramabhadran, A. Krichevsky, and I. Boime. Gonadotropin beta subunits determine the rate of assembly and the oligosaccharide processing of hormone dimer in transfected cells. Journal of Cell Biology, 104(5):1173–??, May 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/104/5/
1173.

Cancedda:1988:DRS

[CMT^{+}88] F. Descalzi Cancedda, P. Manduca, C. Tacchetti, P. Fossa, R. Quarto, and R. Cancedda. Developmentally regulated synthesis of a low molecular weight protein (Ch 21) by differentiating chondrocytes. Journal of Cell Biology, 107(6):2455–??, December 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/107/6/
2455.

Cooper:1985:CDP

[CMTC85] N. G. Cooper, B. J. McLaughlin, E. A. Tallant, and W. Y. Cheung. Calmodulin-dependent protein phosphatase: immunocytochemical localization in chick retina. Journal of Cell Biology, 101 (4):1212–??, October 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/101/4/1212.

Chen:1985:ITL

[CMW^{+}85] J. W. Chen, T. L. Murphy, M. C. Willingham, I. Pastan, and J. T. August. Identification of two lysosomal membrane glycoproteins. Journal of Cell Biology, 101(1):85–??, July 1985. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/101/1/85.

Clawson:1987:EHR

[CMW87] G. A. Clawson, J. R. MacDonald, and C. H. Woo. Early hypomethylation of 2’-O-ribose moieties in hepatocyte cytoplasmic ribosomal RNA underlies the protein synthetic defect produced by CCl4. Journal of Cell Biology, 105(2):705–??, August 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/105/2/

Cleves:1989:MSG

[CNB89] A. E. Cleves, P. J. Novick, and V. A. Bankaitis. Mutations in the SAC1 gene suppress defects in yeast Golgi and yeast actin function. Journal of Cell Biology, 109(6):2939–??, December 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/109/6/
2939.

Critchley:1985:FTT

[CNHF85] D. R. Critchley, P. G. Nelson, W. H. Habig, and P. H. Fishman. Fate of tetanus toxin bound to the surface of primary neurons in culture: evidence for rapid internalization. Journal of Cell Biology, 100(5):1499–??, May 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/100/5/1499.

Church:1986:MBC

[CNL86] K. Church, R. B. Nicklas, and H. P. Lin. Micromanipulated bivalents can trigger mini-spindle formation in Drosophila melanogaster spermatocyte cytoplasm. Journal of Cell Biology, 103(6):2765–??, December 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/103/6/2765.

Collin-Osdoby:1985:CPC

[COA85] P. Collin-Osdoby and W. S. Adair. Characterization of the purified Chlamydomonas minus agglutinin. Journal of Cell Biology, 101(3):1144–??, September 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/101/3/1144.

Coller:1986:AAA

[Col86] B. S. Coller. Activation affects access to the platelet receptor for adhesive glycoproteins. Journal of Cell Biology, 103 (2):451–??, August 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/103/2/451.

Canipari:1987:MOG

[COMS87] R. Canipari, M. L. O’Connell, G. Meyer, and S. Strickland. Mouse ovarian granulosa cells produce urokinase-type plasminogen activator, whereas the corresponding rat cells produce tissuetype plasminogen activator. Journal of Cell Biology, 105(2):

977–??, August 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/2/977.

Cooper:1987:ECP

[Coo87] J. A. Cooper. Effects of cytochalasin and phalloidin on actin. Journal of Cell Biology, 105(4):1473–??, October 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/105/4/1473.

Couchman:1987:HDB

[Cou87] J. R. Couchman. Heterogeneous distribution of a basement membrane heparan sulfate proteoglycan in rat tissues. Journal of Cell Biology, 105(4):1901–??, October 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/105/4/1901.

Cardarelli:1987:IFR

[CP87a] P. M. Cardarelli and M. D. Pierschbacher. Identification of fibronectin receptors on T lymphocytes. Journal of Cell Biology, 105(1):499–??, July 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/1/499.

Chang:1987:CRA

[CP87b] X. J. Chang and G. Piperno. Cross-reactivity of antibodies specific for flagellar tektin and intermediate filament subunits. Journal of Cell Biology, 104(6):1563–??, June 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/104/6/1563.

Castellot:1989:HSI

[CPC^{+}89] J. J. Castellot, L. A. Pukac, B. L. Caleb, T. C. Wright, and M. J. Karnovsky. Heparin selectively inhibits a protein kinase C-dependent mechanism of cell cycle progression in calf aortic smooth muscle cells. Journal of Cell Biology, 109(6): 3147–??, December 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/6/3147.

Claudio:1989:FTT

[CPG^{+}89] T. Claudio, H. L. Paulson, W. N. Green, A. F. Ross, D. S. Hartman, and D. Hayden. Fibroblasts transfected with Torpedo acetylcholine receptor beta-, gamma-, and delta-subunit cDNAs express functional receptors when infected with a retroviral alpha recombinant. Journal of Cell Biology, 108(6):2277–??, June 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/108/6/
2277.

Cheresh:1986:DGG

[CPHM86] D. A. Cheresh, M. D. Pierschbacher, M. A. Herzig, and K. Mujoo. Disialogangliosides GD2 and GD3 are involved in the attachment of human melanoma and neuroblastoma cells to extracellular matrix proteins. Journal of Cell Biology, 102(3):688–??, March 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/102/3/
688.

Craig:1987:SCA

[CPKJ87] R. Craig, R. Padrón, and J. Kendrick-Jones. Structural changes accompanying phosphorylation of tarantula muscle myosin filaments. Journal of Cell Biology, 105(3):1319–??, September 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/105/3/1319.

Cowan:1986:SEI

[CPM86] A. E. Cowan, P. Primakoff, and D. G. Myles. Sperm exocytosis increases the amount of PH-20 antigen on the surface of guinea pig sperm. Journal of Cell Biology, 103(4):1289–??, October 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/103/4/
1289.

Cheresh:1987:AGA

[CPP^{+}87] D. A. Cheresh, R. Pytela, M. D. Pierschbacher, F. G. Klier, E. Ruoslahti, and R. A. Reisfeld. An Arg–Gly–Asp–directed receptor on the surface of human melanoma cells exists in an divalent cation-dependent functional complex with the disialoganglioside GD2. Journal of Cell Biology, 105(3):1163–??, September 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/105/3/
1163.

Cassimeris:1988:RTO

[CPS88] L. Cassimeris, N. K. Pryer, and E. D. Salmon. Real-time observations of microtubule dynamic instability in living cells. Journal of Cell Biology, 107(6):2223–??, December 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/107/6/2223.

Cheng:1985:PCO

[CR85] T. P. Cheng and T. S. Reese. Polarized compartmentalization of organelles in growth cones from developing optic tectum. Journal of Cell Biology, 101(4):1473–??, October 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/101/4/1473.

Conrad:1986:PAF

[CR86] G. W. Conrad and T. J. Rink. Platelet activating factor raises intracellular calcium ion concentration in macrophages. Journal of Cell Biology, 103(2):439–??, August 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/103/2/439.

Carlson:1987:MUI

[CR87] N. Carlson and M. Rechsteiner. Microinjection of ubiquitin: intracellular distribution and metabolism in HeLa cells maintained under normal physiological conditions. Journal of Cell Biology, 104(3):537–??, March 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/104/3/537.

Cardwell:1988:ERD

[CR88a] M. C. Cardwell and L. H. Rome. Evidence that an RGDdependent receptor mediates the binding of oligodendrocytes to a novel ligand in a glial-derived matrix. Journal of Cell Biology, 107(4):1541–??, October 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/107/4/1541.

Cardwell:1988:RCP

[CR88b] M. C. Cardwell and L. H. Rome. RGD-containing peptides inhibit the synthesis of myelin-like membrane by cultured oligodendrocytes. Journal of Cell Biology, 107(4):1551–??, October 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/107/4/

Chicoine:1987:SHA

[CRC^{+}87] L. G. Chicoine, R. Richman, R. G. Cook, M. A. Gorovsky, and C. D. Allis. A single histone acetyltransferase from Tetrahymena macronuclei catalyzes deposition-related acetylation of free histones and transcription-related acetylation of nucleosomal histones. Journal of Cell Biology, 105(1):127–??, July 1987. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/105/1/127.

Carrington:1988:ALC

[CRF^{+}88] J. L. Carrington, A. B. Roberts, K. C. Flanders, N. S. Roche, and A. H. Reddi. Accumulation, localization, and compartmentation of transforming growth factor beta during endochondral bone development. Journal of Cell Biology, 107(5):1969–??, November 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/107/5/
1969.

Chasis:1988:STG

[CRJM88] J. A. Chasis, M. E. Reid, R. H. Jensen, and N. Mohandas. Signal transduction by glycophorin A: role of extracellular and cytoplasmic domains in a modulatable process. Journal of Cell Biology, 107(4):1351–??, October 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/107/4/1351.

Ciment:1986:NIF

[CRLW86] G. Ciment, A. Ressler, P. C. Letourneau, and J. A. Weston. A novel intermediate filament-associated protein, NAPA-73, that binds to different filament types at different stages of nervous system development. Journal of Cell Biology, 102(1):246–??, January 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/102/1/
246.

Carlson:1987:MUC

[CRR87a] N. Carlson, S. Rogers, and M. Rechsteiner. Microinjection of ubiquitin: changes in protein degradation in HeLa cells subjected to heat-shock. Journal of Cell Biology, 104(3):547–??, March 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/104/3/

Chang:1987:ENA

[CRR87b] S. Chang, F. G. Rathjen, and J. A. Raper. Extension of neurites on axons is impaired by antibodies against specific neural cell surface glycoproteins. Journal of Cell Biology, 104(2):355–??, February 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/104/2/
355.

Carey:1987:EIP

[CRT87] D. J. Carey, C. M. Rafferty, and M. S. Todd. Effects of inhibition of proteoglycan synthesis on the differentiation of cultured rat Schwann cells. Journal of Cell Biology, 105(2):1013–??, August 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/105/2/
1013.

Cantino:1986:RMC

[CS86a] M. Cantino and J. Squire. Resting myosin cross-bridge configuration in frog muscle thick filaments. Journal of Cell Biology, 102 (2):610–??, February 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/102/2/610.

Collier:1986:DSH

[CS86b] N. C. Collier and M. J. Schlesinger. The dynamic state of heat shock proteins in chicken embryo fibroblasts. Journal of Cell Biology, 103(4):1495–??, October 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/103/4/1495.

Cooper:1986:MCF

[CS86c] M. S. Cooper and M. Schliwa. Motility of cultured fish epidermal cells in the presence and absence of direct current electric fields. Journal of Cell Biology, 102(4):1384–??, April 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/102/4/1384.

Covault:1986:DCS

[CS86d] J. Covault and J. R. Sanes. Distribution of N-CAM in synaptic and extrasynaptic portions of developing and adult skeletal muscle. Journal of Cell Biology, 102(3):716–??, March 1986. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/102/3/716.

Choi:1987:FEC

[CS87] A. H. Choi and C. H. Siu. Filopodia are enriched in a cell cohesion molecule of Mr 80,000 and participate in cell–cell contact formation in Dictyostelium discoideum. Journal of Cell Biology, 104(5):1375–??, May 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/104/5/1375.

Campbell:1988:KIP

[CS88a] S. C. Campbell and N. B. Schwartz. Kinetics of intracellular processing of chondroitin sulfate proteoglycan core protein and other matrix components. Journal of Cell Biology, 106(6):2191–??, June 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/106/6/
2191.

Caroni:1988:TMP

[CS88b] P. Caroni and M. E. Schwab. Two membrane protein fractions from rat central myelin with inhibitory properties for neurite growth and fibroblast spreading. Journal of Cell Biology, 106 (4):1281–??, April 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/106/4/1281.

Chun:1988:FLM

[CS88c] J. J. Chun and C. J. Shatz. A fibronectin-like molecule is present in the developing cat cerebral cortex and is correlated with subplate neurons. Journal of Cell Biology, 106(3):857–??, March 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/106/3/
857.

Cummings:1988:PKA

[CS88d] A. Cummings and J. Sommerville. Protein kinase activity associated with stored messenger ribonucleoprotein particles of Xenopus oocytes. Journal of Cell Biology, 107(1):45–??, July 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/107/1/45.

Caretta:1989:VCN

[CS89] A. Caretta and H. Saibil. Visualization of cyclic nucleotide binding sites in the vertebrate retina by fluorescence microscopy. Journal of Cell Biology, 108(4):1517–??, April 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/108/4/1517.

Cabib:1989:CSA

[CSBS89] E. Cabib, A. Sburlati, B. Bowers, and S. J. Silverman. Chitin synthase 1, an auxiliary enzyme for chitin synthesis in Saccharomyces cerevisiae. Journal of Cell Biology, 108(5):1665–??, May 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/108/5/
1665.

Colbaugh:1989:ILT

[CSD89] P. A. Colbaugh, M. Stookey, and R. K. Draper. Impaired lysosomes in a temperature-sensitive mutant of Chinese hamster ovary cells. Journal of Cell Biology, 108(6):2211–??, June 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/108/6/
2211.

Chaney:1989:LCG

[CSFS89] W. Chaney, S. Sundaram, N. Friedman, and P. Stanley. The Lec4A CHO glycosylation mutant arises from miscompartmentalization of a Golgi glycosyltransferase. Journal of Cell Biology, 109(5):2089–??, November 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/109/5/2089.

Cole:1985:CSA

[CSG85] G. J. Cole, D. Schubert, and L. Glaser. Cell-substratum adhesion in chick neural retina depends upon protein-heparan sulfate interactions. Journal of Cell Biology, 100(4):1192–??, April 1985. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/100/4/1192.

Chang:1988:MIC

[CSGC88] L. Y. Chang, J. W. Slot, H. J. Geuze, and J. D. Crapo. Molecular immunocytochemistry of the CuZn superoxide dismutase in rat hepatocytes. Journal of Cell Biology, 107(6):2169–??, December 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/107/6/

Capetanaki:1989:OVG

[CSH89] Y. Capetanaki, S. Smith, and J. P. Heath. Overexpression of the vimentin gene in transgenic mice inhibits normal lens cell differentiation. Journal of Cell Biology, 109(4):1653–??, October 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/109/4/
1653.

Charonis:1988:NSP

[CSK^{+}88] A. S. Charonis, A. P. Skubitz, G. G. Koliakos, L. A. Reger, J. Dege, A. M. Vogel, R. Wohlhueter, and L. T. Furcht. A novel synthetic peptide from the B1 chain of laminin with heparinbinding and cell adhesion-promoting activities. Journal of Cell Biology, 107(3):1253–??, September 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/107/3/1253.

Cox:1987:EAT

[CSL87] J. V. Cox, J. H. Stack, and E. Lazarides. Erythroid anion transporter assembly is mediated by a developmentally regulated recruitment onto a preassembled membrane cytoskeleton. Journal of Cell Biology, 105(3):1405–??, September 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/105/3/1405.

Cajot:1989:MCE

[CSM^{+}89] J. F. Cajot, W. D. Schleuning, R. L. Medcalf, J. Bamat, J. Testuz, L. Liebermann, and B. Sordat. Mouse L cells expressing human prourokinase-type plasminogen activator: effects on extracellular matrix degradation and invasion. Journal of Cell Biology, 109(2):915–??, August 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/2/915.

Castellani:1986:THC

[CSR^{+}86] P. Castellani, A. Siri, C. Rosellini, E. Infusini, L. Borsi, and L. Zardi. Transformed human cells release different fibronectin variants than do normal cells. Journal of Cell Biology, 103(5): 1671–??, November 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/103/5/1671.

Carson:1988:VSP

[CTJG88] D. D. Carson, J. P. Tang, J. Julian, and S. R. Glasser. Vectorial secretion of proteoglycans by polarized rat uterine epithelial cells. Journal of Cell Biology, 107(6):2425–??, December 1988. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/107/6/2425.

Carey:1986:SCM

[CTR86] D. J. Carey, M. S. Todd, and C. M. Rafferty. Schwann cell myelination: induction by exogenous basement membrane-like extracellular matrix. Journal of Cell Biology, 102(6):2254–??, June 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/102/6/
2254.

Chen:1987:PGI

[CTS^{+}87] C. S. Chen, P. Thiagarajan, S. M. Schwartz, J. M. Harlan, and R. L. Heimark. The platelet glycoprotein IIb/IIIa-like protein in human endothelial cells promotes adhesion but not initial attachment to extracellular matrix. Journal of Cell Biology, 105 (4):1885–??, October 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/4/1885.

Charonis:1986:ILS

[CTSF86] A. S. Charonis, E. C. Tsilibary, T. Saku, and H. Furthmayr. Inhibition of laminin self-assembly and interaction with type IV collagen by antibodies to the terminal domain of the long arm. Journal of Cell Biology, 103(5):1689–??, November 1986. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/103/5/1689.

Charonis:1985:BLT

[CTYF85] A. S. Charonis, E. C. Tsilibary, P. D. Yurchenco, and H. Furthmayr. Binding of laminin to type IV collagen: a morphological study. Journal of Cell Biology, 100(6):1848–??, June 1985. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/100/6/1848.

Cassab:1987:IED

[CV87a] G. I. Cassab and J. E. Varner. Immunocytolocalization of extensin in developing soybean seed coats by immunogold-silver staining and by tissue printing on nitrocellulose paper. Journal of Cell Biology, 105(6):2581–??, December 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/105/6/2581.

Collins:1987:TDR

[CV87b] C. A. Collins and R. B. Vallee. Temperature-dependent reversible assembly of taxol-treated microtubules. Journal of Cell Biology, 105(6):2847–??, December 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/105/6/2847.

Capps:1989:ECM

[CVWZ89] G. G. Capps, M. Van Kampen, C. L. Ward, and M. C. Zúñiga. Endocytosis of the class I major histocompatibility antigen via a phorbol myristate acetate-inducible pathway is a cell-specific phenomenon and requires the cytoplasmic domain. Journal of Cell Biology, 108(4):1317–??, April 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/108/4/1317.

Carlson:1987:NTA

[CW87] S. S. Carlson and T. N. Wight. Nerve terminal anchorage protein 1 (TAP-1) is a chondroitin sulfate proteoglycan: biochemical and electron microscopic characterization. Journal of Cell Biology, 105(6):3075–??, December 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/105/6/3075.

Cato:1988:MRT

[CW88] A. C. Cato and J. Weinmann. Mineralocorticoid regulation of transcription of transfected mouse mammary tumor virus DNA in cultured kidney cells. Journal of Cell Biology, 106(6):2119–??, June 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/106/6/
2119.

Chappell:1989:GFY

[CW89] T. G. Chappell and G. Warren. A galactosyltransferase from the fission yeast Schizosaccharomyces pombe. Journal of Cell Biology, 109(6):2693–??, December 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/109/6/2693.

Czaja:1989:VVA

[CWF^{+}89] M. J. Czaja, F. R. Weiner, K. C. Flanders, M. A. Giambrone, R. Wind, L. Biempica, and M. A. Zern. In vitro and in vivo association of transforming growth factor-beta 1 with hepatic fibrosis. Journal of Cell Biology, 108(6):2477–??, June 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/108/6/2477.

Chen:1986:RFR

[CWH^{+}86a] W. T. Chen, J. Wang, T. Hasegawa, S. S. Yamada, and K. M. Yamada. Regulation of fibronectin receptor distribution by transformation, exogenous fibronectin, and synthetic peptides. Journal of Cell Biology, 103(5):1649–??, November 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/103/5/1649.

Cremins:1986:NGF

[CWH86b] J. Cremins, J. A. Wagner, and S. Halegoua. Nerve growth factor action is mediated by cyclic AMP- and Ca^{+2}/phospholipiddependent protein kinases. Journal of Cell Biology, 103(3): 887–??, September 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/103/3/887.

Carpentier:1987:DVP

[CWOK87] J. L. Carpentier, M. F. White, L. Orci, and R. C. Kahn. Direct visualization of the phosphorylated epidermal growth factor receptor during its internalization in A-431 cells. Journal of Cell Biology, 105(6):2751–??, December 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/105/6/2751.

Cannon:1986:IRB

[CWR86] S. Cannon, P. Wang, and H. Roy. Inhibition of ribulose bisphosphate carboxylase assembly by antibody to a binding protein. Journal of Cell Biology, 103(4):1327–??, October 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/103/4/1327.

Cassimeris:1986:DMD

[CWS86] L. U. Cassimeris, P. Wadsworth, and E. D. Salmon. Dynamics of microtubule depolymerization in monocytes. Journal of Cell Biology, 102(6):2023–??, June 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/102/6/2023.

Caras:1989:ASA

[CWW89] I. W. Caras, G. N. Weddell, and S. R. Williams. Analysis of the signal for attachment of a glycophospholipid membrane anchor. Journal of Cell Biology, 108(4):1387–??, April 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/108/4/1387.

Carson:1986:ANA

[CWZ86] M. Carson, A. Weber, and S. H. Zigmond. An actinnucleating activity in polymorphonuclear leukocytes is modulated by chemotactic peptides. Journal of Cell Biology, 103(6): 2707–??, December 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/103/6/2707.

Chandler:1989:HMW

[CWZ89] D. E. Chandler, M. Whitaker, and J. Zimmerberg. High molecular weight polymers block cortical granule exocytosis in sea urchin eggs at the level of granule matrix disassembly. Journal of Cell Biology, 109(3):1269–??, September 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/109/3/1269.

Defize:1988:ACV

[DAJJ^{+}88] L. H. Defize, D. J. Arndt-Jovin, T. M. Jovin, J. Boonstra, J. Meisenhelder, T. Hunter, H. T. de Hey, and S. W. de Laat. A431 cell variants lacking the blood group A antigen display increased high affinity epidermal growth factor-receptor number, protein-tyrosine kinase activity, and receptor turnover. Journal of Cell Biology, 107(3):939–??, September 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/107/3/939.

Dedhar:1987:HOC

[DAS^{+}87] S. Dedhar, W. S. Argraves, S. Suzuki, E. Ruoslahti, and M. D. Pierschbacher. Human osteosarcoma cells resistant to detachment by an Arg–Gly–Asp–containing peptide overproduce the fibronectin receptor. Journal of Cell Biology, 105(3):1175– ??, September 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/3/1175.

Dubreuil:1985:MSU

[DB85] R. R. Dubreuil and G. B. Bouck. The membrane skeleton of a unicellular organism consists of bridged, articulating strips. Journal of Cell Biology, 101(5):1884–??, November 1985. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/101/5/1884.

Dubreuil:1987:DSC

[DBB^{+}87] R. Dubreuil, T. J. Byers, D. Branton, L. S. Goldstein, and D. P. Kiehart. Drosophilia spectrin. I. Characterization of the purified protein. Journal of Cell Biology, 105(5):2095–??, November 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/105/5/
2095.

deBernard:1986:BIE

[dBBB^{+}86] B. de Bernard, P. Bianco, E. Bonucci, M. Costantini, G. C. Lunazzi, P. Martinuzzi, C. Modricky, L. Moro, E. Panfili, and P. Pollesello. Biochemical and immunohistochemical evidence that in cartilage an alkaline phosphatase is a Ca^{2+}-binding glycoprotein. Journal of Cell Biology, 103(4):1615–??, October 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/103/4/1615.

Duperray:1987:BPP

[DBC^{+}87] A. Duperray, R. Berthier, E. Chagnon, J. J. Ryckewaert, M. Ginsberg, E. Plow, and G. Marguerie. Biosynthesis and processing of platelet GPIIb–IIIa in human megakaryocytes. Journal of Cell Biology, 104(6):1665–??, June 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/104/6/1665.

Doherty:1989:NPO

[DBD^{+}89] P. Doherty, C. H. Barton, G. Dickson, P. Seaton, L. H. Rowett, S. E. Moore, H. J. Gower, and F. S. Walsh. Neuronal process outgrowth of human sensory neurons on monolayers of cells transfected with cDNAs for five human N-CAM isoforms. Journal of Cell Biology, 109(2):789–??, August 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/109/2/789.

DeBiasio:1987:FPF

[DBE^{+}87] R. DeBiasio, G. R. Bright, L. A. Ernst, A. S. Waggoner, and D. L. Taylor. Five-parameter fluorescence imaging: wound healing of living Swiss 3T3 cells. Journal of Cell Biology, 105(4): 1613–??, October 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/4/1613.

Dorner:1987:RLG

[DBK87] A. J. Dorner, D. G. Bole, and R. J. Kaufman. The relationship of N-linked glycosylation and heavy chain-binding protein association with the secretion of glycoproteins. Journal of Cell Biology, 105(6):2665–??, December 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/105/6/2665.

DeLozanne:1987:EEC

[DBLS87] A. De Lozanne, C. H. Berlot, L. A. Leinwand, and J. A. Spudich. Expression in Escherichia coli of a functional Dictyostelium myosin tail fragment. Journal of Cell Biology, 105 (6):2999–??, December 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/6/2999.

Defize:1989:STE

[DBM^{+}89] L. H. Defize, J. Boonstra, J. Meisenhelder, W. Kruijer, L. G. Tertoolen, B. C. Tilly, T. Hunter, P. M. van Bergen en Henegouwen, W. H. Moolenaar, and S. W. de Laat. Signal transduction by epidermal growth factor occurs through the subclass of high affinity receptors. Journal of Cell Biology, 109(5):2495–??, November 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/109/5/
2495.

Dubreuil:1989:CSD

[DBS^{+}89] R. R. Dubreuil, T. J. Byers, A. L. Sillman, D. Bar-Zvi, L. S. Goldstein, and D. Branton. The complete sequence of Drosophila alpha-spectrin: conservation of structural domains between alpha-spectrins and alpha-actinin. Journal of Cell Biology, 109 (5):2197–??, November 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/5/2197.

Dechesne:1987:VCV

[DBWL87] C. A. Dechesne, P. Bouvagnet, D. Walzthöny, and J. J. Léger. Visualization of cardiac ventricular myosin heavy chain homodimers and heterodimers by monoclonal antibody epitope mapping. Journal of Cell Biology, 105(6):3031–??, December 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/105/6/3031.

Durkin:1988:AAS

[DCB^{+}88] M. E. Durkin, S. Chakravarti, B. B. Bartos, S. H. Liu, R. L. Friedman, and A. E. Chung. Amino acid sequence and domain structure of entactin. Homology with epidermal growth factor precursor and low density lipoprotein receptor. Journal of Cell Biology, 107(6):2749–??, December 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/107/6/2749.

Deatherage:1989:AFC

[DCB89] J. F. Deatherage, N. Q. Cheng, and B. Bullard. Arrangement of filaments and cross-links in the bee flight muscle Z disk by image analysis of oblique sections. Journal of Cell Biology, 108 (5):1775–??, May 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/108/5/1775.

Dejana:1988:FVR

[DCC^{+}88] E. Dejana, S. Colella, G. Conforti, M. Abbadini, M. Gaboli, and P. C. Marchisio. Fibronectin and vitronectin regulate the organization of their respective Arg–Gly–Asp adhesion receptors in cultured human endothelial cells. Journal of Cell Biology, 107(3): 1215–??, September 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/107/3/1215.

deCurtis:1986:PCT

[dCFB86] I. de Curtis, G. Fumagalli, and N. Borgese. Purification and characterization of two plasma membrane domains from ejaculated bull spermatozoa. Journal of Cell Biology, 102(5):1813–??, May 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/102/5/

Dunn:1986:RME

[DCH86] W. A. Dunn, T. P. Connolly, and A. L. Hubbard. Receptormediated endocytosis of epidermal growth factor by rat hepatocytes: receptor pathway. Journal of Cell Biology, 102(1):24–??, January 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
102/1/24.

Dejana:1987:FIA

[DCL^{+}87] E. Dejana, S. Colella, L. R. Languino, G. Balconi, G. C. Corbascio, and P. C. Marchisio. Fibrinogen induces adhesion, spreading, and microfilament organization of human endothelial cells in vitro. Journal of Cell Biology, 104(5):1403–??, May 1987. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/104/5/1403.

Daniloff:1989:ECN

[DCPR^{+}89] J. K. Daniloff, K. L. Crossin, M. Pinçon-Raymond, M. Murawsky, F. Rieger, and G. M. Edelman. Expression of cytotactin in the normal and regenerating neuromuscular system. Journal of Cell Biology, 108(2):625–??, February 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/108/2/625.

Draye:1988:QMT

[DCQB88] J. P. Draye, P. J. Courtoy, J. Quintart, and P. Baudhuin. A quantitative model of traffic between plasma membrane and secondary lysosomes: evaluation of inflow, lateral diffusion, and degradation. Journal of Cell Biology, 107(6):2109–??, December 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/107/6/
2109.

Devore:1989:MAS

[DCR89] J. J. Devore, G. W. Conrad, and R. Rappaport. A model for astral stimulation of cytokinesis in animal cells. Journal of Cell Biology, 109(5):2225–??, November 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/109/5/2225.

Danpure:1989:ETD

[DCWJ89] C. J. Danpure, P. J. Cooper, P. J. Wise, and P. R. Jennings. An enzyme trafficking defect in two patients with primary hyperoxaluria type 1: peroxisomal alanine/glyoxylate aminotransferase rerouted to mitochondria. Journal of Cell Biology, 108 (4):1345–??, April 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/108/4/1345.

Diggins:1987:DAA

[DD87] M. A. Diggins and W. F. Dove. Distribution of acetylated alphatubulin in Physarum polycephalum. Journal of Cell Biology, 104 (2):303–??, February 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/104/2/303.

Drenckhahn:1988:OAF

[DD88] D. Drenckhahn and R. Dermietzel. Organization of the actin filament cytoskeleton in the intestinal brush border: a quantitative and qualitative immunoelectron microscope study. Journal of Cell Biology, 107(3):1037–??, September 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/107/3/1037.

DeLuca:1988:HEC

[DDB^{+}88] M. De Luca, F. D’Anna, S. Bondanza, A. T. Franzi, and R. Cancedda. Human epithelial cells induce human melanocyte growth in vitro but only skin keratinocytes regulate its proper differentiation in the absence of dermis. Journal of Cell Biology, 107 (5):1919–??, November 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/107/5/1919.

Dubois-Dalcq:1986:ETM

[DDBHL86] M. Dubois-Dalcq, T. Behar, L. Hudson, and R. A. Lazzarini. Emergence of three myelin proteins in oligodendrocytes cultured without neurons. Journal of Cell Biology, 102(2):384–??, February 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/102/2/
384.

Duband:1987:AMD

[DDH^{+}87] J. L. Duband, S. Dufour, K. Hatta, M. Takeichi, G. M. Edelman, and J. P. Thiery. Adhesion molecules during somitogenesis in the avian embryo. Journal of Cell Biology, 104(5):1361–??, May 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/104/5/
1361.

Driscoll:1989:GMA

[DDR^{+}89] M. Driscoll, E. Dean, E. Reilly, E. Bergholz, and M. Chalfie. Genetic and molecular analysis of a Caenorhabditis elegans betatubulin that conveys benzimidazole sensitivity. Journal of Cell Biology, 109(6):2993–??, December 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/109/6/2993.

Dentler:1988:FTC

[Den88] W. L. Dentler. Fractionation of Tetrahymena ciliary membranes with Triton X-114 and the identification of a ciliary membrane ATPase. Journal of Cell Biology, 107(6):2679–??, December 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/107/6/
2679.

Dabauvalle:1986:DIS

[DF86a] M. C. Dabauvalle and W. W. Franke. Determination of the intracellular state of soluble macromolecules by gel filtration in vivo in the cytoplasm of amphibian oocytes. Journal of Cell Biology, 102(6):2006–??, June 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/102/6/2006.

Drenckhahn:1986:IAA

[DF86b] D. Drenckhahn and H. Franz. Identification of actin-, alphaactinin-, and vinculin-containing plaques at the lateral membrane of epithelial cells. Journal of Cell Biology, 102(5):1843–??, May 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/102/5/
1843.

Doctrow:1987:PKC

[DF87] S. R. Doctrow and J. Folkman. Protein kinase C activators suppress stimulation of capillary endothelial cell growth by angiogenic endothelial mitogens. Journal of Cell Biology, 104 (3):679–??, March 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/104/3/679.

Duden:1988:ODP

[DF88a] R. Duden and W. W. Franke. Organization of desmosomal plaque proteins in cells growing at low calcium concentrations. Journal of Cell Biology, 107(3):1049–??, September 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/107/3/1049.

Dworetzky:1988:TRC

[DF88b] S. I. Dworetzky and C. M. Feldherr. Translocation of RNAcoated gold particles through the nuclear pores of oocytes. Journal of Cell Biology, 106(3):575–??, March 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/106/3/575.

Denoulet:1989:DAT

[DFdN^{+}89] P. Denoulet, G. Filliatreau, B. de Néchaud, F. Gros, and L. Di Giamberardino. Differential axonal transport of isotubulins in the motor axons of the rat sciatic nerve. Journal of Cell Biology, 108(3):965–??, March 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/108/3/965.

Doctor:1985:PCD

[DFF85] J. Doctor, D. Fristrom, and J. W. Fristrom. The pupal cuticle of Drosophila: biphasic synthesis of pupal cuticle proteins in vivo and in vitro in response to 20-hydroxyecdysone. Journal of Cell Biology, 101(1):189–??, July 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/101/1/189.

Drubin:1985:NGF

[DFSK85] D. G. Drubin, S. C. Feinstein, E. M. Shooter, and M. W. Kirschner. Nerve growth factor-induced neurite outgrowth in PC12 cells involves the coordinate induction of microtubule assembly and assembly-promoting factors. Journal of Cell Biology, 101(5):1799–??, November 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/101/5/1799.

DeLemos-Chiarandini:1987:DMT

[DFSK87] C. De Lemos-Chiarandini, A. B. Frey, D. D. Sabatini, and G. Kreibich. Determination of the membrane topology of the phenobarbital-inducible rat liver cytochrome P-450 isoenzyme PB-4 using site-specific antibodies. Journal of Cell Biology, 104 (2):209–??, February 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/104/2/209.

Dresser:1988:MCB

[DG88] M. E. Dresser and C. N. Giroux. Meiotic chromosome behavior in spread preparations of yeast. Journal of Cell Biology, 106 (3):567–??, March 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/106/3/567.

Dahl:1989:IPM

[DG89] S. C. Dahl and L. B. Grabel. Integrin phosphorylation is modulated during the differentiation of F-9 teratocarcinoma stem cells. Journal of Cell Biology, 108(1):183–??, January 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/108/1/183.

deGroot:1987:PHE

[dGRS87] P. G. de Groot, J. H. Reinders, and J. J. Sixma. Perturbation of human endothelial cells by thrombin or PMA changes the reactivity of their extracellular matrix towards platelets. Journal of Cell Biology, 104(3):697–??, March 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/104/3/697.

Davis:1987:CAA

[DH87] G. L. Davis and E. Hunter. A charged amino acid substitution within the transmembrane anchor of the Rous sarcoma virus envelope glycoprotein affects surface expression but not intracellular transport. Journal of Cell Biology, 105(3):1191–??, September 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/105/3/
1191.

Daleke:1989:EMR

[DH89a] D. L. Daleke and W. H. Huestis. Erythrocyte morphology reflects the transbilayer distribution of incorporated phospholipids. Journal of Cell Biology, 108(4):1375–??, April 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/108/4/1375.

Duncan:1989:PSP

[DH89b] R. F. Duncan and J. W. Hershey. Protein synthesis and protein phosphorylation during heat stress, recovery, and adaptation. Journal of Cell Biology, 109(4):1467–??, October 1989. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/109/4/1467.

Dingus:1986:IMA

[DHB86] J. Dingus, S. Hwo, and J. Bryan. Identification by monoclonal antibodies and characterization of human platelet caldesmon. Journal of Cell Biology, 102(5):1748–??, May 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/102/5/1748.

DeNofrio:1989:FSA

[DHH89] D. DeNofrio, T. C. Hoock, and I. M. Herman. Functional sorting of actin isoforms in microvascular pericytes. Journal of Cell Biology, 109(1):191–??, July 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/109/1/191.

Dale:1985:EEK

[DHK^{+}85] B. A. Dale, K. A. Holbrook, J. R. Kimball, M. Hoff, and T. T. Sun. Expression of epidermal keratins and filaggrin during human fetal skin development. Journal of Cell Biology, 101(4): 1257–??, October 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/101/4/1257.

Dawson:1985:MAI

[DHL85] P. J. Dawson, J. S. Hulme, and C. W. Lloyd. Monoclonal antibody to intermediate filament antigen cross-reacts with higher plant cells. Journal of Cell Biology, 100(5):1793–??, May 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/100/5/1793.

deHostos:1988:PBD

[dHTG88] E. L. de Hostos, R. K. Togasaki, and A. Grossman. Purification and biosynthesis of a derepressible periplasmic arylsulfatase from Chlamydomonas reinhardtii. Journal of Cell Biology, 106 (1):29–??, January 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/106/1/29.

DePasquale:1987:EAC

[DI87] J. A. DePasquale and C. S. Izzard. Evidence for an actincontaining cytoplasmic precursor of the focal contact and the timing of incorporation of vinculin at the focal contact. Journal of Cell Biology, 105(6):2803–??, December 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/105/6/2803.

Duszenko:1988:ITV

[DIF^{+}88] M. Duszenko, I. E. Ivanov, M. A. Ferguson, H. Plesken, and G. A. Cross. Intracellular transport of a variant surface glycoprotein in Trypanosoma brucei. Journal of Cell Biology, 106(1):77–??, January 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/106/1/
77.

Danilov:1989:PEM

[DJ89] Y. N. Danilov and R. L. Juliano. Phorbol ester modulation of integrin-mediated cell adhesion: a postreceptor event. Journal of Cell Biology, 108(5):1925–??, May 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/108/5/1925.

Dorne:1985:LPO

[DJBD85] A. J. Dorne, J. Joyard, M. A. Block, and R. Douce. Localization of phosphatidylcholine in outer envelope membrane of spinach chloroplasts. Journal of Cell Biology, 100(5):1690–??, May 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/100/5/
1690.

deJong:1986:HSR

[dJHMB86] W. W. de Jong, W. A. Hoekman, J. W. Mulders, and H. Bloemendal. Heat shock response of the rat lens. Journal of Cell Biology, 102(1):104–??, January 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/102/1/104.

Disney:1989:HMG

[DJM^{+}89] J. E. Disney, K. R. Johnson, N. S. Magnuson, S. R. Sylvester, and R. Reeves. High-mobility group protein HMG-I localizes to G/ Q- and C-bands of human and mouse chromosomes. Journal of Cell Biology, 109(5):1975–??, November 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/109/5/1975.

Dennerll:1988:TCC

[DJS^{+}88] T. J. Dennerll, H. C. Joshi, V. L. Steel, R. E. Buxbaum, and S. R. Heidemann. Tension and compression in the cytoskeleton of PC-12 neurites. II: Quantitative measurements. Journal of Cell Biology, 107(2):665–??, August 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/107/2/665.

Drubin:1986:TPF

[DK86] D. G. Drubin and M. W. Kirschner. Tau protein function in living cells. Journal of Cell Biology, 103(6):2739–??, December 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/103/6/
2739.

Duncan:1988:IMT

[DK88] J. R. Duncan and S. Kornfeld. Intracellular movement of two mannose 6-phosphate receptors: return to the Golgi apparatus. Journal of Cell Biology, 106(3):617–??, March 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/106/3/617.

Damsky:1985:DCS

[DKB^{+}85] C. H. Damsky, K. A. Knudsen, D. Bradley, C. A. Buck, and A. F. Horwitz. Distribution of the cell substratum attachment (CSAT) antigen on myogenic and fibroblastic cells in culture. Journal of Cell Biology, 100(5):1528–??, May 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/100/5/1528.

Davidson:1985:AWR

[DKH^{+}85] J. M. Davidson, M. Klagsbrun, K. E. Hill, A. Buckley, R. Sullivan, P. S. Brewer, and S. C. Woodward. Accelerated wound repair, cell proliferation, and collagen accumulation are produced by a cartilage-derived growth factor. Journal of Cell Biology, 100(4):1219–??, April 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/100/4/1219.

Doms:1987:RAT

[DKHB87] R. W. Doms, D. S. Keller, A. Helenius, and W. E. Balch. Role for adenosine triphosphate in regulating the assembly and transport of vesicular stomatitis virus G protein trimers. Journal of Cell Biology, 105(5):1957–??, November 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/105/5/1957.

Drubin:1988:RMP

[DKKK88] D. Drubin, S. Kobayashi, D. Kellogg, and M. Kirschner. Regulation of microtubule protein levels during cellular morphogenesis in nerve growth factor-treated PC12 cells. Journal of Cell Biology, 106(5):1583–??, May 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/106/5/1583.

Dennerll:1989:CAE

[DLBH89] T. J. Dennerll, P. Lamoureux, R. E. Buxbaum, and S. R. Heidemann. The cytomechanics of axonal elongation and retraction. Journal of Cell Biology, 109(6):3073–??, December 1989. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/109/6/3073.

Dreesen:1988:DRE

[DLC88] T. D. Dreesen, M. Lezzi, and S. T. Case. Developmentally regulated expression of a Balbiani ring 1 gene for a 180-kD secretory polypeptide in Chironomus tentans salivary glands before larval/pupal ecdysis. Journal of Cell Biology, 106(1):21–??, January 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/106/1/
21.

Dworetzky:1988:EVN

[DLF88] S. I. Dworetzky, R. E. Lanford, and C. M. Feldherr. The effects of variations in the number and sequence of targeting signals on nuclear uptake. Journal of Cell Biology, 107(4):1279–??, October 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/107/4/
1279.

Dubinsky:1989:FAR

[DLFE89] J. M. Dubinsky, D. J. Loftus, G. D. Fischbach, and E. L. Elson. Formation of acetylcholine receptor clusters in chick myotubes: migration or new insertion? Journal of Cell Biology, 109 (4):1733–??, October 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/4/1733.

Daniloff:1986:AEN

[DLG^{+}86a] J. K. Daniloff, G. Levi, M. Grumet, F. Rieger, and G. M. Edelman. Altered expression of neuronal cell adhesion molecules induced by nerve injury and repair. Journal of Cell Biology, 103 (3):929–??, September 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/103/3/929.

deLamirande:1986:EPI

[dLG86b] E. de Lamirande and C. Gagnon. Effects of protease inhibitors and substrates on motility of mammalian spermatozoa. Journal of Cell Biology, 102(4):1378–??, April 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/102/4/1378.

Dejana:1989:WFP

[DLG^{+}89] E. Dejana, M. G. Lampugnani, M. Giorgi, M. Gaboli, A. B. Federici, Z. M. Ruggeri, and P. C. Marchisio. Von Willebrand factor promotes endothelial cell adhesion via an Arg–Gly–Aspdependent mechanism. Journal of Cell Biology, 109(1):367–??, July 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/109/1/
367.

David:1989:MAC

[DLH^{+}89] G. David, V. Lories, A. Heremans, B. Van der Schueren, J. J. Cassiman, and H. Van den Berghe. Membrane-associated chondroitin sulfate proteoglycans of human lung fibroblasts. Journal of Cell Biology, 108(3):1165–??, March 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/108/3/1165.

DiMario:1987:FSY

[DM87] P. J. DiMario and A. P. Mahowald. Female sterile (1) yolkless: a recessive female sterile mutation in Drosophila melanogaster with depressed numbers of coated pits and coated vesicles within the developing oocytes. Journal of Cell Biology, 105(1):199–??, July 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/105/1/
199.

Drubin:1988:YAB

[DMB88] D. G. Drubin, K. G. Miller, and D. Botstein. Yeast actin-binding proteins: evidence for a role in morphogenesis. Journal of Cell Biology, 107(6):2551–??, December 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/107/6/2551.

Drummond:1985:SMM

[DMC85] D. R. Drummond, M. A. McCrae, and A. Colman. Stability and movement of mRNAs and their encoded proteins in Xenopus oocytes. Journal of Cell Biology, 100(4):1148–??, April 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/100/4/1148.

DeCamilli:1986:HDC

[DMD^{+}86] P. De Camilli, M. Moretti, S. D. Donini, U. Walter, and S. M. Lohmann. Heterogeneous distribution of the cAMP receptor protein RII in the nervous system: evidence for its intracellular accumulation on microtubules, microtubule-organizing centers, and in the area of the Golgi complex. Journal of Cell Biology, 103(1):189–??, July 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/103/1/189.

Damiani:1987:CSD

[DMF^{+}87] E. Damiani, A. Margreth, A. Furlan, A. S. Dahms, J. Arnn, and R. A. Sabbadini. Common structural domains in the sarcoplasmic reticulum Ca–ATPase and the transverse tubule Mg– ATPase. Journal of Cell Biology, 104(3):461–??, March 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/104/3/461.

Dotto:1989:TMM

[DMG^{+}89] G. P. Dotto, G. Moellmann, S. Ghosh, M. Edwards, and R. Halaban. Transformation of murine melanocytes by basic fibroblast growth factor cDNA and oncogenes and selective suppression of the transformed phenotype in a reconstituted cutaneous environment. Journal of Cell Biology, 109(6):3115–??, December 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/109/6/
3115.

DiVirgilio:1988:FRM

[DMGS88] F. Di Virgilio, B. C. Meyer, S. Greenberg, and S. C. Silverstein. Fc receptor-mediated phagocytosis occurs in macrophages at exceedingly low cytosolic Ca^{2+} levels. Journal of Cell Biology, 106 (3):657–??, March 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/106/3/657.

Donaldson:1985:LFD

[DMH^{+}85] D. J. Donaldson, J. T. Mahan, D. L. Hasty, J. B. McCarthy, and L. T. Furcht. Location of a fibronectin domain involved in newt epidermal cell migration. Journal of Cell Biology, 101(1):73–??, July 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/101/1/
73.

Dunn:1989:IFR

[DMM89] K. W. Dunn, T. E. McGraw, and F. R. Maxfield. Iterative fractionation of recycling receptors from lysosomally destined ligands in an early sorting endosome. Journal of Cell Biology, 109(6): 3303–??, December 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/6/3303.

Dunia:1987:EMO

[DMRB87] I. Dunia, S. Manenti, A. Rousselet, and E. L. Benedetti. Electron microscopic observations of reconstituted proteoliposomes with the purified major intrinsic membrane protein of eye lens fibers. Journal of Cell Biology, 105(4):1679–??, October 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/105/4/1679.

Doherty:1988:CEN

[DMW88] P. Doherty, D. A. Mann, and F. S. Walsh. Comparison of the effects of NGF, activators of protein kinase C, and a calcium ionophore on the expression of Thy-1 and N-CAM in PC12 cell cultures. Journal of Cell Biology, 107(1):333–??, July 1988. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/107/1/333.

Dunphy:1988:MIF

[DN88] W. G. Dunphy and J. W. Newport. Mitosis-inducing factors are present in a latent form during interphase in the Xenopus embryo. Journal of Cell Biology, 106(6):2047–??, June 1988. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/106/6/2047.

Dean:1988:RFB

[DNB88] D. C. Dean, R. F. Newby, and S. Bourgeois. Regulation of fibronectin biosynthesis by dexamethasone, transforming growth factor beta, and cAMP in human cell lines. Journal of Cell Biology, 106(6):2159–??, June 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/106/6/2159.

Duband:1988:FRE

[DNI^{+}88] J. L. Duband, G. H. Nuckolls, A. Ishihara, T. Hasegawa, K. M. Yamada, J. P. Thiery, and K. Jacobson. Fibronectin receptor exhibits high lateral mobility in embryonic locomoting cells but is immobile in focal contacts and fibrillar streaks in stationary cells. Journal of Cell Biology, 107(4):1385–??, October 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/107/4/1385.

Deitcher:1986:FEP

[DNM86] D. L. Deitcher, M. R. Neutra, and K. E. Mostov. Functional expression of the polymeric immunoglobulin receptor from cloned cDNA in fibroblasts. Journal of Cell Biology, 102(3):911–??, March 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
102/3/911.

Diaz-Nido:1988:CKI

[DNSMA88] J. Díaz-Nido, L. Serrano, E. Méndez, and J. Avila. A casein kinase II-related activity is involved in phosphorylation of microtubule-associated protein MAP-1B during neuroblastoma cell differentiation. Journal of Cell Biology, 106(6):2057–??, June 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/106/6/

Dreyfuss:1988:RPC

[DPM88] G. Dreyfuss, L. Philipson, and I. W. Mattaj. Ribonucleoprotein particles in cellular processes. Journal of Cell Biology, 106 (5):1419–??, May 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/106/5/1419.

Duband:1986:CAM

[DRC^{+}86] J. L. Duband, S. Rocher, W. T. Chen, K. M. Yamada, and J. P. Thiery. Cell adhesion and migration in the early vertebrate embryo: location and possible role of the putative fibronectin receptor complex. Journal of Cell Biology, 102(1):160–??, January 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/102/1/
160.

Dingwall:1988:NNL

[DRD^{+}88] C. Dingwall, J. Robbins, S. M. Dilworth, B. Roberts, and W. D. Richardson. The nucleoplasmin nuclear location sequence is larger and more complex than that of SV-40 large T antigen. Journal of Cell Biology, 107(3):841–??, September 1988. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/107/3/841.

Dudouet:1987:CVS

[DRH^{+}87] B. Dudouet, S. Robine, C. Huet, C. Sahuquillo-Merino, L. Blair, E. Coudrier, and D. Louvard. Changes in villin synthesis and subcellular distribution during intestinal differentiation of HT29-18 clones. Journal of Cell Biology, 105(1):359–??, July 1987. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/105/1/359.

Doms:1988:DEM

[DRM^{+}88] R. W. Doms, A. Ruusala, C. Machamer, J. Helenius, A. Helenius, and J. K. Rose. Differential effects of mutations in three domains on folding, quaternary structure, and intracellular transport of vesicular stomatitis virus G protein. Journal of Cell Biology, 107(1):89–??, July 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/107/1/89.

Drake:1989:FTF

[DRME89] T. A. Drake, W. Ruf, J. H. Morrissey, and T. S. Edgington. Functional tissue factor is entirely cell surface expressed on lipopolysaccharide-stimulated human blood monocytes and a constitutively tissue factor-producing neoplastic cell line. Journal of Cell Biology, 109(1):389–??, July 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/109/1/389.

Dedhar:1987:CSR

[DRP87] S. Dedhar, E. Ruoslahti, and M. D. Pierschbacher. A cell surface receptor complex for collagen type I recognizes the Arg–Gly– Asp sequence. Journal of Cell Biology, 104(3):585–??, March 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/104/3/
585.

Dubreuil:1988:PTM

[DRRB88] R. R. Dubreuil, T. K. Rosiere, M. C. Rosner, and G. B. Bouck. Properties and topography of the major integral plasma membrane protein of a unicellular organism. Journal of Cell Biology, 107(1):191–??, July 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/107/1/191.

Doyle:1985:MCD

[DRSG85] C. Doyle, M. G. Roth, J. Sambrook, and M. J. Gething. Mutations in the cytoplasmic domain of the influenza virus hemagglutinin affect different stages of intracellular transport. Journal of Cell Biology, 100(3):704–??, March 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/100/3/704.

Doms:1989:BRR

[DRY89] R. W. Doms, G. Russ, and J. W. Yewdell. Brefeldin A redistributes resident and itinerant Golgi proteins to the endoplasmic reticulum. Journal of Cell Biology, 109(1):61–??, July 1989. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/109/1/61.

Deshaies:1987:YMD

[DS87] R. J. Deshaies and R. Schekman. A yeast mutant defective at an early stage in import of secretory protein precursors into the endoplasmic reticulum. Journal of Cell Biology, 105(2):633–??, August 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/105/2/
633.

Dustin:1988:LFA

[DS88] M. L. Dustin and T. A. Springer. Lymphocyte functionassociated antigen-1 (LFA-1) interaction with intercellular adhesion molecule-1 (ICAM-1) is one of at least three mechanisms for lymphocyte adhesion to cultured endothelial cells. Journal of Cell Biology, 107(1):321–??, July 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/107/1/321.

Deshaies:1989:SEP

[DS89] R. J. Deshaies and R. Schekman. SEC62 encodes a putative membrane protein required for protein translocation into the yeast endoplasmic reticulum. Journal of Cell Biology, 109(6): 2653–??, December 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/6/2653.

Darr:1986:MAL

[DSA86] S. C. Darr, S. C. Somerville, and C. J. Arntzen. Monoclonal antibodies to the light-harvesting chlorophyll a/b protein complex of photosystem II. Journal of Cell Biology, 103(3):733–??, September 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/103/3/
733.

deSa:1988:APS

[dSdSH^{+}88] M. F. Grossi de Sa, C. Martins de Sa, F. Harper, M. Olink-Coux, M. Huesca, and K. Scherrer. The association of prosomes with some of the intermediate filament networks of the animal cell. Journal of Cell Biology, 107(4):1517–??, October 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/107/4/1517.

Doyle:1986:APD

[DSG86] C. Doyle, J. Sambrook, and M. J. Gething. Analysis of progressive deletions of the transmembrane and cytoplasmic domains of influenza hemagglutinin. Journal of Cell Biology, 103 (4):1193–??, October 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/103/4/1193.

Dixon:1987:TSB

[DSGS87] S. J. Dixon, D. Stewart, S. Grinstein, and S. Spiegel. Transmembrane signaling by the B subunit of cholera toxin: increased cytoplasmic free calcium in rat lymphocytes. Journal of Cell Biology, 105(3):1153–??, September 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/105/3/1153.

Doxsey:1985:EMI

[DSHW85] S. J. Doxsey, J. Sambrook, A. Helenius, and J. White. An efficient method for introducing macromolecules into living cells. Journal of Cell Biology, 101(1):19–??, July 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/101/1/19.

DiStefano:1985:SDN

[DSTJ85] P. S. DiStefano, J. B. Schweitzer, M. Taniuchi, and E. M. Johnson. Selective destruction of nerve growth factor receptor-bearing cells in vitro using a hybrid toxin composed of ricin A chain and a monoclonal antibody against the nerve growth factor receptor. Journal of Cell Biology, 101(3):1107–??, September 1985. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/101/3/1107.

DeRosier:1989:SCP

[DT89] D. J. DeRosier and L. G. Tilney. The structure of the cuticular plate, an in vivo actin gel. Journal of Cell Biology, 109(6): 2853–??, December 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/6/2853.

Davies:1985:MCB

[DTW^{+}85] P. F. Davies, G. A. Truskey, H. B. Warren, S. E. O’Connor, and B. H. Eisenhaure. Metabolic cooperation between vascular endothelial cells and smooth muscle cells in co-culture: changes in low density lipoprotein metabolism. Journal of Cell Biology, 101 (3):871–??, September 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/101/3/871.

Dorel:1989:TPP

[DVHC89] C. Dorel, T. A. Voelker, E. M. Herman, and M. J. Chrispeels. Transport of proteins to the plant vacuole is not by bulk flow through the secretory system, and requires positive sorting information. Journal of Cell Biology, 108(2):327–??, February 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/108/2/
327.

deVitry:1989:PEL

[dVOD^{+}89] C. de Vitry, J. Olive, D. Drapier, M. Recouvreur, and F. A. Wollman. Posttranslational events leading to the assembly of photosystem II protein complex: a study using photosynthesis mutants from Chlamydomonas reinhardtii. Journal of Cell Biology, 109 (3):991–??, September 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/3/991.

Distel:1988:AOE

[DVVT88] B. Distel, I. Van der Leÿ, M. Veenhuis, and H. F. Tabak. Alcohol oxidase expressed under nonmethylotrophic conditions is imported, assembled, and enzymatically active in peroxisomes of Hansenula polymorpha. Journal of Cell Biology, 107(5): 1669–??, November 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/107/5/1669.

Drenckhahn:1986:SFS

[DW86] D. Drenckhahn and J. Wagner. Stress fibers in the splenic sinus endothelium in situ: molecular structure, relationship to the extracellular matrix, and contractility. Journal of Cell Biology, 102(5):1738–??, May 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/102/5/1738.

Davidson:1989:EDP

[DW89] H. W. Davidson and C. Watts. Epitope-directed processing of specific antigen by B lymphocytes. Journal of Cell Biology, 109(1):85–??, July 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/1/85.

DeBiasio:1988:DDF

[DWFT88] R. L. DeBiasio, L. L. Wang, G. W. Fisher, and D. L. Taylor. The dynamic distribution of fluorescent analogues of actin and myosin in protrusions at the leading edge of migrating Swiss 3T3 fibroblasts. Journal of Cell Biology, 107(6):2631–??, December 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/107/6/
2631.

Darlington:1986:MCM

[DWL86] G. J. Darlington, D. R. Wilson, and L. B. Lachman. Monocyteconditioned medium, interleukin-1, and tumor necrosis factor stimulate the acute phase response in human hepatoma cells in vitro. Journal of Cell Biology, 103(3):787–??, September 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/103/3/787.

Detmers:1987:ACR

[DWO^{+}87] P. A. Detmers, S. D. Wright, E. Olsen, B. Kimball, and Z. A. Cohn. Aggregation of complement receptors on human neutrophils in the absence of ligand. Journal of Cell Biology, 105(3): 1137–??, September 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/3/1137.

DelBuono:1988:RBO

[DWS88] B. J. Del Buono, P. L. Williamson, and R. A. Schlegel. Relation between the organization of spectrin and of membrane lipids in lymphocytes. Journal of Cell Biology, 106(3):697–??, March 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/106/3/
697.

Davies:1989:OFA

[DWT^{+}89] J. Davies, J. Warwick, N. Totty, R. Philp, M. Helfrich, and M. Horton. The osteoclast functional antigen, implicated in the regulation of bone resorption, is biochemically related to the vitronectin receptor. Journal of Cell Biology, 109(4):1817–??, October 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/109/4/

Dermietzel:1987:MLK

[DYT^{+}87] R. Dermietzel, S. B. Yancey, O. Traub, K. Willecke, and J. P. Revel. Major loss of the 28-kD protein of gap junction in proliferating hepatocytes. Journal of Cell Biology, 105(4):1925–??, October 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/105/4/
1925.

Daukas:1985:IRM

[DZ85] G. Daukas and S. H. Zigmond. Inhibition of receptor-mediated but not fluid-phase endocytosis in polymorphonuclear leukocytes. Journal of Cell Biology, 101(5):1673–??, November 1985. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/101/5/1673.

Earnshaw:1987:ARN

[Ear87] W. C. Earnshaw. Anionic regions in nuclear proteins. Journal of Cell Biology, 105(4):1479–??, October 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/105/4/1479.

Edwards:1987:CSL

[EB87] H. C. Edwards and A. G. Booth. Calcium-sensitive, lipid-binding cytoskeletal proteins of the human placental microvillar region. Journal of Cell Biology, 105(1):303–??, July 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/105/1/303.

Eldridge:1987:DAR

[EBBW87] C. F. Eldridge, M. B. Bunge, R. P. Bunge, and P. M. Wood. Differentiation of axon-related Schwann cells in vitro. I. Ascorbic acid regulates basal lamina assembly and myelin formation. Journal of Cell Biology, 105(2):1023–??, August 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/105/2/1023.

Epstein:1988:PTF

[EBCO88] H. F. Epstein, G. C. Berliner, D. L. Casey, and I. Ortiz. Purified thick filaments from the nematode Caenorhabditis elegans: evidence for multiple proteins associated with core structures. Journal of Cell Biology, 106(6):1985–??, June 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/106/6/1985.

Eppler:1986:SSF

[EBGS86] C. M. Eppler, H. Bayley, S. M. Greenberg, and J. H. Schwartz. Structural studies on a family of cAMP-binding proteins in the nervous system of Aplysia. Journal of Cell Biology, 102(1): 320–??, January 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/102/1/320.

Eckmiller:1987:COS

[Eck87] M. S. Eckmiller. Cone outer segment morphogenesis: taper change and distal invaginations. Journal of Cell Biology, 105 (5):2267–??, November 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/5/2267.

Entwistle:1986:MFR

[ECZ86] A. Entwistle, D. H. Curtis, and R. J. Zalin. Myoblast fusion is regulated by a prostanoid of the one series independently of a rise in cyclic AMP. Journal of Cell Biology, 103(3):857–??, September 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/103/3/
857.

Engvall:1986:MDH

[EDD^{+}86] E. Engvall, G. E. Davis, K. Dickerson, E. Ruoslahti, S. Varon, and M. Manthorpe. Mapping of domains in human laminin using monoclonal antibodies: localization of the neurite-promoting site. Journal of Cell Biology, 103(6):2457–??, December 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/103/6/2457.

Ebert:1989:DDM

[EFR^{+}89] D. L. Ebert, H. H. Freeze, J. Richardson, R. L. Dimond, and J. A. Cardelli. A Dictyostelium discoideum mutant that missorts and oversecretes lysosomal enzyme precursors is defective in endocytosis. Journal of Cell Biology, 109(4):1445–??, October 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/109/4/
1445.

Einstein:1989:SFA

[EG89] R. Einstein and C. A. Gabel. Serum factors alter the extent of dephosphorylation of ligands endocytosed via the mannose 6phosphate/insulin-like growth factor II receptor. Journal of Cell Biology, 109(3):1037–??, September 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/109/3/1037.

Earnshaw:1985:LTI

[EH85] W. C. Earnshaw and M. M. Heck. Localization of topoisomerase II in mitotic chromosomes. Journal of Cell Biology, 100 (5):1716–??, May 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/100/5/1716.

Earnshaw:1985:TIS

[EHC^{+}85] W. C. Earnshaw, B. Halligan, C. A. Cooke, M. M. Heck, and L. F. Liu. Topoisomerase II is a structural component of mitotic chromosome scaffolds. Journal of Cell Biology, 100(5):1706–??, May 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/100/5/
1706.

Engvall:1986:MAS

[EHK86] E. Engvall, H. Hessle, and G. Klier. Molecular assembly, secretion, and matrix deposition of type VI collagen. Journal of Cell Biology, 102(3):703–??, March 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/102/3/703.

Enders:1986:SCB

[EHM86] G. C. Enders, J. H. Henson, and C. F. Millette. Sertoli cell binding to isolated testicular basement membrane. Journal of Cell Biology, 103(3):1109–??, September 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/103/3/1109.

Eilers:1989:NMA

[EKH89] U. Eilers, J. Klumperman, and H. P. Hauri. Nocodazole, a microtubule-active drug, interferes with apical protein delivery in cultured intestinal epithelial cells (Caco-2). Journal of Cell Biology, 108(1):13–??, January 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/108/1/13.

Edmondson:1988:ANN

[ELKH88] J. C. Edmondson, R. K. Liem, J. E. Kuster, and M. E. Hatten. Astrotactin: a novel neuronal cell surface antigen that mediates neuron-astroglial interactions in cerebellar microcultures. Journal of Cell Biology, 106(2):505–??, February 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/106/2/505.

Epstein:1986:TSD

[EMG86] L. M. Epstein, K. A. Mahon, and J. G. Gall. Transcription of a satellite DNA in the newt. Journal of Cell Biology, 103 (4):1137–??, October 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/103/4/1137.

Estes:1987:MBI

[EMH^{+}87] K. Estes, L. L. Monfalcone, S. R. Hammes, D. Holowka, and B. Baird. Membrane-bound IgE receptor complexes fused with rat basophilic leukemia cells mediate degranulation. Journal of Cell Biology, 105(2):747–??, August 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/105/2/747.

Evans:1985:ICS

[EMK85] L. Evans, T. Mitchison, and M. Kirschner. Influence of the centrosome on the structure of nucleated microtubules. Journal of Cell Biology, 100(4):1185–??, April 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/100/4/1185.

Epstein:1985:MPO

[EMOB85] H. F. Epstein, D. M. Miller, I. Ortiz, and G. C. Berliner. Myosin and paramyosin are organized about a newly identified core structure. Journal of Cell Biology, 100(3):904–??, March 1985. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/100/3/904.

Einspahr:1988:CCD

[EMT88] K. J. Einspahr, M. Maeda, and G. A. Thompson. Concurrent changes in Dunaliella salina ultrastructure and membrane phospholipid metabolism after hyperosmotic shock. Journal of Cell Biology, 107(2):529–??, August 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/107/2/529.

Ertl:1989:EMV

[EMW^{+}89] H. Ertl, R. Mengele, S. Wenzl, J. Engel, and M. Sumper. The extracellular matrix of Volvox carteri: molecular structure of the cellular compartment. Journal of Cell Biology, 109(6): 3493–??, December 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/6/3493.

Eusebi:1985:AAP

[EMZ85] F. Eusebi, M. Molinaro, and B. M. Zani. Agents that activate protein kinase C reduce acetylcholine sensitivity in cultured myotubes. Journal of Cell Biology, 100(4):1339–??, April 1985. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/100/4/1339.

Epstein:1986:AMI

[EOM86] H. F. Epstein, I. Ortiz, and L. A. Mackinnon. The alteration of myosin isoform compartmentation in specific mutants of Caenorhabditis elegans. Journal of Cell Biology, 103(3): 985–??, September 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/103/3/985.

Ecob-Prince:1986:NAM

[EPJBBW86] M. S. Ecob-Prince, M. Jenkison, G. S. Butler-Browne, and R. G. Whalen. Neonatal and adult myosin heavy chain isoforms in a nerve-muscle culture system. Journal of Cell Biology, 103(3): 995–??, September 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/103/3/995.

Eisen:1985:SSC

[ER85] A. Eisen and G. T. Reynolds. Source and sinks for the calcium released during fertilization of single sea urchin eggs. Journal of Cell Biology, 100(5):1522–??, May 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/100/5/1522.

Englander:1987:ARC

[ER87] L. L. Englander and L. L. Rubin. Acetylcholine receptor clustering and nuclear movement in muscle fibers in culture. Journal of Cell Biology, 104(1):87–??, January 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/104/1/87.

Euteneuer:1985:EIA

[ES85] U. Euteneuer and M. Schliwa. Evidence for an involvement of actin in the positioning and motility of centrosomes. Journal of Cell Biology, 101(1):96–??, July 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/101/1/96.

Eskridge:1986:NTP

[ES86] E. M. Eskridge and D. Shields. The NH2 terminus of preproinsulin directs the translocation and glycosylation of a bacterial cytoplasmic protein by mammalian microsomal membranes. Journal of Cell Biology, 103(6):2263–??, December 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/103/6/2263.

Eckstein:1989:LIS

[ES89] D. J. Eckstein and B. D. Shur. Laminin induces the stable expression of surface galactosyltransferase on lamellipodia of migrating cells. Journal of Cell Biology, 108(6):2507–??, June 1989. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/108/6/2507.

Eichner:1986:RKS

[ESA86] R. Eichner, T. T. Sun, and U. Aebi. The role of keratin subfamilies and keratin pairs in the formation of human epidermal intermediate filaments. Journal of Cell Biology, 102(5):1767–??, May 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/102/5/
1767.

Earnshaw:1987:MCC

[ESM^{+}87] W. C. Earnshaw, K. F. Sullivan, P. S. Machlin, C. A. Cooke, D. A. Kaiser, T. D. Pollard, N. F. Rothfield, and D. W. Cleveland. Molecular cloning of cDNA for CENP–B, the major human centromere autoantigen. Journal of Cell Biology, 104(4): 817–??, January 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/104/4/817.

Eichele:1987:CCG

[ET87a] G. Eichele and C. Thaller. Characterization of concentration gradients of a morphogenetically active retinoid in the chick limb bud. Journal of Cell Biology, 105(4):1917–??, October 1987. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/105/4/1917.

Erickson:1987:HPE

[ET87b] H. P. Erickson and H. C. Taylor. Hexabrachion proteins in embryonic chicken tissues and human tumors. Journal of Cell Biology, 105(3):1387–??, September 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/105/3/1387.

Eichele:1985:SMR

[ETA85] G. Eichele, C. Tickle, and B. M. Alberts. Studies on the mechanism of retinoid-induced pattern duplications in the early chick limb bud: temporal and spatial aspects. Journal of Cell Biology, 101(5):1913–??, November 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/101/5/1913.

Edgar:1988:SRS

[ETT88] D. Edgar, R. Timpl, and H. Thoenen. Structural requirements for the stimulation of neurite outgrowth by two variants of laminin and their inhibition by antibodies. Journal of Cell Biology, 106 (4):1299–??, April 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/106/4/1299.

Evans:1989:PVM

[Eva89] R. M. Evans. Phosphorylation of vimentin in mitotically selected cells. In vitro cyclic AMP-independent kinase and calciumstimulated phosphatase activities. Journal of Cell Biology, 108 (1):67–??, January 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/108/1/67.

Emr:1986:ATY

[EVG^{+}86] S. D. Emr, A. Vassarotti, J. Garrett, B. L. Geller, M. Takeda, and M. G. Douglas. The amino terminus of the yeast F1-ATPase beta-subunit precursor functions as a mitochondrial import signal. Journal of Cell Biology, 102(2):523–??, February 1986. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/102/2/523.

Ellis:1985:NGC

[EWAP85] L. Ellis, I. Wallis, E. Abreu, and K. H. Pfenninger. Nerve growth cones isolated from fetal rat brain. IV. Preparation of a membrane subfraction and identification of a membrane glycoprotein expressed on sprouting neurons. Journal of Cell Biology, 101(5): 1977–??, November 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/101/5/1977.

Ewenstein:1987:CWF

[EWHP87] B. M. Ewenstein, M. J. Warhol, R. I. Handin, and J. S. Pober. Composition of the von Willebrand factor storage organelle (Weibel–Palade body) isolated from cultured human umbilical vein endothelial cells. Journal of Cell Biology, 104(5):1423–??, May 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/104/5/
1423.

Entwistle:1988:CCM

[EZBW88] A. Entwistle, R. J. Zalin, S. Bevan, and A. E. Warner. The control of chick myoblast fusion by ion channels operated by prostaglandins and acetylcholine. Journal of Cell Biology, 106 (5):1693–??, May 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/106/5/1693.

Entwistle:1988:RAR

[EZWB88] A. Entwistle, R. J. Zalin, A. E. Warner, and S. Bevan. A role for acetylcholine receptors in the fusion of chick myoblasts. Journal of Cell Biology, 106(5):1703–??, May 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/106/5/1703.

Foe:1985:RCC

[FA85] V. E. Foe and B. M. Alberts. Reversible chromosome condensation induced in Drosophila embryos by anoxia: visualization of interphase nuclear organization. Journal of Cell Biology, 100 (5):1623–??, May 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/100/5/1623.

Franzini-Armstrong:1987:SCT

[FAKVM87] C. Franzini-Armstrong, L. J. Kenney, and E. Varriano-Marston. The structure of calsequestrin in triads of vertebrate skeletal muscle: a deep-etch study. Journal of Cell Biology, 105(1):49–??, July 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/105/1/
49.

Fallon:1985:POC

[Fal85] J. R. Fallon. Preferential outgrowth of central nervous system neurites on astrocytes and Schwann cells as compared with nonglial cells in vitro. Journal of Cell Biology, 100(1):198–??, January 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
100/1/198.

Fox:1988:IMS

[FBB^{+}88] J. E. Fox, J. K. Boyles, M. C. Berndt, P. K. Steffen, and L. K. Anderson. Identification of a membrane skeleton in platelets. Journal of Cell Biology, 106(5):1525–??, May 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/106/5/1525.

Fornieri:1987:LOA

[FBCQPR87] C. Fornieri, M. Baccarani-Contri, D. Quaglino, and I. Pasquali-Ronchetti. Lysyl oxidase activity and elastin/glycosaminoglycan interactions in growing chick and rat aortas. Journal of Cell Biology, 105(3):1463–??, September 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/105/3/1463.

Friedlander:1985:CEF

[FBE85] D. R. Friedlander, R. Brackenbury, and G. M. Edelman. Conversion of embryonic form to adult forms of N-CAM in vitro: results from de novo synthesis of adult forms. Journal of Cell Biology, 101(2):412–??, August 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/101/2/412.

Feige:1989:DEH

[FBF^{+}89] J. J. Feige, J. D. Bradley, K. Fryburg, J. Farris, L. C. Cousens, P. J. Barr, and A. Baird. Differential effects of heparin, fibronectin, and laminin on the phosphorylation of basic fibroblast growth factor by protein kinase C and the catalytic subunit of protein kinase A. Journal of Cell Biology, 109(6):3105–??, December 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
109/6/3105.

Fini:1989:UBC

[FBP89] M. E. Fini, W. G. Bendena, and M. L. Pardue. Unusual behavior of the cytoplasmic transcript of hsr omega: an abundant, stressinducible RNA that is translated but yields no detectable protein product. Journal of Cell Biology, 108(6):2045–??, June 1989. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/108/6/2045.

Ferris:1987:CHO

[FBPS87] A. L. Ferris, J. C. Brown, R. D. Park, and B. Storrie. Chinese hamster ovary cell lysosomes rapidly exchange contents. Journal of Cell Biology, 105(6):2703–??, December 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/105/6/2703.

Fedarko:1986:UHS

[FC86] N. S. Fedarko and H. E. Conrad. A unique heparan sulfate in the nuclei of hepatocytes: structural changes with the growth state of the cells. Journal of Cell Biology, 102(2):587–??, February 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/102/2/
587.

Fontaine:1989:LNA

[FC89a] B. Fontaine and J. P. Changeux. Localization of nicotinic acetylcholine receptor alpha-subunit transcripts during myogenesis and motor endplate development in the chick. Journal of Cell Biology, 108(3):1025–??, March 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/108/3/1025.

Frado:1989:SCI

[FC89b] L. L. Frado and R. Craig. Structural changes induced in Ca^{2+}- regulated myosin filaments by Ca^{2+} and ATP. Journal of Cell Biology, 109(2):529–??, August 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/109/2/529.

Farach-Carson:1989:CBA

[FCCC^{+}89] M. C. Farach-Carson, D. D. Carson, J. L. Collier, W. J. Lennarz, H. R. Park, and G. C. Wright. A calcium-binding, asparaginelinked oligosaccharide is involved in skeleton formation in the sea urchin embryo. Journal of Cell Biology, 109(3):1289–??, September 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/109/3/
1289.

Fessler:1987:DLC

[FCDF87] L. I. Fessler, A. G. Campbell, K. G. Duncan, and J. H. Fessler. Drosophila laminin: characterization and localization. Journal of Cell Biology, 105(5):2383–??, November 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/105/5/2383.

Fuller:1987:GAM

[FCH^{+}87] M. T. Fuller, J. H. Caulton, J. A. Hutchens, T. C. Kaufman, and E. C. Raff. Genetic analysis of microtubule structure: a beta-tubulin mutation causes the formation of aberrant microtubules in vivo and in vitro. Journal of Cell Biology, 104 (3):385–??, March 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/104/3/385.

Fuller:1988:MEP

[FCH^{+}88] M. T. Fuller, J. H. Caulton, J. A. Hutchens, T. C. Kaufman, and E. C. Raff. Mutations that encode partially functional beta 2 tubulin subunits have different effects on structurally different microtubule arrays. Journal of Cell Biology, 107(1):141–??, July 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/107/1/

Faust:1987:RSG

[FCK87] P. L. Faust, J. M. Chirgwin, and S. Kornfeld. Renin, a secretory glycoprotein, acquires phosphomannosyl residues. Journal of Cell Biology, 105(5):1947–??, November 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/105/5/1947.

Fulton:1986:TCN

[FCL86] C. Fulton, K. L. Cheng, and E. Y. Lai. Two calmodulins in Naegleria flagellates: characterization, intracellular segregation, and programmed regulation of mRNA abundance during differentiation. Journal of Cell Biology, 102(5):1671–??, May 1986. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/102/5/1671.

Fife:1985:ILP

[FCM85] R. S. Fife, B. Caterson, and S. L. Myers. Identification of link proteins in canine synovial cell cultures and canine articular cartilage. Journal of Cell Biology, 100(4):1050–??, April 1985. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/100/4/1050.

Ffrench-Constant:1986:MSA

[FCMK^{+}86] C. Ffrench-Constant, R. H. Miller, J. Kruse, M. Schachner, and M. C. Raff. Molecular specialization of astrocyte processes at nodes of Ranvier in rat optic nerve. Journal of Cell Biology, 102(3):844–??, March 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/102/3/844.

Ffrench-Constant:1989:REP

[FCVDH89] C. Ffrench-Constant, L. Van de Water, H. F. Dvorak, and R. O. Hynes. Reappearance of an embryonic pattern of fibronectin splicing during wound healing in the adult rat. Journal of Cell Biology, 109(2):903–??, August 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/109/2/903.

Fisher:1989:ADP

[FCZ^{+}89] S. J. Fisher, T. Y. Cui, L. Zhang, L. Hartman, K. Grahl, G. Y. Zhang, J. Tarpey, and C. H. Damsky. Adhesive and degradative properties of human placental cytotrophoblast cells in vitro.

Journal of Cell Biology, 109(2):891–??, August 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/109/2/891.

Fowler:1985:HEM

[FDB85] V. M. Fowler, J. Q. Davis, and V. Bennett. Human erythrocyte myosin: identification and purification. Journal of Cell Biology, 100(1):47–??, January 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/100/1/47.

Featherstone:1988:MAA

[FDG88] C. Featherstone, M. K. Darby, and L. Gerace. A monoclonal antibody against the nuclear pore complex inhibits nucleocytoplasmic transport of protein and RNA in vivo. Journal of Cell Biology, 107(4):1289–??, October 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/107/4/1289.

Fechheimer:1987:DDD

[Fec87] M. Fechheimer. The Dictyostelium discoideum 30,000-dalton protein is an actin filament-bundling protein that is selectively present in filopodia. Journal of Cell Biology, 104(6):1539–??, June 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/104/6/
1539.

Friederich:1988:ITS

[FFH88] E. Friederich, H. J. Fritz, and W. B. Huttner. Inhibition of tyrosine sulfation in the trans-Golgi retards the transport of a constitutively secreted protein to the cell surface. Journal of Cell Biology, 107(5):1655–??, November 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/107/5/1655.

Fallon:1989:ARM

[FG89] J. R. Fallon and C. E. Gelfman. Agrin-related molecules are concentrated at acetylcholine receptor clusters in normal and aneural developing muscle. Journal of Cell Biology, 108(4):1527–??, April 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/108/4/

Franke:1986:CEC

[FGA86] W. W. Franke, C. Grund, and T. Achtstätter. Co-expression of cytokeratins and neurofilament proteins in a permanent cell line: cultured rat PC12 cells combine neuronal and epithelial features. Journal of Cell Biology, 103(5):1933–??, November 1986. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/103/5/1933.

Friedlander:1986:NGF

[FGE86] D. R. Friedlander, M. Grumet, and G. M. Edelman. Nerve growth factor enhances expression of neuron-glia cell adhesion molecule in PC12 cells. Journal of Cell Biology, 102(2):413–??, February 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
102/2/413.

Finlay:1988:PST

[FGF88] B. B. Finlay, B. Gumbiner, and S. Falkow. Penetration of Salmonella through a polarized Madin–Darby canine kidney epithelial cell monolayer. Journal of Cell Biology, 107(1):221–??, July 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/107/1/
221.

Fairbairn:1985:EMN

[FGO^{+}85] S. Fairbairn, R. Gilbert, G. Ojakian, R. Schwimmer, and J. P. Quigley. The extracellular matrix of normal chick embryo fibroblasts: its effect on transformed chick fibroblasts and its proteolytic degradation by the transformants. Journal of Cell Biology, 101(5):1790–??, November 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/101/5/1790.

Featherstone:1985:NSG

[FGW85] C. Featherstone, G. Griffiths, and G. Warren. Newly synthesized G protein of vesicular stomatitis virus is not transported to the Golgi complex in mitotic cells. Journal of Cell Biology, 101(6): 2036–??, December 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/101/6/2036.

French:1986:LLS

[FH86] P. W. French and D. R. Hewish. Localization of low-sulfur keratin proteins in the wool follicle using monoclonal antibodies. Journal of Cell Biology, 102(4):1412–??, April 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/102/4/1412.

Friedlander:1988:FMC

[FHE88] D. R. Friedlander, S. Hoffman, and G. M. Edelman. Functional mapping of cytotactin: proteolytic fragments active in cellsubstrate adhesion. Journal of Cell Biology, 107(6):2329–??, December 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
107/6/2329.

Festoff:1986:PAM

[FHS^{+}86] B. W. Festoff, D. Hantaï, J. Soria, A. Thomaïdis, and C. Soria. Plasminogen activator in mammalian skeletal muscle: characteristics of effect of denervation on urokinase-like and tissue activator. Journal of Cell Biology, 103(4):1415–??, October 1986. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/103/4/1415.

Farrell:1987:PDM

[FJMW87] K. W. Farrell, M. A. Jordan, H. P. Miller, and L. Wilson. Phase dynamics at microtubule ends: the coexistence of microtubule length changes and treadmilling. Journal of Cell Biology, 104 (4):1035–??, January 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/104/4/1035.

Fisher:1987:AEP

[FJR87] C. Fisher, A. Jones, and D. R. Roop. Abnormal expression and processing of keratins in pupoid fetus (pf/pf) and repeated epilation (Er/Er) mutant mice. Journal of Cell Biology, 105(4): 1807–??, October 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/4/1807.

Firestone:1986:GRG

[FJY86] G. L. Firestone, N. J. John, and K. R. Yamamoto. Glucocorticoidregulated glycoprotein maturation in wild-type and mutant rat cell lines. Journal of Cell Biology, 103(6):2323–??, December 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/103/6/
2323.

Fontaine:1987:CGR

[FKC87] B. Fontaine, A. Klarsfeld, and J. P. Changeux. Calcitonin generelated peptide and muscle activity regulate acetylcholine receptor alpha-subunit mRNA levels by distinct intracellular pathways. Journal of Cell Biology, 105(3):1337–??, September 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/105/3/1337.

Froshauer:1988:ARR

[FKH88] S. Froshauer, J. Kartenbeck, and A. Helenius. Alphavirus RNA replicase is located on the cytoplasmic surface of endosomes and lysosomes. Journal of Cell Biology, 107(6):2075–??, December 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/107/6/
2075.

Friesel:1987:IEC

[FKM87] R. Friesel, A. Komoriya, and T. Maciag. Inhibition of endothelial cell proliferation by gamma-interferon. Journal of Cell Biology, 104(3):689–??, March 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/104/3/689.

Fey:1986:NSN

[FKP86] E. G. Fey, G. Krochmalnic, and S. Penman. The nonchromatin substructures of the nucleus: the ribonucleoprotein (RNP)- containing and RNP-depleted matrices analyzed by sequential fractionation and resinless section electron microscopy. Journal of Cell Biology, 102(5):1654–??, May 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/102/5/1654.

Fath:1988:TCA

[FL88] K. R. Fath and R. J. Lasek. Two classes of actin microfilaments are associated with the inner cytoskeleton of axons. Journal of Cell Biology, 107(2):613–??, August 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/107/2/613.

Feghali:1989:MGC

[FL89] R. Feghali and L. A. Leinwand. Molecular genetic characterization of a developmentally regulated human perinatal myosin heavy chain. Journal of Cell Biology, 108(5):1791–??, May 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/108/5/1791.

Flucher:1986:SPX

[FLBP^{+}86] B. E. Flucher, C. Lenglachner-Bachinger, K. Pohlhammer, H. Adam, and C. Mollay. Skin peptides in Xenopus laevis: morphological requirements for precursor processing in developing and regenerating granular skin glands. Journal of Cell Biology, 103(6):2299–??, December 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/103/6/2299.

Foisner:1988:CAP

[FLH^{+}88] R. Foisner, F. E. Leichtfried, H. Herrmann, J. V. Small, D. Lawson, and G. Wiche. Cytoskeleton-associated plectin: in situ localization, in vitro reconstitution, and binding to immobilized intermediate filament proteins. Journal of Cell Biology, 106 (3):723–??, March 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/106/3/723.

Fakan:1986:IMV

[FLM86] S. Fakan, G. Leser, and T. E. Martin. Immunoelectron microscope visualization of nuclear ribonucleoprotein antigens within spread transcription complexes. Journal of Cell Biology, 103 (4):1153–??, October 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/103/4/1153.

Funderburg:1986:CNS

[FM86] F. M. Funderburg and R. R. Markwald. Conditioning of native substrates by chondroitin sulfate proteoglycans during cardiac mesenchymal cell migration. Journal of Cell Biology, 103(6): 2475–??, December 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/103/6/2475.

Fisher:1989:QAC

[FMG89] P. R. Fisher, R. Merkl, and G. Gerisch. Quantitative analysis of cell motility and chemotaxis in Dictyostelium discoideum by using an image processing system and a novel chemotaxis chamber providing stationary chemical gradients. Journal of Cell Biology, 108(3):973–??, March 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/108/3/973.

Fleming:1989:DTJ

[FMJS89] T. P. Fleming, J. McConnell, M. H. Johnson, and B. R. Stevenson. Development of tight junctions de novo in the mouse early embryo: control of assembly of the tight junction-specific protein, ZO-1. Journal of Cell Biology, 108(4):1407–??, April 1989. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/108/4/1407.

Foley:1986:LDL

[FMK^{+}86] M. Foley, A. N. MacGregor, J. R. Kusel, P. B. Garland, T. Downie, and I. Moore. The lateral diffusion of lipid probes in the surface membrane of Schistosoma mansoni. Journal of Cell Biology, 103(3):807–??, September 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/103/3/807.

Froehner:1987:PMP

[FMT^{+}87] S. C. Froehner, A. A. Murnane, M. Tobler, H. B. Peng, and R. Sealock. A postsynaptic Mr 58,000 (58K) protein concentrated at acetylcholine receptor-rich sites in Torpedo electroplaques and skeletal muscle. Journal of Cell Biology, 104(6):1633–??, June 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/104/6/
1633.

Finlay:1987:IVN

[FNPF87] D. R. Finlay, D. D. Newmeyer, T. M. Price, and D. J. Forbes. Inhibition of in vitro nuclear transport by a lectin that binds to nuclear pores. Journal of Cell Biology, 104(2):189–??, February 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/104/2/

Furst:1988:OTF

[FONW88] D. O. Fürst, M. Osborn, R. Nave, and K. Weber. The organization of titin filaments in the half-sarcomere revealed by monoclonal antibodies in immunoelectron microscopy: a map of ten nonrepetitive epitopes starting at the Z line extends close to the M line. Journal of Cell Biology, 106(5):1563–??, May 1988. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/106/5/1563.

Fuller:1985:EHS

[FOW^{+}85] G. M. Fuller, J. M. Otto, B. M. Woloski, C. T. McGary, and M. A. Adams. The effects of hepatocyte stimulating factor on fibrinogen biosynthesis in hepatocyte monolayers. Journal of Cell Biology, 101(4):1481–??, October 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/101/4/1481.

Furst:1989:MME

[FOW89] D. O. Fürst, M. Osborn, and K. Weber. Myogenesis in the mouse embryo: differential onset of expression of myogenic proteins and the involvement of titin in myofibril assembly. Journal of Cell Biology, 109(2):517–??, August 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/109/2/517.

Flicker:1985:SSI

[FPS^{+}85] P. F. Flicker, G. Peltz, M. P. Sheetz, P. Parham, and J. A. Spudich. Site-specific inhibition of myosin-mediated motility in vitro by monoclonal antibodies. Journal of Cell Biology, 100 (4):1024–??, April 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/100/4/1024.

Frelinger:1986:TCS

[FR86] A. L. Frelinger and U. Rutishauser. Topography of N-CAM structural and functional determinants. II. Placement of monoclonal antibody epitopes. Journal of Cell Biology, 103(5): 1729–??, November 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/103/5/1729.

Frazier:1987:TMA

[Fra87] W. A. Frazier. Thrombospondin: a modular adhesive glycoprotein of platelets and nucleated cells. Journal of Cell Biology, 105(2):625–??, August 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/2/625.

Finidori:1987:IHI

[FRG^{+}87] J. Finidori, L. Rizzolo, A. Gonzalez, G. Kreibich, M. Adesnik, and D. D. Sabatini. The influenza hemagglutinin insertion signal is not cleaved and does not halt translocation when presented to the endoplasmic reticulum membrane as part of a translocating polypeptide. Journal of Cell Biology, 104(6):1705–??, June 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/104/6/
1705.

Fliesler:1985:MMR

[FRH85] S. J. Fliesler, M. E. Rayborn, and J. G. Hollyfield. Membrane morphogenesis in retinal rod outer segments: inhibition by tunicamycin. Journal of Cell Biology, 100(2):574–??, February 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/100/2/
574.

Fritze:1985:AHS

[FRR85] L. M. Fritze, C. F. Reilly, and R. D. Rosenberg. An antiproliferative heparan sulfate species produced by postconfluent smooth muscle cells. Journal of Cell Biology, 100(4):1041–??, April 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/100/4/1041.

Fujimoto:1986:ISE

[FS86a] T. Fujimoto and S. J. Singer. Immunocytochemical studies of endothelial cells in vivo. I. The presence of desmin only, or of desmin plus vimentin, or vimentin only, in the endothelial cells of different capillaries of the adult chicken. Journal of Cell Biology, 103 (6):2775–??, December 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/103/6/2775.

Fuller:1986:TRP

[FS86b] S. D. Fuller and K. Simons. Transferrin receptor polarity and recycling accuracy in “tight” and “leaky” strains of Madin– Darby canine kidney cells. Journal of Cell Biology, 103(5): 1767–??, November 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/103/5/1767.

Fox:1987:DFG

[FS87] L. A. Fox and W. S. Sale. Direction of force generated by the inner row of dynein arms on flagellar microtubules. Journal of Cell Biology, 105(4):1781–??, October 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/105/4/1781.

Fields:1988:MKI

[FS88a] A. P. Fields and J. H. Shaper. A major 62-kD intranuclear matrix polypeptide is a component of metaphase chromosomes. Journal of Cell Biology, 107(3):833–??, September 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/107/3/833.

Forscher:1988:ACO

[FS88b] P. Forscher and S. J. Smith. Actions of cytochalasins on the organization of actin filaments and microtubules in a neuronal growth cone. Journal of Cell Biology, 107(4):1505–??, October 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/107/4/
1505.

Facci:1988:RGA

[FSFL88] L. Facci, S. D. Skaper, M. Favaron, and A. Leon. A role for gangliosides in astroglial cell differentiation in vitro. Journal of Cell Biology, 106(3):821–??, March 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/106/3/821.

Feuerstein:1988:NMP

[FSM88] N. Feuerstein, S. Spiegel, and J. J. Mond. The nuclear matrix protein, numatrin (B23), is associated with growth factorinduced mitogenesis in Swiss 3T3 fibroblasts and with T lymphocyte proliferation stimulated by lectins and anti-T cell antigen receptor antibody. Journal of Cell Biology, 107(5):1629–??, November 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/107/5/
1629.

Frost:1987:MAT

[FTC87] G. H. Frost, W. C. Thompson, and D. H. Carney. Monoclonal antibody to the thrombin receptor stimulates DNA synthesis in combination with gamma-thrombin or phorbol myristate acetate. Journal of Cell Biology, 105(6):2551–??, December 1987. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/105/6/2551.

Flanders:1989:TGF

[FTC^{+}89] K. C. Flanders, N. L. Thompson, D. S. Cissel, E. Van Obberghen-Schilling, C. C. Baker, M. E. Kass, L. R. Ellingsworth, A. B. Roberts, and M. B. Sporn. Transforming growth factor-beta 1: histochemical localization with antibodies to different epitopes. Journal of Cell Biology, 108(2):653–??, February 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/108/2/653.

Fambrough:1988:SLG

[FTLSS88] D. M. Fambrough, K. Takeyasu, J. Lippincott-Schwarz, and N. R. Siegel. Structure of LEP100, a glycoprotein that shuttles between lysosomes and the plasma membrane, deduced from the nucleotide sequence of the encoding cDNA. Journal of Cell Biology, 106(1):61–??, January 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/106/1/61.

Francis:1985:MOC

[FW85] G. R. Francis and R. H. Waterston. Muscle organization in Caenorhabditis elegans: localization of proteins implicated in thin filament attachment and I-band organization. Journal of Cell Biology, 101(4):1532–??, October 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/101/4/1532.

Furukawa:1988:MCP

[FWF88] R. Furukawa, J. E. Wampler, and M. Fechheimer. Measurement of the cytoplasmic pH of Dictyostelium discoideum using a low light level microspectrofluorometer. Journal of Cell Biology, 107 (6):2541–??, December 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/107/6/2541.

Faust:1987:EHC

[FWP^{+}87] P. L. Faust, D. A. Wall, E. Perara, V. R. Lingappa, and S. Kornfeld. Expression of human cathepsin D in Xenopus oocytes: phosphorylation and intracellular targeting. Journal of Cell Biology, 105(5):1937–??, November 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/105/5/1937.

Fisher:1986:EMS

[FY86] K. A. Fisher and K. C. Yanagimoto. Effect of membrane splitting on transmembrane polypeptides. Journal of Cell Biology, 102 (2):551–??, February 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/102/2/551.

Ferguson:1988:LPS

[FYRK88] D. G. Ferguson, E. F. Young, L. Raeymaekers, and E. G. Kranias. Localization of phospholamban in smooth muscle using immunogold electron microscopy. Journal of Cell Biology, 107 (2):555–??, August 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/107/2/555.

Fumagalli:1985:CAP

[FZ85] G. Fumagalli and A. Zanini. In cow anterior pituitary, growth hormone and prolactin can be packed in separate granules of the same cell. Journal of Cell Biology, 100(6):2019–??, June 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/100/6/
2019.

Gilson:1986:RRE

[GAB86] C. A. Gilson, N. Ackland, and B. Burnside. Regulation of reactivated elongation in lysed cell models of teleost retinal cones by cAMP and calcium. Journal of Cell Biology, 102(3):1047–??, March 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
102/3/1047.

Giulian:1986:BPGa

[GABT86] D. Giulian, R. L. Allen, T. J. Baker, and Y. Tomozawa. Brain peptides and glial growth. I. Glia-promoting factors as regulators of gliogenesis in the developing and injured central nervous system. Journal of Cell Biology, 102(3):803–??, March 1986. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/102/3/803.

Goodenough:1985:SCA

[GACOH85] U. W. Goodenough, W. S. Adair, P. Collin-Osdoby, and J. E. Heuser. Structure of the Chlamydomonas agglutinin and related flagellar surface proteins in vitro and in situ. Journal of Cell Biology, 101(3):924–??, September 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/101/3/924.

Gurney:1986:STA

[GAH86] M. E. Gurney, B. R. Apatoff, and S. P. Heinrich. Suppression of terminal axonal sprouting at the neuromuscular junction by monoclonal antibodies against a muscle-derived antigen of 56,000 daltons. Journal of Cell Biology, 102(6):2264–??, June 1986. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/102/6/2264.

Gehlsen:1988:IVT

[GAPR88] K. R. Gehlsen, W. S. Argraves, M. D. Pierschbacher, and E. Ruoslahti. Inhibition of in vitro tumor cell invasion by Arg–Gly– Asp–containing synthetic peptides. Journal of Cell Biology, 106 (3):925–??, March 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/106/3/925.

Gabel:1985:PAL

[GB85a] C. A. Gabel and J. E. Bergmann. Processing of the asparaginelinked oligosaccharides of secreted and intracellular forms of the vesicular stomatitis virus G protein: in vivo evidence of Golgi apparatus compartmentalization. Journal of Cell Biology, 101 (2):460–??, August 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/101/2/460.

Giulian:1985:PRA

[GB85b] D. Giulian and T. J. Baker. Peptides released by ameboid microglia regulate astroglial proliferation. Journal of Cell Biology, 101(6):2411–??, December 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/101/6/2411.

Goldberg:1986:SAF

[GB86a] D. J. Goldberg and D. W. Burmeister. Stages in axon formation: observations of growth of Aplysia axons in culture using videoenhanced contrast-differential interference contrast microscopy. Journal of Cell Biology, 103(5):1921–??, November 1986. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/103/5/1921.

Gundersen:1986:DTN

[GB86b] G. G. Gundersen and J. C. Bulinski. Distribution of tyrosinated and nontyrosinated alpha-tubulin during mitosis. Journal of Cell Biology, 102(3):1118–??, March 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/102/3/1118.

Georgatos:1987:LBC

[GB87a] S. D. Georgatos and G. Blobel. Lamin B constitutes an intermediate filament attachment site at the nuclear envelope. Journal of Cell Biology, 105(1):117–??, July 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/105/1/117.

Georgatos:1987:TDA

[GB87b] S. D. Georgatos and G. Blobel. Two distinct attachment sites for vimentin along the plasma membrane and the nuclear envelope in avian erythrocytes: a basis for a vectorial assembly of intermediate filaments. Journal of Cell Biology, 105(1):105–??, July 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/105/1/
105.

Gorbsky:1989:MKF

[GB89a] G. J. Gorbsky and G. G. Borisy. Microtubules of the kinetochore fiber turn over in metaphase but not in anaphase. Journal of Cell Biology, 109(2):653–??, August 1989. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/109/2/653.

Goslin:1989:EOD

[GB89b] K. Goslin and G. Banker. Experimental observations on the development of polarity by hippocampal neurons in culture. Journal of Cell Biology, 108(4):1507–??, April 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/108/4/1507.

Goldberg:1989:RMG

[GBB89] M. Goldberg, L. S. Belkowski, and B. R. Bloom. Regulation of macrophage growth and antiviral activity by interferon-gamma. Journal of Cell Biology, 109(3):1331–??, September 1989. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/109/3/1331.

Goslin:1989:RCO

[GBBS89] K. Goslin, E. Birgbauer, G. Banker, and F. Solomon. The role of cytoskeleton in organizing growth cones: a microfilamentassociated growth cone component depends upon microtubules for its localization. Journal of Cell Biology, 109(4):1621–??, October 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
109/4/1621.

Golan:1986:SSM

[GBC^{+}86] D. E. Golan, C. S. Brown, C. M. Cianci, S. T. Furlong, and J. P. Caulfield. Schistosomula of Schistosoma mansoni use lysophosphatidylcholine to lyse adherent human red blood cells and immobilize red cell membrane components. Journal of Cell Biology, 103(3):819–??, September 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/103/3/819.

Giugni:1987:EFI

[GBH87] T. D. Giugni, D. L. Braslau, and H. T. Haigler. Electric field-induced redistribution and postfield relaxation of epidermal growth factor receptors on A431 cells. Journal of Cell Biology, 104(5):1291–??, May 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/104/5/1291.

Griffiths:1989:QAE

[GBM89] G. Griffiths, R. Back, and M. Marsh. A quantitative analysis of the endocytic pathway in baby hamster kidney cells. Journal of Cell Biology, 109(6):2703–??, December 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/109/6/2703.

Gramzow:1986:IFC

[GBU^{+}86] M. Gramzow, M. Bachmann, G. Uhlenbruck, A. Dorn, and W. E. Müller. Identification and further characterization of the specific cell binding fragment from sponge aggregation factor. Journal of Cell Biology, 102(4):1344–??, April 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/102/4/1344.

Goff:1987:SCP

[GC87] L. J. Goff and A. W. Coleman. The solution to the cytological paradox of isomorphy. Journal of Cell Biology, 104(3):739–??, March 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
104/3/739.

Gould:1986:PTK

[GCBH86] K. L. Gould, J. A. Cooper, A. Bretscher, and T. Hunter. The protein-tyrosine kinase substrate, p81, is homologous to a chicken microvillar core protein. Journal of Cell Biology, 102(2):660–??, February 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
102/2/660.

Garcia:1989:PDS

[GCC^{+}89] A. Garcia, E. Coudrier, J. Carboni, J. Anderson, J. Vandekerkhove, M. Mooseker, D. Louvard, and M. Arpin. Partial deduced sequence of the 110-kD-calmodulin complex of the avian intestinal microvillus shows that this mechanoenzyme is a member of the myosin I family. Journal of Cell Biology, 109(6): 2895–??, December 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/6/2895.

Grinstein:1985:OPE

[GCGR85] S. Grinstein, S. Cohen, J. D. Goetz, and A. Rothstein. Osmotic and phorbol ester-induced activation of Na+/H+ exchange: possible role of protein phosphorylation in lymphocyte volume regulation. Journal of Cell Biology, 101(1):269–??, July 1985. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/101/1/269.

Gerstenfeld:1988:CEU

[GCK^{+}88] L. C. Gerstenfeld, S. D. Chipman, C. M. Kelly, K. J. Hodgens, D. D. Lee, and W. J. Landis. Collagen expression, ultrastructural assembly, and mineralization in cultures of chicken embryo osteoblasts. Journal of Cell Biology, 106(3):979–??, March 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/106/3/
979.

Grove:1985:MMP

[GCPE85] B. K. Grove, L. Cerny, J. C. Perriard, and H. M. Eppenberger. Myomesin and M-protein: expression of two M-band proteins in pectoral muscle and heart during development. Journal of Cell Biology, 101(4):1413–??, October 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/101/4/1413.

Gajdusek:1986:ACR

[GCR^{+}86] C. Gajdusek, S. Carbon, R. Ross, P. Nawroth, and D. Stern. Activation of coagulation releases endothelial cell mitogens. Journal of Cell Biology, 103(2):419–??, August 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/103/2/419.

Gennarini:1989:MNC

[GCR^{+}89] G. Gennarini, G. Cibelli, G. Rougon, M. G. Mattei, and C. Goridis. The mouse neuronal cell surface protein F3: a phosphatidylinositol-anchored member of the immunoglobulin superfamily related to chicken contactin. Journal of Cell Biology, 109(2):775–??, August 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/2/775.

Giancotti:1986:FPM

[GCT86] F. G. Giancotti, P. M. Comoglio, and G. Tarone. Fibronectinplasma membrane interaction in the adhesion of hemopoietic cells. Journal of Cell Biology, 103(2):429–??, August 1986. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/103/2/429.

Garchon:1989:XGP

[GD89] H. J. Garchon and M. M. Davis. The XLR gene product defines a novel set of proteins stabilized in the nucleus by zinc ions. Journal of Cell Biology, 108(3):779–??, March 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/108/3/779.

Greene:1986:SIR

[GDC^{+}86] L. A. Greene, S. A. Drexler, J. L. Connolly, A. Rukenstein, and S. H. Green. Selective inhibition of responses to nerve growth factor and of microtubule-associated protein phosphorylation by activators of adenylate cyclase. Journal of Cell Biology, 103(5): 1967–??, November 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/103/5/1967.

Geller:1988:AFC

[GDD88] H. M. Geller and M. Dubois-Dalcq. Antigenic and functional characterization of a rat central nervous system-derived cell line immortalized by a retroviral vector. Journal of Cell Biology, 107 (5):1977–??, November 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/107/5/1977.

Gomer:1986:CSD

[GDF86] R. H. Gomer, S. Datta, and R. A. Firtel. Cellular and subcellular distribution of a cAMP-regulated prestalk protein and prespore protein in Dictyostelium discoideum: a study on the ontogeny of prestalk and prespore cells. Journal of Cell Biology, 103(5): 1999–??, November 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/103/5/1999.

Godfrey:1988:ARA

[GDM^{+}88] E. W. Godfrey, M. E. Dietz, A. L. Morstad, P. A. Wallskog, and D. E. Yorde. Acetylcholine receptor-aggregating proteins are associated with the extracellular matrix of many tissues in Torpedo. Journal of Cell Biology, 106(4):1263–??, April 1988. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/106/4/1263.

Galvin:1985:MEM

[GDO^{+}85] N. J. Galvin, V. M. Dixit, K. M. O’Rourke, S. A. Santoro, G. A. Grant, and W. A. Frazier. Mapping of epitopes for monoclonal antibodies against human platelet thrombospondin with electron microscopy and high sensitivity amino acid sequencing. Journal of Cell Biology, 101(4):1434–??, October 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/101/4/1434.

Graeve:1989:PEM

[GDRB89] L. Graeve, K. Drickamer, and E. Rodriguez-Boulan. Polarized endocytosis by Madin–Darby canine kidney cells transfected with functional chicken liver glycoprotein receptor. Journal of Cell Biology, 109(6):2809–??, December 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/109/6/2809.

Griffith:1987:MLC

[GDS87] L. M. Griffith, S. M. Downs, and J. A. Spudich. Myosin light chain kinase and myosin light chain phosphatase from Dictyostelium: effects of reversible phosphorylation on myosin structure and function. Journal of Cell Biology, 104(5):1309–??, May 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/104/5/
1309.

Goodman:1987:TDC

[GDvdM87] S. L. Goodman, R. Deutzmann, and K. von der Mark. Two distinct cell-binding domains in laminin can independently promote nonneuronal cell adhesion and spreading. Journal of Cell Biology, 105(1):589–??, July 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/1/589.

Gething:1986:SMM

[GDYW86] M. J. Gething, R. W. Doms, D. York, and J. White. Studies on the mechanism of membrane fusion: site-specific mutagenesis of the hemagglutinin of influenza virus. Journal of Cell Biology, 102(1):11–??, January 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/102/1/11.

Grumet:1988:NGC

[GE88] M. Grumet and G. M. Edelman. Neuron-glia cell adhesion molecule interacts with neurons and astroglia via different binding mechanisms. Journal of Cell Biology, 106(2):487–??, February 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/106/2/
487.

Gibson:1985:TXC

[GF85] G. J. Gibson and M. H. Flint. Type X collagen synthesis by chick sternal cartilage and its relationship to endochondral development. Journal of Cell Biology, 101(1):277–??, July 1985. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/101/1/277.

Gabel:1986:LET

[GF86a] C. A. Gabel and S. A. Foster. Lysosomal enzyme trafficking in mannose 6-phosphate receptor-positive mouse L-cells: demonstration of a steady state accumulation of phosphorylated acid hydrolases. Journal of Cell Biology, 102(3):943–??, March 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/102/3/943.

Gabel:1986:MPR

[GF86b] C. A. Gabel and S. A. Foster. Mannose 6-phosphate receptormediated endocytosis of acid hydrolases: internalization of betaglucuronidase is accompanied by a limited dephosphorylation. Journal of Cell Biology, 103(5):1817–??, November 1986. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/103/5/1817.

Gabel:1987:PMA

[GF87] C. A. Gabel and S. A. Foster. Postendocytic maturation of acid hydrolases: evidence of prelysosomal processing. Journal of Cell Biology, 105(4):1561–??, October 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/105/4/1561.

Glass:1988:ISD

[GF88] C. Glass and E. Fuchs. Isolation, sequence, and differential expression of a human K7 gene in simple epithelial cells. Journal of Cell Biology, 107(4):1337–??, October 1988. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/107/4/1337.

Ghinea:1988:IAB

[GFA^{+}88] N. Ghinea, A. Fixman, D. Alexandru, D. Popov, M. Hasu, L. Ghitescu, M. Eskenasy, M. Simionescu, and N. Simionescu. Identification of albumin-binding proteins in capillary endothelial cells. Journal of Cell Biology, 107(1):231–??, July 1988. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/107/1/231.

Griffiths:1989:DNG

[GFB^{+}89] G. Griffiths, S. D. Fuller, R. Back, M. Hollinshead, S. Pfeiffer, and K. Simons. The dynamic nature of the Golgi complex. Journal of Cell Biology, 108(2):277–??, February 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/108/2/277.

Godfraind:1989:VAG

[GFHDD89] C. Godfraind, V. L. Friedrich, K. V. Holmes, and M. Dubois-Dalcq. In vivo analysis of glial cell phenotypes during a viral demyelinating disease in mice. Journal of Cell Biology, 109(5): 2405–??, November 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/5/2405.

Glukhova:1989:EED

[GFS^{+}89] M. A. Glukhova, M. G. Frid, B. V. Shekhonin, T. D. Vasilevskaya, J. Grunwald, M. Saginati, and V. E. Koteliansky. Expression of extra domain A fibronectin sequence in vascular smooth muscle cells is phenotype dependent. Journal of Cell Biology, 109(1):357–??, July 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/109/1/357.

Ghitescu:1986:SBS

[GFSS86] L. Ghitescu, A. Fixman, M. Simionescu, and N. Simionescu. Specific binding sites for albumin restricted to plasmalemmal vesicles of continuous capillary endothelium: receptor-mediated transcytosis. Journal of Cell Biology, 102(4):1304–??, April 1986. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/102/4/1304.

Glenney:1985:CCR

[GG85] J. R. Glenney and P. Glenney. Comparison of Ca++-regulated events in the intestinal brush border. Journal of Cell Biology, 100(3):754–??, March 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/100/3/754.

Grimes:1987:CPC

[GG87a] G. W. Grimes and R. H. Gavin. Ciliary protein conservation during development in the ciliated protozoan, Oxytricha. Journal of Cell Biology, 105(6):2855–??, December 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/105/6/2855.

Guidry:1987:HMO

[GG87b] C. Guidry and F. Grinnell. Heparin modulates the organization of hydrated collagen gels and inhibits gel contraction by fibroblasts. Journal of Cell Biology, 104(4):1097–??, January 1987. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/104/4/1097.

Gresham:1989:NMI

[GGA^{+}89] H. D. Gresham, J. L. Goodwin, P. M. Allen, D. C. Anderson, and E. J. Brown. A novel member of the integrin receptor family mediates Arg–Gly–Asp–stimulated neutrophil phagocytosis. Journal of Cell Biology, 108(5):1935–??, May 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/108/5/1935.

Gruenberg:1989:CEE

[GGH89] J. Gruenberg, G. Griffiths, and K. E. Howell. Characterization of the early endosome and putative endocytic carrier vesicles in vivo and with an assay of vesicle fusion in vitro. Journal of Cell Biology, 108(4):1301–??, April 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/108/4/1301.

Green:1987:RBI

[GGJ^{+}87] K. J. Green, B. Geiger, J. C. Jones, J. C. Talian, and R. D. Goldman. The relationship between intermediate filaments and microfilaments before and during the formation of desmosomes and adherens-type junctions in mouse epidermal keratinocytes.

Journal of Cell Biology, 104(5):1389–??, May 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/104/5/1389.

Goodenough:1986:CCR

[GGMH86] U. W. Goodenough, B. Gebhart, R. P. Mecham, and J. E. Heuser. Crystals of the Chlamydomonas reinhardtii cell wall: polymerization, depolymerization, and purification of glycoprotein monomers. Journal of Cell Biology, 103(2):405–??, August 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/103/2/
405.

Geuens:1986:UCT

[GGN^{+}86] G. Geuens, G. G. Gundersen, R. Nuydens, F. Cornelissen, J. C. Bulinski, and M. DeBrabander. Ultrastructural colocalization of tyrosinated and detyrosinated alpha-tubulin in interphase and mitotic cells. Journal of Cell Biology, 103(5):1883–??, November 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/103/5/
1883.

Gottlieb:1986:SEV

[GGR^{+}86] T. A. Gottlieb, A. Gonzalez, L. Rizzolo, M. J. Rindler, M. Adesnik, and D. D. Sabatini. Sorting and endocytosis of viral glycoproteins in transfected polarized epithelial cells. Journal of Cell Biology, 102(4):1242–??, April 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/102/4/1242.

Garcia:1989:ECE

[GGR89] J. V. Garcia, B. D. Gehm, and M. R. Rosner. An evolutionarily conserved enzyme degrades transforming growth factor-alpha as well as insulin. Journal of Cell Biology, 109(3):1301–??, September 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/109/3/
1301.

Goodenough:1985:CCW

[GH85a] U. W. Goodenough and J. E. Heuser. The Chlamydomonas cell wall and its constituent glycoproteins analyzed by the quickfreeze, deep-etch technique. Journal of Cell Biology, 101(4): 1550–??, October 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/101/4/1550.

Goodenough:1985:SID

[GH85b] U. W. Goodenough and J. E. Heuser. Substructure of inner dynein arms, radial spokes, and the central pair/projection complex of cilia and flagella. Journal of Cell Biology, 100(6):2008–??, June 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/100/6/
2008.

Grandori:1988:PSC

[GH88] C. Grandori and H. Hanafusa. p60c-src is complexed with a cellular protein in subcellular compartments involved in exocytosis. Journal of Cell Biology, 107(6):2125–??, December 1988. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/107/6/2125.

Grigoriadis:1988:DMF

[GHA88] A. E. Grigoriadis, J. N. Heersche, and J. E. Aubin. Differentiation of muscle, fat, cartilage, and bone from progenitor cells present in a bone-derived clonal cell population: effect of dexamethasone. Journal of Cell Biology, 106(6):2139–??, June 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/106/6/2139.

Goud:1985:AUP

[GHL85] B. Goud, C. Huet, and D. Louvard. Assembled and unassembled pools of clathrin: a quantitative study using an enzyme immunoassay. Journal of Cell Biology, 100(2):521–??, February 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/100/2/
521.

Geuens:1989:MDI

[GHL^{+}89] G. Geuens, A. M. Hill, N. Levilliers, A. Adoutte, and M. De-Brabander. Microtubule dynamics investigated by microinjection of Paramecium axonemal tubulin: lack of nucleation but proximal assembly of microtubules at the kinetochore during prometaphase. Journal of Cell Biology, 108(3):939–??, March 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/108/3/

Gallagher:1988:ACS

[GHW^{+}88] P. Gallagher, J. Henneberry, I. Wilson, J. Sambrook, and M. J. Gething. Addition of carbohydrate side chains at novel sites on influenza virus hemagglutinin can modulate the folding, transport, and activity of the molecule. Journal of Cell Biology, 107 (6):2059–??, December 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/107/6/2059.

Glasser:1988:DMF

[GJD^{+}88] S. R. Glasser, J. Julian, G. L. Decker, J. P. Tang, and D. D. Carson. Development of morphological and functional polarity in primary cultures of immature rat uterine epithelial cells. Journal of Cell Biology, 107(6):2409–??, December 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/107/6/2409.

Gard:1985:PDI

[GK85] D. L. Gard and M. W. Kirschner. A polymer-dependent increase in phosphorylation of beta-tubulin accompanies differentiation of a mouse neuroblastoma cell line. Journal of Cell Biology, 100(3):764–??, March 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/100/3/764.

Glatzer:1986:DCD

[GK86] K. H. Glätzer and P. M. Kloetzel. Differential chromosomal distribution of ribonucleoprotein antigens in nuclei of Drosophila spermatocytes. Journal of Cell Biology, 103(6):2113–??, December 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/103/6/
2113.

Gard:1987:MAC

[GK87a] D. L. Gard and M. W. Kirschner. Microtubule assembly in cytoplasmic extracts of Xenopus oocytes and eggs. Journal of Cell Biology, 105(5):2191–??, November 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/105/5/2191.

Gard:1987:MAP

[GK87b] D. L. Gard and M. W. Kirschner. A microtubule-associated protein from Xenopus eggs that specifically promotes assembly at the plus-end. Journal of Cell Biology, 105(5):2203–??, November 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/105/5/
2203.

Gundersen:1987:PDA

[GKB87] G. G. Gundersen, S. Khawaja, and J. C. Bulinski. Postpolymerization detyrosination of alpha-tubulin: a mechanism for subcellular differentiation of microtubules. Journal of Cell Biology, 105(1):251–??, July 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/1/251.

Gundersen:1989:GSP

[GKB89] G. G. Gundersen, S. Khawaja, and J. C. Bulinski. Generation of a stable, posttranslationally modified microtubule array is an early event in myogenic differentiation. Journal of Cell Biology, 109 (5):2275–??, November 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/5/2275.

Gruijters:1987:IML

[GKBG87] W. T. Gruijters, J. Kistler, S. Bullivant, and D. A. Goodenough. Immunolocalization of MP70 in lens fiber 16-17-nm intercellular junctions. Journal of Cell Biology, 104(3):565–??, March 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/104/3/
565.

Glass:1985:SEH

[GKF85] C. Glass, K. H. Kim, and E. Fuchs. Sequence and expression of a human type II mesothelial keratin. Journal of Cell Biology, 101 (6):2366–??, December 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/101/6/2366.

Gimlich:1988:SDE

[GKG88] R. L. Gimlich, N. M. Kumar, and N. B. Gilula. Sequence and developmental expression of mRNA coding for a gap junction protein in Xenopus. Journal of Cell Biology, 107(3): 1065–??, September 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/107/3/1065.

Gould:1989:CTS

[GKH^{+}89] S. J. Gould, G. A. Keller, N. Hosken, J. Wilkinson, and S. Subramani. A conserved tripeptide sorts proteins to peroxisomes. Journal of Cell Biology, 108(5):1657–??, May 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/108/5/1657.

Geering:1987:MCA

[GKR87] K. Geering, J. P. Kraehenbuhl, and B. C. Rossier. Maturation of the catalytic alpha-subunit of Na,K–ATPase during intracellular transport. Journal of Cell Biology, 105(6):2613–??, December 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/105/6/
2613.

Genton:1987:PEI

[GKS87a] C. Genton, E. K. Kruithof, and W. D. Schleuning. Phorbol ester induces the biosynthesis of glycosylated and nonglycosylated plasminogen activator inhibitor 2 in high excess over urokinasetype plasminogen activator in human U-937 lymphoma cells. Journal of Cell Biology, 104(3):705–??, March 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/104/3/705.

Gould:1987:IPT

[GKS87b] S. G. Gould, G. A. Keller, and S. Subramani. Identification of a peroxisomal targeting signal at the carboxy terminus of firefly luciferase. Journal of Cell Biology, 105(6):2923–??, December 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/105/6/
2923.

Gould:1988:IPT

[GKS88] S. J. Gould, G. A. Keller, and S. Subramani. Identification of peroxisomal targeting signals located at the carboxy terminus of four peroxisomal proteins. Journal of Cell Biology, 107(3): 897–??, September 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/107/3/897.

Glenney:1989:INM

[GKZ89] J. R. Glenney, M. S. Kindy, and L. Zokas. Isolation of a new member of the S100 protein family: amino acid sequence, tissue, and subcellular distribution. Journal of Cell Biology, 108 (2):569–??, February 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/108/2/569.

Grobner:1985:IAE

[GL85] P. Gröbner and P. Loidl. An immunological approach to enrich a mitotic stimulator and to reveal G2-phase-specific proteins in Physarum polycephalum. Journal of Cell Biology, 100 (6):1930–??, June 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/100/6/1930.

Glabe:1985:ISAa

[Gla85a] C. G. Glabe. Interaction of the sperm adhesive protein, bindin, with phospholipid vesicles. I. Specific association of bindin with gel-phase phospholipid vesicles. Journal of Cell Biology, 100 (3):794–??, March 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/100/3/794.

Glabe:1985:ISAb

[Gla85b] C. G. Glabe. Interaction of the sperm adhesive protein, bindin, with phospholipid vesicles. II. Bindin induces the fusion of mixed-phase vesicles that contain phosphatidylcholine and phosphatidylserine in vitro. Journal of Cell Biology, 100(3):800–??, March 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
100/3/800.

Gu:1988:GAD

[GLC88] W. Gu, S. A. Lewis, and N. J. Cowan. Generation of antisera that discriminate among mammalian alpha-tubulins: introduction of specialized isotypes into cultured cells results in their coassembly without disruption of normal microtubule function. Journal of Cell Biology, 106(6):2011–??, June 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/106/6/2011.

Gruenbaum:1988:DNL

[GLD^{+}88] Y. Gruenbaum, Y. Landesman, B. Drees, J. W. Bare, H. Saumweber, M. R. Paddy, J. W. Sedat, D. E. Smith, B. M. Benton, and P. A. Fisher. Drosophila nuclear lamin precursor Dm0 is translated from either of two developmentally regulated mRNA species apparently encoded by a single gene. Journal of Cell Biology, 106(3):585–??, March 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/106/3/585.

Grupp:1985:ILP

[GLH85] S. A. Grupp, M. A. Lieberman, and J. A. Harmony. Inhibition of lymphocyte proliferation by detergent-solubilized mouse liver membranes. Journal of Cell Biology, 101(2):380–??, August 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/101/2/
380.

Grant:1989:ILC

[GLK^{+}89] D. S. Grant, C. P. Leblond, H. K. Kleinman, S. Inoue, and J. R. Hassell. The incubation of laminin, collagen IV, and heparan sulfate proteoglycan at 35 degrees C yields basement membranelike structures. Journal of Cell Biology, 108(4):1567–??, April 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/108/4/
1567.

Garrigues:1986:MPR

[GLL^{+}86] H. J. Garrigues, M. W. Lark, S. Lara, I. Hellström, K. E. Hellström, and T. N. Wight. The melanoma proteoglycan: restricted expression on microspikes, a specific microdomain of the cell surface. Journal of Cell Biology, 103(5):1699–??, November 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/103/5/1699.

Goldstein:1986:MAP

[GLM86] L. S. Goldstein, R. A. Laymon, and J. R. McIntosh. A microtubule-associated protein in Drosophila melanogaster: identification, characterization, and isolation of coding sequences. Journal of Cell Biology, 102(6):2076–??, June 1986. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/102/6/2076.

Gustafson:1989:HMW

[GLW^{+}89] E. J. Gustafson, H. Lukasiewicz, Y. T. Wachtfogel, K. J. Norton, A. H. Schmaier, S. Niewiarowski, and R. W. Colman. High molecular weight kininogen inhibits fibrinogen binding to cytoadhesins of neutrophils and platelets. Journal of Cell Biology, 109(1):377–??, July 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/1/377.

Georgatos:1985:BVH

[GM85a] S. D. Georgatos and V. T. Marchesi. The binding of vimentin to human erythrocyte membranes: a model system for the study of intermediate filament-membrane interactions. Journal of Cell Biology, 100(6):1955–??, June 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/100/6/1955.

Gottlieb:1985:AMB

[GM85b] R. A. Gottlieb and D. B. Murphy. Analysis of the microtubulebinding domain of MAP-2. Journal of Cell Biology, 101(5): 1782–??, November 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/101/5/1782.

Greenwalt:1985:CAD

[GM85c] D. E. Greenwalt and I. H. Mather. Characterization of an apically derived epithelial membrane glycoprotein from bovine milk, which is expressed in capillary endothelia in diverse tissues. Journal of Cell Biology, 100(2):397–??, February 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/100/2/397.

Goodall:1986:MLJ

[GM86] H. Goodall and B. Maro. Major loss of junctional coupling during mitosis in early mouse embryos. Journal of Cell Biology, 102 (2):568–??, February 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/102/2/568.

Garner:1988:DFM

[GM88] C. C. Garner and A. Matus. Different forms of microtubuleassociated protein 2 are encoded by separate mRNA transcripts. Journal of Cell Biology, 106(3):779–??, March 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/106/3/779.

Georgatos:1989:LLB

[GMB89] S. D. Georgatos, I. Maroulakou, and G. Blobel. Lamin A, lamin B, and lamin B receptor analogues in yeast. Journal of Cell Biology, 108(6):2069–??, June 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/108/6/2069.

Godfrey:1987:SSI

[GMCH87] R. W. Godfrey, R. M. Manzi, M. A. Clark, and S. T. Hoffstein. Stimulus-specific induction of phospholipid and arachidonic acid metabolism in human neutrophils. Journal of Cell Biology, 104 (4):925–??, January 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/104/4/925.

Greenspan:1985:NRS

[GMF85] P. Greenspan, E. P. Mayer, and S. D. Fowler. Nile red: a selective fluorescent stain for intracellular lipid droplets. Journal of Cell Biology, 100(3):965–??, March 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/100/3/965.

Gorvel:1989:CCR

[GML^{+}89] J. P. Gorvel, Z. Mishal, F. Liegey, A. Rigal, and S. Maroux. Conformational change of rabbit aminopeptidase N into enterocyte plasma membrane domains analyzed by flow cytometry fluorescence energy transfer. Journal of Cell Biology, 108(6):2193–??, June 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/108/6/
2193.

Greer:1985:ATA

[GMLR85] K. Greer, H. Maruta, S. W. L’Hernault, and J. L. Rosenbaum. Alpha-tubulin acetylase activity in isolated Chlamydomonas flagella. Journal of Cell Biology, 101(6):2081–??, December 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/101/6/2081.

Goodenough:1989:CAE

[Goo89] U. W. Goodenough. Cyclic AMP enhances the sexual agglutinability of Chlamydomonas flagella. Journal of Cell Biology, 109(1):247–??, July 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/1/247.

Gordon:1985:OAN

[Gor85] R. E. Gordon. Orthogonal arrays in normal and injured respiratory airway epithelium. Journal of Cell Biology, 100(2): 648–??, February 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/100/2/648.

Gordon:1989:IED

[Gor89] J. I. Gordon. Intestinal epithelial differentiation: new insights from chimeric and transgenic mice. Journal of Cell Biology, 108 (4):1187–??, April 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/108/4/1187.

Garcia:1988:THB

[GORW88] P. D. Garcia, J. H. Ou, W. J. Rutter, and P. Walter. Targeting of the hepatitis B virus precore protein to the endoplasmic reticulum membrane: after signal peptide cleavage translocation can be aborted and the product released into the cytoplasm. Journal of Cell Biology, 106(4):1093–??, April 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/106/4/1093.

Greenberger:1986:MGN

[GP86] L. M. Greenberger and K. H. Pfenninger. Membrane glycoproteins of the nerve growth cone: diversity and growth regulation of oligosaccharides. Journal of Cell Biology, 103(4):1369–??, October 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/103/4/
1369.

Goodenough:1988:TDT

[GPJ88] D. A. Goodenough, D. L. Paul, and L. Jesaitis. Topological distribution of two connexin32 antigenic sites in intact and split rodent hepatocyte gap junctions. Journal of Cell Biology, 107 (5):1817–??, November 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/107/5/1817.

Griffiths:1986:VBC

[GPSC86] S. L. Griffiths, R. M. Perkins, C. H. Streuli, and D. R. Critchley. Variants of BALB/c 3T3 cells lacking complex gangliosides retain a fibronectin matrix and spread normally on fibronectincoated substrates. Journal of Cell Biology, 102(2):469–??, February 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/102/2/
469.

Griffiths:1985:ENS

[GPSM85] G. Griffiths, S. Pfeiffer, K. Simons, and K. Matlin. Exit of newly synthesized membrane proteins from the trans cisterna of the Golgi complex to the plasma membrane. Journal of Cell Biology, 101(3):949–??, September 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/101/3/949.

Geoffroy:1989:DLL

[GR89] J. S. Geoffroy and S. D. Rosen. Demonstration that a lectin-like receptor (gp90MEL) directly mediates adhesion of lymphocytes to high endothelial venules of lymph nodes. Journal of Cell Biology, 109(5):2463–??, November 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/109/5/2463.

Green:1986:PCM

[GRCG86] S. H. Green, R. E. Rydel, J. L. Connolly, and L. A. Greene. PC12 cell mutants that possess low- but not high-affinity nerve growth factor receptors neither respond to nor internalize nerve growth factor. Journal of Cell Biology, 102(3):830–??, March 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/102/3/
830.

Grinnell:1986:FAS

[Gri86] F. Grinnell. Focal adhesion sites and the removal of substratumbound fibronectin. Journal of Cell Biology, 103(6):2697–??, December 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
103/6/2697.

Gu:1989:ARC

[GRME^{+}89] Y. Gu, E. Ralston, C. Murphy-Erdosh, R. A. Black, and Z. W. Hall. Acetylcholine receptor in a C2 muscle cell variant is retained in the endoplasmic reticulum. Journal of Cell Biology, 109(2):729–??, August 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/2/729.

Goodman:1989:ESL

[GRvdM89] S. L. Goodman, G. Risse, and K. von der Mark. The E8 subfragment of laminin promotes locomotion of myoblasts over extracellular matrix. Journal of Cell Biology, 109(2):799–??, August 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/109/2/
799.

Ghinea:1985:ACH

[GS85] N. Ghinea and N. Simionescu. Anionized and cationized hemeundecapeptides as probes for cell surface charge and permeability studies: differentiated labeling of endothelial plasmalemmal vesicles. Journal of Cell Biology, 100(2):606–??, February 1985. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/100/2/606.

Gilbert:1986:IML

[GS86a] S. P. Gilbert and R. D. Sloboda. Identification of a MAP 2-like ATP-binding protein associated with axoplasmic vesicles that translocate on isolated microtubules. Journal of Cell Biology, 103 (3):947–??, September 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/103/3/947.

Gumbiner:1986:FAP

[GS86b] B. Gumbiner and K. Simons. A functional assay for proteins involved in establishing an epithelial occluding barrier: identification of a uvomorulin-like polypeptide. Journal of Cell Biology, 102(2):457–??, February 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/102/2/457.

Gilbert:1989:SDI

[GS89] S. P. Gilbert and R. D. Sloboda. A squid dynein isoform promotes axoplasmic vesicle translocation. Journal of Cell Biology, 109(5):

2379–??, November 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/5/2379.

Gonzalez-Sanchez:1985:CMH

[GSB85] A. González-Sánchez and D. Bader. Characterization of a myosin heavy chain in the conductive system of the adult and developing chicken heart. Journal of Cell Biology, 100(1):270–??, January 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/100/1/
270.

Gorbsky:1987:CMP

[GSB87] G. J. Gorbsky, P. J. Sammak, and G. G. Borisy. Chromosomes move poleward in anaphase along stationary microtubules that coordinately disassemble from their kinetochore ends. Journal of Cell Biology, 104(1):9–??, January 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/104/1/9.

Gorbsky:1988:MDC

[GSB88] G. J. Gorbsky, P. J. Sammak, and G. G. Borisy. Microtubule dynamics and chromosome motion visualized in living anaphase cells. Journal of Cell Biology, 106(4):1185–??, April 1988. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/106/4/1185.

Garcia:1989:IEG

[GSDR89] J. V. Garcia, M. P. Stoppelli, S. J. Decker, and M. R. Rosner. An insulin epidermal growth factor-binding protein from Drosophila has insulin-degrading activity. Journal of Cell Biology, 108(1):177–??, January 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/108/1/177.

Gorr:1988:NMG

[GSF^{+}88] S. U. Gorr, B. Stieger, J. A. Fransen, M. Kedinger, A. Marxer, and H. P. Hauri. A novel marker glycoprotein for the microvillus membrane of surface colonocytes of rat large intestine and its presence in small-intestinal crypt cells. Journal of Cell Biology, 106(6):1937–??, June 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/106/6/1937.

Gumbiner:1988:RCA

[GSG88] B. Gumbiner, B. Stevenson, and A. Grimaldi. The role of the cell adhesion molecule uvomorulin in the formation and maintenance of the epithelial junctional complex. Journal of Cell Biology, 107 (4):1575–??, October 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/107/4/1575.

Geuze:1985:PPL

[GSS^{+}85] H. J. Geuze, J. W. Slot, G. J. Strous, A. Hasilik, and K. von Figura. Possible pathways for lysosomal enzyme delivery. Journal of Cell Biology, 101(6):2253–??, December 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/101/6/2253.

Geuze:1987:MSO

[GSS87] H. J. Geuze, J. W. Slot, and A. L. Schwartz. Membranes of sorting organelles display lateral heterogeneity in receptor distribution. Journal of Cell Biology, 104(6):1715–??, June 1987. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/104/6/1715.

Geuze:1988:SMP

[GSS^{+}88] H. J. Geuze, W. Stoorvogel, G. J. Strous, J. W. Slot, J. E. Bleekemolen, and I. Mellman. Sorting of mannose 6-phosphate receptors and lysosomal membrane proteins in endocytic vesicles. Journal of Cell Biology, 107(6):2491–??, December 1988. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/107/6/2491.

Gatchalian:1989:FPN

[GSS89] C. L. Gatchalian, M. Schachner, and J. R. Sanes. Fibroblasts that proliferate near denervated synaptic sites in skeletal muscle synthesize the adhesive molecules tenascin(J1), N-CAM, fibronectin, and a heparan sulfate proteoglycan. Journal of Cell Biology, 108(5):1873–??, May 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/108/5/1873.

Garcia:1987:CDP

[GST^{+}87] J. V. Garcia, M. P. Stoppelli, K. L. Thompson, S. J. Decker, and M. R. Rosner. Characterization of a Drosophila protein that binds both epidermal growth factor and insulin-related growth factors. Journal of Cell Biology, 105(1):449–??, July 1987. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/105/1/449.

Goetinck:1987:TRS

[GSTC87] P. F. Goetinck, N. S. Stirpe, P. A. Tsonis, and D. Carlone. The tandemly repeated sequences of cartilage link protein contain the sites for interaction with hyaluronic acid. Journal of Cell Biology, 105(5):2403–??, November 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/105/5/2403.

Gorza:1986:MTF

[GSTS86] L. Gorza, S. Sartore, L. E. Thornell, and S. Schiaffino. Myosin types and fiber types in cardiac muscle. III. Nodal conduction tissue. Journal of Cell Biology, 102(5):1758–??, May 1986. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/102/5/1758.

Gingell:1985:TCG

[GTB85] D. Gingell, I. Todd, and J. Bailey. Topography of cell-glass apposition revealed by total internal reflection fluorescence of volume markers. Journal of Cell Biology, 100(4):1334–??, April 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/100/4/
1334.

Gibbons:1985:OAS

[GTG85] B. H. Gibbons, W. J. Tang, and I. R. Gibbons. Organic anions stabilize the reactivated motility of sperm flagella and the latency of dynein 1 ATPase activity. Journal of Cell Biology, 101 (4):1281–??, October 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/101/4/1281.

Goldmacher:1986:EPL

[GTN86] V. S. Goldmacher, N. L. Tinnel, and B. C. Nelson. Evidence that pinocytosis in lymphoid cells has a low capacity. Journal of Cell Biology, 102(4):1312–??, April 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/102/4/1312.

Glenney:1987:CTD

[GTP87] J. R. Glenney, B. Tack, and M. A. Powell. Calpactins: two distinct Ca++-regulated phospholipid- and actin-binding proteins isolated from lung and placenta. Journal of Cell Biology, 104(3):503–??, March 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/104/3/503.

Galvin:1987:IHT

[GVD^{+}87] N. J. Galvin, P. M. Vance, V. M. Dixit, B. Fink, and W. A. Frazier. Interaction of human thrombospondin with types I–V collagen: direct binding and electron microscopy. Journal of Cell Biology, 104(5):1413–??, May 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/104/5/1413.

Gown:1985:SMS

[GVGL85] A. M. Gown, A. M. Vogel, D. Gordon, and P. L. Lu. A smooth muscle-specific monoclonal antibody recognizes smooth muscle actin isozymes. Journal of Cell Biology, 100(3):807–??, March 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/100/3/
807.

Geiger:1985:MHA

[GVV85] B. Geiger, T. Volk, and T. Volberg. Molecular heterogeneity of adherens junctions. Journal of Cell Biology, 101(4):1523–??, October 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
101/4/1523.

Gottmann:1986:SME

[GW86] K. Gottmann and C. J. Weijer. In situ measurements of external pH and optical density oscillations in Dictyostelium discoideum aggregates. Journal of Cell Biology, 102(5):1623–??, May 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/102/5/
1623.

Grabel:1987:REM

[GW87] L. B. Grabel and T. D. Watts. The role of extracellular matrix in the migration and differentiation of parietal endoderm from teratocarcinoma embryoid bodies. Journal of Cell Biology, 105(1):441–??, July 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/1/441.

Garcia:1988:FLP

[GW88a] P. D. Garcia and P. Walter. Full-length prepro-alpha-factor can be translocated across the mammalian microsomal membrane only if translation has not terminated. Journal of Cell Biology, 106(4):1043–??, April 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/106/4/1043.

Guthrie:1988:YVF

[GW88b] B. A. Guthrie and W. Wickner. Yeast vacuoles fragment when microtubules are disrupted. Journal of Cell Biology, 107(1):115– ??, July 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
107/1/115.

Grainger:1989:SUM

[GW89] J. L. Grainger and M. M. Winkler. The sea urchin multicatalytic protease: purification, biochemical analysis, subcellular distribution, and relationship to snRNPs. Journal of Cell Biology, 109 (2):675–??, August 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/2/675.

Ghosh:1985:ERC

[GWCW85] P. C. Ghosh, R. B. Wellner, E. J. Cragoe, and H. C. Wu. Enhancement of ricin cytotoxicity in Chinese hamster ovary cells by depletion of intracellular K+: evidence for an Na+/H+ exchange system in Chinese hamster ovary cells. Journal of Cell Biology, 101(2):350–??, August 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/101/2/350.

Gal:1985:PLL

[GWG85] S. Gal, M. C. Willingham, and M. M. Gottesman. Processing and lysosomal localization of a glycoprotein whose secretion is transformation stimulated. Journal of Cell Biology, 100 (2):535–??, February 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/100/2/535.

Grab:1987:SLV

[GWI^{+}87] D. J. Grab, P. Webster, S. Ito, W. R. Fish, Y. Verjee, and J. D. Lonsdale-Eccles. Subcellular localization of a variable surface glycoprotein phosphatidylinositol-specific phospholipase–C in African trypanosomes. Journal of Cell Biology, 105(2):737–??, August 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
105/2/737.

Georgatos:1985:SSV

[GWM85] S. D. Georgatos, D. C. Weaver, and V. T. Marchesi. Site specificity in vimentin-membrane interactions: intermediate filament subunits associate with the plasma membrane via their head domains. Journal of Cell Biology, 100(6):1962–??, June 1985. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/100/6/1962.

Giulian:1986:BPGb

[GY86] D. Giulian and D. G. Young. Brain peptides and glial growth. II. Identification of cells that secrete glia-promoting factors. Journal of Cell Biology, 102(3):812–??, March 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/102/3/812.

Glenney:1989:NTK

[GZ89] J. R. Glenney and L. Zokas. Novel tyrosine kinase substrates from Rous sarcoma virus-transformed cells are present in the membrane skeleton. Journal of Cell Biology, 108(6):2401–??, June 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/108/6/
2401.

Green:1987:KIT

[GZGM87] S. A. Green, K. P. Zimmer, G. Griffiths, and I. Mellman. Kinetics of intracellular transport and sorting of lysosomal membrane and plasma membrane proteins. Journal of Cell Biology, 105(3): 1227–??, September 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/3/1227.

Giebelhaus:1987:CEA

[GZHM87] D. H. Giebelhaus, B. D. Zelus, S. K. Henchman, and R. T. Moon. Changes in the expression of alpha-fodrin during embryonic development of Xenopus laevis. Journal of Cell Biology, 105(2):843–??, August 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/2/843.

Gossett:1988:DDI

[GZO88] L. A. Gossett, W. Zhang, and E. N. Olson. Dexamethasonedependent inhibition of differentiation of C2 myoblasts bearing steroid-inducible N-ras oncogenes. Journal of Cell Biology, 106 (6):2127–??, June 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/106/6/2127.

Heuser:1989:HMI

[HA89] J. E. Heuser and R. G. Anderson. Hypertonic media inhibit receptor-mediated endocytosis by blocking clathrin-coated pit formation. Journal of Cell Biology, 108(2):389–??, February 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/108/2/
389.

Haffar:1988:GRL

[HAB^{+}88] O. K. Haffar, G. W. Aponte, D. A. Bravo, N. J. John, R. T. Hess, and G. L. Firestone. Glucocorticoid-regulated localization of cell surface glycoproteins in rat hepatoma cells is mediated within the Golgi complex. Journal of Cell Biology, 106(5):1463–??, May 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/106/5/
1463.

Huber:1985:MCN

[HABM85] G. Huber, D. Alaimo-Beuret, and A. Matus. MAP3: characterization of a novel microtubule-associated protein. Journal of Cell Biology, 100(2):496–??, February 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/100/2/496.

Humphries:1986:IAS

[HAK^{+}86] M. J. Humphries, S. K. Akiyama, A. Komoriya, K. Olden, and K. M. Yamada. Identification of an alternatively spliced site in human plasma fibronectin that mediates cell type-specific adhesion. Journal of Cell Biology, 103(6):2637–??, December 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/103/6/2637.

Humphries:1988:NEC

[HAK^{+}88] M. J. Humphries, S. K. Akiyama, A. Komoriya, K. Olden, and K. M. Yamada. Neurite extension of chicken peripheral nervous system neurons on fibronectin: relative importance of specific adhesion sites in the central cell-binding domain and the alternatively spliced type III connecting segment. Journal of Cell Biology, 106(4):1289–??, April 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/106/4/1289.

Hortsch:1986:CSP

[HAM86] M. Hortsch, D. Avossa, and D. I. Meyer. Characterization of secretory protein translocation: ribosome-membrane interaction in endoplasmic reticulum. Journal of Cell Biology, 103(1):241–??, July 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/103/1/
241.

Hashimoto:1988:KPP

[Has88] S. Hashimoto. K-252a, a potent protein kinase inhibitor, blocks nerve growth factor-induced neurite outgrowth and changes in the phosphorylation of proteins in PC12h cells. Journal of Cell Biology, 107(4):1531–??, October 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/107/4/1531.

Hatten:1985:NRA

[Hat85] M. E. Hatten. Neuronal regulation of astroglial morphology and proliferation in vitro. Journal of Cell Biology, 100(2):384–??, February 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
100/2/384.

Hatten:1987:NIA

[Hat87] M. E. Hatten. Neuronal inhibition of astroglial cell proliferation is membrane mediated. Journal of Cell Biology, 104(5):1353–??, May 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/104/5/
1353.

Hwo:1986:IIC

[HB86] S. Hwo and J. Bryan. Immuno-identification of Ca^{2+}-induced conformational changes in human gelsolin and brevin. Journal of Cell Biology, 102(1):227–??, January 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/102/1/227.

Hollenbeck:1987:RTO

[HB87] P. J. Hollenbeck and D. Bray. Rapidly transported organelles containing membrane and cytoskeletal components: their relation to axonal growth. Journal of Cell Biology, 105(6): 2827–??, December 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/6/2827.

Hebert:1988:LEP

[HB88a] C. A. Hébert and J. B. Baker. Linkage of extracellular plasminogen activator to the fibroblast cytoskeleton: colocalization of cell surface urokinase with vinculin. Journal of Cell Biology, 106(4):1241–??, April 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/106/4/1241.

Higgins:1988:CAD

[HB88b] L. S. Higgins and D. K. Berg. Cyclic AMP-dependent mechanism regulates acetylcholine receptor function on bovine adrenal chromaffin cells and discriminates between new and old receptors. Journal of Cell Biology, 107(3):1157–??, September 1988. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/107/3/1157.

Higgins:1988:MSA

[HB88c] L. S. Higgins and D. K. Berg. Metabolic stability and antigenic modulation of nicotinic acetylcholine receptors on bovine adrenal chromaffin cells. Journal of Cell Biology, 107(3):1147– ??, September 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/107/3/1147.

Hubbard:1985:IRH

[HBB85] A. L. Hubbard, J. R. Bartles, and L. T. Braiterman. Identification of rat hepatocyte plasma membrane proteins using monoclonal antibodies. Journal of Cell Biology, 100(4):1115–??, April 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/100/4/

Huarte:1987:PAM

[HBB^{+}87] J. Huarte, D. Belin, D. Bosco, A. P. Sappino, and J. D. Vassalli. Plasminogen activator and mouse spermatozoa: urokinase synthesis in the male genital tract and binding of the enzyme to the sperm cell surface. Journal of Cell Biology, 104(5):1281–??, May 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/104/5/
1281.

Henson:1989:CLP

[HBB^{+}89] J. H. Henson, D. A. Begg, S. M. Beaulieu, D. J. Fishkind, E. M. Bonder, M. Terasaki, D. Lebeche, and B. Kaminer. A calsequestrin-like protein in the endoplasmic reticulum of the sea urchin: localization and dynamics in the egg and first cell cycle embryo. Journal of Cell Biology, 109(1):149–??, July 1989. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/109/1/149.

Hamati:1989:IPS

[HBC89] H. F. Hamati, E. L. Britton, and D. J. Carey. Inhibition of proteoglycan synthesis alters extracellular matrix deposition, proliferation, and cytoskeletal organization of rat aortic smooth muscle cells in culture. Journal of Cell Biology, 108(6):2495–??, June 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/108/6/
2495.

Hedin:1988:DEF

[HBF^{+}88] U. Hedin, B. A. Bottger, E. Forsberg, S. Johansson, and J. Thyberg. Diverse effects of fibronectin and laminin on phenotypic properties of cultured arterial smooth muscle cells. Journal of Cell Biology, 107(1):307–??, July 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/107/1/307.

Hurtley:1989:IMI

[HBHL^{+}89] S. M. Hurtley, D. G. Bole, H. Hoover-Litty, A. Helenius, and C. S. Copeland. Interactions of misfolded influenza virus hemagglutinin with binding protein (BiP). Journal of Cell Biology, 108(6):2117–??, June 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/108/6/2117.

Hendershot:1987:ASH

[HBKK87] L. Hendershot, D. Bole, G. Köhler, and J. F. Kearney. Assembly and secretion of heavy chains that do not associate posttranslationally with immunoglobulin heavy chain-binding protein. Journal of Cell Biology, 104(3):761–??, March 1987. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/104/3/761.

Hashimoto:1988:ICL

[HBL^{+}88] S. Hashimoto, B. Bruno, D. P. Lew, T. Pozzan, P. Volpe, and J. Meldolesi. Immunocytochemistry of calciosomes in liver and pancreas. Journal of Cell Biology, 107(6):2523–??, December 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/107/6/
2523.

Hammer:1987:CNS

[HBPK87] J. A. Hammer, B. Bowers, B. M. Paterson, and E. D. Korn. Complete nucleotide sequence and deduced polypeptide sequence of a nonmuscle myosin heavy chain gene from Acanthamoeba: evidence of a hinge in the rodlike tail. Journal of Cell Biology, 105(2):913–??, August 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/2/913.

Hadley:1985:EMR

[HBSQ^{+}85] M. A. Hadley, S. W. Byers, C. A. Suárez-Quian, H. K. Kleinman, and M. Dym. Extracellular matrix regulates Sertoli cell differentiation, testicular cord formation, and germ cell development in vitro. Journal of Cell Biology, 101(4):1511–??, October 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/101/4/
1511.

Hirokawa:1985:CAIa

[HBV85] N. Hirokawa, G. S. Bloom, and R. B. Vallee. Cytoskeletal architecture and immunocytochemical localization of microtubuleassociated proteins in regions of axons associated with rapid axonal transport: the beta,beta’-iminodipropionitrile-intoxicated axon as a model system. Journal of Cell Biology, 101(1):227–??, July 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/101/1/

Herman:1987:HVE

[HBW^{+}87] I. M. Herman, A. M. Brant, V. S. Warty, J. Bonaccorso, E. C. Klein, R. L. Kormos, and H. S. Borovetz. Hemodynamics and the vascular endothelial cytoskeleton. Journal of Cell Biology, 105(1):291–??, July 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/1/291.

Hollenbeck:1986:NMA

[HC86] P. J. Hollenbeck and K. Chapman. A novel microtubuleassociated protein from mammalian nerve shows ATP-sensitive binding to microtubules. Journal of Cell Biology, 103(4):1539–??, October 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
103/4/1539.

Hepler:1987:FCI

[HC87] P. K. Hepler and D. A. Callaham. Free calcium increases during anaphase in stamen hair cells of Tradescantia. Journal of Cell Biology, 105(5):2137–??, November 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/105/5/2137.

Hoffman:1988:MFB

[HCE88] S. Hoffman, K. L. Crossin, and G. M. Edelman. Molecular forms, binding functions, and developmental expression patterns of cytotactin and cytotactin-binding proteoglycan, an interactive pair of extracellular matrix molecules. Journal of Cell Biology, 106 (2):519–??, February 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/106/2/519.

Howell:1987:ESB

[HCG87] T. W. Howell, S. Cockcroft, and B. D. Gomperts. Essential synergy between Ca^{2+} and guanine nucleotides in exocytotic secretion from permeabilized rat mast cells. Journal of Cell Biology, 105(1):191–??, July 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/1/191.

Hoppe:1985:TTP

[HCH85] C. A. Hoppe, T. P. Connolly, and A. L. Hubbard. Transcellular transport of polymeric IgA in the rat hepatocyte: biochemical and morphological characterization of the transport pathway. Journal of Cell Biology, 101(6):2113–??, December 1985. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/101/6/2113.

Hartwig:1989:APF

[HCHK89] J. H. Hartwig, K. A. Chambers, K. L. Hopcia, and D. J. Kwiatkowski. Association of profilin with filament-free regions of human leukocyte and platelet membranes and reversible membrane binding during platelet activation. Journal of Cell Biology, 109(4):1571–??, October 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/109/4/1571.

Honer:1988:MCM

[HCKJJ88] B. Höner, S. Citi, J. Kendrick-Jones, and B. M. Jockusch. Modulation of cellular morphology and locomotory activity by antibodies against myosin. Journal of Cell Biology, 107(6):2181–??, December 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
107/6/2181.

Hartwig:1989:AGA

[HCS89] J. H. Hartwig, K. A. Chambers, and T. P. Stossel. Association of gelsolin with actin filaments and cell membranes of macrophages and platelets. Journal of Cell Biology, 108(2):467–??, February 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/108/2/
467.

Herman:1985:MPC

[HD85] I. M. Herman and P. A. D’Amore. Microvascular pericytes contain muscle and nonmuscle actins. Journal of Cell Biology, 101(1):43–??, July 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/101/1/43.

Hannink:1986:CSE

[HD86] M. Hannink and D. J. Donoghue. Cell surface expression of membrane-anchored v-sis gene products: glycosylation is not required for cell surface transport. Journal of Cell Biology, 103(6): 2311–??, December 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/103/6/2311.

Haas:1988:MSM

[HD88a] C. A. Haas and L. J. DeGennaro. Multiple synapsin I messenger RNAs are differentially regulated during neuronal development. Journal of Cell Biology, 106(1):195–??, January 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/106/1/195.

Hannink:1988:ASV

[HD88b] M. Hannink and D. J. Donoghue. Autocrine stimulation by the vsis gene product requires a ligand-receptor interaction at the cell surface. Journal of Cell Biology, 107(1):287–??, July 1988. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/107/1/287.

Haldar:1989:TFP

[HdAC89] K. Haldar, A. F. de Amorim, and G. A. Cross. Transport of fluorescent phospholipid analogues from the erythrocyte membrane to the parasite in Plasmodium falciparum-infected cells. Journal of Cell Biology, 108(6):2183–??, June 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/108/6/2183.

Haffar:1988:TAH

[HDB88] O. K. Haffar, D. J. Dowbenko, and P. W. Berman. Topogenic analysis of the human immunodeficiency virus type 1 envelope glycoprotein, gp160, in microsomal membranes. Journal of Cell Biology, 107(5):1677–??, November 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/107/5/1677.

Horwitz:1985:CSA

[HDG^{+}85] A. Horwitz, K. Duggan, R. Greggs, C. Decker, and C. Buck. The cell substrate attachment (CSAT) antigen has properties of a receptor for laminin and fibronectin. Journal of Cell Biology, 101 (6):2134–??, December 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/101/6/2134.

Hill:1986:TMD

[HDL^{+}86] C. S. Hill, S. Duran, Z. X. Lin, K. Weber, and H. Holtzer. Titin and myosin, but not desmin, are linked during myofibrillogenesis in postmitotic mononucleated myoblasts. Journal of Cell Biology, 103(6):2185–??, December 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/103/6/2185.

Hammel:1985:DVD

[HDP^{+}85] I. Hammel, A. M. Dvorak, S. P. Peters, E. S. Schulman, H. F. Dvorak, L. M. Lichtenstein, and S. J. Galli. Differences in the volume distributions of human lung mast cell granules and lipid bodies: evidence that the size of these organelles is regulated by distinct mechanisms. Journal of Cell Biology, 100(5):1488–??, May 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/100/5/
1488.

Heck:1986:TIS

[HE86] M. M. Heck and W. C. Earnshaw. Topoisomerase II: a specific marker for cell proliferation. Journal of Cell Biology, 103(6): 2569–??, December 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/103/6/2569.

Hepler:1985:CRP

[Hep85] P. K. Hepler. Calcium restriction prolongs metaphase in dividing Tradescantia stamen hair cells. Journal of Cell Biology, 100(5):1363–??, May 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/100/5/1363.

Hepler:1989:CTD

[Hep89] P. K. Hepler. Calcium transients during mitosis: observations in flux. Journal of Cell Biology, 109(6):2567–??, December 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/109/6/2567.

Herrera:1987:TDE

[HERO^{+}87] V. L. Herrera, J. R. Emanuel, N. Ruiz-Opazo, R. Levenson, and B. Nadal-Ginard. Three differentially expressed Na,K–ATPase alpha subunit isoforms: structural and functional implications. Journal of Cell Biology, 105(4):1855–??, October 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/105/4/1855.

Heuser:1986:DSS

[Heu86] J. E. Heuser. Different structural states of a microtubule crosslinking molecule, captured by quick-freezing motile axostyles in protozoa. Journal of Cell Biology, 103(6):2209–??, December 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/103/6/
2209.

Heuser:1989:CLS

[Heu89a] J. Heuser. Changes in lysosome shape and distribution correlated with changes in cytoplasmic pH. Journal of Cell Biology, 108 (3):855–??, March 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/108/3/855.

Heuser:1989:ECA

[Heu89b] J. Heuser. Effects of cytoplasmic acidification on clathrin lattice morphology. Journal of Cell Biology, 108(2):401–??, February 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/108/2/
401.

Hatzfeld:1985:PFP

[HF85a] M. Hatzfeld and W. W. Franke. Pair formation and promiscuity of cytokeratins: formation in vitro of heterotypic complexes and intermediate-sized filaments by homologous and heterologous recombinations of purified polypeptides. Journal of Cell Biology, 101(5):1826–??, November 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/101/5/1826.

Higashi-Fujime:1985:USM

[HF85b] S. Higashi-Fujime. Unidirectional sliding of myosin filaments along the bundle of F-actin filaments spontaneously formed during superprecipitation. Journal of Cell Biology, 101(6): 2335–??, December 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/101/6/2335.

Horst:1987:CMI

[HFB87] C. J. Horst, D. M. Forestner, and J. C. Besharse. Cytoskeletalmembrane interactions: a stable interaction between cell surface glycoconjugates and doublet microtubules of the photoreceptor connecting cilium. Journal of Cell Biology, 105(6):2973–??, December 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
105/6/2973.

Hudson:1989:IEM

[HFB^{+}89] L. D. Hudson, V. L. Friedrich, T. Behar, M. Dubois-Dalcq, and R. A. Lazzarini. The initial events in myelin synthesis: orientation of proteolipid protein in the plasma membrane of cultured oligodendrocytes. Journal of Cell Biology, 109(2):717–??, August 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/109/2/
717.

Hoffman:1986:DCN

[HFC^{+}86] S. Hoffman, D. R. Friedlander, C. M. Chuong, M. Grumet, and G. M. Edelman. Differential contributions of Ng–CAM and N-CAM to cell adhesion in different neural regions. Journal of Cell Biology, 103(1):145–??, July 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/103/1/145.

Horwich:1985:EAD

[HFF^{+}85] A. L. Horwich, W. A. Fenton, F. A. Firgaira, J. E. Fox, D. Kolansky, I. S. Mellman, and L. E. Rosenberg. Expression of amplified DNA sequences for ornithine transcarbamylase in HeLa cells: arginine residues may be required for mitochondrial import of enzyme precursor. Journal of Cell Biology, 100(5):1515–??, May 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/100/5/
1515.

He:1987:BMA

[HFG87] H. T. He, J. Finne, and C. Goridis. Biosynthesis, membrane association, and release of N-CAM–120, a phosphatidylinositol-linked form of the neural cell adhesion molecule. Journal of Cell Biology, 105(6):2489–??, December 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/105/6/2489.

Haeuptle:1989:TSS

[HFGD89] M. T. Haeuptle, N. Flint, N. M. Gough, and B. Dobberstein. A tripartite structure of the signals that determine protein insertion into the endoplasmic reticulum membrane. Journal of Cell Biology, 108(4):1227–??, April 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/108/4/1227.

Hall:1988:ICL

[HFHR88] D. E. Hall, K. A. Frazer, B. C. Hann, and L. F. Reichardt. Isolation and characterization of a laminin-binding protein from rat and chick muscle. Journal of Cell Biology, 107(2):687–??, August 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/107/2/
687.

Hennessy:1989:CTM

[HFK^{+}89] S. W. Hennessy, B. A. Frazier, D. D. Kim, T. L. Deckwerth, D. M. Baumgartel, P. Rotwein, and W. A. Frazier. Complete thrombospondin mRNA sequence includes potential regulatory sites in the 3’ untranslated region. Journal of Cell Biology, 108 (2):729–??, February 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/108/2/729.

Hashimoto:1987:STS

[HFZM87] S. Hashimoto, G. Fumagalli, A. Zanini, and J. Meldolesi. Sorting of three secretory proteins to distinct secretory granules in acidophilic cells of cow anterior pituitary. Journal of Cell Biology, 105(4):1579–??, October 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/105/4/1579.

Haverstick:1988:VDF

[HG88] D. M. Haverstick and M. Glaser. Visualization of domain formation in the inner and outer leaflets of a phospholipid bilayer. Journal of Cell Biology, 106(6):1885–??, June 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/106/6/1885.

Harris:1989:STM

[HG89] A. K. Harris and S. L. Gewalt. Simulation testing of mechanisms for inducing the formation of the contractile ring in cytokinesis. Journal of Cell Biology, 109(5):2215–??, November 1989. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/109/5/2215.

Heriot:1985:PTS

[HGL85] K. Heriot, P. Gambetti, and R. J. Lasek. Proteins transported in slow components a and b of axonal transport are distributed differently in the transverse plane of the axon. Journal of Cell Biology, 100(4):1167–??, April 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/100/4/1167.

Hull:1988:ISI

[HGL88] J. D. Hull, R. Gilmore, and R. A. Lamb. Integration of a small integral membrane protein, M2, of influenza virus into the endoplasmic reticulum: analysis of the internal signal-anchor domain of a protein with an ectoplasmic NH2 terminus. Journal of Cell Biology, 106(5):1489–??, May 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/106/5/1489.

Hirai:1988:TIM

[HGMS88] M. Hirai, S. Gamou, S. Minoshima, and N. Shimizu. Two independent mechanisms for escaping epidermal growth factormediated growth inhibition in epidermal growth factor receptorhyperproducing human tumor cells. Journal of Cell Biology, 107 (2):791–??, August 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/107/2/791.

Hearing:1989:ATO

[HGS89] J. Hearing, M. J. Gething, and J. Sambrook. Addition of truncated oligosaccharides to influenza virus hemagglutinin results in its temperature-conditional cell-surface expression. Journal of Cell Biology, 108(2):355–??, February 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/108/2/355.

Honig:1986:FCD

[HH86] M. G. Honig and R. I. Hume. Fluorescent carbocyanine dyes allow living neurons of identified origin to be studied in long-term cultures. Journal of Cell Biology, 103(1):171–??, July 1986. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/103/1/171.

Hirokawa:1987:BUB

[HH87] N. Hirokawa and S. Hisanaga. ”Buttonin,” a unique buttonshaped microtubule-associated protein (75 kD) that decorates spindle microtubule surface hexagonally. Journal of Cell Biology, 104(6):1553–??, June 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/104/6/1553.

Horvat:1986:ECM

[HHD^{+}86] R. Horvat, A. Hovorka, G. Dekan, H. Poczewski, and D. Kerjaschki. Endothelial cell membranes contain podocalyxin–the major sialoprotein of visceral glomerular epithelial cells. Journal of Cell Biology, 102(2):484–??, February 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/102/2/484.

Holcomb:1988:SVE

[HHES88] C. L. Holcomb, W. J. Hansen, T. Etcheverry, and R. Schekman. Secretory vesicles externalize the major plasma membrane AT-Pase in yeast. Journal of Cell Biology, 106(3):641–??, March 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/106/3/
641.

Holland:1987:VEM

[HHMH87] P. W. Holland, S. J. Harper, J. H. McVey, and B. L. Hogan. In vivo expression of mRNA for the Ca++-binding protein SPARC (osteonectin) revealed by in situ hybridization. Journal of Cell Biology, 105(1):473–??, July 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/105/1/473.

Hearing:1989:ICH

[HHR^{+}89] J. Hearing, E. Hunter, L. Rodgers, M. J. Gething, and J. Sambrook. Isolation of Chinese hamster ovary cell lines temperature conditional for the cell-surface expression of integral membrane glycoproteins. Journal of Cell Biology, 108(2):339–??, February 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/108/2/

Hornick:1985:ICM

[HHS^{+}85] C. A. Hornick, R. L. Hamilton, E. Spaziani, G. H. Enders, and R. J. Havel. Isolation and characterization of multivesicular bodies from rat hepatocytes: an organelle distinct from secretory vesicles of the Golgi apparatus. Journal of Cell Biology, 100(5):1558–??, May 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/100/5/1558.

Hugle:1985:LRP

[HHSF85] B. Hügle, R. Hazan, U. Scheer, and W. W. Franke. Localization of ribosomal protein S1 in the granular component of the interphase nucleolus and its distribution during mitosis. Journal of Cell Biology, 100(3):873–??, March 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/100/3/873.

Hirano:1988:TSM

[HHY88] T. Hirano, Y. Hiraoka, and M. Yanagida. A temperaturesensitive mutation of the Schizosaccharomyces pombe gene nuc2+ that encodes a nuclear scaffold-like protein blocks spindle elongation in mitotic anaphase. Journal of Cell Biology, 106(4):1171– ??, April 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
106/4/1171.

Hirokawa:1986:MAP

[Hir86] N. Hirokawa. 270K microtubule-associated protein cross-reacting with anti-MAP2 IgG in the crayfish peripheral nerve axon. Journal of Cell Biology, 103(1):33–??, July 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/103/1/33.

Heidemann:1985:SEC

[HJS^{+}85] S. R. Heidemann, H. C. Joshi, A. Schechter, J. R. Fletcher, and M. Bothwell. Synergistic effects of cyclic AMP and nerve growth factor on neurite outgrowth and microtubule stability of PC12 cells. Journal of Cell Biology, 100(3):916–??, March 1985. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/100/3/916.

Hwang:1989:DYM

[HJV^{+}89] S. Hwang, T. Jascur, D. Vestweber, L. Pon, and G. Schatz. Disrupted yeast mitochondria can import precursor proteins directly through their inner membrane. Journal of Cell Biology, 109(2):487–??, August 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/2/487.

Heuser:1988:DEV

[HK88a] J. E. Heuser and J. Keen. Deep-etch visualization of proteins involved in clathrin assembly. Journal of Cell Biology, 107(3): 877–??, September 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/107/3/877.

Huitorel:1988:PSM

[HK88b] P. Huitorel and M. W. Kirschner. The polarity and stability of microtubule capture by the kinetochore. Journal of Cell Biology, 106(1):151–??, January 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/106/1/151.

Heuser:1987:DEV

[HKA^{+}87] J. E. Heuser, J. H. Keen, L. M. Amende, R. E. Lippoldt, and K. Prasad. Deep-etch visualization of 27S clathrin: a tetrahedral tetramer. Journal of Cell Biology, 105(5):1999–??, November 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/105/5/
1999.

Heumann:1987:CNG

[HKBT87] R. Heumann, S. Korsching, C. Bandtlow, and H. Thoenen. Changes of nerve growth factor synthesis in nonneuronal cells in response to sciatic nerve transection. Journal of Cell Biology, 104(6):1623–??, June 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/104/6/1623.

Horwich:1987:OTL

[HKF^{+}87] A. L. Horwich, F. Kalousek, W. A. Fenton, K. Furtak, R. A. Pollock, and L. E. Rosenberg. The ornithine transcarbamylase leader peptide directs mitochondrial import through both its midportion structure and net positive charge. Journal of Cell Biology, 105(2):669–??, August 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/2/669.

Haverty:1988:CRT

[HKH^{+}88] T. P. Haverty, C. J. Kelly, W. H. Hines, P. S. Amenta, M. Watanabe, R. A. Harper, N. A. Kefalides, and E. G. Neilson. Characterization of a renal tubular epithelial cell line which secretes the autologous target antigen of autoimmune experimental interstitial nephritis. Journal of Cell Biology, 107(4):1359–??, October 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/107/4/
1359.

Hagen:1986:CMA

[HKKP86] S. J. Hagen, D. P. Kiehart, D. A. Kaiser, and T. D. Pollard. Characterization of monoclonal antibodies to Acanthamoeba myosin– I that cross-react with both myosin–II and low molecular mass nuclear proteins. Journal of Cell Biology, 103(6):2121–??, December 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
103/6/2121.

Hanai:1987:PPI

[HKM^{+}87] K. Hanai, H. Kato, S. Matsuhashi, H. Morita, E. W. Raines, and R. Ross. Platelet proteins, including platelet-derived growth factor, specifically depress a subset of the multiple components of the response elicited by glutathione in Hydra. Journal of Cell Biology, 104(6):1675–??, June 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/104/6/1675.

Horovitz:1989:ARA

[HKPS89] O. Horovitz, D. Knaack, T. R. Podleski, and M. M. Salpeter. Acetylcholine receptor alpha-subunit mRNA is increased by ascorbic acid in cloned L5 muscle cells: Northern blot analysis and in situ hybridization. Journal of Cell Biology, 108 (5):1823–??, May 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/108/5/1823.

Hirano:1989:ERR

[HKTY89] T. Hirano, G. Konoha, T. Toda, and M. Yanagida. Essential roles of the RNA polymerase I largest subunit and DNA topoisomerases in the formation of fission yeast nucleolus. Journal of Cell Biology, 108(2):243–??, February 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/108/2/243.

Hiebert:1988:CSE

[HL88] S. W. Hiebert and R. A. Lamb. Cell surface expression of glycosylated, nonglycosylated, and truncated forms of a cytoplasmic protein pyruvate kinase. Journal of Cell Biology, 107(3): 865–??, September 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/107/3/865.

Halaban:1988:BFG

[HLB^{+}88] R. Halaban, R. Langdon, N. Birchall, C. Cuono, A. Baird, G. Scott, G. Moellmann, and J. McGuire. Basic fibroblast growth factor from human keratinocytes is a natural mitogen for melanocytes. Journal of Cell Biology, 107(4):1611–??, October 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/107/4/
1611.

Hegmann:1988:MDH

[HLL88] T. E. Hegmann, J. L. Lin, and J. J. Lin. Motility-dependence of the heterogeneous staining of culture cells by a monoclonal anti-tropomyosin antibody. Journal of Cell Biology, 106(2): 385–??, February 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/106/2/385.

Hegmann:1989:PRN

[HLL89] T. E. Hegmann, J. L. Lin, and J. J. Lin. Probing the role of nonmuscle tropomyosin isoforms in intracellular granule movement by microinjection of monoclonal antibodies. Journal of Cell Biology, 109(3):1141–??, September 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/109/3/1141.

Howard:1987:TMA

[HLU^{+}87] R. J. Howard, J. A. Lyon, S. Uni, A. J. Saul, S. B. Aley, F. Klotz, L. J. Panton, J. A. Sherwood, K. Marsh, and M. Aikawa. Transport of an Mr approximately 300,000 Plasmodium falciparum protein (Pf EMP 2) from the intraerythrocytic asexual parasite to the cytoplasmic face of the host cell membrane. Journal of Cell Biology, 104(5):1269–??, May 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/104/5/1269.

Hagestedt:1989:TPB

[HLW^{+}89] T. Hagestedt, B. Lichtenberg, H. Wille, E. M. Mandelkow, and E. Mandelkow. Tau protein becomes long and stiff upon phosphorylation: correlation between paracrystalline structure and degree of phosphorylation. Journal of Cell Biology, 109(4): 1643–??, October 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/4/1643.

Hinshaw:1989:LLH

[HM89a] J. E. Hinshaw and K. R. Miller. Localization of light-harvesting complex II to the occluded surfaces of photosynthetic membranes. Journal of Cell Biology, 109(4):1725–??, October 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/109/4/1725.

Houliston:1989:PMD

[HM89b] E. Houliston and B. Maro. Posttranslational modification of distinct microtubule subpopulations during cell polarization and differentiation in the mouse preimplantation embryo. Journal of Cell Biology, 108(2):543–??, February 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/108/2/543.

Hunziker:1989:EML

[HM89c] W. Hunziker and I. Mellman. Expression of macrophagelymphocyte Fc receptors in Madin–Darby canine kidney cells: polarity and transcytosis differ for isoforms with or without coated pit localization domains. Journal of Cell Biology, 109 (6):3291–??, December 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/6/3291.

Hardeman:1988:VSC

[HMBV^{+}88] E. C. Hardeman, A. Minty, P. Benton-Vosman, L. Kedes, and H. M. Blau. In vivo system for characterizing clonal variation and tissue-specific gene regulatory factors based on function. Journal of Cell Biology, 106(4):1027–??, April 1988. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/106/4/1027.

Homan:1988:MAD

[HMdN^{+}88] W. Homan, A. Musgrave, H. de Nobel, R. Wagter, D. de Wit, A. Kolk, and H. van den Ende. Monoclonal antibodies directed against the sexual binding site of Chlamydomonas eugametos gametes. Journal of Cell Biology, 107(1):177–??, July 1988. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/107/1/177.

Heine:1987:RTG

[HMF^{+}87] U. Heine, E. F. Munoz, K. C. Flanders, L. R. Ellingsworth, H. Y. Lam, N. L. Thompson, A. B. Roberts, and M. B. Sporn. Role of transforming growth factor-beta in the development of the mouse embryo. Journal of Cell Biology, 105(6):2861–??, December 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/105/6/
2861.

Hochstrasser:1986:SOC

[HMG^{+}86] M. Hochstrasser, D. Mathog, Y. Gruenbaum, H. Saumweber, and J. W. Sedat. Spatial organization of chromosomes in the salivary gland nuclei of Drosophila melanogaster. Journal of Cell Biology, 102(1):112–??, January 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/102/1/112.

Ho:1986:ELC

[HMHWS86] S. C. Ho, S. Malek-Hedayat, J. L. Wang, and M. Schindler. Endogenous lectins from cultured soybean cells: isolation of a protein immunologically cross-reactive with seed soybean agglutinin and analysis of its role in binding of Rhizobium japonicum. Journal of Cell Biology, 103(3):1043–??, September 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/103/3/1043.

Huang:1988:MCC

[HML88] B. Huang, A. Mengersen, and V. D. Lee. Molecular cloning of cDNA for caltractin, a basal body-associated Ca^{2+}-binding protein: homology in its protein sequence with calmodulin and the yeast CDC31 gene product. Journal of Cell Biology, 107 (1):133–??, July 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/107/1/133.

Hosokawa:1987:BMC

[HMN87] Y. Hosokawa and T. Miki-Noumura. Bending motion of Chlamydomonas axonemes after extrusion of central-pair microtubules. Journal of Cell Biology, 105(3):1297–??, September 1987. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/105/3/1297.

Hochman:1985:TSS

[HMP85] J. Hochman, N. Mador, and A. Panet. Tubular structures in S49 mouse lymphoma are regulated through in vivo host-cell interaction and in vitro interferon treatment. Journal of Cell Biology, 100(5):1351–??, May 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/100/5/1351.

Herrin:1986:RGE

[HMP86] D. L. Herrin, A. S. Michaels, and A. L. Paul. Regulation of genes encoding the large subunit of ribulose-1,5-bisphosphate carboxylase and the photosystem II polypeptides D-1 and D-2 during the cell cycle of Chlamydomonas reinhardtii. Journal of Cell Biology, 103(5):1837–??, November 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/103/5/1837.

Horowits:1989:EBC

[HMP89] R. Horowits, K. Maruyama, and R. J. Podolsky. Elastic behavior of connectin filaments during thick filament movement in activated skeletal muscle. Journal of Cell Biology, 109(5): 2169–??, November 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/5/2169.

Hynes:1989:IHR

[HMS^{+}89] R. O. Hynes, E. E. Marcantonio, M. A. Stepp, L. A. Urry, and G. H. Yee. Integrin heterodimer and receptor complexity in avian and mammalian cells. Journal of Cell Biology, 109(1):409–??, July 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/109/1/

Herbst:1988:DEL

[HMTF88] T. J. Herbst, J. B. McCarthy, E. C. Tsilibary, and L. T. Furcht. Differential effects of laminin, intact type IV collagen, and specific domains of type IV collagen on endothelial cell adhesion and migration. Journal of Cell Biology, 106(4):1365–??, April 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/106/4/
1365.

Hirano:1987:CDC

[HNH^{+}87] S. Hirano, A. Nose, K. Hatta, A. Kawakami, and M. Takeichi. Calcium-dependent cell–cell adhesion molecules (cadherins): subclass specificities and possible involvement of actin bundles. Journal of Cell Biology, 105(6):2501–??, December 1987. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/105/6/2501.

Hatta:1988:CEC

[HNNT88] K. Hatta, A. Nose, A. Nagafuchi, and M. Takeichi. Cloning and expression of cDNA encoding a neural calcium-dependent cell adhesion molecule: its identity in the cadherin gene family. Journal of Cell Biology, 106(3):873–??, March 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/106/3/873.

Hayashi:1986:DLM

[HNP^{+}86] M. Hayashi, Y. Ninomiya, J. Parsons, K. Hayashi, B. R. Olsen, and R. L. Trelstad. Differential localization of mRNAs of collagen types I and II in chick fibroblasts, chondrocytes, and corneal cells by in situ hybridization using cDNA probes. Journal of Cell Biology, 102(6):2302–??, June 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/102/6/2302.

Hall:1987:ENR

[HNR87] D. E. Hall, K. M. Neugebauer, and L. F. Reichardt. Embryonic neural retinal cell response to extracellular matrix proteins: developmental changes and effects of the cell substratum attachment antibody (CSAT). Journal of Cell Biology, 104 (3):623–??, March 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/104/3/623.

Howard:1985:KCP

[HO85] T. H. Howard and C. O. Oresajo. The kinetics of chemotactic peptide-induced change in F-actin content, F-actin distribution, and the shape of neutrophils. Journal of Cell Biology, 101(3): 1078–??, September 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/101/3/1078.

Hollenbeck:1989:DAS

[Hol89a] P. J. Hollenbeck. The distribution, abundance and subcellular localization of kinesin. Journal of Cell Biology, 108(6):2335–??, June 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/108/6/
2335.

Hollenbeck:1989:TAA

[Hol89b] P. J. Hollenbeck. The transport and assembly of the axonal cytoskeleton. Journal of Cell Biology, 108(2):223–??, February 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/108/2/
223.

Harnik-Ort:1987:ICC

[HOPM^{+}87] V. Harnik-Ort, K. Prakash, E. Marcantonio, D. R. Colman, M. G. Rosenfeld, M. Adesnik, D. D. Sabatini, and G. Kreibich. Isolation and characterization of cDNA clones for rat ribophorin I: complete coding sequence and in vitro synthesis and insertion of the encoded product into endoplasmic reticulum membranes. Journal of Cell Biology, 104(4):855–??, January 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/104/4/855.

Herman:1985:PDG

[HP85a] B. Herman and W. J. Pledger. Platelet-derived growth factorinduced alterations in vinculin and actin distribution in BALB/ c-3T3 cells. Journal of Cell Biology, 100(4):1031–??, April 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/100/4/1031.

Hyman:1985:IRN

[HP85b] C. Hyman and K. H. Pfenninger. Intracellular regulators of neuronal sprouting: calmodulin-binding proteins of nerve growth cones. Journal of Cell Biology, 101(3):1153–??, September 1985.

CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/101/3/1153.

Hepler:1986:MIB

[HP86] P. K. Hepler and B. A. Palevitz. Metabolic inhibitors block anaphase A in vivo. Journal of Cell Biology, 102(6):1995–??, June 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/102/6/
1995.

Hanspal:1987:SAM

[HP87a] M. Hanspal and J. Palek. Synthesis and assembly of membrane skeletal proteins in mammalian red cell precursors. Journal of Cell Biology, 105(3):1417–??, September 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/105/3/1417.

Horowits:1987:PST

[HP87b] R. Horowits and R. J. Podolsky. The positional stability of thick filaments in activated skeletal muscle depends on sarcomere length: evidence for the role of titin filaments. Journal of Cell Biology, 105(5):2217–??, November 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/105/5/2217.

Houliston:1987:RMP

[HPM87] E. Houliston, S. J. Pickering, and B. Maro. Redistribution of microtubules and pericentriolar material during the development of polarity in mouse blastomeres. Journal of Cell Biology, 104 (5):1299–??, May 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/104/5/1299.

Haire:1988:MAI

[HPO88] R. N. Haire, M. S. Peterson, and J. J. O’Leary. Mitogen activation induces the enhanced synthesis of two heat-shock proteins in human lymphocytes. Journal of Cell Biology, 106(3):883–??, March 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
106/3/883.

Hayman:1985:DCC

[HPR85] E. G. Hayman, M. D. Pierschbacher, and E. Ruoslahti. Detachment of cells from culture substrate by soluble fibronectin peptides. Journal of Cell Biology, 100(6):1948–??, June 1985. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/100/6/1948.

Hann:1989:SCS

[HPW89] B. C. Hann, M. A. Poritz, and P. Walter. Saccharomyces cerevisiae and Schizosaccharomyces pombe contain a homologue to the 54-kD subunit of the signal recognition particle that in S. cerevisiae is essential for growth. Journal of Cell Biology, 109 (6):3223–??, December 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/6/3223.

Hall:1987:VNC

[HR87] A. K. Hall and U. Rutishauser. Visualization of neural cell adhesion molecule by electron microscopy. Journal of Cell Biology, 104(6):1579–??, June 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/104/6/1579.

Hart:1989:PIS

[HRBR89] I. K. Hart, W. D. Richardson, S. R. Bolsover, and M. C. Raff. PDGF and intracellular signaling in the timing of oligodendrocyte differentiation. Journal of Cell Biology, 109(6):3411–??, December 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
109/6/3411.

Hinek:1987:CCM

[HRP87] A. Hinek, A. Reiner, and A. R. Poole. The calcification of cartilage matrix in chondrocyte culture: studies of the C-propeptide of type II collagen (chondrocalcin). Journal of Cell Biology, 104(5):1435–??, May 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/104/5/1435.

Haehnel:1989:LDD

[HRR89] W. Haehnel, R. Ratajczak, and H. Robenek. Lateral distribution and diffusion of plastocyanin in chloroplast thylakoids. Journal of Cell Biology, 108(4):1397–??, April 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/108/4/1397.

Hartzell:1985:SCP

[HS85a] H. C. Hartzell and W. S. Sale. Structure of C protein purified from cardiac muscle. Journal of Cell Biology, 100(1):208–??, January 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/100/1/
208.

Heimark:1985:RMM

[HS85b] R. L. Heimark and S. M. Schwartz. The role of membranemembrane interactions in the regulation of endothelial cell growth. Journal of Cell Biology, 100(6):1934–??, June 1985. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/100/6/1934.

Hartwig:1986:AAF

[HS86a] J. H. Hartwig and P. Shevlin. The architecture of actin filaments and the ultrastructural location of actin-binding protein in the periphery of lung macrophages. Journal of Cell Biology, 103(3): 1007–??, September 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/103/3/1007.

Hayward:1986:SEC

[HS86b] L. J. Hayward and R. J. Schwartz. Sequential expression of chicken actin genes during myogenesis. Journal of Cell Biology, 102(4):1485–??, April 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/102/4/1485.

Hochstrasser:1987:TDOa

[HS87a] M. Hochstrasser and J. W. Sedat. Three-dimensional organization of Drosophila melanogaster interphase nuclei. I. Tissuespecific aspects of polytene nuclear architecture. Journal of Cell Biology, 104(6):1455–??, June 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/104/6/1455.

Hochstrasser:1987:TDOb

[HS87b] M. Hochstrasser and J. W. Sedat. Three-dimensional organization of Drosophila melanogaster interphase nuclei. II. Chromosome spatial organization and gene regulation. Journal of Cell Biology, 104(6):1471–??, June 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/104/6/1471.

Heuser:1989:TBK

[HS89] J. Heuser and C. J. Steer. Trimeric binding of the 70-kD uncoating ATPase to the vertices of clathrin triskelia: a candidate intermediate in the vesicle uncoating reaction. Journal of Cell Biology, 109(4):1457–??, October 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/109/4/1457.

Hauri:1985:EIT

[HSB^{+}85] H. P. Hauri, E. E. Sterchi, D. Bienz, J. A. Fransen, and A. Marxer. Expression and intracellular transport of microvillus membrane hydrolases in human intestinal epithelial cells. Journal of Cell Biology, 101(3):838–??, September 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/101/3/838.

Heidaran:1988:CST

[HSK88] M. A. Heidaran, R. M. Showman, and W. S. Kistler. A cytochemical study of the transcriptional and translational regulation of nuclear transition protein 1 (TP1), a major chromosomal protein of mammalian spermatids. Journal of Cell Biology, 106(5):1427–??, May 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/106/5/1427.

Hirokawa:1989:CAP

[HSK^{+}89a] N. Hirokawa, K. Sobue, K. Kanda, A. Harada, and H. Yorifuji. The cytoskeletal architecture of the presynaptic terminal and molecular structure of synapsin 1. Journal of Cell Biology, 108 (1):111–??, January 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/108/1/111.

Hirvonen:1989:MPO

[HSK^{+}89b] H. Hirvonen, M. Sandberg, H. Kalimo, V. Hukkanen, E. Vuorio, T. T. Salmi, and K. Alitalo. The N-myc proto-oncogene and IGF–II growth factor mRNAs are expressed by distinct cells in human fetal kidney and brain. Journal of Cell Biology, 108 (3):1093–??, March 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/108/3/1093.

Huet:1987:AMS

[HSMCL87] C. Huet, C. Sahuquillo-Merino, E. Coudrier, and D. Louvard. Absorptive and mucus-secreting subclones isolated from a multipotent intestinal cell line (HT-29) provide new models for cell polarity and terminal differentiation. Journal of Cell Biology, 105(1):345–??, July 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/1/345.

Howe:1985:MCR

[HSMM85] C. L. Howe, L. M. Sacramone, M. S. Mooseker, and J. S. Morrow. Mechanisms of cytoskeletal regulation: modulation of membrane affinity in avian brush border and erythrocyte spectrins. Journal of Cell Biology, 101(4):1379–??, October 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/101/4/1379.

Hirokawa:1988:TPM

[HSO88] N. Hirokawa, Y. Shiomura, and S. Okabe. Tau proteins: the molecular structure and mode of binding on microtubules. Journal of Cell Biology, 107(4):1449–??, October 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/107/4/1449.

Harrison:1985:ACC

[HSS85] J. J. Harrison, E. Soudry, and R. Sager. Adipocyte conversion of CHEF cells in serum-free medium. Journal of Cell Biology, 100 (2):429–??, February 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/100/2/429.

Holt:1987:NPC

[HSS^{+}87] G. D. Holt, C. M. Snow, A. Senior, R. S. Haltiwanger, L. Gerace, and G. W. Hart. Nuclear pore complex glycoproteins contain cytoplasmically disposed O-linked N-acetylglucosamine. Journal of Cell Biology, 104(5):1157–??, May 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/104/5/1157.

Horovitz:1989:RAR

[HSS89] O. Horovitz, V. Spitsberg, and M. M. Salpeter. Regulation of acetylcholine receptor synthesis at the level of translation in rat primary muscle cells. Journal of Cell Biology, 108(5):1817–??, May 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/108/5/
1817.

Hardwick:1986:ATD

[HSWH86] J. M. Hardwick, K. E. Shaw, J. W. Wills, and E. Hunter. Aminoterminal deletion mutants of the Rous sarcoma virus glycoprotein do not block signal peptide cleavage but can block intracellular transport. Journal of Cell Biology, 103(3):829–??, September 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/103/3/
829.

Haylett:1986:LST

[HT86] T. Haylett and L. Thilo. Limited and selective transfer of plasma membrane glycoproteins to membrane of secondary lysosomes. Journal of Cell Biology, 103(4):1249–??, October 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/103/4/1249.

Huffaker:1988:DEB

[HTB88] T. C. Huffaker, J. H. Thomas, and D. Botstein. Diverse effects of beta-tubulin mutations on microtubule formation and function. Journal of Cell Biology, 106(6):1997–??, June 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/106/6/1997.

Hoffman:1985:CNT

[HTGP85] P. N. Hoffman, G. W. Thompson, J. W. Griffin, and D. L. Price. Changes in neurofilament transport coincide temporally with alterations in the caliber of axons in regenerating motor fibers. Journal of Cell Biology, 101(4):1332–??, October 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/101/4/1332.

Hirokawa:1985:CAIb

[HTH85] N. Hirokawa, R. Takemura, and S. Hisanaga. Cytoskeletal architecture of isolated mitotic spindle with special reference to microtubule-associated proteins and cytoplasmic dynein. Journal of Cell Biology, 101(5):1858–??, November 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/101/5/1858.

Harlan:1986:ATI

[HTRBP86] J. M. Harlan, P. J. Thompson, R. R. Ross, and D. F. Bowen-Pope. Alpha-thrombin induces release of platelet-derived growth factor-like molecule(s) by cultured human endothelial cells. Journal of Cell Biology, 103(3):1129–??, September 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/103/3/1129.

Hamaguchi:1985:DFL

[HTSH85] Y. Hamaguchi, M. Toriyama, H. Sakai, and Y. Hiramoto. Distribution of fluorescently labeled tubulin injected into sand dollar eggs from fertilization through cleavage. Journal of Cell Biology, 100(4):1262–??, April 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/100/4/1262.

Hieda:1989:NHM

[HTT89] Y. Hieda, S. Tsukita, and S. Tsukita. A new high molecular mass protein showing unique localization in desmosomal plaque. Journal of Cell Biology, 109(4):1511–??, October 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/109/4/1511.

Haimann:1985:MQS

[HTTFC85] C. Haimann, F. Torri-Tarelli, R. Fesce, and B. Ceccarelli. Measurement of quantal secretion induced by ouabain and its correlation with depletion of synaptic vesicles. Journal of Cell Biology, 101(5):1953–??, November 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/101/5/1953.

Howard:1986:SMH

[HUA^{+}86] R. J. Howard, S. Uni, M. Aikawa, S. B. Aley, J. H. Leech, A. M. Lew, T. E. Wellems, J. Rener, and D. W. Taylor. Secretion of a malarial histidine-rich protein (Pf HRP II) from Plasmodium falciparum-infected erythrocytes. Journal of Cell Biology, 103 (4):1269–??, October 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/103/4/1269.

Hutchison:1987:LCC

[Hut87] N. Hutchison. Lampbrush chromosomes of the chicken, Gallus domesticus. Journal of Cell Biology, 105(4):1493–??, October 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/105/4/
1493.

Hiriyanna:1988:EMV

[HVBB88] K. T. Hiriyanna, J. Varkey, M. Beer, and R. M. Benbow. Electron microscopic visualization of sites of nascent DNA synthesis by streptavidin-gold binding to biotinylated nucleotides incorporated in vivo. Journal of Cell Biology, 107(1):33–??, July 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/107/1/
33.

Heremans:1989:MAH

[HvdSdC^{+}89] A. Heremans, B. van der Schueren, B. de Cock, M. Paulsson, J. J. Cassiman, H. van den Berghe, and G. David. Matrix-associated heparan sulfate proteoglycan: core protein-specific monoclonal antibodies decorate the pericellular matrix of connective tissue cells and the stromal side of basement membranes. Journal of Cell Biology, 109(6):3199–??, December 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/109/6/3199.

Hinssen:1987:GEE

[HVL87] H. Hinssen, J. Vandekerckhove, and E. Lazarides. Gelsolin is expressed in early erythroid progenitor cells and negatively regulated during erythropoiesis. Journal of Cell Biology, 105(3): 1425–??, September 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/3/1425.

Hollyfield:1985:EDI

[HVR^{+}85] J. G. Hollyfield, H. H. Varner, M. E. Rayborn, G. I. Liou, and C. D. Bridges. Endocytosis and degradation of interstitial retinol-binding protein: differential capabilities of cells that border the interphotoreceptor matrix. Journal of Cell Biology, 100(5):1676–??, May 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/100/5/1676.

Harland:1985:TMI

[HW85a] R. Harland and H. Weintraub. Translation of mRNA injected into Xenopus oocytes is specifically inhibited by antisense RNA. Journal of Cell Biology, 101(3):1094–??, September 1985. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/101/3/1094.

Hoops:1985:BBA

[HW85b] H. J. Hoops and G. B. Witman. Basal bodies and associated structures are not required for normal flagellar motion or phototaxis in the green alga Chlorogonium elongatum. Journal of Cell Biology, 100(1):297–??, January 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/100/1/297.

Hill:1986:MAD

[HW86a] C. Hill and K. Weber. Monoclonal antibodies distinguish titins from heart and skeletal muscle. Journal of Cell Biology, 102 (3):1099–??, March 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/102/3/1099.

Hsu:1986:GCC

[HW86b] Y. M. Hsu and J. L. Wang. Growth control in cultured 3T3 fibroblasts. V. Purification of an Mr 13,000 polypeptide responsible for growth inhibitory activity. Journal of Cell Biology, 102 (2):362–??, February 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/102/2/362.

Hyman:1987:DCD

[HW87] A. A. Hyman and J. G. White. Determination of cell division axes in the early embryogenesis of Caenorhabditis elegans. Journal of Cell Biology, 105(5):2123–??, November 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/105/5/2123.

Hansen:1988:PCT

[HW88] W. Hansen and P. Walter. Prepro-carboxypeptidase Y and a truncated form of pre-invertase, but not full-length pre-invertase, can be posttranslationally translocated across microsomal vesicle membranes from Saccharomyces cerevisiae. Journal of Cell Biology, 106(4):1075–??, April 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/106/4/1075.

Hall:1989:IAN

[HWDC89] A. L. Hall, V. Warren, S. Dharmawardhane, and J. Condeelis. Identification of actin nucleation activity and polymerization inhibitor in ameboid cells: their regulation by chemotactic stimulation. Journal of Cell Biology, 109(5):2207–??, November 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/109/5/2207.

Huang:1988:PCB

[HWLS88] B. Huang, D. M. Watterson, V. D. Lee, and M. J. Schibler. Purification and characterization of a basal body-associated Ca^{2+}- binding protein. Journal of Cell Biology, 107(1):121–??, July 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/107/1/
121.

Hoffman:1989:DSI

[HWSK89] E. P. Hoffman, S. C. Watkins, H. S. Slayter, and L. M. Kunkel. Detection of a specific isoform of alpha-actinin with antisera directed against dystrophin. Journal of Cell Biology, 108(2): 503–??, February 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/108/2/503.

Hyman:1989:CME

[Hym89] A. A. Hyman. Centrosome movement in the early divisions of Caenorhabditis elegans: a cortical site determining centrosome position. Journal of Cell Biology, 109(3):1185–??, September 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/109/3/
1185.

Hayward:1988:CLM

[HZS88] L. J. Hayward, Y. Y. Zhu, and R. J. Schwartz. Cellular localization of muscle and nonmuscle actin mRNAs in chicken primary myogenic cultures: the induction of alpha-skeletal actin mRNA is regulated independently of alpha-cardiac actin gene expression.

Journal of Cell Biology, 106(6):2077–??, June 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/106/6/2077.

Ibrahimi:1987:DTP

[Ibr87a] I. Ibrahimi. Dithiothreitol and the translocation of preprolactin across mammalian endoplasmic reticulum. Journal of Cell Biology, 105(4):1555–??, October 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/105/4/1555.

Ibrahimi:1987:SRP

[Ibr87b] I. Ibrahimi. Signal recognition particle causes a transient arrest in the biosynthesis of prepromelittin and mediates its translocation across mammalian endoplasmic reticulum. Journal of Cell Biology, 104(1):61–??, January 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/104/1/61.

Imam:1985:CCW

[IBSS85] S. H. Imam, M. J. Buchanan, H. C. Shin, and W. J. Snell. The Chlamydomonas cell wall: characterization of the wall framework. Journal of Cell Biology, 101(4):1599–??, October 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/101/4/1599.

Iacopetta:1986:RIC

[ICP^{+}86] B. Iacopetta, J. L. Carpentier, T. Pozzan, D. P. Lew, P. Gorden, and L. Orci. Role of intracellular calcium and protein kinase C in the endocytosis of transferrin and insulin by HL60 cells. Journal of Cell Biology, 103(3):851–??, September 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/103/3/851.

Ingold:1988:IKD

[ICS88] A. L. Ingold, S. A. Cohn, and J. M. Scholey. Inhibition of kinesindriven microtubule motility by monoclonal antibodies to kinesin heavy chains. Journal of Cell Biology, 107(6):2657–??, December 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/107/6/

Ince:1988:PHM

[ICY^{+}88] C. Ince, J. M. Coremans, D. L. Ypey, P. C. Leijh, A. A. Verveen, and R. van Furth. Phagocytosis by human macrophages is accompanied by changes in ionic channel currents. Journal of Cell Biology, 106(6):1873–??, June 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/106/6/1873.

Ishiwata:1985:DAB

[IF85] S. Ishiwata and T. Funatsu. Does actin bind to the ends of thin filaments in skeletal muscle? Journal of Cell Biology, 100 (1):282–??, January 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/100/1/282.

Ingber:1989:MSB

[IF89] D. E. Ingber and J. Folkman. Mechanochemical switching between growth and differentiation during fibroblast growth factorstimulated angiogenesis in vitro: role of extracellular matrix. Journal of Cell Biology, 109(1):317–??, July 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/109/1/317.

Isom:1987:PLS

[IGSGD87] I. Isom, I. Georgoff, M. Salditt-Georgieff, and J. E. Darnell. Persistence of liver-specific messenger RNA in cultured hepatocytes: different regulatory events for different genes. Journal of Cell Biology, 105(6):2877–??, December 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/105/6/2877.

Ishihara:1988:ALR

[IHJ88] A. Ishihara, B. Holifield, and K. Jacobson. Analysis of lateral redistribution of a plasma membrane glycoprotein-monoclonal antibody complex [corrected]. Journal of Cell Biology, 106(2): 329–??, February 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/106/2/329.

Ishigami:1987:DAC

[IKH87] M. Ishigami, K. Kuroda, and S. Hatano. Dynamic aspects of the contractile system in Physarum plasmodium. III. Cyclic contraction-relaxation of the plasmodial fragment in accordance with the generation-degeneration of cytoplasmic actomyosin fibrils. Journal of Cell Biology, 105(1):381–??, July 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/105/1/381.

Isaacs:1989:BTC

[IKSF89] W. B. Isaacs, I. S. Kim, A. Struve, and A. B. Fulton. Biosynthesis of titin in cultured skeletal muscle cells. Journal of Cell Biology, 109(5):2189–??, November 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/109/5/2189.

Isenberg:1987:PFR

[IMP^{+}87] W. M. Isenberg, R. P. McEver, D. R. Phillips, M. A. Shuman, and D. F. Bainton. The platelet fibrinogen receptor: an immunogoldsurface replica study of agonist-induced ligand binding and receptor clustering. Journal of Cell Biology, 104(6):1655–??, June 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/104/6/
1655.

Ip:1988:MDI

[Ip88] W. Ip. Modulation of desmin intermediate filament assembly by a monoclonal antibody. Journal of Cell Biology, 106 (3):735–??, March 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/106/3/735.

Imam:1988:CCW

[IS88] S. H. Imam and W. J. Snell. The Chlamydomonas cell wall degrading enzyme, lysin, acts on two substrates within the framework of the wall. Journal of Cell Biology, 106(6):2211–??, June 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/106/6/
2211.

Imanaka:1987:TAC

[ISL87] T. Imanaka, G. M. Small, and P. B. Lazarow. Translocation of acyl–CoA oxidase into peroxisomes requires ATP hydrolysis but not a membrane potential. Journal of Cell Biology, 105(6): 2915–??, December 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/6/2915.

Irwin:1985:MAA

[ISM85] M. H. Irwin, S. H. Silvers, and R. Mayne. Monoclonal antibody against chicken type IX collagen: preparation, characterization, and recognition of the intact form of type IX collagen secreted by chondrocytes. Journal of Cell Biology, 101(3):814–??, September 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/101/3/
814.

Ito:1988:IET

[ITY^{+}88] T. Ito, T. Tanaka, T. Yoshida, K. Onoda, H. Ohta, M. Hagiwara, Y. Itoh, M. Ogura, H. Saito, and H. Hidaka. Immunocytochemical evidence for translocation of protein kinase C in human megakaryoblastic leukemic cells: synergistic effects of Ca^{2+} and activators of protein kinase C on the plasma membrane association. Journal of Cell Biology, 107(3):929–??, September 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/107/3/
929.

Jacobs:1988:FMS

[JASP88] C. W. Jacobs, A. E. Adams, P. J. Szaniszlo, and J. R. Pringle. Functions of microtubules in the Saccharomyces cerevisiae cell cycle. Journal of Cell Biology, 107(4):1409–??, October 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/107/4/1409.

Jones:1986:FGM

[JAT86] G. E. Jones, R. G. Arumugham, and M. L. Tanzer. Fibronectin glycosylation modulates fibroblast adhesion and spreading. Journal of Cell Biology, 103(5):1663–??, November 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/103/5/1663.

Jochen:1986:CSA

[JB86] A. L. Jochen and P. Berhanu. Chymotrypsin substrate analogues inhibit endocytosis of insulin and insulin receptors in adipocytes. Journal of Cell Biology, 103(5):1807–??, November 1986. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/103/5/1807.

Jacob:1987:EPD

[JB87] M. H. Jacob and D. K. Berg. Effects of preganglionic denervation and postganglionic axotomy on acetylcholine receptors in the chick ciliary ganglion. Journal of Cell Biology, 105(4): 1847–??, October 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/4/1847.

Jalkanen:1987:LRH

[JBdlTB87] S. Jalkanen, R. F. Bargatze, J. de los Toyos, and E. C. Butcher. Lymphocyte recognition of high endothelium: antibodies to distinct epitopes of an 85-95-kD glycoprotein antigen differentially inhibit lymphocyte binding to lymph node, mucosal, or synovial endothelial cells. Journal of Cell Biology, 105(2):983–??, August 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/105/2/
983.

Jingami:1987:PDM

[JBG^{+}87] H. Jingami, M. S. Brown, J. L. Goldstein, R. G. Anderson, and K. L. Luskey. Partial deletion of membrane-bound domain of 3-hydroxy-3-methylglutaryl coenzyme A reductase eliminates sterol-enhanced degradation and prevents formation of crystalloid endoplasmic reticulum. Journal of Cell Biology, 104 (6):1693–??, June 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/104/6/1693.

Jesaitis:1988:LSN

[JBTA88] A. J. Jesaitis, G. M. Bokoch, J. O. Tolley, and R. A. Allen. Lateral segregation of neutrophil chemotactic receptors into actinand fodrin-rich plasma membrane microdomains depleted in guanyl nucleotide regulatory proteins. Journal of Cell Biology, 107(3):921–??, September 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/107/3/921.

Joshi:1989:DUB

[JC89] H. C. Joshi and D. W. Cleveland. Differential utilization of betatubulin isotypes in differentiating neurites. Journal of Cell Biology, 109(2):663–??, August 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/109/2/663.

Joshi:1985:TCC

[JCBH85] H. C. Joshi, D. Chu, R. E. Buxbaum, and S. R. Heidemann. Tension and compression in the cytoskeleton of PC 12 neurites. Journal of Cell Biology, 101(3):697–??, September 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/101/3/697.

Johansen:1989:NGP

[JFAT89] K. M. Johansen, R. G. Fehon, and S. Artavanis-Tsakonas. The notch gene product is a glycoprotein expressed on the cell surface of both epidermal and neuronal precursor cells during Drosophila development. Journal of Cell Biology, 109(5):2427–??, November 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/109/5/
2427.

Joly:1989:MBF

[JFP89] J. C. Joly, G. Flynn, and D. L. Purich. The microtubulebinding fragment of microtubule-associated protein-2: location of the protease-accessible site and identification of an assembly-promoting peptide. Journal of Cell Biology, 109(5): 2289–??, November 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/5/2289.

Jones:1985:IFI

[JG85] J. C. Jones and R. D. Goldman. Intermediate filaments and the initiation of desmosome assembly. Journal of Cell Biology, 101(2):506–??, August 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/101/2/506.

Joiner:1989:OLS

[JGAR89] K. A. Joiner, T. Ganz, J. Albert, and D. Rotrosen. The opsonizing ligand on Salmonella typhimurium influences incorporation of specific, but not azurophil, granule constituents into neutrophil phagosomes. Journal of Cell Biology, 109(6):2771–??, December 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/109/6/

Jones:1985:OFI

[JGYG85] J. C. Jones, A. E. Goldman, H. Y. Yang, and R. D. Goldman. The organizational fate of intermediate filament networks in two epithelial cell types during mitosis. Journal of Cell Biology, 100 (1):93–??, January 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/100/1/93.

Joyce:1985:CPPa

[JHP85a] N. C. Joyce, M. F. Haire, and G. E. Palade. Contractile proteins in pericytes. I. Immunoperoxidase localization of tropomyosin. Journal of Cell Biology, 100(5):1379–??, May 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/100/5/1379.

Joyce:1985:CPPb

[JHP85b] N. C. Joyce, M. F. Haire, and G. E. Palade. Contractile proteins in pericytes. II. Immunocytochemical evidence for the presence of two isomyosins in graded concentrations. Journal of Cell Biology, 100(5):1387–??, May 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/100/5/1387.

Jensen:1986:BPI

[JHPS86] K. H. Jensen, D. L. Herrin, F. G. Plumley, and G. W. Schmidt. Biogenesis of photosystem II complexes: transcriptional, translational, and posttranslational regulation. Journal of Cell Biology, 103(4):1315–??, October 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/103/4/1315.

Johnson:1985:CAB

[JIFH85] C. H. Johnson, S. Inoué, A. Flint, and J. W. Hastings. Compartmentalization of algal bioluminescence: autofluorescence of bioluminescent particles in the dinoflagellate Gonyaulax as studied with image-intensified video microscopy and flow cytometry. Journal of Cell Biology, 100(5):1435–??, May 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/100/5/1435.

Jorgensen:1987:IML

[JJ87] A. O. Jorgensen and L. R. Jones. Immunoelectron microscopical localization of phospholamban in adult canine ventricular muscle.

Journal of Cell Biology, 104(5):1343–??, May 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/104/5/1343.

Jones:1989:ITB

[JJP^{+}89] P. Jones, P. Jackson, G. J. Price, B. Patel, V. Ohanion, A. L. Lear, and D. R. Critchley. Identification of a talin binding site in the cytoskeletal protein vinculin. Journal of Cell Biology, 109 (6):2917–??, December 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/6/2917.

Jacob:1986:SID

[JLB86] M. H. Jacob, J. M. Lindstrom, and D. K. Berg. Surface and intracellular distribution of a putative neuronal nicotinic acetylcholine receptor. Journal of Cell Biology, 103(1):205–??, July 1986. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/103/1/205.

Johnson:1986:LIJ

[JLH^{+}86] K. R. Johnson, P. D. Lampe, K. C. Hur, C. F. Louis, and R. G. Johnson. A lens intercellular junction protein, MP26, is a phosphoprotein. Journal of Cell Biology, 102(4):1334–??, April 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/102/4/1334.

Jaken:1989:ATP

[JLK89] S. Jaken, K. Leach, and T. Klauck. Association of type 3 protein kinase C with focal contacts in rat embryo fibroblasts. Journal of Cell Biology, 109(2):697–??, August 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/109/2/697.

Jacob:1986:EGD

[JM86] J. S. Jacob and K. R. Miller. The effects of galactolipid depletion on the structure of a photosynthetic membrane. Journal of Cell Biology, 103(4):1337–??, October 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/103/4/1337.

Jessen:1985:GEN

[JMBM85] K. R. Jessen, L. Morgan, M. Brammer, and R. Mirsky. Galactocerebroside is expressed by non-myelin-forming Schwann cells in situ. Journal of Cell Biology, 101(3):1135–??, September 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/101/3/1135.

Johnson:1988:AMM

[JMW88] C. S. Johnson, N. M. McKenna, and Y. Wang. Association of microinjected myosin and its subfragments with myofibrils in living muscle cells. Journal of Cell Biology, 107(6):2213–??, December 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/107/6/
2213.

Jones:1985:FRM

[JNA85] D. H. Jones, J. Nusbacher, and C. L. Anderson. Fc receptormediated binding and endocytosis by human mononuclear phagocytes: monomeric IgG is not endocytosed by U937 cells and monocytes. Journal of Cell Biology, 100(2):558–??, February 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/100/2/
558.

Jalkanen:1985:HSP

[JNR^{+}85] M. Jalkanen, H. Nguyen, A. Rapraeger, N. Kurn, and M. Bernfield. Heparan sulfate proteoglycans from mouse mammary epithelial cells: localization on the cell surface with a monoclonal antibody. Journal of Cell Biology, 101(3):976–??, September 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/101/3/
976.

Jones:1985:IIF

[JP85] L. N. Jones and F. M. Pope. Isolation of intermediate filament assemblies from human hair follicles. Journal of Cell Biology, 101 (4):1569–??, October 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/101/4/1569.

Johnson:1985:RCS

[JPH85] M. I. Johnson, K. Paik, and D. Higgins. Rapid changes in synaptic vesicle cytochemistry after depolarization of cultured cholinergic sympathetic neurons. Journal of Cell Biology, 101(1):217–??, July 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/101/1/

Job:1985:GMS

[JPM85] D. Job, M. Pabion, and R. L. Margolis. Generation of microtubule stability subclasses by microtubule-associated proteins: implications for the microtubule “dynamic instability” model. Journal of Cell Biology, 101(5):1680–??, November 1985. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/101/5/1680.

Jaakkola:1989:PCC

[JPU89] S. Jaakkola, J. Peltonen, and J. J. Uitto. Perineurial cells coexpress genes encoding interstitial collagens and basement membrane zone components. Journal of Cell Biology, 108(3):1157–??, March 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
108/3/1157.

Jalkanen:1988:MME

[JRB88] M. Jalkanen, A. Rapraeger, and M. Bernfield. Mouse mammary epithelial cells produce basement membrane and cell surface heparan sulfate proteoglycans containing distinct core proteins. Journal of Cell Biology, 106(3):953–??, March 1988. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/106/3/953.

Jalkanen:1987:CSP

[JRSB87] M. Jalkanen, A. Rapraeger, S. Saunders, and M. Bernfield. Cell surface proteoglycan of mouse mammary epithelial cells is shed by cleavage of its matrix-binding ectodomain from its membraneassociated domain. Journal of Cell Biology, 105(6):3087–??, December 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
105/6/3087.

Jacobs:1986:CON

[JS86a] J. R. Jacobs and J. K. Stevens. Changes in the organization of the neuritic cytoskeleton during nerve growth factor-activated differentiation of PC12 cells: a serial electron microscopic study of the development and control of neurite shape. Journal of Cell Biology, 103(3):895–??, September 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/103/3/895.

Jacobs:1986:EMP

[JS86b] J. R. Jacobs and J. K. Stevens. Experimental modification of PC12 neurite shape with the microtubule-depolymerizing drug Nocodazole: a serial electron microscopic study of neurite shape control. Journal of Cell Biology, 103(3):907–??, September 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/103/3/907.

Jensen:1986:ASO

[JS86c] C. G. Jensen and B. H. Smaill. Analysis of the spatial organization of microtubule-associated proteins. Journal of Cell Biology, 103(2):559–??, August 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/103/2/559.

Jiang:1986:CFI

[JS86d] L. W. Jiang and M. Schindler. Chemical factors that influence nucleocytoplasmic transport: a fluorescence photobleaching study. Journal of Cell Biology, 102(3):853–??, March 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/102/3/853.

Jiang:1988:NTF

[JS88a] L. W. Jiang and M. Schindler. Nuclear transport in 3T3 fibroblasts: effects of growth factors, transformation, and cell shape. Journal of Cell Biology, 106(1):13–??, January 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/106/1/13.

Johansson:1988:IPC

[JS88b] M. W. Johansson and K. Söderhäll. Isolation and purification of a cell adhesion factor from crayfish blood cells. Journal of Cell Biology, 106(5):1795–??, May 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/106/5/1795.

Jorgensen:1989:SDD

[JSA^{+}89] A. O. Jorgensen, A. C. Shen, W. Arnold, A. T. Leung, and K. P. Campbell. Subcellular distribution of the 1,4-dihydropyridine receptor in rabbit skeletal muscle in situ: an immunofluorescence and immunocolloidal gold-labeling study. Journal of Cell Biology, 109(1):135–??, July 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/109/1/135.

Jorgensen:1985:ULC

[JSC85] A. O. Jorgensen, A. C. Shen, and K. P. Campbell. Ultrastructural localization of calsequestrin in adult rat atrial and ventricular muscle cells. Journal of Cell Biology, 101(1):257–??, July 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/101/1/
257.

Joseph-Silverstein:1989:BFG

[JSCLP89] J. Joseph-Silverstein, S. A. Consigli, K. M. Lyser, and C. Ver Pault. Basic fibroblast growth factor in the chick embryo: immunolocalization to striated muscle cells and their precursors. Journal of Cell Biology, 108(6):2459–??, June 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/108/6/2459.

Jentoft:1985:IPD

[JT85] J. E. Jentoft and C. D. Town. Intracellular pH in Dictyostelium discoideum: a 31P nuclear magnetic resonance study. Journal of Cell Biology, 101(3):778–??, September 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/101/3/778.

Jesaitis:1989:RCR

[JTBA89] A. J. Jesaitis, J. O. Tolley, G. M. Bokoch, and R. A. Allen. Regulation of chemoattractant receptor interaction with transducing proteins by organizational control in the plasma membrane of human neutrophils. Journal of Cell Biology, 109(6): 2783–??, December 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/6/2783.

Jerdan:1985:ICF

[JVG^{+}85] J. A. Jerdan, H. H. Varner, J. H. Greenberg, V. J. Horn, and G. R. Martin. Isolation and characterization of a factor from calf serum that promotes the pigmentation of embryonic and transformed melanocytes. Journal of Cell Biology, 100(5):1493–??, May 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/100/5/

Jackson:1985:MPD

[JWH85] R. C. Jackson, K. K. Ward, and J. G. Haggerty. Mild proteolytic digestion restores exocytotic activity to N-ethylmaleimide-inactivated cell surface complex from sea urchin eggs. Journal of Cell Biology, 101(1):6–??, July 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/101/1/6.

Jesaitis:1986:LMK

[JY86] A. J. Jesaitis and J. Yguerabide. The lateral mobility of the (Na+,K+)-dependent ATPase in Madin–Darby canine kidney cells. Journal of Cell Biology, 102(4):1256–??, April 1986. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/102/4/1256.

Joshi:1987:VCD

[JYC87] H. C. Joshi, T. J. Yen, and D. W. Cleveland. In vivo coassembly of a divergent beta-tubulin subunit (c beta 6) into microtubules of different function. Journal of Cell Biology, 105(5): 2179–??, November 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/5/2179.

Jones:1986:FAP

[JYG86] J. C. Jones, K. M. Yokoo, and R. D. Goldman. Further analysis of pemphigus autoantibodies and their use in studies on the heterogeneity, structure, and function of desmosomes. Journal of Cell Biology, 102(3):1109–??, March 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/102/3/1109.

Kabcenell:1985:PRE

[KA85] A. K. Kabcenell and P. H. Atkinson. Processing of the rough endoplasmic reticulum membrane glycoproteins of rotavirus SA11. Journal of Cell Biology, 101(4):1270–??, October 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/101/4/1270.

Karr:1986:OCE

[KA86] T. L. Karr and B. M. Alberts. Organization of the cytoskeleton in early Drosophila embryos. Journal of Cell Biology, 102 (4):1494–??, April 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/102/4/1494.

Kenna:1988:SEC

[KAB88] M. Kenna, E. Amaya, and K. Bloom. Selective excision of the centromere chromatin complex from Saccharomyces cerevisiae. Journal of Cell Biology, 107(1):9–??, July 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/107/1/9.

Kamiya:1988:MTI

[Kam88] R. Kamiya. Mutations at twelve independent loci result in absence of outer dynein arms in Chylamydomonas reinhardtii. Journal of Cell Biology, 107(6):2253–??, December 1988. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/107/6/2253.

Kuriyama:1985:IMC

[KB85] R. Kuriyama and G. G. Borisy. Identification of molecular components of the centrosphere in the mitotic spindle of sea urchin eggs. Journal of Cell Biology, 101(2):524–??, August 1985. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/101/2/524.

Keen:1986:PCM

[KB86] J. H. Keen and M. M. Black. The phosphorylation of coated membrane proteins in intact neurons. Journal of Cell Biology, 102(4):1325–??, April 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/102/4/1325.

Kalomiris:1988:MLC

[KB88] E. L. Kalomiris and L. Y. Bourguignon. Mouse T lymphoma cells contain a transmembrane glycoprotein (GP85) that binds ankyrin. Journal of Cell Biology, 106(2):319–??, February 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/106/2/319.

Knoll:1988:FLP

[KBB^{+}88] G. Knoll, K. N. Burger, R. Bron, G. van Meer, and A. J. Verkleij. Fusion of liposomes with the plasma membrane of epithelial cells: fate of incorporated lipids as followed by freeze fracture and autoradiography of plastic sections. Journal of Cell Biology, 107 (6):2511–??, December 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/107/6/2511.

Kozarsky:1986:UFL

[KBK86] K. F. Kozarsky, H. A. Brush, and M. Krieger. Unusual forms of low density lipoprotein receptors in hamster cell mutants with defects in the receptor structural gene. Journal of Cell Biology, 102(5):1567–??, May 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/102/5/1567.

Kawamoto:1986:ICL

[KBKC86] R. M. Kawamoto, J. P. Brunschwig, K. C. Kim, and A. H. Caswell. Isolation, characterization, and localization of the spanning protein from skeletal muscle triads. Journal of Cell Biology, 103(4):1405–??, October 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/103/4/1405.

Klotz:1986:MAP

[KBL^{+}86] C. Klotz, N. Bordes, M. C. Laine, D. Sandoz, and M. Bornens. Myosin at the apical pole of ciliated epithelial cells as revealed by a monoclonal antibody. Journal of Cell Biology, 103 (2):613–??, August 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/103/2/613.

Klein:1988:GBM

[KBO^{+}88] D. J. Klein, D. M. Brown, T. R. Oegema, P. E. Brenchley, J. C. Anderson, M. A. Dickinson, E. A. Horigan, and J. R. Hassell. Glomerular basement membrane proteoglycans are derived from a large precursor. Journal of Cell Biology, 106(3):963–??, March 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/106/3/
963.

Kaput:1989:VIC

[KBPW89] J. Kaput, M. C. Brandriss, and T. Prussak-Wieckowska. In vitro import of cytochrome c peroxidase into the intermembrane space: release of the processed form by intact mitochondria. Journal of Cell Biology, 109(1):101–??, July 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/109/1/101.

Kachar:1987:DSC

[KBR87] B. Kachar, P. C. Bridgman, and T. S. Reese. Dynamic shape changes of cytoplasmic organelles translocating along microtubules. Journal of Cell Biology, 105(3):1267–??, September 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/105/3/
1267.

Kao:1985:NGS

[KC85] W. Y. Kao and S. T. Case. A novel giant secretion polypeptide in Chironomus salivary glands: implications for another Balbiani ring gene. Journal of Cell Biology, 101(3):1044–??, September 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/101/3/
1044.

Kazlauskas:1988:PKC

[KC88] A. Kazlauskas and J. A. Cooper. Protein kinase C mediates platelet-derived growth factor-induced tyrosine phosphorylation of p42. Journal of Cell Biology, 106(4):1395–??, April 1988. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/106/4/1395.

Kargacin:1989:POC

[KCAF89] G. J. Kargacin, P. H. Cooke, S. B. Abramson, and F. S. Fay. Periodic organization of the contractile apparatus in smooth muscle revealed by the motion of dense bodies in single cells. Journal of Cell Biology, 108(4):1465–??, April 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/108/4/1465.

Keller:1985:RMT

[KCCM85] T. C. Keller, K. A. Conzelman, R. Chasan, and M. S. Mooseker. Role of myosin in terminal web contraction in isolated intestinal epithelial brush borders. Journal of Cell Biology, 100(5):1647–??, May 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/100/5/
1647.

Kaczmarek:1987:CHR

[KCK^{+}87] L. Kaczmarek, B. Calabretta, H. T. Kao, N. Heintz, J. Nevins, and R. Baserga. Control of hsp70 RNA levels in human lymphocytes. Journal of Cell Biology, 104(2):183–??, February 1987.

CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/104/2/183.

Kordeli:1986:EPD

[KCN^{+}86] E. Kordeli, J. Cartaud, H. O. Nghiêm, L. A. Pradel, C. Dubreuil, D. Paulin, and J. P. Changeux. Evidence for a polarity in the distribution of proteins from the cytoskeleton in Torpedo marmorata electrocytes. Journal of Cell Biology, 102(3):748–??, March 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
102/3/748.

Kordeli:1989:AAA

[KCN^{+}89] E. Kordeli, J. Cartaud, H. O. Nghiêm, A. Devillers-Thiéry, and J. P. Changeux. Asynchronous assembly of the acetylcholine receptor and of the 43-kD nu1 protein in the postsynaptic membrane of developing Torpedo marmorata electrocyte. Journal of Cell Biology, 108(1):127–??, January 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/108/1/127.

Knox:1986:RFM

[KCR86] P. Knox, S. Crooks, and C. S. Rimmer. Role of fibronectin in the migration of fibroblasts into plasma clots. Journal of Cell Biology, 102(6):2318–??, June 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/102/6/2318.

Koenig:1987:QSS

[KDB87] C. S. Koenig, M. Dabiké, and M. Bronfman. Quantitative subcellular study of apical pole membranes from chicken oxyntic cells in resting and HCl secretory state. Journal of Cell Biology, 105 (6):2945–??, December 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/6/2945.

Krotoski:1986:DPC

[KDBF86] D. M. Krotoski, C. Domingo, and M. Bronner-Fraser. Distribution of a putative cell surface receptor for fibronectin and laminin in the avian embryo. Journal of Cell Biology, 103(3): 1061–??, September 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/103/3/1061.

Kwan:1989:CST

[KDFG89] A. P. Kwan, I. R. Dickson, A. J. Freemont, and M. E. Grant. Comparative studies of type X collagen expression in normal and rachitic chicken epiphyseal cartilage. Journal of Cell Biology, 109 (4):1849–??, October 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/4/1849.

Kelley:1987:MPC

[KDHS87] C. Kelley, P. D’Amore, H. B. Hechtman, and D. Shepro. Microvascular pericyte contractility in vitro: comparison with other cells of the vascular wall. Journal of Cell Biology, 104(3):483–??, March 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
104/3/483.

Kohtz:1989:GPD

[KDIG89] D. S. Kohtz, N. R. Dische, T. Inagami, and B. Goldman. Growth and partial differentiation of presumptive human cardiac myoblasts in culture. Journal of Cell Biology, 108(3):1067–??, March 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
108/3/1067.

Kan:1986:PAG

[KdS86] F. W. Kan and P. P. da Silva. Preferential association of glycoproteins to the euchromatin regions of cross-fractured nuclei is revealed by fracture-label. Journal of Cell Biology, 102(2): 576–??, February 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/102/2/576.

Kooijman:1989:WGA

[KdWB^{+}89] R. Kooijman, P. de Wildt, S. Beumer, G. van der Vliet, W. Homan, H. Kalshoven, A. Musgrave, and H. van den Ende. Wheat germ agglutinin induces mating reactions in Chlamydomonas eugametos by cross-linking agglutinin-associated glycoproteins in the flagellar membrane. Journal of Cell Biology, 109 (4):1677–??, October 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/4/1677.

Keen:1987:CAP

[Kee87] J. H. Keen. Clathrin assembly proteins: affinity purification and a model for coat assembly. Journal of Cell Biology, 105 (5):1989–??, November 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/5/1989.

Keene:1988:UTV

[KEG88] D. R. Keene, E. Engvall, and R. W. Glanville. Ultrastructure of type VI collagen in human skin and cartilage suggests an anchoring function for this filamentous network. Journal of Cell Biology, 107(5):1995–??, November 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/107/5/1995.

Kiehart:1986:CMD

[KF86] D. P. Kiehart and R. Feghali. Cytoplasmic myosin from Drosophila melanogaster. Journal of Cell Biology, 103(4): 1517–??, October 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/103/4/1517.

Kardami:1989:BFG

[KF89a] E. Kardami and R. R. Fandrich. Basic fibroblast growth factor in atria and ventricles of the vertebrate heart. Journal of Cell Biology, 109(4):1865–??, October 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/109/4/1865.

Kopan:1989:URA

[KF89b] R. Kopan and E. Fuchs. The use of retinoic acid to probe the relation between hyperproliferation-associated keratins and cell proliferation in normal and malignant epidermal cells. Journal of Cell Biology, 109(1):295–??, July 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/109/1/295.

Kellogg:1989:IMA

[KFA89] D. R. Kellogg, C. M. Field, and B. M. Alberts. Identification of microtubule-associated proteins in the centrosome, spindle, and kinetochore of the early Drosophila embryo. Journal of Cell Biology, 109(6):2977–??, December 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/109/6/2977.

Knapp:1986:CNE

[KFH^{+}86] A. C. Knapp, W. W. Franke, H. Heid, M. Hatzfeld, J. L. Jorcano, and R. Moll. Cytokeratin No. 9, an epidermal type I keratin characteristic of a special program of keratinocyte differentiation displaying body site specificity. Journal of Cell Biology, 103(2):657–??, August 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/103/2/657.

Kaufman:1985:EDR

[KFHF85] S. J. Kaufman, R. F. Foster, K. R. Haye, and L. E. Faiman. Expression of a developmentally regulated antigen on the surface of skeletal and cardiac muscle cells. Journal of Cell Biology, 100(6):1977–??, June 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/100/6/1977.

Kucherer:1987:NCE

[KFS87] A. Kücherer, A. Faissner, and M. Schachner. The novel carbohydrate epitope L3 is shared by some neural cell adhesion molecules. Journal of Cell Biology, 104(6):1597–??, June 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/104/6/1597.

Kaufmann:1989:IVU

[KFW^{+}89] R. Kaufmann, D. Frösch, C. Westphal, L. Weber, and C. E. Klein. Integrin VLA-3: ultrastructural localization at cell–cell contact sites of human cell cultures. Journal of Cell Biology, 109 (4):1807–??, October 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/4/1807.

Kato:1985:EEE

[KG85a] Y. Kato and D. Gospodarowicz. Effect of exogenous extracellular matrices on proteoglycan synthesis by cultured rabbit costal chondrocytes. Journal of Cell Biology, 100(2):486–??, February 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/100/2/

Kato:1985:SPS

[KG85b] Y. Kato and D. Gospodarowicz. Sulfated proteoglycan synthesis by confluent cultures of rabbit costal chondrocytes grown in the presence of fibroblast growth factor. Journal of Cell Biology, 100 (2):477–??, February 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/100/2/477.

Kumar:1986:CCH

[KG86] N. M. Kumar and N. B. Gilula. Cloning and characterization of human and rat liver cDNAs coding for a gap junction protein. Journal of Cell Biology, 103(3):767–??, September 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/103/3/767.

Karess:1989:RDG

[KG89] R. E. Karess and D. M. Glover. Rough deal: a gene required for proper mitotic segregation in Drosophila. Journal of Cell Biology, 109(6):2951–??, December 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/109/6/2951.

Khawaja:1988:ESM

[KGB88] S. Khawaja, G. G. Gundersen, and J. C. Bulinski. Enhanced stability of microtubules enriched in detyrosinated tubulin is not a direct function of detyrosination level. Journal of Cell Biology, 106(1):141–??, January 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/106/1/141.

Kohtz:1987:MTF

[KGHK^{+}87] D. S. Kohtz, V. Georgieva-Hanson, J. D. Kohtz, W. J. Schook, and S. Puszkin. Mapping two functional domains of clathrin light chains with monoclonal antibodies. Journal of Cell Biology, 104 (4):897–??, January 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/104/4/897.

Kabat:1985:CAD

[KGRP85] D. Kabat, B. Gliniak, L. Rohrschneider, and E. Polonoff. Cell anchorage determines whether mammary tumor virus glycoproteins are processed for plasma membranes or secretion. Journal of Cell Biology, 101(6):2274–??, December 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/101/6/2274.

Katow:1985:RFP

[KH85a] H. Katow and M. Hayashi. Role of fibronectin in primary mesenchyme cell migration in the sea urchin. Journal of Cell Biology, 101(4):1487–??, October 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/101/4/1487.

Kielian:1985:PIA

[KH85b] M. Kielian and A. Helenius. pH-induced alterations in the fusogenic spike protein of Semliki Forest virus. Journal of Cell Biology, 101(6):2284–??, December 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/101/6/2284.

Kane:1988:CLI

[KHB88] P. M. Kane, D. Holowka, and B. Baird. Cross-linking of IgEreceptor complexes by rigid bivalent antigens greater than 200 A in length triggers cellular degranulation. Journal of Cell Biology, 107(3):969–??, September 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/107/3/969.

Kohorn:1986:FMA

[KHC^{+}86] B. D. Kohorn, E. Harel, P. R. Chitnis, J. P. Thornber, and E. M. Tobin. Functional and mutational analysis of the light-harvesting chlorophyll a/b protein of thylakoid membranes. Journal of Cell Biology, 102(3):972–??, March 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/102/3/972.

Kedersha:1986:SLC

[KHKR86] N. L. Kedersha, D. F. Hill, K. E. Kronquist, and L. H. Rome. Subpopulations of liver coated vesicles resolved by preparative agarose gel electrophoresis. Journal of Cell Biology, 103(1):287– ??, July 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
103/1/287.

Kato:1987:EVC

[KIKS87] Y. Kato, M. Iwamoto, T. Koike, and F. Suzuki. Effect of vanadate on cartilage-matrix proteoglycan synthesis in rabbit costal chondrocyte cultures. Journal of Cell Biology, 104(2): 311–??, February 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/104/2/311.

Kunemund:1988:LHC

[KJF^{+}88] V. Künemund, F. B. Jungalwala, G. Fischer, D. K. Chou, G. Keilhauer, and M. Schachner. The L2/HNK-1 carbohydrate of neural cell adhesion molecules is involved in cell interactions. Journal of Cell Biology, 106(1):213–??, January 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/106/1/213.

Krieg:1989:PTA

[KJW89] U. C. Krieg, A. E. Johnson, and P. Walter. Protein translocation across the endoplasmic reticulum membrane: identification by photocross-linking of a 39-kD integral membrane glycoprotein as part of a putative translocation tunnel. Journal of Cell Biology, 109(5):2033–??, November 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/109/5/2033.

Kwiatkowski:1989:ICF

[KJY89] D. J. Kwiatkowski, P. A. Janmey, and H. L. Yin. Identification of critical functional and regulatory domains in gelsolin. Journal of Cell Biology, 108(5):1717–??, May 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/108/5/1717.

Kistler:1985:IDP

[KKB85] J. Kistler, B. Kirkland, and S. Bullivant. Identification of a 70,000-D protein in lens membrane junctional domains. Journal of Cell Biology, 101(1):28–??, July 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/101/1/28.

Koury:1989:CDF

[KKB89] S. T. Koury, M. J. Koury, and M. C. Bondurant. Cytoskeletal distribution and function during the maturation and enucleation of mammalian erythroblasts. Journal of Cell Biology, 109(6): 3005–??, December 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/6/3005.

Kosher:1986:CGE

[KKG86] R. A. Kosher, W. M. Kulyk, and S. W. Gay. Collagen gene expression during limb cartilage differentiation. Journal of Cell Biology, 102(4):1151–??, April 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/102/4/1151.

Klatte:1989:ICT

[KKGJ89] D. H. Klatte, M. A. Kurpakus, K. A. Grelling, and J. C. Jones. Immunochemical characterization of three components of the hemidesmosome and their expression in cultured epithelial cells. Journal of Cell Biology, 109(6):3377–??, December 1989. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/109/6/3377.

Kubota:1988:RLB

[KKML88] Y. Kubota, H. K. Kleinman, G. R. Martin, and T. J. Lawley. Role of laminin and basement membrane in the morphological differentiation of human endothelial cells into capillary-like structures. Journal of Cell Biology, 107(4):1589–??, October 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/107/4/1589.

Kingsley:1986:TTL

[KKSK86] D. M. Kingsley, K. F. Kozarsky, M. Segal, and M. Krieger. Three types of low density lipoprotein receptor-deficient mutant have pleiotropic defects in the synthesis of N-linked, O-linked, and lipid-linked carbohydrate chains. Journal of Cell Biology, 102 (5):1576–??, May 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/102/5/1576.

Kato:1988:BMP

[KKSK88] M. Kato, Y. Koike, S. Suzuki, and K. Kimata. Basement membrane proteoglycan in various tissues: characterization using monoclonal antibodies to the Engelbreth–Holm–Swarm mouse tumor low density heparan sulfate proteoglycan. Journal of Cell Biology, 106(6):2203–??, June 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/106/6/2203.

Kennedy:1986:EMH

[KKT^{+}86] J. M. Kennedy, S. Kamel, W. W. Tambone, G. Vrbova, and R. Zak. The expression of myosin heavy chain isoforms in normal and hypertrophied chicken slow muscle. Journal of Cell Biology, 103(3):977–??, September 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/103/3/977.

Kalimi:1988:CCG

[KL88] G. H. Kalimi and C. W. Lo. Communication compartments in the gastrulating mouse embryo. Journal of Cell Biology, 107(1):241–??, July 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/107/1/241.

Kalimi:1989:GJC

[KL89a] G. H. Kalimi and C. W. Lo. Gap junctional communication in the extraembryonic tissues of the gastrulating mouse embryo. Journal of Cell Biology, 109(6):3015–??, December 1989. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/109/6/3015.

Katoh:1989:MML

[KL89b] T. Katoh and S. Lowey. Mapping myosin light chains by immunoelectron microscopy. Use of anti-fluorescyl antibodies as structural probes. Journal of Cell Biology, 109(4):1549–??, October 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/109/4/
1549.

Ketis:1988:EHS

[KLHK88] N. V. Ketis, J. Lawler, R. L. Hoover, and M. J. Karnovsky. Effects of heat shock on the expression of thrombospondin by endothelial cells in culture. Journal of Cell Biology, 106(3):893–??, March 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
106/3/893.

Klein:1985:CIR

[KLHS85] C. Klein, J. Lubs-Haukeness, and S. Simons. cAMP induces a rapid and reversible modification of the chemotactic receptor in Dictyostelium discoideum. Journal of Cell Biology, 100 (3):715–??, March 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/100/3/715.

Kensler:1985:EMO

[KLS85] R. W. Kensler, R. J. Levine, and M. Stewart. Electron microscopic and optical diffraction analysis of the structure of scorpion muscle thick filaments. Journal of Cell Biology, 101 (2):395–??, August 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/101/2/395.

Kruskal:1987:CFC

[KM87] B. A. Kruskal and F. R. Maxfield. Cytosolic free calcium increases before and oscillates during frustrated phagocytosis in macrophages. Journal of Cell Biology, 105(6):2685–??, December 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/105/6/
2685.

Kristofferson:1986:DOS

[KMK86] D. Kristofferson, T. Mitchison, and M. Kirschner. Direct observation of steady-state microtubule dynamics. Journal of Cell Biology, 102(3):1007–??, March 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/102/3/1007.

Kim:1987:DDM

[KMKS87] S. Kim, M. Magendantz, W. Katz, and F. Solomon. Development of a differentiated microtubule structure: formation of the chicken erythrocyte marginal band in vivo. Journal of Cell Biology, 104(1):51–??, January 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/104/1/51.

Klein:1988:LRT

[KMM88] R. R. Klein, H. S. Mason, and J. E. Mullet. Light-regulated translation of chloroplast proteins. I. Transcripts of psaA-psaB, psbA, and rbcL are associated with polysomes in dark-grown and illuminated barley seedlings. Journal of Cell Biology, 106 (2):289–??, February 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/106/2/289.

Krawczyk:1988:ETS

[KMP88] Z. Krawczyk, P. Mali, and M. Parvinen. Expression of a testisspecific hsp70 gene-related RNA in defined stages of rat seminiferous epithelium. Journal of Cell Biology, 107(4):1317–??, October 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/107/4/
1317.

Keane:1988:NDN

[KMR^{+}88] R. W. Keane, P. P. Mehta, B. Rose, L. S. Honig, W. R. Loewenstein, and U. Rutishauser. Neural differentiation, NCAMmediated adhesion, and gap junctional communication in neuroectoderm. A study in vitro. Journal of Cell Biology, 106 (4):1307–??, April 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/106/4/1307.

Kwiatkowski:1988:GOB

[KMY88] D. J. Kwiatkowski, R. Mehl, and H. L. Yin. Genomic organization and biosynthesis of secreted and cytoplasmic forms of gelsolin. Journal of Cell Biology, 106(2):375–??, February 1988. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/106/2/375.

Kristensen:1985:IDT

[KNGH^{+}85] P. Kristensen, L. S. Nielsen, J. Grøndahl-Hansen, P. B. Andresen, L. I. Larsson, and K. Danø. Immunocytochemical demonstration of tissue-type plasminogen activator in endocrine cells of the rat pituitary gland. Journal of Cell Biology, 101(1):305–??, July 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/101/1/
305.

Knudsen:1985:CDM

[Knu85] K. A. Knudsen. The calcium-dependent myoblast adhesion that precedes cell fusion is mediated by glycoproteins. Journal of Cell Biology, 101(3):891–??, September 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/101/3/891.

Kaiser:1989:ACD

[KOB89] H. W. Kaiser, E. O’Keefe, and V. Bennett. Adducin: Ca++- dependent association with sites of cell–cell contact. Journal of Cell Biology, 109(2):557–??, August 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/109/2/557.

Keski-Oja:1988:RSA

[KOBLM88] J. Keski-Oja, F. Blasi, E. B. Leof, and H. L. Moses. Regulation of the synthesis and activity of urokinase plasminogen activator in A549 human lung carcinoma cells by transforming growth factorbeta. Journal of Cell Biology, 106(2):451–??, February 1988. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/106/2/451.

Kapprell:1988:IBP

[KOF88] H. P. Kapprell, K. Owaribe, and W. W. Franke. Identification of a basic protein of Mr 75,000 as an accessory desmosomal plaque protein in stratified and complex epithelia. Journal of Cell Biology, 106(5):1679–??, May 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/106/5/1679.

Kolega:1986:EMT

[Kol86] J. Kolega. Effects of mechanical tension on protrusive activity and microfilament and intermediate filament organization in an epidermal epithelium moving in culture. Journal of Cell Biology, 102(4):1400–??, April 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/102/4/1400.

Kozak:1988:PC

[Koz88] M. Kozak. A profusion of controls. Journal of Cell Biology, 107(1):1–??, July 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/107/1/1.

Kozak:1989:SMT

[Koz89] M. Kozak. The scanning model for translation: an update. Journal of Cell Biology, 108(2):229–??, February 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/108/2/229.

Klymkowsky:1985:GAN

[KP85] M. W. Klymkowsky and D. J. Plummer. Giant axonal neuropathy: a conditional mutation affecting cytoskeletal organization.

Journal of Cell Biology, 100(1):245–??, January 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/100/1/245.

Koval:1989:LRB

[KP89] M. Koval and R. E. Pagano. Lipid recycling between the plasma membrane and intracellular compartments: transport and metabolism of fluorescent sphingomyelin analogues in cultured fibroblasts. Journal of Cell Biology, 108(6):2169–??, June 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/108/6/
2169.

Keller:1986:HMC

[KPK86] G. A. Keller, M. Pazirandeh, and S. Krisans. 3-hydroxy-3-methylglutaryl coenzyme A reductase localization in rat liver peroxisomes and microsomes of control and cholestyraminetreated animals: quantitative biochemical and immunoelectron microscopical analyses. Journal of Cell Biology, 103(3):875– ??, September 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/103/3/875.

Kedersha:1986:ICN

[KR86] N. L. Kedersha and L. H. Rome. Isolation and characterization of a novel ribonucleoprotein particle: large structures contain a single species of small RNA. Journal of Cell Biology, 103(3): 699–??, September 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/103/3/699.

Kachar:1988:MCS

[KR88] B. Kachar and T. S. Reese. The mechanism of cytoplasmic streaming in characean algal cells: sliding of endoplasmic reticulum along actin filaments. Journal of Cell Biology, 106(5):1545– ??, May 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
106/5/1545.

Kubalek:1987:LSW

[KRLU87] E. Kubalek, S. Ralston, J. Lindstrom, and N. Unwin. Location of subunits within the acetylcholine receptor by electron image analysis of tubular crystals from Torpedo marmorata. Journal of Cell Biology, 105(1):9–??, July 1987. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/105/1/9.

Kuznetsov:1986:IKP

[KRN^{+}86] S. A. Kuznetsov, V. I. Rodionov, E. S. Nadezhdina, D. B. Murphy, and V. I. Gelfand. Identification of a 34-kD polypeptide as a light chain of microtubule-associated protein-1 (MAP-1) and its association with a MAP-1 peptide that binds to microtubules. Journal of Cell Biology, 102(3):1060–??, March 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/102/3/1060.

Kropf:1986:EPA

[Kro86] D. L. Kropf. Electrophysiological properties of Achlya hyphae: ionic currents studied by intracellular potential recording. Journal of Cell Biology, 102(4):1209–??, April 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/102/4/1209.

Kaplan:1985:ITP

[KS85a] M. R. Kaplan and R. D. Simoni. Intracellular transport of phosphatidylcholine to the plasma membrane. Journal of Cell Biology, 101(2):441–??, August 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/101/2/441.

Kaplan:1985:TCE

[KS85b] M. R. Kaplan and R. D. Simoni. Transport of cholesterol from the endoplasmic reticulum to the plasma membrane. Journal of Cell Biology, 101(2):446–??, August 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/101/2/446.

Koonce:1985:BOT

[KS85c] M. P. Koonce and M. Schliwa. Bidirectional organelle transport can occur in cell processes that contain single microtubules. Journal of Cell Biology, 100(1):322–??, January 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/100/1/322.

Koonce:1986:ROM

[KS86] M. P. Koonce and M. Schliwa. Reactivation of organelle movements along the cytoskeletal framework of a giant freshwater ameba. Journal of Cell Biology, 103(2):605–??, August 1986. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/103/2/605.

Kohno:1988:ASP

[KS88] T. Kohno and T. Shimmen. Accelerated sliding of pollen tube organelles along Characeae actin bundles regulated by Ca^{2+}. Journal of Cell Biology, 106(5):1539–??, May 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/106/5/1539.

Keene:1987:TIC

[KSBB87] D. R. Keene, L. Y. Sakai, H. P. Bächinger, and R. E. Burgeson. Type III collagen can be present on banded collagen fibrils regardless of fibril diameter. Journal of Cell Biology, 105(5): 2393–??, November 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/5/2393.

Keller:1989:SLS

[KSC^{+}89] G. A. Keller, T. J. Scallen, D. Clarke, P. A. Maher, S. K. Krisans, and S. J. Singer. Subcellular localization of sterol carrier protein-2 in rat hepatocytes: its primary localization to peroxisomes. Journal of Cell Biology, 108(4):1353–??, April 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/108/4/1353.

Kaiser:1986:PCT

[KSEP86] D. A. Kaiser, M. Sato, R. F. Ebert, and T. D. Pollard. Purification and characterization of two isoforms of Acanthamoeba profilin. Journal of Cell Biology, 102(1):221–??, January 1986. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/102/1/221.

Kapoor:1988:TVC

[KSF^{+}88] R. Kapoor, L. Y. Sakai, S. Funk, E. Roux, P. Bornstein, and E. H. Sage. Type VIII collagen has a restricted distribution in specialized extracellular matrices. Journal of Cell Biology, 107 (2):721–??, August 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/107/2/721.

Kindle:1989:SNT

[KSFL89] K. L. Kindle, R. A. Schnell, E. Fernández, and P. A. Lefebvre. Stable nuclear transformation of Chlamydomonas using the Chlamydomonas gene for nitrate reductase. Journal of Cell Biology, 109(6):2589–??, December 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/109/6/2589.

Krogstad:1985:AIV

[KSG85] D. J. Krogstad, P. H. Schlesinger, and I. Y. Gluzman. Antimalarials increase vesicle pH in Plasmodium falciparum. Journal of Cell Biology, 101(6):2302–??, December 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/101/6/2302.

Kay:1987:ECB

[KSH87] B. K. Kay, A. J. Shah, and W. E. Halstead. Expression of the Ca^{2+}-binding protein, parvalbumin, during embryonic development of the frog, Xenopus laevis. Journal of Cell Biology, 104 (4):841–??, January 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/104/4/841.

Kartenbeck:1989:ESV

[KSH89] J. Kartenbeck, H. Stukenbrok, and A. Helenius. Endocytosis of simian virus 40 into the endoplasmic reticulum. Journal of Cell Biology, 109(6):2721–??, December 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/109/6/2721.

Kim:1987:RHM

[KSJF87] K. H. Kim, V. Stellmach, J. Javors, and E. Fuchs. Regulation of human mesothelial cell differentiation: opposing roles of retinoids and epidermal growth factor in the expression of intermediate filament proteins. Journal of Cell Biology, 105(6):3039–??, December 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/105/6/
3039.

Kesbeke:1988:STD

[KSJV88] F. Kesbeke, B. E. Snaar-Jagalska, and P. J. Van Haastert. Signal transduction in Dictyostelium fgd A mutants with a defective interaction between surface cAMP receptors and a GTP-binding regulatory protein. Journal of Cell Biology, 107(2):521–??, August 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/107/2/
521.

Kuettel:1985:LNS

[KSKW^{+}85] M. R. Kuettel, S. P. Squinto, J. Kwast-Welfeld, G. Schwoch, J. S. Schweppe, and R. A. Jungmann. Localization of nuclear subunits of cyclic AMP-dependent protein kinase by the immunocolloidal gold method. Journal of Cell Biology, 101(3):965–??, September 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/101/3/
965.

Keene:1987:TVC

[KSL^{+}87] D. R. Keene, L. Y. Sakai, G. P. Lunstrum, N. P. Morris, and R. E. Burgeson. Type VII collagen forms an extended network of anchoring fibrils. Journal of Cell Biology, 104(3):611–??, March 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/104/3/
611.

Knudsen:1989:ICS

[KSM89] K. A. Knudsen, L. Smith, and S. McElwee. Involvement of cell surface phosphatidylinositol-anchored glycoproteins in cell–cell adhesion of chick embryo myoblasts. Journal of Cell Biology, 109 (4):1779–??, October 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/4/1779.

Kreiner:1986:LAN

[KSS86] T. Kreiner, W. Sossin, and R. H. Scheller. Localization of Aplysia neurosecretory peptides to multiple populations of dense core vesicles. Journal of Cell Biology, 102(3):769–??, March 1986. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/102/3/769.

Kelvin:1989:GFSb

[KSSAQC89] D. J. Kelvin, G. Simard, A. Sue-A-Quan, and J. A. Connolly. Growth factors, signaling pathways, and the regulation of proliferation and differentiation in BC3H1 muscle cells. II. Two signaling pathways distinguished by pertussis toxin and a potential role for the ras oncogene. Journal of Cell Biology, 108(1):169–??, January 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/108/1/
169.

Knaack:1986:SEA

[KSSP86] D. Knaack, I. Shen, M. M. Salpeter, and T. R. Podleski. Selective effects of ascorbic acid on acetylcholine receptor number and distribution. Journal of Cell Biology, 102(3):795–??, March 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/102/3/
795.

Kelvin:1989:GFSa

[KST^{+}89] D. J. Kelvin, G. Simard, H. H. Tai, T. P. Yamaguchi, and J. A. Connolly. Growth factors, signaling pathways, and the regulation of proliferation and differentiation in BC3H1 muscle cells. I. A pertussis toxin-sensitive pathway is involved. Journal of Cell Biology, 108(1):159–??, January 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/108/1/159.

Knudson:1985:FMP

[KT85] C. B. Knudson and B. P. Toole. Fluorescent morphological probe for hyaluronate. Journal of Cell Biology, 100(5):1753–??, May 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/100/5/
1753.

King:1987:EMK

[KTF87a] I. A. King, A. Tabiowo, and P. R. Fryer. Evidence that major 78-44-kD concanavalin A-binding glycopolypeptides in pig epidermis arise from the degradation of desmosomal glycoproteins during terminal differentiation. Journal of Cell Biology, 105(6): 3053–??, December 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/6/3053.

Kopan:1987:RIR

[KTF87b] R. Kopan, G. Traska, and E. Fuchs. Retinoids as important regulators of terminal differentiation: examining keratin expression in individual epidermal cells at various stages of keratinization. Journal of Cell Biology, 105(1):427–??, July 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/105/1/427.

Kanai:1989:EMT

[KTO^{+}89] Y. Kanai, R. Takemura, T. Oshima, H. Mori, Y. Ihara, M. Yanagisawa, T. Masaki, and N. Hirokawa. Expression of multiple tau isoforms and microtubule bundle formation in fibroblasts transfected with a single tau cDNA. Journal of Cell Biology, 109(3): 1173–??, September 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/3/1173.

Kuczmarski:1987:EHC

[KTP87] E. R. Kuczmarski, S. R. Tafuri, and L. M. Parysek. Effect of heavy chain phosphorylation on the polymerization and structure of Dictyostelium myosin filaments. Journal of Cell Biology, 105 (6):2989–??, December 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/6/2989.

Komesli:1989:MIC

[KTP^{+}89] S. Komesli, F. Tournier, M. Paintrand, R. L. Margolis, D. Job, and M. Bornens. Mass isolation of calf thymus centrosomes: identification of a specific configuration. Journal of Cell Biology, 109 (6):2869–??, December 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/6/2869.

Kobayashi:1986:SCS

[KTT^{+}86] T. Kobayashi, S. Tsukita, S. Tsukita, Y. Yamamoto, and G. Matsumoto. Subaxolemmal cytoskeleton in squid giant axon. I. Biochemical analysis of microtubules, microfilaments, and their associated high-molecular-weight proteins. Journal of Cell Biology, 102(5):1699–??, May 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/102/5/1699.

Kubai:1987:NCA

[Kub87] D. F. Kubai. Nonrandom chromosome arrangements in germ line nuclei of Sciara coprophila males: the basis for nonrandom chromosome segregation on the meiosis I spindle. Journal of Cell Biology, 105(6):2433–??, December 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/105/6/2433.

Kagan:1986:UIL

[KVB^{+}86] H. M. Kagan, C. A. Vaccaro, R. E. Bronson, S. S. Tang, and J. S. Brody. Ultrastructural immunolocalization of lysyl oxidase in vascular connective tissue. Journal of Cell Biology, 103(3): 1121–??, September 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/103/3/1121.

Kinsella:1986:MSP

[KW86a] M. G. Kinsella and T. N. Wight. Modulation of sulfated proteoglycan synthesis by bovine aortic endothelial cells during migration. Journal of Cell Biology, 102(3):679–??, March 1986. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/102/3/679.

Kmiec:1986:PTF

[KW86b] E. B. Kmiec and A. Worcel. The positive transcription factor of the 5S RNA gene proteolyses during direct exchange between 5S DNA sites. Journal of Cell Biology, 103(3):673–??, September 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/103/3/
673.

King:1988:SGH

[KW88] S. M. King and G. B. Witman. Structure of the gamma heavy chain of the outer arm dynein from Chlamydomonas flagella. Journal of Cell Biology, 107(5):1799–??, November 1988. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/107/5/1799.

Krohne:1989:CCT

[KWH89] G. Krohne, I. Waizenegger, and T. H. Höger. The conserved carboxy-terminal cysteine of nuclear lamins is essential for lamin association with the nuclear envelope. Journal of Cell Biology, 109(5):2003–??, November 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/109/5/2003.

Kamboj:1988:MCB

[KWLS88] R. K. Kamboj, L. M. Wong, T. Y. Lam, and C. H. Siu. Mapping of a cell-binding domain in the cell adhesion molecule gp80 of Dictyostelium discoideum. Journal of Cell Biology, 107(5):

1835–??, November 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/107/5/1835.

Kumar:1989:MIP

[KWPH89] R. V. Kumar, A. Wolfman, R. Panniers, and E. C. Henshaw. Mechanism of inhibition of polypeptide chain initiation in calcium-depleted Ehrlich ascites tumor cells. Journal of Cell Biology, 108(6):2107–??, June 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/108/6/2107.

Kondo:1989:IVD

[KWS^{+}89] K. Kondo, T. Watanabe, H. Sasaki, Y. Uehara, and M. Oishi. Induction of in vitro differentiation of mouse embryonal carcinoma (F9) and erythroleukemia (MEL) cells by herbimycin A, an inhibitor of protein phosphorylation. Journal of Cell Biology, 109(1):285–??, July 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/1/285.

Kaplan:1985:POI

[KWW85] J. Kaplan, D. M. Ward, and H. S. Wiley. Phenylarsine oxideinduced increase in alveolar macrophage surface receptors: evidence for fusion of internal receptor pools with the cell surface. Journal of Cell Biology, 101(1):121–??, July 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/101/1/121.

Kusakabe:1988:NMA

[KYS^{+}88] M. Kusakabe, M. Yokoyama, T. Sakakura, T. Nomura, H. L. Hosick, and Y. Nishizuka. A novel methodology for analysis of cell distribution in chimeric mouse organs using a strain specific antibody. Journal of Cell Biology, 107(1):257–??, July 1988. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/107/1/257.

Lamppa:1987:PWL

[LA87] G. K. Lamppa and M. S. Abad. Processing of a wheat lightharvesting chlorophyll a/b protein precursor by a soluble enzyme from higher plant chloroplasts. Journal of Cell Biology, 105(6): 2641–??, December 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/6/2641.

Linck:1985:LTF

[LAA85] R. W. Linck, L. A. Amos, and W. B. Amos. Localization of tektin filaments in microtubules of sea urchin sperm flagella by immunoelectron microscopy. Journal of Cell Biology, 100(1):126–??, January 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/100/1/
126.

Lothstein:1985:GOP

[LAC^{+}85] L. Lothstein, H. P. Arenstorf, S. Y. Chung, B. W. Walker, J. C. Wooley, and W. M. LeStourgeon. General organization of protein in HeLa 40S nuclear ribonucleoprotein particles. Journal of Cell Biology, 100(5):1570–??, May 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/100/5/1570.

Lefcort:1989:OCE

[LB89] F. Lefcort and D. Bentley. Organization of cytoskeletal elements and organelles preceding growth cone emergence from an identified neuron in situ. Journal of Cell Biology, 108(5):1737–??, May 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/108/5/
1737.

Leube:1988:MCE

[LBB^{+}88] R. E. Leube, B. L. Bader, F. X. Bosch, R. Zimbelmann, T. Achtstaetter, and W. W. Franke. Molecular characterization and expression of the stratification-related cytokeratins 4 and 15. Journal of Cell Biology, 106(4):1249–??, April 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/106/4/1249.

Lotz:1989:CAF

[LBEM89] M. M. Lotz, C. A. Burdsal, H. P. Erickson, and D. R. McClay. Cell adhesion to fibronectin and tenascin: quantitative measurements of initial binding and subsequent strengthening response. Journal of Cell Biology, 109(4):1795–??, October 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/109/4/1795.

Liu:1988:TMI

[LBFS88] X. Q. Liu, A. W. Bell, K. B. Freeman, and G. C. Shore. Topogenesis of mitochondrial inner membrane uncoupling protein. Rerouting transmembrane segments to the soluble matrix compartment. Journal of Cell Biology, 107(2):503–??, August 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/107/2/503.

Lopez:1985:RFM

[LBL^{+}85] L. C. Lopez, E. M. Bayna, D. Litoff, N. L. Shaper, J. H. Shaper, and B. D. Shur. Receptor function of mouse sperm surface galactosyltransferase during fertilization. Journal of Cell Biology, 101 (4):1501–??, October 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/101/4/1501.

Landry:1985:SAC

[LBO^{+}85] J. Landry, D. Bernier, C. Ouellet, R. Goyette, and N. Marceau. Spheroidal aggregate culture of rat liver cells: histotypic reorganization, biomatrix deposition, and maintenance of functional activities. Journal of Cell Biology, 101(3):914–??, September 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/101/3/
914.

Livneh:1986:LDC

[LBP^{+}86] E. Livneh, M. Benveniste, R. Prywes, S. Felder, Z. Kam, and J. Schlessinger. Large deletions in the cytoplasmic kinase domain of the epidermal growth factor receptor do not affect its laternal mobility. Journal of Cell Biology, 103(2):327–??, August 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/103/2/
327.

Lucocq:1986:IOT

[LBR86] J. M. Lucocq, D. Brada, and J. Roth. Immunolocalization of the oligosaccharide trimming enzyme glucosidase II. Journal of Cell Biology, 102(6):2137–??, June 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/102/6/2137.

Lucocq:1989:MGF

[LBW89] J. M. Lucocq, E. G. Berger, and G. Warren. Mitotic Golgi fragments in HeLa cells and their role in the reassembly pathway. Journal of Cell Biology, 109(2):463–??, August 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/109/2/463.

Lewis:1985:GEE

[LC85] S. A. Lewis and N. J. Cowan. Genetics, evolution, and expression of the 68,000-mol-wt neurofilament protein: isolation of a cloned cDNA probe. Journal of Cell Biology, 100(3):843–??, March 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/100/3/
843.

Lopata:1987:VMC

[LC87] M. A. Lopata and D. W. Cleveland. In vivo microtubules are copolymers of available beta-tubulin isotypes: localization of each of six vertebrate beta-tubulin isotypes using polyclonal antibodies elicited by synthetic peptide antigens. Journal of Cell Biology, 105(4):1707–??, October 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/105/4/1707.

Larabell:1988:EMX

[LC88a] C. A. Larabell and D. E. Chandler. The extracellular matrix of Xenopus laevis eggs: a quick-freeze, deep-etch analysis of its modification at fertilization. Journal of Cell Biology, 107 (2):731–??, August 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/107/2/731.

Lewis:1988:CRF

[LC88b] S. A. Lewis and N. J. Cowan. Complex regulation and functional versatility of mammalian alpha- and beta-tubulin isotypes during the differentiation of testis and muscle cells. Journal of Cell Biology, 106(6):2023–??, June 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/106/6/2023.

Larson:1989:PSL

[LCBS89] R. S. Larson, A. L. Corbi, L. Berman, and T. Springer. Primary structure of the leukocyte function-associated molecule-1 alpha subunit: an integrin with an embedded domain defining a protein superfamily. Journal of Cell Biology, 108(2):703–??, February 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/108/2/
703.

Lai:1989:LMA

[LCD^{+}89] W. H. Lai, P. H. Cameron, J. J. Doherty, B. I. Posner, and J. J. Bergeron. Ligand-mediated autophosphorylation activity of the epidermal growth factor receptor during internalization. Journal of Cell Biology, 109(6):2751–??, December 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/109/6/2751.

Lisanti:1989:GMA

[LCDRB89] M. P. Lisanti, I. W. Caras, M. A. Davitz, and E. Rodriguez-Boulan. A glycophospholipid membrane anchor acts as an apical targeting signal in polarized epithelial cells. Journal of Cell Biology, 109(5):2145–??, November 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/109/5/2145.

Levi:1987:ESD

[LCE87] G. Levi, K. L. Crossin, and G. M. Edelman. Expression sequences and distribution of two primary cell adhesion molecules during embryonic development of Xenopus laevis. Journal of Cell Biology, 105(5):2359–??, November 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/105/5/2359.

Levine:1988:AMH

[LCK88] R. J. Levine, P. D. Chantler, and R. W. Kensler. Arrangement of myosin heads on Limulus thick filaments. Journal of Cell Biology, 107(5):1739–??, November 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/107/5/1739.

Landry:1989:HSR

[LCL^{+}89] J. Landry, P. Chrétien, H. Lambert, E. Hickey, and L. A. Weber. Heat shock resistance conferred by expression of the human HSP27 gene in rodent cells. Journal of Cell Biology, 109(1):7–??, July 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/109/1/
7.

Lewandowska:1987:FMA

[LCR^{+}87] K. Lewandowska, H. U. Choi, L. C. Rosenberg, L. Zardi, and L. A. Culp. Fibronectin-mediated adhesion of fibroblasts: inhibition by dermatan sulfate proteoglycan and evidence for a cryptic glycosaminoglycan-binding domain. Journal of Cell Biology, 105 (3):1443–??, September 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/3/1443.

Lai:1989:LMI

[LCW^{+}89] W. H. Lai, P. H. Cameron, I. Wada, J. J. Doherty, D. G. Kay, B. I. Posner, and J. J. Bergeron. Ligand-mediated internalization, recycling, and downregulation of the epidermal growth factor receptor in vivo. Journal of Cell Biology, 109(6): 2741–??, December 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/6/2741.

Lipp:1986:SRP

[LD86] J. Lipp and B. Dobberstein. Signal recognition particledependent membrane insertion of mouse invariant chain: a membrane-spanning protein with a cytoplasmically exposed amino terminus. Journal of Cell Biology, 102(6):2169–??, June 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/102/6/
2169.

Lipp:1988:SMA

[LD88] J. Lipp and B. Dobberstein. Signal and membrane anchor functions overlap in the type II membrane protein I gamma CAT. Journal of Cell Biology, 106(6):1813–??, June 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/106/6/1813.

Larkin:1986:PDA

[LDA86] J. M. Larkin, W. C. Donzell, and R. G. Anderson. Potassiumdependent assembly of coated pits: new coated pits form as planar clathrin lattices. Journal of Cell Biology, 103(6):2619–??, December 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
103/6/2619.

Lazarovici:1987:LTH

[LDKG87] P. Lazarovici, G. Dickens, H. Kuzuya, and G. Guroff. Long-term, heterologous down-regulation of the epidermal growth factor receptor in PC12 cells by nerve growth factor. Journal of Cell Biology, 104(6):1611–??, June 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/104/6/1611.

Liu:1987:VHL

[LDP87] S. C. Liu, L. H. Derick, and J. Palek. Visualization of the hexagonal lattice in the erythrocyte membrane skeleton. Journal of Cell Biology, 104(3):527–??, March 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/104/3/527.

Leivo:1986:CFC

[LE86] I. Leivo and E. Engvall. C3d fragment of complement interacts with laminin and binds to basement membranes of glomerulus and trophoblast. Journal of Cell Biology, 103(3): 1091–??, September 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/103/3/1091.

Lin:1987:SDM

[LEFS^{+}87] Z. X. Lin, J. R. Eshelman, S. Forry-Schaudies, S. Duran, J. L. Lessard, and H. Holtzer. Sequential disassembly of myofibrils induced by myristate acetate in cultured myotubes. Journal of Cell Biology, 105(3):1365–??, September 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/105/3/1365.

Lin:1989:DRM

[LEG^{+}89] Z. X. Lin, J. Eshleman, C. Grund, D. A. Fischman, T. Masaki, W. W. Franke, and H. Holtzer. Differential response of myofibrillar and cytoskeletal proteins in cells treated with phorbol myristate acetate. Journal of Cell Biology, 108(3):1079–??, March 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/108/3/
1079.

LeBaron:1988:AGD

[LEW^{+}88] R. G. LeBaron, J. D. Esko, A. Woods, S. Johansson, and M. Höök. Adhesion of glycosaminoglycan-deficient chinese hamster ovary cell mutants to fibronectin substrata. Journal of Cell Biology, 106(3):945–??, March 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/106/3/945.

Lamb:1988:RAM

[LFC^{+}88] N. J. Lamb, A. Fernandez, M. A. Conti, R. Adelstein, D. B. Glass, W. J. Welch, and J. R. Feramisco. Regulation of actin microfilament integrity in living nonmuscle cells by the cAMPdependent protein kinase and the myosin light chain kinase. Journal of Cell Biology, 106(6):1955–??, June 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/106/6/1955.

Lee:1989:CER

[LFC89] C. Lee, M. Ferguson, and L. B. Chen. Construction of the endoplasmic reticulum. Journal of Cell Biology, 109(5):2045– ??, November 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/5/2045.

Lamb:1989:MVC

[LFFW89] N. J. Lamb, A. Fernandez, J. R. Feramisco, and W. J. Welch. Modulation of vimentin containing intermediate filament distribution and phosphorylation in living fibroblasts by the cAMPdependent protein kinase. Journal of Cell Biology, 108(6):2409– ??, June 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
108/6/2409.

Lipp:1989:SRM

[LFHD89] J. Lipp, N. Flint, M. T. Haeuptle, and B. Dobberstein. Structural requirements for membrane assembly of proteins spanning the membrane several times. Journal of Cell Biology, 109(5): 2013–??, November 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/5/2013.

Lander:1985:PFP

[LFR85] A. D. Lander, D. K. Fujii, and L. F. Reichardt. Purification of a factor that promotes neurite outgrowth: isolation of laminin and associated molecules. Journal of Cell Biology, 101 (3):898–??, September 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/101/3/898.

LeBlanc:1989:AGA

[LFS^{+}89] A. C. LeBlanc, M. Földvári, D. F. Spencer, W. C. Breckenridge, R. G. Fenwick, D. L. Williams, and C. Mezei. The apolipoprotein A–I gene is actively expressed in the rapidly myelinating avian peripheral nerve. Journal of Cell Biology, 109(3): 1245–??, September 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/3/1245.

Laliberte:1988:PAP

[LG88] A. Laliberte and C. Gicquaud. Polymerization of actin by positively charged liposomes. Journal of Cell Biology, 106(4):1221– ??, April 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
106/4/1221.

Ludwig:1989:LPL

[LG89] M. Ludwig and S. P. Gibbs. Localization of phycoerythrin at the lumenal surface of the thylakoid membrane in Rhodomonas lens. Journal of Cell Biology, 108(3):875–??, March 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/108/3/875.

Lightner:1989:THH

[LGBE89] V. A. Lightner, F. Gumkowski, D. D. Bigner, and H. P. Erickson. Tenascin/ hexabrachion in human skin: biochemical identification and localization by light and electron microscopy. Journal of Cell Biology, 108(6):2483–??, June 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/108/6/2483.

Leonard:1988:NGF

[LGC^{+}88] D. G. Leonard, J. D. Gorham, P. Cole, L. A. Greene, and E. B. Ziff. A nerve growth factor-regulated messenger RNA encodes a new intermediate filament protein. Journal of Cell Biology, 106 (1):181–??, January 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/106/1/181.

Linsenmayer:1986:EAC

[LGF86] T. F. Linsenmayer, E. Gibney, and J. M. Fitch. Embryonic avian cornea contains layers of collagen with greater than average stability. Journal of Cell Biology, 103(4):1587–??, October 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/103/4/
1587.

Lewis:1985:GLM

[LGM^{+}85] V. Lewis, S. A. Green, M. Marsh, P. Vihko, A. Helenius, and I. Mellman. Glycoproteins of the lysosomal membrane. Journal of Cell Biology, 100(6):1839–??, June 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/100/6/1839.

Languino:1989:ECU

[LGW^{+}89] L. R. Languino, K. R. Gehlsen, E. Wayner, W. G. Carter, E. Engvall, and E. Ruoslahti. Endothelial cells use alpha 2 beta 1 integrin as a laminin receptor. Journal of Cell Biology, 109(5): 2455–??, November 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/5/2455.

Lawler:1986:SHT

[LH86] J. Lawler and R. O. Hynes. The structure of human thrombospondin, an adhesive glycoprotein with multiple calciumbinding sites and homologies with several different proteins. Journal of Cell Biology, 103(5):1635–??, November 1986. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/103/5/1635.

Lew:1988:MSI

[LHFB88] V. L. Lew, A. Hockaday, C. J. Freeman, and R. M. Bookchin. Mechanism of spontaneous inside-out vesiculation of red cell membranes. Journal of Cell Biology, 106(6):1893–??, June 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/106/6/1893.

Lin:1985:TIC

[LHHC85] J. J. Lin, D. M. Helfman, S. H. Hughes, and C. S. Chou. Tropomyosin isoforms in chicken embryo fibroblasts: purification, characterization, and changes in Rous sarcoma virustransformed cells. Journal of Cell Biology, 100(3):692–??, March 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/100/3/
692.

Laurie:1989:SHR

[LHK^{+}89] G. W. Laurie, S. Horikoshi, P. D. Killen, B. Segui-Real, and Y. Yamada. In situ hybridization reveals temporal and spatial changes in cellular expression of mRNA for a laminin receptor, laminin, and basement membrane (type IV) collagen in the developing kidney. Journal of Cell Biology, 109(3):1351–??, September 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/109/3/
1351.

Lin:1988:DLT

[LHL88] J. J. Lin, T. E. Hegmann, and J. L. Lin. Differential localization of tropomyosin isoforms in cultured nonmuscle cells. Journal of Cell Biology, 107(2):563–??, August 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/107/2/563.

Li:1989:NND

[LHM89] J. M. Li, A. K. Hopper, and N. C. Martin. N2,N2dimethylguanosine-specific tRNA methyltransferase contains both nuclear and mitochondrial targeting signals in Saccharomyces cerevisiae. Journal of Cell Biology, 109(4):1411–??, October 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
109/4/1411.

Lieber:1987:MOT

[LHNF87] M. R. Lieber, J. E. Hesse, J. M. Nickol, and G. Felsenfeld. The mechanism of osmotic transfection of avian embryonic erythrocytes: analysis of a system for studying developmental gene expression. Journal of Cell Biology, 105(3):1055–??, September 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/105/3/
1055.

Lin:1989:PAP

[LHS^{+}89] Z. X. Lin, S. Holtzer, T. Schultheiss, J. Murray, T. Masaki, D. A. Fischman, and H. Holtzer. Polygons and adhesion plaques and the disassembly and assembly of myofibrils in cardiac myocytes. Journal of Cell Biology, 108(6):2355–??, June 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/108/6/2355.

Lemansky:1987:LEP

[LHvF^{+}87] P. Lemansky, A. Hasilik, K. von Figura, S. Helmy, J. Fishman, R. E. Fine, N. L. Kedersha, and L. H. Rome. Lysosomal enzyme precursors in coated vesicles derived from the exocytic and endocytic pathways. Journal of Cell Biology, 104(6):1743–??, June 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/104/6/
1743.

Lind:1987:RBA

[LJC^{+}87] S. E. Lind, P. A. Janmey, C. Chaponnier, T. J. Herbert, and T. P. Stossel. Reversible binding of actin to gelsolin and profilin in human platelet extracts. Journal of Cell Biology, 105 (2):833–??, August 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/2/833.

Longo:1987:BPP

[LKF87] F. J. Longo, G. Krohne, and W. W. Franke. Basic proteins of the perinuclear theca of mammalian spermatozoa and spermatids: a novel class of cytoskeletal elements. Journal of Cell Biology, 105 (3):1105–??, September 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/3/1105.

Lodish:1987:VIT

[LKHR87] H. F. Lodish, N. Kong, S. Hirani, and J. Rasmussen. A vesicular intermediate in the transport of hepatoma secretory proteins from the rough endoplasmic reticulum to the Golgi complex. Journal of Cell Biology, 104(2):221–??, February 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/104/2/221.

Lyons:1988:PAL

[LKOM88] R. M. Lyons, J. Keski-Oja, and H. L. Moses. Proteolytic activation of latent transforming growth factor-beta from fibroblastconditioned medium. Journal of Cell Biology, 106(5):1659–??, May 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/106/5/

Lin:1986:ADI

[LL86] J. J. Lin and J. L. Lin. Assembly of different isoforms of actin and tropomyosin into the skeletal tropomyosin-enriched microfilaments during differentiation of muscle cells in vitro. Journal of Cell Biology, 103(6):2173–??, December 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/103/6/2173.

Lehnert:1988:USD

[LL88] M. E. Lehnert and H. F. Lodish. Unequal synthesis and differential degradation of alpha and beta spectrin during murine erythroid differentiation. Journal of Cell Biology, 107(2):413–??, August 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
107/2/413.

Lankford:1989:ECM

[LL89a] K. L. Lankford and P. C. Letourneau. Evidence that calcium may control neurite outgrowth by regulating the stability of actin filaments. Journal of Cell Biology, 109(3):1229–??, September 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/109/3/
1229.

Lourim:1989:ENL

[LL89b] D. Lourim and J. J. Lin. Expression of nuclear lamin A and muscle-specific proteins in differentiating muscle cells in ovo and in vitro. Journal of Cell Biology, 109(2):495–??, August 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/109/2/495.

Lancillotti:1985:LZD

[LLAS85] F. Lancillotti, M. C. Lopez, C. Alonso, and B. D. Stollar. Locations of Z–DNA in polytene chromosomes. Journal of Cell Biology, 100(5):1759–??, May 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/100/5/1759.

Lewis:1985:FMT

[LLC85] S. A. Lewis, M. G. Lee, and N. J. Cowan. Five mouse tubulin isotypes and their regulated expression during development.

Journal of Cell Biology, 101(3):852–??, September 1985. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/101/3/852.

Lin:1989:ASA

[LLCA89] R. Lin, J. W. Leone, R. G. Cook, and C. D. Allis. Antibodies specific to acetylated histones document the existence of depositionand transcription-related histone acetylation in Tetrahymena. Journal of Cell Biology, 108(5):1577–??, May 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/108/5/1577.

Lebel:1987:LCA

[LLRR87] S. Lebel, C. Lampron, A. Royal, and Y. Raymond. Lamins A and C appear during retinoic acid-induced differentiation of mouse embryonal carcinoma cells. Journal of Cell Biology, 105 (3):1099–??, September 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/3/1099.

Lisanti:1989:SSD

[LLSRB89] M. P. Lisanti, A. Le Bivic, M. Sargiacomo, and E. Rodriguez-Boulan. Steady-state distribution and biogenesis of endogenous Madin–Darby canine kidney glycoproteins: evidence for intracellular sorting and polarized cell surface delivery. Journal of Cell Biology, 109(5):2117–??, November 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/109/5/2117.

Lohka:1985:INE

[LM85a] M. J. Lohka and J. L. Maller. Induction of nuclear envelope breakdown, chromosome condensation, and spindle formation in cell-free extracts. Journal of Cell Biology, 101(2):518–??, August 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/101/2/
518.

Lyon:1985:CLH

[LM85b] M. K. Lyon and K. R. Miller. Crystallization of the lightharvesting chlorophyll a/b complex within thylakoid membranes. Journal of Cell Biology, 100(4):1139–??, April 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/100/4/1139.

Leser:1987:CHN

[LM87] G. P. Leser and T. E. Martin. Changes in heterogeneous nuclear RNP core polypeptide complements during the cell cycle. Journal of Cell Biology, 105(5):2083–??, November 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/105/5/2083.

Langanger:1986:MOM

[LMD^{+}86] G. Langanger, M. Moeremans, G. Daneels, A. Sobieszek, M. De Brabander, and J. De Mey. The molecular organization of myosin in stress fibers of cultured cells. Journal of Cell Biology, 102 (1):200–??, January 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/102/1/200.

Lombardi:1986:EDF

[LMF^{+}86] T. Lombardi, R. Montesano, M. B. Furie, S. C. Silverstein, and L. Orci. Endothelial diaphragmed fenestrae: in vitro modulation by phorbol myristate acetate. Journal of Cell Biology, 102(5):1965–??, May 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/102/5/1965.

Lowe:1988:ESG

[LMK88] A. W. Lowe, L. Madeddu, and R. B. Kelly. Endocrine secretory granules and neuronal synaptic vesicles have three integral membrane proteins in common. Journal of Cell Biology, 106 (1):51–??, January 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/106/1/51.

Lang:1989:SDC

[LMSM89] E. Lang, M. L. Mazauric-Stüker, and A. Maelicke. States of developmental commitment of a mouse embryonal carcinoma cell line differentiating along a neural pathway. Journal of Cell Biology, 109(5):2481–??, November 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/109/5/2481.

Lew:1986:QAC

[LMW^{+}86] P. D. Lew, A. Monod, F. A. Waldvogel, B. Dewald, M. Baggiolini, and T. Pozzan. Quantitative analysis of the cytosolic free calcium dependency of exocytosis from three subcellular compartments in intact human neutrophils. Journal of Cell Biology, 102(6):2197–??, June 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/102/6/2197.

Logsdon:1985:GIA

[LMWG85] C. D. Logsdon, J. Moessner, J. A. Williams, and I. D. Goldfine. Glucocorticoids increase amylase mRNA levels, secretory organelles, and secretion in pancreatic acinar AR42J cells. Journal of Cell Biology, 100(4):1200–??, April 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/100/4/1200.

Lewis:1988:MPV

[LN88] S. E. Lewis and R. A. Nixon. Multiple phosphorylated variants of the high molecular mass subunit of neurofilaments in axons of retinal cell neurons: characterization and evidence for their differential association with stationary and moving neurofilaments. Journal of Cell Biology, 107(6):2689–??, December 1988. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/107/6/2689.

Lica:1986:MSD

[LNH86] L. M. Lica, S. Narayanswami, and B. A. Hamkalo. Mouse satellite DNA, centromere structure, and sister chromatid pairing. Journal of Cell Biology, 103(4):1145–??, October 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/103/4/1145.

Lathrop:1985:CFG

[LOG85] B. Lathrop, E. Olson, and L. Glaser. Control by fibroblast growth factor of differentiation in the BC3H1 muscle cell line. Journal of Cell Biology, 100(5):1540–??, May 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/100/5/1540.

Lynch:1986:ABH

[LOH^{+}86] M. H. Lynch, W. M. O’Guin, C. Hardy, L. Mak, and T. T. Sun. Acidic and basic hair/nail (”hard”) keratins: their colocalization in upper cortical and cuticle cells of the human hair follicle and their relationship to “soft” keratins. Journal of Cell Biology, 103 (6):2593–??, December 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/103/6/2593.

Lipsky:1985:ITF

[LP85a] N. G. Lipsky and R. E. Pagano. Intracellular translocation of fluorescent sphingolipids in cultured fibroblasts: endogenously synthesized sphingomyelin and glucocerebroside analogues pass through the Golgi apparatus en route to the plasma membrane. Journal of Cell Biology, 100(1):27–??, January 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/100/1/27.

Luther:1985:MRS

[LP85b] P. W. Luther and H. B. Peng. Membrane-related specializations associated with acetylcholine receptor aggregates induced by electric fields. Journal of Cell Biology, 100(1):235–??, January 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/100/1/
235.

LeDizet:1986:CMC

[LP86] M. LeDizet and G. Piperno. Cytoplasmic microtubules containing acetylated alpha-tubulin in Chlamydomonas reinhardtii: spatial arrangement and properties. Journal of Cell Biology, 103(1):13–??, July 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/103/1/13.

Lucocq:1987:MFG

[LPBW87] J. M. Lucocq, J. G. Pryde, E. G. Berger, and G. Warren. A mitotic form of the Golgi apparatus in HeLa cells. Journal of Cell Biology, 104(4):865–??, January 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/104/4/865.

Luby-Phelps:1985:BFA

[LPLT85] K. Luby-Phelps, F. Lanni, and D. L. Taylor. Behavior of a fluorescent analogue of calmodulin in living 3T3 cells. Journal of Cell Biology, 101(4):1245–??, October 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/101/4/1245.

Leach:1989:TPK

[LPR^{+}89] K. L. Leach, E. A. Powers, V. A. Ruff, S. Jaken, and S. Kaufmann. Type 3 protein kinase C localization to the nuclear envelope of phorbol ester-treated NIH 3T3 cells. Journal of Cell Biology, 109(2):685–??, August 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/109/2/685.

Luby-Phelps:1986:PSC

[LPTL86] K. Luby-Phelps, D. L. Taylor, and F. Lanni. Probing the structure of cytoplasm. Journal of Cell Biology, 102(6):2015–??, June 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/102/6/
2015.

Landreth:1985:NGF

[LR85a] G. E. Landreth and G. D. Rieser. Nerve growth factor- and epidermal growth factor-stimulated phosphorylation of a PC12 cytoskeletally associated protein in situ. Journal of Cell Biology, 100(3):677–??, March 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/100/3/677.

LHernault:1985:RPM

[LR85b] S. W. L’Hernault and J. L. Rosenbaum. Reversal of the posttranslational modification on Chlamydomonas flagellar alphatubulin occurs during flagellar resorption. Journal of Cell Biology, 100(2):457–??, February 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/100/2/457.

Luca:1989:CPC

[LR89] F. C. Luca and J. V. Ruderman. Control of programmed cyclin destruction in a cell-free system. Journal of Cell Biology, 109(5): 1895–??, November 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/5/1895.

Lai:1988:ATG

[LRF88] E. Y. Lai, S. P. Remillard, and C. Fulton. The alpha-tubulin gene family expressed during cell differentiation in Naegleria gruberi. Journal of Cell Biology, 106(6):2035–??, June 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/106/6/2035.

Liscum:1989:ITL

[LRF89] L. Liscum, R. M. Ruggiero, and J. R. Faust. The intracellular transport of low density lipoprotein-derived cholesterol is defective in Niemann–Pick type C fibroblasts. Journal of Cell Biology, 108(5):1625–??, May 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/108/5/1625.

Levine:1985:CCI

[LS85] J. F. Levine and F. E. Stockdale. Cell–cell interactions promote mammary epithelial cell differentiation. Journal of Cell Biology, 100(5):1415–??, May 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/100/5/1415.

Levin:1987:APA

[LS87a] E. G. Levin and L. Santell. Association of a plasminogen activator inhibitor (PAI-1) with the growth substratum and membrane of human endothelial cells. Journal of Cell Biology, 105 (6):2543–??, December 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/6/2543.

Linck:1987:BCT

[LS87b] R. W. Linck and R. E. Stephens. Biochemical characterization of tektins from sperm flagellar doublet microtubules. Journal of Cell Biology, 104(4):1069–??, January 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/104/4/1069.

Lopez:1987:RMS

[LS87c] L. C. Lopez and B. D. Shur. Redistribution of mouse sperm surface galactosyltransferase after the acrosome reaction. Journal of Cell Biology, 105(4):1663–??, October 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/105/4/1663.

Leyton:1989:EAM

[LS89] L. Leyton and P. Saling. Evidence that aggregation of mouse sperm receptors by ZP3 triggers the acrosome reaction. Journal of Cell Biology, 108(6):2163–??, June 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/108/6/2163.

Laiho:1986:EPE

[LSAKO86] M. Laiho, O. Saksela, P. A. Andreasen, and J. Keski-Oja. Enhanced production and extracellular deposition of the endothelial-type plasminogen activator inhibitor in cultured human lung fibroblasts by transforming growth factor-beta. Journal of Cell Biology, 103(6):2403–??, December 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/103/6/2403.

Lim:1989:PSD

[LSB89] S. S. Lim, P. J. Sammak, and G. G. Borisy. Progressive and spatially differentiated stability of microtubules in developing neuronal cells. Journal of Cell Biology, 109(1):253–??, July 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/109/1/253.

Lewis:1986:CCE

[LSC86] S. A. Lewis, P. Sherline, and N. J. Cowan. A cloned cDNA encoding MAP1 detects a single copy gene in mouse and a brain-abundant RNA whose level decreases during development. Journal of Cell Biology, 102(6):2106–??, June 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/102/6/2106.

Lee:1986:IVG

[LSDP86] S. C. Lee, D. E. Sabath, C. Deutsch, and M. B. Prystowsky. Increased voltage-gated potassium conductance during interleukin 2-stimulated proliferation of a mouse helper T lymphocyte clone. Journal of Cell Biology, 102(4):1200–??, April 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/102/4/1200.

Lehner:1987:DEN

[LSEN87] C. F. Lehner, R. Stick, H. M. Eppenberger, and E. A. Nigg. Differential expression of nuclear lamin proteins during chicken development. Journal of Cell Biology, 105(1):577–??, July 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/105/1/577.

Lippincott-Schwartz:1986:LMD

[LSF86] J. Lippincott-Schwartz and D. M. Fambrough. Lysosomal membrane dynamics: structure and interorganellar movement of a major lysosomal membrane glycoprotein. Journal of Cell Biology, 102(5):1593–??, May 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/102/5/1593.

Lenhard:1989:DDC

[LSF89] J. M. Lenhard, A. Siegel, and S. J. Free. Developing Dictyostelium cells contain the lysosomal enzyme alpha-mannosidase in a secretory granule. Journal of Cell Biology, 109(6):2761– ??, December 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/6/2761.

Lang:1986:MMN

[LSP86] I. Lang, M. Scholz, and R. Peters. Molecular mobility and nucleocytoplasmic flux in hepatoma cells. Journal of Cell Biology, 102(4):1183–??, April 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/102/4/1183.

Li:1989:IHP

[LT89] R. H. Li and J. O. Thomas. Identification of a human protein that interacts with nuclear localization signals. Journal of Cell Biology, 109(6):2623–??, December 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/109/6/2623.

Lathrop:1985:CMD

[LTG85] B. Lathrop, K. Thomas, and L. Glaser. Control of myogenic differentiation by fibroblast growth factor is mediated by position in the G1 phase of the cell cycle. Journal of Cell Biology, 101 (6):2194–??, December 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/101/6/2194.

Lacy:1987:HRA

[LU87] B. E. Lacy and C. B. Underhill. The hyaluronate receptor is associated with actin filaments. Journal of Cell Biology, 105(3): 1395–??, September 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/3/1395.

Lyon:1988:TDS

[LU88] M. K. Lyon and P. N. Unwin. Two-dimensional structure of the light-harvesting chlorophyll a/b complex by cryoelectron microscopy. Journal of Cell Biology, 106(5):1515–??, May 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/106/5/1515.

Lewis:1986:BSE

[LVSC86] S. A. Lewis, A. Villasante, P. Sherline, and N. J. Cowan. Brainspecific expression of MAP2 detected using a cloned cDNA probe. Journal of Cell Biology, 102(6):2098–??, June 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/102/6/2098.

Last:1986:INL

[LW86] R. L. Last and J. L. Woolford. Identification and nuclear localization of yeast pre-messenger RNA processing components: RNA2 and RNA3 proteins. Journal of Cell Biology, 103(6): 2103–??, December 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/103/6/2103.

Lawler:1988:CAT

[LWH88] J. Lawler, R. Weinstein, and R. O. Hynes. Cell attachment to thrombospondin: the role of ARG–GLY–ASP, calcium, and integrin receptors. Journal of Cell Biology, 107(6):2351–??, December 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/107/6/
2351.

Landreth:1985:WGA

[LWM85] G. E. Landreth, L. K. Williams, and C. McCutchen. Wheat germ agglutinin blocks the biological effects of nerve growth factor. Journal of Cell Biology, 101(5):1690–??, November 1985. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/101/5/1690.

Lieberman:1988:ILR

[LWM^{+}88] S. J. Lieberman, W. Wasco, J. MacLeod, P. Satir, and G. A. Orr. Immunogold localization of the regulatory subunit of a type II cAMP-dependent protein kinase tightly associated with mammalian sperm flagella. Journal of Cell Biology, 107(5): 1809–??, November 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/107/5/1809.

Low:1985:SCC

[LWMM^{+}85] R. B. Low, J. Woodcock-Mitchell, J. J. Mitchell, J. Arnold, and P. M. Absher. Synthesis of cytoskeletal and contractile proteins by cultured IMR-90 fibroblasts. Journal of Cell Biology, 101 (2):500–??, August 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/101/2/500.

Landreth:1985:AEG

[LWR85] G. E. Landreth, L. K. Williams, and G. D. Rieser. Association of the epidermal growth factor receptor kinase with the detergentinsoluble cytoskeleton of A431 cells. Journal of Cell Biology, 101 (4):1341–??, October 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/101/4/1341.

LaRochelle:1985:IDA

[LWSF85] W. J. LaRochelle, B. E. Wray, R. Sealock, and S. C. Froehner. Immunochemical demonstration that amino acids 360–377 of the acetylcholine receptor gamma-subunit are cytoplasmic. Journal of Cell Biology, 100(3):684–??, March 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/100/3/684.

Lanni:1985:SOI

[LWT85] F. Lanni, A. S. Waggoner, and D. L. Taylor. Structural organization of interphase 3T3 fibroblasts studied by total internal reflection fluorescence microscopy. Journal of Cell Biology, 100 (4):1091–??, April 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/100/4/1091.

Liu:1987:RBL

[LYE87] Z. Y. Liu, J. I. Young, and E. L. Elson. Rat basophilic leukemia cells stiffen when they secrete. Journal of Cell Biology, 105(6): 2933–??, December 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/6/2933.

Lieska:1985:PIF

[LYG85] N. Lieska, H. Y. Yang, and R. D. Goldman. Purification of the 300K intermediate filament-associated protein and its in vitro recombination with intermediate filaments. Journal of Cell Biology, 101(3):802–??, September 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/101/3/802.

McEwen:1985:PRH

[MA85] B. F. McEwen and W. J. Arion. Permeabilization of rat hepatocytes with Staphylococcus aureus alpha-toxin. Journal of Cell Biology, 100(6):1922–??, June 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/100/6/1922.

Myrdal:1986:AAP

[MA86] S. E. Myrdal and N. Auersperg. An agent or agents produced by virus-transformed cells cause unregulated ruffling in untransformed cells. Journal of Cell Biology, 102(4):1224–??, April 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/102/4/1224.

Moreau:1986:ANT

[MABJ^{+}86] N. Moreau, N. Angelier, M. L. Bonnanfant-Jais, P. Gounon, and P. Kubisz. Association of nucleoplasmin with transcription products as revealed by immunolocalization in the amphibian oocyte. Journal of Cell Biology, 103(3):683–??, September 1986. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/103/3/683.

MacGlashan:1989:SCA

[Mac89a] D. MacGlashan. Single-cell analysis of Ca++ changes in human lung mast cells: graded vs. all-or-nothing elevations after IgEmediated stimulation. Journal of Cell Biology, 109(1):123–??, July 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/109/1/
123.

McCarthy:1989:ICB

[MAC89b] K. J. McCarthy, M. A. Accavitti, and J. R. Couchman. Immunological characterization of a basement membrane-specific chondroitin sulfate proteoglycan. Journal of Cell Biology, 109 (6):3187–??, December 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/6/3187.

Mar:1988:AUR

[MACO88] J. H. Mar, P. B. Antin, T. A. Cooper, and C. P. Ordahl. Analysis of the upstream regions governing expression of the chicken cardiac troponin T gene in embryonic cardiac and skeletal muscle cells. Journal of Cell Biology, 107(2):573–??, August 1988. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/107/2/573.

Mugrauer:1988:MPO

[MAE88] G. Mugrauer, F. W. Alt, and P. Ekblom. N-myc proto-oncogene expression during organogenesis in the developing mouse as revealed by in situ hybridization. Journal of Cell Biology, 107 (4):1325–??, October 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/107/4/1325.

Meyer:1988:ICM

[MAG88] D. J. Meyer, C. L. Afonso, and D. W. Galbraith. Isolation and characterization of monoclonal antibodies directed against plant plasma membrane and cell wall epitopes: identification of a monoclonal antibody that recognizes extensin and analysis of the process of epitope biosynthesis in plant tissues and cell cultures. Journal of Cell Biology, 107(1):163–??, July 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/107/1/163.

Majack:1987:BTT

[Maj87] R. A. Majack. Beta-type transforming growth factor specifies organizational behavior in vascular smooth muscle cell cultures. Journal of Cell Biology, 105(1):465–??, July 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/105/1/465.

Martinuzzi:1986:EMG

[MAK^{+}86] A. Martinuzzi, V. Askanas, T. Kobayashi, W. K. Engel, and S. Di Mauro. Expression of muscle-gene-specific isozymes of phosphorylase and creatine kinase in innervated cultured human muscle. Journal of Cell Biology, 103(4):1423–??, October 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/103/4/1423.

Massague:1985:TGF

[Mas85] J. Massagué. Transforming growth factor-beta modulates the high-affinity receptors for epidermal growth factor and transforming growth factor-alpha. Journal of Cell Biology, 100 (5):1508–??, May 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/100/5/1508.

Minden:1989:DCL

[MASA89] J. S. Minden, D. A. Agard, J. W. Sedat, and B. M. Alberts. Direct cell lineage analysis in Drosophila melanogaster by timelapse, three-dimensional optical microscopy of living embryos. Journal of Cell Biology, 109(2):505–??, August 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/109/2/505.

Matlin:1986:SPP

[Mat86] K. S. Matlin. The sorting of proteins to the plasma membrane in epithelial cells. Journal of Cell Biology, 103(6):2565–??, December 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/103/6/
2565.

May:1989:HDB

[May89] G. S. May. The highly divergent beta-tubulins of Aspergillus nidulans are functionally interchangeable. Journal of Cell Biology, 109(5):2267–??, November 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/109/5/2267.

Majack:1985:HRC

[MB85a] R. A. Majack and P. Bornstein. Heparin regulates the collagen phenotype of vascular smooth muscle cells: induced synthesis of an Mr 60,000 collagen. Journal of Cell Biology, 100 (2):613–??, February 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/100/2/613.

Morgan:1985:NPB

[MB85b] D. A. Morgan and I. Brodsky. Novel peripheral blood-derived human cell lines with properties of megakaryocytes. Journal of Cell Biology, 100(2):565–??, February 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/100/2/565.

Musil:1987:CMS

[MB87] L. S. Musil and J. U. Baenziger. Cleavage of membrane secretory component to soluble secretory component occurs on the cell surface of rat hepatocyte monolayers. Journal of Cell Biology, 104(6):1725–??, June 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/104/6/1725.

Matzuk:1988:RAL

[MB88] M. M. Matzuk and I. Boime. The role of the asparagine-linked oligosaccharides of the alpha subunit in the secretion and assembly of human chorionic gonadotrophin. Journal of Cell Biology, 106(4):1049–??, April 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/106/4/1049.

Madara:1986:ECD

[MBC86] J. L. Madara, D. Barenberg, and S. Carlson. Effects of cytochalasin D on occluding junctions of intestinal absorptive cells: further evidence that the cytoskeleton may influence paracellular permeability and junctional charge selectivity. Journal of Cell Biology, 102(6):2125–??, June 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/102/6/2125.

Mason:1989:CGC

[MBD89] P. W. Mason, J. W. Bigbee, and G. H. DeVries. Cerebellar granule cells contain a membrane mitogen for cultured Schwann cells. Journal of Cell Biology, 108(2):607–??, February 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/108/2/607.

Mostov:1987:AMF

[MBH87] K. E. Mostov, P. Breitfeld, and J. M. Harris. An anchor-minus form of the polymeric immunoglobulin receptor is secreted predominantly apically in Madin–Darby canine kidney cells. Journal of Cell Biology, 105(5):2031–??, November 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/105/5/2031.

Miller:1986:LEG

[MBK^{+}86] K. Miller, J. Beardmore, H. Kanety, J. Schlessinger, and C. R. Hopkins. Localization of the epidermal growth factor (EGF) receptor within the endosome of EGF-stimulated epidermoid carcinoma (A431) cells. Journal of Cell Biology, 102(2):500–??, February 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/102/2/
500.

Miyata:1989:PMA

[MBK89] H. Miyata, B. Bowers, and E. D. Korn. Plasma membrane association of Acanthamoeba myosin I. Journal of Cell Biology, 109 (4):1519–??, October 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/4/1519.

Meda:1986:BCC

[MBKO86] P. Meda, R. Bruzzone, S. Knodel, and L. Orci. Blockage of cell-to-cell communication within pancreatic acini is associated with increased basal release of amylase. Journal of Cell Biology, 103(2):475–??, August 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/103/2/475.

Marazzi:1989:MMA

[MBKR89] G. Marazzi, F. Bard, M. W. Klymkowsky, and L. L. Rubin. Microinjection of a monoclonal antibody against a 37-kD protein (tropomyosin 2) prevents the formation of new acetylcholine receptor clusters. Journal of Cell Biology, 109(5):2337–??, November 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/109/5/
2337.

Mehta:1989:ARC

[MBL89] P. P. Mehta, J. S. Bertram, and W. R. Loewenstein. The actions of retinoids on cellular growth correlate with their actions on gap junctional communication. Journal of Cell Biology, 108 (3):1053–??, March 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/108/3/1053.

Matsuuchi:1988:TSV

[MBLK88] L. Matsuuchi, K. M. Buckley, A. W. Lowe, and R. B. Kelly. Targeting of secretory vesicles to cytoplasmic domains in AtT-20 and PC-12 cells. Journal of Cell Biology, 106(2):239–??, February 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/106/2/
239.

Martin:1987:RNR

[MBLS87] T. E. Martin, S. C. Barghusen, G. P. Leser, and P. G. Spear. Redistribution of nuclear ribonucleoprotein antigens during herpes simplex virus infection. Journal of Cell Biology, 105(5): 2069–??, November 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/5/2069.

Maruyama:1985:TFE

[MBM85] I. Maruyama, C. E. Bell, and P. W. Majerus. Thrombomodulin is found on endothelium of arteries, veins, capillaries, and lymphatics, and on syncytiotrophoblast of human placenta. Journal of Cell Biology, 101(2):363–??, August 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/101/2/363.

Murthy:1986:PLR

[MBSMD86] U. Murthy, M. Basu, A. Sen-Majumdar, and M. Das. Perinuclear location and recycling of epidermal growth factor receptor kinase: immunofluorescent visualization using antibodies directed to kinase and extracellular domains. Journal of Cell Biology, 103(2):333–??, August 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/103/2/333.

Mullock:1989:REL

[MBvS^{+}89] B. M. Mullock, W. J. Branch, M. van Schaik, L. K. Gilbert, and J. P. Luzio. Reconstitution of an endosome-lysosome interaction in a cell-free system. Journal of Cell Biology, 108(6):2093–??, June 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/108/6/

Monteiro:1989:ENN

[MC89a] M. J. Monteiro and D. W. Cleveland. Expression of NF–L and NF–M in fibroblasts reveals coassembly of neurofilament and vimentin subunits. Journal of Cell Biology, 108(2):579–??, February 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/108/2/
579.

Mooseker:1989:KPC

[MC89b] M. S. Mooseker and T. R. Coleman. The 110-kD proteincalmodulin complex of the intestinal microvillus (brush border myosin I) is a mechanoenzyme. Journal of Cell Biology, 108 (6):2395–??, June 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/108/6/2395.

Morrow:1989:ALF

[MCA^{+}89] J. S. Morrow, C. D. Cianci, T. Ardito, A. S. Mann, and M. Kashgarian. Ankyrin links fodrin to the alpha subunit of Na,K– ATPase in Madin–Darby canine kidney cells and in intact renal tubule cells. Journal of Cell Biology, 108(2):455–??, February 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/108/2/
455.

Majack:1985:PDG

[MCB85] R. A. Majack, S. C. Cook, and P. Bornstein. Plateletderived growth factor and heparin-like glycosaminoglycans regulate thrombospondin synthesis and deposition in the matrix by smooth muscle cells. Journal of Cell Biology, 101(3):1059– ??, September 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/101/3/1059.

Milks:1986:ENM

[MCC86] L. C. Milks, G. P. Conyers, and E. B. Cramer. The effect of neutrophil migration on epithelial permeability. Journal of Cell Biology, 103(6):2729–??, December 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/103/6/2729.

Mooseker:1989:CIM

[MCC^{+}89] M. S. Mooseker, K. A. Conzelman, T. R. Coleman, J. E. Heuser, and M. P. Sheetz. Characterization of intestinal microvillar membrane disks: detergent-resistant membrane sheets enriched in associated brush border myosin I (110K-calmodulin). Journal of Cell Biology, 109(3):1153–??, September 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/109/3/1153.

Musil:1988:ARA

[MCCM88] L. S. Musil, C. Carr, J. B. Cohen, and J. P. Merlie. Acetylcholine receptor-associated 43K protein contains covalently bound myristate. Journal of Cell Biology, 107(3):1113–??, September 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/107/3/1113.

Mierendorf:1985:PIT

[MCD85] R. C. Mierendorf, J. A. Cardelli, and R. L. Dimond. Pathways involved in targeting and secretion of a lysosomal enzyme in Dictyostelium discoideum. Journal of Cell Biology, 100(5):1777–??, May 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/100/5/
1777.

Majack:1988:EAC

[MCG^{+}88a] R. A. Majack, C. K. Castle, L. V. Goodman, K. H. Weisgraber, R. W. Mahley, E. M. Shooter, and P. J. Gebicke-Haerter. Expression of apolipoprotein E by cultured vascular smooth muscle cells is controlled by growth state. Journal of Cell Biology, 107 (3):1207–??, September 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/107/3/1207.

Morris:1988:MFL

[MCG88b] L. Morris, P. R. Crocker, and S. Gordon. Murine fetal liver macrophages bind developing erythroblasts by a divalent cationdependent hemagglutinin. Journal of Cell Biology, 106(3): 649–??, March 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/106/3/649.

Maher:1985:ZNH

[MCS85a] P. A. Maher, G. F. Cox, and S. J. Singer. Zeugmatin: a new high molecular weight protein associated with Z lines in adult and early embryonic striated muscle. Journal of Cell Biology, 101 (5):1871–??, November 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/101/5/1871.

Miller:1985:SFM

[MCS85b] J. B. Miller, M. T. Crow, and F. E. Stockdale. Slow and fast myosin heavy chain content defines three types of myotubes in early muscle cell cultures. Journal of Cell Biology, 101(5): 1643–??, November 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/101/5/1643.

Madara:1985:OJS

[MD85] J. L. Madara and K. Dharmsathaphorn. Occluding junction structure-function relationships in a cultured epithelial monolayer. Journal of Cell Biology, 101(6):2124–??, December 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/101/6/2124.

Moehle:1989:PPP

[MDJ89] C. M. Moehle, C. K. Dixon, and E. W. Jones. Processing pathway for protease B of Saccharomyces cerevisiae. Journal of Cell Biology, 108(2):309–??, February 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/108/2/309.

McGraw:1988:PET

[MDM88] T. E. McGraw, K. W. Dunn, and F. R. Maxfield. Phorbol ester treatment increases the exocytic rate of the transferrin receptor recycling pathway independent of serine-24 phosphorylation. Journal of Cell Biology, 106(4):1061–??, April 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/106/4/1061.

Moya:1985:ICP

[MDVG^{+}85] M. Moya, A. Dautry-Varsat, B. Goud, D. Louvard, and P. Boquet. Inhibition of coated pit formation in Hep2 cells blocks the cytotoxicity of diphtheria toxin but not that of ricin toxin.

Journal of Cell Biology, 101(2):548–??, August 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/101/2/548.

Moutsatsos:1986:ELC

[MDW86] I. K. Moutsatsos, J. M. Davis, and J. L. Wang. Endogenous lectins from cultured cells: subcellular localization of carbohydrate-binding protein 35 in 3T3 fibroblasts. Journal of Cell Biology, 102(2):477–??, February 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/102/2/477.

Madreperla:1989:KAT

[MEA89] S. A. Madreperla, M. Edidin, and R. Adler. Na+,k+-adenosine triphosphatase polarity in retinal photoreceptors: a role for cytoskeletal attachments. Journal of Cell Biology, 109(4):1483–??, October 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
109/4/1483.

Mendler:1989:CCM

[MEBV^{+}89] M. Mendler, S. G. Eich-Bender, L. Vaughan, K. H. Winterhalter, and P. Bruckner. Cartilage contains mixed fibrils of collagen types II, IX, and XI. Journal of Cell Biology, 108(1):191–??, January 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/108/1/
191.

Mattsson:1986:MRH

[MER^{+}86] M. E. Mattsson, G. Enberg, A. I. Ruusala, K. Hall, and S. Påhlman. Mitogenic response of human SH–SY5Y neuroblastoma cells to insulin-like growth factor I and II is dependent on the stage of differentiation. Journal of Cell Biology, 102(5):1949–??, May 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/102/5/1949.

Muir:1989:SCD

[MEVM89] D. Muir, E. Engvall, S. Varon, and M. Manthorpe. Schwannoma cell-derived inhibitor of the neurite-promoting activity of laminin. Journal of Cell Biology, 109(5):2353–??, November 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/109/5/

Milligan:1987:SRA

[MF87a] R. A. Milligan and P. F. Flicker. Structural relationships of actin, myosin, and tropomyosin revealed by cryo-electron microscopy. Journal of Cell Biology, 105(1):29–??, July 1987. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/105/1/29.

Mohr:1987:DRB

[MF87b] F. C. Mohr and C. Fewtrell. Depolarization of rat basophilic leukemia cells inhibits calcium uptake and exocytosis. Journal of Cell Biology, 104(3):783–??, March 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/104/3/783.

Miller:1989:ABP

[MFA89] K. G. Miller, C. M. Field, and B. M. Alberts. Actin-binding proteins from Drosophila embryos: a complex network of interacting proteins detected by F-actin affinity chromatography. Journal of Cell Biology, 109(6):2963–??, December 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/109/6/2963.

McCaffrey:1989:TGF

[MFB^{+}89] T. A. McCaffrey, D. J. Falcone, C. F. Brayton, L. A. Agarwal, F. G. Welt, and B. B. Weksler. Transforming growth factorbeta activity is potentiated by heparin via dissociation of the transforming growth factor-beta/alpha 2-macroglobulin inactive complex. Journal of Cell Biology, 109(1):441–??, July 1989. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/109/1/441.

Maratos-Flier:1987:RMV

[MFKVK87] E. Maratos-Flier, C. Y. Kao, E. M. Verdin, and G. L. King. Receptor-mediated vectorial transcytosis of epidermal growth factor by Madin–Darby canine kidney cells. Journal of Cell Biology, 105(4):1595–??, October 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/105/4/1595.

Musil:1989:MKA

[MFM89] L. S. Musil, D. E. Frail, and J. P. Merlie. The mammalian 43-kD acetylcholine receptor-associated protein (RAPsyn) is expressed in some nonmuscle cells. Journal of Cell Biology, 108(5):1833–??, May 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/108/5/
1833.

Miller:1985:GJS

[MG85] T. M. Miller and D. A. Goodenough. Gap junction structures after experimental alteration of junctional channel conductance. Journal of Cell Biology, 101(5):1741–??, November 1985. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/101/5/1741.

Miller:1986:ETP

[MG86a] T. M. Miller and D. A. Goodenough. Evidence for two physiologically distinct gap junctions expressed by the chick lens epithelial cell. Journal of Cell Biology, 102(1):194–??, January 1986. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/102/1/194.

Muller:1986:PPC

[MG86b] W. A. Muller and M. A. Gimbrone. Plasmalemmal proteins of cultured vascular endothelial cells exhibit apical-basal polarity: analysis by surface-selective iodination. Journal of Cell Biology, 103(6):2389–??, December 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/103/6/2389.

Montell:1989:DLS

[MG89] D. J. Montell and C. S. Goodman. Drosophila laminin: sequence of B2 subunit and expression of all three subunits during embryogenesis. Journal of Cell Biology, 109(5):2441–??, November 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/109/5/
2441.

Majack:1988:CST

[MGD88] R. A. Majack, L. V. Goodman, and V. M. Dixit. Cell surface thrombospondin is functionally essential for vascular smooth muscle cell proliferation. Journal of Cell Biology, 106(2):415–??, February 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
106/2/415.

Murphy:1988:DDA

[MGGP88] D. B. Murphy, R. O. Gray, W. A. Grasser, and T. D. Pollard. Direct demonstration of actin filament annealing in vitro. Journal of Cell Biology, 106(6):1947–??, June 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/106/6/1947.

Muller-Glauser:1986:RTI

[MGHG^{+}86] W. Müller-Glauser, B. Humbel, M. Glatt, P. Sträuli, K. H. Winterhalter, and P. Bruckner. On the role of type IX collagen in the extracellular matrix of cartilage: type IX collagen is localized to intersections of collagen fibrils. Journal of Cell Biology, 102(5):1931–??, May 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/102/5/1931.

Messner:1989:ICM

[MGK89] D. J. Messner, G. Griffiths, and S. Kornfeld. Isolation and characterization of membranes from bovine liver which are highly enriched in mannose 6-phosphate receptors. Journal of Cell Biology, 108(6):2149–??, June 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/108/6/2149.

McGraw:1987:FEH

[MGM87] T. E. McGraw, L. Greenfield, and F. R. Maxfield. Functional expression of the human transferrin receptor cDNA in Chinese hamster ovary cells deficient in endogenous transferrin receptor. Journal of Cell Biology, 105(1):207–??, July 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/105/1/207.

Maruta:1986:AAT

[MGR86] H. Maruta, K. Greer, and J. L. Rosenbaum. The acetylation of alpha-tubulin and its relationship to the assembly and disassembly of microtubules. Journal of Cell Biology, 103(2):571–??, August 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/103/2/
571.

McCoy:1988:APS

[MGSG88] K. McCoy, S. Gal, R. H. Schwartz, and M. M. Gottesman. An acid protease secreted by transformed cells interferes with antigen processing. Journal of Cell Biology, 106(6):1879–??, June 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/106/6/
1879.

Murphy:1986:IEB

[MGW86] D. B. Murphy, W. A. Grasser, and K. T. Wallis. Immunofluorescence examination of beta tubulin expression and marginal band formation in developing chicken erythroblasts. Journal of Cell Biology, 102(2):628–??, February 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/102/2/628.

May:1985:IFA

[MGWM85] G. S. May, J. Gambino, J. A. Weatherbee, and N. R. Morris. Identification and functional analysis of beta-tubulin genes by site specific integrative transformation in Aspergillus nidulans. Journal of Cell Biology, 101(3):712–??, September 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/101/3/712.

Mayerson:1986:RRP

[MH86a] P. L. Mayerson and M. O. Hall. Rat retinal pigment epithelial cells show specificity of phagocytosis in vitro. Journal of Cell Biology, 103(1):299–??, July 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/103/1/299.

Mueller:1986:RME

[MH86b] S. C. Mueller and A. L. Hubbard. Receptor-mediated endocytosis of asialoglycoproteins by rat hepatocytes: receptor-positive and receptor-negative endosomes. Journal of Cell Biology, 102 (3):932–??, March 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/102/3/932.

Marcantonio:1988:ACC

[MH88] E. E. Marcantonio and R. O. Hynes. Antibodies to the conserved cytoplasmic domain of the integrin beta 1 subunit react with proteins in vertebrates, invertebrates, and fungi. Journal of Cell Biology, 106(5):1765–??, May 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/106/5/1765.

Moens:1987:SCA

[MHD^{+}87] P. B. Moens, C. Heyting, A. J. Dietrich, W. van Raamsdonk, and Q. Chen. Synaptonemal complex antigen location and conservation. Journal of Cell Biology, 105(1):93–??, July 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/105/1/93.

McCarthy:1986:HFC

[MHF86] J. B. McCarthy, S. T. Hagen, and L. T. Furcht. Human fibronectin contains distinct adhesion- and motility-promoting domains for metastatic melanoma cells. Journal of Cell Biology, 102(1):179–??, January 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/102/1/179.

Mabuchi:1985:AAS

[MHK^{+}85] I. Mabuchi, Y. Hamaguchi, T. Kobayashi, H. Hosoya, S. Tsukita, and S. Tsukita. Alpha-actinin from sea urchin eggs: biochemical properties, interaction with actin, and distribution in the cell during fertilization and cleavage. Journal of Cell Biology, 100 (2):375–??, February 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/100/2/375.

Mackie:1988:ITH

[MHL88] E. J. Mackie, W. Halfter, and D. Liverani. Induction of tenascin in healing wounds. Journal of Cell Biology, 107(6):2757–??, December 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
107/6/2757.

Murray:1986:ASM

[MHP^{+}86] B. A. Murray, J. J. Hemperly, E. A. Prediger, G. M. Edelman, and B. A. Cunningham. Alternatively spliced mRNAs code for different polypeptide chains of the chicken neural cell adhesion molecule (N-CAM). Journal of Cell Biology, 102(1):189–??, January 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/102/1/
189.

Maro:1985:NSM

[MHW85] B. Maro, S. K. Howlett, and M. Webb. Non-spindle microtubule organizing centers in metaphase II-arrested mouse oocytes. Journal of Cell Biology, 101(5):1665–??, November 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/101/5/1665.

Menon:1986:CMT

[MHWB86a] A. K. Menon, D. Holowka, W. W. Webb, and B. Baird. Clustering, mobility, and triggering activity of small oligomers of immunoglobulin E on rat basophilic leukemia cells. Journal of Cell Biology, 102(2):534–??, February 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/102/2/534.

Menon:1986:CLR

[MHWB86b] A. K. Menon, D. Holowka, W. W. Webb, and B. Baird. Crosslinking of receptor-bound IgE to aggregates larger than dimers leads to rapid immobilization. Journal of Cell Biology, 102(2): 541–??, February 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/102/2/541.

Muller:1986:ESS

[MIHS86] H. W. Müller, M. J. Ignatius, D. H. Hangen, and E. M. Shooter. Expression of specific sheath cell proteins during peripheral nerve growth and regeneration in mammals. Journal of Cell Biology, 102(2):393–??, February 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/102/2/393.

Mitchison:1989:PMF

[Mit89] T. J. Mitchison. Polewards microtubule flux in the mitotic spindle: evidence from photoactivation of fluorescence. Journal of Cell Biology, 109(2):637–??, August 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/109/2/637.

Miyazaki:1988:ITI

[Miy88] S. Miyazaki. Inositol 1,4,5-trisphosphate-induced calcium release and guanine nucleotide-binding protein-mediated periodic calcium rises in golden hamster eggs. Journal of Cell Biology, 106 (2):345–??, February 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/106/2/345.

McKenna:1986:FAZ

[MJW86] N. M. McKenna, C. S. Johnson, and Y. L. Wang. Formation and alignment of Z lines in living chick myotubes microinjected with rhodamine-labeled alpha-actinin. Journal of Cell Biology, 103 (6):2163–??, December 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/103/6/2163.

Mitchison:1985:PKVa

[MK85a] T. J. Mitchison and M. W. Kirschner. Properties of the kinetochore in vitro. I. Microtubule nucleation and tubulin binding. Journal of Cell Biology, 101(3):755–??, September 1985. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/101/3/755.

Mitchison:1985:PKVb

[MK85b] T. J. Mitchison and M. W. Kirschner. Properties of the kinetochore in vitro. II. Microtubule capture and ATP-dependent translocation. Journal of Cell Biology, 101(3):766–??, September 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/101/3/
766.

Moore:1985:SPT

[MK85c] H. P. Moore and R. B. Kelly. Secretory protein targeting in a pituitary cell line: differential transport of foreign secretory proteins to distinct secretory pathways. Journal of Cell Biology, 101 (5):1773–??, November 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/101/5/1773.

Matteoni:1987:TCE

[MK87] R. Matteoni and T. E. Kreis. Translocation and clustering of endosomes and lysosomes depends on microtubules. Journal of Cell Biology, 105(3):1253–??, September 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/105/3/1253.

Murofushi:1986:PCK

[MKA^{+}86] H. Murofushi, S. Kotani, H. Aizawa, S. Hisanaga, N. Hirokawa, and H. Sakai. Purification and characterization of a 190-kD microtubule-associated protein from bovine adrenal cortex. Journal of Cell Biology, 103(5):1911–??, November 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/103/5/1911.

Mueller:1989:DCI

[MKD^{+}89] S. C. Mueller, T. Kelly, M. Z. Dai, H. N. Dai, and W. T. Chen. Dynamic cytoskeleton-integrin associations induced by cell binding to immobilized fibronectin. Journal of Cell Biology, 109(6): 3455–??, December 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/6/3455.

Meyer:1988:NMN

[MKH88] S. L. Meyer, S. G. Kaminskyj, and I. B. Heath. Nuclear migration in a nud mutant of Aspergillus nidulans is inhibited in the presence of a quantitatively normal population of cytoplasmic microtubules. Journal of Cell Biology, 106(3):773–??, March 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/106/3/
773.

Mack:1989:APC

[MKK89] D. Mack, B. Kluxen, and J. Kruppa. Accessibility to proteases of the cytoplasmic G protein domain of vesicular stomatitis virus is increased during intracellular transport. Journal of Cell Biology, 109(5):2057–??, November 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/109/5/2057.

Miller:1985:CBM

[ML85] R. H. Miller and R. J. Lasek. Cross-bridges mediate anterograde and retrograde vesicle transport along microtubules in squid axoplasm. Journal of Cell Biology, 101(6):2181–??, December 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/101/6/2181.

Mugnai:1988:MMA

[MLC^{+}88] G. Mugnai, K. Lewandowska, B. Carnemolla, L. Zardi, and L. A. Culp. Modulation of matrix adhesive responses of human neuroblastoma cells by neighboring sequences in the fibronectins. Journal of Cell Biology, 106(3):931–??, March 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/106/3/931.

McKeown-Longo:1987:IFM

[MLE87] P. J. McKeown-Longo and C. A. Etzler. Induction of fibronectin matrix assembly in human fibrosarcoma cells by dexamethasone. Journal of Cell Biology, 104(3):601–??, March 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/104/3/601.

Momayezi:1987:EIP

[MLK^{+}87] M. Momayezi, C. J. Lumpert, H. Kersken, U. Gras, H. Plattner, M. H. Krinks, and C. B. Klee. Exocytosis induction in Paramecium tetraurelia cells by exogenous phosphoprotein phosphatase in vivo and in vitro: possible involvement of calcineurin in exocytotic membrane fusion. Journal of Cell Biology, 105(1):181–??, July 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/105/1/
181.

McKeown-Longo:1985:IMW

[MLM85] P. J. McKeown-Longo and D. F. Mosher. Interaction of the 70,000-mol-wt amino-terminal fragment of fibronectin with the matrix-assembly receptor of fibroblasts. Journal of Cell Biology, 100(2):364–??, February 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/100/2/364.

Mahl:1989:REH

[MLPF89] P. Mähl, Y. Lutz, E. Puvion, and J. P. Fuchs. Rapid effect of heat shock on two heterogeneous nuclear ribonucleoproteinassociated antigens in HeLa cells. Journal of Cell Biology, 109 (5):1921–??, November 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/5/1921.

Madsen:1986:CCL

[MLR^{+}86] O. D. Madsen, L. I. Larsson, J. F. Rehfeld, T. W. Schwartz, A. Lernmark, A. D. Labrecque, and D. F. Steiner. Cloned cell lines from a transplantable islet cell tumor are heterogeneous and express cholecystokinin in addition to islet hormones. Journal of Cell Biology, 103(5):2025–??, November 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/103/5/2025.

Mar:1986:NPR

[MLSW86] J. Mar, J. H. Lee, D. Shea, and C. J. Walsh. New poly(A)+RNAs appear coordinately during the differentiation of Naegleria gruberi amebae into flagellates. Journal of Cell Biology, 102(2): 353–??, February 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/102/2/353.

Molday:1987:DPC

[MM87] R. S. Molday and L. L. Molday. Differences in the protein composition of bovine retinal rod outer segment disk and plasma membranes isolated by a ricin-gold-dextran density perturbation method. Journal of Cell Biology, 105(6):2589–??, December 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/105/6/
2589.

Milarski:1989:MAH

[MM89] K. L. Milarski and R. I. Morimoto. Mutational analysis of the human HSP70 protein: distinct domains for nucleolar localization and adenosine triphosphate binding. Journal of Cell Biology, 109 (5):1947–??, November 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/5/1947.

Mahaffey:1989:CPI

[MMBA89] D. T. Mahaffey, M. S. Moore, F. M. Brodsky, and R. G. Anderson. Coat proteins isolated from clathrin coated vesicles can assemble into coated pits. Journal of Cell Biology, 108(5):1615–??, May 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/108/5/
1615.

Masuda:1988:MAS

[MMC88] H. Masuda, K. L. McDonald, and W. Z. Cande. The mechanism of anaphase spindle elongation: uncoupling of tubulin incorporation and microtubule sliding during in vitro spindle reactivation. Journal of Cell Biology, 107(2):623–??, August 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/107/2/623.

Migliaccio:1989:SLR

[MMK^{+}89] G. Migliaccio, A. R. Migliaccio, B. L. Kreider, G. Rovera, and J. W. Adamson. Selection of lineage-restricted cell lines immortalized at different stages of hematopoietic differentiation from the murine cell line 32D. Journal of Cell Biology, 109 (2):833–??, August 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/2/833.

Morel:1985:LSP

[MMM^{+}85] N. Morel, J. Marsal, R. Manaranche, S. Lazereg, J. C. Mazie, and M. Israel. Large-scale purification of presynaptic plasma membranes from Torpedo marmorata electric organ. Journal of Cell Biology, 101(5):1757–??, November 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/101/5/1757.

Mische:1987:EAC

[MMM87] S. M. Mische, M. S. Mooseker, and J. S. Morrow. Erythrocyte adducin: a calmodulin-regulated actin-bundling protein that stimulates spectrin-actin binding. Journal of Cell Biology, 105(6): 2837–??, December 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/6/2837.

Masumoto:1989:HCA

[MMM^{+}89] H. Masumoto, H. Masukata, Y. Muro, N. Nozaki, and T. Okazaki. A human centromere antigen (CENP–B) interacts with a short specific sequence in alphoid DNA, a human centromeric satellite. Journal of Cell Biology, 109(5): 1963–??, November 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/5/1963.

Malgaroli:1987:FMC

[MMMP87] A. Malgaroli, D. Milani, J. Meldolesi, and T. Pozzan. Fura-2 measurement of cytosolic free Ca^{2+} in monolayers and suspensions of various types of animal cells. Journal of Cell Biology, 105 (5):2145–??, November 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/5/2145.

Malnoe:1988:CAB

[MMR88] P. Malnoë, S. P. Mayfield, and J. D. Rochaix. Comparative analysis of the biogenesis of photosystem II in the wild-type and Y-1 mutant of Chlamydomonas reinhardtii. Journal of Cell Biology, 106(3):609–??, March 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/106/3/609.

Miskimins:1986:CPE

[MMRR86] W. K. Miskimins, A. McClelland, M. P. Roberts, and F. H. Ruddle. Cell proliferation and expression of the transferrin receptor gene: promoter sequence homologies and protein interactions. Journal of Cell Biology, 103(5):1781–??, November 1986. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/103/5/1781.

Mitra:1989:TDS

[MMS89] A. K. Mitra, M. P. McCarthy, and R. M. Stroud. Threedimensional structure of the nicotinic acetylcholine receptor and location of the major associated 43-kD cytoskeletal protein, determined at 22 A by low dose electron microscopy and X-ray diffraction to 12.5 A. Journal of Cell Biology, 109(2):755–??, August 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/109/2/
755.

McNeil:1985:TRC

[MMT85] P. L. McNeil, M. P. McKenna, and D. L. Taylor. A transient rise in cytosolic calcium follows stimulation of quiescent cells with growth factors and is inhibitable with phorbol myristate acetate. Journal of Cell Biology, 101(2):372–??, August 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/101/2/372.

McKenna:1985:IDF

[MMW85a] N. McKenna, J. B. Meigs, and Y. L. Wang. Identical distribution of fluorescently labeled brain and muscle actins in living cardiac fibroblasts and myocytes. Journal of Cell Biology, 100 (1):292–??, January 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/100/1/292.

McKenna:1985:EAA

[MMW85b] N. M. McKenna, J. B. Meigs, and Y. L. Wang. Exchangeability of alpha-actinin in living cardiac fibroblasts and muscle cells. Journal of Cell Biology, 101(6):2223–??, December 1985. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/101/6/2223.

McNeil:1989:GFR

[MMWD89] P. L. McNeil, L. Muthukrishnan, E. Warder, and P. A. D’Amore. Growth factors are released by mechanically wounded endothelial cells. Journal of Cell Biology, 109(2):811–??, August 1989. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/109/2/811.

Matagne:1989:ICM

[MMWM^{+}89] R. F. Matagne, M. R. Michel-Wolwertz, C. Munaut, C. Duyckaerts, and F. Sluse. Induction and characterization of mitochondrial DNA mutants in Chlamydomonas reinhardtii. Journal of Cell Biology, 108(4):1221–??, April 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/108/4/1221.

Matsuzaki:1988:TCT

[MMY88] F. Matsuzaki, S. Matsumoto, and I. Yahara. Truncation of the carboxy-terminal domain of yeast beta-tubulin causes temperature-sensitive growth and hypersensitivity to antimitotic drugs. Journal of Cell Biology, 107(4):1427–??, October 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/107/4/1427.

Makarow:1987:TFM

[MN87] M. Makarow and L. T. Nevalainen. Transport of a fluorescent macromolecule via endosomes to the vacuole in Saccharomyces cerevisiae. Journal of Cell Biology, 104(1):67–??, January 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/104/1/67.

Madsen:1986:MAS

[MNC86] P. Madsen, S. Nielsen, and J. E. Celis. Monoclonal antibody specific for human nuclear proteins IEF 8Z30 and 8Z31 accumulates in the nucleus a few hours after cytoplasmic microinjection of cells expressing these proteins. Journal of Cell Biology, 103 (6):2083–??, December 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/103/6/2083.

Miller:1987:IAT

[MND^{+}87] F. D. Miller, C. C. Naus, M. Durand, F. E. Bloom, and R. J. Milner. Isotypes of alpha-tubulin are differentially regulated during neuronal maturation. Journal of Cell Biology, 105(6): 3065–??, December 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/6/3065.

Matuoka:1987:HSI

[MNM87] K. Matuoka, M. Namba, and Y. Mitsui. Hyaluronate synthetase inhibition by normal and transformed human fibroblasts during growth reduction. Journal of Cell Biology, 104(4):1105–??, January 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/104/4/
1105.

Moon:1985:TSE

[MNWL85] R. T. Moon, J. Ngai, B. J. Wold, and E. Lazarides. Tissuespecific expression of distinct spectrin and ankyrin transcripts in erythroid and nonerythroid cells. Journal of Cell Biology, 100 (1):152–??, January 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/100/1/152.

Markelonis:1985:STR

[MOP^{+}85] G. J. Markelonis, T. H. Oh, L. P. Park, C. Y. Cha, C. A. Sofia, J. W. Kim, and P. Azari. Synthesis of the transferrin receptor by cultures of embryonic chicken spinal neurons. Journal of Cell Biology, 100(1):8–??, January 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/100/1/8.

Murray:1986:CSM

[MOP^{+}86] B. A. Murray, G. C. Owens, E. A. Prediger, K. L. Crossin, B. A. Cunningham, and G. M. Edelman. Cell surface modulation of the neural cell adhesion molecule resulting from alternative mRNA splicing in a tissue-specific developmental sequence. Journal of Cell Biology, 103(4):1431–??, October 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/103/4/1431.

Moscatelli:1988:MRB

[Mos88] D. Moscatelli. Metabolism of receptor-bound and matrix-bound basic fibroblast growth factor by bovine capillary endothelial cells. Journal of Cell Biology, 107(2):753–??, August 1988. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/107/2/753.

Mehta:1985:SSC

[MOT^{+}85] H. Mehta, C. Orphe, M. S. Todd, C. J. Cornbrooks, and D. J. Carey. Synthesis by Schwann cells of basal lamina and membrane-associated heparan sulfate proteoglycans. Journal of Cell Biology, 101(2):660–??, August 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/101/2/660.

Mekada:1988:IDT

[MOU88] E. Mekada, Y. Okada, and T. Uchida. Identification of diphtheria toxin receptor and a nonproteinous diphtheria toxin-binding molecule in Vero cell membrane. Journal of Cell Biology, 107 (2):511–??, August 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/107/2/511.

McCloskey:1986:CIR

[MP86a] M. A. McCloskey and M. M. Poo. Contact-induced redistribution of specific membrane components: local accumulation and development of adhesion. Journal of Cell Biology, 102(6):2185–??, June 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/102/6/
2185.

McCloskey:1986:RMA

[MP86b] M. A. McCloskey and M. M. Poo. Rates of membrane-associated reactions: reduction of dimensionality revisited. Journal of Cell Biology, 102(1):88–??, January 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/102/1/88.

McCurdy:1986:IEM

[MP86c] D. W. McCurdy and L. H. Pratt. Immunogold electron microscopy of phytochrome in Avena: identification of intracellular sites responsible for phytochrome sequestering and enhanced pelletability. Journal of Cell Biology, 103(6):2541–??, December 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/103/6/
2541.

McNiven:1986:MPC

[MP86d] M. A. McNiven and K. R. Porter. Microtubule polarity confers direction to pigment transport in chromatophores. Journal of Cell Biology, 103(4):1547–??, October 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/103/4/1547.

Maher:1988:TPP

[MP88a] P. A. Maher and E. B. Pasquale. Tyrosine phosphorylated proteins in different tissues during chick embryo development. Journal of Cell Biology, 106(5):1747–??, May 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/106/5/1747.

McNiven:1988:OMC

[MP88b] M. A. McNiven and K. R. Porter. Organization of microtubules in centrosome-free cytoplasm. Journal of Cell Biology, 106(5):1593–??, May 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/106/5/1593.

Maher:1989:HSI

[MP89] P. A. Maher and E. B. Pasquale. Heat shock induces protein tyrosine phosphorylation in cultured cells. Journal of Cell Biology, 108(6):2029–??, June 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/108/6/2029.

Murakami:1988:KHS

[MPB88] H. Murakami, D. Pain, and G. Blobel. 70-kD heat shock-related protein is one of at least two distinct cytosolic factors stimulating protein import into mitochondria. Journal of Cell Biology, 107 (6):2051–??, December 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/107/6/2051.

Martin-Perez:1989:TRS

[MPBZBE89] J. Martin-Perez, D. Bar-Zvi, D. Branton, and R. L. Erikson. Transformation by Rous sarcoma virus induces clathrin heavy chain phosphorylation. Journal of Cell Biology, 109(2):577–??, August 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
109/2/577.

Madri:1988:PME

[MPT88] J. A. Madri, B. M. Pratt, and A. M. Tucker. Phenotypic modulation of endothelial cells by transforming growth factor-beta depends upon the composition and organization of the extracellular matrix. Journal of Cell Biology, 106(4):1375–??, April 1988. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/106/4/1375.

Moscatelli:1989:TNC

[MQ89] D. Moscatelli and N. Quarto. Transformation of NIH 3T3 cells with basic fibroblast growth factor or the hst/K-fgf oncogene causes downregulation of the fibroblast growth factor receptor: reversal of morphological transformation and restoration of receptor number by suramin. Journal of Cell Biology, 109(5): 2519–??, November 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/5/2519.

Mitchell:1985:APP

[MR85a] C. K. Mitchell and D. A. Redburn. Analysis of pre- and postsynaptic factors of the serotonin system in rabbit retina. Journal of Cell Biology, 100(1):64–??, January 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/100/1/64.

Mitchell:1985:MCF

[MR85b] D. R. Mitchell and J. L. Rosenbaum. A motile Chlamydomonas flagellar mutant that lacks outer dynein arms. Journal of Cell Biology, 100(4):1228–??, April 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/100/4/1228.

Machamer:1987:STD

[MR87] C. E. Machamer and J. K. Rose. A specific transmembrane domain of a coronavirus E1 glycoprotein is required for its retention in the Golgi region. Journal of Cell Biology, 105(3): 1205–??, September 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/3/1205.

Mercier:1989:MCD

[MRD^{+}89] F. Mercier, H. Reggio, G. Devilliers, D. Bataille, and P. Mangeat. Membrane-cytoskeleton dynamics in rat parietal cells: mobilization of actin and spectrin upon stimulation of gastric acid secretion. Journal of Cell Biology, 108(2):441–??, February 1989. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/108/2/441.

Matthieu:1986:MIO

[MRO^{+}86] J. M. Matthieu, J. M. Roch, F. X. Omlin, I. Rambaldi, G. Almazan, and P. E. Braun. Myelin instability and oligodendrocyte metabolism in myelin-deficient mutant mice. Journal of Cell Biology, 103(6):2673–??, December 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/103/6/2673.

Marchese-Ragona:1988:SMA

[MRWJ88] S. P. Marchese-Ragona, J. S. Wall, and K. A. Johnson. Structure and mass analysis of 14S dynein obtained from Tetrahymena cilia. Journal of Cell Biology, 106(1):127–??, January 1988. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/106/1/127.

Ma:1986:DDC

[MS86a] A. S. Ma and T. T. Sun. Differentiation-dependent changes in the solubility of a 195-kD protein in human epidermal keratinocytes. Journal of Cell Biology, 103(1):41–??, July 1986. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/103/1/41.

Martini:1986:IML

[MS86b] R. Martini and M. Schachner. Immunoelectron microscopic localization of neural cell adhesion molecules (L1, N-CAM, and MAG) and their shared carbohydrate epitope and myelin basic protein in developing sciatic nerve. Journal of Cell Biology, 103 (6):2439–??, December 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/103/6/2439.

Miller:1986:DRM

[MS86c] J. B. Miller and F. E. Stockdale. Developmental regulation of the multiple myogenic cell lineages of the avian embryo. Journal of Cell Biology, 103(6):2197–??, December 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/103/6/2197.

Martini:1988:IML

[MS88a] R. Martini and M. Schachner. Immunoelectron microscopic localization of neural cell adhesion molecules (L1, N-CAM, and myelin-associated glycoprotein) in regenerating adult mouse sciatic nerve. Journal of Cell Biology, 106(5):1735–??, May 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/106/5/1735.

Mercurio:1988:MIL

[MS88b] A. M. Mercurio and L. M. Shaw. Macrophage interactions with laminin: PMA selectively induces the adherence and spreading of mouse macrophages on a laminin substratum. Journal of Cell Biology, 107(5):1873–??, November 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/107/5/1873.

Mikol:1988:PLP

[MS88c] D. D. Mikol and K. Stefansson. A phosphatidylinositol-linked peanut agglutinin-binding glycoprotein in central nervous system myelin and on oligodendrocytes. Journal of Cell Biology, 106 (4):1273–??, April 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/106/4/1273.

Millonig:1988:PAP

[MSA88] R. Millonig, H. Salvo, and U. Aebi. Probing actin polymerization by intermolecular cross-linking. Journal of Cell Biology, 106(3):785–??, March 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/106/3/785.

Meyers:1985:GAA

[MSC^{+}85] M. B. Meyers, B. A. Spengler, T. D. Chang, P. W. Melera, and J. L. Biedler. Gene amplification-associated cytogenetic aberrations and protein changes in vincristine-resistant Chinese hamster, mouse, and human cells. Journal of Cell Biology, 100(2):

588–??, February 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/100/2/588.

Matzuk:1989:MCG

[MSC^{+}89] M. M. Matzuk, M. M. Spangler, M. Camel, N. Suganuma, and I. Boime. Mutagenesis and chimeric genes define determinants in the beta subunits of human chorionic gonadotropin and lutropin for secretion and assembly. Journal of Cell Biology, 109(4): 1429–??, October 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/4/1429.

Martin:1988:IPS

[MSD^{+}88] D. P. Martin, R. E. Schmidt, P. S. DiStefano, O. H. Lowry, J. G. Carter, and E. M. Johnson. Inhibitors of protein synthesis and RNA synthesis prevent neuronal death caused by nerve growth factor deprivation. Journal of Cell Biology, 106(3):829–??, March 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/106/3/
829.

Martins:1989:DIT

[MSE89] T. J. Martins, Y. Sugimoto, and R. L. Erikson. Dissociation of inositol trisphosphate from diacylglycerol production in Rous sarcoma virus-transformed fibroblasts. Journal of Cell Biology, 108 (2):683–??, February 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/108/2/683.

Marsh:1987:RAP

[MSK^{+}87] M. Marsh, S. Schmid, H. Kern, E. Harms, P. Male, I. Mellman, and A. Helenius. Rapid analytical and preparative isolation of functional endosomes by free flow electrophoresis. Journal of Cell Biology, 104(4):875–??, January 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/104/4/875.

Mendoza:1986:EEE

[MSLR^{+}86] S. A. Mendoza, J. A. Schneider, A. Lopez-Rivas, J. W. Sinnett-Smith, and E. Rozengurt. Early events elicited by bombesin and structurally related peptides in quiescent Swiss 3T3 cells. II. Changes in Na+ and Ca^{2+} fluxes, Na+/K+ pump activity, and intracellular pH. Journal of Cell Biology, 102(6):2223–??, June 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/102/6/
2223.

Magill-Solc:1988:MNC

[MSM88] C. Magill-Solc and U. J. McMahan. Motor neurons contain agrinlike molecules. Journal of Cell Biology, 107(5):1825–??, November 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/107/5/
1825.

Marxer:1989:KAP

[MSQ^{+}89] A. Marxer, B. Stieger, A. Quaroni, M. Kashgarian, and H. P. Hauri. (Na^{+} + K^{+})-ATPase and plasma membrane polarity of intestinal epithelial cells: presence of a brush border antigen in the distal large intestine that is immunologically related to beta subunit. Journal of Cell Biology, 109(3):1057–??, September 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/109/3/
1057.

Mandelkow:1986:SLM

[MSR^{+}86] E. M. Mandelkow, R. Schultheiss, R. Rapp, M. Müller, and E. Mandelkow. On the surface lattice of microtubules: helix starts, protofilament number, seam, and handedness. Journal of Cell Biology, 102(3):1067–??, March 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/102/3/1067.

McFadden:1987:BBR

[MSS^{+}87a] G. I. McFadden, D. Schulze, B. Surek, J. L. Salisbury, and M. Melkonian. Basal body reorientation mediated by a Ca^{2+}- modulated contractile protein. Journal of Cell Biology, 105 (2):903–??, August 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/2/903.

Mittal:1987:VML

[MSS87b] B. Mittal, J. M. Sanger, and J. W. Sanger. Visualization of myosin in living cells. Journal of Cell Biology, 105(4):1753–??, October 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
105/4/1753.

Mohraz:1987:TDS

[MSS87c] M. Mohraz, M. V. Simpson, and P. R. Smith. The threedimensional structure of the Na,K–ATPase from electron microscopy. Journal of Cell Biology, 105(1):1–??, July 1987. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/105/1/1.

Minoo:1989:LPP

[MSS^{+}89] P. Minoo, W. Sullivan, L. R. Solomon, T. E. Martin, D. O. Toft, and R. E. Scott. Loss of proliferative potential during terminal differentiation coincides with the decreased abundance of a subset of heterogeneous ribonuclear proteins. Journal of Cell Biology, 109(5):1937–??, November 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/109/5/1937.

Mali:1988:LPM

[MSV^{+}88] P. Mali, M. Sandberg, E. Vuorio, P. C. Yelick, N. B. Hecht, and M. Parvinen. Localization of protamine 1 mRNA in different stages of the cycle of the rat seminiferous epithelium. Journal of Cell Biology, 107(2):407–??, August 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/107/2/407.

McNeil:1986:FRM

[MSW^{+}86] P. L. McNeil, J. A. Swanson, S. D. Wright, S. C. Silverstein, and D. L. Taylor. Fc-receptor-mediated phagocytosis occurs in macrophages without an increase in average [Ca++]i. Journal of Cell Biology, 102(5):1586–??, May 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/102/5/1586.

Matsuda:1987:TCW

[MSY^{+}87] Y. Matsuda, T. Saito, T. Yamaguchi, M. Koseki, and K. Hayashi. Topography of cell wall lytic enzyme in Chlamydomonas reinhardtii: form and location of the stored enzyme in vegetative cell and gamete. Journal of Cell Biology, 104(2):321–??, February 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/104/2/

Maeda:1986:MRP

[MT86] M. Maeda and G. A. Thompson. On the mechanism of rapid plasma membrane and chloroplast envelope expansion in Dunaliella salina exposed to hypoosmotic shock. Journal of Cell Biology, 102(1):289–??, January 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/102/1/289.

Myrdal:1986:CMC

[MTA86] S. E. Myrdal, D. R. Twardzik, and N. Auersperg. Cell-mediated co-action of transforming growth factors: incubation of type beta with normal rat kidney cells produces a soluble activity that prolongs the ruffling response to type alpha. Journal of Cell Biology, 102(4):1230–??, April 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/102/4/1230.

Mackie:1987:TAC

[MTCE87] E. J. Mackie, I. Thesleff, and R. Chiquet-Ehrismann. Tenascin is associated with chondrogenic and osteogenic differentiation in vivo and promotes chondrogenesis in vitro. Journal of Cell Biology, 105(6):2569–??, December 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/105/6/2569.

Mocz:1988:MPT

[MTG88] G. Mocz, W. J. Tang, and I. R. Gibbons. A map of photolytic and tryptic cleavage sites on the beta heavy chain of dynein AT-Pase from sea urchin sperm flagella. Journal of Cell Biology, 106(5):1607–??, May 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/106/5/1607.

Moore:1987:SMN

[MTK^{+}87] S. E. Moore, J. Thompson, V. Kirkness, J. G. Dickson, and F. S. Walsh. Skeletal muscle neural cell adhesion molecule (N-CAM): changes in protein and mRNA species during myogenesis of muscle cell lines. Journal of Cell Biology, 105(3):1377–??, September 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/105/3/

Mignatti:1989:VAH

[MTRR89] P. Mignatti, R. Tsuboi, E. Robbins, and D. B. Rifkin. In vitro angiogenesis on the human amniotic membrane: requirement for basic fibroblast growth factor-induced proteinases. Journal of Cell Biology, 108(2):671–??, February 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/108/2/671.

Mogensen:1989:MPI

[MTS89] M. M. Mogensen, J. B. Tucker, and H. Stebbings. Microtubule polarities indicate that nucleation and capture of microtubules occurs at cell surfaces in Drosophila. Journal of Cell Biology, 108(4):1445–??, April 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/108/4/1445.

Murphy-Ullrich:1989:TMF

[MUH89] J. E. Murphy-Ullrich and M. Höök. Thrombospondin modulates focal adhesions in endothelial cells. Journal of Cell Biology, 109 (3):1309–??, September 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/3/1309.

Murphy-Ullrich:1987:ITE

[MUM87] J. E. Murphy-Ullrich and D. F. Mosher. Interactions of thrombospondin with endothelial cells: receptor-mediated binding and degradation. Journal of Cell Biology, 105(4):1603–??, October 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/105/4/
1603.

Muscarella:1985:RRG

[MVB85] D. E. Muscarella, V. M. Vogt, and S. E. Bloom. The ribosomal RNA gene cluster in aneuploid chickens: evidence for increased gene dosage and regulation of gene expression. Journal of Cell Biology, 101(5):1749–??, November 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/101/5/1749.

Muscarella:1987:CRR

[MVB87] D. E. Muscarella, V. M. Vogt, and S. E. Bloom. Characterization of ribosomal RNA synthesis in a gene dosage mutant: the relationship of topoisomerase I and chromatin structure to transcriptional activity. Journal of Cell Biology, 105(4):1501–??, October 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/105/4/
1501.

Mauch:1988:DCS

[MvdMH^{+}88] C. Mauch, K. van der Mark, O. Helle, J. Mollenhauer, M. Pfäffle, and T. Krieg. A defective cell surface collagen-binding protein in dermatosparactic sheep fibroblasts. Journal of Cell Biology, 106(1):205–??, January 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/106/1/205.

Monier:1988:SIO

[MVK^{+}88] S. Monier, P. Van Luc, G. Kreibich, D. D. Sabatini, and M. Adesnik. Signals for the incorporation and orientation of cytochrome P450 in the endoplasmic reticulum membrane. Journal of Cell Biology, 107(2):457–??, August 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/107/2/457.

Medcalf:1988:GMG

[MVS88] R. L. Medcalf, E. Van den Berg, and W. D. Schleuning. Glucocorticoid-modulated gene expression of tissue- and urinarytype plasminogen activator and plasminogen activator inhibitor 1 and 2. Journal of Cell Biology, 106(3):971–??, March 1988. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/106/3/971.

Metcalf:1986:MAD

[MVSW86] T. N. Metcalf, M. A. Villanueva, M. Schindler, and J. L. Wang. Monoclonal antibodies directed against protoplasts of soybean cells: analysis of the lateral mobility of plasma membrane-bound antibody MVS-1. Journal of Cell Biology, 102(4):1350–??, April 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/102/4/
1350.

Meigs:1986:RAA

[MW86] J. B. Meigs and Y. L. Wang. Reorganization of alpha-actinin and vinculin induced by a phorbol ester in living cells. Journal of Cell Biology, 102(4):1430–??, April 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/102/4/1430.

McNiven:1988:CRP

[MW88a] M. A. McNiven and J. B. Ward. Calcium regulation of pigment transport in vitro. Journal of Cell Biology, 106(1):111–??, January 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/106/1/
111.

Mizzen:1988:CTC

[MW88b] L. A. Mizzen and W. J. Welch. Characterization of the thermotolerant cell. I. Effects on protein synthesis activity and the regulation of heat-shock protein 70 expression. Journal of Cell Biology, 106(4):1105–??, April 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/106/4/1105.

McKenna:1989:FMM

[MWK89] N. M. McKenna, Y. L. Wang, and M. E. Konkel. Formation and movement of myosin-containing structures in living fibroblasts. Journal of Cell Biology, 109(3):1163–??, September 1989. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/109/3/1163.

Milarski:1989:CCD

[MWM89] K. L. Milarski, W. J. Welch, and R. I. Morimoto. Cell cycledependent association of HSP70 with specific cellular proteins. Journal of Cell Biology, 108(2):413–??, February 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/108/2/413.

Mishkind:1985:FDT

[MWS85] M. L. Mishkind, S. R. Wessler, and G. W. Schmidt. Functional determinants in transit sequences: import and partial maturation by vascular plant chloroplasts of the ribulose-1,5-bisphosphate carboxylase small subunit of Chlamydomonas. Journal of Cell Biology, 100(1):226–??, January 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/100/1/226.

Miyatani:1986:SSK

[MWSD86] S. Miyatani, J. A. Winkles, T. D. Sargent, and I. B. Dawid. Stage-specific keratins in Xenopus laevis embryos and tadpoles: the XK81 gene family. Journal of Cell Biology, 103(5):

1957–??, November 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/103/5/1957.

McConnell:1987:HSI

[MWSF87] M. McConnell, A. M. Whalen, D. E. Smith, and P. A. Fisher. Heat shock-induced changes in the structural stability of proteinaceous karyoskeletal elements in vitro and morphological effects in situ. Journal of Cell Biology, 105(3):1087–??, September 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/105/3/
1087.

Milici:1987:TAC

[MWSP87] A. J. Milici, N. E. Watrous, H. Stukenbrok, and G. E. Palade. Transcytosis of albumin in capillary endothelium. Journal of Cell Biology, 105(6):2603–??, December 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/105/6/2603.

Maruyama:1985:CFL

[MYH^{+}85] K. Maruyama, T. Yoshioka, H. Higuchi, K. Ohashi, S. Kimura, and R. Natori. Connectin filaments link thick filaments and Z lines in frog skeletal muscle as revealed by immunoelectron microscopy. Journal of Cell Biology, 101(6):2167–??, December 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/101/6/
2167.

Masaki:1987:CPN

[MYT87] R. Masaki, A. Yamamoto, and Y. Tashiro. Cytochrome P-450 and NADPH-cytochrome P-450 reductase are degraded in the autolysosomes in rat liver. Journal of Cell Biology, 104(5):1207–??, May 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/104/5/
1207.

Nybroe:1985:BNC

[NAD^{+}85] O. Nybroe, M. Albrechtsen, J. Dahlin, D. Linnemann, J. M. Lyles, C. J. Møller, and E. Bock. Biosynthesis of the neural cell adhesion molecule: characterization of polypeptide C. Journal of Cell Biology, 101(6):2310–??, December 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/101/6/2310.

Noe:1986:CPP

[NADS86] B. D. Noe, P. C. Andrews, J. E. Dixon, and J. Spiess. Cotranslational and posttranslational proteolytic processing of preprosomatostatin–I in intact islet tissue. Journal of Cell Biology, 103(4):1205–??, October 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/103/4/1205.

Nguyen:1987:SSD

[NAS^{+}87] M. Nguyen, C. Argan, W. P. Sheffield, A. W. Bell, D. Shields, and G. C. Shore. A signal sequence domain essential for processing, but not import, of mitochondrial pre-ornithine carbamyl transferase. Journal of Cell Biology, 104(5):1193–??, May 1987. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/104/5/1193.

Neely:1988:SCP

[NB88a] M. D. Neely and K. Boekelheide. Sertoli cell processes have axoplasmic features: an ordered microtubule distribution and an abundant high molecular weight microtubule-associated protein (cytoplasmic dynein). Journal of Cell Biology, 107(5): 1767–??, November 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/107/5/1767.

Nickells:1988:RGP

[NB88b] R. W. Nickells and L. W. Browder. A role for glyceraldehyde-3-phosphate dehydrogenase in the development of thermotolerance in Xenopus laevis embryos. Journal of Cell Biology, 107(5): 1901–??, November 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/107/5/1901.

Nakayasu:1989:MRS

[NB89a] H. Nakayasu and R. Berezney. Mapping replicational sites in the eucaryotic cell nucleus. Journal of Cell Biology, 108(1):1–??, January 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/108/1/
1.

Nicchitta:1989:NSC

[NB89b] C. V. Nicchitta and G. Blobel. Nascent secretory chain binding and translocation are distinct processes: differentiation by chemical alkylation. Journal of Cell Biology, 108(3):789–??, March 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/108/3/
789.

Nakano:1988:MGS

[NBS88a] A. Nakano, D. Brada, and R. Schekman. A membrane glycoprotein, Sec12p, required for protein transport from the endoplasmic reticulum to the Golgi apparatus in yeast. Journal of Cell Biology, 107(3):851–??, September 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/107/3/851.

Nguyen:1988:PSB

[NBS88b] M. Nguyen, A. W. Bell, and G. C. Shore. Protein sorting between mitochondrial membranes specified by position of the stoptransfer domain. Journal of Cell Biology, 106(5):1499–??, May 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/106/5/
1499.

Nickells:1988:EHS

[NCB88] R. W. Nickells, M. J. Cavey, and L. W. Browder. The effects of heat shock on the morphology and protein synthesis of the epidermis of Xenopus laevis larvae. Journal of Cell Biology, 106 (3):905–??, March 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/106/3/905.

Navone:1989:MNB

[NDJ^{+}89] F. Navone, G. Di Gioia, R. Jahn, M. Browning, P. Greengard, and P. De Camilli. Microvesicles of the neurohypophysis are biochemically related to small synaptic vesicles of presynaptic nerve terminals. Journal of Cell Biology, 109(6):3425–??, December 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/109/6/
3425.

Newmeyer:1986:VTF

[NFF86] D. D. Newmeyer, D. R. Finlay, and D. J. Forbes. In vitro transport of a fluorescent nuclear protein and exclusion of non-nuclear proteins. Journal of Cell Biology, 103(6):2091–??, December 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/103/6/
2091.

Narusawa:1987:SMD

[NFI^{+}87] M. Narusawa, R. B. Fitzsimons, S. Izumo, B. Nadal-Ginard, N. A. Rubinstein, and A. M. Kelly. Slow myosin in developing rat skeletal muscle. Journal of Cell Biology, 104(3):447–??, March 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/104/3/
447.

Newman:1987:CNM

[NFN87] A. P. Newman and S. Ferro-Novick. Characterization of new mutants in the early part of the yeast secretory pathway isolated by a [3H]mannose suicide selection. Journal of Cell Biology, 105 (4):1587–??, October 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/4/1587.

Nave:1989:VPI

[NFW89] R. Nave, D. O. Fürst, and K. Weber. Visualization of the polarity of isolated titin molecules: a single globular head on a long thin rod as the M band anchoring domain? Journal of Cell Biology, 109(5):2177–??, November 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/109/5/2177.

Nicklas:1985:TLS

[NG85] R. B. Nicklas and G. W. Gordon. The total length of spindle microtubules depends on the number of chromosomes present. Journal of Cell Biology, 100(1):1–??, January 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/100/1/1.

Notario:1989:EFP

[NGI^{+}89] V. Notario, J. S. Gutkind, M. Imaizumi, S. Katamine, and K. C. Robbins. Expression of the fgr protooncogene product as a function of myelomonocytic cell maturation. Journal of Cell Biology, 109(6):3129–??, December 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/109/6/3129.

Nose:1985:CDP

[NGS85] P. S. Nose, L. C. Griffith, and H. Schulman. Ca^{2+}-dependent phosphorylation of tyrosine hydroxylase in PC12 cells. Journal of Cell Biology, 101(4):1182–??, October 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/101/4/1182.

Nakata:1987:CRH

[NH87] T. Nakata and N. Hirokawa. Cytoskeletal reorganization of human platelets after stimulation revealed by the quick-freeze deepetch technique. Journal of Cell Biology, 105(4):1771–??, October 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/105/4/
1771.

Nelson:1989:MCC

[NH89] W. J. Nelson and R. W. Hammerton. A membrane-cytoskeletal complex containing Na+,K+-ATPase, ankyrin, and fodrin in Madin–Darby canine kidney (MDCK) cells: implications for the biogenesis of epithelial cell polarity. Journal of Cell Biology, 108(3):893–??, March 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/108/3/893.

Nyman:1986:ILS

[NHH^{+}86] U. Nyman, H. Hallman, G. Hadlaczky, I. Pettersson, G. Sharp, and N. R. Ringertz. Intranuclear localization of snRNP antigens. Journal of Cell Biology, 102(1):137–??, January 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/102/1/137.

Novak-Hofer:1988:REE

[NHKE88] I. Novak-Hofer, W. Küng, and U. Eppenberger. Role of extracellular electrolytes in the activation of ribosomal protein S6 kinase by epidermal growth factor, insulin-like growth factor 1, and insulin in ZR-75-1 cells. Journal of Cell Biology, 106(2): 395–??, February 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/106/2/395.

Nagata:1985:CCM

[NHOY85] K. Nagata, M. J. Humphries, K. Olden, and K. M. Yamada. Collagen can modulate cell interactions with fibronectin. Journal of Cell Biology, 101(2):386–??, August 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/101/2/386.

Nicklas:1989:MPC

[Nic89] R. B. Nicklas. The motor for poleward chromosome movement in anaphase is in or near the kinetochore. Journal of Cell Biology, 109(5):2245–??, November 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/109/5/2245.

Norgauer:1989:IBC

[NJAS89] J. Norgauer, I. Just, K. Aktories, and L. A. Sklar. Influence of botulinum C2 toxin on F-actin and N-formyl peptide receptor dynamics in human neutrophils. Journal of Cell Biology, 109(3): 1133–??, September 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/3/1133.

Navone:1986:PPI

[NJD^{+}86] F. Navone, R. Jahn, G. Di Gioia, H. Stukenbrok, P. Greengard, and P. De Camilli. Protein p38: an integral membrane protein specific for small vesicles of neurons and neuroendocrine cells. Journal of Cell Biology, 103(6):2511–??, December 1986. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/103/6/2511.

Nash:1987:SUM

[NKA^{+}87] M. A. Nash, S. E. Kozak, L. M. Angerer, R. C. Angerer, H. Schatten, G. Schatten, and W. F. Marzluff. Sea urchin maternal and embryonic U1 RNAs are spatially segregated in early embryos. Journal of Cell Biology, 104(5):1133–??, May 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/104/5/1133.

Newman:1986:SCH

[NKMK86] P. J. Newman, Y. Kawai, R. R. Montgomery, and T. J. Kunicki. Synthesis by cultured human umbilical vein endothelial cells of two proteins structurally and immunologically related to platelet membrane glycoproteins IIb and IIIa. Journal of Cell Biology, 103(1):81–??, July 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/103/1/81.

Nelson:1985:PCM

[NL85] W. J. Nelson and E. Lazarides. Posttranslational control of membrane-skeleton (ankyrin and alpha beta-spectrin) assembly in early myogenesis. Journal of Cell Biology, 100(5):1726–??, May 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/100/5/
1726.

Nixon:1986:MFN

[NL86] R. A. Nixon and K. B. Logvinenko. Multiple fates of newly synthesized neurofilament proteins: evidence for a stationary neurofilament network distributed nonuniformly along axons of retinal ganglion cell neurons. Journal of Cell Biology, 102(2): 647–??, February 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/102/2/647.

Nusse:1988:DEH

[NL88] O. Nüsse and M. Lindau. The dynamics of exocytosis in human neutrophils. Journal of Cell Biology, 107(6):2117–??, December 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/107/6/
2117.

Noble:1989:MBD

[NLC89] M. Noble, S. A. Lewis, and N. J. Cowan. The microtubule binding domain of microtubule-associated protein MAP1B contains a repeated sequence motif unrelated to that of MAP2 and tau. Journal of Cell Biology, 109(6):3367–??, December 1989. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/109/6/3367.

Nakano:1989:NGB

[NM89] A. Nakano and M. Muramatsu. A novel GTP-binding protein, Sar1p, is involved in transport from the endoplasmic reticulum to the Golgi apparatus. Journal of Cell Biology, 109(6): 2677–??, December 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/6/2677.

Neufeld:1985:USD

[NMKS85] E. J. Neufeld, P. W. Majerus, C. M. Krueger, and J. E. Saffitz. Uptake and subcellular distribution of [3H]arachidonic acid in murine fibrosarcoma cells measured by electron microscope autoradiography. Journal of Cell Biology, 101(2):573–??, August 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/101/2/
573.

Nguyen:1986:LHN

[NMM^{+}86] C. Nguyen, M. G. Mattei, J. F. Mattei, M. J. Santoni, C. Goridis, and B. R. Jordan. Localization of the human NCAM gene to band q23 of chromosome 11: the third gene coding for a cell interaction molecule mapped to the distal portion of the long arm of chromosome 11. Journal of Cell Biology, 102(3):711–??, March 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/102/3/
711.

Neufeld:1988:EHB

[NMPG88] G. Neufeld, R. Mitchell, P. Ponte, and D. Gospodarowicz. Expression of human basic fibroblast growth factor cDNA in baby hamster kidney-derived cells results in autonomous cell growth. Journal of Cell Biology, 106(4):1385–??, April 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/106/4/1385.

Nicolas:1987:CBO

[NNJ^{+}87] M. T. Nicolas, G. Nicolas, C. H. Johnson, J. M. Bassot, and J. W. Hastings. Characterization of the bioluminescent organelles in Gonyaulax polyedra (dinoflagellates) after fast-freeze fixation and antiluciferase immunogold staining. Journal of Cell Biology, 105(2):723–??, August 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/2/723.

Napolitano:1985:BIU

[NPCL85] E. W. Napolitano, J. S. Pachter, S. S. Chin, and R. K. Liem. beta-internexin, a ubiquitous intermediate filament-associated protein. Journal of Cell Biology, 101(4):1323–??, October 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/101/4/1323.

Nicolet:1986:GAA

[NPRR86] M. Nicolet, M. Pinçon-Raymond, and F. Rieger. Globular and asymmetric acetylcholinesterase in frog muscle basal lamina sheaths. Journal of Cell Biology, 102(3):762–??, March 1986.

CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/102/3/762.

Navre:1988:GFR

[NR88] M. Navre and G. M. Ringold. A growth factor-repressible gene associated with protein kinase C-mediated inhibition of adipocyte differentiation. Journal of Cell Biology, 107(1):279–??, July 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/107/1/279.

Navre:1989:DEF

[NR89a] M. Navre and G. M. Ringold. Differential effects of fibroblast growth factor and tumor promoters on the initiation and maintenance of adipocyte differentiation. Journal of Cell Biology, 109 (4):1857–??, October 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/4/1857.

Noda:1989:TRO

[NR89b] M. Noda and G. A. Rodan. Transcriptional regulation of osteopontin production in rat osteoblast-like cells by parathyroid hormone. Journal of Cell Biology, 108(2):713–??, February 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/108/2/713.

Ng:1989:IMT

[NRL89a] D. T. Ng, R. E. Randall, and R. A. Lamb. Intracellular maturation and transport of the SV5 type II glycoprotein hemagglutinin-neuraminidase: specific and transient association with GRP78-BiP in the endoplasmic reticulum and extensive internalization from the cell surface. Journal of Cell Biology, 109 (6):3273–??, December 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/6/3273.

Noegel:1989:DGF

[NRL^{+}89b] A. A. Noegel, S. Rapp, F. Lottspeich, M. Schleicher, and M. Stewart. The Dictyostelium gelation factor shares a putative actin binding site with alpha-actinins and dystrophin and also has a rod domain containing six 100-residue motifs that appear to have a cross-beta conformation. Journal of Cell Biology, 109(2):607–??, August 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/2/607.

Nadin:1989:SIV

[NRTE89] C. Y. Nadin, J. Rogers, S. Tomlinson, and J. M. Edwardson. A specific interaction in vitro between pancreatic zymogen granules and plasma membranes: stimulation by G-protein activators but not by Ca^{2+}. Journal of Cell Biology, 109(6):2801–??, December 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/109/6/
2801.

Nathan:1989:CIR

[NSF^{+}89] C. Nathan, S. Srimal, C. Farber, E. Sanchez, L. Kabbash, A. Asch, J. Gailit, and S. D. Wright. Cytokine-induced respiratory burst of human neutrophils: dependence on extracellular matrix proteins and CD11/CD18 integrins. Journal of Cell Biology, 109(3):1341–??, September 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/109/3/1341.

Nitkin:1987:IAS

[NSM^{+}87] R. M. Nitkin, M. A. Smith, C. Magill, J. R. Fallon, Y. M. Yao, B. G. Wallace, and U. J. McMahan. Identification of agrin, a synaptic organizing protein from Torpedo electric organ. Journal of Cell Biology, 105(6):2471–??, December 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/105/6/2471.

Nagata:1986:MCB

[NSY86] K. Nagata, S. Saga, and K. M. Yamada. A major collagenbinding protein of chick embryo fibroblasts is a novel heat shock protein. Journal of Cell Biology, 103(1):223–??, July 1986. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/103/1/223.

Nakamura:1985:CCC

[NT85] S. Nakamura and S. L. Tamm. Calcium control of ciliary reversal in ionophore-treated and ATP-reactivated comb plates of ctenophores. Journal of Cell Biology, 100(5):1447–??, May 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/100/5/

Nose:1986:NCC

[NT86] A. Nose and M. Takeichi. A novel cadherin cell adhesion molecule: its expression patterns associated with implantation and organogenesis of mouse embryos. Journal of Cell Biology, 103 (6):2649–??, December 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/103/6/2649.

Neugebauer:1988:CNI

[NTLR88] K. M. Neugebauer, K. J. Tomaselli, J. Lilien, and L. F. Reichardt. N-cadherin, NCAM, and integrins promote retinal neurite outgrowth on astrocytes in vitro. Journal of Cell Biology, 107(3): 1177–??, September 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/107/3/1177.

Novikoff:1985:ESR

[NTNT85] P. M. Novikoff, O. Touster, A. B. Novikoff, and D. P. Tulsiani. Effects of swainsonine on rat liver and kidney: biochemical and morphological studies. Journal of Cell Biology, 101 (2):339–??, August 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/101/2/339.

Nelson:1986:DMS

[NV86] W. J. Nelson and P. J. Veshnock. Dynamics of membraneskeleton (fodrin) organization during development of polarity in Madin–Darby canine kidney epithelial cells. Journal of Cell Biology, 103(5):1751–??, November 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/103/5/1751.

Nelson:1987:MFM

[NV87] W. J. Nelson and P. J. Veshnock. Modulation of fodrin (membrane skeleton) stability by cell–cell contact in Madin–Darby canine kidney epithelial cells. Journal of Cell Biology, 104 (6):1527–??, June 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/104/6/1527.

Noda:1989:FGF

[NV89] M. Noda and R. Vogel. Fibroblast growth factor enhances type beta 1 transforming growth factor gene expression in osteoblastlike cells. Journal of Cell Biology, 109(5):2529–??, November 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/109/5/
2529.

Neefjes:1988:RGD

[NVB^{+}88] J. J. Neefjes, J. M. Verkerk, H. J. Broxterman, G. A. van der Marel, J. H. van Boom, and H. L. Ploegh. Recycling glycoproteins do not return to the cis–Golgi. Journal of Cell Biology, 107(1):79–??, July 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/107/1/79.

Nomura:1988:DEO

[NWE^{+}88] S. Nomura, A. J. Wills, D. R. Edwards, J. K. Heath, and B. L. Hogan. Developmental expression of 2ar (osteopontin) and SPARC (osteonectin) RNA as revealed by in situ hybridization. Journal of Cell Biology, 106(2):441–??, February 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/106/2/441.

Neighbors:1988:LKC

[NWM88] B. W. Neighbors, R. C. Williams, and J. R. McIntosh. Localization of kinesin in cultured cells. Journal of Cell Biology, 106 (4):1193–??, April 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/106/4/1193.

Nachmias:1987:LPB

[NYG87] V. T. Nachmias, K. Yoshida, and M. C. Glennon. Lowering pH in blood platelets dissociates myosin phosphorylation from shape change and myosin association with the cytoskeleton. Journal of Cell Biology, 105(4):1761–??, October 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/105/4/1761.

Noda:1987:HLM

[NYRK87] M. Noda, K. Yoon, G. A. Rodan, and D. E. Koppel. High lateral mobility of endogenous and transfected alkaline phosphatase: a phosphatidylinositol-anchored membrane protein. Journal of Cell Biology, 105(4):1671–??, October 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/105/4/1671.

Netland:1985:MPB

[NZ85] P. A. Netland and B. R. Zetter. Metastatic potential of B16 melanoma cells after in vitro selection for organ-specific adherence. Journal of Cell Biology, 101(3):720–??, September 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/101/3/720.

ODonnell:1988:MUD

[OB88] P. T. O’Donnell and S. I. Bernstein. Molecular and ultrastructural defects in a Drosophila myosin heavy chain mutant: differential effects on muscle function produced by similar thick filament abnormalities. Journal of Cell Biology, 107(6):2601–??, December 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
107/6/2601.

Oursler:1985:IOS

[OBCO85] M. J. Oursler, L. V. Bell, B. Clevinger, and P. Osdoby. Identification of osteoclast-specific monoclonal antibodies. Journal of Cell Biology, 100(5):1592–??, May 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/100/5/1592.

OKeefe:1987:CIA

[OBH87] E. J. O’Keefe, R. A. Briggaman, and B. Herman. Calciuminduced assembly of adherens junctions in keratinocytes. Journal of Cell Biology, 105(2):807–??, August 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/105/2/807.

Ohkuma:1986:EBS

[OCP86] S. Ohkuma, J. Chudzik, and B. Poole. The effects of basic substances and acidic ionophores on the digestion of exogenous and endogenous proteins in mouse peritoneal macrophages. Journal of Cell Biology, 102(3):959–??, March 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/102/3/959.

Orlidge:1987:ICE

[OD87] A. Orlidge and P. A. D’Amore. Inhibition of capillary endothelial cell growth by pericytes and smooth muscle cells. Journal of Cell Biology, 105(3):1455–??, September 1987. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/105/3/1455.

OShea:1988:UDE

[OD88] K. S. O’Shea and V. M. Dixit. Unique distribution of the extracellular matrix component thrombospondin in the developing mouse embryo. Journal of Cell Biology, 107(6):2737–??, December 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/107/6/
2737.

Olsson:1985:THC

[ODD85] L. Olsson, C. Due, and M. Diamant. Treatment of human cell lines with 5-azacytidine may result in profound alterations in clonogenicity and growth rate. Journal of Cell Biology, 100 (2):508–??, February 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/100/2/508.

Owaribe:1985:IAC

[OE85] K. Owaribe and G. Eguchi. Increase in actin contents and elongation of apical projections in retinal pigmented epithelial cells during development of the chicken eye. Journal of Cell Biology, 101(2):590–??, August 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/101/2/590.

Otto:1987:CRF

[OGF87] J. M. Otto, H. E. Grenett, and G. M. Fuller. The coordinated regulation of fibrinogen gene transcription by hepatocytestimulating factor and dexamethasone. Journal of Cell Biology, 105(3):1067–??, September 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/105/3/1067.

Ouellette:1989:DRC

[OGJ^{+}89] A. J. Ouellette, R. M. Greco, M. James, D. Frederick, J. Naftilan, and J. T. Fallon. Developmental regulation of cryptdin, a corticostatin/defensin precursor mRNA in mouse small intestinal crypt epithelium. Journal of Cell Biology, 108(5):1687–??, May 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/108/5/

Owens:1988:TGF

[OGYK88] G. K. Owens, A. A. Geisterfer, Y. W. Yang, and A. Komoriya. Transforming growth factor-beta-induced growth inhibition and cellular hypertrophy in cultured vascular smooth muscle cells. Journal of Cell Biology, 107(2):771–??, August 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/107/2/771.

Okabe:1988:MDN

[OH88a] S. Okabe and N. Hirokawa. Microtubule dynamics in nerve cells: analysis using microinjection of biotinylated tubulin into PC12 cells. Journal of Cell Biology, 107(2):651–??, August 1988. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/107/2/651.

Olwin:1988:CSF

[OH88b] B. B. Olwin and S. D. Hauschka. Cell surface fibroblast growth factor and epidermal growth factor receptors are permanently lost during skeletal muscle terminal differentiation in culture. Journal of Cell Biology, 107(2):761–??, August 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/107/2/761.

Onuma:1988:EFD

[OH88c] E. K. Onuma and S. W. Hui. Electric field-directed cell shape changes, displacement, and cytoskeletal reorganization are calcium dependent. Journal of Cell Biology, 106(6):2067–??, June 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/106/6/
2067.

Okabe:1989:ITB

[OH89] S. Okabe and N. Hirokawa. Incorporation and turnover of biotinlabeled actin microinjected into fibroblastic cells: an immunoelectron microscopic study. Journal of Cell Biology, 109(4): 1581–??, October 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/4/1581.

Oyler:1989:INS

[OHH^{+}89] G. A. Oyler, G. A. Higgins, R. A. Hart, E. Battenberg, M. Billingsley, F. E. Bloom, and M. C. Wilson. The identification of a novel synaptosomal-associated protein, SNAP-25, differentially expressed by neuronal subpopulations. Journal of Cell Biology, 109(6):3039–??, December 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/109/6/3039.

Oberdorf:1986:CUR

[OHK86] J. A. Oberdorf, J. F. Head, and B. Kaminer. Calcium uptake and release by isolated cortices and microsomes from the unfertilized egg of the sea urchin Strongylocentrotus droebachiensis. Journal of Cell Biology, 102(6):2205–??, June 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/102/6/2205.

Oldberg:1989:RBS

[OJHAH89] A. Oldberg, B. Jirskog-Hed, S. Axelsson, and D. Heinegård. Regulation of bone sialoprotein mRNA by steroid hormones. Journal of Cell Biology, 109(6):3183–??, December 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/109/6/3183.

Okagaki:1986:MSM

[OK86] T. Okagaki and R. Kamiya. Microtubule sliding in mutant Chlamydomonas axonemes devoid of outer or inner dynein arms. Journal of Cell Biology, 103(5):1895–??, November 1986. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/103/5/1895.

Otey:1986:IGI

[OKLB86] C. A. Otey, M. H. Kalnoski, J. L. Lessard, and J. C. Bulinski. Immunolocalization of the gamma isoform of nonmuscle actin in cultured cells. Journal of Cell Biology, 102(5):1726–??, May 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/102/5/
1726.

Owens:1986:ESM

[OLGT86] G. K. Owens, A. Loeb, D. Gordon, and M. M. Thompson. Expression of smooth muscle-specific alpha-isoactin in cultured vascular smooth muscle cells: relationship between growth and cytodifferentiation. Journal of Cell Biology, 102(2):343–??, February 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/102/2/

Osmani:1987:RML

[OMM87] S. A. Osmani, G. S. May, and N. R. Morris. Regulation of the mRNA levels of nimA, a gene required for the G2-M transition in Aspergillus nidulans. Journal of Cell Biology, 104(6):1495–??, June 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/104/6/
1495.

OGuin:1989:ABP

[OMS89] W. M. O’Guin, M. Manabe, and T. T. Sun. Association of a basic 25K protein with membrane coating granules of human epidermis. Journal of Cell Biology, 109(5):2313–??, November 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/109/5/
2313.

Ono:1985:MRM

[OMY^{+}85] M. Ono, K. Mifune, A. Yoshimura, S. Ohnishi, and M. Kuwano. Monensin-resistant mouse Balb/3T3 cell mutant with aberrant penetration of vesicular stomatitis virus. Journal of Cell Biology, 101(1):60–??, July 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/101/1/60.

Olins:1987:VDS

[OO87] D. E. Olins and A. L. Olins. In vitro DNA synthesis in the macronuclear replication band of Euplotes eurystomus. Journal of Cell Biology, 104(5):1125–??, May 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/104/5/1125.

Olins:1989:PCN

[OOCT89] D. E. Olins, A. L. Olins, L. H. Cacheiro, and E. M. Tan. Proliferating cell nuclear antigen/cyclin in the ciliate Euplotes eurystomus: localization in the replication band and in micronuclei. Journal of Cell Biology, 109(4):1399–??, October 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/109/4/1399.

Ohtsubo:1989:RPR

[OON89] M. Ohtsubo, H. Okazaki, and T. Nishimoto. The RCC1 protein, a regulator for the onset of chromosome condensation locates in the nucleus and binds to DNA. Journal of Cell Biology, 109 (4):1389–??, October 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/4/1389.

Okita:1985:AGI

[OPN^{+}85] J. R. Okita, D. Pidard, P. J. Newman, R. R. Montgomery, and T. J. Kunicki. On the association of glycoprotein Ib and actinbinding protein in human platelets. Journal of Cell Biology, 100 (1):317–??, January 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/100/1/317.

Oakley:1985:MNM

[OR85] B. R. Oakley and J. E. Rinehart. Mitochondria and nuclei move by different mechanisms in Aspergillus nidulans. Journal of Cell Biology, 101(6):2392–??, December 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/101/6/2392.

Orci:1986:CPI

[ORA^{+}86] L. Orci, M. Ravazzola, M. Amherdt, O. Madsen, A. Perrelet, J. D. Vassalli, and R. G. Anderson. Conversion of proinsulin to insulin occurs coordinately with acidification of maturing secretory vesicles. Journal of Cell Biology, 103(6):2273–??, December 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/103/6/
2273.

Ojakian:1988:PDA

[OS88a] G. K. Ojakian and R. Schwimmer. The polarized distribution of an apical cell surface glycoprotein is maintained by interactions with the cytoskeleton of Madin–Darby canine kidney cells. Journal of Cell Biology, 107(6):2377–??, December 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/107/6/2377.

Omann:1988:RNS

[OS88b] G. M. Omann and L. A. Sklar. Response of neutrophils to stimulus infusion: differential sensitivity of cytoskeletal activation and oxidant production. Journal of Cell Biology, 107(3): 951–??, September 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/107/3/951.

Oettling:1985:MRC

[OSD85] G. Oettling, H. Schmidt, and U. Drews. The muscarinic receptor of chick embryo cells: correlation between ligand binding and calcium mobilization. Journal of Cell Biology, 100(4):1073–??, April 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/100/4/
1073.

Olson:1986:RMD

[OSH^{+}86] E. N. Olson, E. Sternberg, J. S. Hu, G. Spizz, and C. Wilcox. Regulation of myogenic differentiation by type beta transforming growth factor. Journal of Cell Biology, 103(5):1799–??, November 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/103/5/
1799.

Oliver:1985:SFD

[OSP^{+}85] J. M. Oliver, J. C. Seagrave, J. R. Pfeiffer, M. L. Feibig, and G. G. Deanin. Surface functions during mitosis in rat basophilic leukemia cells. Journal of Cell Biology, 101(6):2156–??, December 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/101/6/
2156.

Ossowski:1988:VIM

[Oss88] L. Ossowski. In vivo invasion of modified chorioallantoic membrane by tumor cells: the role of cell surface-bound urokinase. Journal of Cell Biology, 107(6):2437–??, December 1988. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/107/6/2437.

Olmsted:1989:CCD

[OSS^{+}89] J. B. Olmsted, D. L. Stemple, W. M. Saxton, B. W. Neighbors, and J. R. McIntosh. Cell cycle-dependent changes in the dynamics of MAP 2 and MAP 4 in cultured cells. Journal of Cell Biology, 109(1):211–??, July 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/109/1/211.

Olsnes:1986:PRA

[OTS86] S. Olsnes, T. I. Tønnessen, and K. Sandvig. pH-regulated anion antiport in nucleated mammalian cells. Journal of Cell Biology, 102(3):967–??, March 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/102/3/967.

Oka:1986:BSH

[OW86] J. A. Oka and P. H. Weigel. Binding and spreading of hepatocytes on synthetic galactose culture surfaces occur as distinct and separable threshold responses. Journal of Cell Biology, 103(3): 1055–??, September 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/103/3/1055.

Owen:1985:ETI

[Owe85] N. E. Owen. Effect of TPA on ion fluxes and DNA synthesis in vascular smooth muscle cells. Journal of Cell Biology, 101 (2):454–??, August 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/101/2/454.

Owen:1986:ECH

[Owe86] N. E. Owen. Effect of catecholamines on Na/H exchange in vascular smooth muscle cells. Journal of Cell Biology, 103(5): 2053–??, November 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/103/5/2053.

Poruchynsky:1988:PSD

[PA88] M. S. Poruchynsky and P. H. Atkinson. Primary sequence domains required for the retention of rotavirus VP7 in the endoplasmic reticulum. Journal of Cell Biology, 107(5):1697–??, November 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/107/5/
1697.

Persson:1988:DAS

[PAF88] R. Persson, E. Ahlström, and E. Fries. Differential arrest of secretory protein transport in cultured rat hepatocytes by azide treatment. Journal of Cell Biology, 107(6):2503–??, December 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/107/6/
2503.

Perrot-Applanat:1986:ULP

[PAGPLM86] M. Perrot-Applanat, M. T. Groyer-Picard, F. Logeat, and E. Milgrom. Ultrastructural localization of the progesterone receptor by an immunogold method: effect of hormone administration. Journal of Cell Biology, 102(4):1191–??, April 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/102/4/1191.

Peattie:1989:ULG

[PAHC89] D. A. Peattie, R. A. Alonso, A. Hein, and J. P. Caulfield. Ultrastructural localization of giardins to the edges of disk microribbons of Giarida lamblia and the nucleotide and deduced protein sequence of alpha giardin. Journal of Cell Biology, 109(5): 2323–??, November 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/5/2323.

Palevitz:1987:APB

[Pal87] B. A. Palevitz. Actin in the preprophase band of Allium cepa. Journal of Cell Biology, 104(6):1515–??, June 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/104/6/1515.

Paul:1986:MCC

[Pau86] D. L. Paul. Molecular cloning of cDNA for rat liver gap junction protein. Journal of Cell Biology, 103(1):123–??, July 1986. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/103/1/123.

Pavlath:1986:EMG

[PB86] G. K. Pavlath and H. M. Blau. Expression of muscle genes in heterokaryons depends on gene dosage. Journal of Cell Biology, 102(1):124–??, January 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/102/1/124.

Pumplin:1987:DRA

[PB87] D. W. Pumplin and R. J. Bloch. Disruption and reformation of the acetylcholine receptor clusters of cultured rat myotubes occur in two distinct stages. Journal of Cell Biology, 104(1):97–??, January 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/104/1/
97.

Peng:1986:BIA

[PBB86] I. Peng, L. I. Binder, and M. M. Black. Biochemical and immunological analyses of cytoskeletal domains of neurons. Journal of Cell Biology, 102(1):252–??, January 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/102/1/252.

Papasozomenos:1985:MAP

[PBBP85] S. C. Papasozomenos, L. I. Binder, P. K. Bender, and M. R. Payne. Microtubule-associated protein 2 within axons of spinal motor neurons: associations with microtubules and neurofilaments in normal and beta,beta’-iminodipropionitrile-treated axons. Journal of Cell Biology, 100(1):74–??, January 1985. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/100/1/74.

Pesacreta:1989:DSM

[PBD^{+}89] T. C. Pesacreta, T. J. Byers, R. Dubreuil, D. P. Kiehart, and D. Branton. Drosophila spectrin: the membrane skeleton during embryogenesis. Journal of Cell Biology, 108(5):1697–??, May 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/108/5/
1697.

Penn:1987:SADb

[PBH^{+}87] E. J. Penn, I. D. Burdett, C. Hobson, A. I. Magee, and D. A. Rees. Structure and assembly of desmosome junctions: biosynthesis and turnover of the major desmosome components of Madin–Darby canine kidney cells in low calcium medium. Journal of Cell Biology, 105(5):2327–??, November 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/105/5/2327.

Polla:1988:DDI

[PBK88] B. S. Polla, J. V. Bonventre, and S. M. Krane. 1,25dihydroxyvitamin D3 increases the toxicity of hydrogen peroxide in the human monocytic line U937: the role of calcium and heat shock. Journal of Cell Biology, 107(1):373–??, July 1988. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/107/1/373.

Pitas:1985:UCM

[PBMB85] R. E. Pitas, J. Boyles, R. W. Mahley, and D. M. Bissell. Uptake of chemically modified low density lipoproteins in vivo is mediated by specific endothelial cells. Journal of Cell Biology, 100(1):103–??, January 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/100/1/103.

Plant:1985:CUI

[PBS85] A. L. Plant, D. M. Benson, and L. C. Smith. Cellular uptake and intracellular localization of benzo(a)pyrene by digital fluorescence imaging microscopy. Journal of Cell Biology, 100 (4):1295–??, April 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/100/4/1295.

Poliard:1986:CAS

[PBT^{+}86] A. M. Poliard, D. Bernuau, I. Tournier, L. G. Legrès, D. Schoevaert, G. Feldmann, and J. M. Sala-Trepat. Cellular analysis by in situ hybridization and immunoperoxidase of alpha-fetoprotein and albumin gene expression in rat liver during the perinatal period. Journal of Cell Biology, 103(3):777–??, September 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/103/3/777.

Payne:1988:PTV

[PBvTS88] G. S. Payne, D. Baker, E. van Tuinen, and R. Schekman. Protein transport to the vacuole and receptor-mediated endocytosis by clathrin heavy chain-deficient yeast. Journal of Cell Biology, 106(5):1453–??, May 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/106/5/1453.

Pittenger:1985:RAC

[PC85] M. F. Pittenger and D. W. Cleveland. Retention of autoregulatory control of tubulin synthesis in cytoplasts: demonstration of a cytoplasmic mechanism that regulates the level of tubulin expression. Journal of Cell Biology, 101(5):1941–??, November 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/101/5/
1941.

Picard:1989:IPP

[PCBD89] A. Picard, J. P. Capony, D. L. Brautigan, and M. Dorée. Involvement of protein phosphatases 1 and 2A in the control of M phasepromoting factor activity in starfish. Journal of Cell Biology, 109 (6):3347–??, December 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/6/3347.

Pechak:1985:EMC

[PCC85] D. G. Pechak, D. A. Carrino, and A. I. Caplan. Electron microscopic characterization of chick embryonic skeletal muscle proteoglycans. Journal of Cell Biology, 100(5):1767–??, May 1985. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/100/5/1767.

Parry:1987:RDP

[PCK^{+}87] G. Parry, B. Cullen, C. S. Kaetzel, R. Kramer, and L. Moss. Regulation of differentiation and polarized secretion in mammary epithelial cells maintained in culture: extracellular matrix and membrane polarity influences. Journal of Cell Biology, 105(5): 2043–??, November 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/5/2043.

Peters:1985:EPV

[PCP85] K. R. Peters, W. W. Carley, and G. E. Palade. Endothelial plasmalemmal vesicles have a characteristic striped bipolar surface structure. Journal of Cell Biology, 101(6):2233–??, December 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/101/6/
2233.

Paganetti:1988:GIC

[PCS88] P. A. Paganetti, P. Caroni, and M. E. Schwab. Glioblastoma infiltration into central nervous system tissue in vitro: involvement of a metalloprotease. Journal of Cell Biology, 107(6): 2281–??, December 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/107/6/2281.

Pasternak:1985:LMR

[PE85] C. Pasternak and E. L. Elson. Lymphocyte mechanical response triggered by cross-linking surface receptors. Journal of Cell Biology, 100(3):860–??, March 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/100/3/860.

Peng:1985:APP

[PF85a] H. B. Peng and S. C. Froehner. Association of the postsynaptic 43K protein with newly formed acetylcholine receptor clusters in cultured muscle cells. Journal of Cell Biology, 100(5):1698–??, May 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/100/5/
1698.

Piperno:1985:MAS

[PF85b] G. Piperno and M. T. Fuller. Monoclonal antibodies specific for an acetylated form of alpha-tubulin recognize the antigen in cilia and flagella from a variety of organisms. Journal of Cell Biology, 101(6):2085–??, December 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/101/6/2085.

Payne:1987:ITR

[PF87] R. Payne and A. Fein. Inositol 1,4,5 trisphosphate releases calcium from specialized sites within Limulus photoreceptors. Journal of Cell Biology, 104(4):933–??, January 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/104/4/933.

Pariset:1989:DES

[PFD^{+}89] C. Pariset, J. Feinberg, J. L. Dacheux, O. Oyen, T. Jahnsen, and S. Weinman. Differential expression and subcellular localization for subunits of cAMP-dependent protein kinase during ram spermatogenesis. Journal of Cell Biology, 109(3):1195–??, September 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/109/3/
1195.

Pfeffer:1987:ECM

[Pfe87] S. R. Pfeffer. The endosomal concentration of a mannose 6phosphate receptor is unchanged in the absence of ligand synthesis. Journal of Cell Biology, 105(1):229–??, July 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/105/1/229.

Plow:1986:PSC

[PFPM86] E. F. Plow, D. E. Freaney, J. Plescia, and L. A. Miles. The plasminogen system and cell surfaces: evidence for plasminogen and urokinase receptors on the same cell type. Journal of Cell Biology, 103(6):2411–??, December 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/103/6/2411.

Pasternak:1989:IVI

[PFRS89] C. Pasternak, P. F. Flicker, S. Ravid, and J. A. Spudich. Intermolecular versus intramolecular interactions of Dictyostelium myosin: possible regulation by heavy chain phosphorylation. Journal of Cell Biology, 109(1):203–??, July 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/109/1/203.

Pfeiffer:1985:ISB

[PFS85] S. Pfeiffer, S. D. Fuller, and K. Simons. Intracellular sorting and basolateral appearance of the G protein of vesicular stomatitis virus in Madin–Darby canine kidney cells. Journal of Cell Biology, 101(2):470–??, August 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/101/2/470.

Pfeffer:1989:HAC

[PFTV89] U. Pfeffer, N. Ferrari, F. Tosetti, and G. Vidali. Histone acetylation in conjugating Tetrahymena thermophila. Journal of Cell Biology, 109(3):1007–??, September 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/109/3/1007.

Pagh:1986:MAB

[PG86] K. Pagh and G. Gerisch. Monoclonal antibodies binding to the tail of Dictyostelium discoideum myosin: their effects on antiparallel and parallel assembly and actin-activated ATPase activity. Journal of Cell Biology, 103(4):1527–??, October 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/103/4/1527.

Pasquale:1987:CAF

[PG87] S. M. Pasquale and U. W. Goodenough. Cyclic AMP functions as a primary sexual signal in gametes of Chlamydomonas reinhardtii. Journal of Cell Biology, 105(5):2279–??, November 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/105/5/
2279.

Putnoky:1988:RFG

[PGH^{+}88] P. Putnoky, E. Grosskopf, D. T. Ha, G. B. Kiss, and A. Kondorosi. Rhizobium fix genes mediate at least two communication steps in symbiotic nodule development. Journal of Cell Biology, 106(3):597–??, March 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/106/3/597.

Papadopoulos:1989:ICP

[PH89] V. Papadopoulos and P. F. Hall. Isolation and characterization of protein kinase C from Y-1 adrenal cell cytoskeleton. Journal of Cell Biology, 108(2):553–??, February 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/108/2/553.

Parks:1989:TDB

[PHL89] G. D. Parks, J. D. Hull, and R. A. Lamb. Transposition of domains between the M2 and HN viral membrane proteins results in polypeptides which can adopt more than one membrane orientation. Journal of Cell Biology, 109(5):2023–??, November 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/109/5/
2023.

Primakoff:1985:RMS

[PHM85] P. Primakoff, H. Hyatt, and D. G. Myles. A role for the migrating sperm surface antigen PH-20 in guinea pig sperm binding to the egg zona pellucida. Journal of Cell Biology, 101(6): 2239–??, December 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/101/6/2239.

Pollanen:1988:ULP

[PHN^{+}88] J. Pöllänen, K. Hedman, L. S. Nielsen, K. Danø, and A. Vaheri. Ultrastructural localization of plasma membrane-associated urokinase-type plasminogen activator at focal contacts. Journal of Cell Biology, 106(1):87–??, January 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/106/1/87.

Pfeffer:1986:BSA

[PHR86] S. R. Pfeffer, T. Huima, and C. M. Redman. Biosynthesis of spectrin and its assembly into the cytoskeletal system of Friend erythroleukemia cells. Journal of Cell Biology, 103(1):103–??, July 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/103/1/

Penn:1987:SADa

[PHRM87] E. J. Penn, C. Hobson, D. A. Rees, and A. I. Magee. Structure and assembly of desmosome junctions: biosynthesis, processing, and transport of the major protein and glycoprotein components in cultured epithelial cells. Journal of Cell Biology, 105(1):57–??, July 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/105/1/
57.

Primakoff:1987:IPS

[PHTK87] P. Primakoff, H. Hyatt, and J. Tredick-Kline. Identification and purification of a sperm surface protein with a potential role in sperm-egg membrane fusion. Journal of Cell Biology, 104 (1):141–??, January 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/104/1/141.

Piperno:1988:ISD

[Pip88] G. Piperno. Isolation of a sixth dynein subunit adenosine triphosphatase of Chlamydomonas axonemes. Journal of Cell Biology, 106(1):133–??, January 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/106/1/133.

Perris:1987:ANC

[PJ87] R. Perris and S. Johansson. Amphibian neural crest cell migration on purified extracellular matrix components: a chondroitin sulfate proteoglycan inhibits locomotion on fibronectin substrates. Journal of Cell Biology, 105(6):2511–??, December 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/105/6/
2511.

Paiement:1989:RGA

[PJFB89] J. Paiement, M. Jolicoeur, A. Fazel, and J. J. Bergeron. Reconstitution of the Golgi apparatus after microinjection of rat liver Golgi fragments into Xenopus oocytes. Journal of Cell Biology, 108(4):1257–??, April 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/108/4/1257.

Pizzey:1988:RCI

[PJW88] J. A. Pizzey, G. E. Jones, and F. S. Walsh. Requirements for the Ca^{2+}-independent component in the initial intercellular adhesion of C2 myoblasts. Journal of Cell Biology, 107(6):2307–??, December 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/107/6/
2307.

Palmer:1989:DCM

[PKK89] R. E. Palmer, M. Koval, and D. Koshland. The dynamics of chromosome movement in the budding yeast Saccharomyces cerevisiae. Journal of Cell Biology, 109(6):3355–??, December 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/109/6/3355.

Picone:1989:RUR

[PKN^{+}89] R. Picone, E. L. Kajtaniak, L. S. Nielsen, N. Behrendt, M. R. Mastronicola, M. V. Cubellis, M. P. Stoppelli, S. Pedersen, K. Danø, and F. Blasi. Regulation of urokinase receptors in monocytelike U937 cells by phorbol ester phorbol myristate acetate. Journal of Cell Biology, 108(2):693–??, February 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/108/2/693.

Pennell:1989:FAP

[PKS^{+}89] R. I. Pennell, J. P. Knox, G. N. Scofield, R. R. Selvendran, and K. Roberts. A family of abundant plasma membrane-associated glycoproteins related to the arabinogalactan proteins is unique to flowering plants. Journal of Cell Biology, 108(5):1967–??, May 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/108/5/
1967.

Pachter:1985:AIK

[PL85a] J. S. Pachter and R. K. Liem. alpha-internexin, a 66-kD intermediate filament-binding protein from mammalian central nervous tissues. Journal of Cell Biology, 101(4):1316–??, October 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/101/4/

Perara:1985:FAT

[PL85b] E. Perara and V. R. Lingappa. A former amino terminal signal sequence engineered to an internal location directs translocation of both flanking protein domains. Journal of Cell Biology, 101 (6):2292–??, December 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/101/6/2292.

Patel:1986:FRM

[PL86] V. P. Patel and H. F. Lodish. The fibronectin receptor on mammalian erythroid precursor cells: characterization and developmental regulation. Journal of Cell Biology, 102(2):449–??, February 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/102/2/
449.

Patel:1987:FMR

[PL87] V. P. Patel and H. F. Lodish. A fibronectin matrix is required for differentiation of murine erythroleukemia cells into reticulocytes. Journal of Cell Biology, 105(6):3105–??, December 1987. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/105/6/3105.

Pathak:1986:BCE

[PLA86] R. K. Pathak, K. L. Luskey, and R. G. Anderson. Biogenesis of the crystalloid endoplasmic reticulum in UT-1 cells: evidence that newly formed endoplasmic reticulum emerges from the nuclear envelope. Journal of Cell Biology, 102(6):2158–??, June 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/102/6/
2158.

Piperno:1987:MCA

[PLC87] G. Piperno, M. LeDizet, and X. J. Chang. Microtubules containing acetylated alpha-tubulin in mammalian cells in culture. Journal of Cell Biology, 104(2):289–??, February 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/104/2/289.

Padh:1989:PAV

[PLS89] H. Padh, M. Lavasa, and T. L. Steck. Prelysosomal acidic vacuoles in Dictyostelium discoideum. Journal of Cell Biology, 108 (3):865–??, March 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/108/3/865.

Peters:1987:LCS

[PM87a] D. M. Peters and D. F. Mosher. Localization of cell surface sites involved in fibronectin fibrillogenesis. Journal of Cell Biology, 104(1):121–??, January 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/104/1/121.

Petrucci:1987:SAB

[PM87b] T. C. Petrucci and J. S. Morrow. Synapsin I: an actin-bundling protein under phosphorylation control. Journal of Cell Biology, 105(3):1355–??, September 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/105/3/1355.

Pathak:1988:ILM

[PMC^{+}88] R. K. Pathak, R. K. Merkle, R. D. Cummings, J. L. Goldstein, M. S. Brown, and R. G. Anderson. Immunocytochemical localization of mutant low density lipoprotein receptors that fail to reach the Golgi complex. Journal of Cell Biology, 106(6):1831–??, June 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/106/6/
1831.

Preston:1987:AEF

[PMJ87] R. A. Preston, R. F. Murphy, and E. W. Jones. Apparent endocytosis of fluorescein isothiocyanate-conjugated dextran by Saccharomyces cerevisiae reflects uptake of low molecular weight impurities, not dextran. Journal of Cell Biology, 105(5): 1981–??, November 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/5/1981.

Pierce:1989:PDG

[PML^{+}89] G. F. Pierce, T. A. Mustoe, J. Lingelbach, V. R. Masakowski, G. L. Griffin, R. M. Senior, and T. F. Deuel. Platelet-derived growth factor and transforming growth factor-beta enhance tissue repair activities by unique mechanisms. Journal of Cell Biology, 109(1):429–??, July 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/109/1/429.

Puddington:1986:CDC

[PMR86] L. Puddington, C. E. Machamer, and J. K. Rose. Cytoplasmic domains of cellular and viral integral membrane proteins substitute for the cytoplasmic domain of the vesicular stomatitis virus glycoprotein in transport to the plasma membrane. Journal of Cell Biology, 102(6):2147–??, June 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/102/6/2147.

Presta:1989:MSP

[PMR89] M. Presta, J. A. Maier, and G. Ragnotti. The mitogenic signaling pathway but not the plasminogen activator-inducing pathway of basic fibroblast growth factor is mediated through protein kinase C in fetal bovine aortic endothelial cells. Journal of Cell Biology, 109(4):1877–??, October 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/109/4/1877.

Pon:1989:PIM

[PMV^{+}89] L. Pon, T. Moll, D. Vestweber, B. Marshallsay, and G. Schatz. Protein import into mitochondria: ATP-dependent protein translocation activity in a submitochondrial fraction enriched in membrane contact sites and specific proteins. Journal of Cell Biology, 109(6):2603–??, December 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/109/6/2603.

Philp:1985:CCR

[PN85] N. J. Philp and V. T. Nachmias. Components of the cytoskeleton in the retinal pigmented epithelium of the chick. Journal of Cell Biology, 101(2):358–??, August 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/101/2/358.

Pasdar:1988:KDAa

[PN88a] M. Pasdar and W. J. Nelson. Kinetics of desmosome assembly in Madin–Darby canine kidney epithelial cells: temporal and spatial regulation of desmoplakin organization and stabilization upon cell–cell contact. I. Biochemical analysis. Journal of Cell Biology, 106(3):677–??, March 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/106/3/677.

Pasdar:1988:KDAb

[PN88b] M. Pasdar and W. J. Nelson. Kinetics of desmosome assembly in Madin–Darby canine kidney epithelial cells: temporal and spatial regulation of desmoplakin organization and stabilization upon cell–cell contact. II. Morphological analysis. Journal of Cell Biology, 106(3):687–??, March 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/106/3/687.

Pasdar:1989:RDA

[PN89] M. Pasdar and W. J. Nelson. Regulation of desmosome assembly in epithelial cells: kinetics of synthesis, transport, and stabilization of desmoglein I, a major protein of the membrane core domain. Journal of Cell Biology, 109(1):163–??, July 1989. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/109/1/163.

Picker:1989:MAH

[PNB89] L. J. Picker, M. Nakache, and E. C. Butcher. Monoclonal antibodies to human lymphocyte homing receptors define a novel class of adhesion molecules on diverse cell types. Journal of Cell Biology, 109(2):927–??, August 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/109/2/927.

Powell:1988:ETS

[POCM88] S. K. Powell, L. Orci, C. S. Craik, and H. P. Moore. Efficient targeting to storage granules of human proinsulins with altered propeptide domain. Journal of Cell Biology, 106(6):1843–??, June 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/106/6/
1843.

Prochownik:1989:EAC

[POD89] E. V. Prochownik, K. O’Rourke, and V. M. Dixit. Expression and analysis of COOH-terminal deletions of the human thrombospondin molecule. Journal of Cell Biology, 109(2):843–??, August 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/109/2/

Pollard:1986:RCR

[Pol86] T. D. Pollard. Rate constants for the reactions of ATP- and ADP-actin with the ends of actin filaments. Journal of Cell Biology, 103(6):2747–??, December 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/103/6/2747.

Pollard:1989:BWM

[Pol89] T. D. Pollard. E.b. Wilson Medalist, 1988. Journal of Cell Biology, 108(6):2545–??, June 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/108/6/2545.

Piotrowicz:1988:GII

[PON^{+}88] R. S. Piotrowicz, R. P. Orchekowski, D. J. Nugent, K. Y. Yamada, and T. J. Kunicki. Glycoprotein Ic–IIa functions as an activation-independent fibronectin receptor on human platelets. Journal of Cell Biology, 106(4):1359–??, April 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/106/4/1359.

Palmer:1987:KCP

[POW^{+}87] D. K. Palmer, K. O’Day, M. H. Wener, B. S. Andrews, and R. L. Margolis. A 17-kD centromere protein (CENP–A) copurifies with nucleosome core particles and with histones. Journal of Cell Biology, 104(4):805–??, January 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/104/4/805.

Pauza:1988:HIV

[PP88a] C. D. Pauza and T. M. Price. Human immunodeficiency virus infection of T cells and monocytes proceeds via receptor-mediated endocytosis. Journal of Cell Biology, 107(3):959–??, September 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/107/3/
959.

Pomeroy:1988:NGR

[PP88b] S. L. Pomeroy and D. Purves. Neuron/ glia relationships observed over intervals of several months in living mice. Journal of Cell Biology, 107(3):1167–??, September 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/107/3/1167.

Poole:1989:IMS

[PP89] A. R. Poole and I. Pidoux. Immunoelectron microscopic studies of type X collagen in endochondral ossification. Journal of Cell Biology, 109(5):2547–??, November 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/109/5/2547.

Parton:1989:MAB

[PPB^{+}89] R. G. Parton, K. Prydz, M. Bomsel, K. Simons, and G. Griffiths. Meeting of the apical and basolateral endocytic pathways of the Madin–Darby canine kidney cell in late endosomes. Journal of Cell Biology, 109(6):3259–??, December 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/109/6/3259.

Painter:1985:ISG

[PPG85] R. G. Painter, K. N. Prodouz, and W. Gaarde. Isolation of a subpopulation of glycoprotein IIb–III from platelet membranes that is bound to membrane actin. Journal of Cell Biology, 100 (2):652–??, February 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/100/2/652.

Pratt:1986:SCA

[Pra86] M. M. Pratt. Stable complexes of axoplasmic vesicles and microtubules: protein composition and ATPase activity. Journal of Cell Biology, 103(3):957–??, September 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/103/3/957.

Pizzey:1989:IAM

[PRB^{+}89] J. A. Pizzey, L. H. Rowett, C. H. Barton, G. Dickson, and F. S. Walsh. Intercellular adhesion mediated by human muscle neural cell adhesion molecule: effects of alternative exon use. Journal of Cell Biology, 109(6):3465–??, December 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/109/6/3465.

Price:1987:SCP

[Pri87] M. G. Price. Skelemins: cytoskeletal proteins located at the periphery of M-discs in mammalian striated muscle. Journal of Cell Biology, 104(5):1325–??, May 1987. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/104/5/1325.

Preston:1985:GSD

[PRS^{+}85] S. F. Preston, C. S. Regula, P. R. Sager, C. B. Pearson, L. S. Daniels, P. A. Brown, and R. D. Berlin. Glycosaminoglycan synthesis is depressed during mitosis and elevated during early G1. Journal of Cell Biology, 101(3):1086–??, September 1985. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/101/3/1086.

Postlethwaite:1988:MFF

[PRS^{+}88] A. E. Postlethwaite, R. Raghow, G. P. Stricklin, H. Poppleton, J. M. Seyer, and A. H. Kang. Modulation of fibroblast functions by interleukin 1: increased steady-state accumulation of type I procollagen messenger RNAs and stimulation of other functions but not chemotaxis by human recombinant interleukin 1 alpha and beta. Journal of Cell Biology, 106(2):311–??, February 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/106/2/
311.

Pinol-Roma:1989:NHN

[PRSGD89] S. Piñol-Roma, M. S. Swanson, J. G. Gall, and G. Dreyfuss. A novel heterogeneous nuclear RNP protein with a unique distribution on nascent transcripts. Journal of Cell Biology, 109(6): 2575–??, December 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/6/2575.

Presky:1986:RBS

[PS86] D. H. Presky and A. Schonbrunn. Receptor-bound somatostatin and epidermal growth factor are processed differently in GH4C1 rat pituitary cells. Journal of Cell Biology, 102(3):878–??, March 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/102/3/
878.

Persohn:1987:IML

[PS87] E. Persohn and M. Schachner. Immunoelectron microscopic localization of the neural cell adhesion molecules L1 and N-CAM during postnatal development of the mouse cerebellum. Journal of Cell Biology, 105(1):569–??, July 1987. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/105/1/569.

Paavola:1985:UGL

[PSBN85] L. G. Paavola, J. F. Strauss, C. O. Boyd, and J. E. Nestler. Uptake of gold- and [3H]cholesteryl linoleate-labeled human low density lipoprotein by cultured rat granulosa cells: cellular mechanisms involved in lipoprotein metabolism and their importance to steroidogenesis. Journal of Cell Biology, 100(4):1235–??, April 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/100/4/
1235.

Pepper:1989:JCI

[PSC^{+}89] M. S. Pepper, D. C. Spray, M. Chanson, R. Montesano, L. Orci, and P. Meda. Junctional communication is induced in migrating capillary endothelial cells. Journal of Cell Biology, 109(6): 3027–??, December 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/6/3027.

Pfeiffer:1985:MCC

[PSD^{+}85] J. R. Pfeiffer, J. C. Seagrave, B. H. Davis, G. G. Deanin, and J. M. Oliver. Membrane and cytoskeletal changes associated with IgE-mediated serotonin release from rat basophilic leukemia cells. Journal of Cell Biology, 101(6):2145–??, December 1985. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/101/6/2145.

Pollerberg:1985:SEK

[PSGS85] E. G. Pollerberg, R. Sadoul, C. Goridis, and M. Schachner. Selective expression of the 180-kD component of the neural cell adhesion molecule N-CAM during development. Journal of Cell Biology, 101(5):1921–??, November 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/101/5/1921.

Poltorak:1987:MAG

[PSK^{+}87] M. Poltorak, R. Sadoul, G. Keilhauer, C. Landa, T. Fahrig, and M. Schachner. Myelin-associated glycoprotein, a member of the L2/HNK-1 family of neural cell adhesion molecules, is involved in neuron-oligodendrocyte and oligodendrocyte-oligodendrocyte interaction. Journal of Cell Biology, 105(4):1893–??, October 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/105/4/
1893.

Pieper:1989:TEM

[PSK^{+}89] F. R. Pieper, G. Schaart, P. J. Krimpenfort, J. B. Henderik, H. J. Moshage, A. van de Kemp, F. C. Ramaekers, A. Berns, and H. Bloemendal. Transgenic expression of the muscle-specific intermediate filament protein desmin in nonmuscle cells. Journal of Cell Biology, 108(3):1009–??, March 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/108/3/1009.

Pagano:1989:MTF

[PSM89] R. E. Pagano, M. A. Sepanski, and O. C. Martin. Molecular trapping of a fluorescent ceramide analogue at the Golgi apparatus of fixed cells: interaction with endogenous lipids provides a trans-Golgi marker for both light and electron microscopy. Journal of Cell Biology, 109(5):2067–??, November 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/109/5/2067.

Peltz:1985:MAA

[PSP85] G. Peltz, J. A. Spudich, and P. Parham. Monoclonal antibodies against seven sites on the head and tail of Dictyostelium myosin. Journal of Cell Biology, 100(4):1016–??, April 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/100/4/1016.

Pfaller:1988:IPP

[PSR^{+}88] R. Pfaller, H. F. Steger, J. Rassow, N. Pfanner, and W. Neupert. Import pathways of precursor proteins into mitochondria: multiple receptor sites are followed by a common membrane insertion site. Journal of Cell Biology, 107(6):2483–??, December 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/107/6/
2483.

Pollanen:1987:DLU

[PSS^{+}87] J. Pöllänen, O. Saksela, E. M. Salonen, P. Andreasen, L. Nielsen, K. Danø, and A. Vaheri. Distinct localizations of urokinase-type plasminogen activator and its type 1 inhibitor under cultured human fibroblasts and sarcoma cells. Journal of Cell Biology, 104 (4):1085–??, January 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/104/4/1085.

Pesheva:1989:JJO

[PSS89] P. Pesheva, E. Spiess, and M. Schachner. J1-160 and J1-180 are oligodendrocyte-secreted nonpermissive substrates for cell adhesion. Journal of Cell Biology, 109(4):1765–??, October 1989. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/109/4/1765.

Predescu:1988:BTG

[PSSP88] D. Predescu, M. Simionescu, N. Simionescu, and G. E. Palade. Binding and transcytosis of glycoalbumin by the microvascular endothelium of the murine myocardium: evidence that glycoalbumin behaves as a bifunctional ligand. Journal of Cell Biology, 107(5):1729–??, November 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/107/5/1729.

Paschal:1987:MMA

[PSV87] B. M. Paschal, H. S. Shpetner, and R. B. Vallee. MAP 1C is a microtubule-activated ATPase which translocates microtubules in vitro and has dynein-like properties. Journal of Cell Biology, 105(3):1273–??, September 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/105/3/1273.

Pagliaro:1988:AEB

[PT88] L. Pagliaro and D. L. Taylor. Aldolase exists in both the fluid and solid phases of cytoplasm. Journal of Cell Biology, 107(3): 981–??, September 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/107/3/981.

Poruchynsky:1985:DNT

[PTB^{+}85] M. S. Poruchynsky, C. Tyndall, G. W. Both, F. Sato, A. R. Bellamy, and P. H. Atkinson. Deletions into an NH2-terminal hydrophobic domain result in secretion of rotavirus VP7, a resident endoplasmic reticulum membrane glycoprotein. Journal of Cell Biology, 101(6):2199–??, December 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/101/6/2199.

Pan:1985:EME

[PTW^{+}85] B. T. Pan, K. Teng, C. Wu, M. Adam, and R. M. Johnstone. Electron microscopic evidence for externalization of the transferrin receptor in vesicular form in sheep reticulocytes. Journal of Cell Biology, 101(3):942–??, September 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/101/3/942.

Pumplin:1989:ARC

[Pum89] D. W. Pumplin. Acetylcholine receptor clusters of rat myotubes have at least three domains with distinctive cytoskeletal and membranous components. Journal of Cell Biology, 109 (2):739–??, August 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/2/739.

Prpic:1988:BFR

[PUW^{+}88] V. Prpic, R. J. Uhing, J. E. Weiel, L. Jakoi, G. Gawdi, B. Herman, and D. O. Adams. Biochemical and functional responses stimulated by platelet-activating factor in murine peritoneal macrophages. Journal of Cell Biology, 107(1):363–??, July 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/107/1/363.

Pavco:1985:PGP

[PV85] P. A. Pavco and G. C. Van Tuyle. Purification and general properties of the DNA-binding protein (P16) from rat liver mitochondria. Journal of Cell Biology, 100(1):258–??, January 1985. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/100/1/258.

Pepper:1987:UTP

[PVMO87] M. S. Pepper, J. D. Vassalli, R. Montesano, and L. Orci. Urokinase-type plasminogen activator is induced in migrating capillary endothelial cells. Journal of Cell Biology, 105(6): 2535–??, December 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/6/2535.

Pleshinger:1985:MSI

[PW85] J. Pleshinger and E. Weidner. The microsporidian spore invasion tube. IV. Discharge activation begins with pH-triggered Ca^{2+} influx. Journal of Cell Biology, 100(6):1834–??, June 1985. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/100/6/1834.

Patzak:1986:EER

[PW86] A. Patzak and H. Winkler. Exocytotic exposure and recycling of membrane antigens of chromaffin granules: ultrastructural evaluation after immunolabeling. Journal of Cell Biology, 102 (2):510–??, February 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/102/2/510.

Patton:1987:UPV

[PW87] J. G. Patton and E. D. Wieben. U1 precursors: variant 3’ flanking sequences are transcribed in human cells. Journal of Cell Biology, 104(2):175–??, February 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/104/2/175.

Pfister:1989:MAK

[PWS^{+}89] K. K. Pfister, M. C. Wagner, D. L. Stenoien, S. T. Brady, and G. S. Bloom. Monoclonal antibodies to kinesin heavy and light chains stain vesicle-like structures, but not microtubules, in cultured cells. Journal of Cell Biology, 108(4):1453–??, April 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/108/4/1453.

Painter:1987:RAS

[PZBD87] R. G. Painter, K. Zahler-Bentz, and R. E. Dukes. Regulation of the affinity state of the N-formylated peptide receptor of neutrophils: role of guanine nucleotide-binding proteins and the cytoskeleton. Journal of Cell Biology, 105(6):2959–??, December 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/105/6/
2959.

Quillen:1985:CST

[QCKR85] M. Quillen, C. Castello, A. Krishan, and R. W. Rubin. Cell surface tubulin in leukemic cells: molecular structure, surface binding, turnover, cell cycle expression, and origin. Journal of Cell Biology, 101(6):2345–??, December 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/101/6/2345.

Quinlan:1987:CTM

[QS87] R. A. Quinlan and M. Stewart. Crystalline tubes of myosin subfragment-2 showing the coiled-coil and molecular interaction geometry. Journal of Cell Biology, 105(1):403–??, July 1987. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/105/1/403.

Quaroni:1985:CCD

[Qua85a] A. Quaroni. Crypt cell development in newborn rat small intestine. Journal of Cell Biology, 100(5):1601–??, May 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/100/5/1601.

Quaroni:1985:DFR

[Qua85b] A. Quaroni. Development of fetal rat intestine in organ and monolayer culture. Journal of Cell Biology, 100(5):1611–??, May 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/100/5/
1611.

Ranscht:1988:SCK

[Ran88] B. Ranscht. Sequence of contactin, a 130-kD glycoprotein concentrated in areas of interneuronal contact, defines a new member of the immunoglobulin supergene family in the nervous system. Journal of Cell Biology, 107(4):1561–??, October 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/107/4/1561.

Rapraeger:1989:TGF

[Rap89] A. Rapraeger. Transforming growth factor (type beta) promotes the addition of chondroitin sulfate chains to the cell surface proteoglycan (syndecan) of mouse mammary epithelia. Journal of Cell Biology, 109(5):2509–??, November 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/109/5/2509.

Rodermel:1985:MPP

[RB85] S. R. Rodermel and L. Bogorad. Maize plastid photogenes: mapping and photoregulation of transcript levels during light-induced development. Journal of Cell Biology, 100(2):463–??, February 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/100/2/

Rungger-Brandle:1989:ETC

[RBAF89] E. Rungger-Brändle, T. Achtstätter, and W. W. Franke. An epithelium-type cytoskeleton in a glial cell: astrocytes of amphibian optic nerves contain cytokeratin filaments and are connected by desmosomes. Journal of Cell Biology, 109(2):705–??, August 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/109/2/
705.

Ratner:1985:NCS

[RBG85] N. Ratner, R. P. Bunge, and L. Glaser. A neuronal cell surface heparan sulfate proteoglycan is required for dorsal root ganglion neuron stimulation of Schwann cell proliferation. Journal of Cell Biology, 101(3):744–??, September 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/101/3/744.

Reichmann:1989:NME

[RBGF89] E. Reichmann, R. Ball, B. Groner, and R. R. Friis. New mammary epithelial and fibroblastic cell clones in coculture form structures competent to differentiate functionally. Journal of Cell Biology, 108(3):1127–??, March 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/108/3/1127.

Rattner:1989:KSE

[RBJ89] J. B. Rattner and D. P. Bazett-Jones. Kinetochore structure: electron spectroscopic imaging of the kinetochore. Journal of Cell Biology, 108(4):1209–??, April 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/108/4/1209.

Robinson:1985:RSC

[RBKK85] J. M. Robinson, J. A. Badwey, M. L. Karnovsky, and M. J. Karnovsky. Release of superoxide and change in morphology by neutrophils in response to phorbol esters: antagonism by inhibitors of calcium-binding proteins. Journal of Cell Biology, 101 (3):1052–??, September 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/101/3/1052.

Robinson:1987:CSD

[RBKK87] J. M. Robinson, J. A. Badwey, M. L. Karnovsky, and M. J. Karnovsky. Cell surface dynamics of neutrophils stimulated with phorbol esters or retinoids. Journal of Cell Biology, 105(1):417– ??, July 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
105/1/417.

Ripellino:1988:LEM

[RBMM88] J. A. Ripellino, M. Bailo, R. U. Margolis, and R. K. Margolis. Light and electron microscopic studies on the localization of hyaluronic acid in developing rat cerebellum. Journal of Cell Biology, 106(3):845–??, March 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/106/3/845.

Roos:1985:BSCa

[RC85a] D. S. Roos and P. W. Choppin. Biochemical studies on cell fusion. I. Lipid composition of fusion-resistant cells. Journal of Cell Biology, 101(4):1578–??, October 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/101/4/1578.

Roos:1985:BSCb

[RC85b] D. S. Roos and P. W. Choppin. Biochemical studies on cell fusion. II. Control of fusion response by lipid alteration. Journal of Cell Biology, 101(4):1591–??, October 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/101/4/1591.

Richman:1988:MCG

[RCC^{+}88] R. Richman, L. G. Chicoine, M. P. Collini, R. G. Cook, and C. D. Allis. Micronuclei and the cytoplasm of growing Tetrahymena contain a histone acetylase activity which is highly specific for free histone H4. Journal of Cell Biology, 106(4):1017–??, April 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/106/4/
1017.

Rougon:1986:MAA

[RDB^{+}86] G. Rougon, C. Dubois, N. Buckley, J. L. Magnani, and W. Zollinger. A monoclonal antibody against meningococcus group B polysaccharides distinguishes embryonic from adult N-CAM. Journal of Cell Biology, 103(6):2429–??, December 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/103/6/2429.

Rieder:1986:OMM

[RDJ^{+}86] C. L. Rieder, E. A. Davison, L. C. Jensen, L. Cassimeris, and E. D. Salmon. Oscillatory movements of monooriented chromosomes and their position relative to the spindle pole result from the ejection properties of the aster and half-spindle. Journal of Cell Biology, 103(2):581–??, August 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/103/2/581.

Ruiz:1987:MAB

[RdLB87] F. Ruiz, N. Garreau de Loubresse, and J. Beisson. A mutation affecting basal body duplication and cell shape in Paramecium. Journal of Cell Biology, 104(3):417–??, March 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/104/3/417.

Rieger:1986:NCA

[RDPR^{+}86] F. Rieger, J. K. Daniloff, M. Pincon-Raymond, K. L. Crossin, M. Grumet, and G. M. Edelman. Neuronal cell adhesion molecules and cytotactin are colocalized at the node of Ranvier. Journal of Cell Biology, 103(2):379–??, August 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/103/2/379.

Roossien:1989:ILL

[RdRBR89] F. F. Roossien, D. de Rijk, A. Bikker, and E. Roos. Involvement of LFA-1 in lymphoma invasion and metastasis demonstrated with LFA-1-deficient mutants. Journal of Cell Biology, 108(5):1979–??, May 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/108/5/1979.

Rothblatt:1989:MGR

[RDS^{+}89] J. A. Rothblatt, R. J. Deshaies, S. L. Sanders, G. Daum, and R. Schekman. Multiple genes are required for proper insertion of secretory proteins into the endoplasmic reticulum in yeast. Journal of Cell Biology, 109(6):2641–??, December 1989. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/109/6/2641.

Roth:1986:HTC

[RDSG86] M. G. Roth, C. Doyle, J. Sambrook, and M. J. Gething. Heterologous transmembrane and cytoplasmic domains direct functional chimeric influenza virus hemagglutinins into the endocytic pathway. Journal of Cell Biology, 102(4):1271–??, April 1986. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/102/4/1271.

Ridley:1989:TGF

[RDSL89] A. J. Ridley, J. B. Davis, P. Stroobant, and H. Land. Transforming growth factors-beta 1 and beta 2 are mitogens for rat Schwann cells. Journal of Cell Biology, 109(6):3419–??, December 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/109/6/
3419.

Resh:1985:HSA

[RE85] M. D. Resh and R. L. Erikson. Highly specific antibody to Rous sarcoma virus src gene product recognizes a novel population of pp60v-src and pp60c-src molecules. Journal of Cell Biology, 100 (2):409–??, February 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/100/2/409.

Risau:1986:PHB

[RE86] W. Risau and P. Ekblom. Production of a heparin-binding angiogenesis factor by the embryonic kidney. Journal of Cell Biology, 103(3):1101–??, September 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/103/3/1101.

Ratner:1986:SAL

[REB^{+}86] N. Ratner, A. Elbein, M. B. Bunge, S. Porter, R. P. Bunge, and L. Glaser. Specific asparagine-linked oligosaccharides are not required for certain neuron-neuron and neuron–Schwann cell interactions. Journal of Cell Biology, 103(1):159–??, July 1986. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/103/1/159.

Reeder:1985:MND

[Ree85] R. H. Reeder. Mechanisms of nucleolar dominance in animals and plants. Journal of Cell Biology, 101(5):2013–??, November 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/101/5/
2013.

Role:1987:CNC

[RF87] L. W. Role and G. D. Fischbach. Changes in the number of chick ciliary ganglion neuron processes with time in cell culture. Journal of Cell Biology, 104(2):363–??, February 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/104/2/363.

Rizzolo:1985:BIS

[RFG^{+}85] L. J. Rizzolo, J. Finidori, A. Gonzalez, M. Arpin, I. E. Ivanov, M. Adesnik, and D. D. Sabatini. Biosynthesis and intracellular sorting of growth hormone-viral envelope glycoprotein hybrids. Journal of Cell Biology, 101(4):1351–??, October 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/101/4/1351.

Ratcliffe:1985:PBC

[RFH85] A. Ratcliffe, P. R. Fryer, and T. E. Hardingham. Proteoglycan biosynthesis in chondrocytes: protein A-gold localization of proteoglycan protein core and chondroitin sulfate within Golgi subcompartments. Journal of Cell Biology, 101(6):2355–??, December 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/101/6/
2355.

Roff:1986:CHO

[RFMR86] C. F. Roff, R. Fuchs, I. Mellman, and A. R. Robbins. Chinese hamster ovary cell mutants with temperature-sensitive defects in endocytosis. I. Loss of function on shifting to the nonpermissive temperature. Journal of Cell Biology, 103(6):2283–??, December 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/103/6/
2283.

Roe:1988:EIM

[RFSL88] J. L. Roe, H. A. Farach, W. J. Strittmatter, and W. J. Lennarz. Evidence for involvement of metalloendoproteases in a step in sea urchin gamete fusion. Journal of Cell Biology, 107(2):539–??, August 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/107/2/
539.

Regauer:1985:IFC

[RFV85] S. Regauer, W. W. Franke, and I. Virtanen. Intermediate filament cytoskeleton of amnion epithelium and cultured amnion epithelial cells: expression of epidermal cytokeratins in cells of a simple epithelium. Journal of Cell Biology, 100(4):997–??, April 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/100/4/
997.

Rosa:1985:MTS

[RFZH85] P. Rosa, G. Fumagalli, A. Zanini, and W. B. Huttner. The major tyrosine-sulfated protein of the bovine anterior pituitary is a secretory protein present in gonadotrophs, thyrotrophs, mammotrophs, and corticotrophs. Journal of Cell Biology, 100 (3):928–??, March 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/100/3/928.

Roman:1986:ACD

[RG86a] L. M. Roman and H. Garoff. Alteration of the cytoplasmic domain of the membrane-spanning glycoprotein p62 of Semliki Forest virus does not affect its polar distribution in established lines of Madin–Darby canine kidney cells. Journal of Cell Biology, 103 (6):2607–??, December 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/103/6/2607.

Rotrosen:1986:HTR

[RG86b] D. Rotrosen and J. I. Gallin. Histamine type I receptor occupancy increases endothelial cytosolic calcium, reduces Factin, and promotes albumin diffusion across cultured endothelial monolayers. Journal of Cell Biology, 103(6):2379–??, December 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/103/6/
2379.

Roth:1987:MAR

[RG87] M. B. Roth and J. G. Gall. Monoclonal antibodies that recognize transcription unit proteins on newt lampbrush chromosomes.

Journal of Cell Biology, 105(3):1047–??, September 1987. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/105/3/1047.

Raff:1988:NCM

[RG88] J. W. Raff and D. M. Glover. Nuclear and cytoplasmic mitotic cycles continue in Drosophila embryos in which DNA synthesis is inhibited with aphidicolin. Journal of Cell Biology, 107(6): 2009–??, December 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/107/6/2009.

Rothwell:1987:RCP

[RGBM87] S. W. Rothwell, W. A. Grasser, H. N. Baker, and D. B. Murphy. The relative contributions of polymer annealing and subunit exchange to microtubule dynamics in vitro. Journal of Cell Biology, 105(2):863–??, August 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/2/863.

Rieger:1985:CVN

[RGE85] F. Rieger, M. Grumet, and G. M. Edelman. N-CAM at the vertebrate neuromuscular junction. Journal of Cell Biology, 101(1):285–??, July 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/101/1/285.

Rothwell:1985:DOM

[RGM85] S. W. Rothwell, W. A. Grasser, and D. B. Murphy. Direct observation of microtubule treadmilling by electron microscopy. Journal of Cell Biology, 101(5):1637–??, November 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/101/5/1637.

Rothwell:1986:EEA

[RGM86] S. W. Rothwell, W. A. Grasser, and D. B. Murphy. End-to-end annealing of microtubules in vitro. Journal of Cell Biology, 102 (2):619–??, February 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/102/2/619.

Roth:1987:LED

[RGPRB87] M. G. Roth, D. Gundersen, N. Patil, and E. Rodriguez-Boulan. The large external domain is sufficient for the correct sorting of secreted or chimeric influenza virus hemagglutinins in polarized monkey kidney cells. Journal of Cell Biology, 104(3):769–??, March 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
104/3/769.

Rinnerthaler:1988:CFD

[RGS88] G. Rinnerthaler, B. Geiger, and J. V. Small. Contact formation during fibroblast locomotion: involvement of membrane ruffles and microtubules. Journal of Cell Biology, 106(3):747–??, March 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/106/3/
747.

Rassow:1989:TAR

[RGW^{+}89] J. Rassow, B. Guiard, U. Wienhues, V. Herzog, F. U. Hartl, and W. Neupert. Translocation arrest by reversible folding of a precursor protein imported into mitochondria. A means to quantitate translocation contact sites. Journal of Cell Biology, 109 (4):1421–??, October 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/4/1421.

Riches:1986:BLS

[RH86a] D. W. Riches and P. M. Henson. Bacterial lipopolysaccharide suppresses the production of catalytically active lysosomal acid hydrolases in human macrophages. Journal of Cell Biology, 102 (5):1606–??, May 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/102/5/1606.

Rozdzial:1986:RMM

[RH86b] M. M. Rozdzial and L. T. Haimo. Reactivated melanophore motility: differential regulation and nucleotide requirements of bidirectional pigment granule transport. Journal of Cell Biology, 103(6):2755–??, December 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/103/6/2755.

Rhodes:1987:NSP

[RH87] C. J. Rhodes and P. A. Halban. Newly synthesized proinsulin/ insulin and stored insulin are released from pancreatic B cells predominantly via a regulated, rather than a constitutive, pathway. Journal of Cell Biology, 105(1):145–??, July 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/105/1/145.

Ralston:1989:ISD

[RH89] E. Ralston and Z. W. Hall. Intracellular and surface distribution of a membrane protein (CD8) derived from a single nucleus in multinucleated myotubes. Journal of Cell Biology, 109(5): 2345–??, November 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/5/2345.

Rosenfeld:1985:IPC

[RHD^{+}85] G. C. Rosenfeld, D. C. Hou, J. Dingus, I. Meza, and J. Bryan. Isolation and partial characterization of human platelet vinculin. Journal of Cell Biology, 100(3):669–??, March 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/100/3/669.

Rosa:1985:SIT

[RHL^{+}85] P. Rosa, A. Hille, R. W. Lee, A. Zanini, P. De Camilli, and W. B. Huttner. Secretogranins I and II: two tyrosine-sulfated secretory proteins common to a variety of cells secreting peptides by the regulated pathway. Journal of Cell Biology, 101(5): 1999–??, November 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/101/5/1999.

Rindler:1985:PDV

[RIPS85] M. J. Rindler, I. E. Ivanov, H. Plesken, and D. D. Sabatini. Polarized delivery of viral glycoproteins to the apical and basolateral plasma membranes of Madin–Darby canine kidney cells infected with temperature-sensitive viruses. Journal of Cell Biology, 100 (1):136–??, January 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/100/1/136.

Ris:1985:CFS

[Ris85] H. Ris. The cytoplasmic filament system in critical point-dried whole mounts and plastic-embedded sections. Journal of Cell Biology, 100(5):1474–??, May 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/100/5/1474.

Rindler:1987:MAD

[RIS87] M. J. Rindler, I. E. Ivanov, and D. D. Sabatini. Microtubuleacting drugs lead to the nonpolarized delivery of the influenza hemagglutinin to the cell surface of polarized Madin–Darby canine kidney cells. Journal of Cell Biology, 104(2):231–??, February 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/104/2/
231.

Rapraeger:1986:CSP

[RJB86] A. Rapraeger, M. Jalkanen, and M. Bernfield. Cell surface proteoglycan associates with the cytoskeleton at the basolateral cell surface of mouse mammary epithelial cells. Journal of Cell Biology, 103(6):2683–??, December 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/103/6/2683.

Rheinwald:1987:MSG

[RJH^{+}87] J. G. Rheinwald, J. L. Jorgensen, W. C. Hahn, A. J. Terpstra, T. M. O’Connell, and K. K. Plummer. Mesosecrin: a secreted glycoprotein produced in abundance by human mesothelial, endothelial, and kidney epithelial cells in culture. Journal of Cell Biology, 104(2):263–??, February 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/104/2/263.

Rogelj:1989:BFG

[RKA^{+}89] S. Rogelj, M. Klagsbrun, R. Atzmon, M. Kurokawa, A. Haimovitz, Z. Fuks, and I. Vlodavsky. Basic fibroblast growth factor is an extracellular matrix component required for supporting the proliferation of vascular endothelial cells and the differentiation of PC12 cells. Journal of Cell Biology, 109(2):823–??, August 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/109/2/
823.

Ruohola:1988:RPT

[RKFN88] H. Ruohola, A. K. Kabcenell, and S. Ferro-Novick. Reconstitution of protein transport from the endoplasmic reticulum to the Golgi complex in yeast: the acceptor Golgi compartment is defective in the sec23 mutant. Journal of Cell Biology, 107 (4):1465–??, October 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/107/4/1465.

Rentrop:1986:DLD

[RKWS86] M. Rentrop, B. Knapp, H. Winter, and J. Schweizer. Differential localization of distinct keratin mRNA-species in mouse tongue epithelium by in situ hybridization with specific cDNA probes. Journal of Cell Biology, 103(6):2583–??, December 1986. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/103/6/2583.

Rossi:1989:ILY

[RL89] J. M. Rossi and S. Lindquist. The intracellular location of yeast heat-shock protein 26 varies with metabolism. Journal of Cell Biology, 108(2):425–??, February 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/108/2/425.

Roth:1985:IDE

[RLB85] J. Roth, M. J. Lentze, and E. G. Berger. Immunocytochemical demonstration of ecto-galactosyltransferase in absorptive intestinal cells. Journal of Cell Biology, 100(1):118–??, January 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/100/1/
118.

Roman:1989:FRO

[RLB^{+}89] J. Roman, R. M. LaChance, T. J. Broekelmann, C. J. Kennedy, E. A. Wayner, W. G. Carter, and J. A. McDonald. The fibronectin receptor is organized by extracellular matrix fibronectin: implications for oncogenic transformation and for cell recognition of fibronectin matrices. Journal of Cell Biology, 108 (6):2529–??, June 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/108/6/2529.

Richter-Landsberg:1986:RCN

[RLJ86] C. Richter-Landsberg and B. Jastorff. The role of cAMP in nerve growth factor-promoted neurite outgrowth in PC12 cells. Journal of Cell Biology, 102(3):821–??, March 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/102/3/821.

Roth:1985:LSS

[RLP85] M. Roth, M. Lin, and D. M. Prescott. Large scale synchronous mating and the study of macronuclear development in Euplotes crassus. Journal of Cell Biology, 101(1):79–??, July 1985. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/101/1/79.

Rogers:1987:SIP

[RLP^{+}87] S. L. Rogers, P. C. Letourneau, B. A. Peterson, L. T. Furcht, and J. B. McCarthy. Selective interaction of peripheral and central nervous system cells with two distinct cell-binding domains of fibronectin. Journal of Cell Biology, 105(3):1435–??, September 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/105/3/
1435.

Rifkin:1989:RDC

[RM89a] D. B. Rifkin and D. Moscatelli. Recent developments in the cell biology of basic fibroblast growth factor. Journal of Cell Biology, 109(1):1–??, July 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/1/1.

Rivas:1989:SSR

[RM89b] R. J. Rivas and H. P. Moore. Spatial segregation of the regulated and constitutive secretory pathways. Journal of Cell Biology, 109(1):51–??, July 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/1/51.

Reist:1987:ALM

[RMM87] N. E. Reist, C. Magill, and U. J. McMahan. Agrin-like molecules at synaptic sites in normal, denervated, and damaged skeletal muscles. Journal of Cell Biology, 105(6):2457–??, December 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/105/6/
2457.

Ripellino:1989:IML

[RMM89] J. A. Ripellino, R. U. Margolis, and R. K. Margolis. Immunoelectron microscopic localization of hyaluronic acid-binding region and link protein epitopes in brain. Journal of Cell Biology, 108(5):1899–??, May 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/108/5/1899.

Rauvala:1988:ANP

[RMP^{+}88] H. Rauvala, J. Merenmies, R. Pihlaskari, M. Korkolainen, M. L. Huhtala, and P. Panula. The adhesive and neurite-promoting molecule p30: analysis of the amino-terminal sequence and production of antipeptide antibodies that detect p30 at the surface of neuroblastoma cells and of brain neurons. Journal of Cell Biology, 107(6):2293–??, December 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/107/6/2293.

Runyan:1986:ENE

[RMS86] R. B. Runyan, G. D. Maxwell, and B. D. Shur. Evidence for a novel enzymatic mechanism of neural crest cell migration on extracellular glycoconjugate matrices. Journal of Cell Biology, 102 (2):432–??, February 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/102/2/432.

Ratan:1988:LLR

[RMS88] R. R. Ratan, F. R. Maxfield, and M. L. Shelanski. Long-lasting and rapid calcium changes during mitosis. Journal of Cell Biology, 107(3):993–??, September 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/107/3/993.

Rieger:1988:DRR

[RNPR^{+}88] F. Rieger, M. Nicolet, M. Pinçon-Raymond, M. Murawsky, G. Levi, and G. M. Edelman. Distribution and role in regeneration of N-CAM in the basal laminae of muscle and Schwann cells. Journal of Cell Biology, 107(2):707–??, August 1988. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/107/2/707.

Rothwell:1989:ICG

[RNW89] S. W. Rothwell, J. Nath, and D. G. Wright. Interactions of cytoplasmic granules with microtubules in human neutrophils. Journal of Cell Biology, 108(6):2313–??, June 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/108/6/2313.

Rifkin:1985:CMD

[RO85] J. M. Rifkin and G. E. Olson. Characterization of maturationdependent extrinsic proteins of the rat sperm surface. Journal of Cell Biology, 100(5):1582–??, May 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/100/5/1582.

Robinson:1985:RCE

[Rob85] K. R. Robinson. The responses of cells to electrical fields: a review. Journal of Cell Biology, 101(6):2023–??, December 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/101/6/2023.

Robinson:1987:KCV

[Rob87] M. S. Robinson. 100-kD coated vesicle proteins: molecular heterogeneity and intracellular distribution studied with monoclonal antibodies. Journal of Cell Biology, 104(4):887–??, January 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/104/4/887.

Robinson:1989:CCE

[Rob89] M. S. Robinson. Cloning of cDNAs encoding two related 100-kD coated vesicle proteins (alpha-adaptins). Journal of Cell Biology, 108(3):833–??, March 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/108/3/833.

Robinson:1986:ULA

[ROCK86] J. M. Robinson, T. Okada, J. J. Castellot, and M. J. Karnovsky. Unusual lysosomes in aortic smooth muscle cells: presence in living and rapidly frozen cells. Journal of Cell Biology, 102 (5):1615–??, May 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/102/5/1615.

Rogalski:1987:PMI

[Rog87a] A. A. Rogalski. A plasma membrane integral sialoglycoprotein (Sgp 130) molecularly distinguishes nonjunctional dense plaque sites of microfilament attachment. Journal of Cell Biology, 105 (2):819–??, August 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/2/819.

Rogers:1987:CGN

[Rog87b] J. H. Rogers. Calretinin: a gene for a novel calcium-binding protein expressed principally in neurons. Journal of Cell Biology, 105 (3):1343–??, September 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/3/1343.

Ross:1989:CAD

[Ros89] P. M. Ross. Cellular and adenovirus dl312 DNA metabolism in cycling or mitotic human cultures exposed to supralethal gamma radiation. Journal of Cell Biology, 109(5):1993–??, November 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/109/5/
1993.

Robinson:1986:ILC

[RP86] M. S. Robinson and B. M. Pearse. Immunofluorescent localization of 100K coated vesicle proteins. Journal of Cell Biology, 102(1):48–??, January 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/102/1/48.

Rimm:1989:PCA

[RP89] D. L. Rimm and T. D. Pollard. Purification and characterization of an Acanthamoeba nuclear actin-binding protein. Journal of Cell Biology, 109(2):585–??, August 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/109/2/585.

Roher:1988:ICC

[RPCB88] A. E. Roher, K. C. Palmer, V. Chau, and M. J. Ball. Isolation and chemical characterization of Alzheimer’s disease paired helical filament cytoskeletons: differentiation from amyloid plaque core protein. Journal of Cell Biology, 107(6):2703–??, December 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/107/6/
2703.

Roberts:1989:SBL

[RPRS89] C. J. Roberts, G. Pohlig, J. H. Rothman, and T. H. Stevens. Structure, biosynthesis, and localization of dipeptidyl aminopeptidase B, an integral membrane glycoprotein of the yeast vacuole. Journal of Cell Biology, 108(4):1363–??, April 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/108/4/1363.

Roberts:1986:MCP

[RPW86] T. M. Roberts, F. M. Pavalko, and S. Ward. Membrane and cytoplasmic proteins are transported in the same organelle complex during nematode spermatogenesis. Journal of Cell Biology, 102(5):1787–??, May 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/102/5/1787.

Rivera:1989:LTP

[RPW^{+}89] R. T. Rivera, S. G. Pasion, D. T. Wong, Y. B. Fei, and D. K. Biswas. Loss of tumorigenic potential by human lung tumor cells in the presence of antisense RNA specific to the ectopically synthesized alpha subunit of human chorionic gonadotropin. Journal of Cell Biology, 108(6):2423–??, June 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/108/6/2423.

Rothnagel:1986:TIF

[RR86] J. A. Rothnagel and G. E. Rogers. Trichohyalin, an intermediate filament-associated protein of the hair follicle. Journal of Cell Biology, 102(4):1419–??, April 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/102/4/1419.

Roos:1987:ILF

[RR87] E. Roos and F. F. Roossien. Involvement of leukocyte functionassociated antigen-1 (LFA-1) in the invasion of hepatocyte cultures by lymphoma and T-cell hybridoma cells. Journal of Cell Biology, 105(1):553–??, July 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/105/1/553.

Rasmussen:1988:ASH

[RR88] S. Rasmussen and A. Rapraeger. Altered structure of the hybrid cell surface proteoglycan of mammary epithelial cells in response to transforming growth factor-beta. Journal of Cell Biology, 107 (5):1959–??, November 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/107/5/1959.

Reisine:1986:LDC

[RRB86] T. Reisine, G. Rougon, and J. Barbet. Liposome delivery of cyclic AMP-dependent protein kinase inhibitor into intact cells: specific blockade of cyclic AMP-mediated adrenocorticotropin release from mouse anterior pituitary tumor cells. Journal of Cell Biology, 102(5):1630–??, May 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/102/5/1630.

Role:1987:DAR

[RRF87] L. W. Role, D. G. Roufa, and G. D. Fischbach. The distribution of acetylcholine receptor clusters and sites of transmitter release along chick ciliary ganglion neurite-myotube contacts in culture. Journal of Cell Biology, 104(2):371–??, February 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/104/2/371.

Ross:1988:IPC

[RRP88] A. Ross, M. Rapuano, and J. Prives. Induction of phosphorylation and cell surface redistribution of acetylcholine receptors by phorbol ester and carbamylcholine in cultured chick muscle cells. Journal of Cell Biology, 107(3):1139–??, September 1988. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/107/3/1139.

Rogalski:1985:IGA

[RS85] A. A. Rogalski and S. J. Singer. An integral glycoprotein associated with the membrane attachment sites of actin microfilaments. Journal of Cell Biology, 101(3):785–??, September 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/101/3/785.

Ruzicka:1988:SAA

[RS88] D. L. Ruzicka and R. J. Schwartz. Sequential activation of alphaactin genes during avian cardiogenesis: vascular smooth muscle alpha-actin gene transcripts mark the onset of cardiomyocyte differentiation. Journal of Cell Biology, 107(6):2575–??, December 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/107/6/

Roos:1985:LCS

[RSF^{+}85] E. Roos, H. Spiele, C. A. Feltkamp, H. Huisman, F. A. Wiegant, J. Traas, and D. A. Mesland. Localization of cell surface glycoproteins in membrane domains associated with the underlying filament network. Journal of Cell Biology, 101(5):1817–??, November 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/101/5/
1817.

Rodman:1986:MCC

[RSF86] J. S. Rodman, L. Seidman, and M. G. Farquhar. The membrane composition of coated pits, microvilli, endosomes, and lysosomes is distinctive in the rat kidney proximal tubule cell. Journal of Cell Biology, 102(1):77–??, January 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/102/1/77.

Roberts:1987:PTM

[RSG87] D. D. Roberts, J. A. Sherwood, and V. Ginsburg. Platelet thrombospondin mediates attachment and spreading of human melanoma cells. Journal of Cell Biology, 104(1):131–??, January 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/104/1/
131.

Ruegg:1989:HAS

[RSL^{+}89] M. A. Ruegg, E. T. Stoeckli, R. B. Lanz, P. Streit, and P. Sonderegger. A homologue of the axonally secreted protein axonin-1 is an integral membrane protein of nerve fiber tracts involved in neurite fasciculation. Journal of Cell Biology, 109(5): 2363–??, November 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/5/2363.

Rimm:1989:LHT

[RSP89a] D. L. Rimm, J. H. Sinard, and T. D. Pollard. Location of the head-tail junction of myosin. Journal of Cell Biology, 108 (5):1783–??, May 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/108/5/1783.

Roditi:1989:PGE

[RSP^{+}89b] I. Roditi, H. Schwarz, T. W. Pearson, R. P. Beecroft, M. K. Liu, J. P. Richardson, H. J. Bühring, J. Pleiss, R. Bülow, and R. O. Williams. Procyclin gene expression and loss of the variant surface glycoprotein during differentiation of Trypanosoma brucei. Journal of Cell Biology, 108(2):737–??, February 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/108/2/737.

Roth:1988:CPS

[RSR^{+}88] S. Y. Roth, I. G. Schulman, R. Richman, R. G. Cook, and C. D. Allis. Characterization of phosphorylation sites in histone H1 in the amitotic macronucleus of Tetrahymena during different physiological states. Journal of Cell Biology, 107(6):2473–??, December 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/107/6/
2473.

Rosoff:1986:CIC

[RT86] P. M. Rosoff and G. Terres. Cyclosporine A inhibits Ca^{2+}- dependent stimulation of the Na+/H+ antiport in human T cells. Journal of Cell Biology, 103(2):457–??, August 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/103/2/457.

Randriamampita:1987:ICM

[RT87] C. Randriamampita and A. Trautmann. Ionic channels in murine macrophages. Journal of Cell Biology, 105(2):761–??, August 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/105/2/
761.

Rindler:1988:SSS

[RT88] M. J. Rindler and M. G. Traber. A specific sorting signal is not required for the polarized secretion of newly synthesized proteins from cultured intestinal epithelial cells. Journal of Cell Biology, 107(2):471–??, August 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/107/2/471.

Resendez:1986:CIA

[RTK^{+}86] E. Resendez, J. Ting, K. S. Kim, S. K. Wooden, and A. S. Lee. Calcium ionophore A23187 as a regulator of gene expression in mammalian cells. Journal of Cell Biology, 103(6): 2145–??, December 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/103/6/2145.

Ruddle:1988:BWM

[Rud88] F. Ruddle. E. B. Wilson medalist, 1987: Marilyn Gist Farquhar. Journal of Cell Biology, 106(4):1403–??, April 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/106/4/1403.

Runyan:1988:FDL

[RVS88] R. B. Runyan, J. Versalovic, and B. D. Shur. Functionally distinct laminin receptors mediate cell adhesion and spreading: the requirement for surface galactosyltransferase in cell spreading. Journal of Cell Biology, 107(5):1863–??, November 1988. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/107/5/1863.

Richardson:1988:IEL

[RWE^{+}88] J. M. Richardson, N. A. Woychik, D. L. Ebert, R. L. Dimond, and J. A. Cardelli. Inhibition of early but not late proteolytic processing events leads to the missorting and oversecretion of precursor forms of lysosomal enzymes in Dictyostelium discoideum. Journal of Cell Biology, 107(6):2097–??, December 1988. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/107/6/2097.

Rathjen:1987:MGI

[RWF^{+}87] F. G. Rathjen, J. M. Wolff, R. Frank, F. Bonhoeffer, and U. Rutishauser. Membrane glycoproteins involved in neurite fasciculation. Journal of Cell Biology, 104(2):343–??, February 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/104/2/
343.

Rosa:1989:AAS

[RWP^{+}89] P. Rosa, U. Weiss, R. Pepperkok, W. Ansorge, C. Niehrs, E. H. Stelzer, and W. B. Huttner. An antibody against secretogranin I (chromogranin B) is packaged into secretory granules. Journal of Cell Biology, 109(1):17–??, July 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/109/1/17.

Rutishauser:1985:SAN

[RWS^{+}85] U. Rutishauser, M. Watanabe, J. Silver, F. A. Troy, and E. R. Vimr. Specific alteration of NCAM-mediated cell adhesion by an endoneuraminidase. Journal of Cell Biology, 101(5): 1842–??, November 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/101/5/1842.

Robey:1987:OSR

[RYF^{+}87] P. G. Robey, M. F. Young, K. C. Flanders, N. S. Roche, P. Kondaiah, A. H. Reddi, J. D. Termine, M. B. Sporn, and A. B. Roberts. Osteoblasts synthesize and respond to transforming growth factor-type beta (TGF-beta) in vitro. Journal of Cell Biology, 105(1):457–??, July 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/105/1/457.

Robey:1989:TOD

[RYFM89] P. G. Robey, M. F. Young, L. W. Fisher, and T. D. McClain. Thrombospondin is an osteoblast-derived component of mineralized extracellular matrix. Journal of Cell Biology, 108(2): 719–??, February 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/108/2/719.

Rothman:1989:ALL

[RYR^{+}89] J. H. Rothman, C. T. Yamashiro, C. K. Raymond, P. M. Kane, and T. H. Stevens. Acidification of the lysosome-like vacuole and the vacuolar H+-ATPase are deficient in two yeast mutants that fail to sort vacuolar proteins. Journal of Cell Biology, 109(1):93–??, July 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/1/93.

Riederer:1986:BSB

[RZG86] B. M. Riederer, I. S. Zagon, and S. R. Goodman. Brain spectrin(240/235) and brain spectrin(240/235E): two distinct spectrin subtypes with different locations within mammalian neural cells. Journal of Cell Biology, 102(6):2088–??, June 1986. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/102/6/2088.

Sobel:1985:CDS

[SA85] J. S. Sobel and M. A. Alliegro. Changes in the distribution of a spectrin-like protein during development of the preimplantation mouse embryo. Journal of Cell Biology, 100(1):333–??, January 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/100/1/
333.

Shyy:1989:CCK

[SAA89] T. T. Shyy, B. B. Asch, and H. L. Asch. Concurrent collapse of keratin filaments, aggregation of organelles, and inhibition of protein synthesis during the heat shock response in mammary epithelial cells. Journal of Cell Biology, 108(3):997–??, March 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/108/3/
997.

Singh:1988:MFG

[SAB88] J. P. Singh, L. D. Adams, and P. D. Bonin. Mode of fibroblast growth enhancement by human interleukin-1. Journal of Cell Biology, 106(3):813–??, March 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/106/3/813.

Shirinsky:1989:MCC

[SAB^{+}89] V. P. Shirinsky, A. S. Antonov, K. G. Birukov, A. V. Sobolevsky, Y. A. Romanov, N. V. Kabaeva, G. N. Antonova, and V. N. Smirnov. Mechano-chemical control of human endothelium orientation and size. Journal of Cell Biology, 109(1):331–??, July 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/109/1/
331.

Schulze:1987:PMM

[SABK87] E. Schulze, D. J. Asai, J. C. Bulinski, and M. Kirschner. Posttranslational modification and microtubule stability. Journal of Cell Biology, 105(5):2167–??, November 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/105/5/2167.

Simon-Assmann:1989:ODB

[SABVK89] P. Simon-Assmann, F. Bouziges, M. Vigny, and M. Kedinger. Origin and deposition of basement membrane heparan sulfate proteoglycan in the developing intestine. Journal of Cell Biology, 109(4):1837–??, October 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/109/4/1837.

Stevenson:1988:TJS

[SAGM88] B. R. Stevenson, J. M. Anderson, D. A. Goodenough, and M. S. Mooseker. Tight junction structure and ZO-1 content are identical in two strains of Madin–Darby canine kidney cells which differ in transepithelial resistance. Journal of Cell Biology, 107 (6):2401–??, December 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/107/6/2401.

Sale:1986:AAC

[Sal86] W. S. Sale. The axonemal axis and Ca^{2+}-induced asymmetry of active microtubule sliding in sea urchin sperm tails. Journal of Cell Biology, 102(6):2042–??, June 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/102/6/2042.

Stick:1988:FCN

[SALN88] R. Stick, B. Angres, C. F. Lehner, and E. A. Nigg. The fates of chicken nuclear lamin proteins during mitosis: evidence for a reversible redistribution of lamin B2 between inner nuclear membrane and elements of the endoplasmic reticulum. Journal of Cell Biology, 107(2):397–??, August 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/107/2/397.

Schroeder:1985:SAC

[SB85a] T. E. Schroeder and D. E. Battaglia. ”spiral asters” and cytoplasmic rotation in sea urchin eggs: induction in Strongylocentrotus purpuratus eggs by elevated temperature. Journal of Cell Biology, 100(4):1056–??, April 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/100/4/1056.

Shear:1985:VSD

[SB85b] C. R. Shear and R. J. Bloch. Vinculin in subsarcolemmal densities in chicken skeletal muscle: localization and relationship to intracellular and extracellular structures. Journal of Cell Biology, 101(1):240–??, July 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/101/1/240.

Shen:1985:ISA

[SB85c] S. S. Shen and L. J. Burgart. Intracellular sodium activity in the sea urchin egg during fertilization. Journal of Cell Biology, 101(2):420–??, August 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/101/2/420.

Siegel:1985:PPC

[SB85d] D. L. Siegel and D. Branton. Partial purification and characterization of an actin-bundling protein, band 4.9, from human erythrocytes. Journal of Cell Biology, 100(3):775–??, March 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/100/3/
775.

Soltys:1985:PTV

[SB85e] B. J. Soltys and G. G. Borisy. Polymerization of tubulin in vivo: direct evidence for assembly onto microtubule ends and from centrosomes. Journal of Cell Biology, 100(5):1682–??, May 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/100/5/
1682.

Stratford:1985:IAB

[SB85f] C. A. Stratford and S. S. Brown. Isolation of an actin-binding protein from membranes of Dictyostelium discoideum. Journal of Cell Biology, 100(3):727–??, March 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/100/3/727.

Sefton:1987:CME

[SB87] B. M. Sefton and J. E. Buss. The covalent modification of eukaryotic proteins with lipid. Journal of Cell Biology, 104(6):1449–??, June 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/104/6/
1449.

Saunders:1988:CSP

[SB88] S. Saunders and M. Bernfield. Cell surface proteoglycan binds mouse mammary epithelial cells to fibronectin and behaves as a receptor for interstitial matrix. Journal of Cell Biology, 106 (2):423–??, February 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/106/2/423.

Sztul:1987:LKA

[SBC^{+}87] E. S. Sztul, D. Biemesderfer, M. J. Caplan, M. Kashgarian, and J. L. Boyer. Localization of Na+,K+-ATPase alpha-subunit to the sinusoidal and lateral but not canalicular membranes of rat hepatocytes. Journal of Cell Biology, 104(5):1239–??, May 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/104/5/1239.

Schreiner:1989:ICC

[SBD^{+}89] C. L. Schreiner, J. S. Bauer, Y. N. Danilov, S. Hussein, M. M. Sczekan, and R. L. Juliano. Isolation and characterization of Chinese hamster ovary cell variants deficient in the expression of fibronectin receptor. Journal of Cell Biology, 109(6):3157–??, December 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
109/6/3157.

Standart:1985:SSR

[SBG^{+}85] N. M. Standart, S. J. Bray, E. L. George, T. Hunt, and J. V. Ruderman. The small subunit of ribonucleotide reductase is encoded by one of the most abundant translationally regulated maternal RNAs in clam and sea urchin eggs. Journal of Cell Biology, 100(6):1968–??, June 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/100/6/1968.

Suh:1989:SRM

[SBG89] K. Suh, J. E. Bergmann, and C. A. Gabel. Selective retention of monoglucosylated high mannose oligosaccharides by a class of mutant vesicular stomatitis virus G proteins. Journal of Cell Biology, 108(3):811–??, March 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/108/3/811.

Steitz:1988:RLC

[SBH^{+}88] J. A. Steitz, C. Berg, J. P. Hendrick, H. La Branche-Chabot, A. Metspalu, J. Rinke, and T. Yario. A 5S rRNA/L5 complex is a precursor to ribosome assembly in mammalian cells. Journal of Cell Biology, 106(3):545–??, March 1988. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/106/3/545.

Stern:1986:PEC

[SBHN86] D. Stern, J. Brett, K. Harris, and P. Nawroth. Participation of endothelial cells in the protein C-protein S anticoagulant pathway: the synthesis and release of protein S. Journal of Cell Biology, 102(5):1971–??, May 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/102/5/1971.

Schroer:1985:FAT

[SBK85] T. A. Schroer, S. T. Brady, and R. B. Kelly. Fast axonal transport of foreign synaptic vesicles in squid axoplasm. Journal of Cell Biology, 101(2):568–??, August 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/101/2/568.

Shieh:1987:QAS

[SBS87a] B. H. Shieh, M. Ballivet, and J. Schmidt. Quantitation of an alpha subunit splicing intermediate: evidence for transcriptional activation in the control of acetylcholine receptor expression in denervated chick skeletal muscle. Journal of Cell Biology, 104 (5):1337–??, May 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/104/5/1337.

Swanson:1987:TLA

[SBS87b] J. Swanson, E. Burke, and S. C. Silverstein. Tubular lysosomes accompany stimulated pinocytosis in macrophages. Journal of Cell Biology, 104(5):1217–??, May 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/104/5/1217.

Salisbury:1988:CBC

[SBS88] J. L. Salisbury, A. T. Baron, and M. A. Sanders. The centrinbased cytoskeleton of Chlamydomonas reinhardtii: distribution in interphase and mitotic cells. Journal of Cell Biology, 107 (2):635–??, August 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/107/2/635.

Safranyos:1985:RDF

[SC85a] R. G. Safranyos and S. Caveney. Rates of diffusion of fluorescent molecules via cell-to-cell membrane channels in a developing tissue. Journal of Cell Biology, 100(3):736–??, March 1985. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/100/3/736.

Samuelson:1985:CGS

[SC85b] J. C. Samuelson and J. P. Caulfield. The cercarial glycocalyx of Schistosoma mansoni. Journal of Cell Biology, 100(5):1423–??, May 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/100/5/
1423.

Schibler:1986:TDM

[SC86] M. J. Schibler and F. Cabral. Taxol-dependent mutants of Chinese hamster ovary cells with alterations in alpha- and betatubulin. Journal of Cell Biology, 102(4):1522–??, April 1986. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/102/4/1522.

Siliciano:1987:PSM

[SC87] J. D. Siliciano and S. W. Craig. Properties of smooth muscle meta-vinculin. Journal of Cell Biology, 104(3):473–??, March 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/104/3/
473.

Sherry:1988:CTN

[SC88] B. Sherry and A. Cerami. Cachectin/ tumor necrosis factor exerts endocrine, paracrine, and autocrine control of inflammatory responses. Journal of Cell Biology, 107(4):1269–??, October 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/107/4/
1269.

Saide:1989:CCZ

[SCBHS^{+}89] J. D. Saide, S. Chin-Bow, J. Hogan-Sheldon, L. Busquets-Turner, J. O. Vigoreaux, K. Valgeirsdottir, and M. L. Pardue. Characterization of components of Z-bands in the fibrillar flight muscle of Drosophila melanogaster. Journal of Cell Biology, 109(5): 2157–??, November 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/5/2157.

Schachat:1985:PTS

[SCBR85] F. H. Schachat, A. C. Canine, M. M. Briggs, and M. C. Reedy. The presence of two skeletal muscle alpha-actinins correlates with troponin-tropomyosin expression and Z-line width. Journal of Cell Biology, 101(3):1001–??, September 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/101/3/1001.

Siu:1987:CSG

[SCC87] C. H. Siu, A. Cho, and A. H. Choi. The contact site A glycoprotein mediates cell–cell adhesion by homophilic binding in Dictyostelium discoideum. Journal of Cell Biology, 105(6): 2523–??, December 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/6/2523.

Sutherland:1988:EFC

[SCD88] A. E. Sutherland, P. G. Calarco, and C. H. Damsky. Expression and function of cell surface extracellular matrix receptors in mouse blastocyst attachment and outgrowth. Journal of Cell Biology, 106(4):1331–??, April 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/106/4/1331.

Sorger:1986:CCE

[SCFP86] P. K. Sorger, R. A. Crowther, J. T. Finch, and B. M. Pearse. Clathrin cubes: an extreme variant of the normal cage. Journal of Cell Biology, 103(4):1213–??, October 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/103/4/1213.

Sprague:1985:RLH

[SCGS85] S. G. Sprague, E. L. Camm, B. R. Green, and L. A. Staehelin. Reconstitution of light-harvesting complexes and photosystem II cores into galactolipid and phospholipid liposomes. Journal of Cell Biology, 100(2):552–??, February 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/100/2/552.

Schantz:1985:CFC

[Sch85] A. R. Schantz. Cytosolic free calcium-ion concentration in cleaving embryonic cells of Oryzias latipes measured with calciumselective microelectrodes. Journal of Cell Biology, 100(3):947–??, March 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
100/3/947.

Schlesinger:1986:HSP

[Sch86a] M. J. Schlesinger. Heat shock proteins: the search for functions. Journal of Cell Biology, 103(2):321–??, August 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/103/2/321.

Schlessinger:1986:ARE

[Sch86b] J. Schlessinger. Allosteric regulation of the epidermal growth factor receptor kinase. Journal of Cell Biology, 103(6):2067–??, December 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
103/6/2067.

Sadler:1989:YGI

[SCK^{+}89] I. Sadler, A. Chiang, T. Kurihara, J. Rothblatt, J. Way, and P. Silver. A yeast gene important for protein assembly into the endoplasmic reticulum and the nucleus has homology to DnaJ, an Escherichia coli heat shock protein. Journal of Cell Biology, 109 (6):2665–??, December 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/6/2665.

Salpeter:1986:DRA

[SCLG86] M. M. Salpeter, D. L. Cooper, and T. Levitt-Gilmour. Degradation rates of acetylcholine receptors can be modified in the postjunctional plasma membrane of the vertebrate neuromuscular junction. Journal of Cell Biology, 103(4):1399–??, October 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/103/4/
1399.

Shibayama:1987:AIB

[SCM87] T. Shibayama, J. M. Carboni, and M. S. Mooseker. Assembly of the intestinal brush border: appearance and redistribution of microvillar core proteins in developing chick enterocytes. Journal of Cell Biology, 105(1):335–??, July 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/105/1/335.

Schwerzmann:1986:MAI

[SCOE^{+}86] K. Schwerzmann, L. M. Cruz-Orive, R. Eggman, A. Sänger, and E. R. Weibel. Molecular architecture of the inner membrane of mitochondria from rat liver: a combined biochemical and stereological study. Journal of Cell Biology, 102(1):97–??, January 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/102/1/
97.

Sorensen:1985:AMP

[SCP85] R. A. Sorensen, M. S. Cyert, and R. A. Pedersen. Active maturation-promoting factor is present in mature mouse oocytes. Journal of Cell Biology, 100(5):1637–??, May 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/100/5/1637.

Schulman:1987:TCT

[SCRA87] I. G. Schulman, R. G. Cook, R. Richman, and C. D. Allis. Tetrahymena contain two distinct and unusual high mobility group (HMG)-like proteins. Journal of Cell Biology, 104(6):1485– ??, June 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
104/6/1485.

Stokes:1987:VTA

[SD87] D. L. Stokes and D. J. DeRosier. The variable twist of actin and its modulation by actin-binding proteins. Journal of Cell Biology, 104(4):1005–??, January 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/104/4/1005.

Suprenant:1988:RIM

[SD88] K. A. Suprenant and W. L. Dentler. Release of intact microtubule-capping structures from Tetrahymena cilia. Journal of Cell Biology, 107(6):2259–??, December 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/107/6/2259.

Sher:1989:HNC

[SDDZ^{+}89] E. Sher, S. Denis-Donini, A. Zanini, C. Bisiani, and F. Clementi. Human neuroblastoma cells acquire regulated secretory properties and different sensitivity to Ca^{2+} and alpha-latrotoxin after exposure to differentiating agents. Journal of Cell Biology, 108(6):2291–??, June 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/108/6/2291.

Shimizu:1985:AAM

[SDMF85] T. Shimizu, J. E. Dennis, T. Masaki, and D. A. Fischman. Axial arrangement of the myosin rod in vertebrate thick filaments: immunoelectron microscopy with a monoclonal antibody to light meromyosin. Journal of Cell Biology, 101(3):1115–??, September 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/101/3/
1115.

Serrano:1987:TPC

[SDNWA87] L. Serrano, J. Díaz-Nido, F. Wandosell, and J. Avila. Tubulin phosphorylation by casein kinase II is similar to that found in vivo. Journal of Cell Biology, 105(4):1731–??, October 1987. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/105/4/1731.

Southwick:1989:PLA

[SDPZ89] F. S. Southwick, G. A. Dabiri, M. Paschetto, and S. H. Zigmond. Polymorphonuclear leukocyte adherence induces actin polymerization by a transduction pathway which differs from that used by chemoattractants. Journal of Cell Biology, 109(4):1561–??, October 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/109/4/
1561.

Spiegelman:1988:FPR

[SDR^{+}88] B. M. Spiegelman, R. J. Distel, H. S. Ro, B. S. Rosen, and B. Satterberg. fos protooncogene and the regulation of gene expression in adipocyte differentiation. Journal of Cell Biology, 107 (3):829–??, September 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/107/3/829.

Seibert:1987:SLO

[SDS87] M. Seibert, M. DeWit, and L. A. Staehelin. Structural localization of the O2-evolving apparatus to multimeric (tetrameric) particles on the lumenal surface of freeze-etched photosynthetic membranes. Journal of Cell Biology, 105(5):2257–??, November 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/105/5/
2257.

Schlaeppi:1989:MCP

[SDS89] K. Schlaeppi, J. Deflorin, and T. Seebeck. The major component of the paraflagellar rod of Trypanosoma brucei is a helical protein that is encoded by two identical, tandemly linked genes. Journal of Cell Biology, 109(4):1695–??, October 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/109/4/1695.

Schuster:1986:TIT

[SDSO86] G. Schuster, M. Dewit, L. A. Staehelin, and I. Ohad. Transient inactivation of the thylakoid photosystem II light-harvesting protein kinase system and concomitant changes in intramembrane particle size during photoinhibition of Chlamydomonas reinhardtii. Journal of Cell Biology, 103(1):71–??, July 1986. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/103/1/71.

Strous:1985:ELA

[SDZB^{+}85] G. J. Strous, A. Du Maine, J. E. Zijderhand-Bleekemolen, J. W. Slot, and A. L. Schwartz. Effect of lysosomotropic amines on the secretory pathway and on the recycling of the asialoglycoprotein receptor in human hepatoma cells. Journal of Cell Biology, 101 (2):531–??, August 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/101/2/531.

Silver:1985:RMA

[SE85] G. Silver and J. D. Etlinger. Regulation of myofibrillar accumulation in chick muscle cultures: evidence for the involvement of calcium and lysosomes in non-uniform turnover of contractile proteins. Journal of Cell Biology, 101(6):2383–??, December 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/101/6/

Swezey:1986:RGP

[SE86] R. R. Swezey and D. Epel. Regulation of glucose-6-phosphate dehydrogenase activity in sea urchin eggs by reversible association with cell structural elements. Journal of Cell Biology, 103 (4):1509–??, October 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/103/4/1509.

Snell:1989:RSS

[SEB89] W. J. Snell, W. A. Eskue, and M. J. Buchanan. Regulated secretion of a serine protease that activates an extracellular matrix-degrading metalloprotease during fertilization in Chlamydomonas. Journal of Cell Biology, 109(4):1689–??, October 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/109/4/1689.

Swanson:1986:AER

[SEG^{+}86] M. E. Swanson, A. M. Elste, S. M. Greenberg, J. H. Schwartz, T. H. Aldrich, and M. E. Furth. Abundant expression of ras proteins in Aplysia neurons. Journal of Cell Biology, 103(2):485–??, August 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
103/2/485.

Schauer:1985:ISM

[SEGS85] I. Schauer, S. Emr, C. Gross, and R. Schekman. Invertase signal and mature sequence substitutions that delay intercompartmental transport of active enzyme. Journal of Cell Biology, 100 (5):1664–??, May 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/100/5/1664.

Stow:1987:DPB

[SF87] J. L. Stow and M. G. Farquhar. Distinctive populations of basement membrane and cell membrane heparan sulfate proteoglycans are produced by cultured cell lines. Journal of Cell Biology, 105(1):529–??, July 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/1/529.

Sale:1988:IBH

[SF88a] W. S. Sale and L. A. Fox. Isolated beta-heavy chain subunit of dynein translocates microtubules in vitro. Journal of Cell Biology, 107(5):1793–??, November 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/107/5/1793.

Stollberg:1988:ARC

[SF88b] J. Stollberg and S. E. Fraser. Acetylcholine receptors and concanavalin A-binding sites on cultured Xenopus muscle cells: electrophoresis, diffusion, and aggregation. Journal of Cell Biology, 107(4):1397–??, October 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/107/4/1397.

Smith:1989:IDN

[SF89] D. E. Smith and P. A. Fisher. Interconversion of Drosophila nuclear lamin isoforms during oogenesis, early embryogenesis, and upon entry of cultured cells into mitosis. Journal of Cell Biology, 108(2):255–??, February 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/108/2/255.

Schweizer:1988:IMA

[SFB^{+}88] A. Schweizer, J. A. Fransen, T. Bächi, L. Ginsel, and H. P. Hauri. Identification, by a monoclonal antibody, of a 53-kD protein associated with a tubulo-vesicular compartment at the cis-side of the Golgi apparatus. Journal of Cell Biology, 107(5): 1643–??, November 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/107/5/1643.

Small:1986:LFS

[SFD86a] J. V. Small, D. O. Fürst, and J. De Mey. Localization of filamin in smooth muscle. Journal of Cell Biology, 102(1):210–??, January 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/102/1/
210.

Steffen:1986:AFS

[SFD^{+}86b] W. Steffen, H. Fuge, R. Dietz, M. Bastmeyer, and G. Müller. Aster-free spindle poles in insect spermatocytes: evidence for chromosome-induced spindle formation? Journal of Cell Biology, 102(5):1679–??, May 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/102/5/1679.

Segall:1987:SCM

[SFG87] J. E. Segall, P. R. Fisher, and G. Gerisch. Selection of chemotaxis mutants of Dictyostelium discoideum. Journal of Cell Biology, 104(1):151–??, January 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/104/1/151.

Schiff:1986:HIH

[SFJU86] J. M. Schiff, M. M. Fisher, A. L. Jones, and B. J. Underdown. Human IgA as a heterovalent ligand: switching from the asialoglycoprotein receptor to secretory component during transport across the rat hepatocyte. Journal of Cell Biology, 102(3):920–??, March 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/102/3/920.

Schmid:1989:AES

[SFK^{+}89] S. Schmid, R. Fuchs, M. Kielian, A. Helenius, and I. Mellman. Acidification of endosome subpopulations in wild-type Chinese hamster ovary cells and temperature-sensitive acidificationdefective mutants. Journal of Cell Biology, 108(4):1291–??, April 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/108/4/
1291.

Sanderson:1986:FPF

[SFLM86] R. D. Sanderson, J. M. Fitch, T. R. Linsenmayer, and R. Mayne. Fibroblasts promote the formation of a continuous basal lamina during myogenesis in vitro. Journal of Cell Biology, 102 (3):740–??, March 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/102/3/740.

Shain:1987:MAC

[SFMT87] W. Shain, D. S. Forman, V. Madelian, and J. N. Turner. Morphology of astroglial cells is controlled by beta-adrenergic receptors. Journal of Cell Biology, 105(5):2307–??, November 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/105/5/2307.

Spray:1987:PGI

[SFS^{+}87] D. C. Spray, M. Fujita, J. C. Saez, H. Choi, T. Watanabe, E. Hertzberg, L. C. Rosenberg, and L. M. Reid. Proteoglycans and glycosaminoglycans induce gap junction synthesis and function in primary liver cultures. Journal of Cell Biology, 105(1):541–??, July 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/1/541.

Seagull:1987:MDA

[SFW87] R. W. Seagull, M. M. Falconer, and C. A. Weerdenburg. Microfilaments: dynamic arrays in higher plant cells. Journal of Cell Biology, 104(4):995–??, January 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/104/4/995.

Suprynowicz:1986:FCF

[SG86] F. A. Suprynowicz and L. Gerace. A fractionated cell-free system for analysis of prophase nuclear disassembly. Journal of Cell Biology, 103(6):2073–??, December 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/103/6/2073.

Senior:1988:IMP

[SG88a] A. Senior and L. Gerace. Integral membrane proteins specific to the inner nuclear membrane and associated with the nuclear lamina. Journal of Cell Biology, 107(6):2029–??, December 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/107/6/2029.

Siliciano:1988:LTJ

[SG88b] J. D. Siliciano and D. A. Goodenough. Localization of the tight junction protein, ZO-1, is modulated by extracellular calcium and cell–cell contact in Madin–Darby canine kidney epithelial cells. Journal of Cell Biology, 107(6):2389–??, December 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/107/6/2389.

Siminoski:1986:UTT

[SGB^{+}86] K. Siminoski, P. Gonnella, J. Bernanke, L. Owen, M. Neutra, and R. A. Murphy. Uptake and transepithelial transport of nerve growth factor in suckling rat ileum. Journal of Cell Biology, 103 (5):1979–??, November 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/103/5/1979.

Sammak:1987:MDV

[SGB87] P. J. Sammak, G. J. Gorbsky, and G. G. Borisy. Microtubule dynamics in vivo: a test of mechanisms of turnover. Journal of Cell Biology, 104(3):395–??, March 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/104/3/395.

Smith:1987:BID

[SGBF87] D. E. Smith, Y. Gruenbaum, M. Berrios, and P. A. Fisher. Biosynthesis and interconversion of Drosophila nuclear lamin isoforms during normal growth and in response to heat shock. Journal of Cell Biology, 105(2):771–??, August 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/105/2/771.

Sasse:1987:AAT

[SGBG87] R. Sasse, M. C. Glyn, C. R. Birkett, and K. Gull. Acetylated alpha-tubulin in Physarum: immunological characterization of the isotype and its usage in particular microtubular organelles. Journal of Cell Biology, 104(1):41–??, January 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/104/1/41.

Stoorvogel:1989:RBI

[SGG^{+}89] W. Stoorvogel, H. J. Geuze, J. M. Griffith, A. L. Schwartz, and G. J. Strous. Relations between the intracellular pathways of the receptors for transferrin, asialoglycoprotein, and mannose 6phosphate in human hepatoma cells. Journal of Cell Biology, 108(6):2137–??, June 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/108/6/2137.

Stoorvogel:1988:PET

[SGGS88] W. Stoorvogel, H. J. Geuze, J. M. Griffith, and G. J. Strous. The pathways of endocytosed transferrin and secretory protein are connected in the trans-Golgi reticulum. Journal of Cell Biology, 106(6):1821–??, June 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/106/6/1821.

Sale:1985:SIS

[SGH85] W. S. Sale, U. W. Goodenough, and J. E. Heuser. The substructure of isolated and in situ outer dynein arms of sea urchin sperm flagella. Journal of Cell Biology, 101(4):1400–??, October 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/101/4/
1400.

Spray:1986:EPG

[SGM^{+}86] D. C. Spray, R. D. Ginzberg, E. A. Morales, Z. Gatmaitan, and I. M. Arias. Electrophysiological properties of gap junctions between dissociated pairs of rat hepatocytes. Journal of Cell Biology, 103(1):135–??, July 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/103/1/135.

Solowska:1989:ENM

[SGM^{+}89] J. Solowska, J. L. Guan, E. E. Marcantonio, J. E. Trevithick, C. A. Buck, and R. O. Hynes. Expression of normal and mutant avian integrin subunits in rodent cells. Journal of Cell Biology, 109(2):853–??, August 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/2/853.

Stolpen:1988:TNF

[SGP88] A. H. Stolpen, D. E. Golan, and J. S. Pober. Tumor necrosis factor and immune interferon act in concert to slow the lateral diffusion of proteins and lipids in human endothelial cell membranes. Journal of Cell Biology, 107(2):781–??, August 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/107/2/781.

Sabanero:1985:CDP

[SGRM85] M. Sabanero, A. Gonzalez-Robles, and I. Meza. Characterization of a 36,000-dalton protein from the surface of Madin–Darby canine kidney cells involved in cell attachment and spreading. Journal of Cell Biology, 100(6):2001–??, June 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/100/6/2001.

Schermer:1986:DRE

[SGS86] A. Schermer, S. Galvin, and T. T. Sun. Differentiation-related expression of a major 64K corneal keratin in vivo and in culture suggests limbal location of corneal epithelial stem cells. Journal of Cell Biology, 103(1):49–??, July 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/103/1/49.

Stoorvogel:1987:SET

[SGS87] W. Stoorvogel, H. J. Geuze, and G. J. Strous. Sorting of endocytosed transferrin and asialoglycoprotein occurs immediately after internalization in HepG2 cells. Journal of Cell Biology, 104(5):1261–??, May 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/104/5/1261.

Sugrue:1989:ITX

[SGS^{+}89] S. P. Sugrue, M. K. Gordon, J. Seyer, B. Dublet, M. van der Rest, and B. R. Olsen. Immunoidentification of type XII collagen in embryonic tissues. Journal of Cell Biology, 109(2):939–??, August 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/109/2/
939.

Sugrue:1986:IEM

[SH86] S. P. Sugrue and E. D. Hay. The identification of extracellular matrix (ECM) binding sites on the basal surface of embryonic corneal epithelium and the effect of ECM binding on epithelial collagen production. Journal of Cell Biology, 102(5):1907–??, May 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/102/5/
1907.

Saibil:1987:OTE

[SH87a] H. Saibil and E. Hewat. Ordered transmembrane and extracellular structure in squid photoreceptor microvilli. Journal of Cell Biology, 105(1):19–??, July 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/105/1/19.

Shiomura:1987:CMA

[SH87b] Y. Shiomura and N. Hirokawa. Colocalization of microtubuleassociated protein 1A and microtubule-associated protein 2 on neuronal microtubules in situ revealed with double-label immunoelectron microscopy. Journal of Cell Biology, 104(6):1575–??, June 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/104/6/

Sappino:1989:PAT

[SHBV89] A. P. Sappino, J. Huarte, D. Belin, and J. D. Vassalli. Plasminogen activators in tissue remodeling and invasion: mRNA localization in mouse ovaries and implanting embryos. Journal of Cell Biology, 109(5):2471–??, November 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/109/5/2471.

Salzer:1987:AAS

[SHC87] J. L. Salzer, W. P. Holmes, and D. R. Colman. The amino acid sequences of the myelin-associated glycoproteins: homology to the immunoglobulin gene superfamily. Journal of Cell Biology, 104(4):957–??, January 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/104/4/957.

Senior:1985:DCM

[SHGD85] R. M. Senior, J. S. Huang, G. L. Griffin, and T. F. Deuel. Dissociation of the chemotactic and mitogenic activities of plateletderived growth factor by human neutrophil elastase. Journal of Cell Biology, 100(2):351–??, February 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/100/2/351.

Schindler:1985:LDN

[SHH85] M. Schindler, J. F. Holland, and M. Hogan. Lateral diffusion in nuclear membranes. Journal of Cell Biology, 100(5):1408–??, May 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/100/5/
1408.

Sztul:1987:IRO

[SHK^{+}87] E. S. Sztul, J. P. Hendrick, J. P. Kraus, D. Wall, F. Kalousek, and L. E. Rosenberg. Import of rat ornithine transcarbamylase precursor into mitochondria: two-step processing of the leader peptide. Journal of Cell Biology, 105(6):2631–??, December 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/105/6/2631.

Stamatoglou:1987:RHS

[SHL87] S. C. Stamatoglou, R. C. Hughes, and U. Lindahl. Rat hepatocytes in serum-free primary culture elaborate an extensive extracellular matrix containing fibrin and fibronectin. Journal of Cell Biology, 105(5):2417–??, November 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/105/5/2417.

Sistonen:1989:ANT

[SHL^{+}89] L. Sistonen, E. Hölttä, H. Lehväslaiho, L. Lehtola, and K. Alitalo. Activation of the neu tyrosine kinase induces the fos/ jun transcription factor complex, the glucose transporter and ornithine decarboxylase. Journal of Cell Biology, 109(5):1911– ??, November 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/5/1911.

Schwaiger:1987:CTC

[SHN87] M. Schwaiger, V. Herzog, and W. Neupert. Characterization of translocation contact sites involved in the import of mitochondrial proteins. Journal of Cell Biology, 105(1):235–??, July 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/105/1/
235.

Sztul:1985:BPIa

[SHP85a] E. S. Sztul, K. E. Howell, and G. E. Palade. Biogenesis of the polymeric IgA receptor in rat hepatocytes. I. Kinetic studies of its intracellular forms. Journal of Cell Biology, 100(4):1248–??, April 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
100/4/1248.

Sztul:1985:BPIb

[SHP85b] E. S. Sztul, K. E. Howell, and G. E. Palade. Biogenesis of the polymeric IgA receptor in rat hepatocytes. II. Localization of its intracellular forms by cell fractionation studies. Journal of Cell Biology, 100(4):1255–??, April 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/100/4/1255.

Scullion:1987:EMT

[SHP^{+}87] B. F. Scullion, Y. Hou, L. Puddington, J. K. Rose, and K. Jacobson. Effects of mutations in three domains of the vesicular stomatitis viral glycoprotein on its lateral diffusion in the plasma membrane. Journal of Cell Biology, 105(1):69–??, July 1987. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/105/1/69.

Standart:1986:DAR

[SHR86] N. Standart, T. Hunt, and J. V. Ruderman. Differential accumulation of ribonucleotide reductase subunits in clam oocytes: the large subunit is stored as a polypeptide, the small subunit as untranslated mRNA. Journal of Cell Biology, 103(6): 2129–??, December 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/103/6/2129.

Sjolund:1988:ASM

[SHS^{+}88] M. Sjölund, U. Hedin, T. Sejersen, C. H. Heldin, and J. Thyberg. Arterial smooth muscle cells express platelet-derived growth factor (PDGF) A chain mRNA, secrete a PDGF-like mitogen, and bind exogenous PDGF in a phenotype- and growth statedependent manner. Journal of Cell Biology, 106(2):403–??, February 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
106/2/403.

Swoboda:1989:CHC

[SHSvdM89] B. Swoboda, R. Holmdahl, H. Stöss, and K. von der Mark. Cellular heterogeneity in cultured human chondrocytes identified by antibodies specific for alpha 2(XI) collagen chains. Journal of Cell Biology, 109(3):1363–??, September 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/109/3/1363.

Saito:1988:UCR

[SIR^{+}88] A. Saito, M. Inui, M. Radermacher, J. Frank, and S. Fleischer. Ultrastructure of the calcium release channel of sarcoplasmic reticulum. Journal of Cell Biology, 107(1):211–??, July 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/107/1/211.

Savagner:1986:HHP

[SIYT86] P. Savagner, B. A. Imhof, K. M. Yamada, and J. P. Thiery. Homing of hemopoietic precursor cells to the embryonic thymus: characterization of an invasive mechanism induced by chemotactic peptides. Journal of Cell Biology, 103(6):2715–??, December 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/103/6/

Schindler:1986:NAM

[SJ86] M. Schindler and L. W. Jiang. Nuclear actin and myosin as control elements in nucleocytoplasmic transport. Journal of Cell Biology, 102(3):859–??, March 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/102/3/859.

Schindler:1987:EGF

[SJ87] M. Schindler and L. W. Jiang. Epidermal growth factor and insulin stimulate nuclear pore-mediated macromolecular transport in isolated rat liver nuclei. Journal of Cell Biology, 104 (4):849–??, January 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/104/4/849.

Saunders:1989:MCS

[SJOB89] S. Saunders, M. Jalkanen, S. O’Farrell, and M. Bernfield. Molecular cloning of syndecan, an integral membrane proteoglycan. Journal of Cell Biology, 108(4):1547–??, April 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/108/4/1547.

Shen:1986:UIS

[SJS86] B. W. Shen, R. Josephs, and T. L. Steck. Ultrastructure of the intact skeleton of the human erythrocyte membrane. Journal of Cell Biology, 102(3):997–??, March 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/102/3/997.

Schweitzer:1985:SPH

[SK85] E. S. Schweitzer and R. B. Kelly. Selective packaging of human growth hormone into synaptic vesicles in a rat neuronal (PC12) cell line. Journal of Cell Biology, 101(2):667–??, August 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/101/2/
667.

Schulze:1986:MDI

[SK86] E. Schulze and M. Kirschner. Microtubule dynamics in interphase cells. Journal of Cell Biology, 102(3):1020–??, March 1986. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/102/3/1020.

Schulze:1987:DSP

[SK87] E. Schulze and M. Kirschner. Dynamic and stable populations of microtubules in cells. Journal of Cell Biology, 104(2): 277–??, February 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/104/2/277.

Sellitto:1988:DMC

[SK88] C. Sellitto and R. Kuriyama. Distribution of a matrix component of the midbody during the cell cycle in Chinese hamster ovary cells. Journal of Cell Biology, 106(2):431–??, February 1988. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/106/2/431.

Sibley:1986:TMM

[SKAW86] L. D. Sibley, J. L. Krahenbuhl, G. M. Adams, and E. Weidner. Toxoplasma modifies macrophage phagosomes by secretion of a vesicular network rich in surface proteins. Journal of Cell Biology, 103(3):867–??, September 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/103/3/867.

Sung:1985:IEL

[SKC85a] L. A. Sung, E. A. Kabat, and S. Chien. Interaction energies in lectin-induced erythrocyte aggregation. Journal of Cell Biology, 101(2):652–??, August 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/101/2/652.

Sung:1985:ILM

[SKC85b] L. A. Sung, E. A. Kabat, and S. Chien. Interaction of lectins with membrane receptors on erythrocyte surfaces. Journal of Cell Biology, 101(2):646–??, August 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/101/2/646.

Snell:1986:MAB

[SKC^{+}86] W. J. Snell, M. G. Kosfiszer, A. Clausell, N. Perillo, S. Imam, and G. Hunnicutt. A monoclonal antibody that blocks adhesion of Chlamydomonas mt+ gametes. Journal of Cell Biology, 103 (6):2449–??, December 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/103/6/2449.

Stoler:1988:UMA

[SKDF88] A. Stoler, R. Kopan, M. Duvic, and E. Fuchs. Use of monospecific antisera and cRNA probes to localize the major changes in keratin expression during normal and abnormal epidermal differentiation. Journal of Cell Biology, 107(2):427–??, August 1988. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/107/2/427.

Sakai:1986:FNK

[SKE86] L. Y. Sakai, D. R. Keene, and E. Engvall. Fibrillin, a new 350-kD glycoprotein, is a component of extracellular microfibrils. Journal of Cell Biology, 103(6):2499–??, December 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/103/6/2499.

Schreiber:1985:UFE

[SKK^{+}85] A. B. Schreiber, J. Kenney, J. Kowalski, K. A. Thomas, G. Gimenez-Gallego, M. Rios-Candelore, J. Di Salvo, D. Barritault, J. Courty, and Y. Courtois. A unique family of endothelial cell polypeptide mitogens: the antigenic and receptor crossreactivity of bovine endothelial cell growth factor, brain-derived acidic fibroblast growth factor, and eye-derived growth factor–II. Journal of Cell Biology, 101(4):1623–??, October 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/101/4/1623.

Stewart:1985:TDR

[SKL85] M. Stewart, R. W. Kensler, and R. J. Levine. Three-dimensional reconstruction of thick filaments from Limulus and scorpion muscle. Journal of Cell Biology, 101(2):402–??, August 1985. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/101/2/402.

Sakai:1986:TVC

[SKMB86] L. Y. Sakai, D. R. Keene, N. P. Morris, and R. E. Burgeson. Type VII collagen is a major structural component of anchoring fibrils. Journal of Cell Biology, 103(4):1577–??, October 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/103/4/1577.

Singer:1987:FCA

[SKS^{+}87] I. I. Singer, D. W. Kawka, S. Scott, R. A. Mumford, and M. W. Lark. The fibronectin cell attachment sequence Arg–Gly–Asp– Ser promotes focal contact formation during early fibroblast attachment and spreading. Journal of Cell Biology, 104(3):573–??, March 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
104/3/573.

Stow:1985:HSP

[SKU^{+}85] J. L. Stow, L. Kjéllen, E. Unger, M. Höök, and M. G. Farquhar. Heparan sulfate proteoglycans are concentrated on the sinusoidal plasmalemmal domain and in intracellular organelles of hepatocytes. Journal of Cell Biology, 100(3):975–??, March 1985. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/100/3/975.

Schmid:1985:ILS

[SL85a] T. M. Schmid and T. F. Linsenmayer. Immunohistochemical localization of short chain cartilage collagen (type X) in avian tissues. Journal of Cell Biology, 100(2):598–??, February 1985. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/100/2/598.

Schubert:1985:ICS

[SL85b] D. Schubert and M. LaCorbiere. Isolation of a cell-surface receptor for chick neural retina adherons. Journal of Cell Biology, 100(1):56–??, January 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/100/1/56.

Schubert:1985:IAM

[SL85c] D. Schubert and M. LaCorbiere. Isolation of an adhesionmediating protein from chick neural retina adherons. Journal of Cell Biology, 101(3):1071–??, September 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/101/3/1071.

Segal:1985:PIC

[SL85d] R. A. Segal and D. J. Luck. Phosphorylation in isolated Chlamydomonas axonemes: a phosphoprotein may mediate the Ca^{2+}- dependent photophobic response. Journal of Cell Biology, 101 (5):1702–??, November 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/101/5/1702.

Schwartz:1986:BAA

[SL86a] M. A. Schwartz and E. J. Luna. Binding and assembly of actin filaments by plasma membranes from Dictyostelium discoideum. Journal of Cell Biology, 102(6):2067–??, June 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/102/6/2067.

Staufenbiel:1986:APD

[SL86b] M. Staufenbiel and E. Lazarides. Assembly of protein 4.1 during chicken erythroid differentiation. Journal of Cell Biology, 102 (4):1157–??, April 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/102/4/1157.

Schmit:1987:CDC

[SL87a] A. C. Schmit and A. M. Lambert. Characterization and dynamics of cytoplasmic F-actin in higher plant endosperm cells during interphase, mitosis, and cytokinesis. Journal of Cell Biology, 105 (5):2157–??, November 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/5/2157.

Small:1987:ICT

[SL87b] G. M. Small and P. B. Lazarow. Import of the carboxy-terminal portion of acyl–CoA oxidase into peroxisomes of Candida tropicalis. Journal of Cell Biology, 105(1):247–??, July 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/105/1/247.

Schwartz:1988:HAB

[SL88] M. A. Schwartz and E. J. Luna. How actin binds and assembles onto plasma membranes from Dictyostelium discoideum. Journal of Cell Biology, 107(1):201–??, July 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/107/1/201.

Sawtell:1989:CDS

[SL89] N. M. Sawtell and J. L. Lessard. Cellular distribution of smooth muscle actins during mammalian embryogenesis: expression of the alpha-vascular but not the gamma-enteric isoform in differentiating striated myocytes. Journal of Cell Biology, 109(6): 2929–??, December 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/6/2929.

Schubert:1987:MIH

[SLB87] D. Schubert, N. Ling, and A. Baird. Multiple influences of a heparin-binding growth factor on neuronal development. Journal of Cell Biology, 104(3):635–??, March 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/104/3/635.

Sarndahl:1989:ALR

[SLB^{+}89] E. Särndahl, M. Lindroth, T. Bengtsson, M. Fällman, J. Gustavsson, O. Stendahl, and T. Andersson. Association of ligandreceptor complexes with actin filaments in human neutrophils: a possible regulatory role for a G-protein. Journal of Cell Biology, 109(6):2791–??, December 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/109/6/2791.

Schubert:1986:CNR

[SLE86] D. Schubert, M. LaCorbiere, and F. Esch. A chick neural retina adhesion and survival molecule is a retinol-binding protein. Journal of Cell Biology, 102(6):2295–??, June 1986. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/102/6/2295.

Simon:1988:SHB

[SLG88] K. Simon, V. R. Lingappa, and D. Ganem. Secreted hepatitis B surface antigen polypeptides are derived from a transmembrane precursor. Journal of Cell Biology, 107(6):2163–??, December 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/107/6/
2163.

Schulze-Lohoff:1985:CDP

[SLHvF85] E. Schulze-Lohoff, A. Hasilik, and K. von Figura. Cathepsin D precursors in clathrin-coated organelles from human fibroblasts. Journal of Cell Biology, 101(3):824–??, September 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/101/3/824.

Shapland:1988:INA

[SLL88] C. Shapland, P. Lowings, and D. Lawson. Identification of new actin-associated polypeptides that are modified by viral transformation and changes in cell shape. Journal of Cell Biology, 107(1):153–??, July 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/107/1/153.

Singer:1989:UVC

[SLL89] R. H. Singer, G. L. Langevin, and J. B. Lawrence. Ultrastructural visualization of cytoskeletal mRNAs and their associated proteins using double-label in situ hybridization. Journal of Cell Biology, 108(6):2343–??, June 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/108/6/2343.

Saxton:1987:IMB

[SM87] W. M. Saxton and J. R. McIntosh. Interzone microtubule behavior in late anaphase and telophase spindles. Journal of Cell Biology, 105(2):875–??, August 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/105/2/875.

Salzman:1988:IFS

[SM88] N. H. Salzman and F. R. Maxfield. Intracellular fusion of sequentially formed endocytic compartments. Journal of Cell Biology, 106(4):1083–??, April 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/106/4/1083.

Salzman:1989:FAE

[SM89a] N. H. Salzman and F. R. Maxfield. Fusion accessibility of endocytic compartments along the recycling and lysosomal endocytic pathways in intact cells. Journal of Cell Biology, 109(5): 2097–??, November 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/5/2097.

Sanz:1989:SYP

[SM89b] P. Sanz and D. I. Meyer. Secretion in yeast: preprotein binding to a membrane receptor and ATP-dependent translocation are sequential and separable events in vitro. Journal of Cell Biology, 108(6):2101–??, June 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/108/6/2101.

Smith:1985:TES

[SMA85] R. G. Smith, J. McManaman, and S. H. Appel. Trophic effects of skeletal muscle extracts on ventral spinal cord neurons in vitro: separation of a protein with morphologic activity from proteins with cholinergic activity. Journal of Cell Biology, 101 (4):1608–??, October 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/101/4/1608.

Stieger:1988:DAT

[SMB^{+}88] B. Stieger, K. Matter, B. Baur, K. Bucher, M. Höchli, and H. P. Hauri. Dissection of the asynchronous transport of intestinal microvillar hydrolases to the cell surface. Journal of Cell Biology, 106(6):1853–??, June 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/106/6/1853.

Schindler:1989:RDL

[SMC89] M. Schindler, S. Meiners, and D. A. Cheresh. RGD-dependent linkage between plant cell wall and plasma membrane: consequences for growth. Journal of Cell Biology, 108(5):1955–??, May 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/108/5/
1955.

Sarkar:1989:ISI

[SME^{+}89] D. P. Sarkar, S. J. Morris, O. Eidelman, J. Zimmerberg, and R. Blumenthal. Initial stages of influenza hemagglutinin-induced cell fusion monitored simultaneously by two fluorescent events: cytoplasmic continuity and lipid mixing. Journal of Cell Biology, 109(1):113–??, July 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/1/113.

Salpeter:1988:DEA

[SMH88] M. M. Salpeter, M. Marchaterre, and R. Harris. Distribution of extrajunctional acetylcholine receptors on a vertebrate muscle: evaluated by using a scanning electron microscope autoradiographic procedure. Journal of Cell Biology, 106(6):2087–??, June 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/106/6/

Smit:1987:LSP

[Smi87] J. Smit. Localizing the subunit pool for the temporally regulated polar pili of Caulobacter crescentus. Journal of Cell Biology, 105 (4):1821–??, October 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/4/1821.

Sanger:1986:MLC

[SMPS86] J. M. Sanger, B. Mittal, M. B. Pochapin, and J. W. Sanger. Myofibrillogenesis in living cells microinjected with fluorescently labeled alpha-actinin. Journal of Cell Biology, 102(6):2053–??, June 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/102/6/
2053.

Sluder:1986:RCN

[SMR86] G. Sluder, F. J. Miller, and C. L. Rieder. The reproduction of centrosomes: nuclear versus cytoplasmic controls. Journal of Cell Biology, 103(5):1873–??, November 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/103/5/1873.

Saksela:1987:OEB

[SMR87] O. Saksela, D. Moscatelli, and D. B. Rifkin. The opposing effects of basic fibroblast growth factor and transforming growth factor beta on the regulation of plasminogen activator activity in capillary endothelial cells. Journal of Cell Biology, 105(2):957–??, August 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/105/2/
957.

Stenberg:1985:PAG

[SMS^{+}85] P. E. Stenberg, R. P. McEver, M. A. Shuman, Y. V. Jacques, and D. F. Bainton. A platelet alpha-granule membrane protein (GMP-140) is expressed on the plasma membrane after activation. Journal of Cell Biology, 101(3):880–??, September 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/101/3/880.

Sheehan:1988:SAR

[SMS^{+}88] M. A. Sheehan, A. D. Mills, A. M. Sleeman, R. A. Laskey, and J. J. Blow. Steps in the assembly of replication-competent nuclei in a cell-free system from Xenopus eggs. Journal of Cell Biology, 106(1):1–??, January 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/106/1/1.

Saksela:1988:ECD

[SMSR88] O. Saksela, D. Moscatelli, A. Sommer, and D. B. Rifkin. Endothelial cell-derived heparan sulfate binds basic fibroblast growth factor and protects it from proteolytic degradation. Journal of Cell Biology, 107(2):743–??, August 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/107/2/743.

Salas:1986:MAF

[SMVS^{+}86] P. J. Salas, D. E. Misek, D. E. Vega-Salas, D. Gundersen, M. Cereijido, and E. Rodriguez-Boulan. Microtubules and actin filaments are not critically involved in the biogenesis of epithelial cell surface polarity. Journal of Cell Biology, 102(5):1853–??, May 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/102/5/
1853.

Sporn:1989:DPC

[SMW89] L. A. Sporn, V. J. Marder, and D. D. Wagner. Differing polarity of the constitutive and regulated secretory pathways for von Willebrand factor in endothelial cells. Journal of Cell Biology, 108(4):1283–??, April 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/108/4/1283.

Salminen:1989:SPR

[SN89] A. Salminen and P. J. Novick. The Sec15 protein responds to the function of the GTP binding protein, Sec4, to control vesicular traffic in yeast. Journal of Cell Biology, 109(3):1023–??, September 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/109/3/
1023.

Simons:1986:CVC

[SNC^{+}86] J. W. Simons, S. J. Noga, P. M. Colombani, W. E. Beschorner, D. S. Coffey, and A. D. Hess. Cyclosporine A, an in vitro calmodulin antagonist, induces nuclear lobulations in human T cell lymphocytes and monocytes. Journal of Cell Biology, 102 (1):145–??, January 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/102/1/145.

Saga:1987:PDF

[SNCY87] S. Saga, K. Nagata, W. T. Chen, and K. M. Yamada. pHdependent function, purification, and intracellular location of a major collagen-binding glycoprotein. Journal of Cell Biology, 105(1):517–??, July 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/1/517.

Sato:1985:EUM

[SNNK85] C. Sato, K. Nishizawa, T. Nakayama, and T. Kobayashi. Effect upon mitogenic stimulation of calcium-dependent phosphorylation of cytoskeleton-associated 350,000- and 80,000-mol-wt polypeptides in quiescent 3Y1 cells. Journal of Cell Biology, 100(3):748–??, March 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/100/3/748.

Steinberg:1987:MPP

[SNSS87] T. H. Steinberg, A. S. Newman, J. A. Swanson, and S. C. Silverstein. Macrophages possess probenecid-inhibitable organic anion transporters that remove fluorescent dyes from the cytoplasmic matrix. Journal of Cell Biology, 105(6):2695–??, December 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/105/6/
2695.

Snyder:1989:SPY

[Sny89] M. Snyder. The SPA2 protein of yeast localizes to sites of cell growth. Journal of Cell Biology, 108(4):1419–??, April 1989. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/108/4/1419.

Sandvig:1989:ECP

[SOB^{+}89] K. Sandvig, S. Olsnes, J. E. Brown, O. W. Petersen, and B. van Deurs. Endocytosis from coated pits of Shiga toxin: a glycolipidbinding protein from Shigella dysenteriae 1. Journal of Cell Biology, 108(4):1331–??, April 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/108/4/1331.

Sklar:1985:RAP

[SOP85] L. A. Sklar, G. M. Omann, and R. G. Painter. Relationship of actin polymerization and depolymerization to light scattering in human neutrophils: dependence on receptor occupancy and intracellular Ca++. Journal of Cell Biology, 101(3): 1161–??, September 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/101/3/1161.

Sandvig:1987:ACI

[SOPvD87] K. Sandvig, S. Olsnes, O. W. Petersen, and B. van Deurs. Acidification of the cytosol inhibits endocytosis from coated pits. Journal of Cell Biology, 105(2):679–??, August 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/105/2/679.

Sowers:1986:LLF

[Sow86] A. E. Sowers. A long-lived fusogenic state is induced in erythrocyte ghosts by electric pulses. Journal of Cell Biology, 102 (4):1358–??, April 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/102/4/1358.

Severinsson:1985:ACS

[SP85] L. Severinsson and P. A. Peterson. Abrogation of cell surface expression of human class I transplantation antigens by an adenovirus protein in Xenopus laevis oocytes. Journal of Cell Biology, 101(2):540–??, August 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/101/2/540.

Suzuki:1986:BIB

[SP86] S. Suzuki and G. H. Pollack. Bridgelike interconnections between thick filaments in stretched skeletal muscle fibers observed by the freeze-fracture method. Journal of Cell Biology, 102 (3):1093–??, March 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/102/3/1093.

Schafer:1988:ITI

[SP88] B. W. Schäfer and J. C. Perriard. Intracellular targeting of isoproteins in muscle cytoarchitecture. Journal of Cell Biology, 106 (4):1161–??, April 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/106/4/1161.

Sinard:1989:EHC

[SP89] J. H. Sinard and T. D. Pollard. The effect of heavy chain phosphorylation and solution conditions on the assembly of Acanthamoeba myosin–II. Journal of Cell Biology, 109(4):1529–??, October 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
109/4/1529.

Saraste:1987:ARP

[SPF87] J. Saraste, G. E. Palade, and M. G. Farquhar. Antibodies to rat pancreas Golgi subfractions: identification of a 58-kD cis-Golgi protein. Journal of Cell Biology, 105(5):2021–??, November 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/105/5/
2021.

Spurck:1987:MAE

[SPH87] T. P. Spurck and J. D. Pickett-Heaps. On the mechanism of anaphase A: evidence that ATP is needed for microtubule disassembly and not generation of polewards force. Journal of Cell Biology, 105(4):1691–??, October 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/105/4/1691.

Simon:1987:TGF

[SPL87] K. Simon, E. Perara, and V. R. Lingappa. Translocation of globin fusion proteins across the endoplasmic reticulum membrane in Xenopus laevis oocytes. Journal of Cell Biology, 104(5):1165–??, May 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/104/5/
1165.

Smythe:1989:FCV

[SPLW89] E. Smythe, M. Pypaert, J. Lucocq, and G. Warren. Formation of coated vesicles from coated pits in broken A431 cells. Journal of Cell Biology, 108(3):843–??, March 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/108/3/843.

Sussman:1985:PEG

[SPO85] I. Sussman, R. Prettyman, and T. G. O’Brien. Phorbol esters and gene expression: the role of rapid changes in K^{+} transport in the induction of ornithine decarboxylase by 12-o-tetradecanoylphorbol-13-acetate in BALB/c 3T3 cells and a mutant cell line defective in Na + K + Cl- cotransport. Journal of Cell Biology, 101(6):2316–??, December 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/101/6/2316.

Seilheimer:1989:ALA

[SPS89a] B. Seilheimer, E. Persohn, and M. Schachner. Antibodies to the L1 adhesion molecule inhibit Schwann cell ensheathment of neurons in vitro. Journal of Cell Biology, 109(6):3095–??, December 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/109/6/
3095.

Seilheimer:1989:NCA

[SPS89b] B. Seilheimer, E. Persohn, and M. Schachner. Neural cell adhesion molecule expression is regulated by Schwann cell-neuron interactions in culture. Journal of Cell Biology, 108(5):1909–??, May 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/108/5/
1909.

Stephens:1989:APU

[SPT^{+}89] R. W. Stephens, J. Pöllänen, H. Tapiovaara, K. C. Leung, P. S. Sim, E. M. Salonen, E. Rønne, N. Behrendt, K. Danø, and A. Vaheri. Activation of pro-urokinase and plasminogen on human sarcoma cells: a proteolytic system with surface-bound reactants. Journal of Cell Biology, 108(5):1987–??, May 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/108/5/1987.

Shpetner:1988:CMA

[SPV88] H. S. Shpetner, B. M. Paschal, and R. B. Vallee. Characterization of the microtubule-activated ATPase of brain cytoplasmic dynein (MAP 1C). Journal of Cell Biology, 107(3):1001–??, September 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/107/3/

Stewart:1989:MIP

[SQM89] M. Stewart, R. A. Quinlan, and R. D. Moir. Molecular interactions in paracrystals of a fragment corresponding to the alphahelical coiled-coil rod portion of glial fibrillary acidic protein: evidence for an antiparallel packing of molecules and polymorphism related to intermediate filament structure. Journal of Cell Biology, 109(1):225–??, July 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/109/1/225.

Sluder:1985:CNR

[SR85a] G. Sluder and C. L. Rieder. Centriole number and the reproductive capacity of spindle poles. Journal of Cell Biology, 100 (3):887–??, March 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/100/3/887.

Sluder:1985:ESP

[SR85b] G. Sluder and C. L. Rieder. Experimental separation of pronuclei in fertilized sea urchin eggs: chromosomes do not organize a spindle in the absence of centrosomes. Journal of Cell Biology, 100(3):897–??, March 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/100/3/897.

Snider:1985:IMC

[SR85c] M. D. Snider and O. C. Rogers. Intracellular movement of cell surface receptors after endocytosis: resialylation of asialotransferrin receptor in human erythroleukemia cells. Journal of Cell Biology, 100(3):826–??, March 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/100/3/826.

Shelton:1986:SRB

[SR86a] D. L. Shelton and L. F. Reichardt. Studies on the regulation of beta-nerve growth factor gene expression in the rat iris: the level of mRNA-encoding nerve growth factor is increased in irises placed in explant cultures in vitro, but not in irises deprived of sensory or sympathetic innervation in vivo. Journal of Cell Biology, 102(5):1940–??, May 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/102/5/1940.

Snider:1986:MTA

[SR86b] M. D. Snider and O. C. Rogers. Membrane traffic in animal cells: cellular glycoproteins return to the site of Golgi mannosidase I. Journal of Cell Biology, 103(1):265–??, July 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/103/1/265.

Streeter:1987:FAR

[SR87] H. B. Streeter and D. A. Rees. Fibroblast adhesion to RGDS shows novel features compared with fibronectin. Journal of Cell Biology, 105(1):507–??, July 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/105/1/507.

Sato:1988:AAB

[SR88] Y. Sato and D. B. Rifkin. Autocrine activities of basic fibroblast growth factor: regulation of endothelial cell movement, plasminogen activator synthesis, and DNA synthesis. Journal of Cell Biology, 107(3):1199–??, September 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/107/3/1199.

Sato:1989:IEC

[SR89] Y. Sato and D. B. Rifkin. Inhibition of endothelial cell movement by pericytes and smooth muscle cells: activation of a latent transforming growth factor-beta 1-like molecule by plasmin during co-culture. Journal of Cell Biology, 109(1):309–??, July 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/109/1/
309.

Streeter:1988:IFC

[SRB88] P. R. Streeter, B. T. Rouse, and E. C. Butcher. Immunohistologic and functional characterization of a vascular addressin involved in lymphocyte homing into peripheral lymph nodes. Journal of Cell Biology, 107(5):1853–??, November 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/107/5/1853.

Schmidt:1987:ISD

[SRH^{+}87] G. Schmidt, H. Robenek, B. Harrach, J. Glössl, V. Nolte, H. Hörmann, H. Richter, and H. Kresse. Interaction of small dermatan sulfate proteoglycan from fibroblasts with fibronectin. Journal of Cell Biology, 104(6):1683–??, June 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/104/6/1683.

Skibbens:1989:DEI

[SRM89] J. E. Skibbens, M. G. Roth, and K. S. Matlin. Differential extractability of influenza virus hemagglutinin during intracellular transport in polarized epithelial cells and nonpolar fibroblasts. Journal of Cell Biology, 108(3):821–??, March 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/108/3/821.

Sager:1986:NMS

[SROB86] P. R. Sager, N. L. Rothfield, J. M. Oliver, and R. D. Berlin. A novel mitotic spindle pole component that originates from the cytoplasm during prophase. Journal of Cell Biology, 103(5): 1863–??, November 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/103/5/1863.

Stevens:1986:GDD

[SRPS86] T. H. Stevens, J. H. Rothman, G. S. Payne, and R. Schekman. Gene dosage-dependent secretion of yeast vacuolar carboxypeptidase Y. Journal of Cell Biology, 102(5):1551–??, May 1986. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/102/5/1551.

Sepsenwol:1989:UCA

[SRR89] S. Sepsenwol, H. Ris, and T. M. Roberts. A unique cytoskeleton associated with crawling in the amoeboid sperm of the nematode, Ascaris suum. Journal of Cell Biology, 108(1):55–??, January 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/108/1/
55.

Skalli:1986:MAA

[SRT^{+}86] O. Skalli, P. Ropraz, A. Trzeciak, G. Benzonana, D. Gillessen, and G. Gabbiani. A monoclonal antibody against alpha-smooth muscle actin: a new probe for smooth muscle differentiation. Journal of Cell Biology, 103(6):2787–??, December 1986. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/103/6/2787.

Sweet:1988:CSK

[SRW88] S. C. Sweet, C. M. Rogers, and M. J. Welsh. Calmodulin stabilization of kinetochore microtubule structure to the effect of nocodazole. Journal of Cell Biology, 107(6):2243–??, December 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/107/6/
2243.

Staatz:1989:MGI

[SRW^{+}89] W. D. Staatz, S. M. Rajpara, E. A. Wayner, W. G. Carter, and S. A. Santoro. The membrane glycoprotein Ia–IIa (VLA-2) complex mediates the Mg++-dependent adhesion of platelets to collagen. Journal of Cell Biology, 108(5):1917–??, May 1989. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/108/5/1917.

Sporn:1987:SRA

[SRWdC87] M. B. Sporn, A. B. Roberts, L. M. Wakefield, and B. de Crombrugghe. Some recent advances in the chemistry and biology of transforming growth factor-beta. Journal of Cell Biology, 105(3): 1039–??, September 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/3/1039.

Saitoh:1985:PDS

[SS85] T. Saitoh and J. H. Schwartz. Phosphorylation-dependent subcellular translocation of a Ca^{2+}/calmodulin-dependent protein kinase produces an autonomous enzyme in Aplysia neurons. Journal of Cell Biology, 100(3):835–??, March 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/100/3/835.

Schafer:1987:ISA

[SS87] D. A. Schafer and F. E. Stockdale. Identification of sarcolemmaassociated antigens with differential distributions on fast and slow skeletal muscle fibers. Journal of Cell Biology, 104(4): 967–??, January 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/104/4/967.

Seilheimer:1988:SAM

[SS88a] B. Seilheimer and M. Schachner. Studies of adhesion molecules mediating interactions between cells of peripheral nervous system indicate a major role for L1 in mediating sensory neuron growth on Schwann cells in culture. Journal of Cell Biology, 107(1):341–??, July 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/107/1/341.

Smith:1988:HHH

[SS88b] M. M. Smith and V. B. Stirling. Histone H3 and H4 gene deletions in Saccharomyces cerevisiae. Journal of Cell Biology, 106 (3):557–??, March 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/106/3/557.

Stoller:1988:RME

[SS88c] T. J. Stoller and D. Shields. Retrovirus-mediated expression of preprosomatostatin: posttranslational processing, intracellular storage, and secretion in GH3 pituitary cells. Journal of Cell Biology, 107(6):2087–??, December 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/107/6/2087.

Sanders:1989:CMM

[SS89a] M. A. Sanders and J. L. Salisbury. Centrin-mediated microtubule severing during flagellar excision in Chlamydomonas reinhardtii. Journal of Cell Biology, 108(5):1751–??, May 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/108/5/1751.

Satlin:1989:CRH

[SS89b] L. M. Satlin and G. J. Schwartz. Cellular remodeling of HCO3(-)-secreting cells in rabbit renal collecting duct in response to an acidic environment. Journal of Cell Biology, 109(3): 1279–??, September 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/3/1279.

Shyng:1989:DRA

[SS89c] S. L. Shyng and M. M. Salpeter. Degradation rate of acetylcholine receptors inserted into denervated vertebrate neuromuscular junctions. Journal of Cell Biology, 108(2):647–??, February 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/108/2/

Stoller:1989:PPM

[SS89d] T. J. Stoller and D. Shields. The propeptide of preprosomatostatin mediates intracellular transport and secretion of alphaglobin from mammalian cells. Journal of Cell Biology, 108 (5):1647–??, May 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/108/5/1647.

Sanes:1986:ESA

[SSC86] J. R. Sanes, M. Schachner, and J. Covault. Expression of several adhesive macromolecules (N-CAM, L1, J1, NILE, uvomorulin, laminin, fibronectin, and a heparan sulfate proteoglycan) in embryonic, adult, and denervated adult skeletal muscle. Journal of Cell Biology, 102(2):420–??, February 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/102/2/420.

Snow:1987:MAI

[SSG87] C. M. Snow, A. Senior, and L. Gerace. Monoclonal antibodies identify a group of nuclear pore complex glycoproteins. Journal of Cell Biology, 104(5):1143–??, May 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/104/5/1143.

Salisbury:1987:FRC

[SSH87] J. L. Salisbury, M. A. Sanders, and L. Harpst. Flagellar root contraction and nuclear movement during flagellar regeneration in Chlamydomonas reinhardtii. Journal of Cell Biology, 105(4): 1799–??, October 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/4/1799.

Sas:1985:JBL

[SSJ^{+}85] D. F. Sas, M. J. Sas, K. R. Johnson, A. S. Menko, and R. G. Johnson. Junctions between lens fiber cells are labeled with a monoclonal antibody shown to be specific for MP26. Journal of Cell Biology, 100(1):216–??, January 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/100/1/216.

Singer:1988:CSD

[SSK^{+}88] I. I. Singer, S. Scott, D. W. Kawka, D. M. Kazazis, J. Gailit, and E. Ruoslahti. Cell surface distribution of fibronectin and vitronectin receptors depends on substrate composition and extracellular matrix accumulation. Journal of Cell Biology, 106 (6):2171–??, June 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/106/6/2171.

Szczesna-Skorupa:1989:NTS

[SSK89] E. Szczesna-Skorupa and B. Kemper. NH2-terminal substitutions of basic amino acids induce translocation across the microsomal membrane and glycosylation of rabbit cytochrome P450IIC2. Journal of Cell Biology, 108(4):1237–??, April 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/108/4/1237.

Singer:1989:ASG

[SSKK89] I. I. Singer, S. Scott, D. W. Kawka, and D. M. Kazazis. Adhesomes: specific granules containing receptors for laminin, C3bi/ fibrinogen, fibronectin, and vitronectin in human polymorphonuclear leukocytes and monocytes. Journal of Cell Biology, 109(6): 3169–??, December 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/6/3169.

Stevenson:1986:IZH

[SSMG86] B. R. Stevenson, J. D. Siliciano, M. S. Mooseker, and D. A. Goodenough. Identification of ZO-1: a high molecular weight polypeptide associated with the tight junction (zonula occludens) in a variety of epithelia. Journal of Cell Biology, 103(3):755–??, September 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/103/3/
755.

Silver:1989:YPR

[SSO89] P. Silver, I. Sadler, and M. A. Osborne. Yeast proteins that recognize nuclear localization sequences. Journal of Cell Biology, 109 (3):983–??, September 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/3/983.

Sinard:1989:MAA

[SSP89] J. H. Sinard, W. F. Stafford, and T. D. Pollard. The mechanism of assembly of Acanthamoeba myosin–II minifilaments: minifilaments assemble by three successive dimerization steps. Journal of Cell Biology, 109(4):1537–??, October 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/109/4/1537.

Saadat:1989:CNF

[SSR89] S. Saadat, M. Sendtner, and H. Rohrer. Ciliary neurotrophic factor induces cholinergic differentiation of rat sympathetic neurons in culture. Journal of Cell Biology, 108(5):1807–??, May 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/108/5/
1807.

Schroer:1988:RKO

[SSRS88] T. A. Schroer, B. J. Schnapp, T. S. Reese, and M. P. Sheetz. The role of kinesin and other soluble factors in organelle movement along microtubules. Journal of Cell Biology, 107(5): 1785–??, November 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/107/5/1785.

Sellers:1985:LCP

[SSS85] J. R. Sellers, J. A. Spudich, and M. P. Sheetz. Light chain phosphorylation regulates the movement of smooth muscle myosin on actin filaments. Journal of Cell Biology, 101(5):1897–??, November 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/101/5/
1897.

Schneider:1987:SFM

[SSS^{+}87a] A. Schneider, T. Sherwin, R. Sasse, D. G. Russell, K. Gull, and T. Seebeck. Subpellicular and flagellar microtubules of Trypanosoma brucei brucei contain the same alpha-tubulin isoforms. Journal of Cell Biology, 104(3):431–??, March 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/104/3/431.

Sherwin:1987:DLC

[SSS^{+}87b] T. Sherwin, A. Schneider, R. Sasse, T. Seebeck, and K. Gull. Distinct localization and cell cycle dependence of COOH terminally tyrosinolated alpha-tubulin in the microtubules of Trypanosoma brucei brucei. Journal of Cell Biology, 104(3):439–??, March 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/104/3/

Steinberg:1988:PCS

[SSS88] T. H. Steinberg, J. A. Swanson, and S. C. Silverstein. A prelysosomal compartment sequesters membrane-impermeant fluorescent dyes from the cytoplasmic matrix of J774 macrophages. Journal of Cell Biology, 107(3):887–??, September 1988. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/107/3/887.

Small:1987:ICC

[SSSA87] S. J. Small, G. E. Shull, M. J. Santoni, and R. Akeson. Identification of a cDNA clone that contains the complete coding sequence for a 140-kD rat NCAM polypeptide. Journal of Cell Biology, 105 (5):2335–??, November 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/5/2335.

Simionescu:1985:DML

[SSSP85] M. Simionescu, N. Simionescu, F. Santoro, and G. E. Palade. Differentiated microdomains of the luminal plasmalemma of murine muscle capillaries: segmental variations in young and old animals. Journal of Cell Biology, 100(5):1396–??, May 1985. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/100/5/1396.

Sato:1988:MPB

[SSSP88] M. Sato, W. H. Schwartz, S. C. Selden, and T. D. Pollard. Mechanical properties of brain tubulin and microtubules. Journal of Cell Biology, 106(4):1205–??, April 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/106/4/1205.

Schneider:1985:LTA

[SSV^{+}85] M. D. Schneider, J. R. Sellers, M. Vahey, Y. A. Preston, and R. S. Adelstein. Localization and topography of antigenic domains within the heavy chain of smooth muscle myosin. Journal of Cell Biology, 101(1):66–??, July 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/101/1/66.

Smith:1985:ACC

[SSW^{+}85] H. C. Smith, D. L. Spector, C. L. Woodcock, R. L. Ochs, and J. Bhorjee. Alterations in chromatin conformation are accompanied by reorganization of nonchromatin domains that contain U-snRNP protein p28 and nuclear protein p107. Journal of Cell Biology, 101(2):560–??, August 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/101/2/560.

Schwarzbauer:1989:SSA

[SSW89] J. E. Schwarzbauer, C. S. Spencer, and C. L. Wilson. Selective secretion of alternatively spliced fibronectin variants. Journal of Cell Biology, 109(6):3445–??, December 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/109/6/3445.

Shaafi:1986:RIC

[SSY^{+}86] R. I. Sha’afi, J. Shefcyk, R. Yassin, T. F. Molski, M. Volpi, P. H. Naccache, J. R. White, M. B. Feinstein, and E. L. Becker. Is a rise in intracellular concentration of free calcium necessary or sufficient for stimulated cytoskeletal-associated actin? Journal of Cell Biology, 102(4):1459–??, April 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/102/4/1459.

Severina:1988:CME

[SSZ88] I. I. Severina, V. P. Skulachev, and D. B. Zorov. Coupling membranes as energy-transmitting cables. II. Cyanobacterial trichomes. Journal of Cell Biology, 107(2):497–??, August 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/107/2/497.

Saadat:1986:SIT

[ST86] S. Saadat and H. Thoenen. Selective induction of tyrosine hydroxylase by cell–cell contact in bovine adrenal chromaffin cells is mimicked by plasma membranes. Journal of Cell Biology, 103 (5):1991–??, November 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/103/5/1991.

Stephens:1985:ETC

[Ste85] R. E. Stephens. Evidence for a tubulin-containing lipid-protein structural complex in ciliary membranes. Journal of Cell Biology, 100(4):1082–??, April 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/100/4/1082.

Schittny:1988:HRI

[STE88] J. C. Schittny, R. Timpl, and J. Engel. High resolution immunoelectron microscopic localization of functional domains of laminin, nidogen, and heparan sulfate proteoglycan in epithelial basement membrane of mouse cornea reveals different topological orientations. Journal of Cell Biology, 107(4):1599–??, October 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/107/4/
1599.

Skinner:1985:CBS

[STF85] M. K. Skinner, P. S. Tung, and I. B. Fritz. Cooperativity between Sertoli cells and testicular peritubular cells in the production and deposition of extracellular matrix components. Journal of Cell Biology, 100(6):1941–??, June 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/100/6/1941.

Sperling:1987:CLP

[STGK87] L. Sperling, A. Tardieu, and T. Gulik-Krzywicki. The crystal lattice of Paramecium trichocysts before and after exocytosis by Xray diffraction and freeze-fracture electron microscopy. Journal of Cell Biology, 105(4):1649–??, October 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/105/4/1649.

Sandberg:1989:EMC

[STH^{+}89] M. Sandberg, M. Tamminen, H. Hirvonen, E. Vuorio, and T. Pihlajaniemi. Expression of mRNAs coding for the alpha 1 chain of type XIII collagen in human fetal tissues: comparison with expression of mRNAs for collagen types I, II, and III. Journal of Cell Biology, 109(3):1371–??, September 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/109/3/1371.

Strome:1986:FVD

[Str86] S. Strome. Fluorescence visualization of the distribution of microfilaments in gonads and early embryos of the nematode Caenorhabditis elegans. Journal of Cell Biology, 103(6):2241– ??, December 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/103/6/2241.

Shelly:1989:HDV

[STS^{+}89] L. L. Shelly, W. Tynan, W. Schmid, G. Schütz, and G. C. Yeoh. Hepatocyte differentiation in vitro: initiation of tyrosine aminotransferase expression in cultured fetal rat hepatocytes. Journal of Cell Biology, 109(6):3403–??, December 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/109/6/3403.

Sandberg:1987:LTI

[SV87] M. Sandberg and E. Vuorio. Localization of types I, II, and III collagen mRNAs in developing human skeletal tissues by in situ hybridization. Journal of Cell Biology, 104(4):1077–??, January 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/104/4/
1077.

Skene:1989:PMA

[SV89] J. H. Skene and I. Virág. Posttranslational membrane attachment and dynamic fatty acylation of a neuronal growth cone protein, GAP-43. Journal of Cell Biology, 108(2):613–??, February 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/108/2/
613.

Sage:1989:SSP

[SVF^{+}89] H. Sage, R. B. Vernon, S. E. Funk, E. A. Everitt, and J. Angello. SPARC, a secreted protein associated with cellular proliferation, inhibits cell spreading in vitro and exhibits Ca+2-dependent binding to the extracellular matrix. Journal of Cell Biology, 109(1):341–??, July 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/1/341.

Stoppelli:1986:IUP

[SVG^{+}86] M. P. Stoppelli, P. Verde, G. Grimaldi, E. K. Locatelli, and F. Blasi. Increase in urokinase plasminogen activator mRNA synthesis in human carcinoma cells is a primary effect of the potent tumor promoter, phorbol myristate acetate. Journal of Cell Biology, 102(4):1235–??, April 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/102/4/1235.

Salas:1988:SAS

[SVSH^{+}88] P. J. Salas, D. E. Vega-Salas, J. Hochman, E. Rodriguez-Boulan, and M. Edidin. Selective anchoring in the specific plasma membrane domain: a role in epithelial cell polarity. Journal of Cell Biology, 107(6):2363–??, December 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/107/6/2363.

Siegel:1985:EAP

[SW85] V. Siegel and P. Walter. Elongation arrest is not a prerequisite for secretory protein translocation across the microsomal membrane. Journal of Cell Biology, 100(6):1913–??, June 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/100/6/1913.

Swann:1986:PPI

[SW86] K. Swann and M. Whitaker. The part played by inositol trisphosphate and calcium in the propagation of the fertilization wave in sea urchin eggs. Journal of Cell Biology, 103(6):2333–??, December 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/103/6/
2333.

Shea:1987:MAB

[SW87a] D. K. Shea and C. J. Walsh. mRNAs for alpha- and betatubulin and flagellar calmodulin are among those coordinately regulated when Naegleria gruberi amebae differentiate into flagellates. Journal of Cell Biology, 105(3):1303–??, September 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/105/3/1303.

Stickel:1987:AAC

[SW87b] S. K. Stickel and Y. L. Wang. Alpha-actinin-containing aggregates in transformed cells are highly dynamic structures. Journal of Cell Biology, 104(6):1521–??, June 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/104/6/1521.

Stickel:1988:SPG

[SW88] S. K. Stickel and Y. L. Wang. Synthetic peptide GRGDS induces dissociation of alpha-actinin and vinculin from the sites of focal contacts. Journal of Cell Biology, 107(3):1231–??, September 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/107/3/
1231.

Stirzaker:1987:PRG

[SWC^{+}87] S. C. Stirzaker, P. L. Whitfeld, D. L. Christie, A. R. Bellamy, and G. W. Both. Processing of rotavirus glycoprotein VP7: implications for the retention of the protein in the endoplasmic reticulum. Journal of Cell Biology, 105(6):2897–??, December 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/105/6/
2897.

Stearns:1988:EBM

[SWTB88] M. E. Stearns, M. Wang, K. D. Tew, and L. I. Binder. Estramustine binds a MAP-1-like protein to inhibit microtubule assembly in vitro and disrupt microtubule organization in DU 145 cells. Journal of Cell Biology, 107(6):2647–??, December 1988. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/107/6/2647.

Schroder:1985:PBT

[SWW85] H. C. Schröder, J. Wehland, and K. Weber. Purification of brain tubulin-tyrosine ligase by biochemical and immunological methods. Journal of Cell Biology, 100(1):276–??, January 1985. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/100/1/276.

Slonczewski:1985:PAI

[SWZ85] J. L. Slonczewski, M. W. Wilde, and S. H. Zigmond. Phosphorylase a activity as an indicator of neutrophil activation by chemotactic peptide. Journal of Cell Biology, 101(4):1191–??, October 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/101/4/
1191.

Spiegel:1985:FGP

[SYH^{+}85] S. Spiegel, K. M. Yamada, B. E. Hom, J. Moss, and P. H. Fishman. Fluorescent gangliosides as probes for the retention and organization of fibronectin by ganglioside-deficient mouse cells. Journal of Cell Biology, 100(3):721–??, March 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/100/3/721.

Spiegel:1986:FOF

[SYH^{+}86] S. Spiegel, K. M. Yamada, B. E. Hom, J. Moss, and P. H. Fishman. Fibrillar organization of fibronectin is expressed coordinately with cell surface gangliosides in a variant murine fibroblast. Journal of Cell Biology, 102(5):1898–??, May 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/102/5/1898.

Swanson:1985:PEH

[SYS85] J. A. Swanson, B. D. Yirinec, and S. C. Silverstein. Phorbol esters and horseradish peroxidase stimulate pinocytosis and redirect the flow of pinocytosed fluid in macrophages. Journal of Cell Biology, 100(3):851–??, March 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/100/3/851.

Sam-Yellowe:1988:SPF

[SYSP88] T. Y. Sam-Yellowe, H. Shio, and M. E. Perkins. Secretion of Plasmodium falciparum rhoptry protein into the plasma membrane of host erythrocytes. Journal of Cell Biology, 106(5):1507–??, May 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/106/5/
1507.

Shimoyama:1989:MCH

[SYT^{+}89] Y. Shimoyama, T. Yoshida, M. Terada, Y. Shimosato, O. Abe, and S. Hirohashi. Molecular cloning of a human Ca^{2+}-dependent cell–cell adhesion molecule homologous to mouse placental cadherin: its low expression in human placental tissues. Journal of Cell Biology, 109(4):1787–??, October 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/109/4/1787.

Trapp:1989:MAG

[TACQ89] B. D. Trapp, S. B. Andrews, C. Cootauco, and R. Quarles. The myelin-associated glycoprotein is enriched in multivesicular bodies and periaxonal membranes of actively myelinating oligodendrocytes. Journal of Cell Biology, 109(5):2417–??, November 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/109/5/

Tranqui:1989:DSR

[TAHC^{+}89] L. Tranqui, A. Andrieux, G. Hudry-Clergeon, J. J. Ryckewaert, S. Soyez, A. Chapel, M. H. Ginsberg, E. F. Plow, and G. Marguerie. Differential structural requirements for fibrinogen binding to platelets and to endothelial cells. Journal of Cell Biology, 108(6):2519–??, June 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/108/6/2519.

Townes-Anderson:1985:RCD

[TAMR85] E. Townes-Anderson, P. R. MacLeish, and E. Raviola. Rod cells dissociated from mature salamander retina: ultrastructure and uptake of horseradish peroxidase. Journal of Cell Biology, 100 (1):175–??, January 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/100/1/175.

Teeter:1986:DAM

[TASK86] L. D. Teeter, S. Atsumi, S. Sen, and T. Kuo. DNA amplification in multidrug, cross-resistant Chinese hamster ovary cells: molecular characterization and cytogenetic localization of the amplified DNA. Journal of Cell Biology, 103(4):1159–??, October 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/103/4/
1159.

Titus:1985:IGP

[TB85a] D. E. Titus and W. M. Becker. Investigation of the glyoxysomeperoxisome transition in germinating cucumber cotyledons using double-label immunoelectron microscopy. Journal of Cell Biology, 101(4):1288–??, October 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/101/4/1288.

Torrisi:1985:ISP

[TB85b] M. R. Torrisi and S. Bonatti. Immunocytochemical study of the partition and distribution of Sindbis virus glycoproteins in freezefractured membranes of infected baby hamster kidney cells. Journal of Cell Biology, 101(4):1300–??, October 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/101/4/1300.

Tooze:1987:AAS

[TB87] J. Tooze and B. Burke. Accumulation of adrenocorticotropin secretory granules in the midbody of telophase AtT20 cells: evidence that secretory granules move anterogradely along microtubules. Journal of Cell Biology, 104(4):1047–??, January 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/104/4/1047.

Toschi:1988:CCP

[TB88a] L. Toschi and R. Bravo. Changes in cyclin/proliferating cell nuclear antigen distribution during DNA repair synthesis. Journal of Cell Biology, 107(5):1623–??, November 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/107/5/1623.

Troutt:1988:UMP

[TB88b] L. L. Troutt and B. Burnside. The unusual microtubule polarity in teleost retinal pigment epithelial cells. Journal of Cell Biology, 107(4):1461–??, October 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/107/4/1461.

Taylor:1989:DFM

[TB89a] L. D. Taylor and E. Bandman. Distribution of fast myosin heavy chain isoforms in thick filaments of developing chicken pectoral muscle. Journal of Cell Biology, 108(2):533–??, February 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/108/2/533.

Tombes:1989:IFC

[TB89b] R. M. Tombes and G. G. Borisy. Intracellular free calcium and mitosis in mammalian cells: anaphase onset is calcium modulated, but is not triggered by a brief transient. Journal of Cell Biology, 109(2):627–??, August 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/109/2/627.

Tedeschi:1986:API

[TBK86] B. Tedeschi, J. N. Barrett, and R. W. Keane. Astrocytes produce interferon that enhances the expression of H-2 antigens on a subpopulation of brain cells. Journal of Cell Biology, 102(6):2244–??, June 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/102/6/
2244.

Tsilibary:1986:RMN

[TC86] E. C. Tsilibary and A. S. Charonis. The role of the main noncollagenous domain (NC1) in type IV collagen self-assembly. Journal of Cell Biology, 103(6):2467–??, December 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/103/6/2467.

Tague:1987:PVP

[TC87] B. W. Tague and M. J. Chrispeels. The plant vacuolar protein, phytohemagglutinin, is transported to the vacuole of transgenic yeast. Journal of Cell Biology, 105(5):1971–??, November 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/105/5/
1971.

Thuret-Carnahan:1985:ETS

[TCBF^{+}85] J. Thuret-Carnahan, J. L. Bossu, A. Feltz, K. Langley, and D. Aunis. Effect of taxol on secretory cells: functional, morphological, and electrophysiological correlates. Journal of Cell Biology, 100(6):1863–??, June 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/100/6/1863.

Terasaki:1986:MER

[TCF86] M. Terasaki, L. B. Chen, and K. Fujiwara. Microtubules and the endoplasmic reticulum are highly interdependent structures. Journal of Cell Biology, 103(4):1557–??, October 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/103/4/1557.

Triller:1987:GAA

[TCK87] A. Triller, F. Cluzeaud, and H. Korn. gamma-aminobutyric acidcontaining terminals can be apposed to glycine receptors at central synapses. Journal of Cell Biology, 104(4):947–??, January 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/104/4/
947.

Triller:1985:DGR

[TCP^{+}85] A. Triller, F. Cluzeaud, F. Pfeiffer, H. Betz, and H. Korn. Distribution of glycine receptors at central synapses: an immunoelectron microscopy study. Journal of Cell Biology, 101 (2):683–??, August 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/101/2/683.

Thiery:1985:IAR

[TDGE85] J. P. Thiery, A. Delouvée, M. Grumet, and G. M. Edelman. Initial appearance and regional distribution of the neuron-glia cell adhesion molecule in the chick embryo. Journal of Cell Biology, 100(2):442–??, February 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/100/2/442.

Terranova:1985:HEC

[TDL^{+}85] V. P. Terranova, R. DiFlorio, R. M. Lyall, S. Hic, R. Friesel, and T. Maciag. Human endothelial cells are chemotactic to endothelial cell growth factor and heparin. Journal of Cell Biology, 101 (6):2330–??, December 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/101/6/2330.

Tomaselli:1987:INC

[TDR87a] K. J. Tomaselli, C. H. Damsky, and L. F. Reichardt. Interactions of a neuronal cell line (PC12) with laminin, collagen IV, and fibronectin: identification of integrin-related glycoproteins involved in attachment and process outgrowth. Journal of Cell Biology, 105(5):2347–??, November 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/105/5/2347.

Traas:1987:ANP

[TDR^{+}87b] J. A. Traas, J. H. Doonan, D. J. Rawlins, P. J. Shaw, J. Watts, and C. W. Lloyd. An actin network is present in the cytoplasm throughout the cell cycle of carrot cells and associates with the dividing nucleus. Journal of Cell Biology, 105(1):387–??, July 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/105/1/
387.

Tomaselli:1988:PCM

[TDR88] K. J. Tomaselli, C. H. Damsky, and L. F. Reichardt. Purification and characterization of mammalian integrins expressed by a rat neuronal cell line (PC12): evidence that they function as alpha/ beta heterodimeric receptors for laminin and type IV collagen. Journal of Cell Biology, 107(3):1241–??, September 1988. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/107/3/1241.

Trimmer:1987:CMA

[TES^{+}87] J. S. Trimmer, Y. Ebina, R. W. Schackmann, C. G. Meinhof, and V. D. Vacquier. Characterization of a monoclonal antibody that induces the acrosome reaction of sea urchin sperm. Journal of Cell Biology, 105(3):1121–??, September 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/105/3/1121.

Tyner:1986:EPR

[TF86] A. L. Tyner and E. Fuchs. Evidence for posttranscriptional regulation of the keratins expressed during hyperproliferation and malignant transformation in human epidermis. Journal of Cell Biology, 103(5):1945–??, November 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/103/5/1945.

Tank:1985:EFI

[TFBW85] D. W. Tank, W. J. Fredericks, L. S. Barak, and W. W. Webb. Electric field-induced redistribution and postfield relaxation of low density lipoprotein receptors on cultured human fibroblasts. Journal of Cell Biology, 101(1):148–??, July 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/101/1/148.

Tilney:1987:MAF

[TFD87] L. G. Tilney, Y. Fukui, and D. J. DeRosier. Movement of the actin filament bundle in Mytilus sperm: a new mechanism is proposed. Journal of Cell Biology, 104(4):981–??, January 1987. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/104/4/981.

Thompson:1989:ETG

[TFS^{+}89] N. L. Thompson, K. C. Flanders, J. M. Smith, L. R. Ellingsworth, A. B. Roberts, and M. B. Sporn. Expression of transforming growth factor-beta 1 in specific cells and tissues of adult and neonatal mice. Journal of Cell Biology, 108(2): 661–??, February 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/108/2/661.

Taubman:1987:CCM

[TGNG87] M. B. Taubman, J. W. Grant, and B. Nadal-Ginard. Cloning and characterization of mammalian myosin regulatory light chain (RLC) cDNA: the RLC gene is expressed in smooth, sarcomeric, and nonmuscle tissues. Journal of Cell Biology, 104(6):1505–??, June 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/104/6/
1505.

Takada:1989:PSV

[TH89a] Y. Takada and M. E. Hemler. The primary structure of the VLA-2/collagen receptor alpha 2 subunit (platelet GPIa): homology to other integrins and the presence of a possible collagen-binding domain. Journal of Cell Biology, 109(1):397–??, July 1989. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/109/1/397.

Trybus:1989:MAD

[TH89b] K. M. Trybus and L. Henry. Monoclonal antibodies detect and stabilize conformational states of smooth muscle myosin. Journal of Cell Biology, 109(6):2879–??, December 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/109/6/2879.

Turkewitz:1989:CTR

[TH89c] A. P. Turkewitz and S. C. Harrison. Concentration of transferrin receptor in human placental coated vesicles. Journal of Cell Biology, 108(6):2127–??, June 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/108/6/2127.

Tooze:1987:ASE

[THFB87] J. Tooze, M. Hollinshead, R. Frank, and B. Burke. An antibody specific for an endoproteolytic cleavage site provides evidence that pro-opiomelanocortin is packaged into secretory granules in AtT20 cells before its cleavage. Journal of Cell Biology, 105(1):155–??, July 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/1/155.

Tash:1986:APC

[THM86] J. S. Tash, H. Hidaka, and A. R. Means. Axokinin phosphorylation by cAMP-dependent protein kinase is sufficient for activation of sperm flagellar motility. Journal of Cell Biology, 103 (2):649–??, August 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/103/2/649.

Tsukita:1989:NKB

[THT89] S. Tsukita, Y. Hieda, and S. Tsukita. A new 82-kD barbed endcapping protein (radixin) localized in the cell-to-cell adherens junction: purification and characterization. Journal of Cell Biology, 108(6):2369–??, June 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/108/6/2369.

Tilney:1985:ART

[TI85] L. G. Tilney and S. Inoué. Acrosomal reaction of the Thyone sperm. III. The relationship between actin assembly and water influx during the extension of the acrosomal process. Journal of Cell Biology, 100(4):1273–??, April 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/100/4/1273.

Tilney:1987:FGM

[TI87] L. G. Tilney and S. Inoué. Flagellar gyration and midpiece rotation during extension of the acrosomal process of Thyone sperm: how and why this occurs. Journal of Cell Biology, 104 (3):407–??, March 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/104/3/407.

Takeda:1986:CAM

[TIS^{+}86] M. Takeda, H. Iwata, S. Suzuki, K. S. Brown, and K. Kimata. Correction of abnormal matrix formed by cmd/cmd chondrocytes in culture by exogenously added cartilage proteoglycan. Journal of Cell Biology, 103(4):1605–??, October 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/103/4/1605.

Tsukita:1989:NKP

[TIT89] S. Tsukita, M. Itoh, and S. Tsukita. A new 400-kD protein from isolated adherens junctions: its localization at the undercoat of adherens junctions and at microfilament bundles such as stress fibers and circumferential bundles. Journal of Cell Biology, 109 (6):2905–??, December 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/6/2905.

Taniuchi:1985:CBP

[TJ85] M. Taniuchi and E. M. Johnson. Characterization of the binding properties and retrograde axonal transport of a monoclonal antibody directed against the rat nerve growth factor receptor. Journal of Cell Biology, 101(3):1100–??, September 1985. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/101/3/1100.

Turner:1986:RCV

[TJF86] P. R. Turner, L. A. Jaffe, and A. Fein. Regulation of cortical vesicle exocytosis in sea urchin eggs by inositol 1,4,5-trisphosphate and GTP-binding protein. Journal of Cell Biology, 102(1):70–??, January 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
102/1/70.

Tooze:1989:CSE

[TKFH89] J. Tooze, H. F. Kern, S. D. Fuller, and K. E. Howell. Condensation-sorting events in the rough endoplasmic reticulum of exocrine pancreatic cells. Journal of Cell Biology, 109(1):35– ??, July 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
109/1/35.

Tavassoli:1986:LEM

[TKK86] M. Tavassoli, T. Kishimoto, and M. Kataoka. Liver endothelium mediates the hepatocyte’s uptake of ceruloplasmin. Journal of Cell Biology, 102(4):1298–??, April 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/102/4/1298.

Takei:1987:FEI

[TKK^{+}87] F. Takei, Y. Kikuchi, A. Kikuchi, S. Mizuno, and K. Shimura. Further evidence for importance of the subunit combination of silk fibroin in its efficient secretion from the posterior silk gland cells. Journal of Cell Biology, 105(1):175–??, July 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/105/1/175.

Tash:1988:ICF

[TKP^{+}88] J. S. Tash, M. Krinks, J. Patel, R. L. Means, C. B. Klee, and A. R. Means. Identification, characterization, and functional correlation of calmodulin-dependent protein phosphatase in sperm. Journal of Cell Biology, 106(5):1625–??, May 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/106/5/1625.

Trybus:1987:ASM

[TL87a] K. M. Trybus and S. Lowey. Assembly of smooth muscle myosin minifilaments: effects of phosphorylation and nucleotide binding. Journal of Cell Biology, 105(6):3007–??, December 1987. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/105/6/3007.

Trybus:1987:SEB

[TL87b] K. M. Trybus and S. Lowey. Subunit exchange between smooth muscle myosin filaments. Journal of Cell Biology, 105(6): 3021–??, December 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/6/3021.

Tanner:1989:LTR

[TL89] L. I. Tanner and G. E. Lienhard. Localization of transferrin receptors and insulin-like growth factor II receptors in vesicles from 3T3-L1 adipocytes that contain intracellular glucose transporters. Journal of Cell Biology, 108(4):1537–??, April 1989. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/108/4/1537.

Trueb:1985:TMG

[TLC85] B. Trüeb, J. B. Lewis, and W. G. Carter. Translatable mRNA for GP140 (a subunit of type VI collagen) is absent in SV40 transformed fibroblasts. Journal of Cell Biology, 100(2):638–??, February 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
100/2/638.

Traub:1989:CCK

[TLD^{+}89] O. Traub, J. Look, R. Dermietzel, F. Brümmer, D. Hülser, and K. Willecke. Comparative characterization of the 21-kD and 26kD gap junction proteins in murine liver and cultured hepatocytes. Journal of Cell Biology, 108(3):1039–??, March 1989. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/108/3/1039.

Torrisi:1987:FDI

[TLP^{+}87] M. R. Torrisi, L. V. Lotti, A. Pavan, G. Migliaccio, and S. Bonatti. Free diffusion to and from the inner nuclear membrane of newly synthesized plasma membrane glycoproteins. Journal of Cell Biology, 104(3):733–??, March 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/104/3/733.

Tougard:1985:AAL

[TLPTV85] C. Tougard, D. Louvard, R. Picart, and A. Tixier-Vidal. Antibodies against a lysosomal membrane antigen recognize a prelysosomal compartment involved in the endocytic pathway in cultured prolactin cells. Journal of Cell Biology, 100(3):786–??, March 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
100/3/786.

Tajima:1986:SRP

[TLRW86] S. Tajima, L. Lauffer, V. L. Rath, and P. Walter. The signal recognition particle receptor is a complex that contains two distinct polypeptide chains. Journal of Cell Biology, 103(4): 1167–??, October 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/103/4/1167.

Tranquillo:1988:SML

[TLZ88] R. T. Tranquillo, D. A. Lauffenburger, and S. H. Zigmond. A stochastic model for leukocyte random motility and chemotaxis based on receptor binding fluctuations. Journal of Cell Biology, 106(2):303–??, February 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/106/2/303.

Tokuyasu:1987:ISCa

[TM87a] K. T. Tokuyasu and P. A. Maher. Immunocytochemical studies of cardiac myofibrillogenesis in early chick embryos. I. Presence of immunofluorescent titin spots in premyofibril stages. Journal of Cell Biology, 105(6):2781–??, December 1987. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/105/6/2781.

Tokuyasu:1987:ISCb

[TM87b] K. T. Tokuyasu and P. A. Maher. Immunocytochemical studies of cardiac myofibrillogenesis in early chick embryos. II. Generation of alpha-actinin dots within titin spots at the time of the first myofibril formation. Journal of Cell Biology, 105(6): 2795–??, December 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/6/2795.

Tassin:1985:FMO

[TMB85] A. M. Tassin, B. Maro, and M. Bornens. Fate of microtubuleorganizing centers during myogenesis in vitro. Journal of Cell Biology, 100(1):35–??, January 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/100/1/35.

Tucker:1985:SMD

[TMH^{+}85] J. B. Tucker, S. A. Mathews, K. A. Hendry, J. B. Mackie, and D. L. Roche. Spindle microtubule differentiation and deployment during micronuclear mitosis in Paramecium. Journal of Cell Biology, 101(5):1966–??, November 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/101/5/1966.

Toffenetti:1987:ICG

[TMP87] J. Toffenetti, D. Mischke, and M. L. Pardue. Isolation and characterization of the gene for myosin light chain two of Drosophila melanogaster. Journal of Cell Biology, 104(1):19–??, January 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/104/1/
19.

Tokuyasu:1985:DVD

[TMS85] K. T. Tokuyasu, P. A. Maher, and S. J. Singer. Distributions of vimentin and desmin in developing chick myotubes in vivo. II. Immunoelectron microscopic study. Journal of Cell Biology, 100(4):1157–??, April 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/100/4/1157.

Therrien:1989:GNI

[TN89] S. Therrien and P. H. Naccache. Guanine nucleotide-induced polymerization of actin in electropermeabilized human neutrophils. Journal of Cell Biology, 109(3):1125–??, September 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/109/3/
1125.

Trahair:1989:UTM

[TNG89] J. F. Trahair, M. R. Neutra, and J. I. Gordon. Use of transgenic mice to study the routing of secretory proteins in intestinal epithelial cells: analysis of human growth hormone compartmentalization as a function of cell type and differentiation. Journal of Cell Biology, 109(6):3231–??, December 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/109/6/3231.

Tidball:1986:TMJ

[TOB86] J. G. Tidball, T. O’Halloran, and K. Burridge. Talin at myotendinous junctions. Journal of Cell Biology, 103(4):1465–??, October 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/103/4/
1465.

Tokuyasu:1989:ISC

[Tok89] K. T. Tokuyasu. Immunocytochemical studies of cardiac myofibrillogenesis in early chick embryos. III. Generation of fasciae adherentes and costameres. Journal of Cell Biology, 108 (1):43–??, January 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/108/1/43.

Tooze:1985:BCP

[Too85] J. Tooze. Blocked coated pits in AtT20 cells result from endocytosis of budding retrovirions. Journal of Cell Biology, 101 (5):1713–??, November 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/101/5/1713.

Takano-Ohmuro:1985:ECS

[TOOK^{+}85] H. Takano-Ohmuro, T. Obinata, M. Kawashima, T. Masaki, and T. Tanaka. Embryonic chicken skeletal, cardiac, and smooth muscles express a common embryo-specific myosin light chain.

Journal of Cell Biology, 100(6):2025–??, June 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/100/6/2025.

Toyoshima:1987:CDTa

[Toy87a] Y. Y. Toyoshima. Chymotryptic digestion of Tetrahymena 22S dynein. I. Decomposition of three-headed 22S dynein to one- and two-headed particles. Journal of Cell Biology, 105(2):887–??, August 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/105/2/
887.

Toyoshima:1987:CDTb

[Toy87b] Y. Y. Toyoshima. Chymotryptic digestion of Tetrahymena ciliary dynein. II. Pathway of the degradation of 22S dynein heavy chains. Journal of Cell Biology, 105(2):897–??, August 1987. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/105/2/897.

Tilney:1989:AFG

[TP89] L. G. Tilney and D. A. Portnoy. Actin filaments and the growth, movement, and spread of the intracellular bacterial parasite, Listeria monocytogenes. Journal of Cell Biology, 109(4):1597–??, October 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
109/4/1597.

Tassin:1985:GAR

[TPBB85] A. M. Tassin, M. Paintrand, E. G. Berger, and M. Bornens. The Golgi apparatus remains associated with microtubule organizing centers during myogenesis. Journal of Cell Biology, 101 (2):630–??, August 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/101/2/630.

Tacchetti:1987:VMC

[TQN^{+}87] C. Tacchetti, R. Quarto, L. Nitsch, D. J. Hartmann, and R. Cancedda. In vitro morphogenesis of chick embryo hypertrophic cartilage. Journal of Cell Biology, 105(2):999–??, August 1987. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/105/2/999.

Trapp:1988:DMA

[Tra88] B. D. Trapp. Distribution of the myelin-associated glycoprotein and P0 protein during myelin compaction in quaking mouse peripheral nerve. Journal of Cell Biology, 107(2):675–??, August 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/107/2/
675.

Tomaselli:1986:DMI

[TRB86] K. J. Tomaselli, L. F. Reichardt, and J. L. Bixby. Distinct molecular interactions mediate neuronal process outgrowth on non-neuronal cell surfaces and extracellular matrices. Journal of Cell Biology, 103(6):2659–??, December 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/103/6/2659.

Trugnan:1987:PPS

[TRC^{+}87] G. Trugnan, M. Rousset, I. Chantret, A. Barbat, and A. Zweibaum. The posttranslational processing of sucraseisomaltase in HT-29 cells is a function of their state of enterocytic differentiation. Journal of Cell Biology, 104(5):1199–??, May 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/104/5/
1199.

Taylor:1989:TDIa

[TRCR89a] K. A. Taylor, M. C. Reedy, L. Córdova, and M. K. Reedy. Three-dimensional image reconstruction of insect flight muscle. I. The rigor myac layer. Journal of Cell Biology, 109(3): 1085–??, September 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/3/1085.

Taylor:1989:TDIb

[TRCR89b] K. A. Taylor, M. C. Reedy, L. Córdova, and M. K. Reedy. Three-dimensional image reconstruction of insect flight muscle. II. The rigor actin layer. Journal of Cell Biology, 109(3): 1103–??, September 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/3/1103.

Taraboletti:1987:TIT

[TRL87] G. Taraboletti, D. D. Roberts, and L. A. Liotta. Thrombospondininduced tumor cell migration: haptotaxis and chemotaxis are mediated by different molecular domains. Journal of Cell Biology, 105(5):2409–??, November 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/105/5/2409.

Talamas-Rohana:1988:IBP

[TRM88] P. Talamás-Rohana and I. Meza. Interaction between pathogenic amebas and fibronectin: substrate degradation and changes in cytoskeleton organization. Journal of Cell Biology, 106(5):1787– ??, May 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
106/5/1787.

Terracio:1988:IPD

[TRT^{+}88] L. Terracio, L. Rönnstrand, A. Tingström, K. Rubin, L. Claesson-Welsh, K. Funa, and C. H. Heldin. Induction of platelet-derived growth factor receptor expression in smooth muscle cells and fibroblasts upon tissue culturing. Journal of Cell Biology, 107 (5):1947–??, November 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/107/5/1947.

Trybus:1989:FSM

[Try89] K. M. Trybus. Filamentous smooth muscle myosin is regulated by phosphorylation. Journal of Cell Biology, 109(6):2887–??, December 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
109/6/2887.

Tiercy:1986:ECS

[TS86] J. M. Tiercy and E. M. Shooter. Early changes in the synthesis of nuclear and cytoplasmic proteins are induced by nerve growth factor in differentiating rat PC12 cells. Journal of Cell Biology, 103(6):2367–??, December 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/103/6/2367.

Takata:1988:PMP

[TS88] K. Takata and S. J. Singer. Phosphotyrosine-modified proteins are concentrated at the membranes of epithelial and endothelial cells during tissue development in chick embryos. Journal of Cell Biology, 106(5):1757–??, May 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/106/5/1757.

Tsay:1989:SMD

[TS89] H. J. Tsay and J. Schmidt. Skeletal muscle denervation activates acetylcholine receptor genes. Journal of Cell Biology, 108 (4):1523–??, April 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/108/4/1523.

Takemura:1986:RRC

[TSBW86] R. Takemura, P. E. Stenberg, D. F. Bainton, and Z. Werb. Rapid redistribution of clathrin onto macrophage plasma membranes in response to Fc receptor-ligand interaction during frustrated phagocytosis. Journal of Cell Biology, 102(1):55–??, January 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/102/1/
55.

Taubman:1989:ESM

[TSI^{+}89] M. B. Taubman, C. W. Smith, S. Izumo, J. W. Grant, T. Endo, A. Andreadis, and B. Nadal-Ginard. The expression of sarcomeric muscle-specific contractile protein genes in BC3H1 cells: BC3H1 cells resemble skeletal myoblasts that are defective for commitment to terminal differentiation. Journal of Cell Biology, 108(5):1799–??, May 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/108/5/1799.

Tsui:1988:MBG

[TSK88] H. C. Tsui, D. Schubert, and W. L. Klein. Molecular basis of growth cone adhesion: anchoring of adheron-containing filaments at adhesive loci. Journal of Cell Biology, 106(6):2095–??, June 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/106/6/
2095.

Tufaro:1987:ICM

[TSM87] F. Tufaro, M. D. Snider, and S. L. McKnight. Identification and characterization of a mouse cell mutant defective in the intracellular transport of glycoproteins. Journal of Cell Biology, 105(2):647–??, August 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/2/647.

Tulsiani:1989:NAD

[TSOC89] D. R. Tulsiani, M. D. Skudlarek, and M. C. Orgebin-Crist. Novel alpha–D–mannosidase of rat sperm plasma membranes: characterization and potential role in sperm-egg interactions. Journal of Cell Biology, 109(3):1257–??, September 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/109/3/1257.

Tranter:1989:EDN

[TSS89] M. P. Tranter, S. P. Sugrue, and M. A. Schwartz. Evidence for a direct, nucleotide-sensitive interaction between actin and liver cell membranes. Journal of Cell Biology, 109(6):2833–??, December 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/109/6/
2833.

Tsukita:1985:DCB

[TT85] S. Tsukita and S. Tsukita. Desmocalmin: a calmodulin-binding high molecular weight protein isolated from desmosomes. Journal of Cell Biology, 101(6):2070–??, December 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/101/6/2070.

Thomas:1986:TCM

[TT86a] G. Thomas and G. Thomas. Translational control of mRNA expression during the early mitogenic response in Swiss mouse 3T3 cells: identification of specific proteins. Journal of Cell Biology, 103(6):2137–??, December 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/103/6/2137.

Tooze:1986:CCV

[TT86b] J. Tooze and S. A. Tooze. Clathrin-coated vesicular transport of secretory proteins during the formation of ACTH-containing secretory granules in AtT20 cells. Journal of Cell Biology, 103 (3):839–??, September 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/103/3/839.

Tilney:1988:AFC

[TT88a] L. G. Tilney and M. S. Tilney. The actin filament content of hair cells of the bird cochlea is nearly constant even though the length, width, and number of stereocilia vary depending on the hair cell location. Journal of Cell Biology, 107(6):2563–??, December 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/107/6/
2563.

Toyama:1988:FAB

[TT88b] S. Toyama and S. Toyama. Functional alterations in beta’-actin from a KB cell mutant resistant to cytochalasin B. Journal of Cell Biology, 107(4):1499–??, October 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/107/4/1499.

Tsukita:1989:ICC

[TT89a] S. Tsukita and S. Tsukita. Isolation of cell-to-cell adherens junctions from rat liver. Journal of Cell Biology, 108(1):31–??, January 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/108/1/
31.

Turner:1989:RGC

[TT89b] J. R. Turner and A. M. Tartakoff. The response of the Golgi complex to microtubule alterations: the roles of metabolic energy and membrane traffic in Golgi complex organization. Journal of Cell Biology, 109(5):2081–??, November 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/109/5/2081.

Toda:1987:FRH

[TTBG87] K. Toda, T. L. Tuan, P. J. Brown, and F. Grinnell. Fibronectin receptors of human keratinocytes and their expression during cell culture. Journal of Cell Biology, 105(6):3097–??, December 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/105/6/
3097.

Tilney:1988:AFS

[TTC88] L. G. Tilney, M. S. Tilney, and D. A. Cotanche. Actin filaments, stereocilia, and hair cells of the bird cochlea. V. How the staircase pattern of stereociliary lengths is generated. Journal of Cell Biology, 106(2):355–??, February 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/106/2/355.

Tooze:1987:SPC

[TTF87] J. Tooze, S. A. Tooze, and S. D. Fuller. Sorting of progeny coronavirus from condensed secretory proteins at the exit from the trans-Golgi network of AtT20 cells. Journal of Cell Biology, 105 (3):1215–??, September 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/3/1215.

Torri-Tarelli:1985:TCB

[TTGFC85] F. Torri-Tarelli, F. Grohovaz, R. Fesce, and B. Ceccarelli. Temporal coincidence between synaptic vesicle fusion and quantal secretion of acetylcholine. Journal of Cell Biology, 101(4):1386–??, October 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
101/4/1386.

Tsukita:1986:SCS

[TTKM86] S. Tsukita, S. Tsukita, T. Kobayashi, and G. Matsumoto. Subaxolemmal cytoskeleton in squid giant axon. II. Morphological identification of microtubule- and microfilament-associated domains of axolemma. Journal of Cell Biology, 102(5):1710–??, May 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/102/5/
1710.

Torti:1989:MAD

[TTLR89] F. M. Torti, S. V. Torti, J. W. Larrick, and G. M. Ringold. Modulation of adipocyte differentiation by tumor necrosis factor and transforming growth factor beta. Journal of Cell Biology, 108(3):1105–??, March 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/108/3/1105.

Tsukita:1988:LIS

[TTM88] S. Tsukita, S. Tsukita, and G. Matsumoto. Light-induced structural changes of cytoskeleton in squid photoreceptor microvilli detected by rapid-freeze method. Journal of Cell Biology, 106 (4):1151–??, April 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/106/4/1151.

Tilney:1989:PBC

[TTS^{+}89] M. S. Tilney, L. G. Tilney, R. E. Stephens, C. Merte, D. Drenckhahn, D. A. Cotanche, and A. Bretscher. Preliminary biochemical characterization of the stereocilia and cuticular plate of hair cells of the chick cochlea. Journal of Cell Biology, 109 (4):1711–??, October 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/4/1711.

Turner:1985:MAP

[TTSSH85] R. J. Turner, J. Thompson, S. Sariban-Sohraby, and J. S. Handler. Monoclonal antibodies as probes of epithelial membrane polarization. Journal of Cell Biology, 101(6):2173–??, December 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/101/6/
2173.

Tooze:1988:SAA

[TTW88] S. A. Tooze, J. Tooze, and G. Warren. Site of addition of Nacetyl-galactosamine to the E1 glycoprotein of mouse hepatitis virus–A59. Journal of Cell Biology, 106(5):1475–??, May 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/106/5/1475.

Tao:1988:LTM

[TWB88] W. Tao, R. J. Walter, and M. W. Berns. Laser-transected microtubules exhibit individuality of regrowth, however most free new ends of the microtubules are stable. Journal of Cell Biology, 107 (3):1025–??, September 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/107/3/1025.

Tennekoon:1987:TNR

[TYP^{+}87] G. I. Tennekoon, J. Yoshino, K. W. Peden, J. Bigbee, J. L. Rutkowski, Y. Kishimoto, G. H. DeVries, and G. M. McKhann. Transfection of neonatal rat Schwann cells with SV-40 large T antigen gene under control of the metallothionein promoter. Journal of Cell Biology, 105(5):2315–??, November 1987. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/105/5/2315.

Usdin:1986:PCP

[UF86] T. B. Usdin and G. D. Fischbach. Purification and characterization of a polypeptide from chick brain that promotes the accumulation of acetylcholine receptors in chick myotubes. Journal of Cell Biology, 103(2):493–??, August 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/103/2/493.

Uyeda:1989:EAI

[UF89] T. Q. Uyeda and M. Furuya. Evidence for active interactions between microfilaments and microtubules in myxomycete flagellates. Journal of Cell Biology, 108(5):1727–??, May 1989. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/108/5/1727.

Ueno:1986:IPC

[UK86] T. Ueno and E. D. Korn. Isolation and partial characterization of a 110-kD dimer actin-binding protein. Journal of Cell Biology, 103(2):621–??, August 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/103/2/621.

Uno:1985:OPA

[UMHI85] I. Uno, K. Matsumoto, A. Hirata, and T. Ishikawa. Outer plaque assembly and spore encapsulation are defective during sporulation of adenylate cyclase-deficient mutants of Saccharomyces cerevisiae. Journal of Cell Biology, 100(6):1854–??, June 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/100/6/1854.

Uster:1986:RET

[UP86] P. S. Uster and R. E. Pagano. Resonance energy transfer microscopy: observations of membrane-bound fluorescent probes in model membranes and in living cells. Journal of Cell Biology, 103 (4):1221–??, October 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/103/4/1221.

Urban:1987:CAS

[UPL^{+}87] J. Urban, K. Parczyk, A. Leutz, M. Kayne, and C. Kondor-Koch. Constitutive apical secretion of an 80-kD sulfated glycoprotein complex in the polarized epithelial Madin–Darby canine kidney cell line. Journal of Cell Biology, 105(6):2735–??, December 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/105/6/
2735.

Ueno:1988:DCM

[UPRS88] H. Ueno, M. B. Perryman, R. Roberts, and M. D. Schneider. Differentiation of cardiac myocytes after mitogen withdrawal exhibits three sequential states of the ventricular growth response. Journal of Cell Biology, 107(5):1911–??, November 1988. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/107/5/1911.

Unwin:1988:AAR

[UTK88] N. Unwin, C. Toyoshima, and E. Kubalek. Arrangement of the acetylcholine receptor subunits in the resting and desensitized states, determined by cryoelectron microscopy of crystallized Torpedo postsynaptic membranes. Journal of Cell Biology, 107 (3):1123–??, September 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/107/3/1123.

Unemori:1986:RPA

[UW86] E. N. Unemori and Z. Werb. Reorganization of polymerized actin: a possible trigger for induction of procollagenase in fibroblasts cultured in and on collagen gels. Journal of Cell Biology, 103(3): 1021–??, September 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/103/3/1021.

Valberg:1985:CMR

[VA85] P. A. Valberg and D. F. Albertini. Cytoplasmic motions, rheology, and structure probed by a novel magnetic particle method. Journal of Cell Biology, 101(1):130–??, July 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/101/1/130.

vanAdelsberg:1986:RCP

[vAAA86] J. van Adelsberg and Q. Al-Awqati. Regulation of cell pH by Ca+2-mediated exocytotic insertion of H+-ATPases. Journal of Cell Biology, 102(5):1638–??, May 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/102/5/1638.

Vicentini:1985:MRI

[VAD^{+}85] L. M. Vicentini, A. Ambrosini, F. Di Virgilio, T. Pozzan, and J. Meldolesi. Muscarinic receptor-induced phosphoinositide hydrolysis at resting cytosolic Ca^{2+} concentration in PC12 cells. Journal of Cell Biology, 100(4):1330–??, April 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/100/4/1330.

VanHaastert:1987:DET

[Van87] P. J. Van Haastert. Differential effects of temperature on cAMPinduced excitation, adaptation, and deadaptation of adenylate and guanylate cyclase in Dictyostelium discoideum. Journal of Cell Biology, 105(5):2301–??, November 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/105/5/2301.

Valdivia:1985:FIC

[VB85a] M. M. Valdivia and B. R. Brinkley. Fractionation and initial characterization of the kinetochore from mammalian metaphase chromosomes. Journal of Cell Biology, 101(3):1124–??, September 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/101/3/
1124.

Verner:1985:MCE

[VB85b] K. Verner and A. Bretscher. Microvillus 110K-calmodulin: effects of nucleotides on isolated cytoskeletons and the interaction of the purified complex with F-actin. Journal of Cell Biology, 100(5):1455–??, May 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/100/5/1455.

Virshup:1988:CCV

[VB88a] D. M. Virshup and V. Bennett. Clathrin-coated vesicle assembly polypeptides: physical properties and reconstitution studies with brain membranes. Journal of Cell Biology, 106(1):39–??, January 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/106/1/
39.

Volpi:1988:IEF

[VB88b] M. Volpi and R. D. Berlin. Intracellular elevations of free calcium induced by activation of histamine H1 receptors in interphase and mitotic HeLa cells: hormone signal transduction is altered during mitosis. Journal of Cell Biology, 107(6):2533–??, December 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/107/6/
2533.

Vassalli:1985:CBS

[VBB85] J. D. Vassalli, D. Baccino, and D. Belin. A cellular binding site for the Mr 55,000 form of the human plasminogen activator, urokinase. Journal of Cell Biology, 100(1):86–??, January 1985. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/100/1/86.

Vivarelli:1988:ESM

[VBWC88] E. Vivarelli, W. E. Brown, R. G. Whalen, and G. Cossu. The expression of slow myosin during mammalian somitogenesis and limb bud differentiation. Journal of Cell Biology, 107 (6):2191–??, December 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/107/6/2191.

Vibert:1985:SCO

[VC85] P. Vibert and R. Craig. Structural changes that occur in scallop myosin filaments upon activation. Journal of Cell Biology, 101 (3):830–??, September 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/101/3/830.

Vibert:1989:SAP

[VC89] P. Vibert and L. Castellani. Substructure and accessory proteins in scallop myosin filaments. Journal of Cell Biology, 109 (2):539–??, August 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/2/539.

Vigers:1988:FMB

[VCM88] G. P. Vigers, M. Coue, and J. R. McIntosh. Fluorescent microtubules break up under illumination. Journal of Cell Biology, 107 (3):1011–??, September 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/107/3/1011.

vandenPol:1986:GCL

[vdPdPR86] A. N. van den Pol, U. di Porzio, and U. Rutishauser. Growth cone localization of neural cell adhesion molecule on central nervous system neurons in vitro. Journal of Cell Biology, 102(6):2281–??, June 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/102/6/
2281.

vanDeurs:1988:EAI

[vDSP^{+}88] B. van Deurs, K. Sandvig, O. W. Petersen, S. Olsnes, K. Simons, and G. Griffiths. Estimation of the amount of internalized ricin that reaches the trans-Golgi network. Journal of Cell Biology, 106(2):253–??, February 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/106/2/253.

vanDeurs:1986:RIR

[vDTP^{+}86] B. van Deurs, T. I. Tønnessen, O. W. Petersen, K. Sandvig, and S. Olsnes. Routing of internalized ricin and ricin conjugates to the Golgi complex. Journal of Cell Biology, 102(1):37–??, January 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/102/1/
37.

Varnum:1985:DAA

[VES85] B. Varnum, K. B. Edwards, and D. R. Soll. Dictyostelium amebae alter motility differently in response to increasing versus decreasing temporal gradients of cAMP. Journal of Cell Biology, 101(1):1–??, July 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/101/1/1.

Vaughan:1989:EMD

[VFB^{+}89] D. K. Vaughan, S. K. Fisher, S. A. Bernstein, I. L. Hale, K. A. Linberg, and B. Matsumoto. Evidence that microtubules do not mediate opsin vesicle transport in photoreceptors. Journal of Cell Biology, 109(6):3053–??, December 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/109/6/3053.

Volk:1986:CKRa

[VG86a] T. Volk and B. Geiger. A-CAM: a 135-kD receptor of intercellular adherens junctions. I. Immunoelectron microscopic localization and biochemical studies. Journal of Cell Biology, 103 (4):1441–??, October 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/103/4/1441.

Volk:1986:CKRb

[VG86b] T. Volk and B. Geiger. A-CAM: a 135-kD receptor of intercellular adherens junctions. II. Antibody-mediated modulation of junction formation. Journal of Cell Biology, 103(4):1451–??, October 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/103/4/
1451.

Volberg:1986:CMM

[VGKF86] T. Volberg, B. Geiger, J. Kartenbeck, and W. W. Franke. Changes in membrane-microfilament interaction in intercellular adherens junctions upon removal of extracellular Ca^{2+} ions. Journal of Cell Biology, 102(5):1832–??, May 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/102/5/1832.

vonGrafenstein:1986:KAT

[vGRB86] H. von Grafenstein, C. S. Roberts, and P. F. Baker. Kinetic analysis of the triggered exocytosis/endocytosis secretory cycle in cultured bovine adrenal medullary cells. Journal of Cell Biology, 103 (6):2343–??, December 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/103/6/2343.

Vale:1988:FMN

[VH88] R. D. Vale and H. Hotani. Formation of membrane networks in vitro by kinesin-driven microtubule movement. Journal of Cell Biology, 107(6):2233–??, December 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/107/6/2233.

VanHooff:1989:NGF

[VHO^{+}89] C. O. Van Hooff, J. C. Holthuis, A. B. Oestreicher, J. Boonstra, P. N. De Graan, and W. H. Gispen. Nerve growth factor-induced changes in the intracellular localization of the protein kinase C substrate B-50 in pheochromocytoma PC12 cells. Journal of Cell Biology, 108(3):1115–??, March 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/108/3/1115.

Valtorta:1988:SPF

[VJF^{+}88] F. Valtorta, R. Jahn, R. Fesce, P. Greengard, and B. Ceccarelli. Synaptophysin (p38) at the frog neuromuscular junction: its incorporation into the axolemma and recycling after intense quantal secretion. Journal of Cell Biology, 107(6):2717–??, December 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/107/6/
2717.

Vainio:1989:STE

[VJT89] S. Vainio, M. Jalkanen, and I. Thesleff. Syndecan and tenascin expression is induced by epithelial-mesenchymal interactions in embryonic tooth mesenchyme. Journal of Cell Biology, 108 (5):1945–??, May 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/108/5/1945.

Vance:1989:HSR

[VKB89] B. A. Vance, C. G. Kowalski, and C. E. Brinckerhoff. Heat shock of rabbit synovial fibroblasts increases expression of mRNAs for two metalloproteinases, collagenase and stromelysin. Journal of Cell Biology, 108(6):2037–??, June 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/108/6/2037.

Vandekerckhove:1989:AAP

[VKP89] J. S. Vandekerckhove, D. A. Kaiser, and T. D. Pollard. Acanthamoeba actin and profilin can be cross-linked between glutamic acid 364 of actin and lysine 115 of profilin. Journal of Cell Biology, 109(2):619–??, August 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/109/2/619.

Vantard:1985:CID

[VLD^{+}85] M. Vantard, A. M. Lambert, J. De Mey, P. Picquot, and L. J. Van Eldik. Characterization and immunocytochemical distribution of calmodulin in higher plant endosperm cells: localization in the mitotic apparatus. Journal of Cell Biology, 101(2):488–??, August 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/101/2/

Vanderkooi:1985:MFD

[VME85] J. M. Vanderkooi, G. Maniara, and M. Erecinska. Mobility of fluorescent derivatives of cytochrome c in mitochondria. Journal of Cell Biology, 100(2):435–??, February 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/100/2/435.

Vaughan:1988:DPD

[VMH^{+}88] L. Vaughan, M. Mendler, S. Huber, P. Bruckner, K. H. Winterhalter, M. I. Irwin, and R. Mayne. D-periodic distribution of collagen type IX along cartilage fibrils. Journal of Cell Biology, 106(3):991–??, March 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/106/3/991.

vanMeer:1987:SSE

[vMSvRS87] G. van Meer, E. H. Stelzer, R. W. Wijnaendts van Resandt, and K. Simons. Sorting of sphingolipids in epithelial (Madin– Darby canine kidney) cells. Journal of Cell Biology, 105(4): 1623–??, October 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/4/1623.

Vacquier:1988:MAM

[VMT^{+}88] V. D. Vacquier, G. W. Moy, J. S. Trimmer, Y. Ebina, and D. C. Porter. Monoclonal antibodies to a membrane glycoprotein induce the phosphorylation of histone H1 in sea urchin spermatozoa. Journal of Cell Biology, 107(6):2021–??, December 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/107/6/2021.

Vanderburg:1988:PCE

[VN88] C. R. Vanderburg and M. A. Nathanson. Posttranscriptional control of embryonic rat skeletal muscle protein synthesis. Control at the level of translation by endogenous RNA. Journal of Cell Biology, 107(3):1085–??, September 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/107/3/1085.

VanTuyle:1985:RLM

[VP85] G. C. Van Tuyle and P. A. Pavco. The rat liver mitochondrial DNA-protein complex: displaced single strands of replicative intermediates are protein coated. Journal of Cell Biology, 100 (1):251–??, January 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/100/1/251.

Volonte:1989:DIN

[VRLG89] C. Volonté, A. Rukenstein, D. M. Loeb, and L. A. Greene. Differential inhibition of nerve growth factor responses by purine analogues: correlation with inhibition of a nerve growth factoractivated protein kinase. Journal of Cell Biology, 109(5): 2395–??, November 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/5/2395.

Vogl:1985:APF

[VS85] A. W. Vogl and L. J. Soucy. Arrangement and possible function of actin filament bundles in ectoplasmic specializations of ground squirrel Sertoli cells. Journal of Cell Biology, 100(3):814–??, March 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
100/3/814.

Vestweber:1988:CMP

[VS88a] D. Vestweber and G. Schatz. A chimeric mitochondrial precursor protein with internal disulfide bridges blocks import of authentic precursors into mitochondria and allows quantitation of import sites. Journal of Cell Biology, 107(6):2037–??, December 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/107/6/
2037.

Vestweber:1988:MCI

[VS88b] D. Vestweber and G. Schatz. Mitochondria can import artificial precursor proteins containing a branched polypeptide chain or a carboxy-terminal stilbene disulfonate. Journal of Cell Biology, 107(6):2045–??, December 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/107/6/2045.

Verrall:1989:CTP

[VS89] S. Verrall and N. W. Seeds. Characterization of 125I-tissue plasminogen activator binding to cerebellar granule neurons. Journal of Cell Biology, 109(1):265–??, July 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/109/1/265.

Vilmart-Seuwen:1986:AKE

[VSKSP86] J. Vilmart-Seuwen, H. Kersken, R. Stürzl, and H. Plattner. ATP keeps exocytosis sites in a primed state but is not required for membrane fusion: an analysis with Paramecium cells in vivo and in vitro. Journal of Cell Biology, 103(4):1279–??, October 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/103/4/
1279.

Volloch:1987:SGR

[VSR87] V. Volloch, B. Schweitzer, and S. Rits. Synthesis of globin RNA in enucleated differentiating murine erythroleukemia cells. Journal of Cell Biology, 105(1):137–??, July 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/105/1/137.

Vega-Salas:1987:FAP

[VSSGRB87] D. E. Vega-Salas, P. J. Salas, D. Gundersen, and E. Rodriguez-Boulan. Formation of the apical pole of epithelial (Madin–Darby canine kidney) cells: polarity of an apical protein is independent of tight junctions while segregation of a basolateral marker requires cell–cell interactions. Journal of Cell Biology, 104(4): 905–??, January 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/104/4/905.

Vega-Salas:1987:MEA

[VSSRB87] D. E. Vega-Salas, P. J. Salas, and E. Rodriguez-Boulan. Modulation of the expression of an apical plasma membrane protein of Madin–Darby canine kidney epithelial cells: cell–cell interactions control the appearance of a novel intracellular storage compartment. Journal of Cell Biology, 104(5):1249–??, May 1987. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/104/5/1249.

Vega-Salas:1988:EVA

[VSSRB88] D. E. Vega-Salas, P. J. Salas, and E. Rodriguez-Boulan. Exocytosis of vacuolar apical compartment (VAC): a cell–cell contact controlled mechanism for the establishment of the apical plasma membrane domain in epithelial cells. Journal of Cell Biology, 107 (5):1717–??, November 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/107/5/1717.

Verrey:1987:RAK

[VSZ^{+}87] F. Verrey, E. Schaerer, P. Zoerkler, M. P. Paccolat, K. Geering, J. P. Kraehenbuhl, and B. C. Rossier. Regulation by aldosterone of Na+,K+-ATPase mRNAs, protein synthesis, and sodium transport in cultured kidney cells. Journal of Cell Biology, 104(5):1231–??, May 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/104/5/1231.

Vale:1989:MTP

[VT89] R. D. Vale and Y. Y. Toyoshima. Microtubule translocation properties of intact and proteolytically digested dyneins from Tetrahymena cilia. Journal of Cell Biology, 108(6):2327–??, June 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/108/6/
2327.

Vertel:1989:PCS

[VVL^{+}89] B. M. Vertel, A. Velasco, S. LaFrance, L. Walters, and K. Kaczman-Daniel. Precursors of chondroitin sulfate proteoglycan are segregated within a subcompartment of the chondrocyte endoplasmic reticulum. Journal of Cell Biology, 109 (4):1827–??, October 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/4/1827.

vonZastrow:1987:PSA

[vZC87] M. von Zastrow and J. D. Castle. Protein sorting among two distinct export pathways occurs from the content of maturing exocrine storage granules. Journal of Cell Biology, 105(6): 2675–??, December 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/6/2675.

White:1987:ECV

[WAF87] J. G. White, W. B. Amos, and M. Fordham. An evaluation of confocal versus conventional imaging of biological structures by fluorescence light microscopy. Journal of Cell Biology, 105(1):41– ??, July 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
105/1/41.

White:1989:NST

[WAG^{+}89] E. M. White, C. D. Allis, D. S. Goldfarb, A. Srivastva, J. W. Weir, and M. A. Gorovsky. Nucleus-specific and temporally restricted localization of proteins in Tetrahymena macronuclei and micronuclei. Journal of Cell Biology, 109(5):1983–??, November 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/109/5/
1983.

Wallace:1986:AFT

[Wal86] B. G. Wallace. Aggregating factor from Torpedo electric organ induces patches containing acetylcholine receptors, acetylcholinesterase, and butyrylcholinesterase on cultured myotubes. Journal of Cell Biology, 102(3):783–??, March 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/102/3/783.

Wallace:1988:RAI

[Wal88] B. G. Wallace. Regulation of agrin-induced acetylcholine receptor aggregation by Ca++ and phorbol ester. Journal of Cell Biology, 107(1):267–??, July 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/107/1/267.

Wang:1985:MWP

[Wan85a] E. Wang. A 57,000-mol-wt protein uniquely present in nonproliferating cells and senescent human fibroblasts. Journal of Cell Biology, 100(2):545–??, February 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/100/2/545.

Wang:1985:CBS

[Wan85b] E. Wang. Are cross-bridging structures involved in the bundle formation of intermediate filaments and the decrease in locomotion that accompany cell aging? Journal of Cell Biology, 100(5):1466–??, May 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/100/5/1466.

Wang:1985:RDS

[Wan85c] E. Wang. Rapid disappearance of statin, a nonproliferating and senescent cell-specific protein, upon reentering the process of cell cycling. Journal of Cell Biology, 101(5):1695–??, November 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/101/5/
1695.

Wang:1985:EAS

[Wan85d] Y. L. Wang. Exchange of actin subunits at the leading edge of living fibroblasts: possible role of treadmilling. Journal of Cell Biology, 101(2):597–??, August 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/101/2/597.

Wang:1987:MFA

[Wan87] Y. L. Wang. Mobility of filamentous actin in living cytoplasm. Journal of Cell Biology, 105(6):2811–??, December 1987. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/105/6/2811.

Watts:1985:RET

[Wat85] C. Watts. Rapid endocytosis of the transferrin receptor in the absence of bound transferrin. Journal of Cell Biology, 100(2): 633–??, February 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/100/2/633.

Wong:1985:MBP

[WB85] D. T. Wong and D. K. Biswas. Mechanism of benzo(a)pyrene induction of alpha-human chorionic gonadotropin gene expression in human lung tumor cells. Journal of Cell Biology, 101(6): 2245–??, December 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/101/6/2245.

Waters:1986:SPT

[WB86] M. G. Waters and G. Blobel. Secretory protein translocation in a yeast cell-free system can occur posttranslationally and requires ATP hydrolysis. Journal of Cell Biology, 102(5):1543–??, May 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/102/5/
1543.

Wozniak:1989:PSA

[WBB89] R. W. Wozniak, E. Bartnik, and G. Blobel. Primary structure analysis of an integral membrane glycoprotein of the nuclear pore. Journal of Cell Biology, 108(6):2083–??, June 1989. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/108/6/2083.

Wiegant:1986:EGF

[WBD^{+}86] F. A. Wiegant, F. J. Blok, L. H. Defize, W. A. Linnemans, A. J. Verkley, and J. Boonstra. Epidermal growth factor receptors associated to cytoskeletal elements of epidermoid carcinoma (A431) cells. Journal of Cell Biology, 103(1):87–??, July 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/103/1/87.

Wright:1988:IAH

[WBDR88] R. Wright, M. Basson, L. D’Ari, and J. Rine. Increased amounts of HMG–CoA reductase induce “karmellae”: a proliferation of stacked membrane pairs surrounding the yeast nucleus. Journal of Cell Biology, 107(1):101–??, July 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/107/1/101.

Ward:1985:CAP

[WBGV85] G. E. Ward, C. J. Brokaw, D. L. Garbers, and V. D. Vacquier. Chemotaxis of Arbacia punctulata spermatozoa to resact, a peptide from the egg jelly layer. Journal of Cell Biology, 101(6): 2324–??, December 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/101/6/2324.

Woods:1986:CED

[WBVL86] C. M. Woods, B. Boyer, P. K. Vogt, and E. Lazarides. Control of erythroid differentiation: asynchronous expression of the anion transporter and the peripheral components of the membrane skeleton in AEV- and S13-transformed cells. Journal of Cell Biology, 103(5):1789–??, November 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/103/5/1789.

Weisman:1987:MMV

[WBW87] L. S. Weisman, R. Bacallao, and W. Wickner. Multiple methods of visualizing the yeast vacuole permit evaluation of its morphology and inheritance during the cell cycle. Journal of Cell Biology, 105(4):1539–??, October 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/105/4/1539.

West:1989:DEP

[WBW89] M. A. West, M. S. Bretscher, and C. Watts. Distinct endocytotic pathways in epidermal growth factor-stimulated human carcinoma A431 cells. Journal of Cell Biology, 109(6): 2731–??, December 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/6/2731.

Wayner:1987:IMC

[WC87] E. A. Wayner and W. G. Carter. Identification of multiple cell adhesion receptors for collagen and fibronectin in human fibrosarcoma cells possessing unique alpha and common beta subunits. Journal of Cell Biology, 105(4):1873–??, October 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/105/4/1873.

Waters:1986:PTA

[WCB86] M. G. Waters, W. J. Chirico, and G. Blobel. Protein translocation across the yeast microsomal membrane is stimulated by a soluble factor. Journal of Cell Biology, 103(6):2629–??, December 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/103/6/
2629.

Wilson:1988:IMP

[WCMG88] C. Wilson, T. Connolly, T. Morrison, and R. Gilmore. Integration of membrane proteins into the endoplasmic reticulum requires GTP. Journal of Cell Biology, 107(1):69–??, July 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/107/1/69.

Wayner:1988:FME

[WCPK88] E. A. Wayner, W. G. Carter, R. S. Piotrowicz, and T. J. Kunicki. The function of multiple extracellular matrix receptors in mediating cell adhesion to extracellular matrix: preparation of monoclonal antibodies to the fibronectin receptor that specifically inhibit cell adhesion to fibronectin and react with platelet glycoproteins Ic–IIa. Journal of Cell Biology, 107(5):1881–??, November 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/107/5/

Wischik:1985:SSP

[WCSR85] C. M. Wischik, R. A. Crowther, M. Stewart, and M. Roth. Subunit structure of paired helical filaments in Alzheimer’s disease. Journal of Cell Biology, 100(6):1905–??, June 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/100/6/1905.

Waddell:1986:BSN

[WD86a] D. R. Waddell and K. T. Duffy. Breakdown of self/nonself recognition in cannibalistic strains of the predatory slime mold, Dictyostelium caveatum. Journal of Cell Biology, 102(1):298–??, January 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
102/1/298.

Wagner:1986:NOI

[WD86b] J. A. Wagner and P. A. D’Amore. Neurite outgrowth induced by an endothelial cell mitogen isolated from retina. Journal of Cell Biology, 103(4):1363–??, October 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/103/4/1363.

Woods:1986:TRR

[WDF86] J. W. Woods, M. Doriaux, and M. G. Farquhar. Transferrin receptors recycle to cis and middle as well as trans Golgi cisternae in Ig-secreting myeloma cells. Journal of Cell Biology, 103(1):277–??, July 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/103/1/277.

Wittig:1989:LZD

[WDR89] B. Wittig, T. Dorbic, and A. Rich. The level of Z–DNA in metabolically active, permeabilized mammalian cell nuclei is regulated by torsional strain. Journal of Cell Biology, 108(3):755–??, March 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
108/3/755.

Waddell:1987:CDD

[WDV87] D. R. Waddell, K. Duffy, and G. Vogel. Cytokinesis is defective in Dictyostelium mutants with altered phagocytic recognition, adhesion, and vegetative cell cohesion properties. Journal of Cell Biology, 105(5):2293–??, November 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/105/5/2293.

Wier:1986:ECD

[WE86a] M. L. Wier and M. Edidin. Effects of cell density and extracellular matrix on the lateral diffusion of major histocompatibility antigens in cultured fibroblasts. Journal of Cell Biology, 103(1):215–??, July 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/103/1/215.

Woodman:1986:CFA

[WE86b] P. G. Woodman and J. M. Edwardson. A cell-free assay for the insertion of a viral glycoprotein into the plasma membrane. Journal of Cell Biology, 103(5):1829–??, November 1986. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/103/5/1829.

Wenderoth:1987:INM

[WE87] M. P. Wenderoth and B. R. Eisenberg. Incorporation of nascent myosin heavy chains into thick filaments of cardiac myocytes in thyroid-treated rabbits. Journal of Cell Biology, 105(6): 2771–??, December 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/6/2771.

Waugh:1989:FRM

[WE89] R. E. Waugh and G. Erwin. Flexural rigidity of marginal bands isolated from erythrocytes of the newt. Journal of Cell Biology, 108(5):1711–??, May 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/108/5/1711.

Wallis:1985:ING

[WESP85] I. Wallis, L. Ellis, K. Suh, and K. H. Pfenninger. Immunolocalization of a neuronal growth-dependent membrane glycoprotein. Journal of Cell Biology, 101(5):1990–??, November 1985. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/101/5/1990.

White:1986:EID

[WF86] G. E. White and K. Fujiwara. Expression and intracellular distribution of stress fibers in aortic endothelium. Journal of Cell Biology, 103(1):63–??, July 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/103/1/63.

Wade:1989:TDC

[WFE89] W. F. Wade, J. H. Freed, and M. Edidin. Translational diffusion of class II major histocompatibility complex molecules is constrained by their cytoplasmic domains. Journal of Cell Biology, 109(6):3325–??, December 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/109/6/3325.

Wolfgang:1986:PCD

[WFF86] W. J. Wolfgang, D. Fristrom, and J. W. Fristrom. The pupal cuticle of Drosophila: differential ultrastructural immunolocalization of cuticle proteins. Journal of Cell Biology, 102(1):306–??, January 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/102/1/
306.

Wilson:1987:RLE

[WFL87] P. D. Wilson, R. A. Firestone, and J. Lenard. The role of lysosomal enzymes in killing of mammalian cells by the lysosomotropic detergent N-dodecylimidazole. Journal of Cell Biology, 104(5):1223–??, May 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/104/5/1223.

Watanabe:1986:TCS

[WFR86] M. Watanabe, A. L. Frelinger, and U. Rutishauser. Topography of N-CAM structural and functional determinants. I. Classification of monoclonal antibody epitopes. Journal of Cell Biology, 103(5):1721–??, November 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/103/5/1721.

Warn:1987:IMO

[WFW87] R. M. Warn, L. Flegg, and A. Warn. An investigation of microtubule organization and functions in living Drosophila embryos by injection of a fluorescently labeled antibody against tyrosinated alpha-tubulin. Journal of Cell Biology, 105(4):1721–??, October 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
105/4/1721.

Winkles:1985:DSD

[WG85] J. A. Winkles and R. M. Grainger. Differential stability of Drosophila embryonic mRNAs during subsequent larval development. Journal of Cell Biology, 101(5):1808–??, November 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/101/5/
1808.

Whiteley:1986:EGF

[WG86] B. Whiteley and L. Glaser. Epidermal growth factor (EGF) promotes phosphorylation at threonine-654 of the EGF receptor: possible role of protein kinase C in homologous regulation of the EGF receptor. Journal of Cell Biology, 103(4):1355–??, October 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/103/4/
1355.

Wang:1987:DSD

[WG87a] Y. L. Wang and M. B. Goren. Differential and sequential delivery of fluorescent lysosomal probes into phagosomes in mouse peritoneal macrophages. Journal of Cell Biology, 104(6):1749–??, June 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/104/6/
1749.

Warner:1987:FGJ

[WG87b] A. Warner and J. B. Gurdon. Functional gap junctions are not required for muscle gene activation by induction in Xenopus embryos. Journal of Cell Biology, 104(3):557–??, March 1987. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/104/3/557.

Webster:1988:ICV

[WG88a] P. Webster and D. J. Grab. Intracellular colocalization of variant surface glycoprotein and transferrin-gold in Trypanosoma brucei. Journal of Cell Biology, 106(2):279–??, February 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/106/2/279.

Wong:1988:RMC

[WG88b] M. K. Wong and A. I. Gotlieb. The reorganization of microfilaments, centrosomes, and microtubules during in vitro small wound reendothelialization. Journal of Cell Biology, 107(5):

1777–??, November 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/107/5/1777.

Williams:1985:INC

[WGA85] R. K. Williams, C. Goridis, and R. Akeson. Individual neural cell types express immunologically distinct N-CAM forms. Journal of Cell Biology, 101(1):36–??, July 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/101/1/36.

Webster:1987:ATD

[WGBB87] D. R. Webster, G. G. Gundersen, J. C. Bulinski, and G. G. Borisy. Assembly and turnover of detyrosinated tubulin in vivo. Journal of Cell Biology, 105(1):265–??, July 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/105/1/265.

Wheeler:1985:RTR

[WGC85] M. E. Wheeler, J. M. Gerrard, and R. C. Carroll. Reciprocal transmembranous receptor-cytoskeleton interactions in concanavalin A-activated platelets. Journal of Cell Biology, 101(3): 993–??, September 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/101/3/993.

Wayner:1989:ICL

[WGPH^{+}89] E. A. Wayner, A. Garcia-Pardo, M. J. Humphries, J. A. Mc-Donald, and W. G. Carter. Identification and characterization of the T lymphocyte adhesion receptor for an alternative cell attachment domain (CS-1) in plasma fibronectin. Journal of Cell Biology, 109(3):1321–??, September 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/109/3/1321.

Way:1989:EHP

[WGPW89] M. Way, J. Gooch, B. Pope, and A. G. Weeds. Expression of human plasma gelsolin in Escherichia coli and dissection of actin binding sites by segmental deletion mutagenesis. Journal of Cell Biology, 109(2):593–??, August 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/109/2/593.

Wang:1988:SCM

[WGS^{+}88] S. M. Wang, M. L. Greaser, E. Schultz, J. C. Bulinski, J. J. Lin, and J. L. Lessard. Studies on cardiac myofibrillogenesis with antibodies to titin, actin, tropomyosin, and myosin. Journal of Cell Biology, 107(3):1075–??, September 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/107/3/1075.

Wall:1985:RME

[WH85] D. A. Wall and A. L. Hubbard. Receptor-mediated endocytosis of asialoglycoproteins by rat liver hepatocytes: biochemical characterization of the endosomal compartments. Journal of Cell Biology, 101(6):2104–??, December 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/101/6/2104.

Wolf:1986:CSP

[WHI86] D. E. Wolf, S. S. Hagopian, and S. Ishijima. Changes in sperm plasma membrane lipid diffusibility after hyperactivation during in vitro capacitation in the mouse. Journal of Cell Biology, 102 (4):1372–??, April 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/102/4/1372.

Wolf:1986:LRD

[WHL^{+}86] D. E. Wolf, S. S. Hagopian, R. G. Lewis, J. K. Voglmayr, and G. Fairbanks. Lateral regionalization and diffusion of a maturation-dependent antigen in the ram sperm plasma membrane. Journal of Cell Biology, 102(5):1826–??, May 1986. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/102/5/1826.

Werb:1986:CEM

[WHMA86] Z. Werb, R. M. Hembry, G. Murphy, and J. Aggeler. Commitment to expression of the metalloendopeptidases, collagenase and stromelysin: relationship of inducing events to changes in cytoskeletal architecture. Journal of Cell Biology, 102(3):697–??, March 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
102/3/697.

Wiley:1988:ABE

[Wil88] H. S. Wiley. Anomalous binding of epidermal growth factor to A431 cells is due to the effect of high receptor densities and a saturable endocytic system. Journal of Cell Biology, 107 (2):801–??, August 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/107/2/801.

Walker:1989:ABS

[WIS89] R. A. Walker, S. Inoué, and E. D. Salmon. Asymmetric behavior of severed microtubule ends after ultraviolet-microbeam irradiation of individual microtubules in vitro. Journal of Cell Biology, 108(3):931–??, March 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/108/3/931.

Wu:1988:ECT

[WJSB88] N. W. Wu, S. Jalkanen, P. R. Streeter, and E. C. Butcher. Evolutionary conservation of tissue-specific lymphocyte-endothelial cell recognition mechanisms involved in lymphocyte homing. Journal of Cell Biology, 107(5):1845–??, November 1988. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/107/5/1845.

Wood:1985:TAM

[WK85] L. Wood and A. Kaplan. Transit of alpha-mannosidase during its maturation in Dictyostelium discoideum. Journal of Cell Biology, 101(6):2063–??, December 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/101/6/2063.

Wiedmann:1987:DPI

[WKBR87] M. Wiedmann, T. V. Kurzchalia, H. Bielka, and T. A. Rapoport. Direct probing of the interaction between the signal sequence of nascent preprolactin and the signal recognition particle by specific cross-linking. Journal of Cell Biology, 104(2):201–??, February 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/104/2/
201.

Wataya-Kaneda:1987:MAA

[WKKS^{+}87] M. Wataya-Kaneda, Y. Kaneda, T. Sakurai, H. Sugawa, and T. Uchida. A monoclonal antibody against the nucleus reveals the presence of a common protein in the nuclear envelope, the perichromosomal region, and cytoplasmic vesicles. Journal of Cell Biology, 104(1):1–??, January 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/104/1/1.

Wu:1986:DBH

[WKR86] L. H. Wu, L. Kuehl, and M. Rechsteiner. Dynamic behavior of histone H1 microinjected into HeLa cells. Journal of Cell Biology, 103(2):465–??, August 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/103/2/465.

Weissman:1986:EKC

[WKRH86] A. M. Weissman, R. D. Klausner, K. Rao, and J. B. Harford. Exposure of K562 cells to anti-receptor monoclonal antibody OKT9 results in rapid redistribution and enhanced degradation of the transferrin receptor. Journal of Cell Biology, 102 (3):951–??, March 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/102/3/951.

Weidman:1985:ASU

[WKS85] P. J. Weidman, E. S. Kay, and B. M. Shapiro. Assembly of the sea urchin fertilization membrane: isolation of proteoliaisin, a calcium-dependent ovoperoxidase binding protein. Journal of Cell Biology, 100(3):938–??, March 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/100/3/938.

Wilkins:1986:REI

[WL86] J. A. Wilkins and S. Lin. A re-examination of the interaction of vinculin with actin. Journal of Cell Biology, 102(3):1085–??, March 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
102/3/1085.

Wuestehube:1987:FAB

[WL87] L. J. Wuestehube and E. J. Luna. F-actin binds to the cytoplasmic surface of ponticulin, a 17-kD integral glycoprotein from Dictyostelium discoideum plasma membranes. Journal of Cell Biology, 105(4):1741–??, October 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/105/4/1741.

Winter:1985:CHHa

[WLG85] E. Winter, D. Levy, and J. S. Gordon. Changes in the H-1 histone complement during myogenesis. I. Establishment by differential coupling of H-1 species synthesis to DNA replication. Journal of Cell Biology, 101(1):167–??, July 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/101/1/167.

Wiedenmann:1985:SPB

[WLGB85] B. Wiedenmann, K. Lawley, C. Grund, and D. Branton. Solubilization of proteins from bovine brain coated vesicles by protein perturbants and Triton X-100. Journal of Cell Biology, 101(1):12–??, July 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/101/1/12.

Weltzin:1989:BTT

[WLJM^{+}89] R. Weltzin, P. Lucia-Jandris, P. Michetti, B. N. Fields, J. P. Kraehenbuhl, and M. R. Neutra. Binding and transepithelial transport of immunoglobulins by intestinal M cells: demonstration using monoclonal IgA antibodies against enteric viral proteins. Journal of Cell Biology, 108(5):1673–??, May 1989. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/108/5/1673.

Wistow:1988:TCA

[WLW^{+}88] G. J. Wistow, T. Lietman, L. A. Williams, S. O. Stapel, W. W. de Jong, J. Horwitz, and J. Piatigorsky. Tau-crystallin/alpha-enolase: one gene encodes both an enzyme and a lens structural protein. Journal of Cell Biology, 107(6):2729–??, December 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/107/6/
2729.

Wall:1985:ULC

[WM85] D. A. Wall and I. Meleka. An unusual lysosome compartment involved in vitellogenin endocytosis by Xenopus oocytes. Journal of Cell Biology, 101(5):1651–??, November 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/101/5/1651.

Welch:1988:CTC

[WM88] W. J. Welch and L. A. Mizzen. Characterization of the thermotolerant cell. II. Effects on the intracellular distribution of heat-shock protein 70, intermediate filaments, and small nuclear ribonucleoprotein complexes. Journal of Cell Biology, 106 (4):1117–??, April 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/106/4/1117.

Weidman:1989:BES

[WMBR89] P. J. Weidman, P. Melançon, M. R. Block, and J. E. Rothman. Binding of an N-ethylmaleimide-sensitive fusion protein to Golgi membranes requires both a soluble protein(s) and an integral membrane receptor. Journal of Cell Biology, 108(5):1589–??, May 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/108/5/
1589.

Wordeman:1986:DCM

[WMC86] L. Wordeman, K. L. McDonald, and W. Z. Cande. The distribution of cytoplasmic microtubules throughout the cell cycle of the centric diatom Stephanopyxis turris: their role in nuclear migration and positioning the mitotic spindle during cytokinesis. Journal of Cell Biology, 102(5):1688–??, May 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/102/5/1688.

Weatherbee:1985:IPS

[WMGM85] J. A. Weatherbee, G. S. May, J. Gambino, and N. R. Morris. Involvement of a particular species of beta-tubulin (beta 3) in conidial development in Aspergillus nidulans. Journal of Cell Biology, 101(3):706–??, September 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/101/3/706.

Wrana:1988:DET

[WMH^{+}88] J. L. Wrana, M. Maeno, B. Hawrylyshyn, K. L. Yao, C. Domenicucci, and J. Sodek. Differential effects of transforming growth factor-beta on the synthesis of extracellular matrix proteins by normal fetal rat calvarial bone cell populations. Journal of Cell Biology, 106(3):915–??, March 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/106/3/915.

Wagner:1986:IGA

[WMM86] D. D. Wagner, T. Mayadas, and V. J. Marder. Initial glycosylation and acidic pH in the Golgi apparatus are required for multimerization of von Willebrand factor. Journal of Cell Biology, 102(4):1320–??, April 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/102/4/1320.

Williams:1986:MCE

[WMR86] B. D. Williams, D. R. Mitchell, and J. L. Rosenbaum. Molecular cloning and expression of flagellar radial spoke and dynein genes of Chlamydomonas. Journal of Cell Biology, 103(1):1–??, July 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/103/1/
1.

Winningham-Major:1989:NEN

[WMSB^{+}89] F. Winningham-Major, J. L. Staecker, S. W. Barger, S. Coats, and L. J. Van Eldik. Neurite extension and neuronal survival activities of recombinant S100 beta proteins that differ in the content and position of cysteine residues. Journal of Cell Biology, 109(6):3063–??, December 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/109/6/3063.

Wagner:1985:IDB

[WMUP^{+}85] D. D. Wagner, T. Mayadas, M. Urban-Pickering, B. H. Lewis, and V. J. Marder. Inhibition of disulfide bonding of von Willebrand protein by monensin results in small, functionally defective multimers. Journal of Cell Biology, 101(1):112–??, July 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/101/1/
112.

Ward:1986:PMB

[WMV86] G. E. Ward, G. W. Moy, and V. D. Vacquier. Phosphorylation of membrane-bound guanylate cyclase of sea urchin spermatozoa. Journal of Cell Biology, 103(1):95–??, July 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/103/1/95.

Walworth:1987:PCC

[WN87] N. C. Walworth and P. J. Novick. Purification and characterization of constitutive secretory vesicles from yeast. Journal of Cell Biology, 105(1):163–??, July 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/105/1/163.

Wilson:1988:TSR

[WN88] K. L. Wilson and J. Newport. A trypsin-sensitive receptor on membrane vesicles is required for nuclear envelope formation in vitro. Journal of Cell Biology, 107(1):57–??, July 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/107/1/57.

Wieser:1986:CIG

[WO86] R. J. Wieser and F. Oesch. Contact inhibition of growth of human diploid fibroblasts by immobilized plasma membrane glycoproteins. Journal of Cell Biology, 103(2):361–??, August 1986. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/103/2/361.

Walker:1988:DII

[WOP^{+}88] R. A. Walker, E. T. O’Brien, N. K. Pryer, M. F. Soboeiro, W. A. Voter, H. P. Erickson, and E. D. Salmon. Dynamic instability of individual microtubules analyzed by video light microscopy: rate constants and transition frequencies. Journal of Cell Biology, 107 (4):1437–??, October 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/107/4/1437.

Wieczorek:1985:CEM

[WPBB^{+}85] D. F. Wieczorek, M. Periasamy, G. S. Butler-Browne, R. G. Whalen, and B. Nadal-Ginard. Co-expression of multiple myosin heavy chain genes, in addition to a tissue-specific one, in extraocular musculature. Journal of Cell Biology, 101(2):618–??, August 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/101/2/
618.

Winter:1985:CHHb

[WPW^{+}85] E. Winter, C. M. Palatnik, D. L. Williams, L. S. Coles, J. R. Wells, and J. S. Gordon. Changes in the H-1 histone complement during myogenesis. II. Regulation by differential coupling of H-1 variant mRNA accumulation to DNA replication. Journal of Cell Biology, 101(1):175–??, July 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/101/1/175.

Wang:1988:EGF

[WR88] Y. C. Wang and P. A. Rubenstein. Epidermal growth factor controls smooth muscle alpha-isoactin expression in BC3H1 cells. Journal of Cell Biology, 106(3):797–??, March 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/106/3/797.

Wright:1985:AEF

[Wri85] W. E. Wright. The amplified expression of factors regulating myogenesis in L6 myoblasts. Journal of Cell Biology, 100 (1):311–??, January 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/100/1/311.

Wilkins:1986:SPC

[WRL86] J. A. Wilkins, M. A. Risinger, and S. Lin. Studies on proteins that co-purify with smooth muscle vinculin: identification of immunologically related species in focal adhesions of nonmuscle and Z-lines of muscle cells. Journal of Cell Biology, 103 (4):1483–??, October 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/103/4/1483.

Wittenberg:1987:SLP

[WRR87] C. Wittenberg, S. L. Richardson, and S. I. Reed. Subcellular localization of a protein kinase required for cell cycle initiation in Saccharomyces cerevisiae: evidence for an association between the CDC28 gene product and the insoluble cytoplasmic matrix. Journal of Cell Biology, 105(4):1527–??, October 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/105/4/1527.

Ward:1986:MAR

[WRS^{+}86] S. Ward, T. M. Roberts, S. Strome, F. M. Pavalko, and E. Hogan. Monoclonal antibodies that recognize a polypeptide antigenic determinant shared by multiple Caenorhabditis elegans spermspecific proteins. Journal of Cell Biology, 102(5):1778–??, May 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/102/5/
1778.

Welch:1985:MSM

[WS85] W. J. Welch and J. P. Suhan. Morphological study of the mammalian stress response: characterization of changes in cytoplasmic organelles, cytoskeleton, and nucleoli, and appearance of intranuclear actin filaments in rat fibroblasts after heat-shock treatment. Journal of Cell Biology, 101(4):1198–??, October 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/101/4/
1198.

Wadsworth:1986:ATM

[WS86a] P. Wadsworth and E. D. Salmon. Analysis of the treadmilling model during metaphase of mitosis using fluorescence redistribution after photobleaching. Journal of Cell Biology, 102(3): 1032–??, March 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/102/3/1032.

Welch:1986:CBE

[WS86b] W. J. Welch and J. P. Suhan. Cellular and biochemical events in mammalian cells during and after recovery from physiological stress. Journal of Cell Biology, 103(5):2035–??, November 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/103/5/
2035.

Wier:1986:RTE

[WS86c] M. L. Wier and R. E. Scott. Regulation of the terminal event in cellular differentiation: biological mechanisms of the loss of proliferative potential. Journal of Cell Biology, 102(5):1955–??, May 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/102/5/
1955.

Weidman:1987:REM

[WS87] P. J. Weidman and B. M. Shapiro. Regulation of extracellular matrix assembly: in vitro reconstitution of a partial fertilization envelope from isolated components. Journal of Cell Biology, 105(1):561–??, July 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/1/561.

Winograd:1989:CMR

[WS89] E. Winograd and I. W. Sherman. Characterization of a modified red cell membrane protein expressed on erythrocytes infected with the human malaria parasite Plasmodium falciparum: possible role as a cytoadherent mediating protein. Journal of Cell Biology, 108(1):23–??, January 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/108/1/23.

Williams:1985:ITM

[WSH85] D. B. Williams, S. J. Swiedler, and G. W. Hart. Intracellular transport of membrane glycoproteins: two closely related histocompatibility antigens differ in their rates of transit to the cell surface. Journal of Cell Biology, 101(3):725–??, September 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/101/3/
725.

Wright:1985:NBB

[WSJ85] R. L. Wright, J. Salisbury, and J. W. Jarvik. A nucleus-basal body connector in Chlamydomonas reinhardtii that may function in basal body localization or segregation. Journal of Cell Biology, 101(5):1903–??, November 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/101/5/1903.

Witke:1986:STT

[WSLN86] W. Witke, M. Schleicher, F. Lottspeich, and A. Noegel. Studies on the transcription, translation, and structure of alpha-actinin in Dictyostelium discoideum. Journal of Cell Biology, 103(3): 969–??, September 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/103/3/969.

Wolf:1986:DRL

[WSM86] D. E. Wolf, B. K. Scott, and C. F. Millette. The development of regionalized lipid diffusibility in the germ cell plasma membrane during spermatogenesis in the mouse. Journal of Cell Biology, 103(5):1745–??, November 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/103/5/1745.

Wakefield:1987:DMC

[WSM^{+}87] L. M. Wakefield, D. M. Smith, T. Masui, C. C. Harris, and M. B. Sporn. Distribution and modulation of the cellular receptor for transforming growth factor-beta. Journal of Cell Biology, 105 (2):965–??, August 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/2/965.

Westendorf:1989:RCB

[WSR89] J. M. Westendorf, K. I. Swenson, and J. V. Ruderman. The role of cyclin B in meiosis I. Journal of Cell Biology, 108 (4):1431–??, April 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/108/4/1431.

Wadsworth:1989:BTI

[WSRR89] P. Wadsworth, E. Shelden, G. Rupp, and C. L. Rieder. Biotintubulin incorporates into kinetochore fiber microtubules during early but not late anaphase. Journal of Cell Biology, 109(5): 2257–??, November 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/5/2257.

Wright-Sandor:1989:AHS

[WSRT89] L. G. Wright-Sandor, M. Reichlin, and S. L. Tobin. Alteration by heat shock and immunological characterization of Drosophila small nuclear ribonucleoproteins. Journal of Cell Biology, 108 (6):2007–??, June 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/108/6/2007.

Wasenius:1989:PSB

[WSS^{+}89] V. M. Wasenius, M. Saraste, P. Salvén, M. Erämaa, L. Holm, and V. P. Lehto. Primary structure of the brain alpha-spectrin. Journal of Cell Biology, 108(1):79–??, January 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/108/1/79.

Wessels:1989:CMI

[WSV^{+}89] D. Wessels, N. A. Schroeder, E. Voss, A. L. Hall, J. Condeelis, and D. R. Soll. cAMP-mediated inhibition of intracellular particle movement and actin reorganization in Dictyostelium. Journal of Cell Biology, 109(6):2841–??, December 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/109/6/2841.

Warn:1985:TDD

[WSW85] R. M. Warn, L. Smith, and A. Warn. Three distinct distributions of F-actin occur during the divisions of polar surface caps to produce pole cells in Drosophila embryos. Journal of Cell Biology, 100(4):1010–??, April 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/100/4/1010.

Woodruff:1987:KSP

[WTB87] M. L. Woodruff, J. Theriot, and S. J. Burden. 300-kD subsynaptic protein copurifies with acetylcholine receptor-rich membranes and is concentrated at neuromuscular synapses. Journal of Cell Biology, 104(4):939–??, January 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/104/4/939.

Werb:1989:STT

[WTB^{+}89] Z. Werb, P. M. Tremble, O. Behrendtsen, E. Crowley, and C. H. Damsky. Signal transduction through the fibronectin receptor induces collagenase and stromelysin gene expression. Journal of Cell Biology, 109(2):877–??, August 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/109/2/877.

Whalen:1985:HPD

[WTBBW85] R. G. Whalen, M. Toutant, G. S. Butler-Browne, and S. C. Watkins. Hereditary pituitary dwarfism in mice affects skeletal and cardiac myosin isozyme transitions differently. Journal of Cell Biology, 101(2):603–??, August 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/101/2/603.

Williams:1989:MCS

[WVCR89] B. D. Williams, M. A. Velleca, A. M. Curry, and J. L. Rosenbaum. Molecular cloning and sequence analysis of the Chlamydomonas gene coding for radial spoke protein 3: flagellar mutation pf-14 is an ochre allele. Journal of Cell Biology, 109 (1):235–??, July 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/1/235.

Wang:1986:MBT

[WVLC86] D. Wang, A. Villasante, S. A. Lewis, and N. J. Cowan. The mammalian beta-tubulin repertoire: hematopoietic expression of a novel, heterologous beta-tubulin isotype. Journal of Cell Biology, 103(5):1903–??, November 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/103/5/1903.

Wehland:1987:TTL

[WW87a] J. Wehland and K. Weber. Tubulin-tyrosine ligase has a binding site on beta-tubulin: a two-domain structure of the enzyme. Journal of Cell Biology, 104(4):1059–??, January 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/104/4/1059.

White:1987:APA

[WW87b] J. M. White and I. A. Wilson. Anti-peptide antibodies detect steps in a protein conformational change: low-pH activation of the influenza virus hemagglutinin. Journal of Cell Biology, 105 (6):2887–??, December 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/6/2887.

Wang:1988:ASM

[WW88] K. Wang and J. Wright. Architecture of the sarcomere matrix of skeletal muscle: immunoelectron microscopic evidence that suggests a set of parallel inextensible nebulin filaments anchored at the Z line. Journal of Cell Biology, 107(6):2199–??, December 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/107/6/
2199.

Wolin:1989:SRP

[WW89] S. L. Wolin and P. Walter. Signal recognition particle mediates a transient elongation arrest of preprolactin in reticulocyte lysate. Journal of Cell Biology, 109(6):2617–??, December 1989. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/109/6/2617.

Wilson:1987:IEA

[WWN87] J. M. Wilson, J. A. Whitney, and M. R. Neutra. Identification of an endosomal antigen specific to absorptive cells of suckling rat ileum. Journal of Cell Biology, 105(2):691–??, August 1987. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/105/2/691.

Wallimann:1985:SLC

[WWW^{+}85] T. Wallimann, D. Walzthöny, G. Wegmann, H. Moser, H. M. Eppenberger, and F. J. Barrantes. Subcellular localization of creatine kinase in Torpedo electrocytes: association with acetylcholine receptor-rich membranes. Journal of Cell Biology, 100 (4):1063–??, April 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/100/4/1063.

Yabkowitz:1987:LIS

[YB87] R. Yabkowitz and D. R. Burgess. Low ionic strength solubility of myosin in sea urchin egg extracts is mediated by a myosinbinding protein. Journal of Cell Biology, 105(2):927–??, August 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/105/2/
927.

Yuan:1987:TIM

[YBBS87] L. Yuan, J. G. Barriocanal, J. S. Bonifacino, and I. V. Sandoval. Two integral membrane proteins located in the cis-middle and trans-part of the Golgi system acquire sialylated N-linked carbohydrates and display different turnovers and sensitivity to cAMP-dependent phosphorylation. Journal of Cell Biology, 105(1):215–??, July 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/105/1/215.

Yednock:1987:RIL

[YBSR87] T. A. Yednock, E. C. Butcher, L. M. Stoolman, and S. D. Rosen. Receptors involved in lymphocyte homing: relationship between a carbohydrate-binding receptor and the MEL-14 antigen. Journal of Cell Biology, 104(3):725–??, March 1987. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/104/3/725.

Yannariello-Brown:1988:DKL

[YBWLM88] J. Yannariello-Brown, U. Wewer, L. Liotta, and J. A. Madri. Distribution of a 69-kD laminin-binding protein in aortic and microvascular endothelial cells: modulation during cell attachment, spreading, and migration. Journal of Cell Biology, 106 (5):1773–??, May 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/106/5/1773.

Yechiel:1987:MSD

[YE87] E. Yechiel and M. Edidin. Micrometer-scale domains in fibroblast plasma membranes. Journal of Cell Biology, 105(2):755–??, August 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/105/2/
755.

Yamamoto:1987:TDR

[YF87] K. Yamamoto and H. D. Fahimi. Three-dimensional reconstruction of a peroxisomal reticulum in regenerating rat liver: evidence of interconnections between heterogeneous segments. Journal of Cell Biology, 105(2):713–??, August 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/105/2/713.

Young:1988:HIB

[YHS88] P. R. Young, D. J. Hazuda, and P. L. Simon. Human interleukin 1 beta is not secreted from hamster fibroblasts when expressed constitutively from a transfected cDNA. Journal of Cell Biology, 107(2):447–??, August 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/107/2/447.

Yin:1988:IPM

[YIJ88] H. L. Yin, K. Iida, and P. A. Janmey. Identification of a polyphosphoinositide-modulated domain in gelsolin which binds to the sides of actin filaments. Journal of Cell Biology, 106 (3):805–??, March 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/106/3/805.

Yancey:1989:KPH

[YJL^{+}89] S. B. Yancey, S. A. John, R. Lal, B. J. Austin, and J. P. Revel. The 43-kD polypeptide of heart gap junctions: immunolocalization, topology, and functional domains. Journal of Cell Biology, 108(6):2241–??, June 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/108/6/2241.

Yancey:1988:EGM

[YKCR88] S. B. Yancey, K. Koh, J. Chung, and J. P. Revel. Expression of the gene for main intrinsic polypeptide (MIP): separate spatial distributions of MIP and beta-crystallin gene transcripts in rat lens development. Journal of Cell Biology, 106(3):705–??, March 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/106/3/
705.

Yuspa:1989:EME

[YKSR89] S. H. Yuspa, A. E. Kilkenny, P. M. Steinert, and D. R. Roop. Expression of murine epidermal differentiation markers is tightly regulated by restricted extracellular calcium concentrations in vitro. Journal of Cell Biology, 109(3):1207–??, September 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/109/3/1207.

Yang:1985:MWI

[YLGG85] H. Y. Yang, N. Lieska, A. E. Goldman, and R. D. Goldman. A 300,000-mol-wt intermediate filament-associated protein in baby hamster kidney (BHK-21) cells. Journal of Cell Biology, 100 (2):620–??, February 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/100/2/620.

Yamashiro:1987:AMD

[YM87a] D. J. Yamashiro and F. R. Maxfield. Acidification of morphologically distinct endosomes in mutant and wild-type Chinese hamster ovary cells. Journal of Cell Biology, 105(6):2723–??, December 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/105/6/
2723.

Yamashiro:1987:KEA

[YM87b] D. J. Yamashiro and F. R. Maxfield. Kinetics of endosome acidification in mutant and wild-type Chinese hamster ovary cells. Journal of Cell Biology, 105(6):2713–??, December 1987. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/105/6/2713.

Yoshino:1987:DCM

[YMD87] J. E. Yoshino, P. W. Mason, and G. H. DeVries. Developmental changes in myelin-induced proliferation of cultured Schwann cells. Journal of Cell Biology, 104(3):655–??, March 1987. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/104/3/655.

Yamashiro-Matsumura:1986:ILK

[YMM86] S. Yamashiro-Matsumura and F. Matsumura. Intracellular localization of the 55-kD actin-bundling protein in cultured cells: spatial relationships with actin, alpha-actinin, tropomyosin, and fimbrin. Journal of Cell Biology, 103(2):631–??, August 1986. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/103/2/631.

Yamashiro-Matsumura:1988:CKN

[YMM88] S. Yamashiro-Matsumura and F. Matsumura. Characterization of 83-kilodalton nonmuscle caldesmon from cultured rat cells: stimulation of actin binding of nonmuscle tropomyosin and periodic localization along microfilaments like tropomyosin. Journal of Cell Biology, 106(6):1973–??, June 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/106/6/1973.

Yamamoto:1985:CPT

[YMT85] A. Yamamoto, R. Masaki, and Y. Tashiro. Is cytochrome P-450 transported from the endoplasmic reticulum to the Golgi apparatus in rat hepatocytes? Journal of Cell Biology, 101(5): 1733–??, November 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/101/5/1733.

Yaar:1986:NHE

[YPG86] M. Yaar, A. V. Palleroni, and B. A. Gilchrest. Normal human epidermis contains an interferon-like protein. Journal of Cell Biology, 103(4):1349–??, October 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/103/4/1349.

Yurchenco:1987:BMS

[YR87] P. D. Yurchenco and G. C. Ruben. Basement membrane structure in situ: evidence for lateral associations in the type IV collagen network. Journal of Cell Biology, 105(6):2559–??, December 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/105/6/

Yednock:1987:PDB

[YSR87] T. A. Yednock, L. M. Stoolman, and S. D. Rosen. Phosphomannosylderivatized beads detect a receptor involved in lymphocyte homing. Journal of Cell Biology, 104(3):713–??, March 1987. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/104/3/713.

Yassin:1985:ECF

[YSW^{+}85] R. Yassin, J. Shefcyk, J. R. White, W. Tao, M. Volpi, T. F. Molski, P. H. Naccache, M. B. Feinstein, and R. I. Sha’afi. Effects of chemotactic factors and other agents on the amounts of actin and a 65,000-mol-wt protein associated with the cytoskeleton of rabbit and human neutrophils. Journal of Cell Biology, 101(1):182–??, July 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/101/1/182.

Zieske:1989:EVS

[ZBG89] J. D. Zieske, G. Bukusoglu, and I. K. Gipson. Enhancement of vinculin synthesis by migrating stratified squamous epithelium. Journal of Cell Biology, 109(2):571–??, August 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/109/2/571.

Zijderhand-Bleekemolen:1987:LWB

[ZBSS^{+}87] J. E. Zijderhand-Bleekemolen, A. L. Schwartz, J. W. Slot, G. J. Strous, and H. J. Geuze. Ligand- and weak base-induced redistribution of asialoglycoprotein receptors in hepatoma cells. Journal of Cell Biology, 104(6):1647–??, June 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/104/6/1647.

Zimmer:1988:FNE

[ZDH88] F. J. Zimmer, C. Dreyer, and P. Hausen. The function of the nuclear envelope in nuclear protein accumulation. Journal of Cell Biology, 106(5):1435–??, May 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/106/5/1435.

Zokas:1987:CLC

[ZG87] L. Zokas and J. R. Glenney. The calpactin light chain is tightly linked to the cytoskeletal form of calpactin I: studies using monoclonal antibodies to calpactin subunits. Journal of Cell Biology, 105(5):2111–??, November 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/105/5/2111.

Zuniga:1986:CVC

[ZH86] M. C. Zuniga and L. E. Hood. Clonal variation in cell surface display of an H-2 protein lacking a cytoplasmic tail. Journal of Cell Biology, 102(1):1–??, January 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/102/1/1.

Zervos:1985:PGJ

[ZHE85] A. S. Zervos, J. Hope, and W. H. Evans. Preparation of a gap junction fraction from uteri of pregnant rats: the 28-kD polypeptides of uterus, liver, and heart gap junctions are homologous. Journal of Cell Biology, 101(4):1363–??, October 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/101/4/1363.

Zampighi:1989:SOP

[ZHES89] G. A. Zampighi, J. E. Hall, G. R. Ehring, and S. A. Simon. The structural organization and protein composition of lens fiber junctions. Journal of Cell Biology, 108(6):2255–??, June 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/108/6/2255.

Zampighi:1988:SCG

[ZKR^{+}88] G. Zampighi, M. Kreman, F. Ramón, A. L. Moreno, and S. A. Simon. Structural characteristics of gap junctions. I. Channel number in coupled and uncoupled conditions. Journal of Cell Biology, 106(5):1667–??, May 1988. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/106/5/1667.

Zuk:1989:TCG

[ZMH89] A. Zuk, K. S. Matlin, and E. D. Hay. Type I collagen gel induces Madin–Darby canine kidney cells to become fusiform in shape and lose apical-basal polarity. Journal of Cell Biology, 108(3):903–??, March 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/108/3/903.

Zhang:1989:STD

[ZN89] J. T. Zhang and B. J. Nicholson. Sequence and tissue distribution of a second protein of hepatic gap junctions, Cx26, as deduced from its cDNA. Journal of Cell Biology, 109(6):3391–??, December 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/109/6/
3391.

Zieseniss:1985:SEP

[ZP85] E. Zieseniss and H. Plattner. Synchronous exocytosis in Paramecium cells involves very rapid (less than or equal to 1 s), reversible dephosphorylation of a 65-kD phosphoprotein in exocytosiscompetent strains. Journal of Cell Biology, 101(6):2028–??, December 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
101/6/2028.

Zerwes:1987:PSP

[ZR87] H. G. Zerwes and W. Risau. Polarized secretion of a plateletderived growth factor-like chemotactic factor by endothelial cells in vitro. Journal of Cell Biology, 105(5):2037–??, November 1987. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/105/5/
2037.

Zagouras:1989:CTS

[ZR89] P. Zagouras and J. K. Rose. Carboxy-terminal SEKDEL sequences retard but do not retain two secretory proteins in the endoplasmic reticulum. Journal of Cell Biology, 109(6): 2633–??, December 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/6/2633.

Zanetti:1985:TSD

[ZRW85] M. Zanetti, A. Ratcliffe, and F. M. Watt. Two subpopulations of differentiated chondrocytes identified with a monoclonal antibody to keratan sulfate. Journal of Cell Biology, 101(1):53–??, July 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/101/1/
53.

Zhang:1986:DVS

[ZSB86] Y. Zhang, S. A. Shafiq, and D. Bader. Detection of a ventricularspecific myosin heavy chain in adult and developing chicken heart. Journal of Cell Biology, 102(4):1480–??, April 1986. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/102/4/1480.

Zimmerberg:1985:ESU

[ZSE85] J. Zimmerberg, C. Sardet, and D. Epel. Exocytosis of sea urchin egg cortical vesicles in vitro is retarded by hyperosmotic sucrose: kinetics of fusion monitored by quantitative light-scattering microscopy. Journal of Cell Biology, 101(6):2398–??, December 1985. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/101/6/
2398.

Zachary:1986:EEE

[ZSSR86] I. Zachary, J. W. Sinnett-Smith, and E. Rozengurt. Early events elicited by bombesin and structurally related peptides in quiescent Swiss 3T3 cells. I. Activation of protein kinase C and inhibition of epidermal growth factor binding. Journal of Cell Biology, 102(6):2211–??, June 1986. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/102/6/2211.

Zaner:1989:VFA

[ZV89a] K. S. Zaner and P. A. Valberg. Viscoelasticity of F-actin measured with magnetic microparticles. Journal of Cell Biology, 109 (5):2233–??, November 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/109/5/2233.

Zimmer:1989:ACM

[ZV89b] D. B. Zimmer and L. J. Van Eldik. Analysis of the calciummodulated proteins, S100 and calmodulin, and their target proteins during C6 glioma cell differentiation. Journal of Cell Biology, 108(1):141–??, January 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/108/1/141.

Zeitlin:1989:ASC

[ZWE89] S. Zeitlin, R. C. Wilson, and A. Efstratiadis. Autonomous splicing and complementation of in vivo-assembled spliceosomes. Journal of Cell Biology, 108(3):765–??, March 1989. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/108/3/765.